A Complete Bibliography of Publications in Computer Vision and Image Understanding: CVIU

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

07 June 2024 Version 2.77

Title word cross-reference

(18, 6) [MW00]. + [BCF06]. 1 [AVGASAP15, BDL^{+}06]. 101 [FFFP07]. 16 [MMS97]. 2 [AXSVL14, AVGASAP15, Ano01m, AS08b, ABVC16, AVC19, AM97, BN15, BBC00, BL16, Bd96, BZ99, BCF06, CL18, CFM^{+}13, CC96, DB03, DAM12, DBB13, FPC^{+}08, FAB97, FKL^{+}98, GSPL10, HB98a, HUI16, HB98b, IAP^{+}11, JDP97, JC98, KMB97, KTE^{+}17, KSL^{+}20, KM03, KMN11, KNO^{+}09, Lau97, LST13, LDH^{+}15, LQQS21, LS12, Luc01, Mal21, Mil09, MBMC11, MIP16, NT10, Neg12, NKPT13, NSEA13, ODT17, OJRT08, SYL^{+}24, Ste01, TH04, WCZ02, YGC15]. 2.5 [MCB13, SRHC13, ZP11]. 3 [ACF00, AMNCM16, AXSVL14, ACG^{+}09, ÁB13, ALY^{+}22, AS08b, ABVC16, AVC19, AM97, ARARCE11, ACDB12, BN15, BM99, BB16, BI10, BI11, BCA98, Bar05, BSALF18, BT05, BR95, BY12, BW15, Bd96, BZ99, BAMK18, BB24, BCF06, BGK95, BF05, BS00a, BBH14, BSBW14, BMX22, COW98, CGH08, CLZY15, CM12, CK11, CL18, CS98, CYNO11, CC11, CPPY21, CLCO13, CLO17, CFM^{+}13, CC96, CP20, CG04, CS00, CPS10, DT96b, Dam08, DWB11, Dan97, DWV19, DF01, DMSM21, DSY10, EK98, EOPS22, ES04, FBF08, FF09, FRL^{+}98, FDMA97, FAB97, FKL^{+}98, FL96, FO18, FCOK24, GM19, GFL^{+}19, GGGROE^{+}17, GSPL10, GHMT09, GKBW14, GSV05, GW07, GLZF23, Gui98, Gui99, GPC^{+}10, GML^{+}21, GWFF22, GSK02, HFKN97, HUI16, HRHZ17, HASS10, HRS02, HR99, Hen98, HSS^{+}16, HGSM11, HMB17]. 3 [HG11, HMF10, HCLZ21, HGB98, IAP^{+}11, IDY^{+}18, JZWD16, JRBD^{+}15, Jok98, JSC23, dOSJVBS12, KTE^{+}17, KSL^{+}20, KC22, KSF16, KHH^{+}22, KMA^{+}00, KNO^{+}09, LCT09, LM96, Lau97, LPS^{+}11, LST13, LÁB15, LAFLB16, Lhu23, LS08, LLG^{+}14, LLL^{+}15a, LDH^{+}15, LSGY24, LDCX24, LSHT02, LS12, LMM22, LSTF12, LEA^{+}10, LK00, LDL^{+}19, MS96a, MW00, MSV^{+}20, Mal21, MBD^{+}22, MFJ95, MC09b, MMA06, MOB14, MWTN04, MCT10, Mil09, MBMC11, MKY01, MB95, MJPS16, MIP16, NSK^{+}97, NG98b, NT10, NFA04, NL96, NDO09, NSEA13, OG98, OMBH06, OJRT08, OCVV04, PSR08, PHH^{+}15, PMW05, Pud98, QL96, RAH97, RB18, RZH17, Rem04, RZZ23, Ros10, RT14, SC96, SECS15, STC^{+}16, SCD11, SBIK16, ST96, SCALFG^{+}18, STV09, SS17a, SSHP17, SM06, SN99, Shi99, SKU^{+}09, SBN^{+}24, ST10, SKVS13, SJH17, SYL^{+}24, SPQ^{+}17, SB00, Ste01]. 3 [SWS11, SKBS13, SWMM22, SS11, SB02, TLFM23, TGQ23, THH^{+}23, TB99, TPT15, TPT17, TN05, TN08, TML00, THL03, UK12b, UFF06, VBVB19, VV02, VBT19, VPP^{+}23, VAC16, VKP98, WCZ23, WPS03, WTZ^{+}21, WHJK23, WWLV11, XOF05, XP11, YB07, YQL^{+}23, YHR^{+}05, YZX^{+}17, YHW^{+}23, YT99, YC98, YJC^{+}09, YZL^{+}21, YARL^{+}20, ZW97, ZZK^{+}20, ZSK^{+}23, ZZJS18, ZZC^{+}13, ZT15, ZCLX20, ZLHJ18, ZDZ^{+}23, ZCW24, ZH04, Ziv10]. 3 × 3 [SW04]. 4 [ABB^{+}23, CLZY15, RWWH00, WPI^{+}16]. 5 × 5 × 5 [SB02]. 6 [FPMK19, SIT07, TWQW23, ZC19]. 8 [CPC99]. ^{2} [ZLL^{+}24]. ^{2+} [BVWS21]. α [JZZ23]. L_{2} [GD19]. d [ADC19, Pat13]. F [LMRMJ08]. G [WGAD14]. k [GC19, JLD12]. l_{0} [LLY^{+}18]. l_{1} [DOSD11]. L_{2} [CH11, CZS^{+}20]. l_{p} [QDLB17]. M [HBH11]. n [DSdlH^{+}11, KCD00]. O(1) [PTM20]. P [Loh10]. q [MRW^{+}97]. S^{3} [HDZR24]. t [XQZL23]. Z^{2} [Egg98].

-based [PLLL03]. -D

[LEA^{+}10, BN15, BT05, BGK95, CGH08, CC96, FRL^{+}98, FL96, JDP97, JZWD16, LCT09, LPS^{+}11, LSHT02, MKY01, NT10, Neg12, NL96, Rem04, WCZ02, YHR^{+}05]. -dimensional [KCD00, Pat13]. -disparity [WGAD14]. -DOF [SIT07, FPMK19, TWQW23]. -EGAN [JZZ23]. -Energy [JZZ23]. -estimator [HBH11]. -hidden [XQZL23]. -means [JLD12]. -measure [LMRMJ08]. -Neighborhood [MMS97]. -norm [QDLB17]. -Point [CPC99]. -Series [MRW^{+}97]. -simple [Loh10]. -SLAM [GC19]. -sphere [PHH^{+}15]. -state [Ros10].

/background [ZHS^{+}24].

1 [BVWS21, KLKF20]. 112 [PZ09]. 113 [MBMC11]. 114 [AK11]. 117 [BB15a]. 12th [Rei16]. 1999 [Ros00b].

2.5D [LS09, SCC^{+}22]. 2010 [KB12]. 2019 [Ano19a, Ano19n]. 2020 [Ano20a, Ano20b, Ano20c, Ano20o, Ano20p, Ano20r, Ano20q, Ano20s, Ano20t, Ano20u, Ano20v]. 2021 [Ano21a, Ano21b, Ano21o, Ano21p, Ano21s, Ano21q, Ano21r, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x]. 2022 [Ano22a, Ano22b, Ano22c, Ano22o, Ano22p, Ano22q, Ano22s, Ano22r, Ano22t, Ano22u, Ano22v]. 2023 [Ano23a, Ano23b, Ano23c, Ano23p, Ano23q, Ano23r, Ano23t, Ano23s, Ano23u, Ano23v, Ano23w, Ano23x]. 2024 [Ano24a, Ano24b, Ano24k, Ano24l, Ano24n, Ano24m, Ano24o, Ano24p]. 214 [Oli01]. 230 [MFSB23b]. 2D [BB04].

3d [CKF18, LZWP03]. 3dSOBS [MP14].

6DOF [ANHGS17, SE11].

’95 [Ano95a]. ’97 [Ano96d]. 99 [Ano06h].

AAM [ARARCE11]. AAMs [HDF12]. abandoned [DETE17]. abnormal [XG08a, ZhZFL22]. abnormality [ZhZFL22]. Absolute [DPB00, Kis96b, PZC17, BK07, Dem05, WSFTK18]. Absorbing [PKH23]. abstraction [HMB17, MDFS11a]. AC [BCC^{+}21]. AC-VRNN [BCC^{+}21]. Accelerated [AHDM10]. acceleration [CO16, NHH14]. Access [DCCL99, SGA12]. accumulation [BCM13]. Accuracy [ACB98, LHH^{+}98, Sha06, Tan95, AVGASAP15, BHMB10, GGGROE^{+}17, GBF12, HCC^{+}16, MN06, MM06, MMBG18]. Accuracy-Based [Tan95]. Accurate [AK10, AK11, AS08b, BGK98, CJC01, FS03, KKSC23, KSHE20, Lin02, MC09a, MG95, PYWZ17, TLCH05, WB15, ZHL^{+}20, AMN18, AVC19, CZS^{+}20, Coe12, DSK^{+}20, FDSB22, GBB^{+}18, GZY23, LSKK10, MC20, PZX13, RTM^{+}17, SJSL21, SJH17, SBMM15, WHL^{+}20, WHL^{+}21, XTZZ14]. Accurately [LMC09]. ACCV [Ano95a]. ACCV’16 [LLNS18]. Acknowledgement [Ano15p]. Acknowledgment [Ano12n, Ano13p, Ano14g]. aclets [BPSV16]. Acoustic [CFM02, BN15, NT10]. acquired [PS12]. Acquiring [CH06]. acquisition [GCEC07, WNH05, YAK^{+}08]. across [AVBK10, JSRS08, RDS24, UMH16]. Action [BDT23, BPSV16, EK12, GBB^{+}18, IB01, MU11, RDS24, SCMP14, ZG10, AAASC11, AAL22, ASCF13, ASF14, BGE^{+}17, BFMW23, BAMK18, BB24, CLL^{+}21, CCFC13, EDJ^{+}20, FF23, IZJ^{+}17, JLD12, JLD13, KFSM17, KIS17, Kim17, KRK11, KSF16, KH13, KRS14, LGG^{+}18, LZW^{+}24, LLG^{+}24, LHZY19, LYSS12, MSF^{+}17, MYV19, NF21, NHZ^{+}22, OGB14, PC05, PWL^{+}23, PWWQ16, PKC^{+}18, PCM21, QZY^{+}24, QLY^{+}17, QCXJ19, QCL^{+}23, RG17, RKL^{+}18, SS17b, SZS17, SS21, TKK24, TKL21, TCZ^{+}12, TBC^{+}21, VAC16, VKNK14, WPQ20, WY21, WZQ^{+}23, WRB06, WRB11, WWXK24, XYQE24, XWLY23, YCZ^{+}23, YLK^{+}23, YST21, ZT15, ZTGL18, ZZSD21, ZCWH23]. action-recognition [PC05]. actions [AB18, BAM16, KRG17, LZS16, NY14, PD11, UK12a, WH18, YS06, YS08]. activation [KHG22, KKCK23, QCL^{+}23, SASYGCRG24, ZTGL18]. activation-based [SASYGCRG24, ZTGL18]. Active [BJ14, Car96, CTCG95, DM01, DCTO97, IP98, KR99, LVW97, LSL^{+}18, LSHT02, PK18, SI03, WCH98, YYL96, BH12, CSGM^{+}24, CUAT13, CCD11, DBZ07, GBY21, KSF19, MFB11, MSF^{+}17, MCB13, Mil09, MBMC11, MPPP14, PD05, SB18, TP05, UM05, WB12, WYC15, WWJ13a, XAB07, YLA09, ZWZZ18, TRG^{+}13]. Activities [MBHRC21, WPZ^{+}16, YB99, BKPS15, DIMT12, SG17, SM17, TSD17, VCDS^{+}17, VZP^{+}09, WSY^{+}16]. Activity [ABK16, ACP16, BLH16, CCFC13, CPT07, EDX16, HRC16, HCC^{+}16, HNB04, LWLC22, NN13, OGH04, OVJ^{+}21, PKK^{+}09, RR06, RS03, SOD10, SSdVL06, SAL16, TABK17, VB16, WLM^{+}14, YG17]. actor [FR11, EDJ^{+}20]. Actor-supervision [EDJ^{+}20]. AdaBoost [YCA^{+}10]. AdaBoost-based [YCA^{+}10]. AdaNI [LZQL24]. adaptable [UWH17]. adaptation [BVCP21, CSS^{+}13a, DPRC17, DD11a, DNG^{+}24, HG11, MJ17, PV14, QCH20, RMC^{+}22, SHSJ23, SGZ21, TKDN16, WLSO23, YWH^{+}23, YNCO11, ZdCR^{+}24, PD23]. adapted [BG18, LCSL07, VMP03]. adapters [CAO^{+}23]. Adapting [QT10]. Adaptive [BJS14, CSD^{+}24, CT12, CS04, CWW^{+}22, CLCZ23, CYC10, DD11b, DAL^{+}22, HGS08, JV97, LAFLB16, RRL20, RM02, SvdMH15, SYL^{+}24, Tan95, TZL^{+}22, WH00, WWJ13a, WLL22b, YCKA10, ZJZY16, ZMJ^{+}15, AAL22, BSM10, BÁRM23, CE14, CYD^{+}22, EDB12, FLHK08, GS08, GYCS21, HYJ11, HBB^{+}12, HLL^{+}23, HBG13, JRAJ17, JSZY17, LRW08, LL04, LJZ18, LWW^{+}21, LYG07, MFSB23b, MFSB23a, MTAA11, MJ17, MĆK09, NPM^{+}16, SBD22, TYDH18, TL16, VMN16, WSSS13, WWWF23, WWXK24, XG08a, XYQE24, YFDA17, ZJL23, ZH04, KKP24, LZQL24, PCC13]. adaptive-binning [LL04]. adaptive-resolution [ZH04]. Adding [TLB^{+}15]. Addressing [DFP23]. adipose [TLY^{+}16]. Adjacency [KCD00, YYZL19]. Adjustable [CSS13b]. adjustment [BS05, DSH04, GA09, KSY15]. ADR [KŽ12]. Advanced [ZS11]. Advances [Ano15n, HD07, CH17, GHMT09, dOSJVBS12, KU19, KHA^{+}05, MSM17, MHK06, FHSKP13]. Advantage [FL96]. advantages [KHK10]. Adversarial [BT23, CWLY22, CGL^{+}21, DWL^{+}24, FDSB22, GFL^{+}19, LB19, LCS^{+}21, LHLZ23, LHG^{+}23, MCAF21, MPT21, MC22, WZJ^{+}21, ZZK^{+}20, ZTB20, ZhZFL22, BMvT^{+}19, CYY^{+}23, FFA^{+}19, GKGM20, HSK23, JRS21, JYL23, KKRK23, LWH^{+}23, LZQL24, MFSB23b, MFSB23a, NCDG21, OZT19, OVJ^{+}21, QCH20, SLK23, SB22, SDK22, SXY^{+}23, SAV24, TY22, XSL^{+}23, ZMM^{+}22, HSTL24, ZZS^{+}23]. AdvFAS [CYY^{+}23]. Aerial [BM99, CJC^{+}98, CJC01, FKL^{+}98, FMR01, GN98, May99, PCJC98, WH01, CSK22, JRH03, KSG^{+}19, KSY15, LSC08, RTM^{+}17, SS21, TDWH07, YZ06]. aesthetic [ZSG^{+}20]. AFA [KKP24]. AFA-Net [KKP24]. affect [JGP19]. Affective [LXFM16, TMM16]. affective-interaction [TMM16]. Affine [Ano01m, BH99, Che96, Luc01, NG98a, SBZ97, ACAAC^{+}08, BCP15, BF14, FB12, GHML17, HY11, HN95, HKWC14, SOJ17, WYC15, XSL^{+}23]. affine-invariant [WYC15, XSL^{+}23]. affinities [CU10a, CU10b]. Affinity [CU10a, CU10b, LmCT16, MTR^{+}23, PDTE06, XTZZ14]. affordances [KRK11]. against [CWC^{+}20, CYY^{+}23, CCYC12, JPN^{+}22, LWH^{+}23, RH06, SXY^{+}23, ZTB20]. Age [KdVL99, OTAH20, GBVDC18, HWK^{+}21, KKSH23]. agent [GBVDC18, KK13]. agents [GLMM16, UM05]. Aggarwal [CV13]. aggregated [MYV19]. aggregating [LHG^{+}23]. Aggregation [FKL^{+}98, FBK16, MYLP98, NL23, ŠSJ^{+}20, WHL^{+}21]. aggression [KLK^{+}16]. aging [SB22, XFSC13]. agnostic [ZGC20]. AI [BCHR24]. aided [PYGGLNG17, PGP15, SB13]. aiming [FLB06]. air [BKP10]. airlight [FSI21]. albedo [TS11]. ALCN [RRL20]. Algebra [dSEdSPdMF24]. Algebraic [BGSdVL98, DC01, MNSK98, UTB^{+}11, WXC20]. Algorithm [ACB98, BM98, CPC99, CRC97, CC01, CCS95, CHRM96, DJG01, ER96, FDMA97, GSK02, LM96, LD98, MS96a, MNHO00, NDBT95, PKP97, Pud98, QL96, SCS99, SP97b, SHKP98, TV99, BGPD09, BTB14, CBD^{+}03, CMBV04, CT12, CM16, CCL04, CLL17, CR03, CSMS14, CP20, Cre08, DBF04, Dam08, DBBB14, GOF^{+}15, HDS08, HMCT22, HWW06, HZW^{+}10, Kim17, DFP^{+}13, LZLP10, LPZ08, Loh10, MYZ^{+}24, MP14, MdOBA19, PCC13, RLB17, SAS12, SW17, VRKL13, WSSS13, WK21, YB07, ZSCP08]. Algorithms [BS00b, CKK^{+}12, DRCF95, DUC97, FHP01, LPH01, LHH^{+}98, MW00, Mil99, MWL99, MEDT96, Oli00, Oli01, SUO00, SU01b, SWG02, THT^{+}98, WWW95, BVWS21, Cha21, CX11, CYG16, DSdlH^{+}11, DMSM21, GRGB^{+}13, GSGJ22, HD07, HZLM11, KK17, KBWT16, KL11, KOC17, MUS06, OSM17, PDK96, PV15, PMW05, QKH^{+}12, SW05, SV14, SRS11, SKS11]. Aligned [CSD^{+}24]. Aligning [TAC23, LSL^{+}18, WYX^{+}16]. alignment [ANHGS17, BAP08, CLL^{+}21, CPS10, FHZX23, FR11, GYF18, HJ12, JT17, KA08, LLCY21, LZZ^{+}23a, LH03, MCB13, MPT21, PWL^{+}23, Sav23, SJSL21, SJH17, VBVB19, WLSO23, ZH18, ZHS^{+}24]. all-focused [CMZ24]. allocation [WZX^{+}14]. allowing [KDV12]. Aloimonos [Zha97]. alone [OSM17]. along [GTP18]. alpha [LWZP17]. alternate [ZZ10]. alternating [HMCT22, SWMM22]. Alternative [Mil99, SM13b]. altitude [CSK22]. ambiguities [CLA^{+}17, Neg12]. Ambiguity [CM99a, YK08]. American [VM01]. Amodal [BF05, AKE23]. among [SU01b, UK12a]. Amount [KABP98]. Anabranch [LNN^{+}19]. analyse [AGB^{+}15]. Analysis [ACLS98, AC99, ABW97, Ano96d, ACW^{+}16, BEPW00, CRC97, Che98, Che96, CN95, EK98, GSP01, GPK99, Gav99, GSU00, IF99, JB15, KS95, Kis96a, LZ97a, Muk97, NDN^{+}97, Nis97, Pen99, Ros95, Ros96, Ros97, Ros98, Ros99a, Ros00a, Ros00b, Ros01, RLC^{+}11, SB96a, SP97a, SHKP98, Spi98, TS01, WKI^{+}16, WPZ^{+}16, WW97, WH00, YYL98, AC07, Ang07, AZN11, BCMR16, BC10, BVVMMS15, BCM06, BW15, BRP04, BSBW14, CICN22, CHP^{+}11, CTWH15, CCL^{+}17, CPT07, CE17, CP09, CLCO13, CT13, CC03, CKS^{+}05, DB03, DRK03, DIMT12, FLB06, FB16, FPMK19, GOF^{+}15, GYTL09, Hu08, HW06, HKZ^{+}16, ITNP12, JGP19, JSC23, KFRD^{+}18, KLL^{+}11, KB12, KSG^{+}13, LB14, LFMP13, LL04, LLE^{+}09, LPVM13, LP10, LWH03, MPF07, MVP06, MP09a, MST16, MHK06, MDL^{+}23, MĆK09, NGR24, OH05]. analysis [OTAH20, PE09, PSE^{+}11, PKK^{+}09, Pop07, RZH17, RMN^{+}17, ROGT14, SOK16, SJB20, SBIK16, SPT^{+}18, SJST07, SCR^{+}17, SYK96, SAC^{+}12, SPK23, SSdVL06, SCCP05, TKK24, TPNP15, TCZ^{+}12, TDT12, UTB^{+}11, VMP03, WD14, WY07, WS08, WLI08, WLMG08, XG08b, XWC^{+}23b, YLM11, YSS^{+}14, YHS^{+}20, YSD03, ZZP^{+}16, ZMCA05, ZZJS18, ZG10, ZZP12, NLW13, ZZCL14]. Analysis-by-synthesis [JB15]. analysis-friendly [CTWH15]. analytic [FAZ14, XSD12]. analytical [YSL11]. analytics [GK23]. Analyzing [AM00, Bic98, Bd96, CCR^{+}05, JB23, SP19, CKS^{+}05, FS03, MB05, RSPD12, SRO^{+}19]. Anatomical [HRS02, LSB^{+}00, LK00, MMA06, ZZC^{+}13]. anatomy [EB14]. Anchor [BFD22, CZS^{+}20, PZM^{+}21, SKA23, ZJJ22]. anchor-free [PZM^{+}21, ZJJ22]. ancient [PRG^{+}14]. And-Or [ZWZZ18]. angiograms [LAFLB16, NBDB04]. angiography [BT05]. angle [BPBS13, UWH17]. Angular [APV99, MP20]. Angular-Based [APV99]. Animated [FM99]. Anisotropic [BS00a, BI11, GR05, KGC05, SGS^{+}10]. Annealed [RRR11]. Annealing [BCG95, PB99, JLL13]. Annotated [Ros01, EHG^{+}10]. Annotation [XL98, ABVC16, ABC^{+}03, BCNS15, BSMK13, LTCT14, SS17b, TLWT12, WHM^{+}09, WZCY22, ZTH^{+}14, ZZ20]. annotations [CLFH22, Mah16, SWMM22, ZWZZ18]. annotators [SYPK13]. Announcement [Ano97a, Ano01a, Ano01b, Ano03a, Ano03b, Ano03c, Ano96a]. anomalies [CHP^{+}11, RL13, WK21]. Anomalous [JYTK11, XYRS17]. anomaly [BIG^{+}23, BDS12, FWL^{+}20, KBKS18, LLS21b, LHLZ23, MDK^{+}24, PKH23, SFF^{+}18, SY23, WX16, WGGH24, WLFL21, YGC13, SAK^{+}24]. AnomalY-Detection [SAK^{+}24]. answer [ZWZZ18]. Answering [DAZ^{+}17, OS19, JYY^{+}24, KK17, MLZRK24, RMS^{+}19, WTW^{+}17, XZW^{+}23, YXW^{+}24, CAO^{+}23]. Anthropometry [BK01]. Anti [WYW^{+}22, CYY^{+}23, ZTB20]. Anti-jamming [WYW^{+}22]. anti-spoofing [CYY^{+}23, ZTB20]. anticipation [FF23]. antipodal [LB10]. any [AVBK10]. Anytime [BAP08]. AP [CZ14]. Aperture [SGA12, BSH13]. Apparent [KMB97]. Appearance [BFY00, CW00, HF01, MKK02, SN99, TRG^{+}13, BF10, BMX22, CD13, DZL07, DB03, ESS10, EL07, Gwa17, HFR06, HJZ16, JVD^{+}20, JSRS08, KEG15, LSD^{+}07, LHYK05, LPS^{+}11, LLS21b, LLL15b, MC09a, MCB13, MSW15, MU11, QTLP22, RB16, RRAR^{+}16, SI03, ŠRDC09, TC11, XYRS17, YJ16, YO11, YT13, YG16]. Appearance-Based [CW00, SN99, ESS10, MC09a, RRAR^{+}16, ŠRDC09, TC11]. appearances [BCC^{+}18, GPG^{+}15]. applicability [KHK10]. applicable [Ano17j, Ano17k, Ano17l, Ano18k]. Application [ABK^{+}18, ACF00, AM01, AVC19, GK98, JLD12, KABP98, LSB^{+}00, MCPB00, MAM97, OMLL98, RAC^{+}13, RAP16, RMFB02, SOL16, SRHC13, TW98, TZ00, VMP03, WSKH13, BT17, BvdHL^{+}13, BB13, BB15a, CTCG95, DB14, GCFMT12, GWT09, KGK10, KGFP10, KMBH09, Lhu23, MUS06, Mar07, PWSvdH17, PD14, PMC13, RC03, RCTV12, PBPD^{+}17, SA04, TNO24, WZY13, Ang07, BC10]. Applications [Ano98d, BY98, Gui99, Gui00, HT98, MS96a, MKK02, NPBM22, SU01b, SWG02, TPR^{+}00, WKI^{+}16, CBT^{+}04, DB03, DBBB14, GWFF22, JB23, KLBP11, KPPK09, LL04, MBD^{+}22, MM05, NBFG20, RC13, SC96, Sah05, TGM^{+}17, TMB12, UWH17, WS08, WB12, WTBdB15, XSD12, YJC^{+}09, YG16, ZT09]. Applied [WF02, AGB^{+}15, GGGROE^{+}17, LEE^{+}18, MJ11]. Approach [APV99, AMMV99, BZ99, CH96, CCP97, DGH98, DY98, DC01, FM99, HLF^{+}97, HP96, KW00, LSHT02, MRW^{+}97, MYLP98, NDN^{+}97, OMLL98, PLL00, RJ00, RH95, Tsa96, YB95, ZXK02, ÅS17b, Ano06h, BBSD15, BMJF^{+}17, BCHR24, BT05, BDS12, BPC^{+}17, BCM06, BL16, BAMK18, BNG03, BPB11, BT23, CTM^{+}13, CDT11, CH17, CU20, DK13, DAL^{+}22, FFFP07, FKV^{+}11, FSV07, FF23, GRGB^{+}13, GKK05, GMF14, GC19, HHG^{+}20, HBH10, HDL^{+}20, HRC09, HW07, HC13c, IDY^{+}18, JNLG15, KS15, KDSF20, KL11, KS12, LEE^{+}18, LJHH07, LDH^{+}15, LG17, LS12, LZmC^{+}17, LRD19, MPST08, MNMK16, MHMO09, MMP09, ME18, NHSC09, Nic95, OAGN18, ODT17, PRG^{+}14, PLYW21, PC15, PTE12, QCL^{+}23, RRK13, SM12, Sha06, SCL13, SOJ17, SAC09, SPK14, TMNM09, TH06, THL03, UJ22, VBA19, VMC^{+}16, VJ17, WZT13, WLX^{+}14, WAPB17]. approach [WGGH24, WHY^{+}23, WDB12, WSFTK18, XSD12, XW16, YS08, ZY14, dP10]. Approaches [LCZ^{+}01, RC97, BCF06, DCFM07, GMM15, GJ10, HHWP03, KYM13, KMN11, SJST07]. Approximate [Che96, DBB13, ZCK09, CLL17]. Approximation [BM98, DGH98, JB99, KP97, LM99b, LL97b, Coe12, KA08, KHK10, LRLB11, LRLR15, SZ16, SWMM22]. Approximations [DG01, CDJM14, Pat13]. April [Ano20a, Ano21a, Ano22a, Ano23a, Ano24a]. Arbitrary [ANM98, HDZR24, APB10, Coe12, CDIF14, KK09, ZSL^{+}24]. Arbitrary-scale [HDZR24]. Arc [WWW95, dMFU10]. arc-weight [dMFU10]. architectural [KRBSV17]. architecture [DRAB08, HGP15, LWH^{+}23, MFG10, SB18, SJSL21, SCS14, SIT07, TRPD20, XWC^{+}23b, ZLLP21]. Architectures [TV99, LJJY24]. Arcs [DGH98, HB98b, Lil97]. ARCTIC [QCL^{+}23]. Area [Jok98, KSI98, Mil99, MSW96, CKM11, CCPK16, GE08, KM03, PK18]. Area-Based [Jok98]. Areas [FMR01, YHS^{+}20]. ARG [PLLL03]. Arrays [THT^{+}98, CPT07]. art [JM09b, KTP08, SCD11, SHL18]. Artefacts [PMV00]. arterial [EX17]. artery [LAFLB16]. article [Ano01m]. Articulated [ACLS98, DF01, GESB95, Tay00, BCMCB09, DGC12, HW07, IAP^{+}11, LDL^{+}19, MFB11, Mal21, RRR11]. articulating [NHY10]. artificial [CKF18, CNO^{+}16, FY06, HC13a, MNMK16].

As-planar-as-possible [PY19]. Ascender [CJC^{+}98]. Asian [Ano95a, Rei16]. ASIST [LRF^{+}17]. ASM [CUAT13]. Aspect [Mun95, NWP97, ACDB12, FFA^{+}19]. Aspect-Trees [Mun95]. Aspects [SKOS95, VM01]. ASSD [YWM19]. ASSERT [SBK^{+}99]. Assessing [BVWS21, DNG^{+}24, JOvW^{+}05, CCTCR09, YZY11]. Assessment [BS00a, LKZ20, LLJ^{+}23, OAGN18, SRP10, TPD^{+}16]. assessment-guided [LKZ20]. assignment [Kim17, MEYD11, PTK24]. assistance [HPvB^{+}10, NPM^{+}16, OBTMT15, PBPD^{+}17, WWH07]. assisted [ÁB13, GRMH19, PJW11, YG16, YG17]. assisting [CNO^{+}16]. Assistive [FKL^{+}16b, FKL^{+}16a, CEA16, CSV^{+}16, CMCM16, CC16, LMT^{+}17, MML^{+}16b, PLB16, RRAR^{+}16]. association [LJC^{+}23, WLL^{+}22a, WB16]. Assumption [CM99a]. assumptions [WS06]. astounding [YCZ^{+}23]. asymmetric [EB13, LLNZ22, WWCZ15]. asymmetry [LSCM03]. Asynchronous [JDP97]. atlas [LvdHK^{+}15, ZZC^{+}13]. Atmospheric [ZHZ17]. ATR [LCZ^{+}01]. attachment [CLA^{+}17]. attack [CWC^{+}20, HSTL24, NCDG21, NCDG21]. Attacks [MCAF21, JPN^{+}22, OVJ^{+}21, XSL^{+}23, ZTB20]. Attending [TLMT^{+}05]. attends [LGG^{+}18]. Attention [ABJ^{+}21, DAZ^{+}17, DCTO97, GZY23, GFW13, HRC09, JYL23, KKP24, LZZ^{+}23a, LXS^{+}23, LLG^{+}23, MK24, OS19, QCXJ19, SKOS95, TW98, YWL^{+}20, YH19, ZWW^{+}20, ZSC^{+}23, BBHF10, CXYZ24, CAGN24, DZQ24, DL05, FOCSB^{+}20, Ham05, HBZ^{+}24, IKST05, JZZM23, JYY^{+}24, JOvW^{+}05, KHG22, LBC^{+}21, LHG^{+}23, LWH^{+}23, LD24, LHZY19, LLWX24, LZYW23, LML^{+}23, NF21, QBZ21, QCL^{+}23, RMS^{+}19, SP23a, SZL^{+}23, SvNW23, SVA^{+}22, SFWG08, THH^{+}23, WRKP05, WPQ20, WG23, WLZM20, YK24, YQL^{+}23, ZWLH24, Ano05j, FRNS05, HH05]. Attention-based [MK24, QCL^{+}23, THH^{+}23]. Attention-from-motion [HRC09]. attention-guided [SP23a]. Attention-induced [ZSC^{+}23]. Attentional [MNE00, YYL96]. Attentive [BCC^{+}21, MHX19, XZQJ21, YWM19, CPPY21, CBTC23, ZZSD21]. attraction [RM03]. Attribute [BJ96, GK95, CWLY22, DPCA15, GKH^{+}21, PHHL23, TL15, ZTGL18, ZRKZ^{+}11]. Attributed [CTF^{+}98, PLLL03, SRS11]. Attributes [DFJL15, Hen98, JLY^{+}17, JGP19, LG24, LSTF12, NLW^{+}17, PC15, RFS03, STC14, TESY15, WHGZ20]. audio [CGD^{+}23]. audio-visual [CGD^{+}23]. Audiovisual [DGG08, SKT18]. augmentation [GYWZ23, KACR^{+}23, MLZRK24, SXY^{+}23, TB23, ZZSD21]. augmented [CKM11, GWFF22]. Augmenting [FAZ14]. August [Ano20b, Ano22b, Ano23b, Ano24b]. Aurora [GFL^{+}11]. authentication [DIMT12, PY08, UBEP09]. Author [Ano95b, Ano95c, Ano96b, Ano96c, Ano97b, Ano97c, Ano97d, Ano97e, Ano98a, Ano98b, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01c, Ano01d, Ano01e, Ano01f, Ano02a, Ano02b, Ano02c, Ano02d, Ano03n, Ano03p, Ano03q, Ano04k, Ano04l, Ano04m, Ano04n, Ano05k, Ano05l, Ano05m, Ano05n, Ano06j, Ano06k, Ano06l, Ano06m, Ano03o]. autism [CSV^{+}16]. Auto [ZH18, BB24, WGGH24]. Auto-Encoder [ZH18, BB24]. auto-encoders [WGGH24]. Autocalibration [Bri17]. AutoEncoder [BT23, FWL^{+}20]. autoencoders [UIK22, ZZK^{+}20, HCLZ21]. automata [Ros10]. Automated [CJC^{+}98, DZLH17, ES06, HPvB^{+}10, LSB^{+}00, NJ95, PKD07, RCJ^{+}13, SZ03, SRP10, CYP^{+}10, MO11, TDK10]. Automatic [ARARCE11, BL98b, CNC03, EX17, GN98, GYW^{+}22, HHAE14, KN04, KY06, KB12, KON^{+}17, LNM^{+}21, Lhu08, LSHT02, LRF^{+}17, MG95, May99, MC20, MEDT96, NY14, SCGAF^{+}17, SS17b, Tan11, VV02, XYW^{+}08, YJC^{+}09, ZZZ06, ABK^{+}18, ABVC16, ABC^{+}03, BUD19, BCNS15, BW15, CZ14, CSZ^{+}15, DK13, FFY^{+}04, HDS08, LDH^{+}15, MDdMG09, MCT10, MTC^{+}14, QKH^{+}12, RG16, RMN^{+}17, RC13, USKB10, YZX^{+}20]. automatically [MLB^{+}18]. Automating [MC20]. automation [CMH13]. Automotive [LB19]. Autonomous [KR99, BKP10, JBC08]. Autonomously [KP00]. auxiliary [BW11, CP21, FXWW17]. AVCD [DK13]. AVCD-FRA [DK13]. Average [GMT00]. averaging [MMA06]. avoidance [CSS13b, JM09a]. Avoiding [RKL^{+}18, GB13]. Award [Ano12m, Ano13o, LLNS18, Ano07f, Ano08k]. aware [ABLL19, BSRV17, BMvT^{+}19, Che24, CNO^{+}24, CACB17, DWL^{+}24, FSL24, FDW21, GK23, GWFF22, GL24, GWCO11, HCLZ21, KLO20, LTAA23, LSH19, LLNZ22, LZZ^{+}23b, LXW^{+}23, MiMO^{+}16, PL10, PMCN22, SDK22, TLH22, WCZ23, WKT22, WMZY23, XTZ^{+}18, XWC^{+}23b, YLLG18, ZYQ^{+}23, ZFG^{+}22, ZFW^{+}24]. Axes [SB98c]. axial [PA13]. Axiomatic [SU01a]. Axis [SB96b, PAK19, PCJ14, WHL14].

B [RAH97]. B-Solids [RAH97]. Back [WH18, BK07]. back-off [BK07]. Background [Ant98, DS07, RDS24, SEFV15, YCH07, ZY14, Cha21, Cha24, HKM22, JBR08, LRLB11, OSM16, OSM17, SZ07, SV14, SPK14, TA11, VTRC14, VAWW10, VWMZ15, WZCG24, YSNiT14, ZJZY16, ZZSD21, ZhZFL22, ZHS^{+}24, ZCF13]. background-action [ZZSD21]. Background-subtraction [DS07]. background-weighted [JBR08]. backgrounds [LBNS09]. backlit [LZL^{+}22]. BacklitNet [LZL^{+}22]. Backpack [HCHD01]. Backtracking [KW12]. Backviews [SK02]. Bag [PWWQ16, ADR16, KBMD15, MYV19, RG17, XQZL23, RB18]. bag-of-discriminant-words [MYV19]. bag-of-models [XQZL23]. bag-of-tracklets [ADR16]. bag-of-visual-words [KBMD15, RB18]. bag-of-words [RG17]. bagging [LLP16]. BAIT [YWH^{+}23]. Balanced [JLM22, MNL^{+}17]. Balancing [ZZH23]. Ball [MSSS09, CG09, ROJX09, WASF14, YJC^{+}09]. ball-tracking [WASF14]. Balloon [CM95]. band [Mil09, MBMC11]. bandwidth [CÇ15]. bandwidth-efficient [CÇ15]. bank [TKL^{+}09]. barrier [CSMS14, Liu10, NBFG20, SCMS13]. Base [KPH02]. baseball [GHHX04]. Based [APV99, Ano01m, BGSdVL98, BM98, BS99a, BL00, BL01, Bra97, CFS98, Che00, CCS01, CL97, CW00, DRCF95, DCCL99, DUC97, DTG96, DLHT99, DY98, Egg98, FDMA97, FL96, HTEB11, HR99, HSIW98, HF01, HLF^{+}97, HY98, IF95, JB99, Jok98, JEK98, KW00, KR98, KABP98, KMA^{+}00, KP00, KR99, LL99, LHHC98, LLSV00, LK00, Luc01, MBKB02, MS97a, MS97b, MWL99, MG01, Mok97, Muk97, NPBM22, NK00, Nis97, OG98, PLL00, PBQ99, PM97, PMV00, RWWH00, SK02, SUO00, SYF99, SB98a, SMK02, SLST99, SN99, SBK^{+}99, SPK^{+}02, SHKP98, SLL01, SL96, TI01, Tan95, TY01, TB99, TS01, VKP98, WF02, WW97, YC98, YB01, AAASC11, AYD^{+}18, AQ09, AGB^{+}15, AS09, AMCB20, AM17, AAL22, ACG^{+}09, ABEN09, AK10, AK11, ATG15, AWK04, Ang07, AS08b, AZN11]. based [AO04, AVC19, AYG23, ARARCE11, BI10, BLJ^{+}23, Bar18, BZS08, BLKG21, BY08, BSALF18, BL04, BL09, BM15, BFMW23, BB15b, BAMK18, BB24, BDFG17, BWG17, BBH14, BJS14, BH12, BRPC17, BPB11, CBD^{+}03, CGU11, CPC08, CEA16, CLZY15, CM12, CTM^{+}13, CM16, Cha24, CK11, CCPK16, CL18, CALO20, CBB19, CKPV21, CS10, CHZ^{+}13, CSLX16, CH17, CKF18, CTH20, CLZZ21, CSW^{+}24, CWW^{+}22, CAGN24, CSS13b, CE17, CJL06, CP09, CO16, CFM^{+}23, CT13, CD13, CU10a, CU10b, CNS18, CS20, CMCM16, CBTC23, CG04, CC16, CZZS07, DK13, DETE17, DLC^{+}24, DT10, DLMC16, DWB11, DKG22, DS07, DD11a, DRK03, DLV15, DZJB14, DH19, ESS10, EPH^{+}21, EH21, EDB12, EBN^{+}07, EyGS11, EB14, FPC^{+}08, FMGA^{+}12, FFY^{+}04, FHZX23, FM22, Far11, FBZP15, FB12, FKV^{+}11, FB16, FBS21, FBK16, FAB12, FSV07]. based [FKS10, FK09, FO18, GRGB^{+}13, GB10, GXC23, GRCD18, GSPL10, GBHS06, GBB^{+}18, GRB13, GGMV08, GB13, GH08, GHHX04, GCPF08, GFW13, GYW^{+}22, GZ19, GWFF22, HTNN18, Ham05, HDS08, HD09, HRHZ17, HKM22, HDZR24, HAT^{+}15, HSH07, HSBS16, HGR^{+}13, Hei04, HHWP03, HSKH07, HSK23, HFR06, HCC^{+}16, HNB04, HQN05, Hu08, HC13b, HH19, HSTL24, HMA10, HWW06, HDF12, HYW^{+}24, HGS08, ILRB04, ITNP12, JLY^{+}17, JHA17, JBC08, JBWK11, JLD13, JGM20, JM09a, JMPG11, JSC23, KKRK23, KS15, KdRM^{+}23, KBWT16, KG14, KU19, KK07, KK09, KLL^{+}11, KS12, KY06, KZ05, KDV12, KT07, KGU10, KL10, KY19, KGM19, LvdHK^{+}15, LBK10, LMRMJ08, LY05, LJHH07, LFMP13, LM16, LLG^{+}14, LLL^{+}15a, LDH^{+}15, LSP^{+}16, LJZ18, LWLC22, LLNZ22, LFLZ23, LZ24, LLG^{+}24, LZLP10, LYX^{+}24, LDCX24, LPZ08, LLWX24, LW18, LL12, LFL08, LC09, LLC11, LWLP23]. based [LML^{+}23, LEA^{+}10, LNS14, LRLR15, LBCA10, LAL^{+}10, LN10, LWH03, MT16, ML13, MRH19, MSV^{+}20, MP09a, MC09a, MSG10, MTG07, MdBJG15, MCT10, MHSP10, MdOBA19, MGPP11, MW13, Mig12, Mil09, MBMC11, MIUS16, MHK06, MFP^{+}20, MML^{+}16b, MP09b, MTAA11, MK24, MJ17, NHK08, NRJ11, NPM^{+}16, NZH^{+}23, NWJ15, OM19, OAGN18, OMBH06, ODT17, OSM16, OVJ^{+}21, PW23, PRK19, PLLL03, PT15, PL07, PSR08, PD11, Pen03, PLYW21, PV14, PKK^{+}09, PA10b, PFGG09, PR03, PKvGS16, PS15, PCM21, Pop07, PZV13, PBG04, PTK24, QCH20, QTLP22, QCL^{+}23, RB18, RM03, RB16, REF15, RRAR^{+}16, RSS07, RFS03, SKLM22, SGS^{+}10, SE11, SBB10, SM12, SB18, SOL16, SS17a, SRB21, SI03, ŠRDC09, SHE17, SG11, SLK23, SB22, SZW^{+}21, SW05, SJSL21, SF16, SM24, SASYGCRG24, SYL^{+}24, SPRS23, SM23]. based [SPK14, SH08, SVA^{+}22, SFWG08, SZB^{+}21, SY23, SHS03, SCEvdH14, TABK17, TTXT21, TAK09, TKK24, TYH^{+}21, THH^{+}23, TA13, TPT17, TRPD20, TT16, TBC^{+}21, TB13, TB23, TMN06, TC11, TVE^{+}16, TDZ^{+}20, UBEP09, VBA19, VPL23, VAWW10, VWMZ15, VAC16, WPS03, WLZW04, WZ04, WGAD14, WLX^{+}14, WWCZ15, WSY^{+}16, WW16, WAPB17, WPSL18, WLO^{+}18, WML21, WG23, WGGH24, WM20, WRB11, WS06, WLI08, WR08, WB11, WYX^{+}16, WZWH16, WLL22b, WZCY22, WWWF23, XQZL23, XAB07, XYW^{+}08, XGT^{+}22, XWLY23, XWC^{+}23b, YB07, YQL^{+}23, YHR^{+}05, YCA^{+}10, YGC13, YFX^{+}18, YYZL19, YWL^{+}22, YHW^{+}23, YS24, YSNiT14, YZL^{+}21, YZX^{+}20, YG17, ZJZY16, ZZL13, ZLZH17, ZTGL18, ZWZZ18, ZD18, ZTB20, ZWW^{+}20, ZZ20, ZJJ22, ZLFH23, ZHS^{+}24, ZWLH24, ZZCL14, ZLS^{+}13, ZCF13, ZWL16, ZHZ17, ZUS06, ZCK09, dSdSF^{+}12, dSM14, FRNS05, ZH18]. baseline [LWlZ16, MDK^{+}24, YCZ^{+}23]. baselines [DM24]. Bases [Nis95]. Basic [ME98a]. BasicTAD [YCZ^{+}23]. basis [BSM10, BH12, DLV15, LPR^{+}03, WR08]. basketball [CD10, PKK^{+}09]. Bayesian [AMGG^{+}16, BAPXH16, Car96, CCPK16, CC07, DLF06, FFFP07, JNLG15, KDV12, LWH03, MC09a, MOB14, QC04, RH95, SKLM22, SC00a, SAC09, SPW15, SS11, TS16, TN07, WLW^{+}16, YC98, ZCK09]. be [MRdRGC23]. bead [FLCdA06]. beam [BZP^{+}23, HSTL24]. beauty [LB14]. Beckmann [RH06]. bee [CKF18]. beginning [WH18]. Behavior [GJH01, SC00a, GZJ05, KDV12, PBI16, RFF23, TDT12]. Behaviors [GMW12, SVS97, WWH07]. Behaviour [CX11, CGH08, HFR06, SGH07, WMBY12, XG08a, ZZP^{+}16]. belief [BCMCB09, CS07, PBW14, PL08, TB13]. belief-propagation [PBW14]. Benchmark [DM24, LWlZ16, AA20, CGD^{+}23, EHG^{+}10, LLL^{+}15a, SCR^{+}17, THL13, WDC^{+}20]. Benchmarking [MNCG01, LYBT17]. benchmarks [CH17, DFS08]. benefit [GKGM20]. best [AQ09, TCB^{+}08]. better [NHTG15]. between [Åst97, BS96, BDFG17, CU11, Col97, CDH99, FDC^{+}19, HSK23, HLKK19, KHB01, KŽ12, MGS15, PRW97a, STC14, ÜE01, WDN^{+}12]. Beyond [CM99a, FHSKP13, LCS^{+}21, BCC^{+}18, HD07]. Bi [LDT21, JSZY17, OAGN18, SXZZ24, ZJJ22]. Bi-branch [LDT21]. bi-channel [JSZY17]. bi-directional [OAGN18, SXZZ24, ZJJ22]. Bias [Che98, WH00, FPNK22, RKL^{+}18]. Bias-Reduced [Che98]. Bias-Variance [WH00]. biased [BMX22]. biases [SHSJ23, WGGHvdW21]. Bibliography [Ros01]. Big [MGLB17]. bijection [AXSVL14]. Bilateral [ZW97, FLL^{+}23]. bilinear [NNS^{+}18]. billions [KSG^{+}19]. Bimodal [FRNS05]. bin [MGW10]. binarization [CMH13]. binarized [SJ15a]. Binary [Hei99, JEK98, KD96, LHY14, MW00, RM98, YSX^{+}19, BPBS13, BDHM09, GRGB^{+}13, HQN05, KSF19, MdBJG15, MB11, OEK08, RLB17, SC96, SW05, SM13b, TT16, UWH17, VNNB14, WTBdB15, YZX^{+}17]. binning [LL04]. Binocular [CPC99, WD96, BK16, LS08, LSL^{+}18]. Bio [MNMK16, BC10, BCDH10, BEK18, EK12]. Bio-inspired [MNMK16, BC10, BCDH10, BEK18, EK12]. bioinformatics [BL16]. Biological [SGDP01, FPC^{+}08, MSG10, MNMK16]. Biologically [BL98a, EF14, HL13, MFG10, SM24]. Biologically-inspired [EF14, MFG10]. Biomedical [ABW97, ACW^{+}16, KORC10, SOL16]. Biometric [CR18, DIMT12, HBF09, LFMP13, MKF15, RBC22, WF05]. Biometrics [CGL^{+}21, NPBM22, AZN11, BHF08, HBL^{+}11, HNC05, SR23, YB07, ZBDP15]. bipartite [LCG^{+}24]. BiPR [SXZZ24]. BiPR-RL [SXZZ24]. BIQA [LLJ^{+}23]. Bit [TV99]. Bit-Serial [TV99]. Blackwellized [KLK14]. blended [SSS13]. blending [LJHH07]. Blind [LLJ^{+}23, WPSL18, XTZ^{+}18, AYG23, JHA17, WDC^{+}24, XZQJ21, YHS^{+}20]. blink [FB16, FB18]. blobs [FB12, SI03]. Block [KH15, HMA10, LHG^{+}23, SOL14, SOL16]. block-spin [SOL14, SOL16]. blocks [NHY10]. blood [TDK10]. Blur [LTAA23, LWLT17, SHE17, WZX^{+}24]. blurred [CG09, MNR18]. blurring [JHA17]. BMI [JGM20]. BMVC96 [Ano96a]. Board [Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano17j, Ano17k, Ano18d, ME98a, Ano05f, Ano06g, BL14, GSPL10, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano05e, Ano05g, Ano05h, Ano05i, Ano06c, Ano06d, Ano06e, Ano06f, Ano06a, Ano06b, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h]. Board [Ano11i, Ano11j, Ano11k, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13c, Ano13e, Ano13g, Ano13h, Ano13b, Ano13d, Ano13f, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17a, Ano17b, Ano17c, Ano17d, Ano17l, Ano18a, Ano18b, Ano18c, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i]. Board [Ano19j, Ano19k, Ano19l, Ano19m, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j]. Boards [ME98b]. Bodies [GK98]. body [BCMCB09, CGH08, CFCP11, CPT07, DLC14, DLF06, HUF05, HW07, NZH^{+}23, NESP10, PA06, PT08, PYS03, RRR11, Rem04, SWMM22, UFF06, WPB^{+}14]. Boltzmann [NWJ15]. Bone [MDFS11a, MDFS11b]. Books [Ano97f, Ano98c]. Boolean [GPK99]. boosted [NB20]. Boosting [CWO^{+}11, LL17, RCT14, Sav23, YZL16, YG16]. Bootstrap [KN11, BRP04]. Border [CCP97]. both [YZX^{+}17]. bottleneck [LHG^{+}23, SPK23]. bottom [KMN11, ZWY14]. bottom-up [KMN11, ZWY14]. bottom-up/top-down [KMN11]. Bound [SHKP98, Zha97, Bre03]. Boundaries [WSSD96, BSH13, SKT18, ZYT10, ZS19]. Boundary [ABK16, GJP96, HKS06, KII98, LHHC98, BB16, DCS05, JA16, KA12, LK03, NRJ11, PDK96, RC03, SOD10, YS24, YFDA17, ZSC^{+}23, WPK09]. bounded [ZZ10]. bounding [SJH17]. box [SJH17]. boxing [KFSM17]. Brain [CFYU12, Dav97, GMT00, WPS03, ASFP03, DCS05, LPR^{+}03, MAK^{+}17, MPPP14, ZRL^{+}11, ZU09]. Branch [SHKP98, Bre03, LJJY24, LDT21]. branch-and-bound [Bre03]. branches [SAdB14, WCZ^{+}20]. BRDF [AH08, YSL11]. breakdown [HBH11]. Breaking [TY01]. Breast [KHB01, CSY08, SRP10]. Bregman [RSBA23]. bridging [WM20]. brightness [TLCH05]. British [Ano96a]. Broadband [SM10]. broadcast [DZLH17, MSSS09, WHN08, YJC^{+}09]. broadcasts [DRK03]. bronchoscopy [HSKH07]. browsing [MĆK09]. brushing [MST16]. Bubbles [TK97]. Building [CJC01, DCH12, FMR01, GN98, HB98a, Hen98, LN98, NHTG15, PCJC98, SF95, VV02, Che08, FBS21, HBH10]. Buildings [FKL^{+}98, May99, JRH03, KN04, XHX^{+}19]. built [GKBW14]. Bundle [KSY15, BS05, GA09]. bundles [LAL^{+}10]. Bus [THT^{+}98]. BVS [FHSKP13]. Byzantine [PRG^{+}14].

CAD [CFS98, EFF98, IF95, ZZZ06]. CAD-Based [CFS98, IF95]. Cadastral [OMLL98]. calculation [WGAD14]. Calculations [MMS99]. Calib [RPBK22]. Calibrated [WLD99, PD14, PD17, UWH17]. Calibration [CRC97, DC01, Gui00, PA13, PBSG12, Rob96a, AAB19, BÁRM23, BHSD^{+}13, CXFS06, CF07, CDT11, CZS^{+}20, CSW^{+}24, CP04, CX11, DWW^{+}12, DMW10, FK09, GOF^{+}15, GGO10, HHAE14, HEPH15, JF10, KK09, KGK10, KGFP10, LSKK10, LWLS12, LLWZ21, LP10, MCT10, MM21, NL17, NNT11, QC04, RSL10, RPBK22, SW13, SP06, SJH17, SBMM15, SL16a, SCCP05, TM04, WCF10, YJC^{+}09, ZKRH04]. Call [Ano01k, Ano01l]. calligraphy [WLI08]. Camera [CF07, CRC97, CYP^{+}10, CC00, DT96b, DC01, Gui00, KS95, KK09, Rob96a, SW13, WC99, WCF10, XL98, AMNCM16, ALY^{+}22, BPS10, BCP15, BBH^{+}12, CSGM^{+}24, CKM11, CA10, CGHTK16, CDT11, CSW^{+}24, CKP^{+}19, CSK22, DPRC17, DDLP10, DZJB14, DM24, DSK^{+}20, ES06, FPMK19, GHA10, GB08, Gol05, GGO10, GYF18, GTMR23, HC13c, JSRS08, JB15, JF10, KD10, KSR^{+}12, KGK10, KYYC14, LBK10, LCP13, LM16, LB19, Lhu08, LSL^{+}18, LDD09, LA05, LP10, LDL^{+}19, MFB11, MBD^{+}22, MCT10, MM21, ME18, NNT11, PD17, PYGGLNG17, QC04, RPBK22, RZH17, RCTV12, RBA20, RTM^{+}17, RLC^{+}11, SPC^{+}15, Sav23, SFK24, SP06, SJH17, SST06, SS11, TWQW23, UTB^{+}11, WHL14, YCKA10, YS06, YJC^{+}09, ZY14, ZLL^{+}24, Ziv10]. camera-captured [LDD09]. camera-independent [ME18]. camera-switching [DM24]. Cameras [WLD99, AAB19, AVBK10, BPS10, BCLNG18, BBK15, BYK^{+}18, CMM20, CVP10, CYP^{+}10, CS10, CL17, CKP^{+}19, DWC16, DWW^{+}12, DMW10, GOF^{+}15, HKHE14, HEPH15, KHK10, KBJ^{+}10, LTAA23, LLL^{+}20, LG14, LWLS12, MHSP10, MLH13, MMBG18, NFA04, NL17, PD11, PBSG12, RSL10, ROJX09, SRO^{+}19, SBMM15, SL16a, SCEvdH14, TS17, TM04, UMH16, UWH17, WZ08, ZZ07]. Camouflage [TY01, WF02]. Camouflaged [ZWB^{+}22, LNN^{+}19, TBZ^{+}24, ZSC^{+}23]. Can [FFA^{+}19]. cancellation [CSK22]. candidates [FBK16]. Canonical [DSNN08, LV96]. capability [ZTB20].

Capsule [HCLZ21, TZL^{+}22, MFP^{+}20].

Capsules [BDT23]. captioned [CLA^{+}17, JEF^{+}12]. captioning [Che24, CBTC23, DWLW23, LXW^{+}17, MRdRGC23, MC22, NB20, NLW^{+}17, RYLZ24, WZFL23, WLZM20, YH19]. Capture [MG01, CFCP11, DMSM21, MHK06]. captured [HKHE14, LDD09, PT08]. Capturing [OGB14, WWJ16, WMZY23, RSY22]. Cardiac [RWWH00, GPDR13, TA13, WSKH13, WWJ13b]. cardio [ACC^{+}16]. cardio-metabolic [ACC^{+}16]. caricaturization [SAK15]. Carlo [SOL14, SOL16]. Carrying [HCHD01]. cartilage [LPS^{+}11]. carving [GJMO14]. Cascade [AVBK10, JYY^{+}24, WPQ20, ZP18, DYM14, DZLH17, LLJ^{+}23, SP23a]. Cascaded [ZH18, MDM^{+}21, SJSL21]. Case [MS96b, SU01a, VF96, DBZ07, Got08, VD10]. Cases [Lin02, RLB17, SCCP05]. Cast [SCE04]. Casting [LZ97a, YWH^{+}23]. catadioptric [ALIRT18, BDVK10, BCLNG18, DWW^{+}12, GA09, Lhu08, LNS14, PA13]. Cataract [TAK^{+}22]. categorical [SBM^{+}06]. Categories [SPK^{+}02, FFFP07, FKS10]. Categorization [BKMSR98, MK01, CCSS14, GB10, MDFS11b, MvGS16, TSL14, YZY11, YYH^{+}23, ZG10, vGSV^{+}10]. Categorizing [BKMSR98]. category [GCPF08]. Causal [CBB95, HWT^{+}24, LA05, RDS24]. Celebration [CV13]. cell [CDIF14, KORC10, SH09, KL10, SM10]. Cellular [SC98, Ros10]. Census [PCC13]. Center [OD97, WWW95, Dem05, EK12, GZY23]. center-surround [EK12]. centered [SCL13]. centerness [ZYQ^{+}23]. centerness-aware [ZYQ^{+}23]. Central [DPB00, Bar06, BCLNG18, CMM20, Dem05, DWW^{+}12, PA13, RSL10]. centre [DMW10]. centroids [KŽ12]. Certifiable [GSGJ22, GSGJ24]. cervical [BvdHL^{+}13]. CFA [LPVM13]. Chain [KD96, PKH23]. Chain-Encoded [KD96]. Chains [Cre99]. Challenge [MST00, BVWS21, IZJ^{+}17, PKZP^{+}24, BGPD09]. Challenges [BS99b, dOSJVBS12, BCF06, KK17]. Chamfer [MMS99]. Change [Che00, HKK08, Lai00, Ros02, SB98a, XL98, CCYC12, DWC16, DLBG19, FDC^{+}19, GCS23, HKWC14, KKRK23, MMP09, WMZY23, YCH07, YYH^{+}23]. Changes [BFY00, ASC17, DD11b, WPI^{+}16, XFSC13, YNCO11]. changing [MTVM04]. Channel [YWL^{+}20, ZWW^{+}20, FWZ^{+}24, IJDAB13, JSZY17, LZB^{+}23, LLG^{+}24, LLWX24, NN13, TYH^{+}21, THH^{+}23]. Channel-based [ZWW^{+}20]. channel-wise [LLG^{+}24]. Channels [WHS^{+}24, OGH04, SGS^{+}10]. Character [MLP97, YT13, ZZC^{+}23]. Characteristics [Hod95, IE99, CCR^{+}05, CE17, TG95c]. Characterization [KW99, NSK^{+}97, NS98, SRT01, VMUO95, ADFR18, AQ09, ASFP03, BCM13, BB04, RBA20, TCB^{+}08, Žun03]. characterize [LSP^{+}16]. Characterizing [CZZF97, Kis96b, SC00b, WSY^{+}16]. Checks [KABP98]. chess [BL14, BL14]. chess-board [BL14]. Chessboard [LH99]. chest [BMvT^{+}19]. children [MST16, NKB11]. Chinese [WLI08]. chip [ZZ07]. chorio [KACR^{+}23]. chorio-retinal [KACR^{+}23]. chromatic [GS95, LPVM13, VAWW10]. chrominance [dLAH07]. cine [WWJ13b]. Circle [CL00, PHH^{+}15]. Circles [CC01]. Circuit [ME98b, ME98a]. Circular [CL00, Lil97, Pla96, JSC23, TWQW23]. Cited [Ano07f, Ano08k, Ano12m, Ano13o]. City [SJ01, IZKB12, JBWK11, SOK16, STO17]. city-scale [SOK16]. Class [JLD12, MCPB99, XZX^{+}21, AZP14, AS23, BMvT^{+}19, CKLP09, CP09, GG20, KKCK23, KSF19, MNL^{+}17, PLJS14, Pen03, QCH20, SY20, SGZ21, WLFL21]. class-specific [AZP14]. classes [BVCP21, SG17, TCM18, ZYXZ13]. Classification [ARC14, BBC00, BCC16, CKPV21, DT09, DF02, DH19, GL19, HdVL99, HB98c, KdVL99, LL97b, LCZ^{+}16, MCPB00, SL99, SC98, TS00a, XL98, AMCB20, AMGG^{+}16, BVWS21, BBCF20, BL16, BMvT^{+}19, CSDNR17, CL15, CCPK16, CP21, CNS18, DFJL15, DPCA15, DL10, FFM05, GG20, GHHX04, GBVDC18, HL13, HAT^{+}15, HCC^{+}16, JLZ23, KT15, KHG22, Kim15, KSL^{+}20, KGB17, KORC10, LHL^{+}21, LZ24, LLC11, LCLH18, MRH19, MNL^{+}17, MIP16, MSP^{+}18, NL23, PSR08, PC15, PLKP23, QSX17, RRR11, RLG^{+}14, RSS07, SB13, SYPK13, TRPD20, VPL23, VMP03, WZT13, WHJK23, WLL22b, XZX^{+}21, XQZL23, XMN^{+}15, YSL^{+}14, YG17, ZZL13, ZLL^{+}14, ZDZ^{+}23, ZWN14, dSdSF^{+}12, kCE^{+}18]. Classified [SYF99]. Classifier [GK95, ZGC20, LLC11, PD17]. Classifier-agnostic [ZGC20]. classifiers [DZLH17, HD23, LJJY24]. Classifying [AO04, Ros00a]. clean [CLFH22]. clinical [MBD^{+}22]. Clinically [BCMR16]. cliques [PL08]. Closed [ASS97, KPPK09, ADC19, BGK95, Eva06, NRJ11, WSFTK18]. closed-form [ADC19, WSFTK18]. Closed-world [KPPK09]. closer [LJJY24]. Closest [GSK02]. closing [MTP21]. closure [WWLV11]. cloth [SP23a, UK12b]. clothing [SP23a, WPB^{+}14]. cloud [CALO20, FBZP15, HDZR24, HWL^{+}22, JSC23, KKSC23, Lhu18, LZZ22, MPST08, VBT19]. clouds [ANHGS17, BSALF18, CLK09, CACB17, TLFM23, THH^{+}23, ZZK^{+}20, ZSK^{+}23, ZMM^{+}22]. cloudy [WSJ15]. clues [GSV05, SL16b]. Cluster [FSG22, MJ17, LWLC22, LZLP10, TWW14, WHY^{+}23]. Cluster-based [MJ17]. Cluster-Contrastive [WHY^{+}23]. clustered [TSD17]. Clustering [AW98, LJZ18, PF99, Pha01, TB99, WF02, YYL98, ZWL16, AS09, BDFG17, CSY08, CFYU12, CO16, CD13, DBT^{+}17, FLHK08, GXC23, HHG^{+}20, HF11, KBN12, Kim17, MTG07, MMK04, MdOBA19, Pha17, RM03, TVC09, VAWW10, WSSS13, WWWF23, XXCR15, ZLZH17]. clustering-based [VAWW10]. clusters [SH09, SBPF17]. cluttered [AM04, Ano06h, BAPXH16, BPLT15, GKK05, LBNS09, WRKP05]. CMGNet [RYLZ24]. CNN [AXJE21, CNS18, HSK23, LXW^{+}24, MCM^{+}17, MAK^{+}17, PLKP23, PBPD^{+}17, SBN^{+}24, YSY^{+}18, ZXC^{+}20, ZSDK19, ZTB20, ZS19]. CNN-based [CNS18, HSK23, ZTB20]. CNNs [BCC^{+}18, CICN22, CYD^{+}22, DKG22, KHH^{+}22]. CNP [PLYW21]. Co [DYM14, PA10b, SVA^{+}22, BCC16, GZX^{+}23, LPVM13, WZW17]. Co-occurrence [PA10b, LPVM13]. co-salient [GZX^{+}23]. Co-segmentation [SVA^{+}22, WZW17]. Co-trained [DYM14]. co-training [BCC16]. Coalitional [DPT07]. Coarse [RT14, SY10, TB99, JRH24, ML13, WHY^{+}23, ZIT^{+}13]. Coarse-to-fine [RT14, SY10, JRH24, ML13, WHY^{+}23, ZIT^{+}13]. cocycles [GDIIHK11]. code [LHY14, SGS^{+}10, ST20, YZX^{+}22]. codebook [HSBS16, ZJZY16]. codebook-based [ZJZY16]. codebooks [vGSV^{+}10]. Codes [BBC00]. codeword [ATC^{+}13]. codices [PRG^{+}14]. Coding [YB01, BGE^{+}17, BRSSAL11, CTWH15, KYM13, LTCT14, LLL15b, TD04, TYDH18, VBA19, XWC^{+}23a, XMT22, ZLL^{+}14]. coefficient [FSI21]. cognition [WZQ^{+}24]. Cognitive [BBH^{+}12, Ham05, WWH07]. coherence [MPF07]. coherency [FWG18, RCLS19, VO24]. Coherent [Che24, SCALFG^{+}18, KBD^{+}12]. cohesive [LWLC22]. cohomology [GDIIHK11]. Collaboration [LLLW23]. Collaborative [BB15b, RYLZ24, WZFL23, ZWN14, BIMD23, NAS^{+}17, PYS03, TBZ^{+}24]. collapse [YZX^{+}22]. collection [MSG10]. collections [WL15]. Collective [KS12]. Collective-reward [KS12]. Collinear [Cre99, DT96a, UTB^{+}11]. collinearity [MCCRAC20]. Collineation [CDH99]. collision [YR06]. colony [CKF18]. Color [APV99, BFF97, BK07, BD02, GFS04, GB97, Hen98, IP98, LL97a, LGL15, LPVM13, LPV07, MVP06, MTG07, MKK02, MLJC20, RPTB01, Sap97, SG11, SGK00, VMP03, ÅS17b, AQ09, ASVO12, ALIRT18, BL04, BDFG17, BH12, Dre96, GTP18, GHML17, HKM22, HC13a, HWW06, HSJS10, HKK08, HLKK19, JWG04, JOvW^{+}05, KGU10, LLR10, LL04, LEB07, LMC09, LJZ18, LL08, LN10, MWF07, MYZ^{+}24, MN06, MGPJ11, MGPF08, NN04, OSY18, Pen15, PA10b, PBG04, PS12, QAB^{+}11, SCE04, SCALFG^{+}18, SF07, SKU^{+}09, SAC09, TLEF06, UJ22, VSP06, YZ06, YCL07, ZZ07, ZT09, ZCF13, ZHZ17, PA10b, SCALFG^{+}18]. color-based [BL04, BH12, LN10]. color-plane [ZHZ17]. colored [DR04, OSY18]. colorimetric [BÁRM23]. colorization [BT17, CWL^{+}23, MLJC20]. colors [HGS08]. colour [Ang07, BG09, CT10, CT12, DCFM07, GE08, GD19, HEPH15, Hei04, PKD07, VBS^{+}04]. column [TH06]. column-space [TH06]. columnar [HBZ^{+}24]. combination [GRMH19, YFF^{+}23, KL11]. Combinational [GFLW24]. Combinatorial [KMT11, NKPT13, CRCM16, DSdlH^{+}11, WDN^{+}12]. Combine [Pen15, WLZM20, SM22, YWL^{+}22]. combine-and-conquer [YWL^{+}22]. Combined [BYR17, HYJ11, GHML17, LV11, SKSR08, VRKL13]. Combining [BK16, CKC14, GCPF08, Hei04, Mah16, MTP21, QKH^{+}12, TFD07, TLEF06, ZWY14, GFL^{+}11, GJ10, HDF12, LvdHK^{+}15, LGL15, MMK04, XP11]. commercials [GS06]. committee [MPM16]. common [SRS11, XGT^{+}22]. communicating [UM05]. Communication [FKL^{+}16a, CC16]. Commute [DDWZ12]. Comp [OBH04]. Compact [BRPC17, HB98c, LVS20, CM16, NNS^{+}18, SGS^{+}10, ZZK^{+}20, vGSV^{+}10]. CompactNets [LVS20]. Comparability [Bre01]. Comparative [Che00, LCZ^{+}01, AVGASAP15, BZ14, BSBW14, Cha21, HS06, JM09b, KLY21, LMRMJ08, MRdRGC23, OH05, OTAH20, PSE^{+}11, SCD11, SYPK13, TPD^{+}16]. compare [ZK17]. Comparing [CDJM14, GJ10, Sha11, vGSV^{+}10, CU11, OJRT08, TN05]. Comparison [HSSB98, KLKF20, KYM13, RFC97, SOL14, SGB01, Ste01, FCM20, LLG^{+}14, LLL^{+}15a, MSR07, PBSG12, She16, VTRC14, YARL^{+}20]. compensation [LMP^{+}19]. competition [MMV06]. Complementary [LL97b, AH24, LYX^{+}24, LL08]. Complete [BNG02, DG01, DY98, TG95b, ZZG^{+}24, KM03]. completeness [FB18]. completing [ZMM^{+}22]. Completion [WH96, WZWT99, AKE23, BF05, GWFF22, KKRK23, LA11, LDH^{+}14, ZA22]. Complex [CM95, Jon97, LM99b, MS97b, SP97a, VKP98, BKPS15, BP09, ÇÖD08, CT10, DETE17, FL09, HY11, Hu11, HML15, KV06, KN04, LL12, LCL^{+}17, LCG21, MJ11, MiMO^{+}16, SZ07, SM17, TN07, VB16, XYW11, YR06]. complex-cue [LL12]. complexes [CDIF14, Cou13]. complexity [GMF14, LT05, SJB20, ZZH23]. Component [BZ14, Jon99, BRSSAL11, CCL04, CE17, DB03, HHWP03, HQN05, Nic95, NGR24, Ros08, SVSM15, SHS03, WLMG08]. component-based [HHWP03]. component-labeling [CCL04]. Components [CCS01, AO16, AHDM10, DBB13, WPZ^{+}18]. Composed [LER95, LL12, WB97]. Composite [HZLM11, SL99, LHLZ23, SOJ17]. Compositing [KW99]. composition [CZ14, LRZ^{+}19, ZSG^{+}20]. Compositional [DFH^{+}22, LSW18, LVS20, LZW^{+}24, TLB^{+}15]. compositions [RL13, SZGK24, TLB^{+}15]. compound [BAM16]. Comprehensive [Cha21, PWWQ16, ASVO12, JKW^{+}21, SV14, TZLT21, TPT15, ZCLX20]. Compressed [Spi98, WHL^{+}21, ZSK^{+}23]. Compression [GSK02, JEK98, KDRC98, NK00, BT17, CWC^{+}20, HDL^{+}20, HBL^{+}17, SBS04, TVLS08, TAC23, WLZW04, WHS^{+}24, YWMS08]. Comput [AK11, Ano06h, BB15a, MFSB23b, MBMC11, PZ09]. Computation [BM00, BM02, CM99a, CCP97, CH99, LHKC97, MKY01, Neg96, OD99, SA96, DRAB08, FKV^{+}11, FBK15, Kle13, MSI10, MN06, OH05, TLCH05, XSD12, Ano95e]. Computational [LZ97a, MJS97, SMK02, SAK15, TVY^{+}18, FFY^{+}04, FFL14, KTP08, Pec07, SGA12, VBS^{+}04]. Computer [Ano95a, Ano98d, Ano15n, BY98, BS99b, CFS98, DRDKE13, FKL^{+}16b, FKL^{+}16a, FHP01, GKL^{+}17, HTEB11, HSKH07, LB14, LHKC97, LMT^{+}17, MP09a, MST00, MG01, MTH^{+}17, MT00, Ros95, Ros96, Ros97, Ros98, Ros99a, Ros00a, Ros00b, Ros01, TGM^{+}17, WKI^{+}16, ZXK02, Ano05j, BK15, GRMH19, HBH11, JS07, JNLG15, KPKH07, KMT11, LBK10, MdBJG15, MNMK16, NLM05, PZ08, PZ09, PYS03, Rei16, Sah05, SBB10, SBD22, SVA^{+}22, SFWG08, TCB^{+}08, WKP13, WZ23, ZSSF16, LLE^{+}09, STLH08, BPQ15]. Computer-based [HSKH07]. Computing [Ano98d, AM97, BY98, DT96a, FK00, GK98, LH99, NWP97, TG95c, WZWT99, CKK^{+}12, FYH11, SRS11]. Concept [WTBdB15, HS14, Kim15, KYM13, KM03, THL13, USKB10, WSY^{+}16]. concepts [LDC^{+}13]. Conciliating [IJDAB13]. Concurrent [CTE95]. Condition [RM02]. Conditional [BCC^{+}21, SKM06, CL18, GFL^{+}19, LZZ^{+}23a, MLB^{+}18, RB19, PV13]. Conditional-VRNN [BCC^{+}21]. Conditioning [LG24]. Conditions [OD01, CSV^{+}16, Mal21, OK04, SPK14, ZJ05]. Conference [Ano95a, Ano96d, Rei16, Ano96a]. Confidence [Neg96, KN11, PTM20, PMC13, SvdMH15]. Configuration [OD01]. Configurations [MRF96, TZM98]. confocal [KGK10]. Conforming [Spe97]. COnfusion [RLB17]. Conic [BF14]. conical [LNS14]. Conics [QV98, BA06, KGK10]. Connected [Hei99, Jon99, PC15, SUO00, SU01a, AHDM10, HQN05, HQW^{+}12, Nic95, SH09, SHS03, ZUS06]. connected-component [HQN05, SHS03]. Connectedness [SU01b, CUSZ07, CU10a, CU10b, CU11, MVP06]. connecting [GBL08]. Connection [ZWB^{+}22]. connections [KHH^{+}22]. connectivities [BNG05]. Connectivity [BDHM09, BNG02, WB97, BNG03, LZ24]. Connectivity-preserving [BDHM09]. conquer [BPC^{+}17, YWL^{+}22]. cons [Bor19, Bor22]. Consecutive [Muk97]. Consensus [CM97, LZ97b, MGS15, CLCO19]. Consideration [SKOS95]. Considering [OD02]. Consistency [OMLL98, SF97, CL18, CBT^{+}04, CK09, Lhu18, MM06, PD14, PS22, WCZ23, ZZC^{+}23, ZYZ24, ZFG^{+}22]. consistency-based [CL18]. consistent [CPC08, JLD12, SXZZ24, TY05, UK12b]. Constancy [BFF97, BJ97, CT12, LGL15, SAC09]. Constant [MS96b, SOL14]. constellation [GLM17]. Constituent [LXS^{+}23]. Constrained [IP98, Ols99, ZCL99, CKC14, LPR^{+}03, MFG10, SOJ17, SMD^{+}08, TLP^{+}17, TLY^{+}16, UO16, WYC15, WDC^{+}24, WWJ13b, YZT^{+}13, ZLL^{+}14]. Constraint [BZ99, Jon97, BHMB10, CJYW23, GXC23, MZC^{+}05, PL08, QCL^{+}23, SDK22, XWDL23].

Constraint-Satisfaction [BZ99].

Constraints [DM01, FL96, FB97, Zha97, BF14, CLZY15, FF09, FK09, GYF18, HLKK19, IJDAB13, Lhu18, NNT11, NDO09, OCVV04, RC03, TR09, WDB12]. Constructing [BNG05, WCZ^{+}20, Eva06, LH95]. construction [CACB17, FT23, Sch06, ZZC^{+}13]. consumer [LDL^{+}19]. consumer-grade [LDL^{+}19]. contact [BHBF10, NLM05]. Content [BZS08, BS99a, DCCL99, DRK03, GH08, GWCO11, JEK98, MBKB02, PBQ99, PA10b, SLST99, SBK^{+}99, SPK^{+}02, AO04, CGD^{+}23, CBB19, FT23, Hei04, ILRB04, KMBH09, LJZ18, LL12, MSG10, Pen03, TPNP15, TL16, WZ04, XG08b, YJC^{+}09]. content-adaptive [TL16]. Content-Based [BS99a, DCCL99, JEK98, MBKB02, PBQ99, SLST99, SBK^{+}99, SPK^{+}02, DRK03, GH08, PA10b, Pen03]. Context [AvdWDM18, GB10, GDR04, ODT17, WZ23, BÁRM23, CL08, DLC14, FFL14, HMF10, JYTK11, KK07, LWZC14, LKZ20, LXFM16, LZZ^{+}23b, MT16, PSE^{+}11, PL10, WTZ24, WMBY12, YZY11, ZLLP21]. context-aware [LZZ^{+}23b]. Context-based [ODT17, MT16]. Context-dependent [GDR04]. context-free [ZLLP21]. contexts [FYH11, LXW^{+}24]. Contextual [MC22, DFP^{+}13, SKM06, ZZG^{+}24]. Continual [ZSK^{+}23, DNG^{+}24, SHS^{+}23]. Continuity [CWL^{+}23]. Continuous [AM97, DPRC17, GGR01, HAT^{+}15, KFN15, ZZL13, CGR13, Eva06, PV13, TP14, TMN06, Zac18]. continuous-discrete [PV13]. Contour [AM00, ASZ99a, BM98, CM99a, CS98, Dem96, DY98, LL99, LAL^{+}10, Pet99, BN15, BB03, CCL04, DT09, DS07, GBY21, GTP18, Mig12, PDTE06, WO10, YZX^{+}17, YFF^{+}23, YLA09]. contour-based [DS07]. Contours [DM01, JDP97, KMB97, KD96, Pla96, Sau99, SC00b, VKP98, ZM96, CT13, Mil09, MBMC11, MPPP14, SECS15, SZ07, VRKL13, WYC15, WWJ13a, XAB07, ZZS^{+}23]. Contrast [RRL20, ZCL99, CZL^{+}24, Che24, LGD16, MYZ^{+}24]. contrast-invariant [LGD16]. contrasting [OZT19]. Contrastive [WLXL24, BFMW23, DGRS22, LBP23, LTL^{+}23, YXW^{+}24, ZdCR^{+}24, ZSL^{+}24, ZZHZ23, WHY^{+}23]. contributed [IZKB12]. contribution [JOvW^{+}05]. Control [DCTO97, MGMS01, BBH^{+}12, GM19, Ham05, JZWD16, LWYF24, TM07]. controllable [CBTC23]. controlled [BBB96]. Controlling [WGGHvdW21, WH00]. conventional [BPS10]. Convergence [BVVMMS15, CRC97, GMT00, SK98, YYL96, VWMZ15]. convergent [Bar05, CLL14b, WML21]. conversational [VMC^{+}16]. CONVERSE [EDX16]. Conversions [ÜE01]. Convex [BBH14, GK98, Rob96b, AJ23, AM15, DBB13, HZLM11, LQQS21, MPPP14, QDLB17, SWMM22]. Convexity [Kis96b, LL99, MMS97, TY01, BMJF^{+}17, RM06]. Convexity-Based [TY01]. Convolution [FCM20, ZDZ^{+}23, DWC^{+}24, FLS^{+}14, GY19, GWFF22, KSL^{+}20, LLG^{+}24, MSM17]. Convolutional [FWL^{+}20, LLP16, VPL23, AM17, AAL22, BRPC17, BMX22, CBB19, CWW^{+}22, CJWW22, EXP^{+}20, EH21, FWZ^{+}24, GL19, GGGROE^{+}17, JSC23, KUHY18, LCLH18, NL23, RKKK22, SFF^{+}18, SCC17, SKS^{+}22, TKK24, UIK22, VBA19, VBT19, WWG^{+}18, ZK17]. convolves [LGG^{+}18]. Cooperating [CA97]. cooperation [FOCSB^{+}20]. Cooperative [DC00a, LYA13, MLH13, KON^{+}17, UM05, ZKRH04]. Coordinate [Big97, ÜE01, DSK^{+}20]. coordinated [PKK^{+}09]. coordinates [JF10]. Coordinating [WWH07]. coordination [YCKA10]. Coping [YYH^{+}23]. Coplanar [CRC97, QV98, Bar05, ODD96]. coplanarities [FK09]. CORE [RLB17, KL10]. Cores [MPPG98, PEFM98].

Corneal [GAD01, ZMCA05]. Corner [BY08, FT98, Ros99b, MM06]. Corners [Dem96]. COROLA [SZ16]. coronary [EX17, LAFLB16]. corpora [RLMK15]. Correct [LZ97b, DL10, PXTZ14]. correcting [LLS^{+}23, MC20]. Correction [SKU^{+}09, ABK^{+}18, Che08, FDW21, GML^{+}21, MUS06, SCGAF^{+}17]. corrections [BCP15]. correlated [LZmC^{+}17, MLB^{+}18]. Correlation [KC99, AVGASAP15, AS09, AT17, BDFG17, DFLS23, DZLH17, HTNN18, JVD^{+}20, LRW08, LY06, LZL^{+}17, MCF10, TTXT21, TLH22, WMZY23, ZLL^{+}14, ZWL16]. correlation-guided [DFLS23]. correlogram [ZT09]. Correspondence [Chu02, Jon97, Jur99, KHB01, SA96, CALO20, GKBW14, KY19, LZLP10, LH03, MEYD11, ME18, PMW05, SAS12, TVE^{+}16, XJK12]. Correspondences [CA97, CH99, SBZ97, Tay00, BN15, BF14, CDT11, KSHE20, MGS15, PW06, PZC17, TKV16, TSD17, ZN08]. Corresponding [WB01, Sha11]. corridor [NPM^{+}16]. Corrigendum [AK11, BB15a, BM02, MFSB23b, MBMC11, PZ09]. Cosegmentation [MCL16, CW15]. Cosine [LL08]. cosmetic [BHBF10]. Cost [FK00, SLK23, FSI21, KHH^{+}12, MSI10, MEYD11, ZWZZ18]. Cost-free [SLK23]. cost-sensitive [ZWZZ18]. could [MRdRGC23]. count [HBB^{+}12]. Counterparts [FKW98]. Counting [Mil99, LKZ20, RDSF15]. counts [KRJ^{+}08]. Coupled [CBM01, FDC^{+}19, YS09, AH24, GFW13, MML^{+}16a, MLJC20, SAC^{+}12, TRG^{+}13, WB16, ZLFH23]. coupled-layer [MML^{+}16a]. Coupling [YSL^{+}14, TMN06]. COV2 [Ano07a, Ano07b, Ano07c, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f]. Covariance [FBZP15, IH15, KRS14]. covariances [YO11]. covariant [TBFJ15]. covariates [SBIK16]. Cover [Ano17j, Ano17k, Ano17l, Ano18k, CCPK16]. Coverage [TG95b, ES06, PW23]. coverage-based [PW23]. Covering [CM99a]. covers [Eva06]. CPNet [CWL^{+}23]. CPRNC [WHS^{+}24]. Crease [SLS01]. Creaseness [LLSV00]. created [SYPK13]. creation [CSZ^{+}15, NNN^{+}22]. Crest [MAM97]. CRF [GG20]. CRFs [YHN11]. Criteria [IW97, Kim04]. criterion [GBHS06]. critical [GB10, OBTMT15]. Critique [Oli00, Oli01]. Cross [DWL19, DWLW23, HBKG22, HEPH15, KIS17, LHG^{+}23, LZC^{+}20, LF98, PWL^{+}23, PV14, QZY^{+}24, RB19, SP23a, WZQ^{+}23, YNZ^{+}19, AWK04, BPCT22, CLDP23, DZQ24, EX17, JLM22, KK15, KACR^{+}23, LD24, LCL^{+}17, MLZRK24, MCF10, PS22, SIRS21, SGZ21, TMS20, VJ17, WLSO23, WHN08, YC05, ZRZ^{+}24, ZFG^{+}22, ZJL23]. cross-attention [DZQ24]. Cross-calibration [HEPH15]. cross-class [SGZ21]. cross-correlation [MCF10]. Cross-domain [DWL19, DWLW23, SP23a, WZQ^{+}23, BPCT22, LD24, SIRS21, TMS20, ZFG^{+}22, ZJL23]. cross-entropy [JLM22]. cross-information [ZRZ^{+}24]. cross-lingual [WHN08]. cross-localisation [KACR^{+}23]. Cross-modal [HBKG22, PWL^{+}23, PV14, LCL^{+}17, PS22, VJ17]. Cross-modality [LHG^{+}23, YNZ^{+}19]. cross-ratio [YC05]. Cross-Ratios [LF98]. cross-referencing [AWK04]. cross-region [WLSO23]. cross-sectional [EX17]. Cross-spectral [LZC^{+}20]. Cross-time [QZY^{+}24]. Cross-view [KIS17, RB19]. crowd [JB15, KB12, LKZ20, PBI16, RDSF15, SCR^{+}17, WX16, ZZP12].

CrowdCam [DMAD17]. crowded

[BSZ^{+}21, HHG^{+}20, SFF^{+}18]. crowding [WHS^{+}24]. crowds [CZZS07, GLOC10]. Crowdsourcing [JRBD^{+}15, TMM16]. Crude [VV02]. CT [HRS02, KKKS24, LAFLB16, MDdMG09, SMD^{+}08]. CT-slice [MDdMG09]. CT-VOS [KKKS24]. CTC [ZLLP21]. CTM [QZY^{+}24]. Cube [CHC11]. cubic [SB05]. cubical [Cou13]. CuDi3D [BAMK18]. Cue [KR99, RJ00, RWWH00, EDB12, JC06, LL12]. Cue-Based [RWWH00]. Cues [LL97b, SLST99, ABN^{+}20, CLZZ13, DKG22, GW07, HLB17, KN03, KSR^{+}12, LGL15, Mig12, MAJ16, NT10, RBC22, ZTH^{+}11]. CUFD [XGT^{+}22]. cultural [BÁRM23, dOSJVBS12]. cumulative [PBD20]. CURL [BCC16]. Current [TGM^{+}17, CH17]. curricular [DDZ^{+}23]. Curriculum [SIRS21, HWW23]. Cursive [AHD98]. curtaining [FMS17]. Curvature [DT97, FW97, Kis96b, LW18, LLSV00, MKY01, OD99, SF97, CLL14b, FB12, MSR07]. Curvature-based [LW18, FB12]. Curve [ASS97, Ols99, SB96b, SdB03]. Curved [KHB01, ST96, VKP98]. Curves [Ano95e, BKD01, FAB97, GLR^{+}99, IW97, LM99a, Mok97, HN95, OBH04, OH04, VKNK14]. Curvilinear [BAMK18, HP96, LCZ09]. cut [CUAT13, DK13, GPDR13, KT08]. cut/max [ZSCP08]. Cutout [KKKS24, SXY^{+}23]. Cuts [KBAS16, CPP^{+}11, CLL17, Mah16, RDT^{+}19, SOL14, XAB07, ZSCP08]. CVIU [BK15, DFJL15, SMHH04]. CVIU DL [PSY^{+}21]. Cycle [CLTW23]. Cycle-GAN [CLTW23]. CycleGAN [SZB^{+}21]. Cycles [CM99a]. cyclic [TAK09]. cylindrical [LCP13].

D [Ano01m, AS08b, ABVC16, BCF06, CLZY15, CL18, CFM^{+}13, FAB97, GSPL10, KTE^{+}17, KHH^{+}22, LEA^{+}10, MBMC11, WHJK23, ACF00, AMNCM16, AXSVL14, AVGASAP15, ACG^{+}09, ÁB13, ALY^{+}22, AS08b, ABVC16, AVC19, ABB^{+}23, AM97, ARARCE11, ACDB12, BN15, BM99, BB16, BBC00, BI10, BI11, BCA98, Bar05, BSALF18, BT05, BR95, BL16, BY12, BW15, Bd96, BZ99, BAMK18, BB24, BCF06, BGK95, BF05, BS00a, BDL^{+}06, BBH14, BSBW14, BMX22, COW98, CICN22, CGH08, CLZY15, CM12, CK11, CL18, CS98, CYNO11, CC11, CPPY21, CZHT15, CLCO13, CLO17, CFM^{+}13, CC96, CP20, CG04, CS00, CPS10, DT96b, Dam08, DSdlH^{+}11, DWB11, Dan97, DWV19, DB03, DF01, DTL17, DMSM21, DAM12, DSY10, DBB13, EK98, EOPS22, ES04, FPC^{+}08, FBF08, FF09, FRL^{+}98, FDMA97, FAB97, FKL^{+}98, FL96, FO18]. D [FCOK24, GM19, GFL^{+}19, GGGROE^{+}17, GSPL10, GHMT09, GKBW14, GSV05, GW07, GLZF23, GC19, Gui98, Gui99, GPC^{+}10, GML^{+}21, GWFF22, GSK02, HFKN97, HB98a, HUI16, HRHZ17, HASS10, HRS02, HR99, HB98b, Hen98, HSS^{+}16, HGSM11, HMB17, HG11, HMF10, HCLZ21, HGB98, IAP^{+}11, IDY^{+}18, JDP97, JC98, JZWD16, JRBD^{+}15, Jok98, JSC23, dOSJVBS12, KMB97, KTE^{+}17, KSL^{+}20, KC22, KSF16, KM03, KMA^{+}00, KMN11, KNO^{+}09, LCT09, LM96, Lau97, LPS^{+}11, LST13, LM16, LÁB15, LAFLB16, Lhu23, LS08, LLG^{+}14, LLL^{+}15a, LDH^{+}15, LQQS21, LSGY24, LDCX24, LSHT02, LS12, LMM22, LSTF12, LK00, LDL^{+}19, Luc01, MS96a, MW00, MSV^{+}20, Mal21, MBD^{+}22, MFJ95, MC09b, MCB13, MMA06, MOB14, MWTN04, MCT10, Mil09, MKY01, MB95, MJPS16, MIP16, NSK^{+}97, NG98b, NT10, Neg12, NFA04, NKPT13, NL96, NDO09, NSEA13, OG98]. D [OMBH06, ODT17, OJRT08, OCVV04, PSR08, PYGGLNG17, PMW05, PMCN22, Pud98, QL96, RAH97, RB18, RZH17, RWWH00, Rem04, RXDS22, RZZ23, RT14, SC96, SECS15, STC^{+}16, SCD11, SBIK16, ST96, SCALFG^{+}18, STV09, SS17a, SSHP17, SM06, SN99, Shi99, SKU^{+}09, SBN^{+}24, ST10, SKVS13, SJH17, SYL^{+}24, SPQ^{+}17, SBMM15, SB00, Ste01, SWS11, SRHC13, SKBS13, SWMM22, SS11, SB02, TLFM23, TGQ23, THH^{+}23, TB99, TPT15, TPT17, TPDP20, TS17, TN05, TN08, TML00, TH04, THL03, UK12b, UFF06, VBVB19, VV02, VBT19, VPP^{+}23, VAC16, VKP98, WCZ23, WCZ02, WPS03, WPI^{+}16, WLO^{+}18, WTZ^{+}21, WHJK23, WWLV11, XOF05, XP11, YB07, YQL^{+}23, YHR^{+}05, YZX^{+}17, YHW^{+}23, YT99, YC98, YGC15, YJC^{+}09, YLX^{+}18, YZL^{+}21, YARL^{+}20, ZW97, ZP11, ZZK^{+}20, ZSK^{+}23, ZSCP08, ZZJS18, ZZC^{+}13, ZT15, ZC19, ZCLX20]. D [ZLHJ18, ZDZ^{+}23, ZCW24, ZH04, Ziv10]. D- [FAB97]. D-based [GSPL10]. D-image [LS12]. D-range [LS12]. D-Space [HR99]. D-tracking-based [AVC19]. D/ [ABVC16, CLZY15, CFM^{+}13]. DAAL [ZTGL18]. DAGs [XYZ16]. daily [BKPS15, VCDS^{+}17]. dandelion [LYG07]. dark [LZC^{+}20, TYH^{+}21, LC19]. Dashed [JvdBS99]. Data [BCA98, BL98a, BZ99, BS00a, BS00b, CKB96, GSK02, Jac01, LR02, MAM97, MGLB17, NWP97, RAH97, RF02, SDK^{+}24, SB00, SM97, WLZW04, WALL00, ZOMK00, AM06, BBSD15, BCC^{+}18, BC10, BR12, BYN^{+}04, BSBW14, BJS14, BG18, CLZY15, CH06, CP21, CBT^{+}04, CD10, CP09, CC96, Cre08, DWV19, FLHK08, GLOC10, GYWZ23, HRHZ17, HF11, JBC08, JRBD^{+}15, Kim04, KSHE20, KACR^{+}23, LRFF24, LY13, LSCK15, LZZ22, LSGY24, LPR^{+}03, MRH19, MSR07, MC09b, MFP^{+}20, NY14, NWJ15, ÖÜ20, Pat13, PSB^{+}24, PPT06, PKC^{+}18, QT10, RH06, RKG03, RBC22, STHBH18, SY10, SPT^{+}18, SRB21, Sha11, SKVS13, SRHC13, TG11, TST14, TFL^{+}09, TN05, TN08, TB23, TZY08, VBT19, WS08, WZW17, WLL^{+}22a, WNH05, WB16, YWMS08, YW07, YW16, ZZZ06, ZZ10, ZCWH23]. Data- [CKB96, SM97]. data-driven [BBSD15, TZY08]. data-efficient [ZCWH23]. Database [BS99a, SPK^{+}02, ABVC16, CM21, DR04, MTAA11, YAK^{+}08]. Databases [ADDK99, KAES99, KR98, MK01, SBK^{+}99, GDR04, PA10b, PS15]. dataset [ABB^{+}23, CLFH22, CYG16, KLKF20, LYW^{+}24, LC19, LZL^{+}22, PKZP^{+}24, RFF23, SCR^{+}17, TS24, WZY13, YST21]. Datasets [KK17, BSH22, CCFC13, EDX16, FPNK22, OB14, TZLT21, WTW^{+}17, YGJ^{+}20, YST21]. dating [HSBS16]. day [ASC17]. days [WSJ15]. DCNNs [MTP21]. dead [Gre04]. Dealing [TO99]. debiased [CAO^{+}23]. Deblurring [MRW^{+}97, WZJ^{+}21, HWZ^{+}23, KKP24, KLY21, LDT21, QSXS23, SRM20, WPSL18, XZQJ21]. Decade [Boo97]. December [Ano19a, Ano20c, Ano21b, Ano22c, Ano23c]. decentralized [CÇ15, HML15, HW07]. deception [SL16b]. Deciduous [HdVL99]. Decision [RM98, CKL18, HPvB^{+}10, PTK24]. decoder [DWC^{+}24, XGT^{+}22, ZRZ^{+}24]. decomposable [CKK^{+}12, LHG^{+}23]. Decomposition [LL99, MK01, SW05, ARFF18, AM15, BLJ^{+}23, BLKG21, BFR13, CW15, DKG22, DAM12, HKM22, HML15, JRH24, KRBSV17, LRZ^{+}19, LQQS21, LWFL24, PAK19, RDM^{+}11, SH09, SKS11, UFK20, UIK22, XYW^{+}08, XGT^{+}22, YZL^{+}21, ZLL^{+}14, ZLL^{+}24, ED16]. decomposition-composition [LRZ^{+}19]. decomposition-like [DAM12]. decompositions [EOPS22]. deconvolution [JHA17, LEE^{+}18, ZLL^{+}24]. Decoupled [LLS21b, ANHGS17]. decoupling [BDVK10]. dedicated [YG17]. Deducing [RBC22]. Deep [ALY^{+}22, AYG23, BBCF20, CLCO19, CGL^{+}21, DAZ^{+}17, GFL^{+}19, GKL^{+}17, HH19, LWFL24, MSF^{+}17, MAK^{+}17, NNS^{+}18, NNN^{+}22, NZH^{+}23, RSBA23, SFF^{+}18, SRB21, SWYP00, ST20, SHS^{+}23, TPDP20, WTZ^{+}21, ZK17, ZTGL18, AM17, AXJE21, ABLL19, BCC^{+}18, BCHR24, CTH20, CWW^{+}22, CKL18, DFSC20, EOPS22, FZ20, FSI21, GG20, GLG22, GYW^{+}22, GZ19, HBL^{+}17, HZK19, HSHA20, JCLZ21, JRH24, KdRM^{+}23, KDSF20, LRZ^{+}19, LZZ^{+}21, LLL15b, LLWZ21, LXW^{+}24, MSV^{+}20, MMLC23, MFP^{+}20, MP20, NL23, OTAH20, PKC^{+}18, PLKP23, RCLS19, PBPD^{+}17, SB18, SJB20, SP23b, SBN^{+}24, SHSJ23, SXZZ24, TS24, TAC21, TAC23, VGLP17, WLO^{+}18, WHL^{+}20, WLZ23, WWG^{+}18, XYRS17, XZQJ21, XMT22, YGJ^{+}20, ZWZZ18, CKPV21]. Deep-anomaly [SFF^{+}18]. Deep-STaR [CKPV21]. deepfakes [NNN^{+}22, LSQL24]. deeply [VBVB19]. deeply-initialized [VBVB19]. DeepShoe [ZSDK19]. defend [LWH^{+}23]. Defending [JPN^{+}22]. defense [SLK23]. defined [TWS06]. Defining [CU10b]. Definition [ACF00, SU01a, DBF04, KMBH09, Dam08]. Defocus [ZD01]. Defocused [RC97]. defogging [NGR24]. Deformable [BCA98, CYES00, Dav97, DJG01, FB97, GSP02, LT05, NFSK97, Pet99, RAH97, TI01, TC11, WRH97, BVVMMS15, BM15, BPB13, CMD06, HW06, ML13, MSF^{+}12, RB18, SB18, SI03, SRHC13, TLY^{+}16, WB12, ZZC^{+}13]. Deformation [KMB97, RW97, FPC^{+}08, LPR^{+}03, Mar07, MWTN04, SY10, SKH08, WMZY23, XFP^{+}16]. Deformations [FT98, LHH97, NMP97, ASFP03]. Deformed [Nis97]. Degenerate [TZM98, MC09b]. Degradation [BHBF10, HWZ^{+}23]. degraded [PS12]. degrades [HBF09]. degree [Sha11]. degrees [LWLS12]. degrouping [ABD11]. dehazed [CYD^{+}22]. Dehazing [FSI21, ECC18, GGP23, JSZY17, LZmC^{+}17, SZB^{+}21, TYH^{+}21, YZX^{+}20, ZWW^{+}20]. delay [NSEA13]. Deletable [Che98]. Delineate [AM00]. delineated [Ano06h, GKK05]. Delineation [SU01a, LCZ09]. dementia [HPvB^{+}10]. demodulation [WB11]. demonstration [KRK11]. demosaicing [dLAH07]. demosaicking [ZZ07]. denoiser [ZXC^{+}20]. denoising [BLJ^{+}23, CWW^{+}22, FZ20, HSJS10, LEE^{+}18, LZmC^{+}17, MGPJ11, PYWZ17, SZW^{+}21, TGQ23, XTZ^{+}18, ZD18, ZZHZ23, ZLHJ18]. Dense [FMR01, LSC08, SPK23, TGQ23, XS98, BG16, CM16, CRCM16, FBS21, FLL^{+}23, HF11, IZKB12, LNM^{+}21, MPK24, SHW24, WNH05, XSL^{+}24, ZCW^{+}23]. DenseNet [ZLLP21]. DenseNet-CTC [ZLLP21]. densities [MIP16]. Density [BH99, PV97, YKA01, JHV19, LCZ09, SPK14, SRP10, WHM^{+}09, ZZP12]. Departure [Lee02, LY05]. Departures [SC00b]. dependencies [CHC11, DFP23]. dependency [XYW11]. Dependent [OYTY98, GDR04, TAC21]. Depth [CP04, MNE00, MMBG18, QGZ^{+}23, RC97, ZD01, AAMO16, ALM23, ASF14, BL20, BZP^{+}23, GKGM20, HCC^{+}16, JC06, KK15, KFSM17, KIS17, KKSC23, KLKF20, KY19, LTAA23, LYKY19, LDL^{+}19, MWL^{+}24, PY19, PCR^{+}04, RF23, RA15, SB96a, SSL^{+}12, SRB21, SRO^{+}19, SKBS13, WNH05, ZT15, ZSL^{+}16, ZTGL18, ZZG^{+}24, HBKG22]. depth-encoded [SKBS13]. deraining [DDZ^{+}23, LZZ^{+}23b]. derivatives [MB95]. derived [SCMP14]. Deriving [SYK96]. dermoscopy [BCMR16]. describing [SJ15a]. Description [AYB^{+}18, Ant98, CM95, DG01, KW00, LN98, LL97b, MBHRC21, ASVO12, BGK95, CH09, CNC03, FMGA^{+}12, KN04, STD14, TPNP15, XHJF12, YJA96]. descriptions [Nis96]. Descriptor [DUC97, TRPD20, DLV15, HC13b, HKWC14, KŽ12, LLS21a, TG11, TWS06, UWH17, YWY^{+}16, ZT15, kCE^{+}18]. Descriptor-Based [DUC97]. Descriptors [ANM98, GAD01, AVBK10, ADGB16, BRPC17, FBZP15, HOH^{+}07, KSF16, LL12, MTV17, PZX13, PG13, PS12, RG16, RLB17, SW17, TABK17, ZZJS18, ZZL13, ZCLX20, dSM14, SGMC15]. Design [BS00a, SBB10]. Designing [DUC97, PK18]. designs [LFMP13]. despeckling [DRTT24]. destinations [PHY^{+}11]. Detail [SZW^{+}21, ZKLZ23, LSH19, WGZL20]. detect [AVBK10, SB18, ÜB05]. detected [HBL^{+}11]. Detecting [BL20, BBK14, CHP^{+}11, CC01, DT96a, DMAD17, GWT09, IW97, LB05, MOT17, ST96, SRHC13, SM99, VMC^{+}16, WZ04, XYRS17, ZhZFL22, ZYT10, BLH16, CCF17, HRC09, RL13, SG17, WK21]. Detection [BB04, BCG95, BS00a, BP09, Che98, CBM01, Che00, CYES00, CMG16, DGH98, FD99, FMR01, GMZ^{+}22, GS95, GJP96, HCHD01, HRS02, HL01, JB99, KMA^{+}00, Lee02, LB98, LL97a, LN98, LD98, Loh10, MLB^{+}18, MCAF21, MGK00, NS98, Ols99, PCJC98, RY98, Ros02, Spi98, TW98, TZM98, VMUO95, XL98, YKA01, YW99, ZYQ^{+}23, ABN^{+}20, AZSVK05, ALY^{+}22, ATG15, ALK^{+}09, AHDM10, ABK16, AvdWDM18, BVWS21, BIG^{+}23, BL14, BT05, BDS12, BBC^{+}07, BL09, BM15, BFD22, BPCT22, BAMK18, BDFG17, BWG17, BJS14, CGD^{+}23, CSY08, CVP10, CM16, CGHTK16, CWO^{+}11, CZS^{+}20, CYD^{+}22, CCYC12, CYG16, CZZS07, DLS^{+}09, DK13, DETE17, DZL07, DWC16, DFJL15, DLBG19, DFP23, DLF06, DNG^{+}24, DD11b, DWC^{+}24, DZLH17, EB13, ED16, FWL^{+}20, FM22, FFM05, FBZP15, FLCdA06, FDC^{+}19, FB16, FB18, FOCSB^{+}20, GZP05]. detection [GCS23, GMM15, GBY21, GS06, GZX^{+}23, GSPL10, GG09, GPG^{+}15, GHHX04, GLG22, GZY23, GYW^{+}22, GYCS21, GYWZ23, HHAE14, HD23, HLL^{+}23, HGP15, HWL^{+}22, HKK08, JA16, JWDF05, JYTK11, JYL23, JHK24, KL07, KKRK23, KBKS18, KLL^{+}11, KLO20, KS12, KYM13, KBD^{+}12, KLK^{+}16, KL10, LWlZ16, LMRMJ08, LE09, LRFF24, LTY^{+}15, LLS21b, LHLZ23, LFLZ23, LZB^{+}23, LLS^{+}23, LYX^{+}24, LG14, LmCT16, LRLR15, LAL^{+}10, LCLH18, LCG21, MYC09, ML13, MP14, MDK^{+}24, MAG^{+}16, MC20, MTV17, MTC^{+}14, MMP09, MTAA11, MSP^{+}18, NCDG21, NNN^{+}22, NB10, OK04, ÖÜ20, PDK96, PZX13, PYWZ17, PD17, PKH23, Pen15, PZM^{+}21, PBI16, PYGGLNG17, PL10, PS05, PLB16, LL17, QKH^{+}12, RG16, RZH17, RXDS22, RB16, RAP16, RCTV12, RCT14, RKKK22, SFF^{+}18, SPC^{+}15, SFK18, SJST07, SVSM15, SZ16, SS09, SOD10, SIRS21, SVF^{+}21, SGZ21, SCC^{+}22, SM13b]. detection [SKBS13, SY23, SMHH04, TABK17, TLY^{+}16, TLFM23, TY05, TKL21, TDK10, TP14, THL13, TBC^{+}21, TNO24, TBZ^{+}24, TAK^{+}22, VCDS^{+}17, VCLS19, VSP06, WJ07, WO10, WZY13, WZT13, WGAD14, WX16, WH18, WCZ^{+}20, WHL^{+}21, WGGH24, WTZ24, WDC^{+}20, WMBY12, WBS14, WSKH13, WB16, WLFL21, WZCY22, WZCG24, XG08a, XYQE24, XSK15, YWZ11, YDP^{+}20, YCA^{+}10, YGH11, YHN11, YGC13, YCZ^{+}23, YHW^{+}23, YZ06, YO11, YS24, YSNiT14, YFDA17, YJC^{+}09, YSKL24, YR06, YG16, ZZZ15, ZZH23, ZMJ^{+}15, ZLZH17, ZZ20, ZWB^{+}22, ZhZFL22, ZSC^{+}23, ZS11, ZJL23, ZFW^{+}24, ZJ05, ZWY14, ZJW15, SAK^{+}24]. detection-driven [TLY^{+}16]. detection-localisation-recognition [CGHTK16]. detections [KEG15]. Detector [BKD01, BS00a, CL00, SGB01, FB12, KY06, MFSB23b, MFSB23a, MCM^{+}17, RLF15, YWL^{+}22, YWM19, ZJJ22, MAY^{+}10]. Detectors [HSSB98, KP00, CHH09, MvGS16, MM06, PK18, TL15, USKB10]. determinantal [JHV19, KLO20]. Determination [LF98]. Determining [HC13c]. deterministic [GB13, KL11]. DETRAC [WDC^{+}20]. DeVCEM [MFP^{+}20]. development [Cre08]. developmental [GLMM16]. developments [Bor22]. device [NLM05, SSHP17]. devices [HSH07, MAG^{+}16, SE11]. DF [YHW^{+}23]. DF-FCOS [YHW^{+}23]. DFA [WHJK23]. DFNet [MWL^{+}24]. DFNet-Trans [MWL^{+}24]. DHumanOutfit [ABB^{+}23]. diagnosis [TDK10]. diagnostic [LSP^{+}16]. Diagram [KSI98]. Diagrams [RM98]. diameter [KŽ12]. diamond [BFR13]. diary [RCJ^{+}13]. dictionaries [SBB18, ZLFH23]. dictionaries-based [ZLFH23]. Dictionary [CWH^{+}13, DH19, FDC^{+}19, GCPF08, TSL14, TRPD20, UFK20, WLW^{+}16, XSQZ15, XW16, ZZL13, ZP18]. dictionary-based [ZZL13]. Diff [LZYW23]. diffeomorphisms [Mar07]. Difference [TMNM09]. differences [CE17, FMS17, HSK23]. Different [KHB01, RWV95, Shi99, TS01, BKK11, CU11, FKS10, MOT17]. Differential [GL95, KPH02, TD04, VB98, WW97, ME18, RMD08, SOJ17, TG95c, YS08]. differential-radon [SOJ17]. differentiation [WCZ^{+}20]. differentiators [HTNN18]. differently [WYX^{+}16]. difficulty [LWYF24]. Diffusion [AG00, BABB19, CBM01, KS96, KY19, SLS01, TÉSK11, BI11, GCC24, HYW^{+}24, KGC05, LG24, LYSS12, WWJ13a]. Diffusion-based [KY19]. Digital [Bor96, Bre01, KCD00, Kis96b, NS96, NGR24, Pud98, Rob96b, SB02, WB97, BRSSAL11, BT05, BBK15, Coe12, CLL14b, DBBB14, EL03, Eva06, FLCdA06, LA11, MOT17, NKPT13, SC96, SOJ17, SRP10, VRKL13, ZZ07]. Digitalization [ASS97]. Digitization [GL97]. Digitizations [GL95]. digitized [CSY08]. Digits [Por00]. dilation [HBF09]. dilations [SVF^{+}21]. Dimension [DL97, CP09, Coe12]. Dimensional [LZ97a, MG95, MNHO00, SF95, SCS99, TK97, WD96, ZM96, ACP16, AMCB20, ASVO12, AH08, BEGB13, BKMV07, DBF04, DM12, GHZ^{+}13, Got08, HQN05, KCD00, KON^{+}17, LB08, LSCK15, ML15, NWJ15, PJW11, Pat13, SOL16, SB05, WD14, ZMM^{+}22]. Dimensionality [KAES99, RRR11, LLL13]. Dimensioning [DV98]. Dimensions [Bor96, Jos99, TML00, CBT^{+}04, CDIF14]. Direct [Dre96, GL98, Neg96, WTYC18, BF07, HC13c, KYYC14, PZC17, SCS14]. directed [BI11, DB14, EKY08, WHGZ20]. Direction [PE09, ACAAC^{+}08, CSS^{+}13a, Dre96, GWT09, HQW^{+}12, Lhu23, MC20, YGH11]. Directional [BS00a, FD99, AS08a, DPM14, FMS17, LSPV04, NNBN20, OAGN18, SXZZ24, TKL^{+}09, ZJJ22, kCE^{+}18]. Directions [AT13, AZP14]. Dirichlet [KBKS18, WZX^{+}14]. disaggregation [QLY^{+}17]. disaster [KB12]. disc [QKH^{+}12]. Discontinuity [SP97b, Spe97, VB98]. Discontinuity-Preserving [SP97b, VB98]. discontinuous [KS03]. discounting [BK07, SS11]. Discovering [JEF^{+}12, JRBD^{+}15, LXW^{+}17, BG16, FR11]. discovery [DLMC16, DHP08, FT23, LC09, MGPP11, MJ17, WW16]. Discrete [Ano15n, DRDKE13, GGO10, IE99, KII98, KC99, LL99, MRW^{+}97, MMS97, PZ08, PZ09, AMGG^{+}16, BTB14, CT12, PV13, TMN06, Žun03, LL08]. Discriminant [HH19, ZZCL14, CLZZ21, ITNP12, LZD^{+}14, MYV19, SAC^{+}12, TLH22, WJ07]. discriminants [TAC23]. discriminate [RAP16]. Discriminating [QV98]. Discrimination [AL99, DH00, YZL16]. Discriminative [CZL^{+}24, GYTL09, PS22, SVSM15, SJ15b, XSQZ15, AAL22, DYM14, DZLH17, HJZ16, JNLG15, LLC12, LTCT14, LLL15b, LSTARMB11, TLB^{+}15, TABK17]. discriminatively [VKL18]. discriminator [FWZ^{+}24]. Disentangled [LYW^{+}24]. disentanglement [LLNZ22]. disentangling [GKH^{+}21]. disguise [CM21]. Disparity [BI11, MGMS01, BK16, Gon09, KN03, LZC^{+}20, MSI10, PTM20, WGAD14]. displacement [BAMK18, FWG18]. displacements [ABB^{+}23]. Display [NNT11, CD10, ZLL^{+}24]. Display-camera [NNT11]. Displays [SGDP01]. disrupting [LSQL24]. Dissected [KHH^{+}22]. Dissimilarity [RPTB01]. Distance [ALK99, APV99, ABLL19, Bor96, BM00, BM02, Chu02, CM99b, Egg98, ER96, KSKB95, Kis96a, KŽ12, LHKC97, LH99, MMS99, Mas02, Por00, Pud98, RG16, SWG02, SJ01, SLK23, SB05, SB02, TV99, CCTCR09, CDJM14, CSMS14, CS20, DT10, ET15, GBB^{+}18, GH08, Gre04, JZZ23, MGW10, MK18, NBFG20, NSEA13, PRR03, REF15, SW04, SCvW11, SCMS13, SCEvdH14, WDN^{+}12, YZX^{+}22, dSdSF^{+}12, BT23]. Distance-based [SLK23]. Distance-Ordered [Pud98]. distances [Ang07, ITNP12, NSEA13, YSX^{+}19]. distancing [JPN^{+}22]. Distillation [PZM^{+}21, BIMD23, FM22, FCOK24, HBKG22, HKL^{+}24, LPSK23, LWLP23, MZ21, QCL^{+}23, SJSL21, WZG^{+}24, WZCY22, YSZ23]. distillation-based [FM22]. distinct [SY20]. distinctive [DDLP10, YK08]. distinctiveness [FLS^{+}14]. distinguish [WLX^{+}14]. Distinguishing [CHL05, WWJ16]. distorted [UWH17]. distortion [CP04, GOF^{+}15, KBJ^{+}10, TM04, WHL14, XMN^{+}15]. distortions [SCGAF^{+}17]. Distributed [BPQ15, CSGM^{+}24, OMLL98, GK23, Ham05, IKST05, MCT10, SKS11]. Distribution [HB98c, TML00, CLO17, Coe12, FL09, FHZX23, FS03, JGM20, Kim04, MPT21, PKD07, PTE12, QAB^{+}11, QT10, STHBH18, TS11, YLLG18, YZL^{+}21, YSKL24]. distribution-aware [YLLG18]. distributions [LH95, TP14]. Disturbances [MPPG98]. Divergence [GD19, RSBA23]. diverse [DR04, WGGHvdW21]. Diversified [LZZ^{+}21]. diversity [MGCS17, WGGHvdW21]. Divide [BPC^{+}17]. DLT [GOF^{+}15]. DLT-Lines [GOF^{+}15]. DNNs [CFM^{+}23, HSTL24]. Do [DAZ^{+}17, DLMC16]. Docking [SVS97]. Document [Ano96d, Doe98, KB98, KH96, KDRC98, LPH01, Spi98, CMH13, LDD09]. Documents [BKMSR98, CB98, SHKP98, GRMH19]. Does [Lau97, SL16b]. DOF [SIT07, FPMK19, TWQW23]. Domain [Ano01m, BKMSR98, DAL^{+}22, GL24, HLL^{+}23, LYD24, Luc01, PD23, TS19, ZD01, AT17, BHA24, BPCT22, BVCP21, CZL^{+}24, CNO^{+}24, DWL19, DNG^{+}24, DWLW23, GLG22, Hu11, KG14, LZZ^{+}23a, LWH^{+}23, LD24, LBCA10, MFSB23b, MFSB23a, MJ17, NFSD13, PLYW21, PV14, PTM20, QCH20, RFF23, RMC^{+}22, SP23a, SHSJ23, SCS14, SIRS21, ŠSJ^{+}20, SGZ21, TMS20, TP05, WZQ^{+}23, WLSO23, WDC^{+}24, YWH^{+}23, YSD03, ZdCR^{+}24, ZLFH23, ZFG^{+}22, ZJL23]. Domain-aware [GL24]. domain-shift [KG14]. Domains [DFH^{+}22, MHL14, RDS24]. Dominant [Spi98, KZ05, RCT14]. door [ESS10]. Dot [CCP97]. double [WLZM20, XW16]. double-layer [XW16]. Dougherty [Ano95d]. down [BYJG23, HLB17, KMN11, MAJ16, MSP^{+}18, TY22, ZWY14]. down-up [TY22]. DP [SHKP98]. DRAU [OS19]. Drawing [JV97, SP97a]. Drawings [CLD96, DL97, DV98, LCD97, PC99]. drift [RMD08, SCALFG^{+}18]. Driven [CKB96, IW97, PBPD^{+}17, SM97, WLXL24, ABD11, BUD19, BBSD15, BCM13, CGD^{+}23, CSZ^{+}15, CÇ15, FAB12, MSP^{+}18, RGA10, TLY^{+}16, TZY08, WK21, Wor05, ZIT^{+}13]. driver [CPT07, OBTMT15, TDT12]. driving [FPMK19, RCJ^{+}13]. drone [SS21]. dropout [LPSK23, NL23]. DSDNet [DDZ^{+}23]. Dual [Kim17, ZMM^{+}22, ÇÖD08, CT10, CS04, CYD^{+}22, CLL17, DFP23, DDZ^{+}23, FWZ^{+}24, Hu11, IDY^{+}18, KTP08, LDH^{+}14, LLWZ21, SKS11, WSKH13, OS19]. dual-channel [FWZ^{+}24]. dual-models [LLWZ21]. dual-point [CS04]. dual-source [IDY^{+}18]. dual-space [DFP23]. dual-stream [CYD^{+}22]. dual-tree [ÇÖD08, CT10, Hu11]. dual-view [LDH^{+}14]. due [BHBF10]. duplicate [CHC11, JN09, XTZZ14]. duplicated [ZTH^{+}11]. during [DLS^{+}09]. dyadic [TNO24]. Dynamic [BPBS13, BBHF10, CS07, CC00, GB13, GSK02, HML15, KAES99, LE09, MdOBA19, MS96b, TW98, UFK20, UIK22, WPK09, XQZL23, XST04, YLM11, ZT98, ZKRH04, AAMO16, BMJF^{+}17, Bar05, BDFG17, BBK15, DLC^{+}24, DWV19, DD11a, EL07, FT23, GA13, HKM22, HQW^{+}12, JBC08, JYY^{+}24, KG14, KTP08, LWH03, LDT21, MSI10, MWTN04, MMP09, NNBN20, QSX17, SKLM22, SCL13, SHK11, TS16, TT16, TN07, TMN06, VWMZ15, XG08b, XZQJ21, YJ16, YR06, ZJZY16, ED16]. Dynamically [AH24]. Dynamics [MJS97, LHLZ23, TPD^{+}16, TFD07, UIK22, YG16].

E-ProSRNet [CU20]. ear [AZN11, HNC05]. early [JZZ23, SGS^{+}10, WH18]. eccentricity [IAP^{+}11]. Ed [Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b]. Ed. [Ano07a, Ano07b, Ano07c, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k]. Edge [BKD01, BS00a, CBM01, HSSB98, HLF^{+}97, JB99, MGPJ11, PA10b, PDTE06, RM02, SGB01, ABN^{+}20, BSRV17, DETE17, FZ20, GB22, GMF14, JZL^{+}23, JM09a, KY06, LMDB11, ML13, MLJC20, SS09, WO10, WBS14, WPK09, WZCG24]. edge-avoidance [JM09a]. edge-aware [BSRV17]. Edge-Based [HLF^{+}97, DETE17]. Edge-Preserving [RM02, MGPJ11, GB22]. Edges [LL97b, PE09, TWQW23]. edit [DT10]. editing [CWLY22, CAGN24, WQY^{+}21]. editor [GSST03]. Editorial [Ano01g, Ano05f, Ano05i, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano07d, Ano07e, Ano15n, Ano15o, Ano17j, Ano17k, Ano18d, ACW^{+}16, BCH^{+}18, BK15, BPQ15, CGL^{+}21, GKL^{+}17, JGSP16, Kak95, LLNS18, MYC^{+}14, NPBM22, PSY^{+}21, SUS^{+}15, TVY^{+}18, YLM^{+}17, ZZP^{+}16, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano05e, Ano05g, Ano05h, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13c, Ano13e, Ano13g, Ano13h, Ano13b, Ano13d, Ano13f, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e]. Editorial [Ano14f, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17a, Ano17b, Ano17c, Ano17d, Ano17l, Ano18a, Ano18b, Ano18c, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j]. Editorial [Ano22k, Ano22l, Ano22m, Ano22n, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j]. Editorial- [GKL^{+}17]. EDITORS [DCCL99, MT97, BS99b]. effect [GGGROE^{+}17, YLK^{+}23]. Effective [LDGS^{+}13, LG17, LKZ20, CWO^{+}11, DETE17, NF21, PD17, SSM06, TAC23, BIMD23]. effectiveness [TKDN16, ZBDP15]. effectors [SRHC13]. Effects [CFA98, FT98, MPPG98, FMS17, HC13a, YLLG18]. Efficiency [LHH^{+}98, KTP08]. Efficient [ATG15, Bar18, BSRV17, BM00, BM02, BG16, CC01, CCL^{+}17, CSMS14, CS20, CYES00, DOSD11, DG01, DZJB14, DMW10, DSK^{+}20, FKW98, FN14, HWZ^{+}23, HMB17, HP96, KB00, KRBSV17, LWY^{+}17, LZmC^{+}17, LA05, MNL^{+}17, MK01, MdRNM15, OK04, PZX13, PLJS14, PG13, PL08, REF15, RCTV12, RSS07, SKH08, TSL14, TWQW23, TGSH98, XOF05, XL98, ZRZ^{+}24, AMN18, BB16, CGHTK16, CBT^{+}04, CYNO11, CZ14, CJYW23, CP20, CÇ15, DRTT24, DLV15, GRGB^{+}13, GCS23, HDL^{+}20, HKL^{+}24, KHH^{+}22, LDH^{+}15, LLS21b, LMM22, PD17, RCT14, Sav23, SPK23, TLEF06, VAWW10, WXWC18, WHL^{+}20, WLL22b, XSD12, XWLY23, ZZH23, ZWT^{+}14, ZCW^{+}23, ZCWH23]. EFSCNN [ZDZ^{+}23]. EGAN [JZZ23]. ego [RN12]. ego-motion [RN12]. Egocentric [DLMC16, ADR16, ADFR18, ASC17, AB18, BMB^{+}17, CGHTK16, DBT^{+}17, FF23, LRFF24, LSH19, RFF23, RFMF21, PBPD^{+}17, VCDS^{+}17]. Egomotion [DT96a, DH00]. Eigenimages [LB00]. eigenspaces [BWL04, EKY08]. Eigenvalues [SB98a]. Eigenvector [PLL00]. Eigenvectors [SB98a]. Elastic [ACLS98, AG00, BSH13, BL09, Far11, JKM07, NBDB04, RFS03, WPSL18, WR08, ZP11]. elastic-net [WPSL18]. Elastically [Dav97]. elasticity [LV11]. elderly [MML^{+}16b]. electroencephalogram [HKZ^{+}16]. Element [TGSH98, KRBSV17]. elementary [CKK^{+}12, ZZRC15]. elements [MGS15, SW05, TCZ^{+}12]. Eliminating [Kim04]. Elimination [CM99a]. elliptic [LDGS^{+}13]. Elliptical [DGH98]. EM-ICP [CP20]. Embedded [EA95, AZSVK05, Bar05, CVP10, CKB10, HZW^{+}10, SBB10, VAWW10, YCA^{+}10]. Embedding [BSZ^{+}21, CLZZ21, FKV^{+}11, GHZ^{+}13, GG20, LCP13, LHY14, LZD^{+}14, LLTL14, LLL^{+}14, LZS16, LCG21, SK15, TS19, XHW09, ZRKZ^{+}11]. embeddings [KL07]. emergence [Ham05]. Emerging [LWYF24]. emotion [HKZ^{+}16, NNS^{+}18, LL17, ZMJ^{+}15]. emphasis [CLDP23, SH09]. Empirical [BKD01, CAO^{+}23, FHP01, RPTB01, DAM12, FCM20]. enable [SSdVL06]. enabled [SRO^{+}19]. enables [TFL^{+}09, WRKP05]. Encoded [KD96, Jea11, SKBS13, YLM11, ZDZ^{+}23]. Encoder [ZH18, BB24, XGT^{+}22]. encoder-decoder [XGT^{+}22]. encoders [SFK24, WGGH24]. Encoding [YSX^{+}19, TVLS08]. End [MSV^{+}20, PKH23, RZZ23, SHW24, MWL^{+}24, SRHC13, ZLLP21]. end-effectors [SRHC13]. End-to-end [MSV^{+}20, PKH23, RZZ23, SHW24, MWL^{+}24, ZLLP21]. Endoscope [OD97]. endoscopic [HSKH07]. endoscopy [MFP^{+}20]. Endothelia [GAD01, ZMCA05]. Energy [Ano01m, JZZ23, Luc01, MRF96, ACG^{+}09, Bar18, EyGS11, MAJ16, QTLP22, WAPB17]. energy-based [ACG^{+}09, Bar18, EyGS11]. enforcing [Lhu18]. engine [LEA^{+}10, SM10]. Engineering [DL97, DV98, EFF98, PRW97b, SOJ^{+}95]. Enhance [QDLB17]. Enhanced [AAL22, BSMK13, GSP02, JZWD16, KACR^{+}23, ACDB12, CU20, KGC05, LSD^{+}07, LD24]. Enhancement [AAMO16, SLS01, ZCL99, Ang07, CSD^{+}24, DZQ24, DWL^{+}24, HWW06, HSJS10, JRH24, JZL^{+}23, LYBT17, LSH19, LZB^{+}23, LWFL24, LZL^{+}22, LXW^{+}23, MYZ^{+}24, SRB21, TKL^{+}09, WLZ23, YAK^{+}08]. Enhancing [CE17, Dem96, MAJ16, SASYGCRG24, WGGH24, ZA22, AZ15, WSY^{+}16, YST21]. enlarged [LYW^{+}24]. enrollment [FBF08]. Ensemble [Cha24, KUHY18, JVD^{+}20, LWLP23, PD23, VBVB19, ZWL16]. ensembles [HBL^{+}17, PWSvdH17]. entirely [TN08]. entity [Che24]. entity-aware [Che24]. Entropic [DFSC20]. Entropy [TVE^{+}16, GHHX04, JLM22, PYWZ17, SE11]. Entropy-based [TVE^{+}16]. Envelope [HGB98]. environment [CP09, LY13, ST10, ZKSV18]. environments [AM04, Ano06h, BPLT15, CM12, Cha21, Cha24, CPS10, FPDK12, GKK05, GC19, GPC^{+}10, HCC^{+}16, LS12, LA05, MP09a, NKB11, ROGT14, STC^{+}16]. Epiflow [ZN08]. Epipolar [KHB01, ACAAC^{+}08, BF14, CPC08, CKS^{+}05, LWY^{+}17]. epipolar-based [CPC08]. epipolar-plane-image [CKS^{+}05]. epipole [LB10]. Epipoles [LF98]. Equalization [ZCL99, BK07]. Equation [KS96, BLJ^{+}23, CS10, MZC^{+}05]. Equations [CBM01, VB98, VF96]. equidistant [AXSVL14]. Equivalence [CU10a]. equivalences [CU11]. equivalent [BYJG23, RG17]. eraser [TDZ^{+}20]. erasing [DZZ^{+}23]. Erratum [Ano06h, OBH04]. erroneous [CX11]. Error [BRP04, CACB17, Jur99, KS95, OD02, SRT01, BÁRM23, CPS05, LHY14, QAB^{+}11, RBdDS14, SB96a, UTB^{+}11, WZWH16, ZWN14]. Error-aware [CACB17]. Errors [CFA98, KW99, KB00, LZ97b, RFS03]. estimated [RF23]. Estimates [Mil99, WALL00, DLC14]. Estimating [BK01, BFY00, DGC12, GA09, KRJ^{+}08, Lhu23, MC09b, PBW14, Shi99, SWMM22, TML00, TZM98, TZ00, WSV05, ZL01, LMC09, LYKY19, RN12, RA15, YSL11]. Estimation [Ano01m, ACB98, BA96, BGK98, CSC96, CL00, CFA98, Dan97, DC98, FD99, Imm96, Jos99, LB10, Lin02, Luc01, MS97a, MGMS01, NDBT95, SP97b, Spe97, SJB02, WLD99, WPB^{+}14, ZD01, AJ23, AS08a, AS09, ACG^{+}09, ABVC16, AVC19, AYG23, AH08, BZP^{+}23, BDVK10, BPLT15, BJS14, BG18, CSS^{+}13a, CL18, CS10, CTH20, CPPY21, CLO17, CRCM16, CC16, DM12, DPCA15, DMSM21, DJF14, EBN^{+}07, FL09, FDW21, FPMK19, FSI21, GSGJ24, Gon09, GKGM20, GLZF23, GML^{+}21, HD09, HWK^{+}21, HSH07, HCLZ21, HBH11, HSHA20, HH12, IH15, IDY^{+}18, JGM20, JC06, JF10, KUHY18, KHK10, KYYC14, KLKF20, KGB17, KMN11, LTAA23, LWY^{+}17, LvdHK^{+}15, LSC08, LCZ09, LWLT17, LZZ^{+}21, LZ24, LZC^{+}20, LYA13, MSR07, MWL^{+}24, MSSS09, MP09b, NT10, NWNT17, NZH^{+}23, ODD96, ODT17, OSM16, OSM17, OTAH20, PRK19, PD05]. estimation [PY19, PBT14, PV06, PHH^{+}15, PRCP16, PZC17, RDM^{+}11, RAC^{+}13, SOK16, SECS15, SBIK16, SHE17, SM06, SO07, SM21, SPK14, SRHC13, SM13b, SCEvdH14, TMNM09, TAK09, TST14, TP14, TNO24, TP05, UU18, UTB^{+}11, VBT19, WCZ23, WHM^{+}09, WTZ^{+}21, WDC^{+}24, WSJ15, WCF10, WTYC18, WYW^{+}22, XTZ^{+}18, YCH07, YZT^{+}13, YA12, YC05, ZDLS13, ZEGEJ15, ZSL^{+}16, ZC19, ZIT^{+}13, ZZP12, ZZG^{+}24, ZCW24, ZDF10, ZHZ17, dP10, dMFU10]. Estimator [TZ00, CBT^{+}04, CYC10, Dre96, HBH11, SKLM22, XSL^{+}23]. estimators [CLL14b]. Euclidean [BM02, BI10, BM00, Cou13, CM99b, Egg98, ER96, GBB^{+}18, KGK10, LHKC97, MMS99, PCJ14, SW04]. Euler [IE99]. Evaluate [WZC^{+}21]. evaluated [SV14]. Evaluating [BH12, Ste01, GKBW14]. Evaluation [BKD01, Che00, DL05, FHP01, GAD01, HRS02, LCZ^{+}01, LPH01, PMR17, PR03, RPTB01, WLM^{+}14, AA20, Bor19, Bor22, BZ14, BG09, BT23, Cha21, CZHT15, CCSS14, CYG16, DL10, FF23, GE08, GJMO14, HYJ11, HMC10, HC13b, HWW06, KDT^{+}18, LK03, LFL08, MO11, MSM17, MM06, OAGN18, PD14, RN12, RBdDS14, RDSF15, RLC^{+}11, SJST07, SHL18, SL16b, TPT15, VO24, VD10, WL15, WBS14, WHL14, YAK^{+}08, ZFG08, ZCLX20]. evaluation-based [OAGN18]. Evaluations [RTM^{+}17]. Event [WPZ^{+}16, CGR13, HHM^{+}16, HNB04, JYTK11, LmCT16, LCG21, SM12, Sav23, SM21, SMHH04, YLM11, ZhZFL22]. events [ABI^{+}04, CCF17, DLS^{+}09, HS14, LCSL07, OBTMT15, PSYZ13, RCJ^{+}13, TD04, XYRS17]. everyday [WSY^{+}16]. Evidence [ANM98, BBK15, MYLP98]. Evidence-Gathering [ANM98]. evidences [YSS^{+}14]. Evidential [HHM^{+}16]. Evolution [LL99, DCS05]. Evolutionary [KBD^{+}12, RF02, BPB11, SCD11, LLWX24]. exact [CSMS14, Mal21]. examples [CYY^{+}23, FFFP07, SS21, XST04, ZTB20]. Exclusively [LC19]. exemplar [AYD^{+}18, AZ15, FBK16, OMBH06, ZH18]. exemplar-based [AYD^{+}18, FBK16, OMBH06, ZH18]. exemplars [SBH^{+}17]. Exhaustive [Lin02]. exhibiting [ABB^{+}23]. exit [LJJY24]. exocentric [AB18]. Expansion [VF96, BKK11, TYH^{+}21]. expectation [SBPF17]. Experience [PSB^{+}24]. experiment [LFMP13]. Experimental [LCZ^{+}01, HF11]. experiments [HMEB07, HKA13, CH17]. expert [CSDNR17, Mah16]. experts [EKY08]. explainable [MMLC23]. Explanation [WLXL24]. explanations [VO24]. explicit [NLW^{+}17, QGZ^{+}23]. Explicitly [HFKN97]. exploitation [CP21]. Exploiting [CHC11, DDLP10, LRFF24, PXTZ14, PKC^{+}18, PD23, ROGT14, STC14, Kui08, NY14, YDP^{+}20]. exploration [JZZM23, OMW^{+}07]. Exploring [HSK23, HZK19, Kui08, MBCJ17, WZWL24, YSKL24, ZMM^{+}22]. exposed [WYX^{+}16]. Exposure [YWL^{+}20, ABK^{+}18, LLL^{+}20, MOT17]. expression [CSG^{+}03, DH19, EB14, HOH^{+}07, LY06, LDH^{+}15, LSCM03, LWSC16, MB11, SS17a, SKVS13, SSS13, SASYGCRG24, WY07, XFP^{+}16, ZZC^{+}23]. Expressions [YB01, HKZ^{+}16, SHK11, SSS13, TMM16, WWJ16, ZWLH24]. Expressive [CSV^{+}16]. Extended [CTF^{+}98, KSS97, WB97, ADR16, LCP13]. Extending [GR05, HL23, KKCK23]. Extension [FDMA97, GBB^{+}18, MMV06]. extensions [PRK19]. exteriors [HBH10]. external [MLH13]. extract [MB95]. extracted [BY08]. Extracting [Cre99, CKS^{+}05, FKL^{+}98, SC99, FYH11, XHX^{+}19]. Extraction [ANM98, AMMV99, ADDK99, CCP97, DT96b, GN98, KII98, KZ05, LPH01, LHHC98, May99, MNHO00, Nis95, Rob96a, SCS99, TSP97, UZC97, WH01, ZKLZ23, BB03, CM12, ÇÖD08, CNC03, DBF04, Dam08, DDWZ12, FLCdA06, FS03, GHZ^{+}13, GBY21, GYCS21, HNC05, JZL^{+}23, KA12, LCZ09, LSQL24, LS09, LDL^{+}19, MTG07, MZB^{+}10, MHL14, NY14, ÖÜ20, PQML11, RT14, RC13, SASYGCRG24, Ste13, TRPD20, WHJK23, YHS^{+}20, YT13, YR06, ZZG^{+}24, ZGC20]. extrapolation [Kim04]. extreme [GLG22, HLL^{+}23, MPM16, SPK14]. Extrinsic [LLSV00, PA13]. Eye [FB16, FB18, HP05, KMBH09, MM05, AZSVK05, GZY23, HH07, JWDF05, LSP^{+}16, NNT11, SFK18, SFWG08, TAK^{+}22, WSV05, WJ07, WB15, YC05, ZJ05]. eye-detection [AZSVK05]. eyebrow [LLC13]. eyes [WASF14].

facade [FBS21]. Face [Ano01k, CC03, HHWP03, HL01, JLY^{+}17, JT17, KL07, LY06, MCAF21, MYLP98, MHAF13, OB14, RY98, SSN03, TTH07, VBVB19, YKA01, ZZC^{+}23, ADR16, AM04, AC09a, AC09b, AKC11, ABVC16, ARARCE11, BC10, BCF06, BF10, CH06, CM21, CFB05, CH17, CYY^{+}23, CLTW23, DM12, EKY08, ESS10, ET15, FBF08, GJ10, GZ19, GYCS21, GYWZ23, HWK^{+}21, HASS10, Hu08, Hu11, HH19, HDF12, JLD12, JGM20, KTE^{+}17, KCM^{+}17, KKSH23, KHA^{+}05, KMBH09, LRW08, LB14, LZZ^{+}23a, LG24, LL08, MYK03, MCB13, PY08, PZX13, PBT14, PTE12, LL17, RM03, SECS15, Sav23, SRB21, SB22, SAC^{+}12, SDK22, SSM06, SJSL21, SKVS13, STC14, SBH^{+}17, SM13b, TD04, WJ07, YCA^{+}10, YAK^{+}08, YZL^{+}21, ZZZ15, ZBDP15, ZH18, ZTB20, ZJ05, BGPD09, JGP19]. face-iris [ET15]. Face-SSD [JGP19]. faces [AZP14, BL09, BW15, BSBW14, DBBB03, KCM^{+}17, Kou03, OTAH20, SR23, ZKC03]. Facets [ZT15]. Facial [ÇÖD08, CSG^{+}03, EB14, FM22, KdVL99, LSCM03, MDM^{+}21, TW98, YB01, CWLY22, DB03, DH19, GZJ05, GKH^{+}21, HOH^{+}07, HKZ^{+}16, JB23, JLY^{+}17, JGP19, LC14, LB05, LY06, LDH^{+}15, LSQL24, LZC^{+}20, MB11, RG16, SS17a, SHK11, SSS13, SASYGCRG24, SL16b, TMM16, TLWT12, WY07, YDP^{+}20, YLM11, ZZP^{+}16, ZMJ^{+}15, ZWLH24]. Factorization [GRCD18, SRT01, TI01, ZEGEJ15, AO16, HRC09, KBWT16, KCZ18, LLL13, ZZ10, LLTL14]. factorization-based [KBWT16]. Factorized [GPG^{+}15, GML^{+}21]. Factors [BGPD09, CP09, GML^{+}21]. Fake [GYCS21, GYWZ23]. Fall [GMZ^{+}22, ALK^{+}09, YG16, ZZH23]. FAM [HBZ^{+}24]. family [DBBB14, SKA23]. far [BBC^{+}07]. far-infrared [BBC^{+}07]. Farin [Ano95e]. fascia [TLY^{+}16]. fashion [SP23a]. Fast [BCMCB09, CH11, Coe12, CM99b, Egg98, GK95, HQN05, Imm96, IP98, KBJ^{+}10, LCZ09, LK03, MAP99, MPST08, MMP15, MPPP14, MĆK09, NFSK97, OG98, QLY^{+}17, RM98, SW04, Sup02, VWMZ15, WHC14, WNH05, XTZZ14, YO11, ARARCE11, BPB11, CBT^{+}04, CCYC12, DSK^{+}20, FL09, GBB^{+}18, GC19, HDS08, HMA10, HZW^{+}10, LZLP10, MDdMG09, MU11, MPK24, SFF^{+}18, Tan11, UWH17, WWG^{+}18, YB07, YWL^{+}22, ZDZ^{+}23]. Faster [ZS19, BAP08, MCM^{+}17]. FCOS [YHW^{+}23]. FDGA [ZZH23]. feasibility [WML21]. Feasible [WSSD96]. Feature [BL98b, GHZ^{+}13, HR99, HH19, KSS97, KN99, KKP24, LCD97, LFLZ23, LZ24, MFJ95, MDL^{+}23, NFSD13, Nis95, Nis99, PHHL23, PLL00, PBQ99, PM97, Rob96a, RWV95, SB98a, TGQ23, TS01, TPR^{+}00, WF02, ZWW^{+}20, ZDZ^{+}23, BWG17, CBD^{+}03, CM12, ÇÖD08, CWO^{+}11, CYNO11, CZ14, CZS^{+}20, CLZZ21, CYD^{+}22, CXYZ24, CZHT15, CWW^{+}22, CNO^{+}24, CP09, CK09, DOSD11, DDWZ12, DLV15, DWL^{+}24, DG11, EXP^{+}20, FYH11, FLL^{+}23, GCT^{+}14, HYJ11, HSHA20, HBZ^{+}24, HKL^{+}24, HNC05, JPN^{+}22, JZZM23, JYX^{+}23, JSC23, KGFP10, Kim15, KYM13, LDH^{+}15, LHSG15, LTY^{+}15, LWZP17, LSH19, LZZ22, LZB^{+}23, LK03, LFL08, LZL^{+}17, LS09, LXW^{+}24, MFSB23b, MFSB23a, MPT21, NNS^{+}18, ODD96, OZT19, PZX13, PQML11, Pha17, Pun03, QT10, QLY^{+}17, RMC^{+}22, RG16, RAP16, SB13, SW17, SASYGCRG24, ŠSJ^{+}20, TLH22, TY05, TFD07, TP14, TKAK14, UTB^{+}11, VBA19, WD14, WLX^{+}14, WHL^{+}21, WHJK23]. feature [XYQE24, XZX^{+}21, XMN^{+}15, XGT^{+}22, YSL^{+}14, YDP^{+}20, YZL16, YZX^{+}17, YWL^{+}22, YO11, YSY^{+}18, YLX^{+}18, ZRL^{+}11, ZNG^{+}13, ZLL^{+}24]. Feature-Based [HR99, LDH^{+}15, LFL08]. Feature-domain [NFSD13]. feature-oriented [FYH11]. Features [AM00, COW98, CS98, HdVL99, Jon97, LLZ23, LRLR15, PA00, RY98, SA95, Tsa96, ACP16, AMCB20, BCM13, BL14, BEGB13, BDL^{+}06, CCSS14, CNS18, CR18, CH09, DSNN08, EK12, ET15, FAZ14, FMGA^{+}12, FAB12, FCOK24, GLM17, GTP18, GS95, GLG22, GBL08, GYWZ23, Gwa17, HAT^{+}15, HGP15, JY14, KDT^{+}18, KK11, LXFM16, LYSS12, MU11, MB95, NHK08, PMR17, PTM20, PMCN22, QGZ^{+}23, RDSF15, SCE04, SJB20, SKVS13, SCMP14, SM13b, TLP^{+}17, UMH16, VAC16, WJ07, WZWL24, YHW^{+}23, YFF^{+}23, YG16, YG17, ZMM^{+}22, ZCLX20, ZYS09, dCCP12, AW09, BETV08, LL08, SYZ^{+}15]. February [Ano20o, Ano21o, Ano22o, Ano23p, Ano24k]. FedER} [PSB^{+}24]. Federated [PSB^{+}24, LYD24, WHY^{+}23]. Feedback [MBKB02, MIUS16, XSZ^{+}20, KDV12, MW13, MLK21, Pen03, RGA10, dSdSF^{+}12]. feedback-based [dSdSF^{+}12]. femoral [KNO^{+}09]. Few [CLL^{+}21, ZYQ^{+}23, BZP^{+}23, CWS^{+}24, CLDP23, FFFP07, JHK24, LHL^{+}21, LFLZ23, LZW^{+}24, MHX19, SAK^{+}24, SS21, WZQ^{+}23, WKT22, XWLY23]. few- [SAK^{+}24]. few-beam [BZP^{+}23]. Few-Shot [ZYQ^{+}23, CLL^{+}21, CWS^{+}24, CLDP23, JHK24, LHL^{+}21, LFLZ23, LZW^{+}24, MHX19, WZQ^{+}23, WKT22, XWLY23]. fidelity [MWTN04]. Field [DC98, MCPB00, CMZ24, CMD06, DWC16, FLS^{+}14, HC13b, HW06, HNC05, JC06, KHR^{+}16, KS03, KKCK23, LSCK15, LL12, MLB^{+}18, MHMO09, MJPS16, WB11, XMN^{+}15, ZSL^{+}16, PV13, WKP13]. Fields [BA96, Mas02, MRF96, WW97, WZWT99, WSSD96, AMCB20, BP05, CL18, LPR^{+}03, SBB18, SK15, TWW14, VGR16]. FIFNET [EH21]. Figural [MPPG98, PEFM98]. Figure [AL99]. filling [HKA13]. film [TDK10]. Filter [CGL98, MDL^{+}23, DD11a, DYM14, HBB^{+}12, HSJS10, JVD^{+}20, JHV19, KDV12, LÁB15, MZ20, MHSP10, MiMO^{+}16, TTXT21, TKL^{+}09, WCYS13, YNCO11, RRR11]. filter-based [DD11a]. filtered [PCJ14]. Filtering [Jon99, Ang07, Ano06h, BL09, BKMV07, CNDS13, FMG23, GKK05, GB22, KLK14, KORC10, LAFLB16, MWF07, SBD22]. Filters [Spe97, AS08a, AXJE21, AC09a, BW11, DZLH17, FAZ14, HDF12, Jea11, KG14, LRW08, LST13, LY06, LSPV04, SBB10, SAC09, TLH22, WB15, SC15]. Find [Hob00, MT16]. Finder [PKP97]. Finding [CDH99, GS06, LF96, PF99, SBZ97, WWW95, CSMS14, OGB14]. Fine [GDCM17, KFSM17, OD02, TB99, YDP^{+}20, CHL21, DFLS23, GZL^{+}23, JLZ23, JRH24, KHG22, LYKY19, ML13, PHHL23, QZY^{+}24, RT14, SY10, WHY^{+}23, YNZ^{+}19, YYH^{+}23, ZIT^{+}13]. Fine-grained [KFSM17, YDP^{+}20, CHL21, DFLS23, GZL^{+}23, JLZ23, KHG22, QZY^{+}24, YNZ^{+}19, YYH^{+}23]. fine-scaled [LYKY19]. fine-tuning [PHHL23]. Finger [WF05, ABEN09]. fingerprint [UBEP09]. fingerspelling [KK15]. fingerwriting [CGHTK16]. Finite [EB13, TGSH98]. fire [BJS14]. First [DPB00, RM02, VF96, ACP16, DD11a, RFMF21, RCJ^{+}13]. first-person [ACP16, RCJ^{+}13]. Fish [TML00]. Fisher [MIUS16, XMT22, YZL16]. fisheye [AXSVL14, UWH17]. Fit [BCA98, MB05]. Fitted [Lil97, ZWT^{+}14]. Fitting [BA06, BCLNG18, Jac01, KB00, CC96, LDGS^{+}13, LG17, MBD^{+}22, WCYS13, Ano95d]. Fiume [Ano95e]. Fixation [Dan97]. Fixed [GLR^{+}99, ROJX09, CTWH15]. Fixed-point [CTWH15]. flattenable [RB18]. Flexible [BHSD^{+}13, BS99a, NMP97, AAB19, FMG23, LWYF24, LHJ^{+}09, NS16]. Flight [LSKK10, SLK15, BHMB10, HHAE14, HEPH15, LBK10]. FLIR [LCZ^{+}01]. floating [RLB17]. floating-point [RLB17]. Floor [MCPB00, ES06]. Flow [BA96, DC98, FSA01, LSH19, LHH^{+}98, MNCG01, NDBT95, SP97b, Spe97, SJB02, WALL00, XS98, ADGB16, BL09, CHZ^{+}13, CSS13b, DRAB08, FWG18, FBK15, FBK16, FSV07, GYTL09, GPY^{+}07, Gon09, HMF10, JM09a, KN03, KN11, LNM^{+}21, LS08, LB10, LmCT16, MN06, Mar07, MZC^{+}05, MEYD11, MCF10, PBW14, RDM^{+}11, RPG12, SM06, SM21, TLCH05, TD19, WWJ13a, ZSCP08, ZLS^{+}13]. flow-based [BL09, CHZ^{+}13]. Flow-guided [LSH19]. Flows [WD96, ACG^{+}09, HC13c, LGG^{+}18]. fluctuations [AFMY14]. Fluid [WALL00]. Fluorine [JSWL24]. Fluorine-19 [JSWL24]. fluoroscopic [KNO^{+}09]. fMRI [KGC05]. Focal [Che08, SCCP05]. Focus [PGP15, SKOS95, ALM23, CXFS06, FWZ^{+}24, IKST05, LTAA23, PLYW21, ZLHJ18, DR04]. Focus-aided [PGP15]. focused [CMZ24, SL16a]. Focusing [BM99, May99, WASF14, ZS19]. FOE [Neg96]. following [NPM^{+}16, NMAL23]. Font [CWS^{+}24, KH96]. food [CNS18, FCM20, MPM16]. foot [TDT12]. footage [CSK22]. Force [HNC05, IW97]. Force-Driven [IW97]. Forces [DF01]. fore [ZHS^{+}24]. fore- [ZHS^{+}24]. fore-/background [ZHS^{+}24]. Foreground [FT23, AHDM10, CVP10, CW15, CMG16, DD11b, LRLR15, MCCRAC20, TS24, UFK20, WZCG24, YO11, ZhZFL22]. Forensics [CGL^{+}21]. forest [CFYU12, CZ14, LLJ^{+}23, MRH19, dSdSF^{+}12, CGHTK16]. Foresting [MSF^{+}12]. forests [ZJW15]. forgery [CGD^{+}23]. Form [BSF02, CF01, CS98, FAB97, HS06, MKY01, ADC19, BvdHL^{+}13, Liu10, MFB11, UJ22, WSFTK18]. formal [DAL^{+}22]. Formation [MS97b]. Forms [ÜE01]. Formulation [ACB98]. forward [AT13, FMS17]. four [HF11, HQW^{+}12, KDSF20]. four-connected [HQW^{+}12]. Fourier [ANM98, DUC97, DG01, LEA^{+}10, TKK24, TS00a, ZS11]. Fourier-Mellin [DG01]. Fourth [Ano96d]. Foveated [YYL96].

FPGA [MZB^{+}10, MAY^{+}10]. FPGAs [MZC^{+}05]. FPN [ZJJ22]. FRA [DK13]. fractal [LPZ08]. fractal-based [LPZ08]. fractures [GYW^{+}22]. Fragment [ASZ99a]. Fragments [EDB12, DT09, TS17]. Frame [ADDK99, WLXL24, FAZ14, HG11, KKSC23, PR03, SM21, SVF^{+}21, TY22]. frame-based [PR03]. frame-to-frame [FAZ14]. Frame-wise [WLXL24]. frames [EH21]. framewise [UO16]. Framework [ADDK99, Car96, GGR01, LH95, VM01, ASFP03, BWG17, BYK^{+}18, CGR13, CCPK16, CCF17, CYY^{+}23, CMH13, CNO^{+}16, CL08, CU11, DWB11, FFM05, FKV^{+}11, FMG23, GGP23, GCD^{+}18, GML16, GYWZ23, HKHE14, JLD13, KK15, KBN12, KSR^{+}12, LC11, LV11, LLC13, LWZP17, LHJ^{+}09, LH03, MSV^{+}20, MAJ16, MIP16, MP20, MTR^{+}23, NS16, PJW11, PL10, PLKP23, PMW05, RLS06, RB18, RS03, RA15, ŠRDC09, TÉSK11, TMB12, WML21, WTZ24, WZG^{+}24, XWLY23, YGC13, ZC19, ZZ20, ZDF10]. frameworks [CU11, TPT15]. FReBIR [PFGG09]. Fréchet [BT23]. Free [BvdHL^{+}13, BSF02, CF01, CS98, FAB97, LHSG15, Liu10, MKY01, TML00, UJ22, WRB06, CZS^{+}20, CC16, JGP19, PZM^{+}21, RC03, SS17a, SLK23, YWH^{+}23, ZLLP21, ZJJ22]. Free-Form [BSF02, CF01, CS98, FAB97, MKY01, BvdHL^{+}13, UJ22]. Free-hand [LHSG15]. Free-Swimming [TML00]. freedom [LWLS12, Sha11]. freehand [MJPS16]. Freeman [Kak97]. French [KABP98]. frequencies [SRM20]. Frequency [Ano01m, AT17, Luc01, SDK22, LWH^{+}23, NL23, QGZ^{+}23, QSXS23, SGS^{+}10]. frequency-domain [LWH^{+}23]. FRIDA [RMC^{+}22]. friendly [CPP^{+}11, CTWH15]. fringe [MSV^{+}20]. Front [Ano17j, Ano17k, Ano17l, Ano18k, SK02]. Front- [SK02]. FS [Neg12]. FSpH [ZWT^{+}14]. Full [BR95, WWWF23, LPR^{+}03]. Full-parameter [WWWF23]. Fully [AGL23, ACB98, BW15, CZ14, CBTC23, CJWW22, FWL^{+}20, MS96a, SFF^{+}18]. Fully-attentive [CBTC23]. Function [GK98, GESB95, KH96, BSM10, HL23, KDSF20, PSR08, RSS07, TS16]. function-based [PSR08]. Functional [Hod95, RDR95]. Functionalities [RR95]. Functionality [BB95, Sta95]. Functions [BGSdVL98, AJ23, CGU11, CU10a, CU10b, DLV15, EPH^{+}21, PRR03, WR08]. Fundamental [BGK98, CZZF97, TZM98, ZL01, ASCF13]. fundus [QKH^{+}12]. fuse [UJ22, ZRL^{+}11]. Fusing [BC10, PS12, BKK11, YG16]. Fusion [HSIW98, HH19, HSJS10, LL08, RFL02, WY21, XZW^{+}23, YWL^{+}20, AM06, ABEN09, ALY^{+}22, BK16, BF10, CA10, CYD^{+}22, DS07, EH21, ET15, ES04, FWLQ23, FWZ^{+}24, GLOC10, GTMR23, HD09, HWG21, HGR^{+}13, HYW^{+}24, JBC08, JYX^{+}23, KH23, LvdHK^{+}15, LB08, LB19, LFL08, LDC^{+}13, LBCA10, MZ20, Mig12, NNS^{+}18, PBT14, PWWQ16, PLYW21, RXDS22, SvdMH15, TMB12, VMN16, WZW17, WYW^{+}22, XSL^{+}24, XWC^{+}23a, XGT^{+}22, YW07, YZX^{+}20, YR06, Zac18, ZA22, ZLFH23, ZRZ^{+}24, ZZZP09]. fusion-based [HD09]. FusionDiff [HYW^{+}24]. Future [MBHRC21, BCC^{+}21, KK17, NHZ^{+}22, RFMF21, ZZZ15]. Fuzzy [KW00, KGU10, LSB^{+}00, MWF07, MCPB00, Pha01, RMFB02, SUO00, SU01a, SU01b, SWG02, SB13, TB99, WDB12, ALK^{+}09, BKPS15, CUSZ07, CU10a, CU10b, CU11, DK13, GF15, ITNP12, LMDB11, PFGG09, WSSS13, WWWF23, ZUS06]. fuzzy-connected [ZUS06]. Fuzzy-rough [SB13]. fuzzy-rule-based [DK13].

G [Ano95e]. Gabor [Far11]. GAFL [SBD22]. gain [YCH07]. Gait [AFMY14, CT13, AM17, CR18, CNC03, HWT^{+}24]. gaits [Boy04]. GaitSCM [HWT^{+}24]. GAMA [XWC^{+}23b]. Game [YB95, PKK^{+}09, RMN^{+}17, VMC^{+}16, ZZC^{+}23]. game-theoretic [VMC^{+}16]. games [CL17, KBD^{+}12]. GAN [Bor19, Bor22, CLTW23, JZZ23, KACR^{+}23, LHG^{+}23, SKS^{+}22]. GANs [FSG22, GM19, RB19, YZX^{+}22]. gap [DNG^{+}24, MTP21, WM20]. gated [LWLP23]. Gathering [ANM98]. Gauss [CRC97, JWG04]. Gaussian [CTWH15, AQ09, AMCB20, CE14, EB13, FL09, FWL^{+}20, Jur99, KNL15, KLK14, KKCK23, Kui08, KMN11, LBCA10, MSR07, MRW^{+}97, OD99, PKvGS16, RRR11, Ste13, UK12a, WWCZ15, WLW^{+}16]. Gaussians [SGMC15, VWMZ15]. gaze [CC16, MM05, NMAL23, NKB11, NLM05, WSV05, YC05, ZSSF16]. GC [CUAT13]. GC-ASM [CUAT13]. GCN [WZCY22]. GCN-based [WZCY22]. GEIKD [LWLP23]. Gender [ZSSF16, CSDNR17, GBVDC18]. General [MWL99, MWLA99, CL08, DMW10, DSY10, LC14, RR06, RLC^{+}11, WTZ24]. generalised [BWG17]. generalizable [BHA24]. generalization [CNO^{+}24, GL24]. Generalized [CLCO13, DFH^{+}22, GPY^{+}07, LK97, MUS06, MP09b, MTR^{+}23, Zac18, AH24, CCL^{+}17, EB13, FL09, GML16, LYD24, PW23, ZS11]. Generalizing [OZT19, WO10]. generate [CKLP09]. Generated [MWL99, MWLA99, JWG04, PHY^{+}11, ZCLX20]. Generating [LMDB11, SZGK24, MNR18, YB01, ZT98, ZMM^{+}22]. Generation [EK98, LK00, Mun95, Nis99, OYTY98, SDK^{+}24, WLXL24, BMvT^{+}19, CWS^{+}24, CP09, DM12, KKSC23, LWY^{+}17, LNM^{+}21, LZZ^{+}21, LZZ^{+}23a, LYW^{+}24, LWYF24, LLG^{+}24, LG24, LCG^{+}24, SP06, SPRS23, ZYQ^{+}23]. Generative [BK15, CWLY22, MCB13, MC22, PL07, RMC^{+}22, ZhZFL22, BCMR16, BBCF20, BMvT^{+}19, DYM14, FFM05, FFFP07, JNLG15, Kim15, NWJ15, OZT19, Pec07, RB16, SEFV15, SB22, SDK22, SXY^{+}23, TLB^{+}15, TY22, VKL18, XHW09, AW09, BT23, DWL^{+}24, FDSB22, GFL^{+}19, LB19, LHG^{+}23]. generator [GLZF23]. generators [GDIIHK11]. Generic [ALIRT18, BKMSR98, GESB95, KBAS16, LD98, RSL10, CC03, DMW10, FKV^{+}11, OCVV04, RLS06, RPBK22]. Genetic [DUC97, SCS99, SC98, GRGB^{+}13, HDS08, SW05]. Genetically [HBL^{+}11]. genomics [KFRD^{+}18]. genomics-inspired [KFRD^{+}18]. geo [RTM^{+}17, WCF10]. geo-accurate [RTM^{+}17]. geo-location [WCF10]. Geodesic [HUI16, PD05, RC13, MJ11, PMCN22, YG17]. geodesic-aware [PMCN22]. geodesic-induced [YG17]. geodesics [WPS03]. geographic [CCPK16]. Geometric [AGB^{+}15, BR95, COW98, DUC97, GK98, GBB98, GL95, HSIW98, KT15, KS96, MNSK98, RH95, SLL01, Tsa96, XWC^{+}23b, ÅS17b, Bar06, BPB13, Bre03, CHSV08, CK09, CPS05, FF09, GSV05, JBWK11, KSY15, PXTZ14, PD14, SRHC13, WB12, WZWH16, WMZY23, XFP^{+}16, XHX^{+}19, YS08, ZY14, dSEdSPdMF24]. geometrical [ABD11, Nis96, ZCLX20]. geometrical/statistical [Nis96]. geometrically [KLKF20]. Geometries [LV96]. Geometry [Åst97, Ano95e, Ano15n, BM98, CFA98, Col97, DRDKE13, FL96, GHMQ97, GSK02, KSF19, PRW97a, Sch06, SA02, TZ00, Ver97, WW97, Bar05, CLL14b, HCLZ21, IH15, JY14, LWY^{+}17, ME18, NNT11, PS05, RB19, ROGT14, SSM06, TKAK14, VSP06, VAC16, WPS03, WXZG18]. geometry-aware [HCLZ21]. Geometry-Based [FL96, VAC16]. geometry-guided [RB19]. Geons [NL96]. GeoSay [XHX^{+}19]. Gesture [RLMK15, AAASC11, BMJF^{+}17, DWV19, HMF10, JM09b, LLS21a, MdBJG15, PS15, TD04, TDT12, YS09, ZT15, ZSSF16].

Gestures [ZXK02, CEA16, JB23, LCP13]. Getting [LC19]. Ghost [YWL^{+}20]. Giant [MAY^{+}10]. Gibbs [VGR16]. gist [HL13]. given [KS03]. Gleason [SM13a]. gliomas [RAC^{+}13]. Glitch [CGD^{+}23]. Global [Ano01m, BFMW23, KA08, KB95a, Luc01, LXW^{+}23, SKB96, SBD22, WZG^{+}24, WB16, YZT^{+}13, YSL11, ZM96, CLZY15, GFW13, HHWP03, JA16, LWZC14, MML^{+}16a, MTP21, PW23, PB11, SCMP14, TBZ^{+}24, VNNB14, WWJ16, WAPB17, WZWL24, RK11]. Global-aware [LXW^{+}23]. Global-local [BFMW23]. globally [MPPP14, UO16]. globally-optimal [UO16]. GLocal [YSL^{+}14]. gloss [LMC09]. glossy [PK05]. GMC [WTZ24]. GNNS [MTR^{+}23]. goal [DLS^{+}09, PSYZ13, TABK17]. goal-based [TABK17]. Golay [HTNN18]. GOLD [SGMC15]. good [CH17, DAL^{+}22, PWWQ16]. GPA [CLCO13]. GPS [JF10]. GPU [CPP^{+}11, NHH14]. GPU-friendly [CPP^{+}11]. grade [LDL^{+}19, RAC^{+}13]. Gradient [GCC24, PA10b, WZJ^{+}21, WSSD96, HTNN18, HAT^{+}15, HC13b, JSWL24, KS03, LLY^{+}18, LMDB11, SSL^{+}12, SK15, XSL^{+}23, ZD18, ZLS^{+}13, PE09, SYZ^{+}15]. gradient-based [HAT^{+}15]. gradient-correlation [HTNN18]. Gradient-guided [GCC24]. Gradient-Spatial-Structural [SYZ^{+}15]. gradientlet [MZ20]. gradients [BL04, SKA23, WLL22b]. grading [PKD07, SM13a]. GradPaint [GCC24]. grained [CHL21, DFLS23, GZL^{+}23, JLZ23, KFSM17, KHG22, LLJ^{+}23, QZY^{+}24, YNZ^{+}19, YDP^{+}20, YYH^{+}23]. Grammatical [JvdBS99]. Grand [BGPD09]. granularity [BHA24, DWL^{+}24, LSGY24, LLLW23]. granulometric [ZMCA05]. Granulometries [BJ96]. Graph [AYB^{+}18, BSALF18, BPB11, CALO20, EXP^{+}20, GPDR13, HTEB11, HCLZ21, JO11, JBWK11, KCD00, LEB07, LLG^{+}23, NRJ11, OTO06, TGQ23, VPL23, YYL98, ZRKZ^{+}11, AS09, AAL22, BB16, CHP^{+}11, CPP^{+}11, CK11, CBB19, CUAT13, CLL^{+}14a, CLL17, CFM^{+}23, Far11, FKV^{+}11, GDIIHK11, GML16, HMCT22, JZZM23, KS15, KIS17, KT08, LWZP17, LLG^{+}24, LYX^{+}24, LCG^{+}24, Mah16, MMK04, PLLL03, RDT^{+}19, RYLZ24, RAHT11, SAS12, SOL14, TKK24, TBC^{+}21, UK12a, VPP^{+}23, WW16, WXZ24, XHW09, XYZH11, XAB07, YQL^{+}23, YW16, ZP11, ZSY^{+}19]. Graph-Based [HTEB11, BPB11, JBWK11, AS09, CK11, TBC^{+}21, WW16]. graph-cut [CUAT13]. graph-cuts [CLL17, RDT^{+}19]. Graph-matching-based [CALO20]. graph-partitioning [MMK04]. Graphical [Ano95e, WKI^{+}16, DPCA15, NN13, XG08b]. Graphics [Hob00, TVY^{+}18, Gon09, KLBP11]. Graphs [Bre01, NWP97, NS96, CNDS13, FT23, MDFS11a, MDFS11b, SRS11, SPRS23, ZNG^{+}13, dMFU10]. grasping [LCP13]. Grassmann [LWSC16]. gravity [GSGJ24]. Gray [DG01, PA00, Sha05, WB97, Dem05, KL07]. Gray-Level [DG01, PA00, Dem05]. Grayscale [TSP97, WCZ02, MYZ^{+}24, YCL07]. greedy [KOC17]. Grey [GPK99]. grid [JCLZ21, ZWLH24]. grids [HHAE14, SB05]. grocery [TS19]. Grooves [LKK00]. Ground [AL99, LB98, Cre08, RLMK15, SYPK13]. ground-truth [SYPK13]. grounded [COV^{+}22]. grounding [SHW24]. Group [GZX^{+}23, KC99, SC99, WPZ^{+}16, BGE^{+}17, BSZ^{+}21, MGPF08, UMH16, XSQZ15, AGB^{+}15]. Grouping [ABD11, ASZ99a, CH96, CA97, Hen98, JDP97, KN99, LM99a, MRF96, PF99, PB99, SN99, YJA96, GZP05, LBNS09, YS09, ZZRC15].

Grouping/degrouping [ABD11]. Groupings [CN95]. Groups [MFJ95, MJD^{+}00, SM97, KRJ^{+}08, MCL16, SAL16, VMC^{+}16]. groupwise [GKBW14]. Grow [TAC23]. growing [GFL^{+}19]. growth [RAC^{+}13]. GSNNet [GZX^{+}23]. Guaranteed [SK98]. Guess [EDJ^{+}20]. Guest [Ano01g, BCH^{+}18, CGL^{+}21, MYC^{+}14, NPBM22, TVY^{+}18, YLM^{+}17, GSST03, DCCL99, MT97, BS99b]. guidance [BKP10, DLMC16, HSKH07, LPC^{+}20, NPM^{+}16, PBT14, RTM^{+}17, RGA10]. guide [TCB^{+}08]. Guided [CLCZ23, KGB17, AZN11, ASFP03, BHA24, BIMD23, CLZZ21, DDWZ12, DFLS23, GZX^{+}23, GCC24, JVD^{+}20, JYX^{+}23, JRH24, LWY^{+}17, LSH19, LKZ20, LZZ^{+}23a, LLS^{+}23, PRCP16, RB19, RS03, SP23a, YS24, ZP18]. guiding [KH23, OH05]. gymnast [RZH17].

Hadamard [WML21]. Hadamard-based [WML21]. Hairs [LKK00]. half [JB23]. hallucinate [FFA^{+}19]. Hallucinating [NHZ^{+}22]. hallucination [SDK22]. Hamming [REF15, YSX^{+}19]. Hand [ABEN09, AS17a, CW00, NWNT17, PC99, Por00, SKOS95, ZXK02, BMB^{+}17, DWV19, EBN^{+}07, JM09b, KGB17, LLS21a, LCP13, LHSG15, LSH19, LLL^{+}20, MdBJG15, OTO06, PBT14, PS15, RZZ23, SGH07, ZT15, ZJW15, dP10, DBZ07]. Hand-based [ABEN09]. Hand-Drawings [PC99]. hand-gesture [MdBJG15]. hand-held [LLL^{+}20]. hand-pose [dP10]. Hand-Printed [Por00]. handle [MiMO^{+}16]. handles [VZP^{+}09]. Handling [BVCP21, CH11, FBK16, KFN15, LST13]. handoff [CYP^{+}10]. handwashing [HPvB^{+}10]. Handwriting [AHD98]. Handwritten [DLHT99, HY98, GRMH19]. Hankel [LL17]. haptic [NPM^{+}16, RRAR^{+}16]. Hard [FB97, LBP23, MT16]. hard-to-find [MT16]. Hardware [MZC^{+}05, MNHO00, AK10, AK11, AHDM10, Gon09, MSI10, PCC13]. hardware-based [AK10, AK11]. hardware-oriented [PCC13]. harmonic [HMF10, SGS^{+}10]. Harnessing [VGLP17]. Hash [GK95, FXWW17]. Hashing [RH95, Tsa96, CBS17, CLL^{+}14a, FWG18, HMCT22, HDL^{+}20, JBWK11, LBP23, ML15, SHS^{+}23, WWG^{+}18, ZWT^{+}14]. Haze [LYBT17, ECC18]. hazy [ZHZ17]. Head [CSS^{+}13a, HGP15, PHH^{+}15, TNO24, ABVC16, AVC19, CC16, DPCA15, HDG^{+}14, HCLZ21, MBD^{+}22, TST14, WPQ20, YWZ11, YC05, ZWLH24]. heading [RS03]. heading-guided [RS03]. Heads [FM99]. healing [LZTX24]. health [RBC22]. Healthcare [NPBM22]. Heart [LSB^{+}00, WYW^{+}22]. Heat [KS96]. heatmap [SJSL21]. heatmap-based [SJSL21]. heavily [BPLT15]. heavy [LG17, MSSS09]. HECOL [CPC08]. Height [SF16, ATG15, BABB19, CH06, LSC08, Mas09]. heightfields [EMMV19]. held [LLL^{+}20]. help [MST16]. hemispherical [GHA10]. hepatic [ARC14]. Herb [Kak97]. heritage [BÁRM23]. heritages [dOSJVBS12]. hermeneutics [GMW12]. Hessian [LTCT14]. Heterogeneous [DWV19, GBL08, HH19, PZX13, WLW^{+}16]. Heteroscedastic [KB00]. Heuristic [KVdG^{+}97]. Hi [GTMR23]. Hi-ROS [GTMR23]. Hidden [Che98, KABP98, BCM06, CL17, CLCO13, NN13, VMN16, XQZL23, ZYXZ13]. hiding [YCL07]. Hierarchical [BAM16, CWH^{+}13, CN95, DPCA15, FWG18, FKL^{+}98, FMG23, HUF05, HP96, KBKS18, KD96, LZW^{+}24, LXW^{+}17, LVS20, ML13, NN13, PCR^{+}04, SL96, SPW15, Tan95, TGFF15, YZ06, YNCO11, YW99, YSY^{+}18, ZWB^{+}22, BPC^{+}17, CL15, CZ14, CDIF14, Cou13, HBH10, JEF^{+}12, KS15, KSF16, LRD19, TLB^{+}15, TS19, XSQZ15, ZWN14].

Hierarchy [Jon97, SN99, MdRNM15, NFA04, PCJ14]. High [AM15, CJL06, CJC01, DT96b, EA95, EPH^{+}21, HSHA20, MCPB99, PCJC98, SM21, UO16, BC10, BEGB13, BKMV07, BBK15, CBT^{+}04, DLC^{+}24, DRAB08, HBH11, JLY^{+}17, JPP^{+}14, KA08, LGL15, LGD16, MWTN04, NWJ15, QGZ^{+}23, QSXS23, RMN^{+}17, RT14, SRM20, SP06, SL16b, MNR18, VGR16, WD14, YAK^{+}08, ZYT10]. high-dimensional [BEGB13, BKMV07, NWJ15, WD14]. high-frequencies [SRM20]. high-frequency [QGZ^{+}23, QSXS23]. High-level [EPH^{+}21, JLY^{+}17, RMN^{+}17, ZYT10]. High-order [UO16, JPP^{+}14, KA08, LGD16, VGR16]. high-performance [DRAB08]. High-Resolution [MCPB99, PCJC98, SP06]. High-Speed [DT96b, HSHA20]. high-stakes [SL16b]. Higher [KSRS16, SJ15a, She16, ZZP12, PL08]. Higher-Order [SJ15a, KSRS16, ZZP12, PL08]. highlight [GCD^{+}18, GHHX04, WXZG18]. Highlights [CTE95, MS00, ABC^{+}03]. Highly [SM10, HHG^{+}20]. hippocampus [XFSC13]. Histogram [MGW10, MAP99, WCZ02, ZT15, ZCL99, BK07, CKC14, KGU10, MHSP10]. histogram-based [KGU10, MHSP10]. histogram-wise [CKC14]. histograms [JWG04, KBMD15, LLR10, LL04, LL12, LDC^{+}13, LmCT16, NHY10, STD14, SM17, TLB^{+}15, ZD18, PA10b]. histology [SM13a, Tan11]. Historical [HSBS16]. history [WRB06]. HMI [FKL^{+}16a]. Hock [SCR^{+}17]. HOG [AT17, HC13b]. holes [CHSV08]. Holistic [VCLS19, ZC19, WCZ23]. Homeostatic [FY06]. homogeneity [KLL^{+}11, MVP06]. homogeneous [BFR13, YZX^{+}22]. homographies [CPS05, SCEvdH14]. homography [GYF18, CPC08]. Homotopic [Pud98]. Hopfield [BBB96]. Horizon [ABN^{+}20, MAL10]. Hough [CGHTK16, CGR13, CS04, CL95, DGH98, FS03, GLR^{+}99, GRB13, KB00, KBD^{+}12, LY05, MGK00, MNHO00, MAK^{+}17, Ols99, PKP97, SYK96, Sha06, SK98, SKBS13, dSM14]. Hough-based [GRB13]. Hough-CNN [MAK^{+}17]. houses [ÜB05]. HRCT [SBK^{+}99]. HSGAN [YZX^{+}22]. HtHT [KB00]. HTS [dSM14]. HTSn [dSM14]. hull [BL08, MHL14]. Human [AC99, BDT23, BL01, BDFG17, CFCP11, CMBP09, DAZ^{+}17, DLF06, FCM20, GCS23, Gav99, GBB^{+}18, GMW12, GAD01, LWlZ16, LRD99, LLC13, LSW18, LSTF12, MDK^{+}24, MdBJG15, MYLP98, MG01, NMAL23, PC05, SBIK16, SPK^{+}02, SS21, YG16, ZXK02, Ano06h, ABB^{+}23, BCM13, BSZ^{+}21, BB24, BSBW14, CGH08, CL18, CCFC13, CYNO11, CTH20, CPPY21, CLO17, CNC03, DPM14, DMSM21, DIMT12, FFY^{+}04, FOCSB^{+}20, GKK05, GBY21, GMZ^{+}22, GLZF23, HRC16, HUF05, HCC^{+}16, HWW06, ITNP12, IDY^{+}18, JS07, KV06, KIS17, Kim17, KC22, KRK11, KPKH07, KLK^{+}16, Kou03, LE09, LRFF24, LSCM03, LWH03, LYA13, MML^{+}16a, MFB11, Mal21, MHK06, MdRNM15, NFM08, NLM05, OMBH06, OVJ^{+}21, PT08, PDS^{+}07, PQML11, PKC^{+}18, PYS03, Pop07, Rem04, RSPD12, RR06, ROGT14, RS03, SKM06, SH08, SP19, SRHC13, TR09, UU18, UFF06]. human [VAC16, VGSMN16, VKNK14, WCZ23, WS08, WH18, WLO^{+}18, WTZ^{+}21, WPB^{+}14, YO11, YS08, YST21, ZMCA05, ZT15, ZSSF16, ZKC03, ZCW24, ZDF10, Ziv10, BCDH10, CEA16, HG11]. Human-computer [MdBJG15, ZSSF16]. human-delineated [Ano06h, GKK05]. human-human [SP19]. human-object [LRFF24]. Human-Scene [MDK^{+}24]. humanoid [ZMJ^{+}15]. Humans [DAZ^{+}17, RFF23]. Hybrid [BCHR24, CC96, FLS^{+}14, SOK16, DWW^{+}12, FN14, KSR^{+}12, KL11, LLF18, LZWX21, LXW^{+}24, MK18, VMP03]. hypercomplex [AS09]. Hypercube [DRCF95, LHKC97]. hypergraph [YYZL19]. hypergraphs [BB13, BB15a, DB14]. hyperquadric [CC96]. Hyperspectral [ZXC^{+}20, GL19, RRK13, TLH22, VBA19, XWC^{+}23a]. hypersphere [MIP16]. Hypotheses [MS97b]. Hypothesis [LVW97, BT17, JHV19, LWY^{+}17].

I-Learn [DLMC16]. IAPR [EHG^{+}10]. Iberian [CCR^{+}05]. IBims [KLKF20]. IBims-1 [KLKF20]. ICA [DBBB03, Hu08]. ICA-based [Hu08]. ICDAR [Ano96d]. Iconic [CBD^{+}03]. ICP [CP20, FDMA97, PLH04, YB07]. ICP-based [YB07]. ICycleGAN [SZB^{+}21]. ideal [LMP^{+}19]. identical [HBL^{+}11]. Identification [CTE95, GLR^{+}99, KH96, LCD97, LLZ23, TN08, ZKLZ23, ABEN09, ABC^{+}03, BCC^{+}18, BRA^{+}10, BCM13, BHA24, CTM^{+}13, CH17, CL08, CKP^{+}19, DPRC17, FHZX23, GZL^{+}23, HBKG22, ILRB04, JRAJ17, KU19, KDSF20, LY05, LLCY21, LLLW23, LZYW23, LSCM03, LML^{+}23, LYD24, LN10, LCG21, ML13, MKF15, PWSvdH17, PGGM04, RCTV12, SYZ^{+}15, ŠSJ^{+}20, TDK10, UMH16, VCDS^{+}17, WHY^{+}23, WPK09, WWG^{+}18, XYZH11, DAL^{+}22, HH05, WZWL24]. identification-verification [KDSF20]. identifier [WF05]. Identifying [HHG^{+}20, KEG15, PRG^{+}14, TN05, TESY15, GS06, PXTZ14]. identity [GFY^{+}14, ZZC^{+}23]. if [Ano17j, Ano17k, Ano17l, Ano18k]. IFS [BBC00]. IFTrace [MSF^{+}12]. IGMG [BHA24]. II [CU10b]. IIANet [JYL23]. Illuminant [DC98, AJ23, DJF14]. illuminants [APB10]. Illumination [ADGB16, BFF97, BWL04, FW97, GG09, Lai00, LZ97a, MCF10, OD99, OD01, ASC17, AC09a, AC09b, AZP14, ARARCE11, CCYC12, DD11b, DL10, Hu11, Jea11, KTE^{+}17, LMP^{+}19, LCT09, LY06, MTVM04, OK04, TD19, WLZ23, YWZ11]. illumination-based [ARARCE11]. illumination-encoded [Jea11]. illumination-invariant [AC09a, TD19]. Illumination-robust [MCF10]. Image [AYB^{+}18, AK11, ABW97, APV99, Ano95d, Ano01l, Ano06h, AKE23, ACW^{+}16, BK01, BS99a, BPQ15, BCC16, BFY00, BB15a, BHF08, CGL98, CXYZ24, CLCZ23, CM97, CH09, CC00, CL97, Cre08, CW00, DT96a, DF02, DCCL99, DPB00, DH00, DG01, DSH04, EK98, EA95, FRL^{+}98, FL96, GFS04, GB17, GGMV08, GMW12, GHS95, GRMH19, GGR01, GKH^{+}21, HR99, HWZ16, HLF^{+}97, HMA10, IP98, JWG04, JSZY17, KB98, KSS97, Kis96a, KD96, KVdG^{+}97, KGM19, Lai00, LN98, LDH^{+}14, LLE^{+}09, MBKB02, MAP99, MFSB23b, MKK02, MS97b, MK01, MSW15, MBMC11, MYLP98, MPPG98, MGLB17, MLK21, NDN^{+}97, NVWV97, NLW13, OD97, OTL96, OYTY98, OBH04, PZ09, PF99, PBQ99, PM97, PMV00, RWWH00, RC03, RM98, Ros95, Ros96, Ros97, Ros98, Ros99a, Ros00a, Ros01, Ros10, SUO00]. Image [SU01b, ST96, SC99, SLST99, SF95, Shi99, SBK^{+}99, SPK^{+}02, SL99, Ste01, TVLS08, TS00a, Tay00, TZ00, THT^{+}98, UZC97, VKP98, WN99, WLD99, WD96, WCZ02, WZX^{+}14, WKI^{+}16, WQY^{+}21, WALL00, YGC15, YB95, YZX^{+}20, YFZ98, ZW97, ZL01, ZFG08, ZLL^{+}14, ZCL99, ÅS17b, AGL23, AM06, AH24, AA20, AQ09, Ang07, Ano17j, Ano17k, Ano17l, Ano18k, AC09a, AO04, AMGG^{+}16, AYG23, AM15, ASFP03, ATC^{+}13, BUD19, BT17, BL20, BLJ^{+}23, BLKG21, BK07, BP05, BF07, BCDH10, BT05, BvdHL^{+}13, BB04, BSMK13, BCA16, BPB13, BRPC17, CG09, CFYU12, CH06, CT10, CEO18, CM16, CL15, CBB19, CKPV21, CYNO11, CUAT13, CLZZ13, CH17, CWC^{+}20, CHL21, CLZZ21, Che24, CWW^{+}22, CAGN24, CE17, CLO17, CFM^{+}13, CU20, CU10a, CU10b, CU11, CCSS14, CBTC23, CG04, CKS^{+}05, DBF04]. image [Dam08, DZQ24, DR04, DKG22, Dem05, DSNN08, DAM12, DWL^{+}24, DCS05, DJF14, DZLH17, DDZ^{+}23, DWLW23, DB14, DSK^{+}20, EPH^{+}21, ECC18, FPC^{+}08, FDW21, FZ20, FWLQ23, FWZ^{+}24, FY06, FFL14, FBS21, FAB12, FYH11, FMG23, GRGB^{+}13, GFL^{+}11, GY19, GBY21, GGP23, GSS12, GKBW14, GH08, GSST03, GS08, GCPF08, GDR04, GDCM17, GYWZ23, HTNN18, HDS08, HMC10, HJ12, HC13a, Hei04, HMCT22, HC13b, HWW06, HYW^{+}24, HKL^{+}24, HGS08, JKW^{+}21, JHA17, JCLZ21, JYX^{+}23, JRH24, JMPG11, JHK24, KS15, KK13, KHG22, KA08, KN03, KHH^{+}12, KH15, Kim15, KKSC23, KH23, KKP24, KLY21, KSF19, KMT11, LT05, LEE^{+}18, LC11, LH95, LSC08, LC14, LEB07, LLTL14, LSP^{+}16, LWLT17, LRZ^{+}19, LKZ20, LPC^{+}20, LZZ^{+}21, LWW^{+}21, LHL^{+}21, LLNZ22, LZZ^{+}23a, LHG^{+}23, LWFL24, LWYF24, LPZ08, LL12, LFL08, LLC11, LS12, LTCT14, LCL^{+}14, LGL15, LLL15b, LZmC^{+}17, LZL^{+}17]. image [LZWX21, LLJ^{+}23, LZZ^{+}23b, LPV07, LZL^{+}22, MWF07, MYYY17, MZ20, MYZ^{+}24, MVP06, MUS06, MSR07, Mah16, MNL^{+}17, MSG10, MMV06, MMK04, Mas09, MdOBA19, MGPP11, MCL16, MvGS16, MB05, MLJC20, ME18, MTAA11, MGPJ11, MK24, MTR^{+}23, NHK08, NL23, NHTG15, NHZ^{+}22, NGR24, OTO06, OM19, OAGN18, OK04, OSY18, PJW11, PSY^{+}21, PSE^{+}11, PLJS14, PNSF21, Pen03, PV15, PLYW21, PV14, PC15, PA10b, PFGG09, PG13, PLKP23, PD23, PBG04, Pun03, QAB^{+}11, QBZ21, QTLP22, RDM^{+}11, RRK13, RB19, Rem04, RZZ23, RLG^{+}14, RBA20, RFS03, Sah05, SCD11, SRB21, SEFV15, SCC17, SG11, SGMC15, SB13, SKH08, SZL^{+}23, SZW^{+}21, SKU^{+}09, SAV24, SA15, SCvW11, SASYGCRG24, ST20, SHS^{+}23, SYL^{+}24, SPRS23, SM23, SZB^{+}21, TYH^{+}21, TLEF06, TS16, TBFJ15, TMB12, UJ22, VCLS19, VPL23, VMP03, WLZW04, WZ04, WO10, WSSS13]. image [WKP13, WHC14, WPSL18, WHL^{+}20, WGGHvdW21, WML21, WLZ23, WDC^{+}24, WM20, WLZM20, WGZL20, WK21, WWJ13a, WLL22b, WWWF23, XTZZ14, XZX^{+}21, XWC^{+}23a, XYW^{+}08, XSZ^{+}20, XGT^{+}22, YZT^{+}13, YSL^{+}14, YGH11, YH19, YS24, YCL07, YSY^{+}18, YSZ23, ZK17, ZXC^{+}20, ZZZ06, ZTH^{+}11, ZYXZ13, ZTH^{+}14, ZD18, ZWW^{+}20, ZLFH23, ZHS^{+}24, ZYZ24, ZZCL14, ZHL^{+}20, ZSG^{+}20, ZZHZ23, ZIT^{+}13, ZCW^{+}23, ZLS^{+}13, ZLHJ18, ZLL^{+}24, ZUS06, ZU09, dMFU10, MSF^{+}12, Ros00b]. image- [SAV24]. Image-Based [FL96, CG04, FPC^{+}08, LSP^{+}16, SASYGCRG24, SYL^{+}24, SM23, WLZW04, WM20]. image-guided [ASFP03]. image-language [JHK24]. Image-Pair [DH00]. Image-speech [GRMH19]. image-text [LZL^{+}17]. image-to-image [WGGHvdW21, ZYZ24]. ImageCLEF [THL13]. Imaged [CB98]. Imagenet [MSM17]. Imagery [Ano15n, BM99, CJC01, DRDKE13, May99, MNSK98, MCPB00, NK00, PCJC98, DZL07, DS07, HOH^{+}07, KFSM17, OM19, PSR08, PK18, STO17, SSN03, YCH07, ZZZP09]. Images [AG00, Ano95d, Big97, Boo97, BM97, CA97, CM95, CJC^{+}98, Dav97, DUC97, Doe98, FKL^{+}98, FMR01, FM99, GPK99, GSU00, GBB98, GN98, GJP96, HdVL99, HRS02, Hei99, JV97, JB99, JEK98, KW99, KCD00, KDRC98, KS96, KSI98, KMA^{+}00, KdVL99, LF96, MW00, MS97a, MGMS01, MY95, Mas02, MCPB99, MWL99, MWLA99, ME98a, MAM97, Muk97, NMP97, NL96, OD99, OD02, PF99, Pud98, RC97, RY98, RFL02, RMFB02, SA96, SF97, Spi98, SPQ^{+}17, SB02, SM99, TSP97, TK97, WB97, WH01, ZT98, dCCP12, ÁB13, ATG15, ALIRT18, BB16, BI10, BCMR16, BDHM09, BSH13, CCTCR09, CCR^{+}05, CTM^{+}13, CSS^{+}13a, CYD^{+}22, CNS18, DMAD17, DFSC20, DCFM07, DWC^{+}24, ET15, EX17, FMGA^{+}12, FL09, FDC^{+}19, GL19, GE08, GCEC07, GBVDC18, GML16, HWK^{+}21, HHAE14, HQN05, HSJS10, HLKK19]. images [IDY^{+}18, JEF^{+}12, JRH03, JGM20, JZL^{+}23, KL07, KSG^{+}19, KN04, KS12, Kou03, KSY15, KNO^{+}09, KSG^{+}13, LLNC20, LJHH07, LPS^{+}11, LB05, LYKY19, LDD09, LS09, LC19, LMDB11, LBCA10, LP10, LYSK17, MLB^{+}18, MN06, MOT17, MJ11, MAL10, Mig12, MB95, MGPF08, MHAF13, NKPT13, NBFG20, NZH^{+}23, NHTG15, OJRT08, PE09, PL10, Pey09, PMCN22, PS12, PCR^{+}04, QKH^{+}12, RF23, RSS07, RBdDS14, RLF15, RTM^{+}17, SOL16, Sch06, SJ15a, SBH^{+}17, SS11, SZGK24, SdB03, TAK09, TA13, TS11, TRPD20, TGFF15, TP05, TAK^{+}22, ÜB05, VBA19, VMC^{+}16, VJ17, VGLP17, WBS14, WPK09, WLI08, WB11, WYX^{+}16, XHX^{+}19, YHR^{+}05, YWMS08, YZ06, YT13, YLX^{+}18, ZMCA05, ZSCP08, ZRL^{+}11, ZC19, ZHZ17]. ImageWeb [XTZZ14]. Imaging [SGK00, AZP14, BN15, BK15, CKF18, DLC^{+}24, GHA10, GCD^{+}18, GHMT09, GPC^{+}10, HGSM11, KLL^{+}11, KLBP11, SGA12, WAPB17]. impact [TM04]. impaired [CNO^{+}16, LM16]. impairment [MAG^{+}16]. Imperfect [DY98]. Implementation [Bre03, GLR^{+}99, LHHC98, MNHO00, MSI10, MFB11, MZC^{+}05, MAY^{+}10, NN04, SBB10, SM10, dLAH07]. implementing [KL10]. Implicit [HSIW98, LDPD97, LSB^{+}00, RAH97, ÜE01, ZOMK00, CLL^{+}21, HUF05, QGZ^{+}23, WSKH13]. importance [AXJE21]. Imposing [FB97]. Improve [ACB98, ZW97, BVWS21, FBF08, KBMD15, LZQL24, WZC^{+}21, dSdSF^{+}12]. Improved [AM17, CM12, GPC^{+}10, MFSB23b, MFSB23a, Mil99, MB05, OEK08, VCDS^{+}17, GYW^{+}22, HH07, HWZ16, HL23, KDSF20, Sav23, SZ07, STC14, SVF^{+}21, SYPK13, WLZ23, ZSDK19, SPK23]. improved-variation [HWZ16]. improvement [SHE17, TVE^{+}16]. improves [BHMB10]. Improving [CL17, FSL24, GBF12, HCC^{+}16, HBZ^{+}24, JZZM23, LvdHK^{+}15, LCG^{+}24, RPG12, RF23, TL15, WASF14, XSL^{+}23, XJK12, YAK^{+}08, BSH13, CCPK16, CE17, GMM15, KKP24, QWHW20, SXY^{+}23]. Improvisation [Hod95]. impulsive [MGPF08]. IMU [GYF18]. IMU-camera [GYF18]. in-the-wild [JT17]. In-vehicle [OBTMT15]. inaccurate [KEG15]. including [NL17, WR08]. Incompatibility [Åst97, Col97, PRW97a]. incomplete [KBN12, MYC09]. incompressible [ACG^{+}09]. inconsistent [LPC08]. Incorporating [ALM23, GW07, LHH97, dSdSF^{+}12, CSY08, PYWZ17]. increasing [ZBDP15]. increment [NFM08]. Incremental [DHP08, GB08, HRC16, IT15, PZM^{+}21, XG08a, ZYQ^{+}23, Dam08, FFFP07, JLM22, MZ21, RMC^{+}22]. independence [YZL^{+}21]. Independent [BKMSR98, DT96a, FD99, NFM08, EKY08, LT05, ME18, MDL^{+}23, NGR24]. independently [OCVV04]. Index [Ano95b, Ano95c, Ano96b, Ano96c, Ano97b, Ano97c, Ano97d, Ano97e, Ano98a, Ano98b, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01c, Ano01d, Ano01e, Ano01f, Ano02a, Ano02b, Ano02c, Ano02d, Ano03n, Ano03p, Ano03q, Ano04k, Ano04l, Ano04m, Ano04n, Ano05k, Ano05l, Ano05m, Ano05n, Ano06j, Ano06k, Ano06l, Ano06m, WCZ02, Ano03o, BJS14, CLZY15, LZWP03, PBG04]. index-based [CLZY15]. Indexing [BGSdVL98, CS98, CS00, DvLV08, Doe98, GFS04, MAP99, MLP97, Nis99, YC98, BZS16, BL04, JN09, MTC^{+}14, MYC^{+}14, Pha17, QT10, TKAK14]. indicators [CH06]. Individual [WPZ^{+}16, XFSC13]. individuals [CSV^{+}16]. Indoor [KM17, LYSK17, SDK^{+}24, SPQ^{+}17, ANHGS17, CGU11, DWB11, DPM14, DTL17, GC19, KPPK09, RRAR^{+}16, TS17, YHS^{+}20]. indoor-sports [KPPK09]. induced [YG17, ZSC^{+}23]. Induction [PC99, VBS^{+}04]. inductive [SHSJ23]. Industrial [SOJ^{+}95, LRFF24, RFF23, ZZZ06]. industrial-like [RFF23]. inexact [CFM^{+}23]. inextensible [BBH14]. Inference [AS17a, JvdBS99, SB95, WKI^{+}16, BBK14, BCA16, CKP^{+}19, GF15, Ham05, HHM^{+}16, JNLG15, PBW14, SCC17, WKP13, WW16]. Inferring [KMB97, LCG^{+}24, OGH04, KRK11]. Infimal [GY19]. Inflating [CM95]. Influence [HFKN97, BGPD09, GZP05]. influences [LWLP23]. influences-based [LWLP23]. Information [BEGB13, Boo97, CM97, HB98a, Hob00, JYL23, PMV00, SB02, ZWB^{+}22, ALY^{+}22, BKPS15, BSZ^{+}21, CSY08, EF14, FWLQ23, GH08, Hei04, KK07, KT07, LWZC14, LZZ^{+}21, LL12, QSXS23, SPC^{+}15, SKU^{+}09, WSSS13, WHGZ20, WWWF23, ZRZ^{+}24, ZYT10, ZZG^{+}24, ZWY14]. Information-Based [PMV00]. Information-theoretic [BEGB13, WSSS13]. Informative [BMvT^{+}19, DL10, JZZM23]. informed [BCHR24, JNLG15]. Infrared [FWLQ23, KH23, MZ20, QCL^{+}23, WB15, BBC^{+}07, CWL^{+}23, DZL07, EB13, FHZX23, GFY^{+}14, GZL^{+}23, HASS10, JYL23, KHA^{+}05, SRO^{+}19, SSN03, TAK^{+}22, XGT^{+}22, ZLFH23]. infrared-enabled [SRO^{+}19]. infrequently [PK18]. inherently [BMX22]. inheritance [LLWX24]. inhibition [ZHL^{+}20]. inhomogeneity [MUS06]. Inhomogeneous [GSP02, YHN11, KSHE20]. Initial [HSSB98]. Initialization [CYES00, NFSK97, SKSR08]. initialized [VBVB19]. initiative [MLK21]. Injection [LZQL24]. inlier [HWL^{+}22]. inpaint [UJ22]. inpainting [BR12, BABB19, CHSV08, CXYZ24, GCC24, JKW^{+}21, JLY^{+}17, QBZ21, UJ22]. Inscribed [BM98]. inscriptions [PRG^{+}14]. insensitive [BWL04, GJ10, NB10, PV06]. insertion [YJC^{+}09]. Inspection [COW98, MG95, MEDT96, ME98b, NJ95, SOJ^{+}95, TG95a, TG95b, LA11]. inspired [BCMR16, BC10, BCDH10, BEK18, EF14, EK12, HL13, KFRD^{+}18, MNMK16, MFG10, SM24, SVA^{+}22]. Instabilities [ASZ99b]. instability [SXY^{+}23]. Instance [ABJ^{+}21, BHA24, LYX^{+}21, WPZ^{+}18, BYJG23, FBF08, GLG22, GKH^{+}21, HWG21, KLO20, PHH^{+}15, YGC13, ZZ20, ZS19]. instance-aware [KLO20]. Instance-guided [BHA24]. Instance-level [LYX^{+}21, BYJG23, GKH^{+}21]. instances [MT16]. instantaneous [PV06]. Instantiating [WRH97]. instrumental [BKPS15]. Integrability [FW97, KS03]. integral [CYG16]. Integrated [BL09, LD98, SA95, VZP^{+}09, ASFP03, CNO^{+}16, GGP23, PBG04, SCS14, TMB12, TG95a]. Integrating [BZ99, DCTO97, MNE00, SSdVL06, TCZ^{+}12, NT10, Nis96, WLM^{+}14, eGZW07]. Integration [DL97, KMN11, MFJ95, Mas02, BCHR24, CUAT13, CJL06, DGG08, EDB12, dOSJVBS12, RFS03, SSL^{+}12, TLP^{+}17, VSP06]. Intelligent [SO07, YHS^{+}20, MFG10, RGA10, Tho10, VD10, Jon08]. intensities [WQY^{+}21]. Intensity [CW00, FDMA97, GJP96, LN98, ZU09, AS08b, CD13, HKWC14, JC06, RG16, RBA20, SM21, SKU^{+}09, SKSR08]. Intensity-Based [FDMA97]. intent [PSYZ13]. Inter [PZM^{+}21, GB08, JSRS08, TLY^{+}16]. inter-camera [JSRS08]. inter-muscular [TLY^{+}16]. Inter-related [PZM^{+}21]. Interacting [PDS^{+}07, JBC08, KPPK09, PA06]. Interaction [LLZ23, ZXK02, DLMC16, EK12, FR11, GZX^{+}23, HSH07, JS07, JZWD16, JRBD^{+}15, KPKH07, LRFF24, LXFM16, MdBJG15, NMAL23, PYS03, RKKK22, SA04, SVSM15, TMM16, TNO24, WHC14, ZSSF16, ZSC^{+}23, CEA16]. interactions [PT08, SP19, TBC^{+}21, ZNG^{+}13]. Interactive [BB95, GK95, MBKB02, PMG^{+}23, PZV13, QTLP22, VGSMN16, BCNS15, CG04, DWB11, FN14, GRMH19, GML16, HSS^{+}16, MO11, MM05, SBS04, THL03, WWH07, WG23, WWLV11, dMFU10]. Interactively [PC99]. Interactivity [JYL23]. interconnected [PBW14]. Interdigital [MKF15]. Interdisciplinary [MST00]. interest [BL20, CHMG12, GG09, ILRB04, KL10]. interest-based [ILRB04]. interface [NLM05, RRAR^{+}16]. interfaces [MĆK09]. interference [SRO^{+}19]. interferometric [WB11]. Interframe [AM01]. intermediate [SJB20, YDP^{+}20]. intermediate-layer [SJB20]. International [Ano96d]. Internet [WL15]. interpolated [EH21, TVE^{+}16, ZS11]. Interpolation [AM01, BS96, GL98, PMV00, FWG18, Kim04, SBB18, TY22]. interpretability [OVJ^{+}21]. Interpretation [DUC97, DTG96, HB98a, MS00, Mun95, OMLL98, SB00, Ste01, TN07, ARARCE11, BC10, KK07, LWH03, SM06, SCS14, VZP^{+}09, XP11]. interpretations [OTO06]. Interval [VB16]. interventions [RDS24]. intra [ASFP03]. intra-surgical [ASFP03]. intraoperative [LPR^{+}03]. Intrinsic [DKG22, DAM12, AAB19, BLKG21, LC11]. introducing [EDX16]. Introduction [Ano95e, BS99b, CFS98, DFJL15, LLE^{+}09, NPBM22, BK15, BPQ15, GSST03, DCCL99, MT97]. intrusive [YC05]. Invariance [Chu02, SC00b, RBA20]. invariances [ZCWH23]. Invariant [DG01, GDIIHK11, KR98, KORC10, MPPG98, PEFM98, SSS13, VKP98, YWY^{+}16, ADGB16, AMCB20, AC09a, AKC11, ASCF13, ASF14, BT05, FB12, FCOK24, HWK^{+}21, HAT^{+}15, HMF10, LRF^{+}17, LSCM03, LGD16, MTV17, OMBH06, OBH04, OH04, Pun03, ROGT14, SCE04, SAC^{+}12, TS19, TD19, TVC09, WCZ^{+}07, WYC15, XSL^{+}23, XYZH11, YYH^{+}23, YLX^{+}18, ZSDK19, ZZL13]. Invariant-Based [KR98, VKP98]. Invariants [Che96, KPH02, NG98b, QV98, RW97, SLL01, BG09, GHML17, GBB98, HN95, MTVM04, PC05, WHL14, ZCF13]. Investigating [OVJ^{+}21]. Investigation [RWV95, LL12]. invisible [ZZS^{+}23]. Involving [KW00]. IP [ZIT^{+}13]. IP-driven [ZIT^{+}13]. IR [CFB05, LCP13, LLZ23, MNSK98]. Iris [BKK11, Far11, GRGB^{+}13, BHBF10, BHF08, ET15, HBF09, HBL^{+}11, LMP^{+}19, LDGS^{+}13, NFSD13, PS12, CJL06]. irises [HBL^{+}11]. irregular [GDIIHK11, KA12, VRKL13]. Irregularly [GSP01, PPT06, TN05]. irrelevant [GZL^{+}23]. Islamic [AGB^{+}15]. isointensity [TG95c]. Isolated [BBC00, NS98, Sup02]. Isolated-Object [BBC00]. Isolating [MGPF08]. isometric [BBH14, KY19, RB18, SB18]. isothetic [DBBB14]. Issue [Ano01k, Ano01l, Ano15o, ACW^{+}16, CFS98, DRDKE13, FKL^{+}16a, FHP01, KB98, MZL^{+}16, NPBM22, RFL02, SPQ^{+}17, WPZ^{+}16, Ano05j, BK15, BPS10, BPQ15, CA10, CKB10, DFJL15, FKL^{+}16b, FPDK12, FYH11, GHMT09, HMC10, HTEB11, HGSM11, JWDF05, Jon08, KPKH07, KLBP11, LBK10, LLE^{+}09, MPF07, MYK03, MYC^{+}14, NLW13, STV09, SST06, SMHH04, THL13, Tho10, ZZP^{+}16]. Iterative [CH99, CUSZ07, DH19, GM19, GSK02, ODD96, AYG23, CO16, CBTC23, HQN05, LTL^{+}23, LBNS09, SZB^{+}21, TMB12]. IterGANs [GM19]. IVIS [TG95a].

J [Ano95d, CV13]. jamming [WYW^{+}22]. January [Ano20p, Ano21p, Ano22p, Ano22q, Ano23q, Ano23r, Ano24l]. jersey [GLM17]. jigsaw [YSKL24]. Johansson [SGDP01]. Joining [NHK08]. Joint [CKP^{+}19, CLA^{+}17, GFY^{+}14, KDSF20, KGFP10, LG14, MS97a, MMA06, QV98, SM06, WZCG24, ZDLS13, ZLFH23, Gon09, HUF05, HDL^{+}20, HH19, JLD13, MSF^{+}17, NGR24, SCEvdH14, TLH22, YO11, ZZ07, ZEGEJ15]. Journal [BPQ15, Par16]. JPEG [CWC^{+}20]. JPEG2000 [BRSSAL11, TVLS08]. JSNet [CWC^{+}20]. July [Ano20q, Ano21q, Ano21r, Ano22r, Ano23s, Ano24m]. Junction [ÁB13, LL97b]. Junctions [Dem96, LM99a, BB04]. June [Ano20r, Ano21s, Ano22s, Ano23t, Ano24n].

Kalman [Ano06h, GKK05, YNCO11]. Kalman-particle [YNCO11]. Keeping [Gui99]. Kernel [LTY^{+}15, MIUS16, ZRL^{+}11, AYG23, BB13, BB15a, CKC14, DWC^{+}24, GBB^{+}18, GGMV08, GCPF08, JSC23, KSF16, LHSG15, LWLT17, LHG^{+}23, SPK14, WHM^{+}09, WDC^{+}24, YG17, ZLL^{+}24, ZCK09, DT10]. kernel-based [GCPF08, JSC23, ZCK09]. Kernel-edit [DT10]. kernel-predictability [GGMV08]. kernels [BPSV16, JBR08, KSL^{+}20, TBFJ15]. Key [ADDK99, LJC^{+}24, PR03, SVSM15, WZG^{+}24]. key-component [SVSM15]. keyframe [DZJB14]. keyframe-based [DZJB14]. keypoints [JZZM23]. Kinect [SLK15]. Kinematic [ZDF10]. kinematical [FLB06]. kinematics [NZH^{+}23]. Kinship [MK18, KKSH23, PMR17]. Kirchhoff [RH06]. knee [LPS^{+}11]. know [LC19]. Knowledge [CL97, DTG96, MZ21, OD99, AZP14, BCHR24, BIMD23, FM22, LPSK23, LXW^{+}17, LWLP23, QCL^{+}23, WZG^{+}24, WZCY22, XZX^{+}21, XP11, YSZ23]. Knowledge-Based [CL97, DTG96]. known [CSW^{+}24, STC^{+}16, WXZG18]. Korean [SHKP98].

L [Ano95d]. Label [BSH22, BBK14, CSLX16, DRTT24, DFSC20, GKPS15, JGM20, Kim15, LvdHK^{+}15, LLLW23, LLS^{+}23, LWLP23, MRH19, MSF^{+}17, SOL14, ST20, SHS^{+}23, TPT17, TL16, XYZ16, XWC^{+}23a, Zac18, ZZCL14]. label-efficient [DRTT24]. Labeled [CYG16, LZZ22, MFP^{+}20, SS17b, WDN^{+}12]. Labeling [YB95, CPC08, CCL04, EyGS11, GLMM16, HAM^{+}16, JLL13, Nic95, SMD^{+}08, SHS03, TLY^{+}16, WZQ^{+}24, DAL^{+}22]. Labelled [MRF96]. Labelling [GLR^{+}99, AHDM10, HQN05, SRS11, ZJW15]. labels [LLWZ21, SYPK13]. laboratory [TN08]. lacunarity [QSX17]. lags [FTT15]. LAMP [ZH04]. Land [CCPK16]. Land-Cover [CCPK16]. Landmark [CLZY15, TW98, DDLP10, FM22, GSS12, LSQL24, Mal21, RFS03, TLWT12, WL15, WR08, YDP^{+}20]. LandmarkBreaker [LSQL24]. Landmarks [HRS02, HS06, MLB^{+}18, MDM^{+}21, SSM06]. Lane [Gui99, Lee02, LY05, PLB16]. Lane-Departure [Lee02, LY05]. Language [BKMSR98, YLM^{+}17, CLA^{+}17, GFLW24, JHK24, KFN15, LBC^{+}21, LXW^{+}17, OTO06, SM22, TLP^{+}17, WCZ^{+}07, WXZ24, VM01]. Laplacian [DvLV08, GXC23]. Large [CEO18, CGR13, CL15, FPDK12, IZKB12, Mar07, PKvGS16, SSHP17, SA02, SPQ^{+}17, TTN17, ANHGS17, ABB^{+}23, BPC^{+}17, CGD^{+}23, CCPK16, CP20, CPS10, DLBG19, DWC^{+}24, FWG18, FTT15, GG20, GC19, GML16, GDCM17, HMCT22, HBH10, KSR^{+}12, KFN15, KON^{+}17, LLL^{+}15a, LHG^{+}23, MNL^{+}17, MPST08, MYC^{+}14, STC^{+}16, SP23b, TS17, TKAK14, WL15, YWZ11, YSS^{+}14, YC05, ZTH^{+}11]. Large-Scale [SPQ^{+}17, CEO18, FPDK12, IZKB12, SSHP17, ANHGS17, BPC^{+}17, CPS10, DLBG19, GDCM17, HMCT22, KON^{+}17, LLL^{+}15a, MNL^{+}17, SP23b, TS17, TKAK14, WL15, YWZ11, ZTH^{+}11]. Laser [CZZS07, FK09, ZG06, FRNS05]. Laser-based [CZZS07, FRNS05]. LASIESTA [CYG16]. lata [TLY^{+}16]. late [LDC^{+}13]. latent [CP21, MJ17, SAC^{+}12, WZX^{+}14, YZX^{+}22, ZG10]. lateral [ZHL^{+}20]. Lattice [Car96]. Lattices [BNG02, Ang07]. Laurent [Ano95d]. layer [HH19, LWZP17, MML^{+}16a, SJB20, SBD22, XW16]. Layered [OGH04, YK24, ZH04]. layering [CLZZ13]. Layers [MDL^{+}23, CKS^{+}05]. Layout [Hob00, ES06, KM17, NHH14]. Lazy [KBAS16, LK03]. LBO [MIP16]. LBO-Shape [MIP16]. LBPE [LY05]. LDA [TAC21]. Leading [Lin02]. leaf [KT15, LZD^{+}14, NHK08]. Learn [BRET19, GM19, LHL^{+}21, MST16, VKL18, DLMC16]. Learnable [LGD16, HKL^{+}24, WML21, WGGH24]. Learned [KP00, NMP97, GCT^{+}14, LLG^{+}24, MLB^{+}18, TMQM13, ZZRC15]. learners [CWO^{+}11]. Learning [BBC00, BYJG23, BCC16, COW98, CWH^{+}13, CGL^{+}21, CHL21, CKLP09, CZ18, DFH^{+}22, DC00b, DH19, FZ20, FFFP07, FO18, GJH01, GKL^{+}17, GK95, HDZR24, HH19, KN99, KLO20, KC22, KSF16, LHL^{+}21, LJC^{+}24, LZTX24, LYSS12, LLL15b, LWSC16, LML^{+}23, LXW^{+}24, MYYY17, NLW^{+}17, PSR08, PSYZ13, PBQ99, PSB^{+}24, PTM20, PMCN22, RAHT11, RKL^{+}18, SFK24, SM22, SA15, SCvW11, ŠSJ^{+}20, SC98, TMN06, USKB10, VKL18, WKI^{+}16, WLL^{+}22a, WLXL24, XYZH11, XYZ16, XYW11, XSZ^{+}20, YYH^{+}23, ZWZ^{+}16, ZCWH23, ABN^{+}20, AYG23, BIG^{+}23, BCHR24, BFMW23, BSMK13, BAM16, BPCT22, BIMD23, Cha24, CL15, CCPK16, CP21, CC11, CTH20, CZHT15, CMH13, CFM^{+}13, DFSC20, DLBG19, DGRS22, DWL19, DD11b, DFLS23, EKY08, EXP^{+}20, EL07, EB13, EOPS22, FHZX23, FDC^{+}19, FKS10, FLHK08, FCOK24, GBB^{+}18, GB17, GLG22, GCPF08, GYW^{+}22, GLZF23, GZ19, Gwa17, HRC16, HOH^{+}07]. learning [HBL^{+}17, HWW23, HWT^{+}24, IT15, JRS21, JPN^{+}22, JLZ23, JRAJ17, JYL23, JLM22, KKRK23, KdRM^{+}23, KG14, KDSF20, KSF19, KRG17, KOC17, LBP23, LHSG15, LPC^{+}20, LLS21b, LZZ22, LLS^{+}23, LZ24, LDCX24, LTL^{+}23, LYD24, LCL^{+}17, ML13, MSV^{+}20, Mah16, MK18, MNL^{+}17, MHX19, MFSB23b, MFSB23a, MPM16, MPT21, MdOBA19, MZ21, MAK^{+}17, MFP^{+}20, MP20, MTR^{+}23, NWNT17, NNN^{+}22, NZH^{+}23, OGH04, OTAH20, PWSvdH17, PHHL23, PS22, PKH23, PZM^{+}21, PBD20, PLKP23, PD23, PK18, RCLS19, RSBA23, RL13, SB18, SP23b, SRB21, SXZZ24, SIRS21, TS24, TSL14, TPDP20, TCM18, TBC^{+}21, TA11, UFK20, VPL23, VGSMN16, WRKP05, WS08, WKP13, WLW^{+}16, WLO^{+}18, WLZ23, WTZ24, WM20, WHY^{+}23, WKT22, WZCG24, XZX^{+}21, XSL^{+}24, XST04, XSQZ15, XW16, XYRS17, YGJ^{+}20, YGC13, YSS^{+}14, YXW^{+}24, YGC15, ZSK^{+}23, ZdCR^{+}24, ZP18, ZTGL18, ZZ20, ZSL^{+}24, ZRKZ^{+}11]. learning [ZSG^{+}20, ZZHZ23, dSdSF^{+}12, RG16, WPZ^{+}18]. Learning-based [TMN06, AYG23, Cha24, CTH20, ML13, MSV^{+}20, NZH^{+}23, SRB21]. learnt [CGH08]. Least [ADC19, FM99, GSV05, HLKK19, MP09b, ZZ10]. Least-Squares [FM99, ADC19, GSV05]. leaves [CTM^{+}13]. Left [BMB^{+}17, WSKH13, WWJ13b]. Left/right [BMB^{+}17]. Legal [KABP98]. legend [Ano17j, Ano17k, Ano17l, Ano18k]. Legendre [KP97]. LeMéHauté [Ano95d]. Length [GJH01, Kis96b, LL97b, Che08, Kle13, SGH07, SCCP05]. lens [dSEdSPdMF24, WHL14]. lenses [BHBF10]. lesions [ARC14]. less [Pen15]. Level [DPB00, DG01, KSKB95, KB95b, LLSV00, ME98b, PA00, ZOMK00, AA20, AZ15, BC10, BCDH10, BB03, BYJG23, CICN22, CU11, DFJL15, DGC12, Dem05, DCS05, EPH^{+}21, FPC^{+}08, GKH^{+}21, HWZ16, HGP15, JLY^{+}17, KK13, KHG22, KYM13, KS04, LBC^{+}21, LYX^{+}21, LFL08, LGL15, MMV06, NLW^{+}17, PSE^{+}11, PD05, RMN^{+}17, STO17, SM06, SB22, WZ04, WLZM20, YWL^{+}22, YFF^{+}23, ZYT10, ZJW15]. Level-Set [LLSV00, FPC^{+}08]. levelings [AHM17]. levels [FKS10, SSdVL06]. levelsets [TRG^{+}13]. leverage [KH23]. Leveraging [KTV17, MSI10, WPI^{+}16]. LHS [SJ15a]. Libraries [DCCL99]. license [KKP24, LYW^{+}24]. LIDAR [GDCM17, SPT^{+}18, SO07, ALY^{+}22, BZP^{+}23, BABB19]. LiDAR-camera [ALY^{+}22]. LiDARTouch [BZP^{+}23]. lie [SL16b]. lifelogs [WSY^{+}16]. Ligature [ASZ99b]. Light [CVP10, LZ97a, OD97, OD01, WZCY22, XMN^{+}15, AZP14, BHSD^{+}13, CF07, CSD^{+}24, CFB05, CMZ24, CMD06, DZQ24, DWC16, Dre96, HASS10, HSTL24, JRH24, KHR^{+}16, LF08, LWFL24, LC19, MC20, MHL14, RBA20, SLK15, SBB18, SW13, SF16, TMNM09, WLZ23, WNH05, YHS95, ZSL^{+}16, ZHZ17]. light-based [HSTL24]. light-field [CMD06]. Light-weight [CVP10, WZCY22]. Lighting [Bic98, GJ10, LCT09, LC14, MC20, ZJ05]. lights [MAG^{+}16]. Lightweight [CMZ24, LWW^{+}21, LZB^{+}23, LCLH18, ZKLZ23, XSZ^{+}20, ZHL^{+}20]. like [DAM12, RFF23, XHJF12]. Likelihood [CHRM96, HH07, KNL15]. likelihoods [JPP^{+}14]. Limb [UZC97]. Limb/Terminator [UZC97]. Limbs [LRD99]. Limited [SMD^{+}08, CD10]. limits [HUF05, PV15]. Line [AHD98, CA97, CH99, DLHT99, GBB98, JV97, JB99, KB00, KP00, LD98, PKP97, PLL00, Rob96b, SP97a, SM97, Tsa96, ABN^{+}20, AAB19, BAPXH16, BCLNG18, CDT11, FS03, HMB17, KM17, NDO09, PYWZ17, PBD20, PZC17, RL13, Sha06, SW17, WXC20, XSK15, YGH11, ZRRK18, ZS11]. Line-Drawing [SP97a]. line-pairs [ZRRK18]. line-scan [AAB19]. Linear [AM01, BS96, BEPW00, Jac01, NN04, PRK19, SHS03, WZWT99, AC09b, AM15, Bar05, BBK15, CCL04, CSS13b, CO16, CP20, GTP18, ITNP12, KL07, KORC10, LY05, LDH^{+}14, MM21, PXTZ14, PL08, PZC17, QAB^{+}11, ZZCL14]. Linear-Time [WZWT99, SHS03, CCL04]. Lines [GL97, JvdBS99, KHB01, MGK00, MAM97, SLL01, BA06, BS05, LJC^{+}24, Sch06, Ste13, WZWH16, GOF^{+}15]. lingual [WHN08]. Linguistic [ALK^{+}09]. linguistics [JN09]. Link [GZY23]. linked [AKC11]. Linking [KVdG^{+}97]. Lip [LmCT16, CZ18, JB23, NN18]. LIP-signature [NN18]. Literature [Ros00a, SBIK16]. live [KK15]. living [BKPS15, YG16, YG17]. LKDA [LHG^{+}23]. LKDA-GAN [LHG^{+}23]. LMMSE [dLAH07]. Lmser [LZTX24]. Lmser-pix2seq [LZTX24]. lobe [YSL11]. Lobula [MAY^{+}10]. Local [GBB98, KP00, LCSL07, LS09, Mil99, MB11, PA00, RRL20, SGMC15, SZL^{+}23, SKVS13, TG11, TGQ23, TS00b, TBZ^{+}24, VNNB14, WTBdB15, YSX^{+}19, ZCL99, kCE^{+}18, BCM13, BFMW23, BB15b, BG09, CLZY15, CH06, CHC11, CK09, ESS10, FBK16, GKPS15, GCFMT12, HBG13, HSJS10, JBR08, KYYC14, KKSC23, LPS^{+}11, LLF18, LZS16, LXW^{+}23, MML^{+}16a, MdBJG15, MTP21, PXTZ14, PV06, PG13, PTE12, PMCN22, REF15, RLB17, Sah05, SBB18, SJ15a, SW17, SHS03, TLP^{+}17, TCZ^{+}12, TS11, TT16, WPS03, WZWL24, WYX^{+}16, WHGZ20, WWWF23, XYW11, YZT^{+}13, YGC13, YZX^{+}17, ZZL13, ZC19, ZCLX20, RK11, SJ15a]. local-aware [LXW^{+}23]. local-global [MML^{+}16a]. local-ternary-pattern [WYX^{+}16]. localisation [AW09, CGHTK16, KACR^{+}23]. Locality [BGE^{+}17]. Localization [CYES00, GMZ^{+}22, HR99, LSB^{+}00, RAH97, STC^{+}16, BBSD15, BDS12, BYK^{+}18, CGD^{+}23, CLZY15, EDJ^{+}20, FWL^{+}20, GYW^{+}22, JLD13, KA12, KMBH09, LTL^{+}23, LYSK17, MLB^{+}18, MN06, MDM^{+}21, NHH14, RAC^{+}13, ŠRDC09, SM22, SBN^{+}24, SIT07, WB15, WR08, WXZ24, ZKSV18, ZZSD21]. localizations [WLM^{+}14]. Localized [SB00, XFSC13]. Localizing [GF15, SAL16, MAL10, TSD17]. Locally [CSD^{+}24, FLHK08, SKS^{+}22, KL07, LvdHK^{+}15, LZD^{+}14, LLC11, PK05, dCCP12]. Locally-Aligned [CSD^{+}24]. Locate [HdVL99, CHL21]. Locating [Kou03, SZ07, CCF17]. Location [AW98, FTT15, Shi99, PBG04, SZ03, SM13b, WCF10, XWDL23]. loci [SWS11]. LocoGAN [SKS^{+}22]. locomotion [LE09]. LOFReg [KdRM^{+}23]. Log [MGMS01, Mas09, Sch06, SCS14, TP05, GBB^{+}18]. Log-Euclidean [GBB^{+}18]. Log-Polar [MGMS01, Mas09, Sch06, SCS14, TP05]. logarithm [Hu11]. Logic [MCPB00, ALK^{+}09, BKPS15, XP11]. logo [PA10b, SGZ21]. Logotype [Spi98]. loin [CCR^{+}05]. Long [NB20, TKL21, CRCM16, DFP23, GBF12, MBCJ17, PA10a, TTN17, WHL^{+}21, YAK^{+}08]. long-range [DFP23]. long-term [CRCM16, MBCJ17, PA10a]. longer [CRCM16]. Look [DAZ^{+}17, AB18, CL17, LJJY24]. Lookahead [JRS21]. Looking [BCC^{+}18]. Looming [RJ00]. Loop [SBK^{+}99, WWLV11]. Loss [HH19, BRPC17, DFSC20, EPH^{+}21, GL24, KDSF20, MDK^{+}24, MP20, SKA23, SXY^{+}23, WGZL20, XWDL23]. lossy [CWC^{+}20, YWMS08]. Loveparade [KB12]. Low [AA20, ASVO12, DPB00, LN10, WLZ23, ARFF18, BCDH10, CSD^{+}24, CSS^{+}13a, DZQ24, DGC12, Dem05, ED16, GXC23, GF15, JRH24, KHR^{+}16, KMBH09, KACR^{+}23, LBC^{+}21, LHY14, LWFL24, LGL15, LmCT16, LC19, MHAF13, RAC^{+}13, SZ16, TS24, WZ04, YFDA17, ZXC^{+}20, ZLL^{+}14, ZLZH17, ZD18, ZMM^{+}22, ZZ10, ZYT10]. low- [ZYT10]. low-data [KACR^{+}23]. Low-dimensional [ASVO12, ZMM^{+}22]. low-grade [RAC^{+}13]. Low-Level [DPB00, AA20, DGC12, Dem05, LBC^{+}21, WZ04]. Low-light [WLZ23, CSD^{+}24, DZQ24, JRH24, LWFL24, LC19]. low-rank [ARFF18, ED16, GXC23, KHR^{+}16, SZ16, TS24, ZXC^{+}20, ZLL^{+}14, ZLZH17, ZD18, ZZ10]. Low-resolution [LN10]. Lowe [ACB98]. Lower [Zha97, JB23]. LSS [TB13]. LSS-based [TB13]. LSTM [BSZ^{+}21, JVD^{+}20]. LSTMs [SKT18]. Luca [Ano01m]. Lucchese [Ano01m]. luggage [DETE17]. luminance [dLAH07]. Lungs [LSB^{+}00].

M2FINet [LLZ23]. Ma [Loh10]. Machine [Ano96a, BD02, FHSKP13, Lee02, Boy04, NWJ15, YHS95, YG17]. machines [CMBP09, CEO18, SB13]. macro [SOK16]. macro-micro [SOK16]. Macrofeature [NHH14]. macula [QKH^{+}12]. made [ROGT14]. MAE [SAK^{+}24]. MAEDAY [SAK^{+}24]. Magnetic [RMFB02, CCR^{+}05]. magnification [YAK^{+}08]. magnitudes [LMDB11]. MAIN [ABJ^{+}21]. maintenance [ZJZY16]. make [ZZS^{+}23]. making [CKL18]. MAL [GZY23]. MAL-Net [GZY23]. malaria [TDK10]. Mallat [AM00]. mammograms [CSY08, SRP10]. mammography [BRSSAL11, LZ24, RC03]. man [ROGT14]. man-made [ROGT14]. Manage [SB95]. management [SPC^{+}15]. maneuvering [MC09a]. Manhattan [KM17]. Manifold [DH19, LY13, Pey09, AC09b, EL07, GFY^{+}14, LCP13, LWSC16, MdRNM15, SM13a, VPL23, YG17]. manifold-valued [YG17]. manifolds [AAASC11, GCT^{+}14, KG14, LHYK05, WS08, YG16, ZEGEJ15]. manipulation [DJF14, dSEdSPdMF24, GYCS21, GWCO11, SGA12]. manipulator [JZWD16]. manner [YXW^{+}24]. manual [BCNS15, KSG^{+}13]. manuscript [HSBS16]. Many [Lau97, DOSD11]. many-to-many [DOSD11]. Map [LK97, OMLL98, BI11, BB03, BR12, GMF14, JC06, KGC05, KORC10, LSC08, PY19, WLZ23, YWL^{+}22, ZGC20, CMBV04, DBZ07]. mapped [LLNC20]. Mapping [CGL98, SWYP00, ABK^{+}18, BZS16, BEK18, CKM11, DWL19, KKCK23, LJZ18, LZL^{+}17, OMW^{+}07, ŠRDC09]. Maps [DTG96, GSV00, HB98c, Jok98, KSKB95, OMLL98, Cou13, DSdlH^{+}11, DDLP10, GWT09, JBWK11, JRBD^{+}15, KIS17, LYKY19, LYSS12, MRH19, Mas09, PMC13, PCR^{+}04, SSL^{+}12, SY20, TÉSK11, TC11, WDN^{+}12]. March [Ano20s, Ano21t, Ano22t, Ano23u, Ano24o]. marching [HMA10]. Margin [CLZZ21, CGR13, CL15, GHZ^{+}13, KSR^{+}12, LLC11]. Margin-based [CLZZ21]. marginal [PRK19]. maritime [Cha24]. Markerless [KV06, SHK11, DMSM21, JBWK11]. Markov [PKH23, AMCB20, BP05, BCM06, CL17, GJH01, HPvB^{+}10, KABP98, MCPB00, MJPS16, NN13, PJW11, SGH07, VGR16, VMN16, WKP13, WB11, XQZL23]. Markovian [MCPB99, PC15, PCR^{+}04, RMFB02]. Mars [OMW^{+}07, SB13]. Mask [CLCZ23, TDZ^{+}20, ZS19]. mask-based [TDZ^{+}20]. masked [RCT14, SR23]. masks [LG24]. mass [CSY08, Dem05, LZ24]. massive [CACB17]. Match [GBB98, Shi99, TKV16, TBFJ15]. matches [DLS^{+}09, PXTZ14]. Matching [AM01, AG00, BR95, BDL^{+}06, COW98, CTF^{+}98, DC00a, GGR01, HB98b, IAP^{+}11, Jok98, KC99, Lai00, Mas02, NG98a, NMP97, PLL00, PC99, PM97, RH95, SHKP98, SA95, THT^{+}98, VKNK14, WYC15, WCH98, YS06, ARC14, AKC11, BZS08, BL09, Bre03, CM12, CDJM14, CK11, CALO20, CC07, CK09, CWLJ13, CFM^{+}23, CR03, DOSD11, DLV15, DSH04, Far11, Goh08, GS95, GDR04, HBG13, HQW^{+}12, HZW^{+}10, JKM07, JZZM23, KD10, KM17, KZ05, KMBH09, LLS21a, LLC13, LZLP10, LZC^{+}20, LS09, MHX19, MAL10, MMP15, MK24, OBH04, OH04, PD14, PLLL03, PFGG09, PMW05, PDTE06, QCL^{+}23, RDA^{+}15, SAS12, SZ03, SKH08, SW17, SBM^{+}06, SK15, SY11, TZY08, UBEP09, WPS03, WHGZ20, WXZ24, XHW09, YS09, YWY^{+}16, YYZL19, YK08, YW16, YSY^{+}18, ZP11, ZSY^{+}19, PE09, STLH08]. Matching-constrained [WYC15]. matching-recognizing [LLC13]. matchings [CKC14]. material [XYZ16]. Mathematical [Ano95d, BB13, BB15a]. Mathematics [Åst97, Col97, PEFM98, PRW97a, PRW97b]. matrices [Gol05, LPVM13, LL17]. Matrix [BGK98, CZZF97, LLTL14, SB98b, TI01, TZM98, ZL01, AO16, ARFF18, CLZZ21, GF15, KK15, LLL13, MSI10, ZZ10, CGD^{+}23]. MATTE [SvNW23]. matter [WZWL24]. matting [HKS06, JYX^{+}23, LWZP17, YS24, YZX^{+}20]. max [CGR13]. max-flow [ZSCP08]. max-margin [CGR13]. maximization [FWLQ23, SBPF17]. Maximizing [WCZ02]. Maximum [CHRM96, GHHX04, CKK^{+}12, LLC11, PW23, PYWZ17, She16]. Maxshift [TVLS08]. May [Ano20t, Ano21u, Ano24p]. MC [RPBK22]. MC-Calib [RPBK22]. MCMC [JSWL24]. mdBRIEF [UWH17]. MDC [WDC^{+}24]. MDC-Net [WDC^{+}24]. MDS [Mig12]. MDS-based [Mig12]. me [SL16b]. Mean [LLR10, MHMO09, ZLS^{+}13, HW06, MSR07, ZYS09]. means [BBC^{+}07, HS06, JLD12, LLF18, MJ11]. Measure [ALK99, APV99, KN11, LMRMJ08, MGW10, PDK96, PTM20, RBdDS14, RM06, Ros08, TH04, WDN^{+}12, YK08]. Measurement [OD02, SGK00, TI01, NN18, SJH17, XFSC13, ZZZ06]. measurements [ATG15, BHMB10, WLM^{+}14]. Measures [Neg96, RPTB01, SB98a, YYL96, Bor19, Bor22, BAP08, KY06, MM06, RKG03, SvdMH15, Got08]. Measuring [Car01, CK11, KT08, Ros99b, RŽ05, WHN08]. MECCANO [RFF23]. Mechanical [CLD96, LCD97, AAB19]. mechanism [GS08, HBZ^{+}24, WWXK24, YK24]. Mechanisms [YYL96, SM24]. media [FSI21, NHTG15]. Medial [SB98c, CLK09, CK11, PAK19, PCJ14, SWS11, MDFS11a]. median [FKV^{+}11]. Medical [AMGG^{+}16, Boo97, BM97, DUC97, MAM97, NLW13, SPK^{+}02, TK97, AH24, BK15, BCA16, CUAT13, EPH^{+}21, KLBP11, KSHE20, KSG^{+}13, MLB^{+}18, Mah16, MJ11, WPK09, YZT^{+}13]. medium [CSK22]. Meet [Ano15o, CICN22]. meets [KKRK23]. MEG [CSDNR17]. Mellin [DG01]. Membranes [Pen99]. Memory [NB20, ZCW^{+}23, JLM22, NKL24, WGGH24]. memory-based [WGGH24]. Memory-efficient [ZCW^{+}23]. merge [DWLW23, LK03]. Merging [BL00, BS00b, SCvW11]. MERLIN [DRTT24]. MERLIN-Seg [DRTT24]. Mesh [LHKC97, TGSH98, BSRV17, dOSJVBS12, MWTN04, RZZ23, SY10, SWMM22, TGQ23, TPT15, ZZC^{+}13]. Meshes [MKY01, Tan95, WH00, CL95, MSR07, RT14, VPP^{+}23, WTBdB15]. meshSIFT [SKVS13]. message [MTR^{+}23]. meta [BPCT22, TFL^{+}09, YST21]. meta-data [TFL^{+}09]. meta-learning [BPCT22]. metabolic [ACC^{+}16]. MetaVD [YST21]. Method [Cre99, HY98, KB95b, KB00, MY95, OD02, PM97, SRT01, TB99, ZOMK00, AAB19, AGB^{+}15, ACG^{+}09, AVC19, BLJ^{+}23, BYN^{+}04, Cha24, CSW^{+}24, CE17, DETE17, DMW10, Eva06, FL09, FWZ^{+}24, GYW^{+}22, HDS08, HMA10, JGM20, KK13, LSL^{+}18, LSQL24, Liu10, MCT10, MM21, MMP15, MJ17, NWJ15, PD14, PW06, PT15, RR06, RL13, RLMK15, SAS12, SSL^{+}12, SOL14, SCCP05, TM07, WGAD14, WWCZ15, WYX^{+}16, WHGZ20, XSK15, YHS^{+}20, YCL07, YZL^{+}21, ZLFH23, ZS11, ZCF13]. methodologies [TPT15]. Methodology [HSSB98, AC09a, DL10, LMRMJ08, LFMP13]. Methods [Car01, FKW98, HdVL99, NPBM22, RFC97, AYD^{+}18, BSALF18, BSH22, Bre03, BBH14, CCTCR09, CTH20, CMH13, CU11, DFS08, DSY10, EK14, GBB^{+}18, HNB04, JKW^{+}21, KLKF20, LLG^{+}14, LLL^{+}15a, MSR07, OEK08, PD05, PWWQ16, PS15, PBSG12, RN12, RDSF15, SCD11, WRB11, WTW^{+}17, XYZH11, YGJ^{+}20, YARL^{+}20, ZFG08, ZCK09, RC13]. Metric [BCP15, KK11, Por00, RG16, ARC14, ALIRT18, BZP^{+}23, CGU11, FLHK08, FK09, JRAJ17, KdRM^{+}23, LTAA23, LFLZ23, LFL08, MYYY17, MTG07, PWSvdH17, SMD^{+}08, SCvW11, WZWH16, ZZZ06]. metric-based [MTG07]. Metrically [KP00]. Metrics [Ste01, CS20, HSK23, KLKF20, SAV24]. MF [WHJK23]. MF-DFA [WHJK23]. MFMAM [CXYZ24]. MGRF [LGD16]. micro [SOK16, TDWH07, XFP^{+}16]. micro-expression [XFP^{+}16]. Microbathymetric [SWYP00]. micrographs [IT15]. microscopy [ZMCA05]. Microstructure [WH01]. Mid [DFJL15, PCJC98, KYM13, LGL15, NLW^{+}17, ZYT10]. Mid- [PCJC98, ZYT10]. Mid-level [DFJL15, KYM13, NLW^{+}17]. min [ZSCP08]. min-cut [ZSCP08]. min-cut/max-flow [ZSCP08]. minima [PV06]. Minimal [GYF18, NSEA13, IH15, KBJ^{+}10, LZZ22]. Minimal-delay [NSEA13]. minimization [LLY^{+}18, MAJ16, QDLB17, SE11, WAPB17]. Minimum [BÁRM23, LL97b, MRF96, CSMS14, Kle13, MEYD11, NBFG20, SCMS13, YYH^{+}23]. minimum-cost [MEYD11]. Minimum-Energy [MRF96]. minimum-length [Kle13]. Mining [TABK17, ZWZZ18, FSL24, GB17, GYWZ23, JYX^{+}23, PHY^{+}11, ZSY^{+}19]. Minutiae [UBEP09]. Minutiae-based [UBEP09]. MIRFLICKR [THL13]. MIRFLICKR/ImageCLEF [THL13]. mirror [LNS14, PA13, ACC^{+}16]. Missing [Jac01, MC09b, PWL^{+}23, ZZ10]. Mixed [SHKP98, LTY^{+}15, MBD^{+}22, MLK21, PV13]. Mixture [CTWH15, FWL^{+}20, MK01, CE14, CLO17, EKY08, EB13, FL09, HL23, JWG04, KLK14, VWMZ15, XQZL23, ZLY^{+}20, AQ09]. mixtures [KNL15, VKNK14]. MLESAC [TZ00]. MLGPnet [LSGY24]. mobile [CSGM^{+}24, DWC16, GLOC10, HSH07, MAG^{+}16, MLH13, SSHP17, ST10, ZKRH04]. Mobility [FKL^{+}16a, BVWS21]. MoCap [BB24]. modal [ABI^{+}04, BCF06, CA10, CM21, HBKG22, HKZ^{+}16, KLK^{+}16, LCL^{+}17, MLZRK24, MML^{+}16b, NT10, PS22, PWL^{+}23, PNSF21, PV14, RYLZ24, RKG03, VJ17]. modalities [CR18, LHJ^{+}09, SM22, WHN08]. Modality [LLZ23, AMGG^{+}16, LHG^{+}23, PWL^{+}23, YNZ^{+}19]. Modality-shared [LLZ23]. Modality-specific [LLZ23]. Mode [ED16, DAM12, HKM22, UFK20, UIK22, YZX^{+}22]. Model [BCA98, BR95, BS00b, CKB96, Car96, CM95, CFM^{+}23, CG04, CC16, FBS21, GPK99, GBB98, GL97, Gui99, HY98, Jur99, KABP98, KMA^{+}00, LZ97a, LK97, LHHC98, MS97a, MWLA99, Muk97, RH95, SK02, SB18, SMK02, SHE17, SLL01, SH08, SM97, TW98, TKDN16, VV02, WC99, WLI08, YC98, YB01, AC09b, AZN11, BAPXH16, BB16, BCMR16, BEK18, BvdHL^{+}13, BCM06, BPB13, BH12, CLZY15, CTM^{+}13, CUAT13, CE14, CL17, CP09, CLO17, CC03, CC96, DBF04, Dam08, DD11a, DPCA15, DH19, EyGS11, FMGA^{+}12, FFY^{+}04, FMS17, FAB12, FOCSB^{+}20, FO18, GFL^{+}19, GBY21, GF15, GBHS06, GHHX04, GPDR13, HL13, HH07, HSS^{+}16, HG11, HBL^{+}17, HKK08, JVD^{+}20, JSWL24, KHG22, KBMD15, KK07, KHH^{+}12, KNO^{+}09, LT05, LA11, Lhu23, LG17, LSL^{+}18, LWFL24, LYG07, LNS14, LBCA10, LN10, LPR^{+}03]. model [ML13, MML^{+}16a, MAY^{+}10, Mig12, MFP^{+}20, NAS^{+}17, PE09, PW23, PL07, PBW14, QSXS23, QTLP22, RH06, RB18, RF23, RLC^{+}11, STHBH18, SOL14, SOL16, SS17a, SLK23, SKH08, SKU^{+}09, SJ15b, SF16, SJH17, SYL^{+}24, SM13a, SFWG08, SZB^{+}21, TLB^{+}15, TLY^{+}16, VAWW10, VMN16, WB12, WMBY12, WCYS13, WWJ13b, WHS^{+}24, XHW09, XSL^{+}24, YZY11, ZZRC15, ZMM^{+}22, ZHZ17, AQ09, CTWH15, HH05]. Model-Based [HY98, KMA^{+}00, MS97a, SK02, SLL01, YC98, YB01, CFM^{+}23, CG04, CC16, SB18, SHE17, SH08, WLI08, AZN11, CTM^{+}13, FAB12, GBHS06, GHHX04, KK07, LBCA10, RB18, SF16]. Model-Driven [CKB96, SM97]. model-free [SS17a]. Model-image [FBS21]. Modelbase [SB98b]. Modeling [ACF00, CJC^{+}98, EK98, FPDK12, GA13, HF01, HFR06, JSRS08, LSB^{+}00, LB98, LSP^{+}16, LCZ^{+}16, Mas02, MKK02, MCPB00, NLW13, PF01, RWV95, SC00a, SL96, SPQ^{+}17, TS17, TDT12, TGSH98, WPI^{+}16, YB99, ZTH^{+}11, ZNG^{+}13, AAASC11, BN15, BCDH10, CLCO13, CD13, CSG^{+}03, ES04, FF09, FBK15, GHMT09, HKM22, HJZ16, KON^{+}17, LD24, MMP09, NWJ15, REF15, STO17, SCD11, SEFV15, SPK14, TKL21, TÉSK11, THL03, TA11, UFK20, UIK22, WY07, WKP13, XFP^{+}16, YJ16, YT13]. Modelled [HFKN97]. modelling [AAL22, HGSM11, KMN11, LRLB11, PZV13, SKBS13, TPD^{+}16, VWMZ15, VGR16, WX16]. Models [ACW^{+}16, BL98a, BD02, Dav97, DF01, DUC97, EFF98, FB97, GJH01, GSP02, GMT00, HB98a, IP98, KVdG^{+}97, LVW97, LK00, LT97, NFSK97, Nis97, Nis99, Pha01, SF95, SP97a, SRS11, SB00, TML00, TS01, TGSH98, WKI^{+}16, WRH97, YKA01, ÁB13, ARARCE11, BK15, BVVMMS15, BBCF20, BSH13, BF10, CGH08, CFCP11, CHSV08, CSS13b, CMD06, CTCG95, CNC03, DPRC17, DCH12, DB03, DSY10, ESS10, EB13, EK14, Eva06, FSL24, FFFP07, GKBW14, GCFMT12, GCC24, HRC16, HSK23, HYW^{+}24, JEF^{+}12, JNLG15, JBC08, JB15, KG14, KLK14, Kim15, KCM^{+}17, KDV16, LSD^{+}07, LSCK15, LG24, LGD16, LLWZ21, MMLC23, MGCS17, MJ11, MCB13, MMA06, MTP21, MSW15, NN13, OJRT08, Pec07, Pey09, QAB^{+}11, QWHW20, RB16, RDSF15, SEFV15, SI03, SVSM15, SKM06, SGH07, SPW15, SRHC13, TS16]. models [TVE^{+}16, UK12a, UFF06, VTRC14, VKL18, WPI^{+}16, XG08b, XQZL23, YSNiT14, ZKSV18, ZZC^{+}13, ZWZ^{+}16, ZDZ^{+}23, DGG08, TRG^{+}13]. modes [DLMC16, OGB14]. modification [Dre96]. modifications [CDIF14]. Modified [LLF18, GBB^{+}18, KK15, MAY^{+}10]. MODS [MMP15]. module [CXYZ24, JCLZ21, JZL^{+}23, QZY^{+}24, SVA^{+}22]. mold [MYZ^{+}24]. Moment [DPB00, MTVM04, GHML17, SM22]. Momental [NNBN20]. Moments [SC99, Dem05]. monitoring [ACC^{+}16, ESS10, HMEB07, HCC^{+}16]. Monocular [BZP^{+}23, BBH14, CTH20, CN95, GML^{+}21, SGDP01, WN99, WLD99, ÁB13, CC03, GKGM20, KM17, KLKF20, RSPD12, ROGT14, UFF06, ZZG^{+}24, dP10]. monocularly [RF23]. monotonic [HKWC14]. Monte [SOL14, SOL16]. morphable [GFL^{+}19]. morphing [XS04]. Morphological [Ang07, CNDS13, GHS95, Hei99, JC98, SH09, CE17, SW05]. Morphology [Ano95d, BB13, BB15a, GE08, XWC^{+}23a]. Morphometric [Boo97, Sah05]. Morse [AC07]. mosaic [AWK04, SP06]. mosaic-based [AWK04]. Mosaicing [LDD09, CPS10]. Mosaics [GSV00, AGB^{+}15]. Most [Ano12m, Ano13o, Ano07f, Ano08k]. Motion [ACLS98, AC99, AS09, BDVK10, BEPW00, Bri17, CSC96, DT96a, Dan97, DH00, DC98, DC00a, FD99, GB97, IF99, Jac01, KN03, KC99, Lin02, LHHC98, MNE00, MS97a, MG01, MS96b, NK00, Oli00, Oli01, Pen99, SA96, SP97b, SGDP01, SF97, SBZ97, TO99, TS01, VF96, WLD99, WF02, WD96, WY21, XL98, ACP16, AMN18, AS08a, ACG^{+}09, ABB^{+}23, BS05, BF07, BC10, BT05, BPC^{+}17, BW15, CG09, CMBV04, CFCP11, CMBP09, CT13, CRCM16, CSK22, DGC12, DMSM21, EMMV19, EF14, EH21, ED16, FDW21, FLB06, FB16, GZP05, GRCD18, GBHS06, GW07, GWT09, Gwa17, HSH07, HWZ^{+}23, HHG^{+}20, HMF10, HGP15, HRC09, HC13c, KBN12, KBWT16, KHK10, KYYC14, KC22, KL10, KRS14, LCSL07, LMRMJ08, Lhu08, LLS21b, LZWP03, LWH03, LYA13, MPF07, Mal21]. motion [MST16, MU11, MHK06, MP09b, NFM08, NT10, Neg12, NWJ15, NHZ^{+}22, OGB14, PD05, PW06, PT15, PV06, PRCP16, Pop07, QSXS23, QWHW20, RDA^{+}15, RLS06, RN12, RSPD12, ROGT14, SHE17, SOJ17, SKM06, SCS14, MNR18, TMQM13, TPD^{+}16, TPNP15, TYDH18, TGFF15, TP05, TR09, TLMT^{+}05, UK12a, UFF06, VSP06, WLO^{+}18, WRB06, WS06, XYW11, XYRS17, XZQJ21, XWC^{+}23b, YWZ11, YNZ^{+}19, YS06, YNCO11, YC05, YSD03, YR06, YG16, ZDLS13, ZT09, LY13]. motion-aware [XWC^{+}23b]. Motion-Based [NK00, WF02, EH21, KL10]. motion-blurred [CG09]. Motion-Egomotion [DH00]. Motion-Model-Based [LHHC98]. Motions [BA96, Bar05, KV06, RRR11, RAP16].

Motivated [BL98a]. mounted [JZWD16]. mouse [TTH07]. Movement [BL01, Gav99, HF01, HFR06, ITNP12, LSP^{+}16, PQML11, WS08, MAY^{+}10]. Movements [KS95, SFWG08]. moves [CLL17, Zac18]. movies [SZ03]. Moving [SMK02, WD96, AMNCM16, BP09, CYC10, CCYC12, CYG16, DMAD17, HLKK19, JKM07, MP14, MOT17, OCVV04, QC04, SZ16, WZT13, XWC^{+}23b, ZY14]. MPEG [ADDK99]. MPM [CMBV04]. MR [BvdHL^{+}13, CFYU12, DCS05, HRS02, LPS^{+}11, LSB^{+}00, ZHL^{+}20, ZU09]. MR-image [CFYU12]. MRF [BBK14, GJP96, KL11, SKH08]. MRFs [AKC11, KTP08]. MRI [GPDR13, JSWL24, MAK^{+}17, MPPP14, RAH97, WSKH13, WWJ13b, ZRL^{+}11]. MSR [CLFH22]. MSR-Video [CLFH22]. MTCD [TAK^{+}22]. MTRNet [TDZ^{+}20]. Multi [ADR16, ABJ^{+}21, AMMV99, BDS12, BF10, CPT07, CRCM16, CPS10, GXC23, GMZ^{+}22, Gwa17, HH19, HKZ^{+}16, HJZ16, ITNP12, KK13, KCM^{+}17, KLK^{+}16, LS08, LLL^{+}20, LJC^{+}23, LSGY24, MFB11, Pat13, PNSF21, Pen03, PLYW21, PMC13, QSXS23, QBZ21, SKA23, Sav23, SZL^{+}23, SCL13, SvNW23, SGZ21, VBT19, WJ07, WZY13, WDC^{+}24, WZQ^{+}24, XSZ^{+}20, YK24, YFF^{+}23, YS24, ZSDK19, ZKLZ23, ACP16, ABI^{+}04, Ano06h, AKC11, ABB^{+}23, AS23, BAPXH16, BYR17, BIG^{+}23, BKK11, BCC^{+}21, BHA24, BSMK13, BBK14, BCF06, BG16, CICN22, CSGM^{+}24, CSDNR17, CA10, CDJM14, CPP^{+}11, CM21, CD10, CWO^{+}11, CSLX16, CZS^{+}20, CXYZ24, Che24, CLL^{+}14a, CACB17, DR04, DPRC17, DD11b, DWL^{+}24, DCS05, DWLW23, DSK^{+}20, EXP^{+}20, EOPS22, FBF08, FWZ^{+}24, FN14, FSI21, GKK05, GCEC07, GBVDC18, GTMR23, GML^{+}21, HWG21, HWZ^{+}23, HDG^{+}14, HGP15, HSHA20, HC13c, IJDAB13]. multi [JRAJ17, JCLZ21, JB15, JHV19, KD10, Kim15, KSF19, KW12, KL10, LTAA23, LWY^{+}17, LvdHK^{+}15, LHSG15, LKZ20, LWLC22, LJC^{+}24, LG14, LJJY24, LZmC^{+}17, LLJ^{+}23, LZZ^{+}23b, LZS16, LBNS09, LYSK17, MNL^{+}17, MSW15, MCM^{+}17, MML^{+}16b, MB11, NAS^{+}17, NN13, NT10, NL17, NZH^{+}23, OSY18, PLJS14, PLKP23, RPBK22, RYLZ24, RM03, RXDS22, RB16, RCTV12, RKG03, RTM^{+}17, SKLM22, SSL^{+}12, SOL14, SB22, SFK24, SOJ17, ST20, SHS^{+}23, SY23, TPT17, UM05, VRKL13, VMN16, WPQ20, WLL^{+}22a, WTZ24, WDC^{+}20, WCYS13, XYZ16, XSL^{+}24, XWC^{+}23a, YWZ11, YGC13, YWY^{+}16, YJ16, YCKA10, ZRL^{+}11, ZZRC15, ZWLH24, ZYZ24, ZCW^{+}23, ZH04, ZNG^{+}13, LLLW23]. Multi-agent [KK13, GBVDC18]. multi-atlas [LvdHK^{+}15]. Multi-Attention [ABJ^{+}21]. Multi-camera [MFB11, CA10, DPRC17, GTMR23, HC13c, JB15, KD10, RPBK22, RCTV12, YCKA10]. multi-cameras [NL17]. multi-channel [IJDAB13, NN13]. Multi-class [Pen03, AS23, KSF19, MNL^{+}17, PLJS14]. multi-colored [DR04, OSY18]. multi-constrained [SOJ17]. multi-core [KL10]. multi-dimensional [ACP16]. Multi-domain [WDC^{+}24]. multi-exit [LJJY24]. multi-expert [CSDNR17]. Multi-exposure [LLL^{+}20]. Multi-face [ADR16]. multi-feature [CWO^{+}11, CZS^{+}20]. Multi-focus [PLYW21, FWZ^{+}24, LTAA23]. multi-future [BCC^{+}21]. multi-grained [LLJ^{+}23]. Multi-granularity [LSGY24, BHA24, DWL^{+}24, LLLW23]. multi-graph [CLL^{+}14a]. Multi-guided-based [YS24]. multi-head [WPQ20, ZWLH24]. Multi-human [GMZ^{+}22]. multi-image [Che24]. multi-instance [FBF08, YGC13]. multi-Kalman [Ano06h, GKK05]. multi-kernel [LHSG15]. multi-label [BBK14, CSLX16, Kim15, SOL14, ST20, SHS^{+}23, TPT17, XYZ16, XWC^{+}23a]. Multi-layer [HH19]. Multi-layered [YK24]. Multi-level [YFF^{+}23, CICN22, SB22]. Multi-modal [HKZ^{+}16, KLK^{+}16, PNSF21, ABI^{+}04, BCF06, CA10, CM21, MML^{+}16b, NT10, RYLZ24, RKG03]. multi-module [JCLZ21]. Multi-object [Gwa17, HJZ16, LJC^{+}23, SCL13, EOPS22, LJC^{+}24, MCM^{+}17, NAS^{+}17, RB16, SKLM22, WLL^{+}22a, WDC^{+}20, ZNG^{+}13]. multi-output [DSK^{+}20]. Multi-patch [QSXS23]. Multi-person [VBT19, BAPXH16, GML^{+}21, HSHA20, LWLC22, LG14, YJ16]. Multi-perspective [SGZ21, CPT07, ZH04]. multi-phase [DCS05, IJDAB13]. multi-prior [HWG21]. Multi-reference [CRCM16]. multi-resolution [AKC11, LKZ20]. Multi-resolutive [Pat13]. Multi-Scale [XSZ^{+}20, ZKLZ23, AMMV99, BDS12, LS08, QBZ21, SZL^{+}23, BKK11, CDJM14, CXYZ24, EXP^{+}20, LZmC^{+}17, LBNS09, MSW15, QSXS23, RXDS22, RTM^{+}17, SSL^{+}12, SvNW23, VRKL13, YWY^{+}16, ZYZ24, ZCW^{+}23]. multi-scale/irregular [VRKL13]. multi-scan [CACB17]. multi-scene [SFK24]. multi-sensored [CD10]. Multi-spectral [CPT07, WZY13, GCEC07, ZRL^{+}11]. multi-stage [HWZ^{+}23, LZZ^{+}23b, WTZ24]. multi-start [FN14]. multi-structure [LWY^{+}17, WCYS13]. multi-style [DWLW23]. multi-subject [ABB^{+}23]. multi-subspace [DD11b]. Multi-target [PMC13, BG16, CSGM^{+}24, CSLX16, KW12, UM05, YCKA10, ZZRC15]. Multi-Task [ZSDK19, SvNW23, BIG^{+}23, BSMK13, JRAJ17, XSL^{+}24]. multi-template [FN14]. Multi-timescale [Sav23, SY23]. multi-tracker [VMN16]. multi-user [YWZ11]. Multi-view [BF10, CPS10, GXC23, ITNP12, KCM^{+}17, WJ07, WZQ^{+}24, BYR17, CPP^{+}11, FSI21, HDG^{+}14, LZS16, LYSK17, MB11, NZH^{+}23, RM03]. multi-view-CNN [PLKP23]. multibox [YWM19]. multibranching [MWL^{+}24]. Multicamera [Mur95, TWW14, TA11]. multichannel [RDM^{+}11]. Multicolored [MS00]. multicuts [KSRS16]. multidimensional [BVVMMS15, MJ11]. Multifactor [PQML11]. Multifingered [SKOS95]. Multiframe [TO99, EH21]. multigranularity [HLL^{+}23]. Multigrid [CLL14b]. multilabel [CLL17]. multilayer [TRPD20]. multilayered [KK07]. Multilevel [OMLL98, HDS08, KMT11]. Multilocal [LLSV00]. multimedia [MYC^{+}14, YSS^{+}14, STLH08]. Multimodal [JS07, LBC^{+}21, LDC^{+}13, MKK02, PY08, YKA01, FDC^{+}19, JZWD16, KT07, LRFF24, LLL^{+}15a, LDH^{+}15, LXW^{+}17, MK24, MLK21, NNS^{+}18, OH05, RFF23, WZT13]. Multimodel [ZHS^{+}24]. Multiocular [LRD99]. Multipart [BLP95]. multipath [OSM16]. Multiperson [IB01]. Multiphase [WSKH13, MPPP14, NHSC09]. Multiple [BT17, BA96, CFM02, CM95, CCS01, CJC^{+}98, CM99b, EFF98, FW97, FMR01, GLG22, GK95, HH12, JRAJ17, Jok98, Kim15, LV96, MFJ95, MY95, Mas02, MS97b, MKY01, Nis95, OD99, OD02, PA10a, SU01a, SU01b, SCS99, Spi98, SA95, WD96, WH01, WB01, WPZ^{+}18, YSD03, ZLHJ18, AJ23, AZP14, BYR17, BL09, BPB13, BYK^{+}18, CKM11, CHH09, CWLY22, CW15, CYP^{+}10, CS10, CCF17, CH11, CKP^{+}19, CUSZ07, CZZS07, DFP23, Gol05, HKHE14, JRH03, JBC08, KV06, KN03, KN04, KHK10, KFN15, KEG15, KPPK09, KON^{+}17, LF08, LLR10, LWLT17, LHJ^{+}09, LML^{+}23, Mah16, MMV06, MMA06, Mas09, MOB14, MCL16, MGS15, MBCJ17, OGH04, PA06, PT08, PD11, PTK24, ROJX09, RZZ23, SPC^{+}15, SW17, SSdVL06, SBH^{+}17, SYPK13, SH08, SCEvdH14, TB13, TRG^{+}13, UK12a, VGSMN16, WRKP05, WDB12, WSJ15, WHN08]. multiple [WB16, XST04, YQL^{+}23, YSS^{+}14, YSL11, dSdSF^{+}12, KBKS18]. Multiple-Attribute [GK95]. Multiple-concept [Kim15]. multiple-lobe [YSL11]. multiple-view [CH11]. Multiregion [MMV06]. Multiresolution [CKB96, FKW98, SL96, TW98, YW99]. Multiscale [BM98, DT97, GJP96, Hu11, KVdG^{+}97, Mok97, NDN^{+}97, NVWV97, PB99, AA20, BNG03, BNG05, DAM12, NBDB04, SH09, GZY23]. multisensory [ACC^{+}16]. Multispectral [AM06, RKKK22, PCR^{+}04, ÜB05]. Multitask [DLBG19]. multitouch [JRBD^{+}15]. multivariate [NBFG20, PC15, TLEF06, AQ09]. Multiview [DF01, LTCT14, TP14, BFMW23, BY12, LYA13, MK18, UFF06, RG16]. Mumford [SOL14, SOL16]. murky [TKDN16]. muscle [SASYGCRG24]. muscular [TLY^{+}16]. music [BLH16]. musicians [BLH16]. Mutual [KT07, LLWZ21, PMV00, WLL^{+}22a, ZhZFL22, EF14, FWLQ23, GKPS15, LZZ^{+}21, PC05, SRO^{+}19, WYX^{+}16, ZKRH04]. myopic [SPC^{+}15].

N [ZSCP08]. naïve [CH17]. Narrow [AS08a, Mil09, MBMC11, LLL^{+}14]. Natural [HWW06, CTM^{+}13, JYX^{+}23, LBNS09, Mig12, MLJC20, TRPD20, YWMS08]. Naturally [GHML17]. naturalness [LLNC20]. Navigation [GSV00, KR99, RJ00, ILRB04, LBC^{+}21, LM16, PLB16, RRAR^{+}16, ŠRDC09, TDWH07]. Navigational [RR95]. NCMS [CZS^{+}20]. Near [LQQS21, MBHRC21, CHC11, HASS10, JRS21, JN09, RB18, TMNM09, TAK^{+}22, XTZZ14, ZTH^{+}11]. Near-convex [LQQS21]. near-duplicate [CHC11, JN09, XTZZ14]. near-duplicated [ZTH^{+}11]. Near-Future [MBHRC21]. near-isometric [RB18]. Nearest [CGU11, GKPS15, KHH^{+}12, LZS16].

Nearest-neighbor [CGU11]. Necklaces [GSP02]. negative [AO16, LLL13, ZLL^{+}14]. neglect [HH05]. Neighbor [ZWB^{+}22, CGU11, GZX^{+}23, KHH^{+}12, TCM18]. Neighborhood [MMS97, MKK02, ADGB16, GHZ^{+}13, Hu08, NSEA13, SW04]. neighborhood-sequence [NSEA13]. Neighborhoods [CM99b, HUI16]. neighbors [GKPS15]. neighbour [LZS16]. Neighbourhoods [SB02]. Neon [HSTL24]. Nested [TS00b, VGR16]. Net [WRH97, LLP16, WPSL18, GZY23, KKP24, THH^{+}23, WDC^{+}24, ZLL^{+}24]. Nets [AMMV99, MAM97, TLEF06]. Network [ABJ^{+}21, CGL98, CWLY22, CLCZ23, JYL23, LLZ23, MDK^{+}24, TGQ23, TZL^{+}22, WZJ^{+}21, ZWW^{+}20, ZKLZ23, AVBK10, AM17, AAL22, BB24, BRPC17, CBB19, CWC^{+}20, CWW^{+}22, CLDP23, CKP^{+}19, DLC^{+}24, DWC^{+}24, EXP^{+}20, EH21, FWZ^{+}24, GZX^{+}23, GFW13, GYCS21, HWZ^{+}23, HLL^{+}23, HYW^{+}24, HL23, JCLZ21, JYX^{+}23, JB15, JSC23, KH23, LNN^{+}19, LRZ^{+}19, LYKY19, LWW^{+}21, LFLZ23, LZB^{+}23, LYW^{+}24, LSGY24, LYX^{+}24, LMM22, LZZ^{+}23b, LCG^{+}24, LDT21, LZL^{+}22, LXW^{+}23, LXW^{+}24, MHX19, MWL^{+}24, MSM17, NKL24, PTK24, QBZ21, QCH20, RYLZ24, RXDS22, RG17, RMS^{+}19, RKKK22, SP23a, SKLM22, SFF^{+}18, SDK22, SZL^{+}23, SPK23, ST20, SYL^{+}24, THH^{+}23, TAC23, WHL^{+}20, WLZ23, WGGH24, WDC^{+}24, WGZL20, WYW^{+}22, WKT22, WXZ24, XYQE24, XSZ^{+}20, XZQJ21, XGT^{+}22, YQL^{+}23, YSZ23, ZSC^{+}23, ZWLH24, ZHL^{+}20, ZCW^{+}23, Ziv10, CWL^{+}23, GZY23, HDZR24, KKP24, LHG^{+}23, MC22, NHTG15, VB16, WZF^{+}24, ZDZ^{+}23]. Networks [BPQ15, DAZ^{+}17, FCM20, LCS^{+}21, LVS20, SB95, SC00a, SKT18, SC98, VPL23, ZhZFL22, AH24, AMGG^{+}16, BSM10, BPS10, BBB96, BSZ^{+}21, BMvT^{+}19, BMX22, CSGM^{+}24, CLCO19, CBTC23, CÇ15, DFSC20, DDLP10, FM22, FFA^{+}19, GL19, GGGROE^{+}17, HZK19, KUHY18, KLY21, LBC^{+}21, LLWZ21, LWLP23, LCLH18, LWH03, MDM^{+}21, MCT10, MP20, NB20, NL23, OBTMT15, OZT19, PKC^{+}18, PCM21, RCLS19, RTM^{+}17, SKLM22, SCC17, SB22, SXY^{+}23, SST06, TKK24, TAC21, TN07, TY22, ÜB05, VKL18, VBT19, WPQ20, WWG^{+}18, WLL22b, YFX^{+}18, ZK17, BT23, DWL^{+}24, FDSB22, GFL^{+}19, LB19, LLG^{+}23, ZH18]. Neural [CGL98, FCM20, PMG^{+}23, SC98, TGQ23, WRH97, ZDZ^{+}23, AM17, BBB96, BRPC17, DFSC20, EXP^{+}20, EH21, FM22, GL19, GGGROE^{+}17, GFW13, HZK19, KUHY18, KLY21, LLP16, LSGY24, LLWZ21, LCLH18, MDM^{+}21, MSM17, NL23, RG17, RKKK22, SP23a, SKLM22, SFF^{+}18, SCC17, TKK24, TLEF06, TAC21, VBT19, WZC^{+}21, WWG^{+}18, WLL22b, ZK17]. neuromimetic [SCS14]. neuron [WHS^{+}24]. Neuroprostheses [PBPD^{+}17]. neutrosophic [SG11]. news [WHN08]. night [ASC17, MCCRAC20]. Nighttime [TYH^{+}21]. No [MvGS16, MYYY17, RDS24, SAV24]. no-reference [MYYY17, SAV24]. nodes [PL08]. Noise [Imm96, LZQL24, TO99, AYG23, DFSC20, GGGROE^{+}17, LG17, LWLP23, MRH19, MGPF08, RK11, WLW^{+}16, WWWF23, XTZ^{+}18]. Noisy [LR02, BTB14, KGC05, LPC^{+}20, LLWZ21, LBCA10, VRKL13, VGLP17]. Non [BSH22, BY12, CMD06, JHA17, LMP^{+}19, LBCA10, PRR03, QDLB17, SPC^{+}15, SS17a, TS16, AMNCM16, AMN18, AO16, AM15, BHBF10, BPS10, BCLNG18, BDS12, CR03, CP20, DPRC17, FB05, GRB13, GW07, HHG^{+}20, HMCT22, HSJS10, HC13c, JSRS08, KORC10, LNM^{+}21, LJHH07, LÁB15, LLL13, LLF18, LW18, Loh10, MMK04, MC20, NLM05, PW23, PA13, RKG03, SCALFG^{+}18, Sha06, SJ15a, SKH08, SZL^{+}23, SAC09, SB05, TMQM13, TLCH05, TWW14, UMH16, WWCZ15, WLW^{+}16, WR08, WWWF23, YC05, ZZZ06, ZLL^{+}14, ZDZ^{+}23]. non-alternating [HMCT22]. non-binarized [SJ15a]. Non-blind [JHA17]. non-central [BCLNG18, PA13]. non-contact [NLM05]. non-conventional [BPS10]. Non-convex [QDLB17, AM15]. non-cosmetic [BHBF10]. non-cubic [SB05]. Non-Gaussian [LBCA10]. Non-ideal [LMP^{+}19]. non-intrusive [YC05]. non-linear [AM15, CP20, KORC10]. non-local [HSJS10, LLF18, SZL^{+}23, WWWF23]. non-metric [ZZZ06]. non-model-based [PW23]. non-motion [GW07]. Non-myopic [SPC^{+}15]. non-negative [AO16, LLL13, ZLL^{+}14]. non-overlapped [LJHH07]. non-overlapping [HC13c, JSRS08, TWW14, UMH16]. Non-parametric [BSH22, CMD06, TS16, BDS12, HHG^{+}20, MMK04, WLW^{+}16]. non-playing [BLH16]. non-radial [WR08]. non-redundant [DPRC17]. Non-rigid [BY12, PRR03, SS17a, AMNCM16, AMN18, CR03, GRB13, LNM^{+}21, LÁB15, LW18, RKG03, SCALFG^{+}18, SKH08, TMQM13, WWCZ15, ZDZ^{+}23]. non-SVP [FB05]. non-topology [Loh10]. non-uniform [MC20, SAC09, TLCH05]. non-voting [Sha06]. Nonanalytic [SCS99]. noncentral [GA09]. Nonconvex [Bd96, BBH14, GGH^{+}24]. Noncoplanar [CRC97]. Nonfuzziness [WCZ02]. Nonlinear [CRC97, CBM01, EL07, KS96, NVWV97, TGSH98, BLJ^{+}23, DAM12, HLKK19, KG14, LV11, PW06, SCvW11, ZP18]. nonlocal [GGH^{+}24]. nonnegative [GXC23]. Nonparametric [GKPS15, PF99, ZOMK00, BCMCB09, TL16, YHN11]. Nonrigid [ACLS98, Ano01l, FDMA97, FT98, GSST03, LPR^{+}03, Pen99, TGSH98, CBD^{+}03, CALO20, SK15]. norm [CZS^{+}20, DOSD11, QDLB17]. normal [CLO17, HC13c, LHLZ23, YA12]. Normalization [RRL20, RY98, CM12, Hu11, KTE^{+}17, LDGS^{+}13, WLFL21, XMT22]. normalized [GH08]. normals [MC20]. normative [WPI^{+}16]. nose [NB10]. Note [Ano01h, Ano01i, Ano01j, Ano03m, Ano06i]. Novel [APV99, CCP97, KR99, ABVC16, BYJG23, CSW^{+}24, CKLP09, CU10b, DK13, GCD^{+}18, GLZF23, HSK23, KBN12, LLS21a, LZYW23, MYZ^{+}24, MDK^{+}24, PRG^{+}14, PCC13, RBdDS14, SP23a, SZGK24, TT16, WGAD14, WXC20, XW16, YWL^{+}22, YC05, YLX^{+}18, ZSCP08, ZCF13]. novelty [WHN08, WLFL21]. November [Ano19n, Ano20u, Ano21v, Ano22u, Ano23v]. NSCT [PLYW21]. NSST [LZWX21]. Number [Ano01m, Oli01, APB10, GLM17]. numbers [HY11]. Numeral [HY98]. Numerical [DFS08, KBJ^{+}10]. NURB [Ano95e].

Object [ACF00, AW09, AW98, BBC00, BB03, BZ99, BSF02, CF01, CGL98, CS98, CS00, DUC97, DCTO97, DC00b, GBL08, GK95, GCT^{+}14, HR99, Hod95, HP96, ILRB04, KMB97, KP00, Lau97, LD98, LLC12, LWH03, MDFS11b, MFJ95, Mas02, MKK02, May99, MNSK98, NG98b, NCDG21, OG98, PRCP16, PS05, QV98, RW97, PBPD^{+}17, SU01a, SF95, SN99, SGB01, SLL01, Sta95, SKBS13, TPNP15, WZW17, WZWL24, WPZ^{+}18, XAB07, YQL^{+}23, YT99, YC98, YSNiT14, ZYQ^{+}23, ZZZP09, ZYS09, ACAAC^{+}08, ALY^{+}22, AT13, AHDM10, BN15, BSM10, BVWS21, BL04, BM15, BPB13, BFD22, BSH13, BH12, CICN22, CZL^{+}24, CHH09, CS04, CWO^{+}11, CZS^{+}20, CJYW23, CSZ^{+}15, CZHT15, CL08, CYC10, CCYC12, CPO16, CYG16, DLC14, DFJL15, DTL17, DFP23, DNG^{+}24, DWC^{+}24, DHP08, DBBB14, EB13, EOPS22, ES04, FFM05, FBZP15, FFFP07, FLCdA06]. object [FR11, GM19, GB10, GGGROE^{+}17, GRCD18, GZX^{+}23, GPG^{+}15, GLG22, Gwa17, HYJ11, HD23, HML15, HJZ16, JEF^{+}12, JBR08, JYL23, JHV19, JHK24, KG14, KLO20, KKKS24, KRK11, KBD^{+}12, KS04, KH13, LMRMJ08, LNN^{+}19, LRFF24, LWZC14, LYX^{+}21, LFLZ23, LJC^{+}23, LLS^{+}23, LJC^{+}24, LYX^{+}24, LL12, LC09, LAL^{+}10, LLG^{+}23, MT16, MW22, MP14, MSF^{+}17, MGCS17, MFSB23b, MFSB23a, MHSP10, MCM^{+}17, MTR^{+}23, NAS^{+}17, NMAL23, NDO09, ODT17, ÖÜ20, PE09, PW23, PSE^{+}11, PSR08, PZM^{+}21, PL10, PKvGS16, PZV13, PTK24, QWHW20, RXDS22, RB16, RCT14, RLF15, SKLM22, SPC^{+}15, SB18, STV09, SAdB14, SZ16, SZ07, SCL13, ST10, SIRS21, SG17, SIT07, SFWG08, TG11, TTXT21, TAK09, TLH22, TFD07, TP14, TPT17, TPDP20, TESY15, TBZ^{+}24, TC11, TL15, VCDS^{+}17, VCLS19, VGSMN16, WW16, WCZ^{+}20, WHL^{+}21, WLL^{+}22a, WHJK23]. object [WG23, WDC^{+}20, WDB12, WB16, WZCG24, XYZH11, XYZ16, XST04, XMN^{+}15, XWC^{+}23b, YZY11, YZL16, YWL^{+}22, YHW^{+}23, YNCO11, YJA96, ZKSV18, ZEGEJ15, ZLZH17, ZWZZ18, ZC19, ZZ20, ZJJ22, ZWB^{+}22, ZSC^{+}23, ZWZ^{+}16, ZYT10, ZJL23, ZNG^{+}13, ZCK09]. object-action [KRK11]. Object-based [LWH03, LMRMJ08]. object-centered [SCL13]. Object-guided [PRCP16]. Object-level [BB03, PSE^{+}11]. Object-Process [LD98]. object-specific [XYZ16]. Objective [SJST07, SYPK13]. objectives [AM15]. ObjectPatchNet [ZTH^{+}14]. Objects [BLP95, BH99, CM95, GESB95, HCHD01, IE99, KII98, LF96, LM99b, LK00, MS97b, MS00, NL96, SK02, SU01b, SMK02, SCS99, Tay00, TGSH98, VKP98, WD96, AXSVL14, AVBK10, Ano06h, BBK14, BL08, BPLT15, BP09, CKLP09, CUSZ07, CMG16, DLMC16, DR04, DGC12, DBB13, GKK05, GB08, GRB13, HRC09, JKM07, KS12, KEG15, LA11, MSV^{+}20, MOT17, MHMO09, MWL^{+}24, MSF^{+}12, MBCJ17, OSM16, OCVV04, PA10a, PLLL03, Pen15, PK18, RB18, TS24, TLFM23, VZP^{+}09, WRKP05, WTYC18, XOF05, YHN11]. oblique [LSC08]. observable [HPvB^{+}10, ZT09]. observation [KTV17]. Obstacle [LB98, BSZ^{+}21, CSS13b, MTAA11, WGAD14]. obstruct [LSQL24]. Obtain [Che98, SSL^{+}12]. Obtaining [KM03]. Occluded [HFKN97, WH96, FFA^{+}19, LML^{+}23, LZS16, OBH04, OH04, PLLL03]. Occluders [ASZ99a]. Occluding [Sau99, ZM96, BN15, SECS15]. Occlusion [CLZZ13, CTE95, CN95, FK00, HKA13, Lai00, CH11, FBK16, GGGROE^{+}17, HH12, LST13, LSH19, MiMO^{+}16, MSSS09]. occlusion-aware [MiMO^{+}16]. occlusions [MiMO^{+}16, PA10a]. occurrence [LPVM13, PA10b]. occurring [PK18]. Ocean [SWYP00]. OCR [CB98, LZ97b]. OCT [KACR^{+}23]. October [Ano20v, Ano21w, Ano22v, Ano23w]. octree [LMM22]. odometry [PYGGLNG17]. odometry-aided [PYGGLNG17]. Off [AHD98, DLHT99, BK07, KK11, WASF14]. Off-Line [AHD98, DLHT99]. off-the-plane [KK11]. offensive [AO04]. office [OGH04]. offline [KSR^{+}12, YWH^{+}23]. offs [LHH^{+}98]. omics [KFRD^{+}18]. omni [DPM14]. omni-directional [DPM14]. Omnidirectional [BI10, OYTY98, SS09, BPS10, CYP^{+}10, NL17, PBSG12, WHL14, SST06]. on-board [GSPL10]. on-line [BAPXH16, NDO09, PBD20, RL13]. One [TDZ^{+}20, WLFL21, BPCT22, GBY21, GSV05, WSV05, WZQ^{+}24, YWL^{+}22, Eva06]. One-class [WLFL21]. one-level [YWL^{+}22]. one-shot [BPCT22, GBY21, WZQ^{+}24]. One-stage [TDZ^{+}20]. ongoing [WH18]. Online [BSM10, CBS17, CSK22, FXWW17, KBWT16, KG14, KRS14, NAS^{+}17, NHY10, PBI16, RB16, VMN16, WG23, WWLV11, BBCF20, BAM16, BIMD23, BAMK18, CJWW22, FSL24, GB17, KHH^{+}22, LLP16, MML^{+}16a, QDLB17, TMQM13, TPD^{+}16, USKB10, UWH17, WLL^{+}22a, YZL16, YJ16, YWH^{+}23, YCKA10, ZZRC15, WPZ^{+}18]. online-adaptable [UWH17]. only [YWL^{+}22, YCZ^{+}23]. Opaque [Sau99]. Open [GTMR23, TMS20, Cha21, Cha24, DSdlH^{+}11, NRJ11, SGZ21, XWDL23, ZdCR^{+}24]. open-set [ZdCR^{+}24]. Open-source [GTMR23]. OpenCV [SM10]. Openings [BJ96]. operation [ST20]. Operations [NK00, SHS03]. operator [ZSL^{+}16]. Operators [GHS95, HRS02, Hei99, Ang07, CSDNR17, GR05, VBS^{+}04]. opti [BN15, NT10]. opti-acoustic [BN15, NT10]. optic [CSS13b, Mar07, QKH^{+}12]. Optical [FSA01, FBK15, FSV07, Jea11, JM09a, LHH^{+}98, MNCG01, Muk97, NDBT95, RDM^{+}11, SP97b, Spe97, SB00, TS00a, XS98, ADGB16, BL09, DRAB08, DWC^{+}24, FWG18, FBK16, GYTL09, GPY^{+}07, HMF10, KGM19, KN11, LNM^{+}21, LB10, LmCT16, MN06, MZC^{+}05, MCF10, RPG12, SCGAF^{+}17, SM06, SM21, TLCH05, TDWH07, TD19, WHL14]. Optical-flow [JM09a, DRAB08]. Optical-Model-Based [Muk97]. optics [FB05]. Optimal [AZP14, ADDK99, ACDB12, BR95, DFSC20, Jos99, LH99, MEYD11, PV06, THT^{+}98, YHS95, DBF04, SS11, UO16, WLMG08]. Optimal-flow [MEYD11]. optimally [HKK08]. optimisation [KGB17, RRK13]. Optimising [XG08b]. Optimization [DH19, FB97, IW97, Jon97, LPS^{+}11, TGSH98, AS09, BRA^{+}10, BPB11, BYK^{+}18, CLL^{+}21, CSLX16, CKF18, CMH13, CKC14, GKBW14, HG11, HZLM11, KL11, KLBP11, OEK08, PB11, PZ08, PZ09, PW06, SLK23, TPT17, WB16, YSL11, Zac18]. optimization-based [CKF18]. Optimized [SBB18, ET15, Pha17, SM10, ZJJ22]. Optimizing [CW15, PKP97, KTP08]. optimum [CFYU12, dSdSF^{+}12]. optimum-path [CFYU12, dSdSF^{+}12]. options [TVLS08]. ORASSYLL [KP00]. Order [RM02, SJ15a, VF96, AM15, DD11a, JPP^{+}14, KSRS16, KA08, LEE^{+}18, LGD16, PL08, She16, UO16, VGR16, ZZP12]. Ordered [Pud98, Ang07]. Ordering [MMS99]. Ordinal [TCM18]. Ordinary [FM99]. Organ [NSK^{+}97, BvdHL^{+}13]. Organization [ACF00, ASZ99b, BS99b, BSF02, SB98a, SMK02, Sau99, HGS08, LRD19]. Organized [KP00]. organizing [TLEF06]. Orientation [AS17a, CF07, Dre96, PBT14, QTLP22, RCT14, RFS03, WZ04, WHGZ20]. orientation-from-color [Dre96]. orientations [ZJ05]. Oriented [BB16, FYH11, GZJ05, HL13, LCL^{+}14, LmCT16, PCC13, QCH20]. Orthogonal [CL00, FB97, LZD^{+}14, KA12, LFMP13, YGH11, ZRRK18]. orthogonally [DBB13]. orthographic [LCT09]. orthoimages [BABB19]. oscillations [Boy04]. Other [WZWL24, WZX^{+}24]. out-of-distribution [YSKL24]. outdated [MRH19]. Outdoor [BD02, SPQ^{+}17, CPC08, DM24, SSHP17]. outfits [ABB^{+}23]. Outlier [DF02, Bar18, KdRM^{+}23, LE09]. outlier-based [KdRM^{+}23]. outliers [LG17]. outlines [Got08, LYG07]. output [DSK^{+}20]. over-exposure [ABK^{+}18]. over-segmentation [BSALF18, KS15]. overfit [KKSC23]. overhead [PE09]. Overlap [MSW96, PKvGS16]. overlap-based [PKvGS16]. overlapped [CL17, LJHH07]. Overlapping [NS98, EKY08, Gol05, HC13c, JSRS08, LG14, TWW14, UMH16]. overlay [XZX^{+}21]. overview [Pop07, TPT15, YGJ^{+}20].

P [Ano95d]. P.-J [Ano95d]. PA [TYDH18]. PA-Search [TYDH18]. paced [DDZ^{+}23, SIRS21]. Packet [TS00a]. paddlers [DZLH17]. Page [Ano17j, Ano17k, Ano17l, Ano18k, Ant98, KSI98]. Page/Cover [Ano17j, Ano17k, Ano17l, Ano18k]. pages [Ano01m, CMCM16, Oli01]. pain [CCF17, LL17, RG16]. paintbrush [ZG06]. Painting [PMG^{+}23]. paintings [CHL05]. Pair [DF02, DH00, DH19, Pha17, SA96, CLL^{+}21, KDT^{+}18, MM21]. Pair-Wise [DF02]. Pair-wisely [Pha17]. Pairs [RFC97, KH15, ZRRK18]. pairwise [Gol05, KBMD15, RM03, YSX^{+}19]. palm [ABEN09, MKF15]. Pan [CC00, SP06, DDLP10, SPC^{+}15, MNR18]. pan-shots [MNR18]. Pan-tilt-zoom [SP06, SPC^{+}15]. panoptic [BCHR24]. panorama [Che08, DWB11, WZT13, ZH04]. panoramas [BDL^{+}06, CACB17]. Panoramic [FB05, KW99, AAB19, CMM20, MAL10, ZKRH04]. Paper [Ano07f, Ano08k, Ano12m, BKMSR98, Ano13o]. Papers [Ano01k, Ano01l, LLNS18]. parabolic [Ste13]. paracatadioptric [BA06]. paradigm [KFRD^{+}18, ZN08]. parallax [ZHS^{+}24]. parallax-tolerant [ZHS^{+}24]. Parallel [AW98, BCG95, Che98, CCS95, DRCF95, ER96, IW97, KSS97, LHKC97, LH99, MS96a, MW00, MNHO00, RF02, SKS11, SM97, Tan95, THT^{+}98, MHSP10, NB20, SHW24]. parallelogram [ZSL^{+}16]. parallelograms [KK09]. Parameter [SC00a, SCS99, HD09, LLWX24, Sah05, SS11, UTB^{+}11, WWWF23]. parameterization [CHZ^{+}13, PHH^{+}15, YNZ^{+}19]. parameterizations [NESP10]. Parameterized [WSSD96, YB99, DB03]. Parameterizing [ANM98]. Parameters [CSC96, CL00, AAB19, BF07, BJS14, GA09, KY06, LMC09, PA13, RRK13, RAC^{+}13, STHBH18, TA11]. Parametric [BCA98, BA96, DM01, GBHS06, Gui99, LVW97, QAB^{+}11, ÜE01, WF02, BUD19, BVVMMS15, BDS12, BSH22, CMD06, FBK16, HHG^{+}20, KA08, KGC05, KNO^{+}09, LWFL24, MMK04, MP09b, TS16, WLW^{+}16]. Parametrization [BGK95]. Paraperspective [Chu02]. paraphrases [COV^{+}22]. Parasite [TDK10]. park [CPC08]. parsing [BSH22, DGG08, MDFS11a, PSYZ13, SP23a, TL16, VB16, YGJ^{+}20, YW16, ZLY^{+}20]. Part [AZ15, BM15, KS04, TPT17, YH19, BB15b, CWO^{+}11, FO18, LLCY21, LAL^{+}10, MvGS16, PS05, SJ15b, WZQ^{+}24, YG17, ZWZ^{+}16, ZJW15, FKL^{+}16b]. Part-based [BM15, TPT17, BB15b, FO18, YG17]. part-in-whole [FO18]. Part-level [KS04]. Part-of-Speech [YH19]. part-sense [CWO^{+}11]. Partial [Lai00, Pla96, QCH20, KS03, LPR^{+}03, MB95, SKVS13, XOF05]. partial-surface [XOF05]. Partially [HFKN97, GB13, HPvB^{+}10, OBH04, OH04, PLLL03, SR23]. Particle [DD11a, LST13, BW11, BL09, BKMV07, DYM14, HBB^{+}12, KG14, KLK14, LÁB15, LAFLB16, MEYD11, MHSP10, MiMO^{+}16, RDA^{+}15, SBB10, YNCO11, RRR11, SC15]. ParticleAugment [TB23]. Particular [Lin02]. Partition [CCTCR09, ABD11, BW11, MWF07]. Partition-distance [CCTCR09]. partitioned [WDB12]. Partitioning [SB98b, DBB13, MMV06, MMK04]. partly [WSJ15]. Parts [DFJL15, LF96, RDR95, DHP08, LLC12, MvGS16, PA06, PYS03, SAdB14, ZZZ06]. PASHA [CBD^{+}03]. Pass [CCS95]. passers [MLH13]. passers-by [MLH13]. passing [MTR^{+}23]. Past [ZZZ15, JKW^{+}21]. Patch [CAGN24, DH19, VV02, DZQ24, GFL^{+}11, PBW14, QSXS23, SZW^{+}21, SXY^{+}23, ZC19]. Patch-based [CAGN24, DH19, SZW^{+}21]. patch-loss [SXY^{+}23]. patch-wise [DZQ24]. Patches [BM97, KBMD15, KYYC14, PZV13, XYW11, ZK17]. Path [DJG01, SU01a, YYL96, CFYU12, CS20, GTP18, MZB^{+}10, WZQ^{+}24, dSdSF^{+}12]. path-based [CS20]. pathological [KSHE20, WPI^{+}16]. Pathology [MFP^{+}20]. Pathology-sensitive [MFP^{+}20]. paths [DBBB14]. Pathway [ZZSD21]. pathways [HHG^{+}20]. Pattern [Big97, CCP97, HB98c, KC99, MT00, ADFR18, BRP04, HSBS16, MGPP11, TT16, WYX^{+}16, YR06, kCE^{+}18]. Patterns [Bd96, ME98a, Nis97, YSX^{+}19, BHSD^{+}13, GWT09, Gwa17, LSP^{+}16, MdBJG15, MB05, MB11, NNBN20, SJ15a, WWJ16, WTBdB15, YLM11, AGB^{+}15]. PCA [BZ14, DBBB03, GGH^{+}24, QDLB17]. PCB [MEDT96]. PDE [MPST08]. peaks [FS03]. Pedestrian [BBC^{+}07, DZL07, JB15, PLB16, YHS^{+}20, CSK22, GSPL10, KRJ^{+}08, NHH14, RKKK22, SPT^{+}18]. pedestrians [MAG^{+}16]. peeling [FMG23]. peer [MGPF08]. pelvis [CZ14]. Penalizing [HD23]. Pentland [Dre96]. People [HCHD01, HF01, MJD^{+}00, PF01, UMH16, CHP^{+}11, CZZS07, FFA^{+}19, GMM15, GLOC10, GTMR23, HRHZ17, HFR06, HH12, DFP^{+}13, PMC13, TMB12, TB13]. perceived [LWYF24]. Perception [MJS97, SGDP01, Boy04, CSGM^{+}24, FY06, MML^{+}16b, OH05, SB96a, WGZL20]. Perceptual [ASZ99b, BS99b, CH96, CCP97, JDP97, SB95, SMK02, Sau99, SN99, SPK^{+}02, WH96, GZP05, KH23, LSP^{+}16, LBNS09]. Perceptually [IW97, SM99]. Perfecting [CLD96]. Performance [BS00a, BG09, Car01, FPMK19, KTP08, LPH01, MM06, PDK96, SGB01, TCB^{+}08, TS01, VD10, Ano05j, BHBF10, BGPD09, DRAB08, FBF08, GMM15, HBF09, HC13b, KDT^{+}18, LvdHK^{+}15, PV15, QWHW20, RZH17, TPT15, WBS14, ZZH23]. Periocular [SR23, PMR17]. Period [GLR^{+}99]. periodic [RSPD12]. permutation [TAK09]. persistency [She16]. Persistent [JY14, MiMO^{+}16]. Person [HF01, LLCY21, LLZ23, LYD24, ZKLZ23, ACP16, ALK^{+}09, BAPXH16, BHA24, CKP^{+}19, DPRC17, FHZX23, GK23, GZL^{+}23, GML^{+}21, HBKG22, HFR06, HSHA20, JRAJ17, KDSF20, KT07, LWLC22, LLLW23, LG14, LZYW23, LML^{+}23, LXW^{+}23, LXW^{+}24, PWSvdH17, PY08, RFMF21, RCJ^{+}13, VBT19, VZP^{+}09, WHY^{+}23, WWG^{+}18, YJ16]. Personal [RCJ^{+}13, MFS^{+}07]. Personality [SCC17]. Personalized [CD10]. Persons [WN99, HPvB^{+}10, MW13, PA06]. Perspective [BR95, Che96, Gui99, BYJG23, CPT07, DWW^{+}12, HN95, MOB14, SCGAF^{+}17, SGZ21, WXWC18, WXC20, YHR^{+}05, YLX^{+}18, ZH04]. perspective-three-line [WXC20]. perspective-three-point [WXWC18]. Perturbations [LCS^{+}21]. pervasive [SFK18]. PET [LWLT17]. PGF [LLJ^{+}23]. PGF-BIQA [LLJ^{+}23]. Phase [AVGASAP15, AS09, AT17, DCS05, HTNN18, IJDAB13, LSCK15, PYWZ17, WB11]. phase-based [HTNN18]. phase-field [LSCK15]. phase-preserving [PYWZ17]. phenotyping [WM20]. Phong [RF23]. photo [ADR16, ADFR18, CLTW23, DBT^{+}17, JRBD^{+}15, WL15]. photo-sketch [CLTW23]. photo-streams [ADR16, ADFR18]. photo-textured [JRBD^{+}15]. photogrammetric [Lhu23]. photographs [ABK^{+}18, Che08, CHL05, WLX^{+}14]. Photography [TVY^{+}18, KHR^{+}16, NFA04]. Photometric [APB10, CMM20, KP97, NG98b, OD01, RBA20, Atk17, GCFMT12, HASS10, HJ12, JC06, JMPG11, OSY18, SF16, TKDN16, YA12]. photomontage [LLL^{+}20]. Photomotion [ZTS96]. photos [IZKB12, PHY^{+}11]. Physical [DF01, Hod95, RWV95, HSTL24]. Physician [SBK^{+}99]. Physician-in-the-Loop [SBK^{+}99]. Physics [BLKG21, Bra97, MS97b, WR08, DKG22].

Physics-Based

[Bra97, MS97b, BLKG21, WR08, DKG22]. physiology [PDS^{+}07]. PICASO [TKV16]. Pick [NCDG21]. Pick-Object-Attack [NCDG21]. pictogram [BRA^{+}10]. Pictorial [KR98]. Picture [Bic98, LRD19]. Piecewise [BS96, BA96, Bar07, BL08, KCZ18, MJPS16, PZV13, SOL14]. Piecewise-Linear [BS96]. Piecewise-Smooth [BA96]. piles [TN08]. Pipelined [OTL96]. pitted [PK05]. PIV [ACG^{+}09]. pix2seq [LZTX24]. Pixel [Che98, AVGASAP15, ACDB12, CKC14, DZQ24, GBF12, GGO10, HUI16, HWZ^{+}23, JLL13, LFL08, SJ15a, VMP03, XJK12, ZLZH17, ZJW15, TKV16]. pixel-labeling [JLL13]. pixel-level [LFL08, ZJW15]. pixel-wise [CKC14, DZQ24, HWZ^{+}23]. pixels [MGPF08]. Pizlo [HM97, May97, Ver97]. place [LD24]. Placement [MG95, CYP^{+}10]. plan [ES06]. plan-specific [ES06]. Planar [BH99, GBB98, MS96b, NG98a, ST96, SY11, ACAAC^{+}08, Bar07, GSGJ22, GSGJ24, HY11, KCZ18, PAK19, PY19, PZV13, WTYC18]. planarity [RF23]. Plane [LB98, CKS^{+}05, HN95, KK11, Neg12, OK04, ZHZ17]. planes [KK11]. Planetary [UZC97]. Planned [IB01]. Planning [SKOS95, TG95b, YT99, PW23, ZKRH04]. plant [LZD^{+}14, WM20]. plate [KKP24, LYW^{+}24]. platform [KK15, MZB^{+}10, RMN^{+}17]. platforms [BVWS21, VAWW10]. Plausibility [CPC99]. plausible [FFA^{+}19]. Play [GB22, WASF14]. playback [SBS04]. player [GLM17, LCLH18, MEM17]. players [FLB06, PD17]. playing [BLH16]. playing/non [BLH16]. playing/non-playing [BLH16]. plenoptic [LTAA23, MMBG18, SL16a]. Plug [GB22]. Plug-and-Play [GB22]. POCS [AM06]. POEM [CLDP23]. Point [CPC99, GSP02, GSK02, HRS02, LK00, OD97, RKG03, SCALFG^{+}18, SBZ97, Tay00, TML00, TS01, WB01, ADC19, ANHGS17, ABD11, ATC^{+}13, BHSD^{+}13, BSALF18, BWG17, CLK09, CALO20, CDT11, CS04, CK09, CR03, CP20, CACB17, FBZP15, FCOK24, GG09, GDCM17, HDZR24, HY11, HWL^{+}22, JHV19, JSC23, KDT^{+}18, Kim04, KLO20, KKSC23, Lhu18, LZZ22, LZB^{+}23, LZLP10, MLB^{+}18, PD14, PB11, RAC^{+}13, RLB17, SAS12, TLFM23, THH^{+}23, WWCZ15, WH18, WXWC18, WHGZ20, YK08, ZZK^{+}20, ZSK^{+}23, ZMJ^{+}15, ZMM^{+}22, HDZR24, CTWH15]. Point-Based [LK00]. Point-Enhanced [GSP02]. point-set [SAS12]. Point-Voxel [HDZR24]. pointed [PBT14]. Pointer [DRCF95]. Pointer-Based [DRCF95]. Pointly [ZLY^{+}20]. Pointly-supervised [ZLY^{+}20]. Points [DT96a, FT98, OG98, PM97, Shi99, SLL01, ZL01, ATG15, BL20, CHMG12, FM22, Kui08, LLL^{+}14, LLY^{+}18, LB10, Loh10, MPST08, ODD96, TY05, UTB^{+}11]. Polar [MGMS01, ÜE01, KORC10, Mas09, Sch06, SCS14, TP05, LMP^{+}19]. Polar/Spherical [ÜE01]. polarimetric [ZZZP09]. Polarisation [Atk17, AH08]. Polarization [LL97a, WAPB17]. policies [OH05]. Polygon [LR02]. Polygonal [BS96, HB98b]. Polygons [BM98, MSW96, Kle13]. Polyhedra [SP97a, KM03]. Polyhedral [KCD00]. Polynomial [DSdlH^{+}11]. Polynomials [KP97, KA12]. pool [JVD^{+}20]. Pooling [ATC^{+}13, KYM13, NNS^{+}18]. popular [CH17]. population [Ham05]. population-based [Ham05]. pork [CCR^{+}05]. Portable [HT98, RZH17, STC^{+}16]. Portrait [SXZZ24]. Pose [AKC11, ACB98, AW98, BK01, CS10, CH99, CS00, HWK^{+}21, HDF12, Jos99, Jur99, LSW18, NB10, RY98, ÁB13, AC09b, ABVC16, AVC19, BPLT15, CLCO19, CL18, CDT11, CYNO11, CTH20, CPPY21, CLO17, CC16, DLC14, DGC12, DPCA15, DMSM21, DLF06, EDX16, EBN^{+}07, FPMK19, GSGJ24, GLZF23, GML^{+}21, HF11, HCLZ21, HSHA20, HH12, IDY^{+}18, KTE^{+}17, KUHY18, KZ05, KGB17, KMN11, LST13, LY06, LSTF12, MML^{+}16a, NWNT17, ODD96, PBT14, PD11, PHH^{+}15, PDTE06, PZC17, SBIK16, Sav23, SO07, SAC^{+}12, SFK24, SRHC13, TAK09, TST14, TWQW23, TPD^{+}16, TP14, TNO24, UU18, VBT19, WCZ23, WXWC18, WXC20, WTZ^{+}21, WSFTK18, WTYC18, ZEGEJ15, ZC19, ZIT^{+}13, ZCW24, ZDF10, Ziv10, dP10]. pose-based [PD11]. pose-contour [PDTE06]. Pose-Estimation [ACB98]. pose-free [CC16]. Pose-insensitive [NB10]. Pose-invariant [AKC11]. pose-wise [AC09b]. posed [WWJ16]. PoseGU [GLZF23]. poses [DLC14, MdRNM15]. position [PA13]. positioning [AVC19, YHS95]. positive [BB13, BB15a]. possible [PY19]. Post [GMM15]. Post-processing [GMM15]. posture [HCC^{+}16, WPB^{+}14]. Potential [BS99b, GESB95]. Potentials [RM02]. Power [QV98, TLB^{+}15]. PPformer [DZQ24]. Practical [Ano95e, SBMM15, dLAH07]. practice [PWWQ16, PBSG12]. practices [DAL^{+}22, TCB^{+}08]. PRCG [WLX^{+}14]. Precise [GCEC07, AAMO16, AS08b, dOSJVBS12, RTM^{+}17, WCZ^{+}20]. Precondition [YLK^{+}23]. preconditioners [KMT11]. predict [CCR^{+}05]. predictability [GGMV08]. Predicting [RFMF21, TYDH18, GML^{+}21]. Prediction [MBHRC21, RWV95, TS01, BMJF^{+}17, BCC^{+}21, BSZ^{+}21, DSK^{+}20, EMMV19, FLL^{+}23, HWZ^{+}23, KC22, KKKS24, LLCY21, LHLZ23, LZWX21, MSF^{+}17, NMAL23, PT15, PSYZ13, QAB^{+}11, SY23, TLFM23, TDT12, TKDN16, XSL^{+}24, YNZ^{+}19, ZSG^{+}20]. Predictive [LG17, SYF99, OBTMT15]. predictors [BB15b]. prejudice [RKL^{+}18]. prepositional [CLA^{+}17]. Preprocessing [RY98, BYN^{+}04, CE17]. prescription [BHBF10]. presence [CXFS06, LF08, PA10a, YS06]. present [JKW^{+}21, ZZZ15]. presentation [TD04]. Preservation [ASS97, CWL^{+}23, BÁRM23, Loh10, ZLFH23]. preserved [ZZC^{+}13]. Preserving [GL95, RM02, SP97b, SBZ97, VB98, BDHM09, CK09, GB22, Hu08, LLL13, MLJC20, MGPJ11, PYWZ17, PSB^{+}24, SZW^{+}21, TGQ23, ZSCP08, ZD18]. Presmoothing [HC13a]. Primal [eGZW07, CLL17, KTP08]. primal-dual [CLL17, KTP08]. Primary [WW16]. principal [CLL14b, BZ14]. Printed [ME98b, ME98a, Por00]. Prior [WZJ^{+}21, ALM23, AZP14, BCHR24, EPH^{+}21, FZ20, GSGJ24, HWG21, PLJS14, TYH^{+}21, TMQM13, WYC15, WPSL18, WZX^{+}24, WSKH13, YZT^{+}13, ZXC^{+}20]. prior-based [EPH^{+}21]. priority [BRSSAL11]. priors [CC11, EOPS22, GRCD18, JSZY17]. prism [KGM19]. prism-based [KGM19]. privacy [GK23, PSB^{+}24]. privacy-aware [GK23]. privacy-preserving [PSB^{+}24]. proactive [LSQL24]. Probabilistic [ACW^{+}16, AS17a, CH96, Cre99, GGR01, HD09, HSS^{+}16, HLKK19, KD10, KHH^{+}12, LT97, MGK00, ÖÜ20, PBQ99, Tsa96, WC99, ZKC03, CDT11, FSV07, GRGB^{+}13, GL19, HNB04, HW07, HYW^{+}24, KSHE20, KMN11, LHYK05, PJW11, PLLL03, SM12, SYK96, TFD07, ZG10, TC11]. probabilities [LPS^{+}11]. Probability [JHV19, LH95, LLJ^{+}23, TC11, XP11]. probes [BFR13]. Problem [Jur99, KB95b, KB00, Dre96, GSGJ22, IKST05, NESP10, PW23, WXWC18, WXC20]. problems [CLL17, JLL13, KL11, KMT11, KBJ^{+}10, MJ17, OEK08, WSFTK18]. Procedural [WLXL24, TZLT21]. Procedure [OG98, JM09a]. Process [IF99, LD98, MRF96, PKvGS16, ABD11, HPvB^{+}10, JHA17, UK12a, RRR11]. Processes [CA97, SB95, JHV19, KLO20, LSP^{+}16, NFM08, KBKS18]. Processing [CKB96, DRCF95, DPB00, KDRC98, LH99, RM98, UZC97, AC09a, ALIRT18, BCDH10, Dem05, FFY^{+}04, GMM15, Ham05, Jea11, JM09b, KHH^{+}22, KMT11, LC11, LEB07, LMM22, LPV07, PC15, Ros10, Sah05, WQY^{+}21, Ano95d]. Processor [OTL96, THT^{+}98]. procrustean [CLO17]. Procrustes [CCL^{+}17, CLCO13]. produce [KGC05]. Product [LPC^{+}20]. production [CD10]. products [PG13, TS19]. Professor [CV13]. Profile [Jok98]. profiles [CD13, Ste13]. profiling [MSV^{+}20, SCC17]. programmable [Gon09]. Programming [BEPW00, OTL96, HQW^{+}12, LZLP10, MSI10, PRK19]. programming-based [PRK19]. Progress [CFS98, IF95]. Progressive [AM01, CWLY22, DZZ^{+}23, JEK98, MGK00, RXDS22, RG10, CU20, GFL^{+}19, SZL^{+}23, ZZ20, WZF^{+}24]. projecting [BHSD^{+}13]. Projection [Chu02, Gui99, OD97, ZT98, AXSVL14, Bar06, DMW10, Gol05, LZLP10, MSV^{+}20, TCM18, WLL22b]. Projections [Ano01m, Luc01, ADC19, BA06, BCLNG18, BTB14, HN95, TP05]. Projective [Ano95e, ACB98, CDH99, FAB97, GHMQ97, LV96, RFC97, ZL01, ASF14, OBH04, OH04, PD14, ROGT14, SY11, TH06, LLTL14]. Propagation [CM99b, Egg98, BCMCB09, CS07, FF09, GKPS15, LSH19, PBW14, PL08, TB13, XTZZ14]. proper [MST16]. Properties [ASS97, GL95, Kis96a, OD02, Ros99b, ABD11, BY08, CKS^{+}05, Eva06, GFW13, MVP06]. Property [OD99, SB98b]. Proposal [LCS^{+}21, HLL^{+}23, PKvGS16, SZS17]. proposals [AvdWDM18, DTL17, HD23, MT16, MGCS17, MYV19]. Pros [Bor19, Bor22]. ProSRNet [CU20]. prostate [SM13a]. prostatic [TRG^{+}13]. prosthetic [HAM^{+}16]. protocol [WDC^{+}20]. prototype [CLDP23, XWDL23]. prototypes [LWSC16, RAHT11]. Protuberance [BL20]. provide [RGA10]. Proximal [KCZ18]. proximity [JN09]. proxy [SKA23]. prune [TAC23]. Pruning [AXJE21, MDL^{+}23, SB98c, BFD22, TAC21, WHS^{+}24]. PS [MFP^{+}20]. PS-DeVCEM [MFP^{+}20]. Pseudo [LLLW23, BBCF20, LLS^{+}23, DAL^{+}22]. pseudo-generative [BBCF20]. Pseudo-label [LLLW23]. pseudo-label-guided [LLS^{+}23]. Pseudo-Labeling [DAL^{+}22]. PSTG [CSLX16]. PSTG-based [CSLX16]. Psychological [CPC99]. PTZ [WZ08]. Publisher [Ano03m, Ano06i]. Pulmonary [WW97]. pulse [GFW13, SVF^{+}21]. punches [KFSM17]. Pupil [HBF09, KA12, SFK18, YWZ11]. puppet [MML^{+}16a]. pure [ECC18, SFK18]. Purely [CMCM16]. purification [TBZ^{+}24]. purposes [CNC03]. pursuit [LmCT16, BZ14]. push [TAC23]. Pushing [SKT18]. puzzles [YSKL24]. PVNet [HDZR24]. Pyramid [WZJ^{+}21, WZWT99, ZWW^{+}20, CWLJ13, HGP15, LSGY24, XYQE24, YSY^{+}18, ZWLH24]. pyramids [BBB96, GDIIHK11].

Quadra [LHY14]. Quadra-embedding [LHY14]. Quadratic [BM97, BPB11, LZLP10, OEK08]. Quadtrees [DRCF95]. Qualitative [Got08, FMGA^{+}12]. Quality [DT96b, KLL^{+}11, LKZ20, LLJ^{+}23, MYYY17, OAGN18, OSM17, SAV24, MNR18, TPD^{+}16, WZC^{+}21, WLM^{+}14, ZZC^{+}13]. quality-sensitive [KLL^{+}11]. quantification [LSCM03, TLY^{+}16]. Quantifying [JSWL24, AXJE21]. Quantitative [SB98a, LYBT17, LFL08, VO24, ZCLX20]. quantity [WLM^{+}14]. Quantization [SYF99, CS07, HDL^{+}20, JO11, JWG04, LHY14, WZY14]. quantized [WLL22b]. quartet [KDSF20]. Quasi [IE99, Por00]. Quasi-Metric [Por00]. Quasi-Objects [IE99]. Quaternion [HKM22, SF07]. Quaternion-based [HKM22]. quaternionic [DCFM07]. Quaternions [HB98b]. queries [LLL^{+}15a]. query [FPMK19, JRAJ17]. Querying [SL99]. Question [CAO^{+}23, DAZ^{+}17, OS19, JYY^{+}24, KK17, MLZRK24, RMS^{+}19, WTW^{+}17, XZW^{+}23, YXW^{+}24, ZWZZ18]. question-answer-based [ZWZZ18]. Quick [BL14].

R [Ano95d, MCM^{+}17, ZS19]. R-CNN [MCM^{+}17, ZS19]. R3DG [VAC16]. racquet [LHJ^{+}09]. radar [LB19, OVJ^{+}21]. radar-based [OVJ^{+}21]. Radial [Ano01m, Luc01, WHL14, BSM10, GOF^{+}15, KBJ^{+}10, TM04, WR08]. radiance [RH06]. radiographs [FLCdA06]. Radiological [PV97, OTO06]. radiometric [KGFP10]. radon [SOJ17, TWS06, ZS11]. rain [JCLZ21, LRZ^{+}19, ZCW^{+}23]. ramp [SA15]. Random [DB14, IF99, MCPB00, MRF96, NL23, PV13, WKP13, AMCB20, Bar07, CICN22, CL18, CZ14, CJL06, MRH19, MLB^{+}18, MJPS16, VGR16, WB11, ZSK^{+}23]. randomization [RG10]. Randomized [CC01, ED16]. Range [BLP95, BR12, BS00b, CFM02, CM95, DF02, EFF98, GJP96, HBH10, JB99, LF96, MY95, Mas02, Mur95, NL96, OD02, RF02, RFL02, SA96, ST96, SF97, SJB02, SPQ^{+}17, SB00, ASFP03, BBK15, CLZY15, CKF18, DLC^{+}24, DFP23, FK09, GBF12, HF11, HSJS10, LQQS21, LSKK10, LS12, LS09, MSR07, Mas09, MB05, MMBG18, RSS07, SY10, SLK15, SKU^{+}09, SKSR08, TG11, TST14, TS11, WB15, YAK^{+}08, YW07, ZG06]. range-sensing [ASFP03]. rank [ARFF18, ED16, GXC23, GF15, KHR^{+}16, LC14, LmCT16, LCL^{+}17, SZ16, TS24, TR09, WPSL18, YFDA17, ZXC^{+}20, ZLL^{+}14, ZLZH17, ZD18, ZZ10]. ranked [WDB12]. ranking [LWH^{+}23, PLJS14, SZS17]. RANSAC [CCL^{+}17, FWG18, LMP^{+}19, LG17]. Rao [KLK14]. rapid [AC09a, YCH07]. rare [LCG^{+}24]. rate [SM21, SVF^{+}21, TVC09, WYW^{+}22]. rates [ZBDP15]. ratio [ACDB12, SF16, YC05]. rationale [Pec07]. Ratios [LF98, ASCF13]. ray [AS08b, GYW^{+}22]. Rays [KHB01, BMvT^{+}19, CZ14]. Re [JHK24, LLZ23, LYD24, WHY^{+}23, ZKLZ23, BCC^{+}18, BCM13, BHA24, CKP^{+}19, FHZX23, GZL^{+}23, HBKG22, JHA17, JRAJ17, KU19, KDSF20, LLCY21, LLLW23, LZYW23, LML^{+}23, PWSvdH17, ŠSJ^{+}20, UMH16, WWG^{+}18, DAL^{+}22, WZWL24]. re-blurring [JHA17]. Re-Identification [LLZ23, ZKLZ23, LYD24, WHY^{+}23, BCC^{+}18, BCM13, BHA24, CKP^{+}19, FHZX23, GZL^{+}23, HBKG22, JRAJ17, KU19, KDSF20, LLCY21, LLLW23, LZYW23, LML^{+}23, PWSvdH17, ŠSJ^{+}20, UMH16, WWG^{+}18, DAL^{+}22, WZWL24]. Re-scoring [JHK24]. re-weighting [JRAJ17]. reactive [TM07]. read [CZ18]. Reading [KABP98, JB23]. Real [AMNCM16, BEPW00, BPQ15, BPLT15, CGH08, CKL18, GK23, Gon09, HT98, LC14, LÁB15, LB98, LHHC98, MWTN04, MTAA11, OYTY98, PGGM04, RZH17, UM05, WHL^{+}21, ZZH23, ZXK02, AM04, BCMCB09, BDS12, CEA16, CSK22, DLS^{+}09, DPCA15, DZJB14, FFM05, GTMR23, HWZ^{+}23, HWL^{+}22, HZW^{+}10, JRS21, DFP^{+}13, LLS21a, MZB^{+}10, MFS^{+}07, Nic95, Pen15, PBI16, RSS07, RL13, SM12, STC^{+}16, SFK18, SV14, SHS^{+}23, SGH07, SIT07, TKV16, UWH17, WX16, WWLV11, YWZ11, YZX^{+}17, YSZ23, ZJ05, Ziv10].

Real-Time

[BEPW00, HT98, LB98, LHHC98, OYTY98, ZXK02, AMNCM16, BPLT15, CGH08, CKL18, GK23, Gon09, LC14, MWTN04, MTAA11, RZH17, UM05, WHL^{+}21, AM04, BCMCB09, BDS12, CEA16, CSK22, DZJB14, GTMR23, HWZ^{+}23, HZW^{+}10, JRS21, LLS21a, MZB^{+}10, MFS^{+}07, Pen15, PBI16, RL13, SM12, STC^{+}16, SFK18, SGH07, SIT07, TKV16, UWH17, WX16, WWLV11, YWZ11, ZJ05, Ziv10]. real-valued [YZX^{+}17]. Real-World [BPQ15, ZZH23, DPCA15, HWL^{+}22, SHS^{+}23]. Realistic [GL97, YB01]. reality [CKM11, GWFF22, MBD^{+}22]. realize [TBZ^{+}24]. Reasoning [GESB95, KN99, AYB^{+}18, DFP^{+}13, LSP^{+}16, LYX^{+}24, SM23, YLK^{+}23]. Received [Ano97f, Ano98c]. receptive [KKCK23, LL12]. reckoning [Gre04]. Recognising [LZS16, SM17]. Recognition [AHD98, Ano96d, Ano01k, Ano15o, BH99, Big97, BB95, BZ99, BSF02, CF01, CGL98, CTF^{+}98, CS98, CCS01, CS00, CW00, DL97, DCTO97, DV98, DC00b, DT97, GBB^{+}18, GESB95, GK95, HR99, Hod95, JRH03, KH96, KABP98, KP00, LB00, LVS20, MFJ95, MLP97, MCAF21, MKK02, MNSK98, MYLP98, MT00, NSK^{+}97, NG98b, NMP97, PLLL03, Pla96, QV98, RDR95, RW97, SN99, Shi99, SGB01, SLL01, Sta95, VKP98, YB99, YC98, YFZ98, ZXK02, AAASC11, ACP16, AM17, AAL22, AT13, AFMY14, AC09a, AC09b, AKC11, ASCF13, ASF14, BGE^{+}17, BHBF10, BMJF^{+}17, BRA^{+}10, BKK11, BDT23, BL04, BFMW23, BWL04, BAM16, BRP04, BEGB13, BB24, BCF06, BPSV16, BH12, CICN22, CGU11, CMBP09, CLL^{+}21, CGR13, CGHTK16, CCFC13, CS04, CFB05, CHL21, CSZ^{+}15, CZHT15, CKLP09, CT13, CSG^{+}03, CR18]. recognition [CNC03, DT10, DFJL15, DWV19, DH19, EKY08, EK12, EB14, FBF08, FFY^{+}04, Far11, FBZP15, FLCdA06, FTT15, FR11, FAB12, FCM20, GFLW24, GGGROE^{+}17, GLM17, GFY^{+}14, GJ10, GBL08, GZJ05, GZ19, HHWP03, HOH^{+}07, HMF10, HNB04, Hu08, Hu11, HH19, HWT^{+}24, IZJ^{+}17, ITNP12, JLD12, JLD13, JM09b, KTE^{+}17, KK15, KFSM17, KIS17, KCM^{+}17, KKP24, KRK11, KFN15, KHA^{+}05, KSF16, KDV12, KS04, KRS14, LRW08, LCSL07, LLS21a, LHYK05, LZD^{+}14, LY06, LLC13, LDH^{+}15, LHSG15, LGG^{+}18, LPC^{+}20, LWLC22, LYW^{+}24, LZW^{+}24, LSGY24, LLG^{+}24, LD24, LXFM16, LHZY19, LL12, LL08, LYSS12, LLC12, LDC^{+}13, LGD16, LWSC16, MW22, MSF^{+}17, MdBJG15, MPM16, MYK03, MU11, MTVM04, MAJ16, MB11, MHAF13, NFM08, NN13, NFSD13, NNS^{+}18, NNBN20, Nis96, NHZ^{+}22, NDO09, OB14, OGB14, OVJ^{+}21, PC05, PWL^{+}23, PQML11, PWWQ16, PPT06]. recognition [PS05, PKC^{+}18, PS15, PCM21, PTE12, LL17, PS12, QZY^{+}24, QCXJ19, QCL^{+}23, RDS24, RAHT11, RM03, RG17, RR06, PBPD^{+}17, RS03, RLMK15, RKL^{+}18, RCJ^{+}13, SM12, STV09, SPT^{+}18, SS17a, SVSM15, SAC^{+}12, SSM06, SJ15b, SKVS13, SKM06, SSN03, SSS13, SASYGCRG24, SCMP14, SM23, SKT18, SS21, TG11, TKK24, TPDP20, TFL^{+}09, TESY15, TT16, TS19, TL15, VAC16, VKNK14, WRKP05, WY07, WCZ^{+}07, WS08, WH18, WLO^{+}18, WPQ20, WY21, WZQ^{+}23, WRB06, WRB11, WL15, WWXK24, XWDL23, XYZ16, YS09, YFF^{+}23, YAK^{+}08, YLK^{+}23, YST21, YSX^{+}19, ZLLP21, ZMJ^{+}15, ZEGEJ15, ZT15, ZSSF16, ZTGL18, ZTB20, ZZCL14, ZKC03, ZCWH23, BGPD09, TFL^{+}09]. Recognizing [BKPS15, DBBB03, IB01, LZL^{+}17, Por00, VM01, ZWLH24, CU10b, HS14, LLC13, PD11]. recombination [SZS17]. Recommendations [HS14]. Reconfigurable [THT^{+}98, CL95]. Reconstruct [Lau97]. reconstructed [RBdDS14]. Reconstructing [Gol05, KS03, OCVV04, RSPD12]. Reconstruction [BM99, BL01, CFM02, CPC99, CCS01, DG01, DC00a, FW97, FRL^{+}98, FKW98, Gui98, Gui99, GJP96, Hen98, LDPD97, LSHT02, OG98, OD97, PCJC98, RFC97, Tan95, Tay00, VB98, ZW97, ZRRK18, ZM96, ZOMK00, AMNCM16, AYG23, BYR17, BI10, BLKG21, BR12, BSRV17, BBK15, BBH14, CLK09, CPP^{+}11, CC11, CC03, CCD11, DWB11, FPC^{+}08, FB05, GRGB^{+}13, GSV05, GPC^{+}10, HLB17, HDG^{+}14, IZKB12, JRH03, JPP^{+}14, dOSJVBS12, KK11, KH15, KCZ18, KNO^{+}09, LB08, LY13, Lhu18, LLL^{+}14, LSCK15, LLY^{+}18, LFLZ23, MPST08, MWTN04, MJPS16, OSM16, PW23, PCR^{+}04, RDT^{+}19, Rem04, RZZ23, SY10, SSHP17, SCL13, SHK11, SMD^{+}08, SH08, SS11, TTXT21, TH06, Tan11, TTN17, UK12b, VNNB14, WZT13, YHR^{+}05, YW07, ZD18, Ziv10]. Reconstructions [CDH99, GJMO14, HASS10, LDH^{+}14, RTM^{+}17]. Recover [FL96, GR05]. Recovering [ACAAC^{+}08, CG09, LR02, MT16, Mur95, SP97a, WD96, WC99, WALL00]. Recovery [CJC01, DC98, RC97, SF97, SA02, TI01, YFZ98, BF07, CYNO11, GF15, KLL^{+}11, KM17, KZ05, LC14, Mal21, RRK13, SKBS13, TGFF15, TWW14, WML21, ZXC^{+}20]. rectangular [KZ05]. rectification [CCD11, KGM19]. rectifying [MDK^{+}24]. rectilinearity [RŽ05, Ros08]. Recurrent [LZZ^{+}23b, OS19, WZF^{+}24, FOCSB^{+}20, NKL24, RG17, YFX^{+}18]. recursion [HQN05]. Recursive [CSC96, DC98, HDG^{+}14, Kle13, LMM22, TMQM13, FKV^{+}11, NHSC09]. Reduced [Che98]. Reducing [RMD08, YZX^{+}22]. REduction [RLB17, BL98a, KAES99, PA00, CP09, GML16, LLL13, RRR11, WXZG18, ZWN14]. Redundancy [CM99a, WHN08]. redundant [DPRC17]. Reference [UK12b, CRCM16, LLR10, MYYY17, SAV24]. referencing [AWK04]. Refinement [DPM14, BBSD15, BI11, BG18, GFL^{+}19, LK03, WZX^{+}14]. Refining [SRM20]. Reflectance [LK97, OD99, OD01, PK05, SP97a, BABB19, GCD^{+}18, LMC09, YA12]. reflection [AO16, RRK13, RF23]. reflections [LF08, NNT11, SW13]. refractive [BK16]. Region [BL00, CWH^{+}13, IP98, KLL^{+}11, LCS^{+}21, OM19, PM97, PBG04, SYF99, SL99, CM16, CKK^{+}12, CBTC23, DTL17, ECC18, EyGS11, FLS^{+}14, IJDAB13, LML^{+}23, MMV06, MJ11, Mil09, MBMC11, MKF15, PFGG09, QCL^{+}23, SI03, SO07, SCvW11, VWMZ15, WLSO23, KL10]. Region-Based [PM97, SYF99, KLL^{+}11, OM19, PBG04, CBTC23, Mil09, MBMC11, SI03, VWMZ15]. region-labeling [EyGS11]. region-merging [SCvW11]. Regional [CD13, LmCT16, MSW15]. Regions [DAZ^{+}17, GSP01, LM99a, PF99, Rob96b, SM99, ABD11, CKM11, CCPK16, DMAD17, GS95, JRBD^{+}15, MAK^{+}17, PD05, SH09, TN05]. Registering [BLP95, TS11]. Registration [Ano01l, CFM02, DF02, Dav97, EFF98, FDMA97, FAB97, HLF^{+}97, JGP19, Jok98, KPH02, KSHE20, MY95, Mas02, OD02, PMV00, PLH04, RC03, RF02, RFL02, SK02, SKSR08, TB99, VV02, WB01, WHGZ20, ADC19, ASC17, AS08b, AT17, ASFP03, BI10, BT05, BvdHL^{+}13, BW15, CBD^{+}03, CALO20, Che08, CHZ^{+}13, CKF18, CFM^{+}13, CR03, CP20, FBS21, GGMV08, GSST03, GDCM17, HTNN18, HY11, HWL^{+}22, JBWK11, KKSC23, KT07, LV11, Liu10, LS12, LPR^{+}03, MMA06, Mas09, MOB14, MDdMG09, NESP10, NBDB04, OM19, PB11, PRR03, RKG03, RFS03, SCD11, SCALFG^{+}18, SS17a, Tan11, TA13, TMB12, TB13, TZY08, WWCZ15, WR08, XOF05, ZIT^{+}13]. Registration-free [JGP19]. Regression [AS17a, LSW18, ABLL19, CZ14, CLZZ21, CFM^{+}13, KGB17, LY05, LTY^{+}15, LJC^{+}23, OZT19, RDSF15, SFK24, SHW24, VBVB19, YGC15]. Regular [BM98]. regularised [VWMZ15]. regulariser [KdRM^{+}23]. Regularity [Kis96a]. Regularization [DH19, RM02, ÅS17b, ALM23, AZ15, GY19, GGH^{+}24, JHA17, LEB07, PV14, QCXJ19, SM13a, ZA22, ZYZ24, ZFG^{+}22]. regularizations [LWLT17]. regularized [BGE^{+}17, BvdHL^{+}13, DBT^{+}17, WZX^{+}14, YLA09, ZXC^{+}20]. regularizing [AM15]. rehearsal [ZSK^{+}23]. Reillumination [Wor05]. Reillumination-driven [Wor05]. reinforced [CKL18]. reinforcement [SP23b, SXZZ24]. Rejection [OSM16, Bar18]. Related [GK98, Ros00a, PZM^{+}21]. relation [FO18, LCG^{+}24, OVJ^{+}21, QCL^{+}23, SM23]. Relational [COW98, CS00, Gwa17, ODT17, PLLL03]. relations [FAB12]. relationship [STC14, SCC^{+}22]. Relationships [KW00, BCHR24, JSRS08]. Relative [Chu02, SU01b, VAC16, CUSZ07, GSGJ24, OGB14, RA15, SM17]. relaxation [GL19, LC14, LPZ08, OEK08]. relaxed [WS06]. Relevance [MBKB02, MIUS16, PBQ99, FSL24, MdOBA19, MW13, Pen03, RLG^{+}14, SR23]. relevance-aware [FSL24]. Relevant [JDP97, KLKF20, NY14]. Reliable [CDT11, LRW08, LCG21, WPZ^{+}18]. relighting [SXZZ24, WLZW04]. relocalization [DSK^{+}20]. remote [CP21, CBB19, DFSC20, DWC^{+}24, FDC^{+}19, JZL^{+}23, MRH19, OM19, ÖÜ20, SVF^{+}21, XHX^{+}19]. Removal [CLCZ23, FMS17, YWL^{+}20, JCLZ21, LRZ^{+}19, LWLP23, SBN^{+}24, WAPB17, WZF^{+}24, WZX^{+}24, ZCW^{+}23]. removing [CYC10, LB05]. Rendering [EK98, CMZ24, CACB17, RLF15]. Reparative [YH19]. Repeated [CCS01, GS06, PGGM04]. replay [RMC^{+}22, ZTB20, PSB^{+}24]. Reply [Åst97, Col97, HM97, May97, Ver97]. Report [BVWS21]. Representation [BCC16, BB95, CF01, CWH^{+}13, CM99a, DT97, GK98, HGB98, KCD00, KD96, Mok97, WLL^{+}22a, ZSG^{+}20, ZT98, ZXK02, AQ09, AWK04, ATC^{+}13, Bar06, BYJG23, BFMW23, BSMK13, CPP^{+}11, CDIF14, CG04, DBF04, Dam08, DFJL15, DGRS22, FPC^{+}08, GXC23, GZL^{+}23, HRHZ17, HNB04, HWT^{+}24, JSC23, KM03, LLL15b, NLW^{+}17, PHHL23, PD11, PLKP23, RK11, REF15, STV09, SGMC15, SZW^{+}21, SBM^{+}06, SSS13, SY11, SWS11, TST14, TPD^{+}16, TCM18, VBS^{+}04, VGLP17, WWCZ15, WSY^{+}16, WRB11, XW16, XMT22, YWY^{+}16, ZLZH17, ZT09, ZH04, BS05]. representational [ZCWH23]. Representations [Ano15o, FPDK12, GK98, GJP96, HTEB11, KP00, LV96, NVWV97, ÜE01, AXJE21, BKK11, CKPV21, HS06, LZW^{+}24, LZTX24, NHTG15, OGH04, RSBA23, SCMP14, VAC16, XYRS17, YDP^{+}20, YZX^{+}17, YYH^{+}23, ZZK^{+}20]. representative [DPRC17, GDIIHK11, LLL15b]. Representing [NL96, TAK09, YS08]. reproduction [LMC09]. repulsion [RM03]. requirements [ES06]. resample [CKF18]. research [TGM^{+}17]. residential [ÜB05]. Residual [CLCZ23, LHZY19, SKT18, WGZL20, XSZ^{+}20, PKC^{+}18, RBdDS14, SKLM22, YFX^{+}18, ZCW^{+}23]. residuals [ZP18]. resistant [HKWC14, RK11]. Resolution [CJC01, MCPB99, PE09, PCJC98, WZWT99, AM06, AAMO16, AKC11, AYG23, CSS^{+}13a, CD10, CWW^{+}22, CLA^{+}17, CU20, EH21, FSV07, FDSB22, GB22, HSJS10, HL23, KKP24, LT05, LEE^{+}18, LLF18, LKZ20, LWW^{+}21, LZWX21, LN10, MYYY17, MHAF13, NFSD13, QGZ^{+}23, RT14, SA15, SRM20, SP06, TDV15, WDC^{+}24, WGZL20, XWC^{+}23a, XSZ^{+}20, YFX^{+}18, YGC15, YSZ23, ZHL^{+}20, ZH04]. resolutive [Pat13]. resolved [JC06]. Resolving [CLA^{+}17]. Resonance [RMFB02, CCR^{+}05]. resource [MFG10]. resource-constrained [MFG10]. respect [BFR13]. response [TS16]. Rest [RM02]. restoration [AGL23, CWC^{+}20, GY19, GGP23, HMA10, LWLT17, MWF07, PSY^{+}21, SZL^{+}23, WHL^{+}20, ZXC^{+}20, ZLL^{+}24]. restricted [LWLS12, NWJ15]. Results [BNG02]. ResUNet [JZL^{+}23]. retargeting [OAGN18, ZDF10]. Rethink [ZSL^{+}24]. retina [BEK18]. retinal [KACR^{+}23, NBDB04]. Retinex [LWFL24, NGR24, TYH^{+}21]. Retrieval [APV99, BS99a, Car01, Doe98, GFS04, JEK98, KB98, KR98, MBKB02, MKK02, MK01, PBQ99, SLST99, SBK^{+}99, SPK^{+}02, Sup02, ÁB13, ABI^{+}04, BRPC17, CEO18, CBB19, CHC11, CWLJ13, CNS18, DSY10, FSL24, FLHK08, FO18, GSS12, GH08, GCPF08, HMC10, Hei04, HC13b, HGS08, ILRB04, JWG04, JN09, KHH^{+}12, KSL^{+}20, LLG^{+}14, LLL^{+}15a, LLNZ22, LDCX24, LW18, LK03, LZWP03, LC09, MSG10, MIUS16, MLK21, NHK08, Pen03, PV14, PA10b, PFGG09, PR03, PBG04, Pun03, QLY^{+}17, RB18, SP23a, SLS03, ST20, SHS^{+}23, SYL^{+}24, TLEF06, TPT17, TBFJ15, YWY^{+}16, YARL^{+}20, ZSDK19, ZTH^{+}11, ZYXZ13, ZTH^{+}14, ZZCL14, ZDZ^{+}23]. Retrieving [LF08]. Retrospective [KW12]. Reverse [EFF98, SOJ^{+}95, WHS^{+}24]. Review [AC99, Ano95d, Ano95e, Ano97f, BL98a, BSBW14, SDK^{+}24, BZ14, DMSM21, EBN^{+}07, HRHZ17, JKW^{+}21, KHA^{+}05, MRdRGC23, PS15, RN12, SBIK16, SV14, WTZ^{+}21, Ano98c]. Reviewer [Ano12n, Ano13p, Ano14g, Ano15p]. Reviewing [Jon97]. Revisiting [BIG^{+}23, JRH24, PAK19]. revolution [BCLNG18]. reward [KS12]. Reweighted [XXCR15]. RFCNet [ZA22]. RFID [GLOC10]. RGB [BÁRM23, CICN22, CZHT15, DTL17, DSK^{+}20, GC19, HBKG22, KSF16, LM16, LYKY19, LLZ23, PYGGLNG17, PMCN22, QCL^{+}23, RXDS22, RZZ23, SBMM15, TPDP20, TS17, WLO^{+}18, YCZ^{+}23, YLX^{+}18, ZC19]. RGB-D [CICN22, CZHT15, DTL17, GC19, KSF16, LM16, PYGGLNG17, PMCN22, RXDS22, SBMM15, TPDP20, TS17, YLX^{+}18, ZC19]. RGB-D-based [WLO^{+}18]. RGB-depth [HBKG22]. RGB-IR [LLZ23]. RGB-only [YCZ^{+}23]. RGB-to-Infrared [QCL^{+}23]. RGBD [KCM^{+}17]. Ribbon [MWL99]. Ribbon-Based [MWL99]. Ribbons [MWLA99]. Ricci [CHZ^{+}13]. richness [EK12]. Riemann [Lil97]. Riemannian [AAASC11, KG14, YG16, YG17, ZRKZ^{+}11]. rig [HC13c, KD10]. right [BMB^{+}17]. Rigid [LHH97, AMNCM16, AMN18, BY12, BPLT15, CR03, GRB13, LNM^{+}21, LST13, LÁB15, LW18, NKPT13, NESP10, PRR03, PV06, RKG03, SCALFG^{+}18, SS17a, SKH08, TMQM13, TS17, WWCZ15, ZDZ^{+}23]. RIMOC [RAP16]. risk [ACC^{+}16, BJS14]. RKLT [SYF99]. RL [SXZZ24]. RNN [ZLLP21]. RNN-free [ZLLP21]. RNNs [CICN22]. Road [BW11, Gui98, Gui99, Gui00, AS23, BRA^{+}10, FFY^{+}04, JZL^{+}23, ZKSV18]. road-sign [BRA^{+}10]. roadway [MZB^{+}10]. Robot [CEA16, SIT07, GLOC10, MFS^{+}07, MLH13, PBT14, ST10, ZMJ^{+}15]. Robot-vision [SIT07]. Robotic [BL98a, JZWD16, MML^{+}16b]. Robotics [FKL^{+}16b, FKL^{+}16a]. robots [KON^{+}17, ZKRH04]. Robust [ACP16, AM04, Ano01m, AVC19, ARFF18, BA96, BGK98, BZ14, CSY08, CYD^{+}22, CTE95, CK09, CCYC12, DB03, DG01, FR11, GTP18, HWL^{+}22, JBR08, JSC23, KGC05, KK07, KB00, Lai00, LB00, LWH^{+}23, LLS^{+}23, Lin02, Luc01, MAG^{+}16, MLZRK24, MY95, MGK00, MK01, MFS^{+}07, MST00, NDBT95, PYS03, STHBH18, SFK18, SMK02, SAC^{+}12, TB99, TZM98, TZ00, VSP06, WLD99, WGAD14, WLW^{+}16, WWW95, XFSC13, YWZ11, YGH11, YZL16, ZYXZ13, ZSL^{+}16, ZJ05, BSM10, BI10, BL14, CWC^{+}20, CYY^{+}23, Cou13, DLC14, EF14, FN14, FLL^{+}23, FS03, GGH^{+}24, GG09, GCFMT12, GD19, HBH10, HBH11, HDF12, KLK14, KKSH23, KCZ18, KSHE20, KBJ^{+}10, KTV17, LRW08, LJC^{+}23, LLG^{+}23, MML^{+}16a, MMP15, MCF10, NAS^{+}17, PB11, QDLB17, RPBK22, SSL^{+}12, VMC^{+}16, WB12, WWCZ15, WCZ^{+}20, WLL^{+}22a, WCYS13, YXW^{+}24, ZH18, BETV08]. Robustly [BFY00, TS11]. robustness [HSK23, LZQL24, MN06, RPG12, SLK23, SAV24, TKK24, XSL^{+}23]. ROC [BKD01, SJST07]. rock [TN08]. rocks [TN08]. RocNet [LMM22]. ROI [BRSSAL11, TVLS08]. ROIs [RSY22]. Role [Hen98, Ham05]. Rolling [FDW21, NL17]. Rolling-Shutter-stereo-aware [FDW21]. room [GPC^{+}10]. ROS [GTMR23]. Rosenfeld [HM97, May97, Ver97]. rotating [TAK09, TM04]. Rotation [AMCB20, EA95, Pun03, TBFJ15, BDVK10, BYJG23, FCOK24, HAT^{+}15, LCP13, SBPF17, ZZL13]. Rotation-invariant [Pun03, FCOK24]. Rotational [YZX^{+}17, GYF18]. Rotationally [SK02]. rotations [CSW^{+}24, OK04]. roto [ANHGS17]. roto-translation [ANHGS17]. rough [AZP14, SB13]. route [MSSS09, MRdRGC23]. RTI [MC20]. Rule [DY98, KW00, LL99, DK13, JZZ23]. Rule-Based [DY98, KW00]. Rules [BS00b, BDFG17, SYK96]. running [LWlZ16].

S [CHC11, SCR^{+}17]. S-Cube [CHC11]. S-Hock [SCR^{+}17]. saddle [Kui08]. safe [NPM^{+}16]. safety [OBTMT15]. Saliency [BSF02, PBPD^{+}17, AvdWDM18, BWG17, FXWW17, LTY^{+}15, MMLC23, MSP^{+}18, REF15, SY20, WZY13, XHX^{+}19, ZWY14, ZGC20, ZWY14]. saliency-based [BWG17]. Salient [CM99a, PF99, SM99, ZLZH17, BB15b, CVP10, CM16, DFP23, DWC^{+}24, GZX^{+}23, JRBD^{+}15, LYX^{+}21, LYX^{+}24, RXDS22, WZCG24]. Same [DAZ^{+}17]. Sample [CM99a, BMvT^{+}19, HBB^{+}12, NAS^{+}17, WCYS13]. sample-and-filter [WCYS13]. Sampled [SWS11, LLWX24, PPT06]. sampler [JNLG15]. samples [LWH^{+}23, YZX^{+}22]. Sampling [IF99, STHBH18, Tan95, TB23, BW11, Bar07, CCD11, HMA10, KL11, MT16, SBB18, WDB12]. Sampling-based [TB23]. Sampson [SCEvdH14]. SAPS [ZZSD21]. SAR [HMEB07, RDT^{+}19]. SAR-Theory [HMEB07]. Satellite [MAM97, KSG^{+}19, PK18, QAB^{+}11, SO07, ÜB05]. Satisfaction [BZ99]. satisfy [ES06]. Savitzky [HTNN18]. SCA [THH^{+}23]. SCA-Net [THH^{+}23]. scaffold [CLK09]. scaffolds [CK11]. Scalable [KOC17, WM20, AMN18, CFCP11, CLL^{+}14a, GB08, MĆK09, NS16, ŠRDC09, ZTH^{+}14]. Scale [FT98, JC98, PCJ14, SUO00, SA02, SPQ^{+}17, TWW14, XHJF12, XSZ^{+}20, ZKLZ23, ANHGS17, AMMV99, ALIRT18, BKK11, BDS12, BPC^{+}17, BDL^{+}06, CGD^{+}23, CDJM14, CEO18, CGR13, CXYZ24, CHC11, CPS10, DLBG19, DSH04, EXP^{+}20, FPDK12, FMG23, GE08, GG20, GPY^{+}07, GDCM17, HDZR24, HMCT22, IZKB12, KL07, Kui08, KON^{+}17, LS08, LLL^{+}15a, LZmC^{+}17, LBNS09, MUS06, MNL^{+}17, MSW15, MYC^{+}14, OB14, PKvGS16, QSXS23, QBZ21, RXDS22, RTM^{+}17, Sah05, SOK16, SSL^{+}12, SP23b, SSHP17, SZL^{+}23, SvNW23, TTN17, TS17, TKAK14, TY22, TL15, WL15, XSD12, YWZ11, YSS^{+}14, YWY^{+}16, ZTH^{+}11, ZYQ^{+}23, ZYZ24, ZCW^{+}23, ZUS06]. scale- [ZYQ^{+}23]. Scale-Based [SUO00, ZUS06]. Scale-space [XHJF12, ALIRT18, BDL^{+}06, FMG23]. scale-spaces [GE08]. scale/irregular [VRKL13]. scaled [IH15, LYKY19]. Scales [BL98b, MKY01, LML^{+}23]. Scan [JB99, YYL96, AAB19, CACB17, NESP10]. scanner [FK09, ZG06]. scanning [LCT09, SO07, WWLV11, YGH11]. Scans [SPQ^{+}17, CPS10, NB10, SW04, SKSR08]. scanty [VGSMN16]. Scattered [OG98, Kim04]. scattering [FSI21]. scenario [LRFF24]. scenarios [CEA16]. Scene [AYB^{+}18, Bic98, CFM02, Che00, CBB95, DC00b, HFKN97, KW00, MDK^{+}24, MNE00, MJS97, MMP09, PD17, SB00, Ste01, TY05, TL16, WSJ15, WWXK24, XL98, YW16, ZT98, BKPS15, Bar07, BC10, BSH22, BYJG23, BCM06, CICN22, CGU11, CSS^{+}13a, CLZZ13, CFM^{+}23, CG04, DFJL15, DCH12, DZZ^{+}23, DSK^{+}20, EOPS22, GF15, GDM14, HUI16, HL13, HMB17, JY14, KK07, Lhu08, LS08, LRF^{+}17, LCG^{+}24, LDT21, MCM^{+}17, MAJ16, PGP15, PBW14, STV09, SFK24, SPK23, SPRS23, SPW15, SZGK24, TL15, TDZ^{+}20, VCDS^{+}17, YGJ^{+}20, YT13, YARL^{+}20, ZLY^{+}20, ZH04, XP11]. Scene-Based [Che00]. Scene-consistent [TY05]. scene-specialized [MCM^{+}17]. Scene-specific [PD17]. Scenes [BM99, BFF97, CCS01, FRL^{+}98, HGB98, SA02, SPQ^{+}17, AAMO16, AS23, BAPXH16, Bar05, BSZ^{+}21, BSRV17, BP09, CLA^{+}17, DWB11, DTL17, HHG^{+}20, HLL^{+}23, HML15, MTC^{+}14, MMP09, PLB16, SFF^{+}18, SCL13, TS17, TN07, TD19, WRKP05, XZQJ21, YR06]. Scheme [SYF99, YW99, FF23, GBY21, KKSH23, LZYW23, LDC^{+}13, LBNS09, NHK08, NBDB04, TT16, WNH05, ZJZY16, ZZ07]. Schrödinger [BLJ^{+}23]. Schumaker [Ano95d]. Science [Åst97, Col97, PRW97a, PRW97b]. Scientific [Ano95e]. score [XMT22]. scoring [GMF14, JHK24, PKvGS16]. Scribble [LYX^{+}24]. Scribble-based [LYX^{+}24]. script [SYZ^{+}15]. scripted [RLMK15]. SDART [BTB14]. SdcNet [MW22]. SE [ADC19, ARFF18]. sea [Cha21, Cha24]. seam [ZHS^{+}24]. seam-based [ZHS^{+}24]. Search [AM01, YT99, YLA09, CALO20, CLL^{+}14a, FN14, GK23, HMCT22, KSG^{+}19, KHH^{+}12, LWLC22, LCL^{+}14, LXW^{+}23, LXW^{+}24, MU11, RSS07, ST10, SM13b, TMS20, TYDH18, VJ17, WZY14, WZQ^{+}24, XTZZ14, XST04, ZWT^{+}14, LLWX24, LEA^{+}10, TYDH18, ZZSD21]. Searching [HP96, KAES99, MRF96, DR04]. Second [Ano95a, RM02, LEE^{+}18]. secret [CJL06]. Secrets [HBG13]. Section [CV13, FHSKP13, FFL14, VTRC14, YSS^{+}14]. sectional [EX17]. sections [NRJ11, Tan11]. security [CJL06]. seedling [KM03]. Seeds [SU01a, CUSZ07]. Seeing [RG10]. Seg [DRTT24]. Segment [MNHO00, FS03, IT15, LK03, XSK15, DGG08]. Segmentation [Ant98, BM98, BL00, BS00b, CM97, DH00, DV98, DCS05, HGR^{+}13, HY98, Jon99, KSI98, KVdG^{+}97, LM99b, LL97b, MNE00, MGCS17, MY95, MS97b, MS00, MCPB99, ME98a, NVWV97, PF99, PB99, RWWH00, RMFB02, SUO00, SU01b, SMK02, SA95, SBPF17, SC98, TK97, WF02, WWJ13b, YHN11, YYL98, AH24, AA20, ABJ^{+}21, AS09, ABEN09, AHDM10, ABLL19, ASFP03, AS23, BYR17, BUD19, BB16, Bar07, BSALF18, BP05, BCHR24, BvdHL^{+}13, BMB^{+}17, BCA16, BPB13, BSH13, BVCP21, BP09, BF10, CMBV04, CFYU12, CT10, CP21, CUAT13, CZ14, CJYW23, CLDP23, CE17, CO16, CLA^{+}17, CU10a, CU10b, CU11, CMCM16, Cre08, DRTT24, DBZ07, DPM14, DBT^{+}17, DFLS23, DB14, EPH^{+}21, EF14, ECC18, EX17, FLS^{+}14, FBS21, FAB12, GFL^{+}11, GBHS06, GKBW14, GCEC07, GB13, GBL08]. segmentation [GDR04, GDM14, GPDR13, GW07, GML16, GWFF22, HDS08, HWG21, HWZ16, HC13a, HSS^{+}16, HBH10, HBL^{+}17, HWW23, IJDAB13, JRS21, JSWL24, JLD13, JMPG11, KS15, KSRS16, KBN12, KK13, KKKS24, KSF19, KGU10, KACR^{+}23, LMP^{+}19, LvdHK^{+}15, LNN^{+}19, LV11, LPS^{+}11, LAFLB16, LWLT17, LSH19, LYX^{+}21, LZZ22, ML13, MVP06, Mah16, MMK04, MCCRAC20, MTP21, MO11, MSW15, MGPP11, MZ21, Mig12, Mil09, MBMC11, MAK^{+}17, MB05, MSF^{+}12, MPK24, MPPP14, MTR^{+}23, NRJ11, NF21, NHSC09, NN04, PJW11, PYWZ17, PLJS14, PNSF21, PV15, PGP15, PCR^{+}04, QAB^{+}11, QTLP22, RDA^{+}15, RBdDS14, STHBH18, SCE04, SOL14, SOL16, SM06, SG11, Sha05, SF07, SY20, SMD^{+}08, SCvW11, SVA^{+}22, TTN17, TA13, TPT15, TN08, TRG^{+}13, TC11, VMP03, VPP^{+}23, WO10, WSSS13, WHC14, WW16, WZW17, WLSO23, WRB11]. segmentation [WS06, WSKH13, WWJ13a, WWWF23, XWC^{+}23a, XST04, XAB07, XYW11, XWLY23, XWC^{+}23b, YK24, YZT^{+}13, YWMS08, YGC13, YJA96, ZDLS13, ZSCP08, ZFG08, ZRL^{+}11, ZA22, ZRZ^{+}24, ZLS^{+}13, ZFG^{+}22, ZUS06, ZU09, dMFU10]. Segmentation-based [HGR^{+}13]. segmentations [CCTCR09, KSG^{+}13, LH95]. Segmented [Pla96, EHG^{+}10]. segmenting [BBK14]. Segments [Cre99, GBB98, HMB17]. SegNet [AH24]. Segregation [JKM07]. Seidel [CRC97]. Sejong [CM21]. Selectable [DT96b]. selected [HKK08]. Selection [BL98b, BS00b, ET15, LSPV04, SM97, BPBS13, BEGB13, CYNO11, CZ14, CZS^{+}20, CLZZ21, DPRC17, GBHS06, GFW13, HG11, KY06, LvdHK^{+}15, LK03, NAS^{+}17, NHH14, PZX13, SO07, SB13, SF16, TG11, TKV16, TKAK14, YSL^{+}14, YZL16, ZRL^{+}11]. Selective [CHMG12, HH05, OH05, PZM^{+}21, WRKP05, DL05, GZJ05, LDC^{+}13, MTG07]. Self [BPCT22, CXFS06, CPPY21, DRTT24, DWW^{+}12, DC01, FCOK24, HDZR24, LPSK23, LZZ22, LLS^{+}23, LWLS12, LWLP23, NL17, PBD20, ZYZ24, BIG^{+}23, CE14, DDZ^{+}23, DZZ^{+}23, FPMK19, FK09, GB13, JLZ23, KKKS24, MDK^{+}24, MM21, NF21, PD23, QC04, RSL10, RSBA23, SIRS21, TLEF06, TM04, WK21, YK24, YXW^{+}24, ZZG^{+}24, ZDF10, ZZSD21]. self-adaptive [CE14]. self-attention [NF21, YK24]. Self-attentive [CPPY21, ZZSD21]. self-avoiding [GB13]. Self-Calibration [DC01, CXFS06, DWW^{+}12, LWLS12, NL17, FK09, MM21, QC04, RSL10, TM04]. Self-correcting [LLS^{+}23]. self-distillation [FCOK24]. self-driving [FPMK19]. Self-knowledge [LPSK23, LWLP23]. self-organizing [TLEF06]. self-paced [DDZ^{+}23, SIRS21]. self-rectifying [MDK^{+}24]. Self-Supervised [HDZR24, DRTT24, FCOK24, LZZ22, PBD20, ZYZ24, BIG^{+}23, JLZ23, KKKS24, PD23, RSBA23, WK21, YXW^{+}24, ZZG^{+}24]. Self-supervision [BPCT22, DZZ^{+}23]. Semantic [ABC^{+}03, CP21, DBT^{+}17, dSEdSPdMF24, GMW12, GLMM16, GDM14, HAM^{+}16, TDV15, ZKLZ23, ZZS^{+}23, ZFW^{+}24, ABI^{+}04, ABLL19, AS23, BVCP21, CL15, CLDP23, CFM^{+}23, COV^{+}22, DRTT24, DLBG19, DCH12, DFLS23, FSL24, FBS21, GYTL09, GZX^{+}23, GWFF22, HBL^{+}17, HWW23, ILRB04, IJDAB13, JRS21, JSWL24, JN09, LG24, LYSS12, LZL^{+}17, LSTARMB11, LCG21, MLZRK24, MTP21, MZ21, MYC^{+}14, PSE^{+}11, PS22, PLJS14, PNSF21, SM12, SDK22, SY20, SYL^{+}24, TLP^{+}17, VPP^{+}23, VZP^{+}09, WLSO23, XST04, YK24, YSY^{+}18, ZG10, ZTH^{+}11, ZTH^{+}14, ZSC^{+}23, ZRZ^{+}24, ZYZ24, ZFG^{+}22]. Semantic-aware [ZFW^{+}24]. semantic-based [SM12]. semantic-guided [GZX^{+}23]. Semantically [CSZ^{+}15, FDSB22, LRF^{+}17, WK21, MC22]. Semantically-driven [CSZ^{+}15]. semantics [BYJG23, FYH11, PV14]. Semi [CLTW23, CLL^{+}14a, CZHT15, JA16, TLWT12, UU18, WHM^{+}09, ZJL23, BUD19, BCNS15, CJYW23, DWB11, DB14, HD23, KS12, KACR^{+}23, LHL^{+}21, LLS^{+}23, Mah16, MPT21, NN13, NWNT17, OZT19, VPL23]. Semi- [UU18]. semi-automatic [BUD19, BCNS15]. semi-interactive [DWB11]. Semi-supervised [CLTW23, CLL^{+}14a, CZHT15, TLWT12, WHM^{+}09, ZJL23, CJYW23, DB14, HD23, KACR^{+}23, LHL^{+}21, LLS^{+}23, Mah16, MPT21, NWNT17, OZT19, VPL23]. semi-transparent [KS12]. semisupervised [MP20]. sense [CWO^{+}11]. sensing [ASFP03, CBB19, DFSC20, DWC^{+}24, FDC^{+}19, GZJ05, JZL^{+}23, LSKK10, MRH19, OM19, OH05, ÖÜ20, SB96a, SLK15, XHX^{+}19]. sensitive [FWG18, KLL^{+}11, MFP^{+}20, SPT^{+}18, ZWZZ18]. Sensitivity [LFMP13, LP10]. Sensor [MG95, TG95b, YT99, AZSVK05, CA10, CP21, CÇ15, GTMR23, HCC^{+}16, LSKK10, SPC^{+}15, TDWH07, TMB12, YHS95]. sensor-based [HCC^{+}16]. sensored [CD10]. sensorial [CCR^{+}05]. sensorimotor [TPDP20]. sensors [IKST05, STC^{+}16, SM21]. sensory [OGH04]. sentence [WLZM20]. sentence-level [WLZM20]. sentimental [RMS^{+}19]. separated [ZhZFL22]. Separation [AO16, AS09, ZZZP09]. September [Ano21x, Ano23x]. Sequence [CA97, LCZ^{+}16, LZ97b, NDN^{+}97, WALL00, XS98, FR11, GS06, JM09b, NSEA13, PGGM04, Rem04, ZZZ06]. Sequences [ALK99, CW00, FRL^{+}98, GMW12, GHS95, IP98, KSS97, PM97, PF01, RWWH00, SF95, SBZ97, TPR^{+}00, WN99, WLD99, ZW97, ABB^{+}23, BYR17, BF07, BPSV16, CXFS06, CSG^{+}03, DCS05, DZLH17, DHP08, GBY21, HJ12, HDG^{+}14, KIS17, LSC08, LS08, LWH03, MC09b, NT10, Neg12, OSM16, PBI16, RM03, SM24, TY05, TS16, TVC09, VMC^{+}16]. Sequential [BSF02, FAB12, HW06, SYK96, SZ16, SAC09, SHS03, WS08, ABK16, VB16].

Serial [TV99, Tan11]. Series [MRW^{+}97, CKPV21, LEA^{+}10, MOT17, TYH^{+}21]. service [MFS^{+}07]. Set [ACF00, Bic98, GAD01, LLSV00, TLFM23, TS00b, ZOMK00, CDT11, CBT^{+}04, CH17, CU11, DM12, FPC^{+}08, FCOK24, HWZ16, KK13, MMV06, PB11, PD05, SAS12, SG11, SRS11, WWCZ15, WHGZ20, XWDL23, ZdCR^{+}24]. Set2Model [VKL18]. Sets [DL97, KSKB95, KB95b, LER95, NG98a, Shi99, WB97, WB01, ADC19, BFR13, CSZ^{+}15, CP20, Cre08, DCS05, GDCM17, HY11, KKSC23, MGS15, SM06, Sha11, WK21, dCCP12]. Setting [KTP08]. settings [KACR^{+}23]. setups [FPMK19]. Seven [SOD10]. Seventh [Ano96a]. several [SKA23]. severe [WZX^{+}24]. SFM [CX11, FAZ14, CCL^{+}17]. Shading [BHMB10, KP97, KB95a, KB95b, LK97, OD97, SKB96, BLKG21, DFS08, KN03, Wor05]. shadow [CYC10, SCE04, SBN^{+}24, WZF^{+}24, WCF10, WZCY22, YZ06, ZFW^{+}24]. shadows [CF07, JF10]. Shah [SOL14, SOL16]. Shape [Ano15o, ASZ99b, BH99, BCG95, Boo97, COW98, Car01, CPC99, CCP97, CTF^{+}98, CFA98, CCD11, DT10, DM01, DC98, DY98, DT97, FW97, HF01, Hob00, JC98, JEK98, JMPG11, KP97, KB95a, KB95b, KR98, LPC08, LL99, LK97, LYG07, LK00, Mas02, Mok97, MPPG98, NSK^{+}97, NN18, Nis96, Nis99, OD97, OBH04, OH04, PEFM98, PV97, SKB96, SP97a, TI01, TSP97, TFL^{+}09, TZY08, YFZ98, ZOMK00, AAASC11, ALM23, BF07, BvdHL^{+}13, BL16, BY12, BGK95, BSBW14, BF10, CLZY15, CH06, CK11, CC11, CUAT13, CZ14, CPPY21, CL08, CLCO13, CT13, Coe12, CTCG95, DZL07, DFS08, EL07, EOPS22, EK14, FPC^{+}08, Goh08, GKBW14, GHML17, GPDR13, GWFF22, HFR06, HG11, HC13c, KK15, KSL^{+}20, KŽ12, KNO^{+}09, KRS14, LLS21a, LE09, LPS^{+}11, LC14, LLG^{+}14, LLL^{+}15a, LQQS21, LSGY24, LDCX24, LPZ08]. shape [LW18, Liu10, MDFS11a, MC09b, MWTN04, MIP16, NHK08, ÖÜ20, PAK19, Pen15, PBG04, PS12, RK11, RAHT11, Rem04, SECS15, SPT^{+}18, SBM^{+}06, SK15, SM13a, SY11, SH08, SWS11, SKBS13, TG11, TWS06, TMQM13, TÉSK11, TH04, TC11, WB12, WYC15, WSKH13, WSJ15, Wor05, WWJ13b, WPB^{+}14, YB07, YZT^{+}13, YWY^{+}16, YZX^{+}17, YFF^{+}23, YLA09, YZL^{+}21, YARL^{+}20, YG16, ZZC^{+}13, dSM14, MIP16, NLW13]. Shape-based [JMPG11]. shape-color [GHML17]. shape-constrained [WWJ13b]. Shape-from-recognition [TFL^{+}09]. shape-from-shading [DFS08]. shape-texture [HG11]. Shaped [GSP01, TA13]. shaped-based [TA13]. Shapes [ANM98, KS96, NWP97, Pla96, ST96, Sup02, AMNCM16, AC07, BSH13, CDJM14, CKK^{+}12, FO18, GR05, HW06, IAP^{+}11, LBNS09, Sha05]. Shared [ASZ99a, KSL^{+}20, LLZ23, QCH20]. Sharing [MvGS16]. sharper [SRM20]. sharpness [RF23]. shearlet [GY19]. sheetmetal [ZZZ06]. shift [KG14, ZYS09, ZLS^{+}13, LLR10]. shifts [GLG22]. shoe [ZSDK19]. shop [ZSDK19]. shorelines [BKP10]. Short [NB20, SPK23, WHL^{+}21, WB15]. Short-Term [NB20, SPK23, WHL^{+}21]. Shortest [DJG01, DBBB14]. Shot [Che00, YFDA17, YW99, ZYQ^{+}23, BPCT22, CLL^{+}21, CWS^{+}24, CLDP23, DWL19, DFH^{+}22, GBY21, JGP19, JHK24, LHL^{+}21, LLNZ22, LFLZ23, LZW^{+}24, LCG21, MHX19, PHHL23, SAK^{+}24, SOD10, WZQ^{+}23, WZQ^{+}24, WKT22, XZX^{+}21, XWLY23, YWM19, STD14]. shots [NY14, MNR18]. should [CL17]. Shutter [FDW21, NL17]. siamese [YQL^{+}23, CBB19, HDZR24, JLZ23, LLG^{+}23, MP20]. SID [PZM^{+}21]. SIERRA [FLL^{+}23]. SIFNet [UJ22]. SIFT [LS09, XHJF12, ZYS09]. SIFT-like [XHJF12]. Sign [CW00, OD99, VM01, BRA^{+}10, FFY^{+}04, GFLW24, KFN15, WCZ^{+}07, YS09]. signal [Jea11]. signals [Pey09]. Signature [DLHT99, MKK02, NN18]. Signatures [Hob00, SC00b, LW18, PG13, STD14, YZX^{+}17]. Signed [Mas02, Gre04]. signers [KFN15]. significance [OVJ^{+}21]. Silhouette [AAASC11, BL01, ES04, CT13, DPM14, LPC08, LYG07]. Silhouette-based [AAASC11]. Silhouettes [HCHD01, Lau97, DT09, KK15, SY10, YW07]. SIMD [MHSP10, TV99]. SIMD-based [MHSP10]. similar [KBMD15, MHMO09]. similarities [PG13]. Similarity [BJ97, Car01, Hen98, KAES99, STLH08, TP05, YK08, BB13, BB15a, BAP08, CLL^{+}21, CK11, CL15, CLL^{+}14a, DL05, EK14, FLHK08, GKPS15, GCPF08, Got08, HBL^{+}11, JHK24, KH23, MGW10, NHK08, PKZP^{+}24, RKG03, SKLM22, SvdMH15, SZW^{+}21, TH04, WZY14, ZWT^{+}14]. similarity-based [NHK08]. similes [LWSC16]. Simple [ASS97, ASZ99a, DWC16, LCL^{+}17, YXW^{+}24, CO16, KA12, Loh10]. Simplicity [LM96]. simplified [BC10, LSL^{+}18]. Simplifies [Dan97, ZU09]. Simplifying [AM97, SdB03, ZdCR^{+}24]. Simulated [BCG95]. Simulating [HH05]. simulation [CWC^{+}20, JB15, PT15, SOL14, SOL16]. simulations [HMEB07]. Simultaneous [DC98, EFF98, HDL^{+}20, Jok98, JC06, Jur99, LEE^{+}18, LWLT17, LWLC22, LM99b, PA06, TRG^{+}13, VM01, WB01, CHH09, TTN17, WCYS13, ZZHZ23]. Single [AJ23, BK01, CC11, CLO17, CCS95, Gui98, HR99, JGP19, JCLZ21, KLY21, LA11, LRZ^{+}19, LN98, LZWX21, MBD^{+}22, OSY18, SZB^{+}21, Tay00, ZRRK18, AMNCM16, ATG15, AZP14, BM15, CG09, CH06, CWW^{+}22, CU20, DDZ^{+}23, DMW10, DSK^{+}20, FPMK19, GGP23, HJ12, HQW^{+}12, IDY^{+}18, KSR^{+}12, KCM^{+}17, KTP08, KS12, KM03, LC14, LYKY19, LWW^{+}21, LZmC^{+}17, LZZ^{+}23b, LDL^{+}19, MYYY17, MDdMG09, PD17, RRK13, RZZ23, RBA20, SPC^{+}15, SRB21, SFK24, SA15, SZGK24, WHC14, WGZL20, WHL14, XYW^{+}08, XSZ^{+}20, YWM19, ZZ07, ZZ20, ZIT^{+}13, ZCW^{+}23]. Single-camera [MBD^{+}22]. single-direction [HQW^{+}12]. Single-image [KLY21, MYYY17]. single-instance [ZZ20]. single-optical-axis [WHL14]. Single-Pass [CCS95]. single-query [FPMK19]. single-stage [RZZ23]. single-touch [WHC14]. Single-View [ZRRK18, HJ12, KM03]. singular [SCCP05]. Sinusoidal [GLR^{+}99]. Site [CJC^{+}98]. sites [AO04]. six [Sha11]. size [MGW10]. Sizes [Shi99]. sizing [TN08]. skeletal [DWV19, HRHZ17, PKC^{+}18, TH04, TVE^{+}16, VAC16]. Skeleton [PCM21, YH19, AAL22, BDT23, BFMW23, GBB^{+}18, LLG^{+}24, LDL^{+}19, RT14, SAdB14, TKK24, XWLY23]. Skeleton-based [PCM21, AAL22, BFMW23, GBB^{+}18, LLG^{+}24, TKK24, XWLY23]. Skeletonization [KSKB95, Pud98]. Skeletons [AM97, Che98, NSK^{+}97, TSP97, Cou13, GCS23, Goh08, Sha05, SdB03]. Sketch [LDCX24, BM15, BRPC17, CLTW23, eGZW07, HC13b, LLG^{+}14, LHSG15, LLNZ22, LZTX24]. Sketch-based [LDCX24, BRPC17, LLG^{+}14, LLNZ22]. Skew [Spi98]. Skewed [VMUO95]. skiing [DM24]. skill [LSP^{+}16]. skills [LWZC14]. skin [BDFG17, SJST07, XYW^{+}08]. skip [KHH^{+}22]. slalom [DZLH17]. SLAM [GC19, KD10, SE11, TWW14]. Slice [TST14, LSCK15, MDdMG09]. Slices [BS96]. slime [MYZ^{+}24]. SlowFastFormer [ZCW24]. Small [FT98, SJSL21, YHW^{+}23, CDT11, JYL23]. Smart [BKMV07, ACC^{+}16, CVP10, GPC^{+}10, HCC^{+}16, MCT10, MHSP10, NS16, WMBY12, Ziv10]. smart-room [GPC^{+}10]. Smartphone [LRD19]. smartphones [JRBD^{+}15]. SMC [MCM^{+}17]. smiles [JGP19]. smoke [BJS14, YSX^{+}19]. Smooth [BA96, NWP97, BL08, GR05, MJPS16, UK12a]. smoother [LV11]. Smoothing [CBM01, JC98, BI11, GS08, HWZ16]. smoothness [CL17, UO16]. smoothness-constrained [UO16]. SMPL [SWMM22]. Snake [Pet99, WWJ13b]. Snakes [RAH97, Sap97, BUD19, SZ07]. SnapshotNet [LZZ22]. snooker [DRK03]. SnooperText [MTC^{+}14]. Snow [CLCZ23]. Soccer [GLM17, ABC^{+}03, CL17, DLS^{+}09, FLB06, MEM17, MSSS09, ROJX09, VMP03]. Social [LCL^{+}14, SCC17, ADFR18, LLTL14, NHTG15, SM23]. Social-oriented [LCL^{+}14]. Soft [ZZCL14, KBMD15, Kim17, LBP23, YLM11, ZBDP15, TKV16]. soft-to-hard [LBP23]. Softassign [SAS12]. soften [ZJJ22]. softmax [JLM22]. solar [CF07, JF10]. Solids [RAH97]. Solution [Jur99, AMN18, DK13, Dre96, KKRK23, PW23, SZ16, WXWC18, WXC20]. Solutions [FKL^{+}16a, OD01, GYF18, KT08, KBJ^{+}10, LPR^{+}03]. solvers [IH15, KMT11]. Solving [KB95b]. SOM [MdOBA19]. SOM-based [MdOBA19]. Some [GK98]. Sonar [MCPB99, MCPB00, TS00a, TPR^{+}00, BSH13, Neg12, ZZHZ23]. Sonka [Loh10]. Sort [LK03]. Sort-Merge [LK03]. sound [LTL^{+}23]. Source [OD97, OD01, CF07, Dre96, GTMR23, IDY^{+}18, RAC^{+}13, SF16, TMNM09, YWH^{+}23, YHS95]. source-free [YWH^{+}23]. Sources [LZ97a, LF08]. Space [Åst97, BL98a, Col97, FT98, HRHZ17, HR99, HGB98, JC98, LL97a, Mok97, NKL24, Pet99, PRW97a, PRW97b, RC97, SC00a, SCS99, ZL01, AQ09, ALIRT18, BT05, BDL^{+}06, CHC11, DFP23, FMG23, FS03, GPY^{+}07, HKK08, JSRS08, KH13, Kui08, KDV16, LH95, LL08, LZL^{+}17, LN10, MHL14, SAC^{+}12, TH06, TCM18, VMP03, WMBY12, XHJF12, YHS^{+}20, ZMM^{+}22]. Space-time [HRHZ17, NKL24].

Space-Variant [BL98a, RC97]. spaceborne [HMEB07]. spaced [TN05]. spaces [BSBW14, CS07, EL03, Eva06, GE08, LTY^{+}15, QT10, WD14, dSdSF^{+}12, dLAH07]. spacetime [GBY21]. SPAMM [RAH97]. spare [MvGS16]. Sparse [CWH^{+}13, KP00, WXZ24, XWC^{+}23a, XMT22, ZKSV18, AO16, ARFF18, BGE^{+}17, BR12, CC11, CZ14, CLZZ21, CS07, DPRC17, FB12, JZZM23, KHR^{+}16, LY13, Lhu18, LDH^{+}14, LTCT14, Mal21, Pat13, REF15, SZW^{+}21, SCMP14, TS24, VBA19, WHL^{+}20, XXCR15, XW16, ZLL^{+}14, ZLFH23]. sparsely [PPT06]. Sparsity [CÇ15, QDLB17, RLG^{+}14, TLY^{+}16, XSQZ15, YSL^{+}14]. sparsity-constrained [TLY^{+}16]. Sparsity-driven [CÇ15]. Spatial [BL98b, CGL98, CA97, CJYW23, Dav97, DCFM07, KW00, KBMD15, PA00, Pha01, SYZ^{+}15, THH^{+}23, VBA19, WF02, WLL^{+}22a, XWDL23, ZD01, ALIRT18, BCHR24, BJS14, CSY08, CCTCR09, CHC11, DFLS23, FMGA^{+}12, FAB12, GLM17, GWFF22, Hei04, HGS08, HLKK19, KM17, KY06, LWZC14, LLL15b, MPF07, MP20, PSE^{+}11, PCM21, TLH22, TP05, WSSS13, WWJ16, WDB12, WYW^{+}22, WWWF23, YFX^{+}18, YSD03, ZTH^{+}11, ZWLH24, ZCLX20]. spatial-aware [GWFF22]. spatial-color [ALIRT18]. spatial-domain [TP05]. Spatial-Feature [WF02]. spatial-scale [CHC11]. Spatial-spectral [VBA19, MP20, TLH22]. Spatial-Temporal [WLL^{+}22a, WYW^{+}22, YFX^{+}18]. Spatially [Lai00, SPT^{+}18, ABLL19, KNL15, MLB^{+}18, SB96a]. spatially-aware [ABLL19]. Spatio [KYYC14, NDO09, Pet99, WX16, XZW^{+}23, CHMG12, CKPV21, CWLJ13, DLF06, FXWW17, LCSL07, LTY^{+}15, LXFM16, MTV17, NNS^{+}18, RL13, SA04, SCMP14, TKL21, TBC^{+}21, XYW11, CGHTK16]. Spatio-Temporal [XZW^{+}23, KYYC14, NDO09, WX16, CHMG12, CKPV21, CWLJ13, DLF06, FXWW17, LCSL07, LTY^{+}15, LXFM16, MTV17, NNS^{+}18, RL13, SCMP14, TKL21, TBC^{+}21, XYW11, CGHTK16]. Spatio-Velocity [Pet99, SA04]. spatiograms [MdBJG15]. Spatiotemporal [DIMT12, QSX17, TI01, WZJ^{+}21, BZS08, EDJ^{+}20, JYTK11, YSNiT14]. SpATr [BB24]. Special [Ano01k, Ano01l, Ano05j, Ano15o, ACW^{+}16, BPS10, CFS98, CA10, CKB10, CV13, DRDKE13, FHSKP13, FKL^{+}16b, FKL^{+}16a, FFL14, FHP01, FPDK12, FYH11, GHMT09, HMC10, HTEB11, HGSM11, JWDF05, Jon08, KB98, KPKH07, KLBP11, LBK10, MPF07, MYK03, MZL^{+}16, MYC^{+}14, NPBM22, NLW13, RFL02, STV09, SPQ^{+}17, SST06, THL13, Tho10, VTRC14, WPZ^{+}16, YSS^{+}14, BK15, BPQ15, DFJL15, LLE^{+}09, SMHH04, ZZP^{+}16]. specialized [AM17, MCM^{+}17]. species [CTM^{+}13, WM20]. Specific [DC00b, AZP14, ES06, LLZ23, NCDG21, NY14, PD17, SY20, XYZ16]. Specification [LD98]. specified [GS95]. specimen [MSG10]. specimens [KORC10]. spectator [SCR^{+}17]. Spectra [SB98b, DvLV08]. Spectral [BL04, SK15, BLJ^{+}23, BEGB13, CHP^{+}11, CPT07, DCFM07, GCEC07, HL23, KIS17, KC22, LW18, LZC^{+}20, MP20, OEK08, PTE12, TLH22, VBA19, WZY13, YSD03, ZRL^{+}11, ZWT^{+}14, ZZZP09]. Spectrometry [SGK00]. Spectrum [FHSKP13, GZL^{+}23, CSV^{+}16, HD07, QSX17, WB15]. Spectrum-irrelevant [GZL^{+}23]. Specular [CTE95, WXZG18, CKS^{+}05, LF08, ZMCA05]. Specularities [LKK00, LB05, OJRT08]. Specularity [LL97a, WAPB17, DJF14]. Speech [YH19, CSV^{+}16, GRMH19, PY08]. Speechreading [LT97]. Speed [DT96b, EA95, THT^{+}98, HSHA20, SJH17]. Speed-Up [THT^{+}98]. Speeded [BETV08, SW17]. Speeded-Up [BETV08, SW17]. Speedup [HBL^{+}17]. Spetsakis [Zha97]. Spetsakis-Aloimonos [Zha97]. Sphere [Lil97, ZDZ^{+}23, BBHF10, PHH^{+}15, SW13, TMNM09]. spheres [LP10, SBMM15]. Spherical [KHK10, ÜE01, AXSVL14, BI10, CHZ^{+}13, CPS10, KH15, RDM^{+}11, WLZW04]. spin [SOL14, SOL16]. spinning [ZSL^{+}16]. spiral [BB24]. Spline [LMP^{+}19, RFS03, LZD^{+}14]. Spline-based [RFS03]. split [UJ22]. split-inpaint-fuse [UJ22]. splitting [HZLM11]. Spontaneous [XFP^{+}16, LWSC16, WWJ16, ZZP^{+}16]. spoofing [CYY^{+}23, ZTB20]. SPORE [AS17a]. sport [DM24, MP09a, PD17]. Sports [MTH^{+}17, HKHE14, KPPK09, LHJ^{+}09, LWH03, MBCJ17, TGM^{+}17]. Spots [NS98]. Spotting [ZXK02, XFP^{+}16]. square [ZZ10]. Squares [FM99, ADC19, GSV05, HLKK19, MP09b]. squares-based [MP09b]. SR [AAMO16, DBT^{+}17]. SR-clustering [DBT^{+}17]. SSBD [ABK16]. SSD [JGP19]. SSDA [ZJL23]. SSDA-YOLO [ZJL23]. SSMTL [BIG^{+}23]. Stability [FT98, QV98]. Stabilization [CC00, KYYC14, SOJ17]. Stabilizing [FF09]. stable [LZTX24]. Stacked [ZH18, CWLY22, HCLZ21]. stacks [ZLHJ18]. Stage [SP97b, HWZ^{+}23, KSY15, LZZ^{+}23b, RZZ23, TDZ^{+}20, WTZ24, WLMG08, XZW^{+}23]. Staged [AS17a]. Stairs [PYGGLNG17]. stakes [SL16b]. Stand [OSM17]. Stand-alone [OSM17]. standard [KMBH09]. standardization [ZU09]. standing [TTN17]. STaR [CKPV21]. start [FN14]. Starting [WH18]. State [Par16, JM09b, KTP08, KDV16, LN10, Ros10, SOK16, SCD11, SHL18]. state-of-the-art [JM09b, SCD11, SHL18]. state-space [LN10]. states [FR11]. Static [LLG^{+}24, WY07, Bar05, CSG^{+}03, HKHE14, JY14, LLS21a, NHZ^{+}22, Rem04]. stationary [CMG16, OSM16, RSPD12].

Statistical [ABK16, KSG^{+}13, LK00, SECS15, SM13a, WZY14, BvdHL^{+}13, BW15, BSBW14, BF10, CLZY15, GMF14, GKBW14, GPDR13, HKK08, KGC05, KFN15, KY06, Nis96, SPT^{+}18, WLX^{+}14, WBS14, WS06, XSK15, YG16]. Statistics [FSA01, SJ15a, TLEF06, dSM14]. Status [BS99b]. steerable [AS08a]. steganographic [YCL07]. step [BYN^{+}04]. Stepwise [SL16a]. Stereo [AM01, BK16, BM99, CN95, CHRM96, DC00a, HLB17, HQW^{+}12, JPP^{+}14, KS95, KP97, LL97a, LSHT02, MS97a, Mur95, OD01, PW06, WZ08, AK10, AK11, APB10, Atk17, BN15, BCMCB09, BBC^{+}07, CMM20, CPP^{+}11, CC07, DBZ07, ES04, FDW21, FB05, FSI21, GBF12, HASS10, HBG13, HZW^{+}10, HKA13, JMPG11, KN03, KGFP10, KH15, KT07, LS08, LZC^{+}20, MSI10, MCT10, NT10, OSY18, PD14, SE11, SvdMH15, TPNP15, TB13, TKDN16, YA12, YK08, ZN08, ZKRH04]. stereo-based [MCT10, SE11]. Stereo-Motion [DC00a]. Stereoscopic [Jon97, KGM19]. stereotactic [MDdMG09]. stereovision [PCC13]. still [PL10]. stimulations [DDZ^{+}23]. stitching [ZHS^{+}24]. Stochastic [ADDK99, LRLB11, PB11, VB98, WZWT99, CAGN24, JSWL24, KK13, KL11, LLWX24, LRLR15, MSW15]. stopping [JZZ23, SYK96]. Straight [GL97, Sch06, Sha06, ZS11]. Straightness [Kis96b, MMS97]. Strategies [Goh08, LVW97, CUAT13, KTP08, KYM13, YLA09]. Strategy [BM99, DH19, YB95, Bar07, CKF18, CRCM16, DLV15, FSL24, GCPF08, JRH24, MFB11, SJSL21, TBZ^{+}24, WCYS13, WZCY22, YZX^{+}20]. stream [CYD^{+}22, KDSF20, WHL^{+}20, WZFL23]. streamed [LYKY19]. Streaming [FF23, FT23]. Streams [DH00, OYTY98, ADR16, ADFR18, DBT^{+}17, GGO10, PBD20]. street [STO17, ÜB05, YW16, ZSDK19]. street-to-shop [ZSDK19]. street-view [YW16]. Strength [SU01a]. String [CTF^{+}98, ZLLP21, ZNG^{+}13]. Strings [HY98]. Structural [MLP97, Nis95, Nis97, Nis99, SM23, WCH98, ALM23, ALY^{+}22, AM15, BEGB13, FLS^{+}14, KRBSV17, Nis96, SDK22, TAC23, YSL^{+}14, ZG10, SYZ^{+}15]. Structure [BS05, Bri17, CNO^{+}24, CJC01, DT96b, Jac01, KMB97, LLL13, LPH01, MS97a, MS96b, Oli00, Oli01, SBZ97, TO99, WD96, XS98, ZD18, AMN18, BPC^{+}17, GRCD18, eGZW07, KD10, KBWT16, KN03, KGK10, Kui08, LWY^{+}17, Lhu08, LCZ09, LLNZ22, MSI10, MBCJ17, NKPT13, PXTZ14, QCXJ19, RLS06, SZW^{+}21, TMQM13, TN07, TGFF15, WCYS13, WMZY23, XYZH11, YZT^{+}13, YT13, ZDLS13, ZLFH23, LY13]. structure-and-motion [TGFF15]. Structure-aware [CNO^{+}24, LLNZ22]. Structure-from-Motion [Jac01, Oli00, Oli01, BS05, BPC^{+}17, RLS06, LY13]. Structured [PWSvdH17, SLK15, WWG^{+}18, ZJW15, BHSD^{+}13, BB03, CCL^{+}17, GYWZ23, HW06, LCT09, VB16, WNH05, XSQZ15]. Structured-light [SLK15, BHSD^{+}13]. Structures [JDP97, KMA^{+}00, LHH97, FPC^{+}08, FAB12, KZ05, KSG^{+}13, RC13, YJA96, ZWZZ18]. structuring [BB16, SW05]. student [LDCX24, XQZL23]. Study [DF02, GMT00, HSSB98, LLNC20, LCZ^{+}01, Lin02, NESP10, AVGASAP15, CAO^{+}23, DBZ07, GGGROE^{+}17, GCFMT12, HS06, HF11, JM09b, PSE^{+}11, PWWQ16, SCD11, SYPK13, TPD^{+}16, VD10, ZK17, ZZJS18]. studying [CU11]. STURE [WLL^{+}22a]. style [DWLW23, GKH^{+}21, HKL^{+}24, WZC^{+}21, ZSL^{+}24]. stylization [CNO^{+}24]. Stylizing [SBH^{+}17]. Stylus [MWL99, MWLA99]. Stylus-Generated [MWL99, MWLA99]. Sub [YYZL19, AVGASAP15, GBF12, NRJ11, XJK12].

Sub-hypergraph [YYZL19]. sub-pixel [AVGASAP15, GBF12, XJK12]. sub-sections [NRJ11]. subband [BLJ^{+}23]. subclustering [BJ14]. subdomains [MJ17]. subgraph [CM16]. subgraphs [BG16]. Subgroup [HB98b]. subisomorphism [DSdlH^{+}11]. subject [ABB^{+}23, LY06]. subjects [SSS13]. Submarine [CC00]. Submersible [NK00]. submersion [ZRKZ^{+}11]. submodular [PRK19]. Subpattern [ME98b]. Subpixel [CL00, HTNN18, AT17]. Subsampling [CO16]. Subsampling-based [CO16]. Subsea [TPR^{+}00]. subsequent [DPCA15]. subset [MVP06, YO11]. subsets [BRP04, DSNN08]. Subspace [DSY10, TTXT21, DD11b, FLHK08, MMP09, XXCR15]. Subspaces [FB97, BDFG17]. Substrate [HT98]. subtraction [BT05, Cha21, Cha24, DS07, OSM17, SV14, ZY14, ZJZY16, ZCF13]. Successive [MDL^{+}23, SWMM22]. Successively [ZZ10]. sudden [MOT17]. Sufficient [Egg98, YYH^{+}23]. suitable [HZW^{+}10]. sumD [LWZP17]. summaries [AWK04]. Summarization [CB98, ALK^{+}09, BZS16, LHJ^{+}09, SLS03]. summarize [CH09]. Summarizing [PHY^{+}11]. summation [WB15]. super [AM06, AAMO16, AYG23, CWW^{+}22, CU20, EH21, FSV07, FDSB22, GB22, HL23, JC06, KKP24, LEE^{+}18, LLF18, LWW^{+}21, LZWX21, MYYY17, NFSD13, QGZ^{+}23, SA15, SRM20, TDV15, WDC^{+}24, WGZL20, XWC^{+}23a, XSZ^{+}20, YFX^{+}18, YGC15, YSZ23, ZLZH17, ZHL^{+}20]. super-pixel [ZLZH17]. super-resolution [AM06, AAMO16, AYG23, CWW^{+}22, CU20, EH21, FSV07, FDSB22, GB22, HL23, KKP24, LLF18, LWW^{+}21, LZWX21, MYYY17, NFSD13, QGZ^{+}23, SA15, SRM20, TDV15, WGZL20, XWC^{+}23a, XSZ^{+}20, YFX^{+}18, YGC15, YSZ23, ZHL^{+}20]. super-resolved [JC06]. Superclass [WKT22]. Superclass-aware [WKT22]. supercoupling [AKC11]. Superpipelined [DRAB08]. superpixel [CO16]. Superpixels [SHL18, GTP18, JSZY17]. superquadrics [KS04]. superresolution [BR12]. superresolution-inpainting [BR12]. Supervised [HDZR24, LCZ^{+}16, ABN^{+}20, AS23, BIG^{+}23, CBS17, CLZZ21, CJYW23, CLTW23, CLL^{+}14a, CZHT15, CCSS14, DRTT24, DFLS23, DB14, EOPS22, FCOK24, GLG22, HWG21, HD23, HWW23, JLZ23, KHG22, KKKS24, KRG17, KACR^{+}23, LHL^{+}21, LZZ22, LLS^{+}23, LYX^{+}24, LXW^{+}24, Mah16, MDK^{+}24, MPM16, MPT21, MTP21, NWNT17, NF21, OZT19, PKH23, PBD20, PD23, RDA^{+}15, RZZ23, RSBA23, RKL^{+}18, SY20, SCvW11, TLWT12, UU18, VPL23, WHM^{+}09, WZW17, WK21, YK24, YXW^{+}24, ZKSV18, ZZ20, ZZSD21, ZYZ24, ZLY^{+}20, ZJL23, ZZG^{+}24]. supervision [BPCT22, DZZ^{+}23, EDJ^{+}20, FKS10, HH19, LPC^{+}20, SG17, VGSMN16]. supervisor [BIMD23]. Support [GK98, CMBP09, CEO18, HGR^{+}13, HBG13, SB13, VJ17]. supporting [LLL^{+}15a, OTO06]. SURF [BETV08]. Surface [Ano95d, BSF02, BM97, CLK09, FW97, FKW98, GL98, HB98a, HSIW98, KP97, KPH02, LSB^{+}00, Lhu18, LLL^{+}14, LLY^{+}18, LM99b, Mil99, OG98, OD99, OD01, QL96, SA96, SL96, SF97, VB98, WH01, WH00, YA12, ZM96, BI11, BSRV17, BBH14, CHSV08, CHZ^{+}13, GBHS06, HUF05, LÁB15, LY13, MPST08, MMA06, MC20, MB05, MB95, PMW05, PBW14, PZV13, RDT^{+}19, SY10, STD14, SKVS13, TN05, TN08, UK12b, WPS03, WXZG18, WF05, XOF05, YW07, ZZJS18]. Surface-Based [HSIW98, OG98]. Surfaces [Ano95e, FAB97, FL96, LKK00, NFSK97, Sau99, WH96, AZP14, BGK95, Eva06, KS03, LC11, LYA13, Mil09, MBMC11, OSY18, PJW11, PK05, SAK15, TG95c]. Surfaces-From [Ano95e]. surfel [CPP^{+}11]. surgery [ASFP03]. surgical [ASFP03]. Surround [LCT09, EK12]. surveillance [BZS16, BZ14, CPC08, CHH09, CTWH15, DETE17, GCS23, GMW12, GWT09, HHM^{+}16, LLS21b, MFB11, MW13, NS16, OBTMT15, RAP16, RCTV12, SJH17, TYDH18, TMB12, VD10, WMBY12, YCKA10, Jon08]. Survey [CF01, CH17, CL97, Doe98, Gav99, HL01, JB23, JT17, LYBT17, May99, MG01, MEDT96, MPK24, NJ95, AYD^{+}18, AKE23, BS19, BCF06, BHF08, CCFC13, CTH20, CR18, CMG16, DFS08, EPH^{+}21, FPNK22, FBK15, GB10, GZ19, HS06, JS07, KU19, KLY21, LB14, MEM17, MHK06, NNN^{+}22, RFMF21, SR23, SP19, TZLT21, TA13, WKP13, WLO^{+}18, WZ23, WRB11, WTW^{+}17, ZZZ15, ZFG08]. Surveying [EDX16]. Suspension [EK14]. suspicious [WMBY12]. svd [YFDA17, SZW^{+}21, ZZP12]. svd-updating [YFDA17]. SVM [MJ17]. SVMs [AZ15, BRA^{+}10]. SVP [FB05]. swapping [ZZC^{+}23]. swarms [GA13]. Swimming [TML00]. Swin [JZL^{+}23]. Swin-ResUNet [JZL^{+}23]. switching [DM24, KDV16]. Sylvester [CS10]. Symbolic [Ano95e, KDRC98, KP00]. Symmetric [SK02, GXC23, LA11, RM06]. symmetrical [YJA96]. Symmetries [Big97, ST96]. Symmetry [BCM13, Rob96b, TS00b, VMUO95, YHR^{+}05, ZW97, BCLNG18, HZK19, AGB^{+}15]. Symmetry-based [YHR^{+}05]. Symmetry-driven [BCM13]. symphonic [BLH16]. Synchronization [Boy04, ARFF18, NL17, TR09]. synergies [PT08]. Synergistic [CUAT13, dMFU10, BEK18, MNMK16]. synergy [WLZM20]. synonyms [GSS12]. synopsis [BS19]. syntactic [IJDAB13]. Synthesis [Boo97, Nis97, AYD^{+}18, CCD11, DWL^{+}24, HKS06, JB15, LHG^{+}23, PSB^{+}24, RB19, SHK11, UBEP09, YLLG18]. synthesize [BRET19]. synthesizing [LPR^{+}03]. Synthetic [BCC^{+}18, SDK^{+}24, AGL23, BSH13, BG18, DM12, DLV15, LRFF24, RLF15, SV14]. System [BKMSR98, BS99a, CN95, CJC^{+}98, FCM20, Lee02, MFJ95, ME98b, SBK^{+}99, THT^{+}98, YYL96, ABI^{+}04, AZSVK05, ALY^{+}22, ACC^{+}16, BMJF^{+}17, CEA16, CJL06, DLS^{+}09, DR04, ESS10, FFY^{+}04, FY06, FLCdA06, GSPL10, GBVDC18, HSKH07, HWW06, ILRB04, KGFP10, LM16, Lhu08, LNS14, MSG10, MTC^{+}14, MML^{+}16b, NKB11, PFGG09, RGA10, TKDN16, ÜB05, VD10, VZP^{+}09, YH19, BCDH10, FRNS05, TG95a]. Systematic [MSM17, SDK^{+}24, LS12]. Systems [BBC00, CL97, EA95, KS95, LH99, SC00a, Bar06, BHSD^{+}13, BRP04, CYP^{+}10, GF15, GA09, GYF18, HD07, HZW^{+}10, KFN15, KGM19, LFMP13, OBTMT15, OH05, OVJ^{+}21, PLYW21, PA13, PV14, RPBK22, SBB10, Tho10, TA11, WMBY12, YCA^{+}10]. Systolic [Nic95].

TAB [MYV19]. Table [GK95, CXFS06]. tablets [JRBD^{+}15]. tag [BBSD15, LDH^{+}14, WZX^{+}14, ZWY14]. Tag-Saliency [ZWY14]. Tagging [CWH^{+}13, KKKS24, LLTL14]. tailored [JPN^{+}22]. Take [Lau97, WASF14, WZX^{+}24]. Taking [EMMV19, FL96]. tampering [KLL^{+}11]. Tangential [LKK00]. Target [IKST05, MYC09, TLH22, BG16, BVCP21, CSGM^{+}24, CSLX16, GFY^{+}14, JBC08, KW12, LSL^{+}18, PMC13, UM05, VSP06, YCKA10, ZZRC15]. Target-aware [TLH22]. targets [BYR17, BYK^{+}18, KPPK09, MC09a, PBT14]. Task [DC00b, GZJ05, SGB01, TAC21, WCZ^{+}20, XSL^{+}24, ZSDK19, BIG^{+}23, BRA^{+}10, BSMK13, ES06, FCM20, HL13, HML15, JRAJ17, RGA10, SvNW23]. task-driven [RGA10]. Task-Specific [DC00b, ES06]. Tasks [KR99, AGL23, CP21, CCF17, MdOBA19, SVA^{+}22, WTZ24, WZX^{+}24]. taxonomy [TESY15]. Taylor [BKK11, TYH^{+}21]. TBS [PT08]. TC [EHG^{+}10]. TC-12 [EHG^{+}10]. TCLR [DGRS22]. teach [LHL^{+}21]. Teacher [BIMD23, EKY08, LLS^{+}23, LDCX24]. Teacher-directed [EKY08]. teacher-student [LDCX24]. team [HKHE14, PD17, PKK^{+}09, WASF14]. TECD Attention [LD24]. Technical [OMLL98]. Technique [Ano01m, BL01, Luc01, OD97, PLL00, CCL04, DM12, HBL^{+}17, KA12, MWF07, RC03, YW07]. Techniques [Ano98d, BY98, BS00b, CF01, MAP99, MNSK98, AS09, Bre03, FK09, HBG13, JB23, JM09b, MPK24, MGPF08, MM05, OTO06, PSE^{+}11, PR03, SM13b, TA13]. technologies [LMT^{+}17]. technology [CSV^{+}16, CMCM16, RMN^{+}17]. Telepresence [OYTY98]. tell [Che24]. tells [YSL^{+}14]. Template [CYES00, THT^{+}98, BBH14, FN14, SBPF17, UBEP09, AW09]. template-based [BBH14]. Templates [DJG01, LSB^{+}00, SL99, DLF06, GRGB^{+}13, RCT14]. Temporal [BZS16, CA97, DGRS22, KHH^{+}22, MIUS16, STO17, SC15, SA04, UFF06, WY21, WLL^{+}22a, XYQE24, XZW^{+}23, YJ16, AAL22, CLL^{+}21, CHMG12, CKPV21, CWLJ13, CSG^{+}03, DPCA15, DLF06, FXWW17, HSBS16, HDF12, KYYC14, LCSL07, LTY^{+}15, LLG^{+}24, LXFM16, MTV17, MYV19, NNS^{+}18, NDO09, PCM21, QZY^{+}24, RCLS19, RL13, SM22, SCMP14, ŠSJ^{+}20, SVF^{+}21, TKL21, TBC^{+}21, WZT13, WX16, WYW^{+}22, WXZ24, XYW11, XWLY23, YFX^{+}18, YCZ^{+}23, CGHTK16]. Temporally [MYV19]. tennis [DGG08, RMN^{+}17, YJC^{+}09]. Tensor [AG00, GGH^{+}24, KKRK23, KHR^{+}16, LLC11, Sah05, XSD12, GYTL09, LBNS09, MGPJ11, Nor09, PG13, RPG12, YYZL19, YGC15, ZXC^{+}20]. Tensor-based [LLC11]. Term [NB20, CRCM16, MBCJ17, PA10a, SPK23, TKL21, WHL^{+}21]. Terminator [UZC97]. Terms [Kis96b]. ternary [WYX^{+}16, kCE^{+}18]. terrain [LPZ08, OMW^{+}07]. terrestrial [RTM^{+}17]. Test [LM96]. tested [FFFP07]. Testing [RH06, EK14]. tests [WBS14]. Text [BKMSR98, DWL^{+}24, DV98, Hob00, WLXL24, YLLG18, YT13, CSV^{+}16, DZZ^{+}23, HLL^{+}23, LZZ^{+}21, LZL^{+}17, MTG07, MTC^{+}14, MAJ16, PV14, SBN^{+}24, TESY15, TDZ^{+}20, CLFH22]. text-based [PV14]. text-to-image [LZZ^{+}21]. text-to-speech [CSV^{+}16]. texton [SPK14, ZZL13]. texton-based [SPK14]. textons [XHJF12]. texts [GF15, VJ17]. Textual [SLST99, LDC^{+}13]. Textural [AM00, CE17]. Texture [BUD19, CSDNR17, GSP01, GPK99, LSD^{+}07, LD24, MCCRAC20, PPT06, PB99, RPTB01, SA02, SM99, SC98, VGR16, WH01, AYD^{+}18, AMCB20, ASVO12, CE17, CCD11, DL10, FLS^{+}14, GFL^{+}11, GB13, eGZW07, HAT^{+}15, HOH^{+}07, HG11, HBL^{+}11, KORC10, LF08, LGD16, LPVM13, MSW15, MGPP11, Mig12, NNBN20, Pen15, Pun03, QAB^{+}11, QSX17, STD14, SG11, SF07, TT16, VBS^{+}04, WX16, XTZ^{+}18, XQZL23, XHJF12, YLLG18, ZZJS18, ZZL13, ZLFH23, kCE^{+}18]. texture-aware [XTZ^{+}18]. texture-based [MGPP11]. Texture-driven [BUD19]. Texture-enhanced [LD24]. texture-less [Pen15]. textured [JRBD^{+}15, TD19, WBS14]. texturing [BI10]. TFUT [XSL^{+}24]. TGV [GY19]. Their [NSK^{+}97, SC00b, CTCG95, CKS^{+}05, DLMC16, FLB06, GCFMT12, KEG15, OVJ^{+}21, SSM06]. theorem [BFR13]. theorems [She16]. theoretic [BEGB13, SPC^{+}15, VMC^{+}16, WSSS13]. Theory [HKA13, Mok97, SUO00, SU01b, SWG02, WKI^{+}16, AGB^{+}15, AC07, BBK15, DB03, KLBP11, NRJ11, NGR24, XP11, HMEB07, KGK10, MUS06]. There [Ver97, AQ09]. thermal [DS07, HOH^{+}07, MHAF13, SSN03, TMB12, TB13, YCH07]. thermal-visible [TMB12, TB13]. Thermophysical [MNSK98]. thickness [Coe12]. thigh [TLY^{+}16]. thighbone [GYW^{+}22]. Thin [AMMV99, MAM97, TDK10]. Thinning [Che98, CCS95, MS96a, MW00, MWL99, Pud98]. Thinnings [BJ96]. Thoracic [LSB^{+}00, ML13]. thoroughly [PK05]. Threat [KR99]. Three [Bor96, Jos99, LSCK15, LWZP17, MNHO00, MCPB99, OD01, SF95, TK97, WD96, ZM96, AMCB20, CH17, HQN05, KON^{+}17, LB08, PJW11, SOL16, SB05, WXWC18, WXC20, WZFL23]. Three-Class [MCPB99]. Three-Dimensional [MNHO00, SF95, TK97, WD96, ZM96, LSCK15, HQN05, LB08, PJW11, SOL16, SB05]. Three-layer [LWZP17]. Three-Light-Source [OD01]. three-stream [WZFL23]. Thresholding [Ros02, WCZ02, GFL^{+}11, HDS08]. THUMOS [IZJ^{+}17]. Tighter [Zha97]. Tilings [Mil99]. Tilt [CC00, DDLP10, SPC^{+}15, SP06]. Time [BEPW00, CBM01, HT98, LB98, LSKK10, LHHC98, OYTY98, SKOS95, SLK15, WZWT99, ZXK02, AMNCM16, AM04, BT05, BCMCB09, BDS12, BHMB10, BPLT15, CGH08, CEA16, CCL04, CKPV21, CKL18, CSK22, DLS^{+}09, DDWZ12, DZJB14, FFM05, FTT15, GK23, Gon09, GTMR23, HRHZ17, HHAE14, HEPH15, HWZ^{+}23, HZW^{+}10, JRS21, JSRS08, DFP^{+}13, LLS21a, LC14, LÁB15, MZB^{+}10, MWTN04, MFS^{+}07, MHL14, MTAA11, NKL24, Nic95, Pen15, PBI16, PGGM04, QZY^{+}24, RZH17, RAC^{+}13, RL13, SM12, STC^{+}16, SFK18, SGH07, SIT07, SHS03, TKV16, UM05, UWH17, WX16, WHL^{+}21, WWLV11, YWZ11, ZJ05, Ziv10, LBK10]. Time-of-Flight [LSKK10, SLK15, BHMB10, HHAE14, HEPH15, LBK10]. Time-Varying [CBM01, SKOS95]. times [MOT17]. timescale [Sav23, SY23]. tissue [CFYU12, DCS05, SRP10]. TMF [WY21]. TOF [NB10, GPC^{+}10]. TOF-scans [NB10]. together [CLA^{+}17]. tokens [WZWL24]. tolerant [MRH19, ZHS^{+}24]. tomographic [VNNB14]. tomography [BPBS13, BTB14, RDT^{+}19, RBdDS14]. tone [ABK^{+}18, BEK18, LLNC20, LJZ18]. tone-mapped [LLNC20]. tone-mapping [ABK^{+}18]. tool [BCNS15, DAM12]. toolbox [RPBK22]. tools [RLMK15]. tooth [MST16]. Toothbrush [MST16]. Top [MSP^{+}18, BYJG23, CFM^{+}23, HLB17, MAJ16, ZWY14]. Top-down [MSP^{+}18, BYJG23, HLB17, KMN11, MAJ16, ZWY14]. Topic [NHTG15]. topics [TGM^{+}17]. topographic [WY07]. Topological [ACF00, ASS97, AC07, CDIF14, Cou13, DBF04, Dam08, Eva06, GL95, GJMO14, ABD11, Bar18, GFW13, WD14, ZZJS18]. Topologies [EL03]. Topology [Bre01, DM01, NS96, ZSCP08, CKP^{+}19, FFL14, Lhu18, LLG^{+}24, Loh10, MdOBA19, SC96]. Torsion [Mok97]. Torsion-Based [Mok97]. torus [LNS14]. Total [Kis96b, MLJC20]. totally [Ang07]. touch [WHC14]. TouchCut [WHC14]. tourist [PHY^{+}11]. tower [XP11]. traced [NRJ11]. traces [GYCS21]. tracing [CCL04, MW13, WPK09]. Track [MW13, AVBK10, PT08, BVWS21]. Tracker [KSS97, TS01, AM04, MiMO^{+}16, SKLM22, SGH07, VMN16]. trackers [DYM14, TMN06]. Tracking [BL98b, DLC14, DF01, Dem96, DJG01, FLB06, HFKN97, IP98, KS95, KB95b, KH13, KDV16, LCP13, LRD99, MJ11, MJD^{+}00, MZL^{+}16, PV13, Pet99, PF01, QL96, RAH97, ROJX09, TPR^{+}00, WN99, WS06, ADR16, Ano06h, ABVC16, AVC19, BAPXH16, BYR17, BSM10, BW11, BBH^{+}12, BCMCB09, BL09, BY12, BBK14, BB15b, BG16, BKMV07, BYK^{+}18, CZL^{+}24, CGH08, CSGM^{+}24, CKM11, CYP^{+}10, CSLX16, CPT07, CKC14, CKL18, CÇ15, CSK22, CZZS07, CJWW22, DZL07, DBZ07, DD11a, DZJB14, DG11, DPT07, DZLH17, DM24, EDB12, FXWW17, FN14, GKK05, GLOC10, GB08, GRB13, GFY^{+}14, GCFMT12, GTMR23, GCT^{+}14, Gwa17, HD09, HYJ11, HP05, HH07, HGR^{+}13, HUF05, HML15, HW07, HDF12, HJZ16, HH12, IKST05, JVD^{+}20, JSRS08, JBR08, JWDF05, JBC08, JY14, JB15, JHV19, KBN12, KNL15, KV06, KG14, KSR^{+}12, KGFP10, KLK14]. tracking [KW12, KPPK09, KT07, KTV17, DFP^{+}13, LHYK05, LST13, LLR10, LÁB15, LWZC14, LLP16, LG17, LSL^{+}18, LWLC22, LJC^{+}23, LJC^{+}24, LG14, LSTF12, LA05, LN10, LLG^{+}23, MYC09, ML15, MML^{+}16a, MC09a, MEM17, MZB^{+}10, MEYD11, MHSP10, MHMO09, MLH13, MBCJ17, MM05, MdRNM15, NAS^{+}17, NHY10, NKB11, NLM05, OMBH06, PA10a, PD05, PA06, PMC13, PYS03, PTK24, QWHW20, RMD08, RRR11, RB16, RCTV12, SPC^{+}15, SC15, STC^{+}16, SFK18, SA04, SHE17, SM24, TTXT21, TLH22, TWQW23, TFD07, TKV16, TMB12, TM07, TP05, TTH07, UM05, UO16, UFF06, VSP06, WASF14, WLL^{+}22a, WG23, WDC^{+}20, WDB12, WB16, WPZ^{+}18, YWZ11, YQL^{+}23, YZL16, YJ16, YNCO11, YJC^{+}09, ZN08, ZZRC15, ZT09, ZWZ^{+}16, ZYS09, ZJ05, ZWL16, ZCK09]. tracklet [HHG^{+}20]. tracklets [ADR16, SM17]. Tractable [SP23b]. Trade [LHH^{+}98]. Trade-offs [LHH^{+}98]. trademarks [PA10b]. Traffic [HMEB07, SJH17, HLL^{+}23, KBKS18, MAG^{+}16]. trained [DYM14]. training [AGL23, BCC^{+}18, BCC16, CHH09, CSZ^{+}15, CTCG95, FSL24, FFFP07, GKGM20, JRH24, KKSC23, LKZ20, LLWX24, LLWZ21, MRH19, RLF15, SLK23, SS21, ZS19]. trajectories [AAASC11, CHP^{+}11, KBN12, OCVV04, TSD17, WCF10]. Trajectory [LB08, BCC^{+}21, BSZ^{+}21, PKK^{+}09, SY23, YGC13, YHS^{+}20]. trajectory-based [PKK^{+}09]. Trans [MWL^{+}24]. transcription [GRMH19]. transcripts [KRG17]. transductive [CEO18, DWL19, WW16]. Transfer [ZWB^{+}22, AZ15, BSH22, BAM16, GDM14, GD19, GKH^{+}21, HSHA20, HKL^{+}24, HLKK19, JPN^{+}22, KOC17, PKD07, SYL^{+}24, TFL^{+}09, TL16, WZC^{+}21, YLLG18, ZSL^{+}24]. Transferable [LCS^{+}21]. transferring [LXW^{+}17]. Transform [AM00, BM00, BM02, Che24, DGH98, DG01, KB00, LHKC97, LH99, MGK00, MNHO00, PKP97, SWG02, SJ01, SK98, TV99, TS00a, AKC11, ABLL19, ÇÖD08, CT10, CT12, CS04, CL95, GY19, Gre04, Hu11, IAP^{+}11, KC22, LY05, LWH^{+}23, NSEA13, SA04, SYK96, TWS06, YLX^{+}18, ZS11, dSM14, MSF^{+}12, PCC13, Sha06]. Transformation [CM99b, Dav97, ER96, GLR^{+}99, LB98, CGR13, CLO17, CS20, DDWZ12, GM19, GCD^{+}18, HKWC14, HKL^{+}24, IH15, KKSH23, LRF^{+}17, OBH04, OH04, PHHL23, PZC17, RK11, SC96, SOL14, SOL16, SG11, SW04, SA15, SY11]. Transformations [Ano01m, Big97, Egg98, Kis96a, Luc01, SC99, BDHM09, DL05, GHML17, NKPT13, NESP10, RSY22]. Transformed [RSY22]. Transformer [CSD^{+}24, DWC^{+}24, PTK24, SPRS23, ZFW^{+}24, BB24, CWS^{+}24, FCOK24, HBZ^{+}24, LML^{+}23, LXW^{+}24, PCM21, SFK24, TLFM23, VPP^{+}23, XSL^{+}24, ZSL^{+}24]. Transformer-based [PTK24, SPRS23, LML^{+}23]. Transformers [LXS^{+}23, WZWL24, JYY^{+}24, WZFL23]. Transforming [ZL01, CLK09]. Transforms [Bor96, Ols99, SB02, Nis96, SB05]. Transition [YW99]. transitions [UK12a]. transitive [PS22]. Translated [MSW96]. Translating [DT96b]. Translation [WC99, ANHGS17, BDVK10, CLTW23, PWL^{+}23, TBFJ15, WGGHvdW21, ZYZ24]. Translational [HJ12]. translations [CSW^{+}24, PD23]. transmission [YZX^{+}20]. transparent [KS12, MWL^{+}24, XMN^{+}15]. Transport [DFSC20, HHM^{+}16].

TransRPN [LCS^{+}21]. trapping [CPO16]. TRASMIL [YGC13]. travelogues [PHY^{+}11]. traversal [PYGGLNG17]. TRECVid [SOD10]. Tree [WW97, ÇÖD08, CT10, CTM^{+}13, CCL^{+}17, Hu11, HQW^{+}12, JLD13, LZWP03, Pha17, RC13, TN07]. tree-based [JLD13]. tree-structure [TN07]. tree-structured [CCL^{+}17]. Trees [HdVL99, Jon99, LHKC97, Mun95, MNL^{+}17, MU11, QT10, VBVB19]. Tri [XS04]. Tri-view [XS04]. triangles [Žun03]. triangular [MSR07, WTBdB15]. Triangulated [KPH02]. Triangulation [HS97, SL96, Tan95, WZWH16, BS05, CH11, GSGJ22, Nor09, WCZ23]. Triangulations [WCH98]. Tribute [Kak97]. Trilinear [Zha97]. Trimap [JYX^{+}23]. Trimap-guided [JYX^{+}23]. Triplanar [KSL^{+}20]. Triple [RMS^{+}19]. Triplet [QV98, BP05, BRPC17, GL24]. truly [CU10b]. truth [Cre08, SYPK13]. truthing [RLMK15]. Tubular [KMA^{+}00]. Tumor [RAC^{+}13, LWLT17, ZRL^{+}11]. tuning [PHHL23]. tunnel [RCTV12]. turn [CXFS06]. turn-table [CXFS06]. Tutor [FKS10]. Tutor-based [FKS10]. TV [ACDB12]. Twin [AH24]. Twin-SegNet [AH24]. Two [AH08, CDH99, DM12, Egg98, Jos99, ML15, QWHW20, SP97b, SA95, WHL^{+}20, WLMG08, XZW^{+}23, ACAAC^{+}08, BI10, BYN^{+}04, DBF04, GHZ^{+}13, GSGJ22, Got08, JM09b, KHG22, KSY15, KNO^{+}09, LYKY19, MMP15, Ros08, Sha11, SW04, SCCP05, WZ08, WCF10, YGH11]. two-component [Ros08]. Two-dimensional [AH08, DBF04, GHZ^{+}13, Got08]. two-level [KHG22]. two-orthogonal [YGH11]. Two-Stage [SP97b, WLMG08, XZW^{+}23, KSY15]. two-step [BYN^{+}04]. Two-stream [WHL^{+}20]. two-streamed [LYKY19]. two-view [GSGJ22, MMP15]. Type [NCDG21, GY19]. Type-specific [NCDG21]. Types [RWV95, SKA23]. typical [MB95]. typology [COV^{+}22].

U [HDZR24]. U-PVNet [HDZR24]. UA [WDC^{+}20]. UA-DETRAC [WDC^{+}20]. UG [BVWS21]. Ultimate [AHM17]. ultrasound [MAK^{+}17, MJPS16, ZIT^{+}13]. Unbiased [Ste13, GLZF23, LZ24]. Uncalibrated [BK01, Tay00, VF96, SCEvdH14, TGFF15]. Uncertain [KN99, NHZ^{+}22, PS05]. uncertainties [WR08]. Uncertainty [CZZF97, GOF^{+}15, Shi99, ZFG^{+}22, BÁRM23, CP04, CC03, DD11a, JSWL24, KT08, KTV17, KN11, SS11, TNO24, TM07, VNNB14, ZLY^{+}20]. Uncertainty-aware [ZFG^{+}22]. unconstrained [BVWS21, DCH12, NKB11, PA10b]. Under-display [ZLL^{+}24]. Underst [MFSB23b]. Understand [MBMC11]. Understanding [AYB^{+}18, AK11, Ano06h, BPQ15, BB15a, Bra97, CGL98, CTM^{+}13, CBB95, CL97, DC00b, GMW12, HF01, KB98, MGLB17, OBH04, PZ09, PT08, TSD17, ZT98, BHF08, CFM^{+}23, HUI16, HFR06, RFF23, SCC17, SPW15, VCLS19, WKP13, WZ23, LLE^{+}09, BPQ15]. Underwater [CFM02, ECC18, GSV00, MCPB00, MT00, NK00, SWYP00, MN06]. Unified [BYK^{+}18, CWH^{+}13, RJ00, HYW^{+}24, JLD13, LLTL14, LYW^{+}24, LH03, MIP16, YZY11, ZLZH17]. uniform [LLWX24, MC20, SAC09, TLCH05]. Unifying [SLST99, SVF^{+}21, Bar06]. Unique [STD14, RAC^{+}13, XGT^{+}22]. Uniqueness [CM99a, OD01, DLV15]. Unit [HB98b, LHZY19]. Unitary [LNS14]. units [TYDH18, OS19]. universal [WSFTK18]. Unknown [FW97, OD99, BBK14, GS06, LC14, SSS13]. unlabeled [CHH09, WZQ^{+}23]. Unmanned [NK00]. unordered [MAL10]. Unorganized [ZOMK00, LLL^{+}14, LLY^{+}18]. Unpaired [ZZHZ23]. unprepared [LA05].

Unscented [DG11, IH15]. unseen [KKRK23, RG10]. Unstructured [BCA98, CPS10, PLB16, RAP16]. Unsupervised [BP05, BCC16, BCM06, CHH09, CT10, DTL17, DAL^{+}22, GMF14, LBP23, LTL^{+}23, MGPP11, MHL14, NHSC09, PB99, RM03, RCLS19, SZS17, SY23, TS24, TVC09, TA11, WLSO23, WHY^{+}23, YWMS08, YSZ23, BPCT22, CCSS14, DLMC16, FWZ^{+}24, FDC^{+}19, GCEC07, HDL^{+}20, LLLW23, PC15, SYL^{+}24, SPW15, XW16, ZFG08, ZYZ24, PD23]. untextured [ÁB13]. UP-SR [AAMO16]. up/top [KMN11]. Updating [MS96b, YFDA17]. upsampler [FLL^{+}23]. upsampling [AAMO16, HDZR24, XJK12]. upward [XSL^{+}24]. Urban [BM99, FRL^{+}98, FMR01, HB98a, RDT^{+}19, SPQ^{+}17, BSRV17, CM12, GDCM17, LS12, MTC^{+}14, ZA22]. Usage [NSK^{+}97]. Use [BBC00, CN95, EFF98, GPK99, RWV95, SGB01, CU11, CCSS14, Loh10, NF21, REF15, Ano95e]. Useful [GHMQ97, TDV15]. User [CYES00, IZKB12, KDV12, PJW11, PHY^{+}11, RTM^{+}17, YWZ11]. user-assisted [PJW11]. user-contributed [IZKB12]. user-generated [PHY^{+}11]. users [CNO^{+}16]. Using [APV99, Ant98, AMMV99, BKP10, BCDH10, BH99, BKD01, COW98, CM95, CS98, Che98, CL00, CM99b, DT96a, DT96b, DZQ24, Dav97, DUC97, DJG01, FBF08, FD99, FKL^{+}98, GKBW14, GBB98, GJP96, GSK02, HB98a, HCHD01, HR99, HB98b, Hob00, HN95, HLF^{+}97, Jon99, Jur99, KP97, KSI98, KHB01, LVW97, LB00, LL97a, LSHT02, LL97b, LZ97b, LF98, MBKB02, MGK00, MS97b, MK01, MB95, Mur95, NG98b, NMP97, NL96, Nis95, OJRT08, PKP97, PA00, PC99, RM98, SYF99, SB95, SC00a, SB98b, SP97a, SPK^{+}02, SHKP98, SL99, SLL01, SF97, Spe97, SYPK13, SB02, SM97, SC98, TML00, Tsa96, ÜE01, VB98, WW97, WZWT99, YKA01, YC98, ZW97, ZOMK00, AJ23, ARC14, AYB^{+}18, AM06, ABN^{+}20, ADC19, AS09, ADGB16, AW09, AC07, ABEN09, ALK^{+}09, AC09a]. using [AC09b, AZP14, AT17, AMGG^{+}16, ASCF13, ASF14, AM15, ABK16, ARARCE11, BW11, BKPS15, BCMR16, BMJF^{+}17, BS05, BRA^{+}10, BZS08, BP05, BBCF20, BL09, BCC16, BWL04, BBK14, BB15b, BMvT^{+}19, BPSV16, BRPC17, BF10, CGH08, CHP^{+}11, CLZY15, CFCP11, CMBP09, CH06, CKM11, ÇÖD08, CT10, CT12, CEO18, CGR13, CCL04, CPP^{+}11, CL17, CE17, CLO17, CAO^{+}23, CFM^{+}13, CC03, Cre08, CKS^{+}05, DK13, DZL07, DT09, DBZ07, DKG22, DM12, DGC12, DS07, DWV19, DLF06, DCS05, Dre96, DZLH17, DH19, EKY08, ESS10, EOPS22, EF14, EH21, ET15, Eva06, FPC^{+}08, FM22, FB05, FN14, FKS10, FK09, FCOK24, GHZ^{+}13, GS06, GLM17, GB22, GBHS06, GBB^{+}18, GKPS15, GTP18, Goh08, GA09, GDIIHK11, GFW13, GYF18, GPC^{+}10, GCT^{+}14, HLB17, HTNN18, HKHE14, HASS10, HWG21, HD23]. using [HWZ16, HSK23, HY11, HPvB^{+}10, HMB17, HBL^{+}17, HMF10, HGP15, Hu11, HQW^{+}12, HC13c, HKK08, IAP^{+}11, JKM07, JSWL24, JHA17, JWG04, JBC08, JYTK11, JBWK11, JY14, JZWD16, JSZY17, JGM20, JHV19, JC06, JPP^{+}14, JHK24, KL07, KK15, KS03, KNL15, KIS17, Kim04, KLL^{+}11, KLO20, KKSC23, KKCK23, KH23, KKSH23, KLKF20, KM03, KS04, KACR^{+}23, KSY15, KMN11, KNO^{+}09, KON^{+}17, KRS14, LMP^{+}19, LEE^{+}18, LRW08, DFP^{+}13, LHYK05, LCP13, LÁB15, LB19, LY06, Lhu08, LCZ09, LWZC14, LSCK15, LWLT17, LKZ20, LQQS21, LCS^{+}21, LZZ22, LSGY24, LXFM16, LB10, LYG07, LHJ^{+}09, Liu10, LLC12, LDC^{+}13, LmCT16, LLWZ21, LZS16, LPVM13, LAL^{+}10, LDL^{+}19, LCG^{+}24, LT97, LYA13, MGW10, ML13, MSI10, MDFS11b, MLB^{+}18, Mah16, MK18, MDM^{+}21, MdBJG15, MZC^{+}05, MSF^{+}12, MM06, MCF10, MJPS16, MdRNM15, MP20, MC22, MTR^{+}23, NB20, NHH14, NN18, NNT11, NGR24]. using [ODD96, OZT19, OTAH20, OCVV04, PY08, PZX13, PYWZ17, PRR03, PC05, PLLL03, PW06, PKH23, PA10b, PG13, PKD07, PBW14, PL08, PBG04, PZC17, RB18, RRR11, RCLS19, RB19, ROJX09, RF23, RL13, Ros10, SKLM22, SY10, STO17, SCE04, STC^{+}16, SAS12, SJB20, SCALFG^{+}18, SvdMH15, SBB18, SJST07, SCC17, SZ16, SAC^{+}12, SW04, SZ07, SY20, SKU^{+}09, ST10, SAC09, SCMP14, SKT18, SBMM15, SGH07, SKS11, SRHC13, SM13b, SS21, TLB^{+}15, TLP^{+}17, TLFM23, TYH^{+}21, TS11, TS17, TN07, TB13, TRG^{+}13, TR09, TKL^{+}09, TL15, UJ22, UMH16, UWH17, VBVB19, VBT19, WZ08, WJ07, WHC14, WPSL18, WHJK23, WRB06, WMBY12, WSKH13, WR08, WWJ13b, WYW^{+}22, XSL^{+}23, XYZH11, XAB07, YGH11, YC05, YW16, YSY^{+}18, YSKL24, ZK17, ZRRK18, ZZC^{+}13, ZA22, ZT09, ZYT10, ZS11, ZYS09, ZLHJ18]. using [ZNG^{+}13, dLAH07, dMFU10]. Utility [DTG96]. utilization [WTZ24]. utilizing [KK11].

Validation [SUO00, BY08, SC15]. valued [YZX^{+}17, YG17]. vanishing [ATG15]. variability [Dem05]. Variable [GJH01, KB00, MGW10, SGH07, ZJ05]. Variable-Length [GJH01, SGH07]. variables [BW11, CLCO13]. Variance [Imm96, WH00, CCPK16]. variance-based [CCPK16]. Variant [BL98a, RC97]. variants [HF11, KK15, RH06]. Variation [MIUS16, GHZ^{+}13, HWZ16, MLJC20]. Variation-based [MIUS16]. Variational [BCA16, FWL^{+}20, FKW98, ZOMK00, BAPXH16, CHSV08, FMS17, HW06, LEE^{+}18, LJHH07, MCF10, RPG12, TD19, dP10]. variations [LY06, SKVS13, TLCH05]. Various [RWV95, LCP13, YWZ11, ZJ05]. Varying [BFF97, Bic98, CBM01, Lai00, SKOS95, ABB^{+}23, DL10, OK04, SB96a, WQY^{+}21]. Vascular [WW97]. Vector [APV99, Che98, SYF99, SJ01, WW97, WSSD96, CMBP09, CEO18, FLS^{+}14, JWG04, LSPV04, MWF07, SB13, VJ17, ZLS^{+}13]. Vector-Based [APV99]. Vector-City [SJ01]. Vectorial [ZUS06, MLJC20]. Vectorization [JV97, VRKL13]. Vectorized [CLD96, DL97, LCD97]. vectors [FB16]. Vehicle [KS95, BKP10, KU19, OBTMT15, PT15, RCTV12, SJH17, WZT13]. Vehicles [HFKN97, NK00, SWYP00, JBC08, MFG10, TDWH07]. Velcro [NFSK97]. Velocity [Pet99, LCSL07, SA04]. velocity-adapted [LCSL07]. venation [NHK08]. ventricle [WSKH13, WWJ13b]. Vergence [CTE95, MGMS01, SB96a]. Verification [DLHT99, LVW97, ABEN09, CJL06, DM12, KDSF20, KKSH23, KSY15, MK18, PMR17, RSS07, SAV24, SKSR08, STC14, ZBDP15]. versa [AB18]. versatile [MZB^{+}10]. versus [HHWP03, KŽ12, SLK15]. vertebra [ML13]. vertical [Lhu23]. Vessel [TKL^{+}09, PYWZ17]. via [AAASC11, ANM98, AXJE21, ARFF18, BI11, BIMD23, BZ14, BG16, CFYU12, CZ14, CYD^{+}22, CXYZ24, DFP23, DFLS23, EK12, FWL^{+}20, FWLQ23, FWZ^{+}24, GFL^{+}19, GWT09, GYCS21, GML^{+}21, HDZR24, HJZ16, HWW23, HKL^{+}24, IH15, JPN^{+}22, JCLZ21, JSC23, KSRS16, KHG22, KA08, KM17, KKSC23, KSKB95, KORC10, LPSK23, LRZ^{+}19, LZZ^{+}21, LHG^{+}23, LSQL24, LDCX24, LLWX24, LDH^{+}14, LTL^{+}23, LG24, LYSS12, LCZ^{+}16, LZL^{+}17, LZWX21, LLJ^{+}23, LYSK17, LCG21, MZ20, MMS99, MFSB23b, MFSB23a, MLZRK24, MSW15, NAS^{+}17, PBT14, PZM^{+}21, PD23, PCM21, QLY^{+}17, QDLB17, QCL^{+}23, SP23a, SM21, SZW^{+}21, SHW24, SJSL21, SMD^{+}08, SBH^{+}17, SXZZ24, SWMM22, TPT17, TCM18, TY22, TAK^{+}22, TGSH98, UFK20, UIK22, WW16, WZW17, WZJ^{+}21, WLSO23, WPZ^{+}18, WLFL21, WZCY22, WHS^{+}24, XFP^{+}16, XTZ^{+}18, YWL^{+}20, YWMS08, YLLG18, YS24, YGC15, YFDA17, YSZ23, YG16, ZXC^{+}20, ZSL^{+}16, ZWZZ18, ZWB^{+}22, ZLFH23, ZRKZ^{+}11]. via [ZHL^{+}20]. vice [AB18]. Vide [KFRD^{+}18]. Vide-omics [KFRD^{+}18]. Video [ALK99, ASC17, AWK04, ADDK99, BS19, BPQ15, DCCL99, FWL^{+}20, GSV00, HR99, HNB04, LC09, LCZ^{+}16, MBHRC21, MSF^{+}12, MRdRGC23, MC22, MGLB17, NB20, NK00, OYTY98, PF01, SLS03, SOD10, TBC^{+}21, TY22, TR09, TPR^{+}00, WPZ^{+}16, WZJ^{+}21, WLXL24, XL98, YGJ^{+}20, YFX^{+}18, YYL98, YST21, YW99, ABJ^{+}21, ABI^{+}04, ALK^{+}09, Ano06h, AHDM10, AC09b, AS23, BYR17, BZS16, BVWS21, BIG^{+}23, BZS08, BCNS15, BY12, BZ14, CSD^{+}24, CHH09, CCFC13, CTWH15, CCF17, CJYW23, CPT07, CWL^{+}23, CWLJ13, CC03, CSG^{+}03, CBTC23, DK13, DLMC16, DGRS22, DCH12, DGG08, DRK03, DHP08, ESS10, ECC18, FSL24, FHZX23, FYH11, GK23, GKK05, GYTL09, GS06, GB22, GB17, GMW12, GLMM16, GDM14, GWCO11, HS14, HMC10, HDG^{+}14, HPvB^{+}10, HHM^{+}16, HKL^{+}24, JN09, JYTK11, JYY^{+}24, JB15, KFRD^{+}18, KYYC14, Kim17, KKKS24, KHH^{+}22]. video [KB12, KGU10, LLS21b, LHLZ23, LK03, LHJ^{+}09, LLE^{+}09, LLC11, LXW^{+}17, LWH03, MDK^{+}24, MWTN04, MIUS16, MSSS09, MFP^{+}20, MPK24, MĆK09, MTR^{+}23, NS16, NY14, NLW^{+}17, OBTMT15, OSM16, PKH23, PSYZ13, PBI16, PBD20, PR03, PKZP^{+}24, PGGM04, QWHW20, RYLZ24, RAP16, RR06, RA15, SM12, SC15, SHW24, SM22, SYZ^{+}15, SOJ17, SAV24, SBS04, SM24, SYPK13, SVA^{+}22, SY23, SMHH04, TD04, TZLT21, TLH22, TY05, TPNP15, TYDH18, TMB12, TVC09, USKB10, VD10, WHM^{+}09, WHC14, WHL^{+}21, WZFL23, WGGH24, WLM^{+}14, XG08b, XTZ^{+}18, XZW^{+}23, YSL^{+}14, YNZ^{+}19, YHS^{+}20, YJC^{+}09, ZdCR^{+}24, ZhZFL22, ZKC03, CLFH22]. Video-based [HNB04, DLMC16, ESS10, FHZX23, SVA^{+}22]. video-hermeneutics [GMW12]. video-quality [SAV24]. video-surveillance [GMW12, RAP16]. VideoLSTM [LGG^{+}18]. Videos [GMZ^{+}22, ABC^{+}03, BBSD15, BLH16, BMB^{+}17, CCTCR09, CD10, CPPY21, CZ18, DETE17, DPM14, DPCA15, DM24, FT23, GCS23, GBL08, HKM22, HRC16, IZJ^{+}17, KKRK23, KM17, KT07, LLF18, LSH19, LYSK17, LYA13, MEM17, MCCRAC20, MW13, MBCJ17, NMAL23, NDO09, QLY^{+}17, QCL^{+}23, RSY22, RCLS19, PBPD^{+}17, RL13, RCJ^{+}13, SV14, SS17b, SS21, SAL16, SZGK24, MNR18, TD04, TB13, UFK20, UIK22, WW16, WZQ^{+}23, WXZ24, XYRS17, YG16, ZTGL18]. View [ASCF13, ASF14, EK98, Gui98, HMF10, KHB01, OD02, OYTY98, WCZ23, ZRRK18, ZSDK19, ATC^{+}13, BYR17, BF10, CPP^{+}11, CC11, CH11, CCD11, CPS10, EKY08, FSI21, GXC23, GSGJ22, GFY^{+}14, HJ12, HKS06, HDG^{+}14, HDF12, ITNP12, KIS17, KCM^{+}17, KM03, LSL^{+}18, LDH^{+}14, LZS16, LYSK17, MMP15, MB11, NZH^{+}23, PW23, PLKP23, RM03, RB19, ROGT14, SBB18, SMD^{+}08, TAK09, TWW14, TVC09, WJ07, WZQ^{+}24, XS04, YW16, ZEGEJ15, ZLHJ18, ZKRH04]. view-based [HDF12, TAK09]. View-Dependent [OYTY98]. view-identity [GFY^{+}14]. view-independent [EKY08]. View-invariant [HMF10, ZSDK19, ROGT14]. view-object [ZEGEJ15]. Viewing [CFA98, Chu02]. Viewpoint [BG18, DCTO97, OMBH06, WCZ^{+}07, CM12, DL10, LA11, MTVM04, ODT17, WRB06]. Viewpoints [RWV95]. Views [BGSdVL98, BLP95, CFM02, EFF98, LV96, MFJ95, RFC97, SA95, ACAAC^{+}08, CKLP09, Gol05, GSV05, JSRS08, KV06, MOB14, PT08, RSPD12, SH08, SCCP05]. vignetting [RBA20]. violence [GCS23, RAP16]. Virtual [EK98, Mur95, BEK18, CCD11, HSKH07, YJC^{+}09, ZKRH04, FPDK12]. virtual-endoscopic [HSKH07]. Vis [AK11, BB15a, MFSB23b, MBMC11, PZ09]. visibility [GGP23, Lhu18, LYBT17, LQQS21]. Visible [FHSKP13, GL98, RWV95, CFB05, DS07, FWLQ23, FHZX23, GZL^{+}23, HD07, HASS10, KH23, MZ20, PS12, SSN03, TN08, TMB12, TB13, XGT^{+}22, ZLFH23]. visible-infrared [FHZX23, GZL^{+}23, ZLFH23]. Vision [Ano96a, Ano98d, Ano06h, BPQ15, BL98a, BY98, BS99b, BD02, CFS98, EBN^{+}07, FHSKP13, FKL^{+}16b, FKL^{+}16a, FHP01, GLOC10, GKL^{+}17, HT98, HTEB11, HSH07, HF01, HFR06, IF95, JBC08, KR99, LVW97, Lee02, LRD99, LXS^{+}23, LSHT02, LLE^{+}09, MST00, MG01, MTH^{+}17, MPPG98, MT00, NPM^{+}16, OBH04, PEFM98, Pop07, Ros95, Ros96, Ros97, Ros98, Ros99a, Ros00a, Ros00b, Ros01, SB95, SC00a, TLP^{+}17, TVY^{+}18, Ver97, WKI^{+}16, YYL96, YLM^{+}17, ACP16, AYB^{+}18, ASC17, AK10, AK11, Ano05j, BK15, BPS10, BDVK10, BC10, BBC^{+}07, CKB10, CNO^{+}16, CLA^{+}17, CMCM16, DBZ07, Ham05, HD07, HAM^{+}16, HBH11, HBZ^{+}24, JB23, JNLG15, JZWD16, KU19, KPKH07, KLBP11, KMT11, LBC^{+}21, LBK10, LMT^{+}17, LXW^{+}17, MP09a, MNMK16, MFS^{+}07, MFG10, MHK06, PZ08, PZ09, PL07, PS15, Rei16, RFMF21, SGS^{+}10]. vision [Sah05, SBB10, SBD22, SKS11, SST06, SVA^{+}22, SIT07, SFWG08, TCB^{+}08, TGM^{+}17, Tho10, UM05, VCDS^{+}17, VAWW10, VZP^{+}09, WWH07, WZ08, WKP13, WZ23, WRB11, YHS95, ZKRH04, Ano95a, STLH08, WZWL24]. vision-and-language [LBC^{+}21]. Vision-Based [HF01, KR99, MG01, EBN^{+}07, HSH07, HFR06, NPM^{+}16, Pop07, CMCM16, KU19, MHK06, PS15, WRB11]. Vision-language [TLP^{+}17]. Visual [Åst97, Ano98d, Ano15o, BY98, Bra97, Col97, CPO16, DAZ^{+}17, DM24, FCM20, Gav99, GSS12, GSV00, GAD01, HOH^{+}07, JN09, JGM20, KK17, KNL15, KSG^{+}19, KRK11, KR99, LHYK05, LWZC14, LVS20, MZL^{+}16, Neg12, NJ95, OMW^{+}07, OS19, PBT14, PRW97a, PRW97b, RB18, RJ00, SVS97, SJB20, SLST99, ST10, SM24, Sup02, TW98, TY01, WS08, WL15, WTW^{+}17, YH19, YR06, ZSY^{+}19, AXJE21, ATC^{+}13, BBH^{+}12, BBHF10, BL08, BF05, BJS14, CGD^{+}23, CSV^{+}16, CGR13, CYNO11, CKL18, CÇ15, DLS^{+}09, DDLP10, DD11a, EMMV19, FPNK22, FMGA^{+}12, FFFP07, FAB12, FKS10, FLHK08, GLMM16, GCPF08, GBL08, HD09, HYJ11, HH05, HWW06, ILRB04, JLZ23, JOvW^{+}05, KD10, KBMD15, KLK14, KHA^{+}05, KYM13, KTV17, LLP16, LD24, LDC^{+}13, LCL^{+}14, LSTARMB11, LN10, LLG^{+}23, LCG21, ML15, MPF07, MMLC23, MdBJG15, MAG^{+}16]. visual [MLZRK24, MHL14, MSP^{+}18, NT10, NHY10, PY08, PWWQ16, PL10, RSY22, REF15, RMS^{+}19, STHBH18, SOK16, SJ15b, SCC^{+}22, SFWG08, TSL14, THL13, TMS20, TESY15, TLMT^{+}05, TTH07, VO24, WRKP05, WZ04, WSY^{+}16, WTZ24, XZX^{+}21, YXW^{+}24, YYH^{+}23, YSX^{+}19, vGSV^{+}10, BCDH10, CAO^{+}23, Jon08, NHTG15]. visual-context-aware [PL10]. visual-object-based [SFWG08]. visual-semantic [LCG21]. Visualization [CC00, ACDB12, CBT^{+}04, CG04, HKHE14, Lhu23, MWTN04]. visualizing [TN05]. visually [CNO^{+}16, COV^{+}22, LM16]. ViT [HSK23]. ViT- [HSK23]. vocabularies [HS14]. vocabulary [KFN15, LSTARMB11]. Volume [Ano95b, Ano95c, Ano96b, Ano96c, Ano97b, Ano97c, Ano97d, Ano97e, Ano98a, Ano98b, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01c, Ano01d, Ano01e, Ano01f, Ano01m, Ano02a, Ano02b, Ano02c, Ano02d, Ano03n, Ano03o, Ano03p, Ano03q, Ano04k, Ano04l, Ano04m, Ano04n, Ano05k, Ano05l, Ano05m, Ano05n, Ano06j, Ano06k, Ano06l, Ano06m, BM97, BYN^{+}04, BF05, FSI21, GJMO14, LB08, LLL^{+}14, LSCK15, LPR^{+}03, SdB03, Tan11, Oli01]. Volumes [FDMA97, LSB^{+}00, BZS08, WRB06]. Volumetric [GSU00, NWP97, SBS04, TG95a, TK97, AMCB20, MdBJG15, THL03, YW07]. Voronoi [BBB96, KSI98, NSK^{+}97]. VOS [KKKS24]. Voting [IF99, LZ97b, LBNS09, MGPJ11, RPG12, RC13, Sha06, SKBS13]. voxel [ALK^{+}09, GJMO14, HDZR24]. voxels [SB05]. VRML [FPDK12]. VRNN [BCC^{+}21]. vs [FCM20, KTP08, LHH^{+}98, TS00a].

walks [DB14, GB13]. Warping [YFZ98, LJHH07, SOJ17]. watching [CZ18]. Water [MTV17, PCR^{+}04, TKDN16]. watermarking [CWC^{+}20]. Watershed [BL00]. Watershed-Based [BL00]. Wave [ACF00]. Waved [WB15]. wavelengths [PS12]. Wavelet [AM00, DLC^{+}24, DLHT99, MAP99, SB22, TS00a, ÇÖD08, CT10, CT12, CE17, Hu11, LZmC^{+}17, LBCA10, MIP16, SG11]. Wavelet-Based [DLHT99, DLC^{+}24, SB22, CE17]. Wavelets [Ano95d, Far11, WLZW04]. Weak [SG17]. Weakly [AS23, DFLS23, EOPS22, HWG21, KHG22, KRG17, LCZ^{+}16, NF21, RDA^{+}15, SY20, GLG22, LYX^{+}24, LXW^{+}24, MDK^{+}24, MTP21, MFP^{+}20, NWNT17, PD14, PKH23, RZZ23, SS17b, TD19, UU18, WZW17, YK24, ZKSV18, ZZ20, ZZSD21]. weakly-supervised [NWNT17, RZZ23, UU18, ZZSD21]. wearable [LM16, NKB11, PYGGLNG17, SE11]. Web [SLST99, BBSD15, CMCM16, TL15, NY14, OB14, VGLP17]. web-scale [TL15, OB14]. Webly [HWW23, RKL^{+}18]. Webly-supervised [HWW23, RKL^{+}18]. Weibo [LZL^{+}17]. weight [CVP10, CKK^{+}12, HBG13, HZK19, WLL22b, WZCY22, ZYQ^{+}23, dMFU10]. Weighted [DH19, ASCF13, HQW^{+}12, JBR08, KNL15, LWW^{+}21, LDC^{+}13, LSPV04, ZWZ^{+}16, dCCP12]. weighting [JRAJ17, KYYC14]. weights [ET15, JRAJ17]. Weiss [HM97, May97, Ver97]. weld [FLCdA06]. Well [LER95, WB97]. Well-Composed [LER95, WB97]. What-and-Where [CGL98]. wheelchair [JZWD16, NPM^{+}16]. wheelchair-mounted [JZWD16]. Where [CGL98, CL17, EDJ^{+}20, MRdRGC23, TDV15, VZP^{+}09]. which [MRdRGC23]. while [TZM98]. whole [FO18, NZH^{+}23, YHS^{+}20]. whole-body [NZH^{+}23]. Wide [CKM11, SLST99, CCPK16, PK18, UWH17]. wide-angle [UWH17]. Wide-area [CKM11, CCPK16, PK18]. wild [CLTW23, HWK^{+}21, HP05, IZJ^{+}17, JGP19, JT17, OGB14, PMR17, SB18, SBN^{+}24, ZZZ15]. win [KKRK23]. win-win [KKRK23]. Window [JLL13, GS08, YHN11]. windows [TL16]. Wireless [Ziv10, LWLS12]. Wise [DF02, AC09b, CKC14, DZQ24, HWZ^{+}23, LLG^{+}24, WLXL24]. wisely [Pha17]. within [Kou03]. without [CB98, CYES00, JLM22, OD99, PLH04, Rob96a, RKL^{+}18, SLK23, SWMM22, YHS^{+}20]. Wize [ACC^{+}16]. WMCP [GGP23]. WMCP-EM [GGP23]. Word [KH96, KABP98, JN09, SKT18, WLZM20]. word-level [WLZM20]. words [CZ18, KBMD15, MYV19, PWWQ16, RG17, RB18]. workflow [ZZH23]. workflows [KDV12]. workspace [RGA10]. World [BPQ15, LSHT02, SLST99, DPCA15, HWL^{+}22, KH15, KPPK09, SHS^{+}23, ZZH23]. Wrinkles [YB01]. writer [PRG^{+}14].

X [AS08b, BMvT^{+}19, CZ14, GYW^{+}22, HT98, KHB01]. X-ray [AS08b, GYW^{+}22]. X-rays [BMvT^{+}19, CZ14].

YCb [BDFG17]. YCr [BDFG17]. Years [AT13, SOD10]. YIQ [LL08]. YOLO [ZJL23]. You-Do [DLMC16]. yourself [ZZS^{+}23].

ZDF [DBZ07]. Zero [DFH^{+}22, LLNZ22, DWL19, PHHL23, SAK^{+}24, XZX^{+}21]. Zeroshot [DFH^{+}22, LLNZ22, DWL19, PHHL23, SAK^{+}24, XZX^{+}21]. Zeta [DJF14]. Zetaimage [DJF14]. zones [TRG^{+}13]. Zoom [MPPG98, PEFM98, CXFS06, DDLP10, SPC^{+}15, SP06, SSdVL06, TM07]. Zoom-Invariant [MPPG98, PEFM98]. Zooming [LDPD97, ZZ07].

References

Akbas:2020:LLM

[AA20] Emre Akbas and Narendra Ahuja. Low-level multiscale image segmentation and a benchmark for its evaluation. Computer Vision and Image Understanding: CVIU, 199(??):Article 103026, October 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300837.

Abdelkader:2011:SBG

[AAASC11] Mohamed F. Abdelkader, Wael Abd-Almageed, Anuj Srivastava, and Rama Chellappa. Silhouette-based gesture and action recognition via modeling trajectories on Riemannian shape manifolds. Computer Vision and Image Understanding: CVIU, 115(3):439– 455, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ait-Aider:2019:FCM

[AAB19] Omar Ait-Aider and François Berry. A flexible calibration method for the intrinsic and mechanical parameters of panoramic linescan cameras. Computer Vision and Image Understanding: CVIU, 180(??): 47–58, March 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219300049.

Alsarhan:2022:EDG

[AAL22] Tamam Alsarhan, Usman Ali, and Hongtao Lu. Enhanced discriminative graph convolutional network with adaptive temporal modelling for skeleton-based action recognition. Computer Vision and Image Understanding: CVIU, 216(??): ??, February 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422100179X.

AlIsmaeil:2016:EDD

[AAMO16] Kassem Al Ismaeil, Djamila Aouada, Bruno Mirbach, and Björn Ottersten. Enhancement of dynamic depth scenes by upsampling for precise super-resolution (UP-SR). Computer Vision and Image Understanding: CVIU, 147(??):38– 49, June 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216300303.

Alvarez:2013:JAP

[ÁB13] Hugo Álvarez and Diego Borro. Junction assisted 3D pose retrieval of untextured 3D models in monocular images. Computer Vision and Image Understanding: CVIU, 117(10):1204– 1214, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001671.

Ardeshir:2018:ELE

[AB18] Shervin Ardeshir and Ali Borji. An exocentric look at egocentric actions and vice versa. Computer Vision and Image Understanding: CVIU, 171(??):61– 68, June 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218300717.

Armando:2023:MSD

[ABB^{+}23] Matthieu Armando, Laurence Boissieux, Edmond Boyer, Jean-Sébastien Franco, Martin Humenberger, Christophe Legras, Vincent Leroy, Mathieu Marsot, Julien Pansiot, Sergi Pujades, Rim Rekik, Grégory Rogez, Anilkumar Swamy, and Stefanie Wuhrer. 4DHumanOutfit: a multi-subject 4D dataset of human motion sequences in varying outfits exhibiting large displacements. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002163.

Assfalg:2003:SAS

[ABC^{+}03] Jürgen Assfalg, Marco Bertini, Carlo Colombo, Alberto Del Bimbo, and Walter Nunziati. Semantic annotation of soccer videos: automatic highlights identification. Computer Vision and Image Understanding: CVIU, 92(2–3):285– 305, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Alata:2011:GDP

[ABD11] O. Alata, S. Burg, and A. Dupas. Grouping/ degrouping point process, a point process driven by geometrical and topological properties of a partition in regions. Computer Vision and Image Understanding: CVIU, 115(9):1324– 1339, September 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001251.

Amayeh:2009:HBV

[ABEN09] Gholamreza Amayeh, George Bebis, Ali Erol, and Mircea Nicolescu. Hand-based verification and identification using palm–finger segmentation and fusion. Computer Vision and Image Understanding: CVIU, 113(4): 477–501, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Amir:2004:MMS

[ABI^{+}04] Arnon Amir, Sankar Basu, Giridharan Iyengar, Ching-Yung Lin, Milind Naphade, John R. Smith, Savitha Srinivasan, and Belle Tseng. A multi-modal system for the retrieval of semantic video events. Computer Vision and Image Understanding: CVIU, 96(2):216–236, November 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Alcazar:2021:MMA

[ABJ^{+}21] Juan León Alcázar, María A. Bravo, Guillaume Jeanneret, Ali K. Thabet, Thomas Brox, Pablo Arbeláez, and Bernard Ghanem. MAIN: Multi-attention instance network for video segmentation. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000849.

Avgerinakis:2016:ADU

[ABK16] Konstantinos Avgerinakis, Alexia Briassouli, and Yiannis Kompatsiaris. Activity detection using Sequential Statistical Boundary Detection (SSBD). Computer Vision and Image Understanding: CVIU, 144(??): 46–61, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002337.

Abebe:2018:TAC

[ABK^{+}18] Mekides Assefa Abebe, Alexandra Booth, Jonathan Kervec, Tania Pouli, and Mohamed-Chaker Larabi.

Towards an automatic correction of over-exposure in photographs: Application to tone-mapping. Computer Vision and Image Understanding: CVIU, 168(??): 3–20, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300954.

Audebert:2019:DTR

[ABLL19] Nicolas Audebert, Alexandre Boulch, Bertrand Le Saux, and Sébastien Lefèvre. Distance transform regression for spatially-aware deep semantic segmentation. Computer Vision and Image Understanding: CVIU, 189(??):Article 102809, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314219301201.

Ahmad:2020:HLD

[ABN^{+}20] Touqeer Ahmad, George Bebis, Monica Nicolescu, Ara Nefian, and Terry Fong. Horizon line detection using supervised learning and edge cues. Computer Vision and Image Understanding: CVIU, 191(??):Article 102879, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218302030.

Ariz:2016:NDA

[ABVC16] Mikel Ariz, José J. Bengoechea, Arantxa Villanueva, and Rafael Cabeza. A novel 2D/3D database with automatic face annotation for head tracking and pose estimation. Computer Vision and Image Understanding: CVIU, 148(??):201– 210, July 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215000934.

Amini:1997:BIA

[ABW97] Amir A. Amini, Fred L. Bookstein, and David C. Wilson. Biomedical image analysis. Computer Vision and Image Understanding: CVIU, 66(2): 95–96, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0616/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0616/production/pdf.

Aggarwal:1999:HMA

[AC99] J. K. Aggarwal and Q. Cai. Human motion analysis: a review. Computer Vision and Image Understanding: CVIU, 73(3):428– 440, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0744/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0744/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0744/production/ref.

Allili:2007:TAS

[AC07] M. Allili and D. Corriveau. Topological analysis of shapes using Morse theory. Computer Vision and Image Understanding: CVIU, 105(3):188– 199, March 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Arandjelovic:2009:MRI

[AC09a] Ognjen Arandjelović and Roberto Cipolla. A methodology for rapid illuminationinvariant face recognition using image processing filters. Computer Vision and Image Understanding: CVIU, 113(2):159–171, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Arandjelovic:2009:PWL

[AC09b] Ognjen Arandjelović and Roberto Cipolla. A posewise linear illumination manifold model for face recognition using video. Computer Vision and Image Understanding: CVIU, 113 (1):113–125, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Alberich-Carraminana:2008:RED

[ACAAC^{+}08] Maria Alberich-Carramiñana, Guillem Alenyà, Juan Andrade-Cetto, Elisa Martínez, and Carme Torras. Recovering epipolar direction from two affine views of a planar object. Computer Vision and Image Understanding: CVIU, 112(2):195–209, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Araujo:1998:FPF

[ACB98] Helder Araújo, Rodrigo L. Carceroni, and Christopher M. Brown. A fully projective formulation to improve the accuracy of Lowe’s pose-estimation algorithm. Computer Vision and Image Understanding: CVIU, 70 (2):227–238, May 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0632/production;   http:
//www.idealibrary.com/

links/artid/cviu.1997.
0632/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0632/production/ref.

Andreu:2016:WMS

[ACC^{+}16] Yasmina Andreu, Franco Chiarugi, Sara Colantonio, Giorgos Giannakakis, Daniela Giorgi, Pedro Henriquez, Eleni Kazantzaki, Dimitris Manousos, Kostas Marias, Bogdan J. Matuszewski, Maria Antonietta Pascali, Matthew Pediaditis, Giovanni Raccichini, and Manolis Tsiknakis. Wize Mirror — a smart, multisensory cardiometabolic risk monitoring system. Computer Vision and Image Understanding: CVIU, 148(??): 3–22, July 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216300224.

Azari:2012:OPA

[ACDB12] Hossein Azari, Irene Cheng, Kostas Daniilidis, and Anup Basu. Optimal pixel aspect ratio for enhanced 3D TV visualization. Computer Vision and Image Understanding: CVIU, 116(1): 38–53, January 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001871.

Adan:2000:MWS

[ACF00] Antonio Adán, Carlos Cerrada, and Vicente Feliu. Modeling wave set: Definition and application of a new topological organization for 3D object modeling. Computer Vision and Image Understanding: CVIU, 79(2):281–307, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2000.0855;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0855/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0855/ref.

Alvarez:2009:NEB

[ACG^{+}09] L. Alvarez, C. A. Castaño, M. García, K. Krissian, L. Mazorra, A. Salgado, and J. Sánchez. A new energybased method for 3D motion estimation of incompressible PIV flows. Computer Vision and Image Understanding: CVIU, 113(7):802– 810, July 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Aggarwal:1998:NMA

[ACLS98] J. K. Aggarwal, Q. Cai, W. Liao, and B. Sabata. Nonrigid motion analysis: Articulated and elastic motion. Computer Vision and Image Understanding: CVIU, 70(2): 142–156, May 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0620/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0620/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0620/production/ref.

Abebe:2016:RMD

[ACP16] Girmaw Abebe, Andrea Cavallaro, and Xavier Parra. Robust multi-dimensional motion features for firstperson vision activity recognition. Computer Vision and Image Understanding: CVIU, 149(??):229– 248, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002350.

Arbel:2016:ESI

[ACW^{+}16] Tal Arbel, M. Jorge Cardoso, William Wells III, Albert C. S. Chung, and Doina Precup. Editorial on special issue on probabilistic models for biomedical image analysis. Computer Vision and Image Understanding: CVIU, 151(??): 1–2, October 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301199.

Ahmed:2019:LSR

[ADC19] Sk. Miraj Ahmed, Niladri Ranjan Das, and Kunal Narayan Chaudhury. Least-squares registration of point sets over SE(d) using closed-form projections. Computer Vision and Image Understanding: CVIU, 183 (??):20–32, June 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219300451.

Avrithis:1999:SFO

[ADDK99] Yannis S. Avrithis, Anastasios D. Doulamis, Nikolaos D. Doulamis, and Stefanos D. Kollias. A stochastic framework for optimal key frame extraction from MPEG video databases. Computer Vision and Image Understanding: CVIU, 75(1–2):3–24, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.

0761/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0761/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0761/production/ref.

Aghaei:2018:TSP

[ADFR18] Maedeh Aghaei, Mariella Dimiccoli, Cristian Canton Ferrer, and Petia Radeva. Towards social pattern characterization in egocentric photo-streams. Computer Vision and Image Understanding: CVIU, 171(??): 104–117, June 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300675.

Ali:2016:IIO

[ADGB16] Sharib Ali, Christian Daul, Ernest Galbrun, and Walter Blondel. Illumination invariant optical flow using neighborhood descriptors. Computer Vision and Image Understanding: CVIU, 145(??):95– 110, April 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002672.

Aghaei:2016:MFT

[ADR16] Maedeh Aghaei, Mariella Dimiccoli, and Petia Radeva.

Multi-face tracking by extended bag-of-tracklets in egocentric photo-streams. Computer Vision and Image Understanding: CVIU, 149 (??):146–156, August 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314216000679.

Aqmar:2014:GRF

[AFMY14] Muhammad Rasyid Aqmar, Yusuke Fujihara, Yasushi Makihara, and Yasushi Yagi. Gait recognition by fluctuations. Computer Vision and Image Understanding: CVIU, 126(??): 38–52, September 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421400112X.

Alexander:2000:EMD

[AG00] Daniel C. Alexander and James C. Gee. Elastic matching of diffusion tensor images. Computer Vision and Image Understanding: CVIU, 77(2): 233–250, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0817;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.

1999.0817/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0817/ref.

Albert:2015:NMA

[AGB^{+}15] F. Albert, J. M. Gómis, J. Blasco, J. M. Valiente, and N. Aleixos. A new method to analyse mosaics based on Symmetry Group theory applied to Islamic Geometric Patterns. Computer Vision and Image Understanding: CVIU, 130(??):54–70, January 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001842.

Achddou:2023:FST

[AGL23] Raphaël Achddou, Yann Gousseau, and Saïd Ladjal. Fully synthetic training for image restoration tasks. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001030.

Atkinson:2008:TDB

[AH08] Gary A. Atkinson and Edwin R. Hancock. Twodimensional BRDF estimation from polarisation. Computer Vision and Image Understanding: CVIU, 111 (2):126–141, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ahmed:2024:TSD

[AH24] Shahed Ahmed and Md. Kamrul Hasan. Twin-SegNet: Dynamically coupled complementary segmentation networks for generalized medical image segmentation. Computer Vision and Image Understanding: CVIU, 240(??):??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002904.

Abuhaiba:1998:RLC

[AHD98] I. S. I. Abuhaiba, M. J. J. Holt, and S. Datta. Recognition of off-line cursive handwriting. Computer Vision and Image Understanding: CVIU, 71(1): 19–38, July 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0629/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0629/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0629/production/ref.

Appiah:2010:AHV

[AHDM10] Kofi Appiah, Andrew Hunter, Patrick Dickinson, and Hongying Meng. Accelerated hardware video object segmentation: From foreground detection to connected components labelling. Computer Vision and Image Understanding: CVIU, 114(11):1282–1291, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Alves:2017:UL

[AHM17] Wonder A. L. Alves, Ronaldo F.

Hashimoto, and Beatriz Marcotegui. Ultimate levelings. Computer Vision and Image Understanding: CVIU, 165(??):60–74, December 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301224.

Abedini:2023:SMI

[AJ23] Zeinab Abedini and Mansour Jamzad. Single and multiple illuminant estimation using convex functions. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000917.

Ambrosch:2010:AHB

[AK10] Kristian Ambrosch and Wilfried Kubinger. Accurate hardware-based stereo vision. Computer Vision and Image Understanding: CVIU, 114(11):1303–1316, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ambrosch:2011:CAH

[AK11] Kristian Ambrosch and Wilfried Kubinger. Corrigendum to “Accurate hardware-based stereo vision” [Comput. Vis. Image Understanding 114 (2010) 1303–1316]. Computer Vision and Image Understanding: CVIU, 115(2):287, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Arashloo:2011:PIF

[AKC11] Shervin Rahimzadeh Arashloo, Josef Kittler, and William J. Christmas. Pose-invariant face recognition by matching on multi-resolution MRFs linked by supercoupling transform. Computer Vision and Image Understanding: CVIU, 115(7): 1073–1083, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211000749.

Ao:2023:IAC

[AKE23] Jiayang Ao, Qiuhong Ke, and Krista A. Ehinger. Image amodal completion: a survey. Computer Vision and Image Understanding: CVIU, 229(??): ??, March 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000413.

Amir:1999:GFD

[AL99] Arnon Amir and M. Lindenbaum. Ground from figure discrimination. Computer Vision and Image Understanding: CVIU, 76(1): 7–18, October 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0786/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0786/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0786/production/ref;
http://www.idealibrary.
com/links/artid/cviu.1999.
0797/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0797/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0797/production/ref.

Aziz:2018:GSC

[ALIRT18] Fatima Aziz, Ouiddad Labbani-Igbida, Amina Radgui, and Ahmed Tamtaoui. Generic spatial-color metric for scale-space processing of catadioptric images. Computer Vision and Image Understanding: CVIU, 176–177(??):54–69, November/December 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218302479.

Adjeroh:1999:DMV

[ALK99] Donald A. Adjeroh, M. C. Lee, and Irwin King. A distance measure for video sequences. Computer Vision and Image Understanding: CVIU, 75(1–2):25–45, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0764/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0764/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0764/production/ref.

Anderson:2009:LSV

[ALK^{+}09] Derek Anderson, Robert H. Luke, James M. Keller, Marjorie Skubic, Marilyn Rantz, and Myra Aud. Linguistic summarization of video for fall detection using voxel person and fuzzy logic. Computer Vision and Image Understanding: CVIU, 113(1):80–89, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ali:2023:ISP

[ALM23] Usman Ali, Ik Hyun Lee, and Muhammad Tariq Mahmood. Incorporating structural prior for depth regularization in shape from focus. Computer Vision and Image Understanding: CVIU, 227 (??):??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001977.

An:2022:DSI

[ALY^{+}22] Pei An, Junxiong Liang, Kun Yu, Bin Fang, and Jie Ma. Deep structural information fusion for 3D object detection on LiDARcamera system. Computer Vision and Image Understanding: CVIU, 214(??): ??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001399.

Attali:1997:CSC

[AM97] Dominique Attali and Annick Montanvert. Computing and simplifying 2D and 3D continuous skeletons. Computer Vision and Image Understanding: CVIU, 67(3):261–273, September 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0536/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0536/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0536/production/ref.

Angel:2000:AMW

[AM00] Paul Angel and Colin Morris. Analyzing the Mallat wavelet transform to delineate contour and textural features. Computer Vision and Image Understanding: CVIU, 80(3): 267–288, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0877;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0877/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0877/ref.

Aguado:2001:PLS

[AM01] Alberto S. Aguado and Eugenia Montiel. Progressive linear search for stereo matching and its application to interframe interpolation. Computer Vision and Image Understanding: CVIU, 81(1):46–71, January 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0886;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0886/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0886/ref.

Anderson:2004:RRT

[AM04] Keith Anderson and Peter W. McOwan. Robust real-time face tracker for cluttered environments. Computer Vision and Image Understanding: CVIU, 95(2):184–200, August 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Aguena:2006:MID

[AM06] Marcia L. S. Aguena and Nelson D. A. Mascarenhas. Multispectral image data fusion using POCS and superresolution. Computer Vision and Image Understanding: CVIU, 102(2):178– 187, May 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Atto:2015:HOS

[AM15] Abdourrahmane Mahamane Atto and Grégoire Mercier. High order structural image decomposition by using non-linear and nonconvex regularizing objectives. Computer Vision and Image Understanding: CVIU, 138(??):38–50, September 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215000867.

Alotaibi:2017:IGR

[AM17] Munif Alotaibi and Ausif Mahmood. Improved gait recognition based on specialized deep convolutional neural network. Computer Vision and Image Understanding: CVIU, 164(??): 103–110, November 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314217301674.

Almakady:2020:RIF

[AMCB20] Yasseen Almakady, Sasan Mahmoodi, Joy Conway, and Michael Bennett. Rotation invariant features based on three dimensional Gaussian Markov random fields for volumetric texture classification. Computer Vision and Image Understanding: CVIU, 194(??):Article 102931, May 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300205.

Arias:2016:MIM

[AMGG^{+}16] Jacinto Arias, Jesus Martínez-Gómez, Jose A. Gámez, Alba G. Seco de Herrera, and Henning Müller. Medical image modality classification using discrete Bayesian networks. Computer Vision and Image Understanding: CVIU, 151 (??):61–71, October 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314216300261.

Armande:1999:TNE

[AMMV99] N. Armande, P. Montesinos, O. Monga, and Guy Vaysseix. Thin nets extraction using a multi-scale approach. Computer Vision and Image Understanding: CVIU, 73(2):248– 257, February 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0658/production;   http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0658/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0658/production/ref.

Agudo:2018:SEA

[AMN18] Antonio Agudo and Francesc Moreno-Noguer. A scalable, efficient, and accurate solution to non-rigid structure from motion. Computer Vision and Image Understanding: CVIU, 167(??):121– 133, February 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421830002X.

Agudo:2016:RTR

[AMNCM16] Antonio Agudo, Francesc Moreno-Noguer, Begoña Calvo, and J. M. M. Montiel. Real-time 3D reconstruction of non-rigid shapes with a single moving camera. Computer Vision and Image Understanding: CVIU, 153(??):37–54, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216300522.

Angulo:2007:MCO

[Ang07] Jesús Angulo. Morphological colour operators in totally ordered lattices based on distances: Application to image filtering, enhancement and analysis. Computer Vision and Image Understanding: CVIU, 107(1– 2):56–73, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Al-Nuaimi:2017:DRT

[ANHGS17] Anas Al-Nuaimi, Sebastian Hilsenbeck, Adrian Garcea, and Eckehard Steinbach. 6DOF decoupled roto-translation alignment of large-scale indoor point clouds. Computer Vision and Image Understanding: CVIU, 157(??): 72–89, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301126.

Aguado:1998:PAS

[ANM98] Alberto S. Aguado, Mark S.

Nixon, and M. Eugenia Montiel. Parameterizing arbitrary shapes via Fourier descriptors for evidence-gathering extraction. Computer Vision and Image Understanding: CVIU, 69(2):202–221, February 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.

                                        0558/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0558/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0558/production/ref.

Anonymous:1995:ASA

[Ano95a] Anonymous. ACCV ’95 Second Asian Conference on Computer Vision. Computer Vision and Image Understanding: CVIU, 61(1): 151, January 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1011/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1011/production/pdf.

Anonymous:1995:AIVa

[Ano95b] Anonymous. Author index for volume 61. Computer Vision and Image Understanding: CVIU, 61 (3):475, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1036/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1036/production/pdf.

Anonymous:1995:AIVb

[Ano95c] Anonymous. Author index for volume 62. Computer Vision and Image Understanding: CVIU, 62(3): 392, November 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1063/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1063/production/pdf.

Anonymous:1995:RMM

[Ano95d] Anonymous. Review of Mathematical Morphology in Image Processing, by E. R. Dougherty and review of Wavelets, Images and Surface Fitting, by P.-J. Laurent, A. LeMéHauté, and L. L. Schumaker. Computer Vision and Image Understanding: CVIU, 61(2): 292, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1022/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1022/production/pdf.

Anonymous:1995:RNC

[Ano95e] Anonymous. Review of NURB Curves and Surfaces-From Projective Geometry to Practical Use, by G. Farin and review of An Introduction to Scientific, Symbolic, and Graphical Computation, by E. Fiume. Computer Vision and Image Understanding: CVIU, 62 (1):144, July 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1045/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1045/production/pdf.

Anonymous:1996:ABS

[Ano96a] Anonymous. ANNOUNCE-MENT: BMVC96 Seventh British Machine Vision Conference. Computer Vision and Image Understanding: CVIU, 63(2): 397, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0030/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0030/production/pdf.

Anonymous:1996:AIVa

[Ano96b] Anonymous. Author index for volume 63. Computer Vision and Image Understanding: CVIU, 63 (3):613, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/

links/artid/cviu.1996.
0042/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0042/production/pdf.

Anonymous:1996:AIVb

[Ano96c] Anonymous. Author index for volume 64. Computer Vision and Image Understanding: CVIU, 64(3): 443, November 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0071/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0071/production/pdf.

Anonymous:1996:IFI

[Ano96d] Anonymous. ICDAR ’97:

Fourth international conference on document analysis and recognition. Computer Vision and Image Understanding: CVIU, 64(2):303– 304, September 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0061/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0061/production/pdf.

Anonymous:1997:A

[Ano97a]  Anonymous.  Announce-

ment. Computer Vision and Image Understanding: CVIU, 67(3):324–??, September 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0642/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0642/production/pdf.

Anonymous:1997:AIVa

[Ano97b] Anonymous. Author index for volume 65. Computer Vision and Image Understanding: CVIU, 65(3): 455, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0622/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0622/production/pdf.

Anonymous:1997:AIVb

[Ano97c] Anonymous. Author index for volume 66. Computer Vision and Image Understanding: CVIU, 66(3): 347–??, June 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0636/production;   http:
//www.idealibrary.com/

links/artid/cviu.1997.
0636/production/pdf.

Anonymous:1997:AIVc

[Ano97d] Anonymous. Author index for volume 67. Computer Vision and Image Understanding: CVIU, 67(3):325– ??, September 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0654/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0654/production/pdf.

Anonymous:1997:AIVd

[Ano97e] Anonymous. Author index for volume 68. Computer Vision and Image Understanding: CVIU, 68(3):363– ??, December 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0671/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0671/production/pdf.

Anonymous:1997:BRR

[Ano97f]  Anonymous.  Books re-

ceived for review. Computer Vision and Image Understanding: CVIU, 65(1): 109, January 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0588/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0588/production/pdf.

Anonymous:1998:AIVa

[Ano98a] Anonymous. Author index for volume 71. Computer Vision and Image Understanding: CVIU, 71(3): 449, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0725/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0725/production/pdf.

Anonymous:1998:AIVb

[Ano98b] Anonymous. Author index for volume 72. Computer Vision and Image Understanding: CVIU, 72(3): 414, December 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0740/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0740/production/pdf.

Anonymous:1998:BRR

[Ano98c] Anonymous. BOOKS RE-CEIVED FOR REVIEW.

Computer Vision and Image Understanding: CVIU, 70(1):120, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0677/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0677/production/pdf.

Anonymous:1998:CVV

[Ano98d]  Anonymous.  Computer

vision for visual computing: Techniques and applications. Computer Vision and Image Understanding: CVIU, 71(2): 153, August 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0714/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0714/production/pdf.

Anonymous:1999:AIVa

[Ano99a] Anonymous. Author index for volume 73. Computer Vision and Image Understanding: CVIU, 73(3): 455, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0749/production;   http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0749/production/pdf.

Anonymous:1999:AIVb

[Ano99b] Anonymous. Author index for volume 74. Computer Vision and Image Understanding: CVIU, 74 (3):236, June 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0774/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0774/production/pdf.

Anonymous:1999:AIVc

[Ano99c] Anonymous. Author index for volume 75. Computer Vision and Image Understanding: CVIU, 75(3): 319, September 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0796/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0796/production/pdf.

Anonymous:1999:AIVd

[Ano99d] Anonymous. Author index for volume 76. Computer Vision and Image Understanding: CVIU, 76(3): 298, December 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0820/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0820/production/pdf.

Anonymous:2000:AIVa

[Ano00a] Anonymous. Author index for volume 77. Computer Vision and Image Understanding: CVIU, 77(3): 388, March 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0839;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0839/pdf.

Anonymous:2000:AIVb

[Ano00b] Anonymous. Author index for volume 78. Computer Vision and Image Understanding: CVIU, 78 (3):374, June 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0863;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0863/pdf.

Anonymous:2000:AIVc

[Ano00c] Anonymous. Author index for volume 79. Computer Vision and Image Understanding: CVIU, 79 (3):442, September 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2000.0873;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0873/pdf.

Anonymous:2000:AIVd

[Ano00d] Anonymous. Author index for volume 80. Computer Vision and Image Understanding: CVIU, 80 (3):384, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2000.0898;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0898/pdf.

Anonymous:2001:Aa

[Ano01a]  Anonymous.  Announce-

ment. Computer Vision and Image Understanding: CVIU, 83(2):??, August 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:Ab

[Ano01b]  Anonymous.  Announce-

ment. Computer Vision and Image Understanding:

CVIU, 84(1):??, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:AIVa

[Ano01c] Anonymous. Author index for volume 81. Computer Vision and Image Understanding: CVIU, 81(3): 446, March 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0911;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0911/pdf.

Anonymous:2001:AIVb

[Ano01d] Anonymous. Author index for volume 82. Computer Vision and Image Understanding: CVIU, 82 (3):255, June 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0928;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0928/pdf.

Anonymous:2001:AIVc

[Ano01e] Anonymous. Author index for volume 83. Computer Vision and Image Understanding: CVIU, 83 (3):296, September 2001.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2001.0942;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0942/pdf.

Anonymous:2001:AIVd

[Ano01f] Anonymous. Author index for volume 84. Computer Vision and Image Understanding: CVIU, 84(3):409, December 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:GE

[Ano01g] Anonymous. Guest editorial. Computer Vision and Image Understanding: CVIU, 84(1):??, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:Na

[Ano01h] Anonymous. Note. Computer Vision and Image Understanding: CVIU, 82(3): ??, June 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:Nb

[Ano01i] Anonymous. Note. Computer Vision and Image Understanding: CVIU, 83(1): ??, July 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:Nc

[Ano01j] Anonymous. Note. Computer Vision and Image Understanding: CVIU, 83(3): ??, September 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:SIF

[Ano01k] Anonymous. Special issue on face recognition: Call for papers. Computer Vision and Image Understanding: CVIU, 84 (1):198–199, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0947;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0947/pdf.

Anonymous:2001:SIN

[Ano01l] Anonymous. Special issue on nonrigid image registration: Call for papers. Computer Vision and Image Understanding: CVIU, 83(2):192–193, August 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0929;   http:

                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0929/pdf.

Anonymous:2001:VNA

[Ano01m] Anonymous. Volume 81, number 1 (2001), in the article “A Frequency Domain Technique Based on Energy Radial Projections for Robust Estimation of Global 2D Affine Transformations,” by Luca Lucchese, pages 72–116. Computer Vision and Image Understanding: CVIU, 82(1): 82–83, April 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2001.0913;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0913/pdf.

Anonymous:2002:AIVa

[Ano02a] Anonymous. Author index for volume 85. Computer Vision and Image Understanding: CVIU, 85(3): 232, March 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2002:AIVb

[Ano02b] Anonymous. Author index for volume 86. Computer Vision and Image Understanding: CVIU, 86(3): 191, June 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2002:AIVc

[Ano02c] Anonymous. Author index for volume 87. Computer Vision and Image Understanding: CVIU, 87(1–3): 131, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2002:AIVd

[Ano02d] Anonymous. Author index for volume 88. Computer Vision and Image Understanding: CVIU, 88(3):189, December 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:Aa

[Ano03a]  Anonymous.  Announce-

ment. Computer Vision and Image Understanding: CVIU, 89(1):108, January 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:Ab

[Ano03b]  Anonymous.  Announce-

ment. Computer Vision and Image Understanding: CVIU, 89(2–3):318, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:Ac

[Ano03c]  Anonymous.  Announce-

ment. Computer Vision and Image Understanding: CVIU, 90(1):128, April 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBa

[Ano03d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 89(1):C02, January 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBb

[Ano03e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 89(2–3):C02, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBc

[Ano03f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 90(1):C02, April 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBd

[Ano03g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 90(2):C02, May 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBe

[Ano03h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 90(3):C02, June 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBf

[Ano03i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 91(1–2):C02, July/ August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBg

[Ano03j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 91(3):C02, September 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBh

[Ano03k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 92(1):C02, October 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBi

[Ano03l]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 92(2–3):CO2, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:PN

[Ano03m]  Anonymous.  Publisher’s

note. Computer Vision and Image Understanding: CVIU, 89(1):iii–iv, January 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:VAIa

[Ano03n] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 89(2–3):319, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:VAIb

[Ano03o] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 90(3):313, June 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:VAIc

[Ano03p] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 91(3):368, September 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:VAId

[Ano03q] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 92(2–3):306, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBg

[Ano04a] Anonymous. Ed. board.

Computer Vision and Image Understanding: CVIU, 95(3):CO2, September 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBh

[Ano04b] Anonymous. Ed. board.

Computer Vision and Image Understanding: CVIU, 96(1):CO2, October 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBi

[Ano04c] Anonymous. Ed. board.

Computer Vision and Image Understanding: CVIU, 96(2):CO2, November 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBj

[Ano04d] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 96(3):CO2, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBa

[Ano04e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 93(1):CO2, January 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBb

[Ano04f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 93(2):CO2, February 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBc

[Ano04g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 93(3):CO2, March 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBd

[Ano04h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 94(1–3):CO2, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBe

[Ano04i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 95(1):CO2, July 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBf

[Ano04j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 95(2):CO2, August 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:VAIa

[Ano04k] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 93(3):347, March 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:VAIb

[Ano04l] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 94(1–3):311, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:VAIc

[Ano04m] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 95(3):354, September 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:VAId

[Ano04n] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 96(3):472, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBa

[Ano05a] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 97(1):CO2, January 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBb

[Ano05b] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 97(2):CO2, February 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBc

[Ano05c] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 97 (3):CO2, March 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBd

[Ano05d] Anonymous. Ed. board.

Computer Vision and Image Understanding: CVIU, 98(1):CO2, April 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBe

[Ano05e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 98(2):CO2, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBf

[Ano05f]  Anonymous.  Editorial

board. Computer Vision and Image Understanding: CVIU, 98(3):CO2, June 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBg

[Ano05g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 99(1):CO2, July 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBh

[Ano05h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 99(2):CO2, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBi

[Ano05i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 100(1–2):??, October/November 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:SIA

[Ano05j] Anonymous. Special issue: Attention and performance in computer vision. Computer Vision and Image Understanding: CVIU, 100(1– 2):1–2, October/November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:VAIa

[Ano05k] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 97(3):384, March 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:VAIb

[Ano05l] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 98(3):513, June 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:VAIc

[Ano05m] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 99(3):527, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:VAId

[Ano05n] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 100(3):458, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBf

[Ano06a] Anonymous. Ed. Board.

Computer Vision and Image Understanding: CVIU, 103 (2):??, August 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBg

[Ano06b] Anonymous. Ed. Board.

Computer Vision and Image Understanding: CVIU, 104 (1):??, October 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBa

[Ano06c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 101(1):??, January 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBb

[Ano06d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 101(2):??, February 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBc

[Ano06e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 102(1):??, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBd

[Ano06f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 102(2):??, May 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBe

[Ano06g]  Anonymous.  Editorial

board. Computer Vision and Image Understanding: CVIU, 103(1):??, July 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EMK

[Ano06h] Anonymous. Erratum to “A multi-Kalman filtering approach for video tracking of human-delineated objects in cluttered environments” [Comput. Vision Image Understanding 99 (2005) 1–57]. Computer Vision and Image Understanding: CVIU, 102(3): 259, June 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See [GKK05].

Anonymous:2006:PN

[Ano06i]  Anonymous.  Publisher’s

Note. Computer Vision and Image Understanding: CVIU, 101(1):??, January 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:VAIa

[Ano06j] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 101(3):204, March 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:VAIb

[Ano06k] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 102(3):260– 316, June 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:VAIc

[Ano06l] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 103(3):229, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:VAId

[Ano06m] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 104(2–3):258, November/December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:CEBa

[Ano07a] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 105(1):??, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:CEBb

[Ano07b] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 107(1–2):??, July/ August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:CEBc

[Ano07c] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 108(1–2):??, October/November 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:EBa

[Ano07d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 105(2):??, February 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:EBb

[Ano07e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 106(1):??, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:MCP

[Ano07f] Anonymous. Most Cited Paper Award. Computer Vision and Image Understanding: CVIU, 106(1):1– 2, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBa

[Ano08a] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding:

CVIU, 109(1):??, January 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBb

[Ano08b] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 109(2):??, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBc

[Ano08c] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 110(1):??, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBd

[Ano08d] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 110(2):??, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBe

[Ano08e] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 111(1):??, July 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBf

[Ano08f] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 111(2):??, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBg

[Ano08g] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 111(3):??, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBh

[Ano08h] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 112(1):??, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBi

[Ano08i] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 112(2):??, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBj

[Ano08j] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 112(3):??, December 2008. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:MCP

[Ano08k] Anonymous. Most Cited Paper Award. Computer Vision and Image Understanding: CVIU, 111(3):247–248, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBa

[Ano09a] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(1):??, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBb

[Ano09b] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(2):??, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBc

[Ano09c] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(3):??, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBd

[Ano09d] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(4):??, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBe

[Ano09e] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 113(5):??, May 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBf

[Ano09f] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 113(6):??, June 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBg

[Ano09g] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 113(7):??, July 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBh

[Ano09h] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 113(8):??, August 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBi

[Ano09i] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(9):??, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBj

[Ano09j] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(10):??, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBk

[Ano09k] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(11):??, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBa

[Ano10a] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(1):??, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBb

[Ano10b] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(2):??, February 2010. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBc

[Ano10c] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 114(3):??, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBd

[Ano10d] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 114(4):??, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBe

[Ano10e] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 114(5):??, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBf

[Ano10f] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 114(6):??, June 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBg

[Ano10g] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding:

CVIU, 114(7):??, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBh

[Ano10h] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(8):??, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBi

[Ano10i] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(9):??, September 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBj

[Ano10j] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(10):??, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBk

[Ano10k] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(11):??, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBa

[Ano11a]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(1):??, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBb

[Ano11b]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(2):??, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBc

[Ano11c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(3):??, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBd

[Ano11d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(4):??, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBe

[Ano11e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(5):??, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBf

[Ano11f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(7): ??, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211001111.

Anonymous:2011:EBg

[Ano11g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(8): ??, August 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001299.

Anonymous:2011:EBh

[Ano11h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(9): ??, September 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001469.

Anonymous:2011:EBi

[Ano11i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(10): ??, October 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421100172X.

Anonymous:2011:EBj

[Ano11j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(11): ??, November 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001901.

Anonymous:2011:EBk

[Ano11k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 115(12):??, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002189.

Anonymous:2012:EBa

[Ano12a]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(1): ??, January 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002426.

Anonymous:2012:EBb

[Ano12b]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(2): ??, February 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211002657.

Anonymous:2012:EBc

[Ano12c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(3): ??, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314212000057.

Anonymous:2012:EBd

[Ano12d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(4): ??, April 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314212000240.

Anonymous:2012:EBe

[Ano12e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(5): ??, May 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212000409.

Anonymous:2012:EBf

[Ano12f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(6): ??, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212000550.

Anonymous:2012:EBg

[Ano12g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(7): ??, July 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212000689.

Anonymous:2012:EBh

[Ano12h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(8): ??, August 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000811.

Anonymous:2012:EBi

[Ano12i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(9): ??, September 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212000926.

Anonymous:2012:EBm

[Ano12j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(10): ??, October 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001087.

Anonymous:2012:EBq

[Ano12k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(11): ??, November 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421200121X.

Anonymous:2012:EBr

[Ano12l]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 116(12):??, December 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001373.

Anonymous:2012:MCP

[Ano12m] Anonymous. Most cited paper award. Computer Vision and Image Understanding: CVIU, 116(5): 661–662, May 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000343.

Anonymous:2012:RA

[Ano12n] Anonymous. Reviewer acknowledgment. Computer Vision and Image Understanding: CVIU, 116(1): ??, January 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002542.

Anonymous:2013:EBa

[Ano13a]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(1): ??, January 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001567.

Anonymous:2013:EBh

[Ano13b]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(2): C2, February 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001993.

Anonymous:2013:EBb

[Ano13c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(2): C2, February 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001993.

Anonymous:2013:EBi

[Ano13d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(3): ??, March 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213000064.

Anonymous:2013:EBc

[Ano13e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(3): C2, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000064.

Anonymous:2013:EBj

[Ano13f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(4): C2, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000180.

Anonymous:2013:EBd

[Ano13g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(4): C2, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000180.

Anonymous:2013:EBe

[Ano13h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(5): C2, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000398.

Anonymous:2013:EBl

[Ano13i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(6): C2, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421300057X.

Anonymous:2013:EBn

[Ano13j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(7): C2, July 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213000817.

Anonymous:2013:EBo

[Ano13k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(8): C2, August 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213000982.

Anonymous:2013:EBp

[Ano13l]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(9): C2, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001112.

Anonymous:2013:EBq

[Ano13m]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(10):C2, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001434.

Anonymous:2013:EBr

[Ano13n]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 117(11):C2, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421300163X.

Anonymous:2013:MCP

[Ano13o] Anonymous. Most cited paper award. Computer Vision and Image Understanding: CVIU, 117(5): 571, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000441.

Anonymous:2013:RA

[Ano13p] Anonymous. Reviewer acknowledgment. Computer Vision and Image Understanding: CVIU, 117(1): ??, January 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001543.

Anonymous:2014:EBa

[Ano14a]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 124(??): ifc, July 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001076.

Anonymous:2014:EBb

[Ano14b]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 125(??): ifc, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214001192.

Anonymous:2014:EBc

[Ano14c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 126(??): ifc, September 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421400143X.

Anonymous:2014:EBd

[Ano14d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 127(??): ifc, October 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001635.

Anonymous:2014:EBe

[Ano14e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 128(??):IFC, November 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001738.

Anonymous:2014:EBf

[Ano14f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 129(??):IFC, December 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001957.

Anonymous:2014:RA

[Ano14g] Anonymous. Reviewer acknowledgment. Computer Vision and Image Understanding: CVIU, 118(??): i–iii, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421300218X.

Anonymous:2015:CEBa

[Ano15a] Anonymous. COV2: Editorial Board. Computer Vision and Image Understanding: CVIU, 131(??): ifc, February 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214002288.

Anonymous:2015:CEBb

[Ano15b] Anonymous. COV2: Editorial Board. Computer Vision and Image Understanding: CVIU, 132(??): IFC, March 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215000053.

Anonymous:2015:CEBc

[Ano15c] Anonymous. COV2: Editorial Board. Computer Vision and Image Understanding: CVIU, 133(??): IFC, April 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000302.

Anonymous:2015:CEBd

[Ano15d] Anonymous. COV2: Editorial Board. Computer Vision and Image Understanding: CVIU, 134(??): ifc, May 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000570.

Anonymous:2015:CEBe

[Ano15e] Anonymous. COV2: Editorial Board. Computer Vision and Image Understanding: CVIU, 135(??): ifc, June 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000740.

Anonymous:2015:CEBf

[Ano15f] Anonymous. COV2: Editorial Board. Computer Vision and Image Understanding: CVIU, 136(??): ifc, July 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000995.

Anonymous:2015:EBa

[Ano15g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 130(??):IFC, January 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214002148.

Anonymous:2015:EBb

[Ano15h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 137(??): ifc, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001101.

Anonymous:2015:EBc

[Ano15i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 138(??): ifc, September 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001435.

Anonymous:2015:EBd

[Ano15j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 139(??): ifc, October 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001678.

Anonymous:2015:EBe

[Ano15k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 140(??): ifc, November 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001903.

Anonymous:2015:EBf

[Ano15l]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 141(??):ifc, December 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002234.

Anonymous:2015:EBg

[Ano15m]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 142(??): ifc, ???? 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002398.

Anonymous:2015:EAD

[Ano15n] Anonymous. Editorial of “Advances in Discrete Geometry for Computer Imagery”. Computer Vision and Image Understanding: CVIU, 138(??): I, September 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001526.

Anonymous:2015:ESI

[Ano15o] Anonymous. Editorial of special issue on shape representations meet visual recognition. Computer Vision and Image Understanding: CVIU, 139(??): 88, October 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001654.

Anonymous:2015:RA

[Ano15p] Anonymous. Reviewer acknowledgement. Computer Vision and Image Understanding: CVIU, 130(??): i–v, January 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214002197.

Anonymous:2016:EBa

[Ano16a]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 143(??): ifc, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000047.

Anonymous:2016:EBb

[Ano16b]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 144(??): ifc, March 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216000229.

Anonymous:2016:EBc

[Ano16c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 145(??): ifc, April 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216000588.

Anonymous:2016:EBd

[Ano16d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 146(??): ifc, May 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216300133.

Anonymous:2016:EBe

[Ano16e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 147(??): ifc, June 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421630039X.

Anonymous:2016:EBf

[Ano16f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 148(??): ifc, July 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216300595.

Anonymous:2016:EBg

[Ano16g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 149(??): ifc, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300698.

Anonymous:2016:EBh

[Ano16h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 150(??): ifc, September 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300893.

Anonymous:2016:EBi

[Ano16i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 151(??): ifc, October 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301278.

Anonymous:2016:EBj

[Ano16j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 152(??): ifc, November 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301497.

Anonymous:2016:EBk

[Ano16k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 153(??):ifc, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421630176X.

Anonymous:2017:EBl

[Ano17a]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 154(??): ifc, January 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301849.

Anonymous:2017:EBm

[Ano17b]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 155(??): ifc, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300073.

Anonymous:2017:EBn

[Ano17c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 156(??): ifc, March 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217300176.

Anonymous:2017:EBo

[Ano17d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 157(??): ifc, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421730036X.

Anonymous:2017:EBa

[Ano17e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 158(??): ifc, May 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300498.

Anonymous:2017:EBb

[Ano17f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 159(??): ifc, June 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300851.

Anonymous:2017:EBc

[Ano17g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 160(??): ifc, July 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421730098X.

Anonymous:2017:EBd

[Ano17h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 161(??): ifc, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421730125X.

Anonymous:2017:EBe

[Ano17i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 162(??): ifc, September 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301510.

Anonymous:2017:IFCa

[Ano17j] Anonymous. Inside front cover — editorial board page/cover image legend if applicable. Computer Vision and Image Understanding: CVIU, 163(??): ifc, October 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301807.

Anonymous:2017:IFCb

[Ano17k] Anonymous. Inside front cover — editorial board page/cover image legend if applicable. Computer Vision and Image Understanding: CVIU, 164(??): ifc, November 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301911.

Anonymous:2017:IFCc

[Ano17l] Anonymous. Inside front cover — Editorial Board page/cover image legend if applicable. Computer Vision and Image Understanding: CVIU, 165(??): ifc, December 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217302084.

Anonymous:2018:EBa

[Ano18a]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 167(??): ii, February 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300213.

Anonymous:2018:EBb

[Ano18b]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 168(??): ii, ???? 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218300316.

Anonymous:2018:EBc

[Ano18c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 169(??): ii, April 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218300456.

Anonymous:2018:EBd

[Ano18d]  Anonymous.  Editorial

board. Computer Vision and Image Understanding: CVIU, 170(??): ii, May 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218300614.

Anonymous:2018:EBe

[Ano18e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 171(??): ii, June 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218300985.

Anonymous:2018:EBf

[Ano18f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 172(??): ii, July 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218301048.

Anonymous:2018:EBg

[Ano18g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 173(??): ii, August 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218302200.

Anonymous:2018:EBh

[Ano18h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 174(??): ii, September 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421830434X.

Anonymous:2018:EBi

[Ano18i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 175(??): ii, October 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218304405.

Anonymous:2018:EBj

[Ano18j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 176–177(??):ii, November/December 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314218304557.

Anonymous:2018:IFC

[Ano18k] Anonymous. Inside front cover — Editorial Board page/cover image legend if applicable. Computer Vision and Image Understanding: CVIU, 166(??): ifc, January 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217302151.

Anonymous:2019:D

[Ano19a]  Anonymous.  December

2019. Computer Vision and Image Understanding: CVIU, 189(??):??, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2019:EBa

[Ano19b]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 178(??): ii, January 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218304612.

Anonymous:2019:EBb

[Ano19c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 179(??): ii, February 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219300098.

Anonymous:2019:EBc

[Ano19d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 180(??): ii, March 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219300256.

Anonymous:2019:EBd

[Ano19e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 181(??): ii, April 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219300360.

Anonymous:2019:EBe

[Ano19f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 182(??): ii, May 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219300499.

Anonymous:2019:EBf

[Ano19g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 183(??): ii, June 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219300700.

Anonymous:2019:EBg

[Ano19h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 184(??): ii, July 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S1077314219300827.

Anonymous:2019:EBh

[Ano19i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 185(??): ii, August 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S1077314219300888.

Anonymous:2019:EBi

[Ano19j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 186(??):ii, September 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S1077314219301110.

Anonymous:2019:EBj

[Ano19k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 187(??):Article 102813, October 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1077314219301249.

Anonymous:2019:EBk

[Ano19l]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 188(??):Article 102833, November 2019.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314219301377.

Anonymous:2019:EBl

[Ano19m]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 189(??):Article 102858, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314219301523.

Anonymous:2019:N

[Ano19n]  Anonymous.  November

2019. Computer Vision and Image Understanding: CVIU, 188(??):??, November 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:Aa

[Ano20a] Anonymous. April 2020. Computer Vision and Image Understanding: CVIU, 193 (??):??, April 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:Ab

[Ano20b] Anonymous. August 2020. Computer Vision and Image Understanding: CVIU, 197– 198(??):??, August 2020.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:D

[Ano20c]  Anonymous.  December

2020. Computer Vision and Image Understanding: CVIU, 201(??):??, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:EBa

[Ano20d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 190(??):Article 102870, January 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219301614.

Anonymous:2020:EBb

[Ano20e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 191(??):Article 102891, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219301730.

Anonymous:2020:EBc

[Ano20f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 192(??):Article 102911, March 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300072.

Anonymous:2020:EBd

[Ano20g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 193(??):Article 102939, April 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300266.

Anonymous:2020:EBe

[Ano20h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 194(??):Article 102957, May 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300394.

Anonymous:2020:EBf

[Ano20i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 195(??):Article 102975, June 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300527.

Anonymous:2020:EBg

[Ano20j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 196(??):Article 102997, July 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300679.

Anonymous:2020:EBh

[Ano20k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 103054, August 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300990.

Anonymous:2020:EBi

[Ano20l]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 199(??):Article 103091, October 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301223.

Anonymous:2020:EBj

[Ano20m]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 200(??):Article 103117, November 2020.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314220301387.

Anonymous:2020:EBk

[Ano20n]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 201(??):Article 103144, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314220301636.

Anonymous:2020:F

[Ano20o] Anonymous. February 2020. Computer Vision and Image Understanding: CVIU, 191(??):??, February 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:Ja

[Ano20p] Anonymous. January 2020. Computer Vision and Image Understanding: CVIU, 190(??):??, January 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:Jc

[Ano20q] Anonymous. July 2020. Computer Vision and Image Understanding: CVIU, 196 (??):??, July 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:Jb

[Ano20r] Anonymous. June 2020.

Computer Vision and Image Understanding: CVIU, 195 (??):??, June 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:Ma

[Ano20s] Anonymous. March 2020.

Computer Vision and Image Understanding: CVIU, 192 (??):??, March 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:Mb

[Ano20t] Anonymous. May 2020.

Computer Vision and Image Understanding: CVIU, 194 (??):??, May 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:N

[Ano20u]  Anonymous.  November

2020. Computer Vision and Image Understanding: CVIU, 200(??):??, November 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2020:O

[Ano20v] Anonymous. October 2020.

Computer Vision and Image Understanding: CVIU, 199(??):??, October 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:A

[Ano21a] Anonymous. April 2021. Computer Vision and Image Understanding: CVIU, 205 (??):??, April 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:D

[Ano21b]  Anonymous.  December

2021. Computer Vision and Image Understanding: CVIU, 213(??):??, December 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:EBa

[Ano21c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 202(??):Article 103151, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301703.

Anonymous:2021:EBb

[Ano21d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 203(??):Article 103160, February 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000047.

Anonymous:2021:EBc

[Ano21e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 204(??):Article 103177, March 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000217.

Anonymous:2021:EBd

[Ano21f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 205(??):Article 103192, April 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000369.

Anonymous:2021:EBe

[Ano21g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 206(??):Article 103198, May 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000424.

Anonymous:2021:EBf

[Ano21h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 207(??): ??, June 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314221000576.

Anonymous:2021:EBg

[Ano21i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 208–209(??): ??, July 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314221000783.

Anonymous:2021:EBh

[Ano21j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 208–209(??): ??, July 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314221000783.

Anonymous:2021:EBi

[Ano21k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001090.

Anonymous:2021:EBj

[Ano21l]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 211(??): ??, October 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001272.

Anonymous:2021:EBk

[Ano21m]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 212(??): ??, November 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001351.

Anonymous:2021:EBl

[Ano21n]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 213(??):??, December 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221001533.

Anonymous:2021:F

[Ano21o] Anonymous. February 2021. Computer Vision and Image Understanding: CVIU, 203(??):??, February 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:Ja

[Ano21p] Anonymous. January 2021. Computer Vision and Image Understanding: CVIU, 202(??):??, January 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:Jc

[Ano21q] Anonymous. July 2021. Computer Vision and Image Understanding: CVIU, 208–209(??):??, July 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:Jd

[Ano21r] Anonymous. July 2021. Computer Vision and Image Understanding: CVIU, 208–209(??):??, July 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:Jb

[Ano21s] Anonymous. June 2021. Computer Vision and Image Understanding: CVIU, 207 (??):??, June 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:Ma

[Ano21t] Anonymous. March 2021.

Computer Vision and Image Understanding: CVIU, 204 (??):??, March 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:Mb

[Ano21u] Anonymous. May 2021.

Computer Vision and Image Understanding: CVIU, 206 (??):??, May 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:N

[Ano21v]  Anonymous.  November

2021. Computer Vision and Image Understanding: CVIU, 212(??):??, November 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:O

[Ano21w] Anonymous. October 2021.

Computer Vision and Image Understanding: CVIU, 211(??):??, October 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2021:S

[Ano21x]  Anonymous.  September

2021. Computer Vision and Image Understanding:

CVIU, 210(??):??, September 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Computer Vision and Image Understanding: CVIU, 218 (??):??, April 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. Computer Vision and Image Understanding: CVIU, 221 (??):??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:D

[Ano22c]  Anonymous.  December

2022. Computer Vision and Image Understanding: CVIU, 225(??):??, December 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:EBa

[Ano22d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 214(??): ??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001636.

Anonymous:2022:EBb

[Ano22e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 215(??): ??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000066.

Anonymous:2022:EBc

[Ano22f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 216(??): ??, February 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000182.

Anonymous:2022:EBd

[Ano22g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 217(??): ??, March 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000273.

Anonymous:2022:EBe

[Ano22h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 218(??): ??, April 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000479.

Anonymous:2022:EBf

[Ano22i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 219(??): ??, June 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000698.

Anonymous:2022:EBg

[Ano22j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 220(??): ??, July 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000777.

Anonymous:2022:EBh

[Ano22k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 221(??): ??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000844.

Anonymous:2022:EBi

[Ano22l]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 223(??): ??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001436.

Anonymous:2022:EBj

[Ano22m]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 224(??): ??, November 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001539.

Anonymous:2022:EBk

[Ano22n]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 225(??):??, December 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222001709.

Anonymous:2022:F

[Ano22o] Anonymous. February 2022.

Computer Vision and Image Understanding: CVIU, 216(??):??, February 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:Ja

[Ano22p] Anonymous. January 2022. Computer Vision and Image Understanding: CVIU, 214(??):??, January 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:Jb

[Ano22q] Anonymous. January 2022. Computer Vision and Image Understanding: CVIU, 215(??):??, January 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:Jd

[Ano22r] Anonymous. July 2022. Computer Vision and Image Understanding: CVIU, 220 (??):??, July 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:Jc

[Ano22s] Anonymous. June 2022. Computer Vision and Image Understanding: CVIU, 219 (??):??, June 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:Ma

[Ano22t] Anonymous. March 2022. Computer Vision and Image Understanding: CVIU, 217 (??):??, March 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:N

[Ano22u]  Anonymous.  November

2022. Computer Vision and Image Understanding: CVIU, 224(??):??, November 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2022:O

[Ano22v] Anonymous. October 2022.

Computer Vision and Image Understanding: CVIU, 223(??):??, October 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023.

Computer Vision and Image Understanding: CVIU, 230 (??):??, April 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023.

Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:D

[Ano23c]  Anonymous.  December

2023. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:EBa

[Ano23d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 226(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001850.

Anonymous:2023:EBb

[Ano23e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 227(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000073.

Anonymous:2023:EBc

[Ano23f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 228(??): ??, February 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422300019X.

Anonymous:2023:EBd

[Ano23g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 229(??): ??, March 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000486.

Anonymous:2023:EBe

[Ano23h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 230(??): ??, April 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000553.

Anonymous:2023:EBf

[Ano23i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 231(??): ??, June 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000796.

Anonymous:2023:EBg

[Ano23j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 232(??): ??, July 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223001108.

Anonymous:2023:EBh

[Ano23k]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 233(??): ??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001352.

Anonymous:2023:EBi

[Ano23l]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001728.

Anonymous:2023:EBj

[Ano23m]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002059.

Anonymous:2023:EBk

[Ano23n]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002278.

Anonymous:2023:EBl

[Ano23o]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002503.

Anonymous:2023:F

[Ano23p] Anonymous. February 2023.

Computer Vision and Image Understanding: CVIU, 228(??):??, February 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:Ja

[Ano23q] Anonymous. January 2023.

Computer Vision and Image Understanding: CVIU, 226(??):??, January 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:Jb

[Ano23r] Anonymous. January 2023. Computer Vision and Image Understanding: CVIU, 227(??):??, January 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:Jd

[Ano23s] Anonymous. July 2023. Computer Vision and Image Understanding: CVIU, 232 (??):??, July 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:Jc

[Ano23t] Anonymous. June 2023. Computer Vision and Image Understanding: CVIU, 231 (??):??, June 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:M

[Ano23u] Anonymous. March 2023. Computer Vision and Image Understanding: CVIU, 229 (??):??, March 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:N

[Ano23v]  Anonymous.  November

2023. Computer Vision and Image Understanding: CVIU, 236(??):??, November 2023. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:O

[Ano23w] Anonymous. October 2023.

Computer Vision and Image Understanding: CVIU, 235(??):??, October 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2023:S

[Ano23x]  Anonymous.  September

2023. Computer Vision and Image Understanding: CVIU, 234(??):??, September 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024.

Computer Vision and Image Understanding: CVIU, 241 (??):??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. August 2024.

Computer Vision and Image Understanding: CVIU, 245 (??):??, August 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2024:EBa

[Ano24c]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002758.

Anonymous:2024:EBb

[Ano24d]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 239(??): ??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000043.

Anonymous:2024:EBc

[Ano24e]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000444.

Anonymous:2024:EBd

[Ano24f]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000699.

Anonymous:2024:EBe

[Ano24g]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 242(??): ??, May 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000857.

Anonymous:2024:EBf

[Ano24h]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 243(??): ??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224001012.

Anonymous:2024:EBg

[Ano24i]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 244(??): ??, July 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224001206.

Anonymous:2024:EBh

[Ano24j]  Anonymous.  Editorial

Board. Computer Vision and Image Understanding: CVIU, 245(??): ??, August 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422400136X.

Anonymous:2024:F

[Ano24k] Anonymous. February 2024. Computer Vision and Image Understanding: CVIU, 239(??):??, February 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2024:Ja

[Ano24l] Anonymous. January 2024. Computer Vision and Image Understanding: CVIU, 238(??):??, January 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2024:Jc

[Ano24m] Anonymous. July 2024. Computer Vision and Image Understanding: CVIU, 244 (??):??, July 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2024:Jb

[Ano24n] Anonymous. June 2024. Computer Vision and Image Understanding: CVIU, 243 (??):??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2024:Ma

[Ano24o] Anonymous. March 2024.

Computer Vision and Image Understanding: CVIU, 240 (??):??, March 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2024:Mb

[Ano24p] Anonymous. May 2024.

Computer Vision and Image Understanding: CVIU, 242 (??):??, May 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Antonacopoulos:1998:PSU

[Ant98] Apostolos Antonacopoulos.

Page segmentation using the description of the background. Computer Vision and Image Understanding: CVIU, 70(3): 350–369, June 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0691/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0691/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0691/production/ref.

Arentz:2004:COS

[AO04] Will Archer Arentz and Bjørn Olstad. Classifying offensive sites based on image content. Computer Vision and Image Understanding: CVIU, 94(1–3):295– 310, April/June 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Akashi:2016:SRC

[AO16] Yasushi Akashi and Takayuki Okatani. Separation of reflection components by sparse non-negative matrix factorization. Computer Vision and Image Understanding: CVIU, 146(??): 77–85, May 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001873.

Argyriou:2010:PSA

[APB10] Vasileios Argyriou, Maria Petrou, and Svetlana Barsky. Photometric stereo with an arbitrary number of illuminants. Computer Vision and Image Understanding: CVIU, 114(8):887–900, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Androutsos:1999:NVB

[APV99] D. Androutsos, K. N. Plataniotis, and A. N. Venetsanopoulos. A novel vectorbased approach to color image retrieval using a vector angular-based distance measure. Computer Vision and Image Understanding: CVIU, 75(1–2):46–58, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0767/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0767/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0767/production/ref.

Alata:2009:TBC

[AQ09] Olivier Alata and Ludovic Quintard. Is there a best color space for color image characterization or representation based on Multivariate Gaussian Mixture Model? Computer Vision and Image Understanding: CVIU, 113(8):867–877, August 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ayala-Raggi:2011:AFI

[ARARCE11] Salvador E. Ayala-Raggi, Leopoldo Altamirano-Robles, and Janeth Cruz-Enriquez. Automatic face interpretation using fast 3 D illumination-based AAM models. Computer Vision and Image Understanding: CVIU, 115(2):194–210, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Adcock:2014:CHL

[ARC14] Aaron Adcock, Daniel Rubin, and Gunnar Carlsson. Classification of hepatic lesions using the matching metric. Computer Vision and Image Understanding: CVIU, 121(??):36– 42, April 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213002221.

Arrigoni:2018:RSS

[ARFF18] Federica Arrigoni, Beatrice Rossi, Pasqualina Fragneto, and Andrea Fusiello. Robust synchronization in SO(3) and SE(3) via lowrank and sparse matrix decomposition. Computer Vision and Image Understanding: CVIU, 174(??):95–113, September 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218301528.

Alexiadis:2008:NDS

[AS08a] Dimitrios S. Alexiadis and George D. Sergiadis. Narrow directional steerable filters in motion estimation. Computer Vision and Image Understanding: CVIU, 110(2):192–211, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Aouadi:2008:APR

[AS08b] Souha Aouadi and Laurent Sarry. Accurate and precise 2D–3D registration based on X-ray intensity. Computer Vision and Image Understanding: CVIU, 110(1): 134–151, April 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Alexiadis:2009:MES

[AS09] Dimitrios S. Alexiadis and George D. Sergiadis. Motion estimation, segmentation and separation, using hypercomplex phase correlation, clustering techniques and graph-based optimization. Computer Vision and Image Understanding: CVIU, 113(2):212–234, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Asad:2017:SSP

[AS17a] Muhammad Asad and Greg Slabaugh. SPORE: Staged probabilistic regression for hand orientation inference. Computer Vision and Image Understanding: CVIU, 161(??):114– 129, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300942.

Aastrom:2017:GAC

[ÅS17b] Freddie Åström and Christoph Schnörr. A geometric approach for color image regularization. Computer Vision and Image Understanding: CVIU, 165(??):43–59, December 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217301765.

Awan:2023:WSM

[AS23] Mehwish Awan and Jitae Shin. Weakly supervised multi-class semantic video segmentation for road scenes. Computer Vision and Image Understanding: CVIU, 230(??): ??, April 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223000449.

Alletto:2017:VRE

[ASC17] Stefano Alletto, Giuseppe Serra, and Rita Cucchiara. Video registration in egocentric vision under day and night illumination changes. Computer Vision and Image Understanding: CVIU, 157(??): 274–283, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421630145X.

Ashraf:2013:VIA

[ASCF13] Nazim Ashraf, Yuping Shen, Xiaochun Cao, and Hassan Foroosh. View invariant action recognition using weighted fundamental ratios. Computer Vision and Image Understanding: CVIU, 117(6):587– 602, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000131.

Ashraf:2014:VIA

[ASF14] Nazim Ashraf, Chuan Sun, and Hassan Foroosh. View invariant action recognition using projective depth. Computer Vision and Image Understanding: CVIU, 123 (??):41–52, June 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000563.

Audette:2003:IRS

[ASFP03] Michel A. Audette, Kaleem Siddiqi, Frank P. Ferrie, and Terry M. Peters. An integrated range-sensing, segmentation and registration framework for the characterization of intrasurgical brain deformations in image-guided surgery. Computer Vision and Image Understanding: CVIU, 89(2–3):226–251, February/ March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Agrawal:1997:PTP

[ASS97] R. C. Agrawal, S. C. Sahasrabudhe, and R. K. Shevgaonkar. Preservation of topological properties of a simple closed curve under digitalization. Computer Vision and Image Understanding: CVIU, 67(2): 99–111, August 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0514/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0514/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0514/production/ref.

Aastrom:1997:GVS

[Åst97] Kalle Åström. The geometry of visual space: About the incompatibility between science and mathematics — reply. Computer Vision and Image Understanding: CVIU, 65(3): 436–438, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0494/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0494/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0494/production/ref. See

[PRW97a].

Alvarez:2012:LDC

[ASVO12] Susana Alvarez, Anna Salvatella, Maria Vanrell, and Xavier Otazu. Lowdimensional and comprehensive color texture description. Computer Vision and Image Understanding: CVIU, 116(1):54–67, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421100186X.

August:1999:CFG

[ASZ99a] Jonas August, Kaleem Siddiqi, and Steven W. Zucker. Contour fragment grouping and shared, simple occluders. Computer Vision and Image Understanding: CVIU, 76(2):146– 162, November 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0795/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0795/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0795/production/ref.

August:1999:LIP

[ASZ99b] Jonas August, Kaleem Siddiqi, and Steven W. Zucker. Ligature instabilities in the perceptual organization of shape. Computer Vision and Image Understanding: CVIU, 76(3):231–243, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0802/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0802/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0802/production/ref.

Andreopoulos:2013:YOR

[AT13] Alexander Andreopoulos and John K. Tsotsos. 50 years of object recognition: Directions forward. Computer Vision and Image Understanding: CVIU, 117(8): 827–891, August 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421300091X.

Argyriou:2017:FDS

[AT17] Vasileios Argyriou and Georgios Tzimiropoulos. Frequency domain subpixel registration using HOG phase correlation. Computer Vision and Image Understanding: CVIU, 155(??): 70–82, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301722.

Avila:2013:PIR

[ATC^{+}13] Sandra Avila, Nicolas Thome, Matthieu Cord, Eduardo Valle, and Arnaldo de A. Araújo. Pooling in image representation: the visual codeword point of view. Computer Vision and Image Understanding: CVIU, 117(5): 453–465, May 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001737.

Andalo:2015:EHM

[ATG15] Fernanda A. Andaló, Gabriel Taubin, and Siome Goldenstein. Efficient height measurements in single images based on the detection of vanishing points. Computer Vision and Image Understanding: CVIU, 138(??): 51–60, September 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215000855.

Atkinson:2017:PPS

[Atk17] Gary A. Atkinson. Polarisation photometric stereo. Computer Vision and Image Understanding: CVIU, 160(??):158–167, July 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314217300760.

Alahi:2010:CDD

[AVBK10] Alexandre Alahi, Pierre Vandergheynst, Michel Bierlaire, and Murat Kunt. Cascade of descriptors to detect and track objects across any network of cameras. Computer Vision and Image Understanding: CVIU, 114(6): 624–640, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ariz:2019:RAT

[AVC19] Mikel Ariz, Arantxa Villanueva, and Rafael Cabeza. Robust and accurate 2Dtracking-based 3D positioning method: Application to head pose estimation. Computer Vision and Image Understanding: CVIU, 180(??):13– 22, March 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219300025.

Azaza:2018:CPS

[AvdWDM18] Aymen Azaza, Joost van de Weijer, Ali Douik, and Marc Masana. Context proposals for saliency detection. Computer Vision and Image Understanding: CVIU, 174(??):1–11, September 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218300778.

Alba:2015:PCS

[AVGASAP15] Alfonso Alba, J. Flavio Vigueras-Gomez, Edgar R. Arce-Santana, and Ruth M. Aguilar-Ponce. Phase correlation with sub-pixel accuracy: a comparative study in 1D and 2D. Computer Vision and Image Understanding: CVIU, 137(??): 76–87, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000685.

Austin:1998:OLP

[AW98] W. J. Austin and A. M.

Wallace. Object location by parallel pose clustering. Computer Vision and Image Understanding: CVIU, 72(3):304–327, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0672/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0672/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0672/production/ref.

Allan:2009:OLU

[AW09] Moray Allan and Christopher K. I. Williams. Object localisation using the Generative Template of Features. Computer Vision and Image Understanding: CVIU, 113(7):824–838, July 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Aner-Wolf:2004:VSC

[AWK04] Aya Aner-Wolf and John R. Kender. Video summaries and cross-referencing through mosaic-based representation. Computer Vision and Image Understanding: CVIU, 95(2):201–237, August 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Alqahtani:2021:PCF

[AXJE21] Ali Alqahtani, Xianghua Xie, Mark W. Jones, and Ehab Essa. Pruning CNN filters via quantifying the importance of deep visual representations. Computer Vision and Image Understanding: CVIU, 208–209 (??):??, July 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000643.

Ahmad:2014:BSO

[AXSVL14] Aamir Ahmad, João Xavier, José Santos-Victor, and Pedro Lima. 3D to 2D bijection for spherical objects under equidistant fisheye projection. Computer Vision and Image Understanding: CVIU, 125(??):172– 183, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000824.

Aditya:2018:IUU

[AYB^{+}18] Somak Aditya, Yezhou Yang, Chitta Baral, Yiannis Aloimonos, and Cornelia Fermüller. Image understanding using vision and reasoning through scene description graph. Computer Vision and Image Understanding: CVIU, 173(??):

33–45, August 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217302291.

Akl:2018:SEB

[AYD^{+}18] Adib Akl, Charles Yaacoub, Marc Donias, Jean-Pierre Da Costa, and Christian Germain. A survey of exemplar-based texture synthesis methods. Computer Vision and Image Understanding: CVIU, 172(??): 12–24, July 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218300523.

Ates:2023:DLB

[AYG23] Hasan F. Ates, Suleyman Yildirim, and Bahadir K. Gunturk. Deep learningbased blind image superresolution with iterative kernel reconstruction and noise estimation. Computer Vision and Image Understanding: CVIU, 233(??): ??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422300098X.

Aytar:2015:PLT

[AZ15] Yusuf Aytar and Andrew Zisserman. Part level transfer regularization for enhancing exemplar SVMs. Computer Vision and Image Understanding: CVIU, 138(??):114– 123, September 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000880.

Arbab-Zavar:2011:GMB

[AZN11] Banafshe Arbab-Zavar and Mark S. Nixon. On guided model-based analysis for ear biometrics. Computer Vision and Image Understanding: CVIU, 115(4):487– 502, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Argyriou:2014:OID

[AZP14] Vasileios Argyriou, Stefanos Zafeiriou, and Maria Petrou. Optimal illumination directions for faces and rough surfaces for single and multiple light imaging using class-specific prior knowledge. Computer Vision and Image Understanding: CVIU, 125(??):16– 36, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000575.

Amir:2005:ESE

[AZSVK05] Arnon Amir, Lior Zimet, Alberto Sangiovanni-Vincentelli, and Sean Kao. An embedded system for an eyedetection sensor. Computer Vision and Image Understanding: CVIU, 98(1):104– 123, April 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Black:1996:REM

[BA96] Michael J. Black and P. Anandan. The robust estimation of multiple motions: Parametric and piecewise-smooth flow fields. Computer Vision and Image Understanding: CVIU, 63(1):75– 104, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0006/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0006/production/pdf.

Barreto:2006:FCP

[BA06] João P. Barreto and Helder Araujo. Fitting conics to paracatadioptric projections of lines. Computer Vision and Image Understanding: CVIU, 101(3):151– 165, March 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Biasutti:2019:DIR

[BABB19] Pierre Biasutti, Jean-François Aujol, Mathieu Brédif, and Aurélie Bugeau. Diffusion and inpainting of reflectance and height LiDAR orthoimages. Computer Vision and Image Understanding: CVIU, 179(??):31–40, February 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218304302.

Bloom:2016:HTL

[BAM16] Victoria Bloom, Vasileios Argyriou, and Dimitrios Makris. Hierarchical transfer learning for online recognition of compound actions. Computer Vision and Image Understanding: CVIU, 144(??):62– 72, March 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002647.

Boulahia:2018:CCD

[BAMK18] Said Yacine Boulahia, Eric Anquetil, Franck Multon, and Richard Kulpa. CuDi3D: Curvilinear displacement based approach for online 3D action detection. Computer Vision and Image Understanding: CVIU, 174(??):57–69, September 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218301127.

Brooks:2008:ASM

[BAP08] Rupert Brooks, Tal Arbel, and Doina Precup. Anytime similarity measures for faster alignment. Computer Vision and Image Understanding: CVIU, 110(3): 378–389, June 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ba:2016:LVB

[BAPXH16] Sileye Ba, Xavier Alameda-Pineda, Alessio Xompero, and Radu Horaud. An on-line variational Bayesian model for multi-person tracking from cluttered scenes. Computer Vision and Image Understanding: CVIU, 153(??):64–76, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216301047.

Bartoli:2005:GDS

[Bar05] Adrien Bartoli. The geometry of dynamic scenes— on coplanar and convergent linear motions embedded in 3D static scenes. Computer Vision and Image Understanding: CVIU, 98(2):223– 238, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barreto:2006:UGR

[Bar06] João P. Barreto. A unifying geometric representation for central projection systems. Computer Vision and Image Understanding: CVIU, 103(3):208–217, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bartoli:2007:RSS

[Bar07] Adrien Bartoli. A random sampling strategy for piecewise planar scene segmentation. Computer Vision and Image Understanding: CVIU, 105(1):42–59, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barath:2018:EEB

[Bar18] Dániel Baráth. Efficient energy-based topological outlier rejection. Computer Vision and Image Understanding: CVIU, 174(??): 70–81, September 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218301152.

Barbero-Alvarez:2023:MEA

[BÁRM23] Miguel Antonio Barbero-Álvarez, Juan Antonio Rodrigo, and José Manuel Menéndez. Minimum error adaptive RGB calibration in a context of colorimetric uncertainty for cultural heritage preservation. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002151.

Bogoni:1995:IRR

[BB95] Luca Bogoni and Ruzena Bajcsy. Interactive recognition and representation of functionality. Computer Vision and Image Understanding: CVIU, 62(2):194– 214, September 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1050/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1050/production/pdf.

Bergevin:2003:OLS

[BB03] R. Bergevin and A. Bubel. Object-level structured contour map extraction. Computer Vision and Image Understanding: CVIU, 91(3): 302–334, September 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bergevin:2004:DCJ

[BB04] R. Bergevin and A. Bubel.

Detection and characterization of junctions in a 2D image. Computer Vision and Image Understanding: CVIU, 93(3):288– 309, March 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bloch:2013:MMH

[BB13] Isabelle Bloch and Alain Bretto. Mathematical morphology on hypergraphs, application to similarity and positive kernel. Computer Vision and Image Understanding: CVIU, 117(4): 342–354, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001828.

See corrigendum [BB15a].

Bloch:2015:CMM

[BB15a] Isabelle Bloch and Alain Bretto. Corrigendum to “Mathematical morphology on hypergraphs, application to similarity and positive kernel” [Comput. Vis.

Image Understanding 117 (2013) 342–354]. Computer Vision and Image Understanding: CVIU, 132(??): 1–2, March 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002410.

See [BB13].

Bouachir:2015:CPB

[BB15b] Wassim Bouachir and Guillaume-Alexandre Bilodeau. Collaborative part-based tracking using salient local predictors. Computer Vision and Image Understanding: CVIU, 137(??):88– 101, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000673.

Baldacci:2016:OBG

[BB16] Fabien Baldacci and Achille Braquelaire. Oriented boundary graph: an efficient structuring model for segmentation of 3D images. Computer Vision and Image Understanding: CVIU, 143(??):92–103, February 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002131.

Bouzid:2024:SMH

[BB24] Hamza Bouzid and Lahoucine Ballihi. SpATr: MoCap 3D human action recognition based on spiral auto-encoder and transformer network. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000559.

Bertin:1996:VPC

[BBB96] Etienne Bertin, Horst Bischof, and Pascal Bertolino. Voronoi pyramids controlled by Hopfield neural networks. Computer Vision and Image Understanding: CVIU, 63(3): 462–475, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0035/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0035/production/pdf.

Baldoni:2000:UIC

[BBC00] Matteo Baldoni, Cristina Baroglio, and Davide Cavagnino. Use of IFS codes for learning 2D isolated-object classification systems. Computer Vision and Image Understanding: CVIU, 77 (3):371–387, March 2000.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0823;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0823/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0823/ref.

Bertozzi:2007:PDM

[BBC^{+}07] M. Bertozzi, A. Broggi, C. Caraffi, M. Del Rose, M. Felisa, and G. Vezzoni. Pedestrian detection by means of far-infrared stereo vision. Computer Vision and Image Understanding: CVIU, 106(2–3): 194–204, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Besedin:2020:DOC

[BBCF20] Andrey Besedin, Pierre Blanchart, Michel Crucianu, and Marin Ferecatu. Deep online classification using pseudo-generative models. Computer Vision and Image Understanding: CVIU, 201(??):Article 103048, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220300941.

Bellotto:2012:CVT

[BBH^{+}12] Nicola Bellotto, Ben Benfold, Hanno Harland, Hans-Hellmut Nagel, Nicola Pirlo, Ian Reid, Eric Sommerlade, and Chuan Zhao. Cognitive visual tracking and camera control. Computer Vision and Image Understanding: CVIU, 116(3): 457–471, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421100213X.

Brunet:2014:MTB

[BBH14] F. Brunet, A. Bartoli, and R. I. Hartley. Monocular template-based 3D surface reconstruction: Convex inextensible and nonconvex isometric methods. Computer Vision and Image Understanding: CVIU, 125(??):138– 154, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000812.

Bogdanova:2010:DVA

[BBHF10] Iva Bogdanova, Alexandre Bur, Heinz Hügli, and Pierre-André Farine. Dynamic visual attention on the sphere. Computer Vision and Image Understanding: CVIU, 114(1):100–110, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bjorkman:2014:DST

[BBK14] Mårten Björkman, Niklas Bergström, and Danica Kragic. Detecting, segmenting and tracking unknown objects using multi-label MRF inference. Computer Vision and Image Understanding: CVIU, 118(??): 111–127, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421300194X.

Bringier:2015:ETH

[BBK15] Benjamin Bringier, Alexandre Bony, and Majdi Khoudeir. Evidence theory for high dynamic range reconstruction with linear digital cameras. Computer Vision and Image Understanding: CVIU, 133(??): 90–101, April 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421400188X.

Ballan:2015:DDA

[BBSD15] Lamberto Ballan, Marco Bertini, Giuseppe Serra, and Alberto Del Bimbo. A data-driven approach for tag refinement and localization in web videos. Computer Vision and Image Understanding: CVIU, 140(??): 58–67, November 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001204.

Benoit:2010:FBI

[BC10] A. Benoit and A. Caplier.

Fusing bio-inspired vision data for simplified high level scene interpretation: Application to face motion analysis. Computer Vision and Image Understanding: CVIU, 114(7):774– 789, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bardinet:1998:PDM

[BCA98] Eric Bardinet, Laurent D.

Cohen, and Nicholas Ayache. A parametric deformable model to fit unstructured 3D data. Computer Vision and Image Understanding: CVIU, 71(1): 39–54, July 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0595/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0595/production/pdf;
http://www.idealibrary.

                                        com/links/artid/cviu.1997.
                                        0595/production/ref.

Blaiotta:2016:VIM

[BCA16] Claudia Blaiotta, M. Jorge Cardoso, and John Ashburner. Variational inference for medical image segmentation. Computer Vision and Image Understanding: CVIU, 151(??): 14–28, October 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300285.

Bianco:2016:CIC

[BCC16] Simone Bianco, Gianluigi Ciocca, and Claudio Cusano. CURL: Image classification using co-training and unsupervised representation learning. Computer Vision and Image Understanding: CVIU, 145(??): 15–29, April 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216000151.

Barbosa:2018:LBA

[BCC^{+}18] Igor Barros Barbosa, Marco Cristani, Barbara Caputo, Aleksander Rognhaugen, and Theoharis Theoharis. Looking beyond appearances: Synthetic training data for deep CNNs in reidentification. Computer Vision and Image Understanding: CVIU, 167(??): 50–62, February 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217302254.

Bertugli:2021:AVA

[BCC^{+}21] Alessia Bertugli, Simone Calderara, Pasquale Coscia, Lamberto Ballan, and Rita Cucchiara. AC-VRNN: Attentive conditional-VRNN for multi-future trajectory prediction. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000898.

Benoit:2010:UHV

[BCDH10] A. Benoit, A. Caplier, B. Durette, and J. Herault. Using Human Visual System modeling for bio-inspired low level image processing. Computer Vision and Image Understanding: CVIU, 114(7):758– 773, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bowyer:2006:SAC

[BCF06] Kevin W. Bowyer, Kyong Chang, and Patrick Flynn.

A survey of approaches and challenges in 3D and multi-modal 3D + 2D face recognition. Computer Vision and Image Understanding: CVIU, 101(1):1–15, January 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bongiovanni:1995:PSA

[BCG95] G. Bongiovanni, P. Crescenzi, and C. Guerra. Parallel simulated annealing for shape detection. Computer Vision and Image Understanding: CVIU, 61(1):60– 69, January 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1005/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1005/production/pdf.

Ballan:2018:GE

[BCH^{+}18] Lamberto Ballan, Shih-Fu Chang, Gang Hua, Thomas Mensink, Greg Mori, and Rahul Sukthankar. Guest editorial. Computer Vision and Image Understanding: CVIU, 173(??): 1, August 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421830465X.

Benkirane:2024:HAP

[BCHR24] Fatima Ezzahra Benkirane, Nathan Crombez, Vincent Hilaire, and Yassine Ruichek. Hybrid AI for panoptic segmentation: an informed deep learning approach with integration of prior spatial relationships knowledge. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002898.

Bermudez-Cameo:2018:FLP

[BCLNG18] Jesus Bermudez-Cameo, Gonzalo Lopez-Nicolas, and Jose J. Guerrero. Fitting line projections in noncentral catadioptric cameras with revolution symmetry. Computer Vision and Image Understanding: CVIU, 167(??):134– 152, February 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300031.

Bicego:2006:USA

[BCM06] Manuele Bicego, Marco Cristani, and Vittorio Murino. Unsupervised scene analysis: a hidden Markov model approach. Computer Vision and Image Understanding: CVIU, 102(1):22– 41, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bazzani:2013:SDA

[BCM13] Loris Bazzani, Marco Cristani, and Vittorio Murino. Symmetrydriven accumulation of local features for human characterization and reidentification. Computer Vision and Image Understanding: CVIU, 117(2):130–144, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314212001464.

Bernier:2009:FNB

[BCMCB09] Olivier Bernier, Pascal Cheung-Mon-Chan, and Arnaud Bouguet. Fast nonparametric belief propagation for real-time stereo articulated body tracking. Computer Vision and Image Understanding: CVIU, 113(1):29–47, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barata:2016:CIA

[BCMR16] Catarina Barata, M. Emre Celebi, Jorge S. Marques, and Jorge Rozeira. Clinically inspired analysis of dermoscopy images using a generative model. Computer Vision and Image Understanding: CVIU, 151(??): 124–137, October 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002064.

Bianco:2015:ITM

[BCNS15] Simone Bianco, Gianluigi Ciocca, Paolo Napoletano, and Raimondo Schettini. An interactive tool for manual, semi-automatic and automatic video annotation. Computer Vision and Image Understanding: CVIU, 131(??):88–99, February 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001544.

Bartoli:2015:MCA

[BCP15] Adrien Bartoli, Toby Collins, and Daniel Pizarro. Metric corrections of the affine camera. Computer Vision and Image Understanding: CVIU, 135(??):141– 156, June 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000478.

Borgefors:1996:ANP

[Bd96] Gunilla Borgefors and Gabriella Sanniti di Baja. Analyzing nonconvex 2D and 3D patterns. Computer Vision and Image Understanding: CVIU, 63(1):145– 157, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0010/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0010/production/pdf.

Buluswar:2002:CMO

[BD02] Shashi D. Buluswar and Bruce A. Draper. Color models for outdoor machine vision. Computer Vision and Image Understanding: CVIU, 85(2):71– 99, February 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Brancati:2017:HSD

[BDFG17] Nadia Brancati, Giuseppe De Pietro, Maria Frucci, and Luigi Gallo. Human skin detection through correlation rules between the YCb and YCr subspaces based on dynamic color clustering. Computer Vision and Image Understanding: CVIU, 155(??):33–42, February 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301989.

Bose:2009:CPT

[BDHM09] Prosenjit Bose, Vida Dujmović, Ferran Hurtado, and Pat Morin. Connectivitypreserving transformations of binary images. Computer Vision and Image Understanding: CVIU, 113(10): 1027–1038, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Briggs:2006:MSS

[BDL^{+}06] Amy J. Briggs, Carrick Detweiler, Yunpeng Li, Peter C. Mullen, and Daniel Scharstein. Matching scalespace features in 1D panoramas. Computer Vision and Image Understanding: CVIU, 103(3):184–195, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bertini:2012:MSR

[BDS12] Marco Bertini, Alberto Del Bimbo, and Lorenzo Seidenari. Multi-scale and realtime non-parametric approach for anomaly detection and localization. Computer Vision and Image Understanding: CVIU, 116(3): 320–329, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211002104.

Bavil:2023:ACH

[BDT23] Ali Farajzadeh Bavil, Hamed Damirchi, and Hamid D. Taghirad. Action capsules: Human skeleton action recognition. Computer Vision and Image Understanding: CVIU, 233(??): ??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001029.

Bazin:2010:MED

[BDVK10] J. C. Bazin, C. Demonceaux, P. Vasseur, and I. S. Kweon. Motion estimation by decoupling rotation and translation in catadioptric vision. Computer Vision and Image Understanding: CVIU, 114(2):254–273, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bonev:2013:ITS

[BEGB13] Boyan Bonev, Francisco Escolano, Daniela Giorgi, and Silvia Biasotti. Informationtheoretic selection of highdimensional spectral features for structural recognition. Computer Vision and Image Understanding: CVIU, 117(3):214– 228, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001919.

Benzi:2018:BIS

[BEK18] Marco Benzi, María-José Escobar, and Pierre Kornprobst. A bio-inspired synergistic virtual retina model for tone mapping. Computer Vision and Image Understanding: CVIU, 168(??): 21–36, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217302047.

Ben-Ezra:2000:RTM

[BEPW00] Moshe Ben-Ezra, Shmuel Peleg, and Michael Werman. Real-time motion analysis with linear programming. Computer Vision and Image Understanding: CVIU, 78(1):32–52, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0826;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0826/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0826/ref.

Bay:2008:SRF

[BETV08] Herbert Bay, Andreas Ess, Tinne Tuytelaars, and Luc Van Gool. Speeded-up Robust Features (SURF). Computer Vision and Image Understanding: CVIU, 110(3):346–359, June 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Breckon:2005:AVC

[BF05] Toby P. Breckon and Robert B. Fisher. Amodal volume completion: 3D visual completion. Computer Vision and Image Understanding: CVIU, 99(3):499– 526, September 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Benoit:2007:TDR

[BF07] Stephen Benoit and Frank P. Ferrie. Towards direct recovery of shape and motion parameters from image sequences. Computer Vision and Image Understanding: CVIU, 105(2):145–165, February 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Butakoff:2010:MVF

[BF10] Constantine Butakoff and Alejandro F. Frangi. Multiview face segmentation using fusion of statistical shape and appearance models. Computer Vision and Image Understanding: CVIU, 114(3):311– 321, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bentolila:2014:CEC

[BF14] Jacob Bentolila and Joseph M.

Francos. Conic epipolar constraints from affine correspondences. Computer Vision and Image Understanding: CVIU, 122(??): 105–114, May 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000307.

Bonnaerens:2022:APO

[BFD22] Maxim Bonnaerens, Matthias Freiberger, and Joni Dambre. Anchor pruning for object detection. Computer Vision and Image Understanding: CVIU, 221(??): ??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000601.

Barnard:1997:CCS

[BFF97] Kobus Barnard, Graham Finlayson, and Brian Funt. Color constancy for scenes with varying illumination. Computer Vision and Image Understanding: CVIU, 65 (2):311–321, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0567/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0567/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0567/production/ref.

Bian:2023:GLC

[BFMW23] Cunling Bian, Wei Feng, Fanbo Meng, and Song Wang. Global-local contrastive multiview representation learning for skeletonbased action recognition. Computer Vision and Image Understanding: CVIU, 229 (??):??, March 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000358.

Battaglino:2013:DTH

[BFR13] D. Battaglino, A. Frosini, and S. Rinaldi. A decomposition theorem for homogeneous sets with respect to diamond probes. Computer Vision and Image Understanding: CVIU, 117(4): 319–325, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001804.

Black:2000:REC

[BFY00] Michael J. Black, David J.

Fleet, and Yaser Yacoob. Robustly estimating changes in image appearance. Computer Vision and Image Understanding: CVIU, 78(1): 8–31, April 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0825;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0825/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0825/ref.

Burghouts:2009:PEL

[BG09] Gertjan J. Burghouts and Jan-Mark Geusebroek. Performance evaluation of local colour invariants. Computer Vision and Image Understanding: CVIU, 113(1):48– 62, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bozorgtabar:2016:EMT

[BG16] Behzad Bozorgtabar and Roland Goecke. Efficient multi-target tracking via discovering dense subgraphs. Computer Vision and Image Understanding: CVIU, 144(??):205– 216, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421500257X.

Busto:2018:VRE

[BG18] Pau Panareda Busto and Juergen Gall. Viewpoint refinement and estimation with adapted synthetic data. Computer Vision and Image Understanding: CVIU, 169(??): 75–89, April 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300055.

Bagheri:2017:LRG

[BGE^{+}17] Mohammad Ali Bagheri, Qigang Gao, Sergio Escalera, Thomas B. Moeslund, Huamin Ren, and Elham Etemad. Locality regularized group sparse coding for action recognition. Computer Vision and Image Understanding: CVIU, 158(??):106– 114, May 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300425.

Brechbuhler:1995:PCS

[BGK95] Ch. Brechbühler, G. Gerig, and O. Kübler. Parametrization of closed surfaces for 3-D shape description. Computer Vision and Image Understanding: CVIU, 61(2): 154–170, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1013/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1013/production/pdf.

Bober:1998:ARE

[BGK98] M. Bober, N. Georgis, and J. Kittler. On accurate and robust estimation of fundamental matrix. Computer Vision and Image Understanding: CVIU, 72(1):39– 53, October 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0670/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0670/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0670/production/ref.

Beveridge:2009:FIA

[BGPD09] J. Ross Beveridge, Geof H.

Givens, P. Jonathon Phillips, and Bruce A. Draper. Factors that influence algorithm performance in the Face Recognition Grand Challenge. Computer Vision and Image Understanding: CVIU, 113(6):750–762, June 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bebis:1998:IBA

[BGSdVL98] George Bebis, Michael Georgiopoulos, Mubarak Shah, and Niels da Vitoria Lobo. Indexing based on algebraic functions of views. Computer Vision and Image Understanding: CVIU, 72(3):360–378, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0679/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0679/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0679/production/ref.

Berthilsson:1999:RPO

[BH99] Rikard Berthilsson and Anders Heyden. Recognition of planar objects using the density of affine shape. Computer Vision and Image Understanding: CVIU, 76(2):135– 145, November 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0781/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0781/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0781/production/ref.

Bui:2012:ECB

[BH12] T. T. Quyen Bui and Keum-Shik Hong. Evaluating a color-based active basis model for object recognition. Computer Vision and Image Understanding: CVIU, 116(11):1111–1120, November 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421200104X.

Bhuiyan:2024:IIG

[BHA24] Amran Bhuiyan, Jimmy Xiangji Huang, and Aijun An. IGMG: Instance-guided multi-granularity for domain generalizable person re-identification. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002850.

Baker:2010:DIR

[BHBF10] Sarah E. Baker, Amanda Hentz, Kevin W. Bowyer, and Patrick J. Flynn. Degradation of iris recognition performance due to non-cosmetic prescription contact lenses. Computer Vision and Image Understanding: CVIU, 114(9): 1030–1044, September 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bowyer:2008:IUI

[BHF08] Kevin W. Bowyer, Karen Hollingsworth, and Patrick J. Flynn. Image understanding for iris biometrics: a survey. Computer Vision and Image Understanding: CVIU, 110(2):281–307, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bohme:2010:SCI

[BHMB10] Martin Böhme, Martin Haker, Thomas Martinetz, and Erhardt Barth. Shading constraint improves accuracy of time-of-flight measurements. Computer Vision and Image Understanding: CVIU, 114(12):1329– 1335, December 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ben-Hamadou:2013:FCS

[BHSD^{+}13] Achraf Ben-Hamadou, Charles Soussen, Christian Daul, Walter Blondel, and Didier Wolf. Flexible calibration of structured-light systems projecting point patterns. Computer Vision and Image Understanding: CVIU, 117(10):1468– 1481, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001069.

Banno:2010:OTB

[BI10] Atsuhiko Banno and Katsushi Ikeuchi. Omnidirectional texturing based on robust 3D registration through Euclidean reconstruction from two spherical images. Computer Vision and Image Understanding: CVIU, 114(4):491– 499, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Banno:2011:DMR

[BI11] Atsuhiko Banno and Katsushi Ikeuchi. Disparity map refinement and 3D surface smoothing via directed anisotropic diffusion. Computer Vision and Image Understanding: CVIU, 115(5): 611–619, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bichsel:1998:ASP

[Bic98] Martin Bichsel. Analyzing a scene’s picture set under varying lighting. Computer Vision and Image Understanding: CVIU, 71(3):271– 280, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0627/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0627/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0627/production/ref.

Bigun:1997:PRI

[Big97] Josef Bigün. Pattern recognition in images by symmetries and coordinate transformations. Computer Vision and Image Understanding: CVIU, 68(3):290– 307, December 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0556/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0556/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0556/production/ref.

Barbalau:2023:SRS

[BIG^{+}23] Antonio Barbalau, Radu Tudor Ionescu, Mariana-Iuliana Georgescu, Jacob Dueholm, Bharathkumar Ramachandra, Kamal Nasrollahi, Fahad Shahbaz Khan, Thomas B. Moeslund, and Mubarak Shah. SSMTL++: Revisiting selfsupervised multi-task learning for video anomaly detection. Computer Vision and Image Understanding: CVIU, 229(??): ??, March 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731422300036X.

Borza:2023:TSE

[BIMD23] Diana Laura Borza, Tudor Alexandru Ileni, Alexandru Ion Marinescu, and Sergiu Adrian Darabant. Teacher or supervisor? Effective online knowledge distillation via guided collaborative learning. Computer Vision and Image Understanding: CVIU, 228(??): ??, February 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000127.

Breen:1996:AOT

[BJ96] Edmond J. Breen and Ronald Jones. Attribute openings, thinnings, and granulometries. Computer Vision and Image Understanding: CVIU, 64(3):377– 389, November 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0066/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0066/production/pdf.

Basri:1997:CS

[BJ97] Ronen Basri and David Jacobs. Constancy and similarity. Computer Vision and Image Understanding: CVIU, 65(3):447– 449, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0497/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0497/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0497/production/ref.

Biswas:2014:AS

[BJ14] Arijit Biswas and David Jacobs. Active subclustering. Computer Vision and Image Understanding: CVIU, 125(??):72– 84, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000745.

Bugaric:2014:AEV

[BJS14] Marin Bugarić, Toni Jakovcević, and Darko Stipanicev. Adaptive estimation of visual smoke detection parameters based on spatial data and fire risk index. Computer Vision and Image Understanding: CVIU, 118(??): 184–196, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001860.

Barron:2001:EAP

[BK01] Carlos Barrón and Ioannis A. Kakadiaris. Estimating anthropometry and pose from a single uncalibrated image. Computer Vision and Image Understanding: CVIU, 81 (3):269–284, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0888;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0888/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0888/ref.

Bassiou:2007:CIH

[BK07] Nikoletta Bassiou and Constantine Kotropoulos. Color image histogram equalization by absolute discounting back-off. Computer Vision and Image Understanding: CVIU, 107(1–2):108– 122, July/August 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barbu:2015:EIC

[BK15] Adrian Barbu and Iasonas Kokkinos. Editorial introduction to the CVIU special issue on “Generative models in computer vision and medical imaging”. Computer Vision and Image Understanding: CVIU, 136 (??):1–2, July 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421500106X.

Baek:2016:SFC

[BK16] Seung-Hwan Baek and Min H. Kim. Stereo fusion: Combining refractive and binocular disparity. Computer Vision and Image Understanding: CVIU, 146 (??):52–66, May 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216000515.

Bowyer:2001:EDE

[BKD01] Kevin Bowyer, Christine Kranenburg, and Sean Dougherty. Edge detector evaluation using empirical ROC curves. Computer Vision and Image Understanding: CVIU, 84 (1):77–103, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0931;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0931/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0931/ref.

Bastys:2011:IRF

[BKK11] Algirdas Bastys, Justas Kranauskas, and Volker Krüger. Iris recognition by fusing different representations of multi-scale Taylor expansion. Computer Vision and Image Understanding: CVIU, 115(6):804– 816, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bayer:1998:CPD

[BKMSR98] T. Bayer, U. Kressel, H. Mogg-Schneider, and I. Renz. Categorizing paper documents. A generic system for domain and language independent text categorization. Computer Vision and Image Understanding: CVIU, 70(3):299–??, ???? 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bray:2007:SPF

[BKMV07] M. Bray, E. Koller-Meier, and L. Van Gool. Smart particle filtering for highdimensional tracking. Computer Vision and Image Understanding: CVIU, 106(1): 116–129, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Baker:2010:USA

[BKP10] Patrick Baker and Behrooz Kamgar-Parsi. Using shorelines for autonomous air vehicle guidance. Computer Vision and Image Understanding: CVIU, 114(6): 723–729, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Banerjee:2015:RCI

[BKPS15] Tanvi Banerjee, James M. Keller, Mihail Popescu, and Marjorie Skubic. Recognizing complex instrumental activities of daily living using scene information and fuzzy logic. Computer Vision and Image Understanding: CVIU, 140(??): 68–82, November 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000892.

Bolduc:1998:RBM

[BL98a] Marc Bolduc and Martin D. Levine. A review of biologically motivated space-variant data reduction models for robotic vision. Computer Vision and Image Understanding: CVIU, 69(2):170– 184, February 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0560/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0560/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0560/production/ref.

Bretzner:1998:FTA

[BL98b] Lars Bretzner and Tony Lindeberg. Feature tracking with automatic selection of spatial scales. Computer Vision and Image Understanding: CVIU, 71(3):385– 392, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0650/production;   http:

//www.idealibrary.com/
links/artid/cviu.1998.
0650/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0650/production/ref.

Bleau:2000:WBS

[BL00] André Bleau and L. Joshua Leon. Watershed-based segmentation and region merging. Computer Vision and Image Understanding: CVIU, 77(3):317– 370, March 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0822;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0822/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0822/ref.

Bottino:2001:SBT

[BL01] Andrea Bottino and Aldo Laurentini. A silhouette based technique for the reconstruction of human movement. Computer Vision and Image Understanding: CVIU, 83(1): 79–95, July 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0918;   http:
//www.idealibrary.com/

links/doi/10.1006/cviu.
2001.0918/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0918/ref.

Berwick:2004:SGC

[BL04] Daniel Berwick and Sang Wook Lee. Spectral gradients for color-based object recognition and indexing. Computer Vision and Image Understanding: CVIU, 94(1– 3):28–43, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bottino:2008:VHP

[BL08] Andrea Bottino and Aldo Laurentini. The visual hull of piecewise smooth objects. Computer Vision and Image Understanding: CVIU, 110 (1):7–18, April 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bhandarkar:2009:IDT

[BL09] Suchendra M. Bhandarkar and Xingzhi Luo. Integrated detection and tracking of multiple faces using particle filtering and optical flow-based elastic matching. Computer Vision and Image Understanding: CVIU, 113(6):708–725, June 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bennett:2014:CQR

[BL14] Stuart Bennett and Joan Lasenby. ChESS — quick and robust detection of chess-board features. Computer Vision and Image Understanding: CVIU, 118(??):197– 210, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001999.

Bicego:2016:BAS

[BL16] Manuele Bicego and Pietro Lovato. A bioinformatics approach to 2D shape classification. Computer Vision and Image Understanding: CVIU, 145(??): 59–69, April 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002556.

Ban:2020:PDD

[BL20] Yuseok Ban and Sangyoun Lee. Protuberance of depth: Detecting interest points from a depth image. Computer Vision and Image Understanding: CVIU, 194(??):Article 102927, May 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300187.

Bazzica:2016:DPN

[BLH16] Alessio Bazzica, Cynthia C. S. Liem, and Alan Hanjalic. On detecting the playing/non-playing activity of musicians in symphonic music videos. Computer Vision and Image Understanding: CVIU, 144(??):188– 204, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002040.

Bao:2023:IDM

[BLJ^{+}23] Fangxun Bao, Yifan Lei, Yiqiao Jia, Hongwei Du, Chengyong Gao, and Yunfeng Zhang. An image denoising method based on the nonlinear Schrödinger equation and spectral subband decomposition. Computer Vision and Image Understanding: CVIU, 237(??): ??, December 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002205.

Baslamisli:2021:PBS

[BLKG21] Anil S. Baslamisli, Yang Liu, Sezer Karaoglu, and Theo Gevers. Physicsbased shading reconstruction for intrinsic image decomposition. Computer Vision and Image Understanding: CVIU, 205(??):Article 103183, April 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000278.

Bergevin:1995:RRV

[BLP95] Robert Bergevin, Denis Laurendeau, and Denis Poussart. Registering range views of multipart objects. Computer Vision and Image Understanding: CVIU, 61 (1):1–16, January 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1001/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1001/production/pdf.

Bricault:1997:VMI

[BM97] Ivan Bricault and Olivier Monga. From volume medical images to quadratic surface patches. Computer Vision and Image Understanding: CVIU, 67(1): 24–38, July 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0501/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0501/production/pdf;
http://www.idealibrary.

com/links/artid/cviu.1996.
0501/production/ref.

Bergevin:1998:MCS

[BM98] Robert Bergevin and Marielle Mokhtari. Multiscale contour segmentation and approximation: An algorithm based on the geometry of regular inscribed polygons. Computer Vision and Image Understanding: CVIU, 71(1):55–73, July 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0634/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0634/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0634/production/ref.

Baillard:1999:RUS

[BM99] C. Baillard and H. Maître.

3D reconstruction of urban scenes from aerial stereo imagery: a focusing strategy. Computer Vision and Image Understanding: CVIU, 76(3):244– 258, December 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0793/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.

                                        0793/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0793/production/ref.

Boxer:2000:ECE

[BM00] Laurence Boxer and Russ Miller. Efficient computation of the Euclidean distance transform. Computer Vision and Image Understanding: CVIU, 80(3): 379–383, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2000.0880;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0880/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0880/ref. See corri-

gendum [BM02].

Boxer:2002:CEC

[BM02] Laurence Boxer and Russ Miller. Corrigendum to “Efficient Computation of the Euclidean Distance Transform”. Computer Vision and Image Understanding: CVIU, 86(2):137–140, May 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See [BM00].

Bhattacharjee:2015:PBD

[BM15] Sreyasee Das Bhattacharjee and Anurag Mittal.

Part-based deformable object detection with a single sketch. Computer Vision and Image Understanding: CVIU, 139(??):73–87, October 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215001332.

Betancourt:2017:LRH

[BMB^{+}17] Alejandro Betancourt, Pietro Morerio, Emilia Barakova, Lucio Marcenaro, Matthias Rauterberg, and Carlo Regazzoni. Left/right hand segmentation in egocentric videos. Computer Vision and Image Understanding: CVIU, 154(??):73–81, January 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301394.

Barros:2017:DGR

[BMJF^{+}17] Pablo Barros, Nestor T.

Maciel-Junior, Bruno J. T. Fernandes, Byron L. D. Bezerra, and Sergio M. M. Fernandes. A dynamic gesture recognition and prediction system using the convexity approach. Computer Vision and Image Understanding: CVIU, 155(??):139– 149, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421630159X.

Bozorgtabar:2019:ISG

[BMvT^{+}19] Behzad Bozorgtabar, Dwarikanath Mahapatra, Hendrik von Teng, Alexander Pollinger, Lukas Ebner, Jean-Phillipe Thiran, and Mauricio Reyes. Informative sample generation using class aware generative adversarial networks for classification of chest X-rays. Computer Vision and Image Understanding: CVIU, 184(??): 57–65, July 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421930061X.

Byvshev:2022:CNI

[BMX22] Petr Byvshev, Pascal Mettes, and Yu Xiao. Are 3D convolutional networks inherently biased towards appearance? Computer Vision and Image Understanding: CVIU, 220(??):??, July 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222000534.

Babaee:2015:DOM

[BN15] M. Babaee and S. Negahdaripour. 3-D object modeling from 2-D occluding contour correspondences by opti-acoustic stereo imaging. Computer Vision and Image Understanding: CVIU, 132(??): 56–74, March 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002112.

Braga-Neto:2002:CCL

[BNG02] Ulisses Braga-Neto and John Goutsias. Connectivity on complete lattices: New results. Computer Vision and Image Understanding: CVIU, 85(1):22– 53, January 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Braga-Neto:2003:MAC

[BNG03] Ulisses Braga-Neto and John Goutsias. A multiscale approach to connectivity. Computer Vision and Image Understanding: CVIU, 89(1):70–107, January 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Braga-Neto:2005:CMC

[BNG05] Ulisses Braga-Neto and John Goutsias. Constructing multiscale connectivities. Computer Vision and Image Understanding: CVIU, 99(1):126–150, July 2005. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Bookstein:1997:SIM

[Boo97] Fred L. Bookstein. Shape and the information in medical images: a decade of the morphometric synthesis. Computer Vision and Image Understanding: CVIU, 66 (2):97–118, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0607/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0607/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0607/production/ref.

Borgefors:1996:DDT

[Bor96] Gunilla Borgefors. On digital distance transforms in three dimensions. Computer Vision and Image Understanding: CVIU, 64(3):368– 376, November 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0065/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0065/production/pdf.

Borji:2019:PCG

[Bor19] Ali Borji. Pros and cons of GAN evaluation measures. Computer Vision and Image Understanding: CVIU, 179(??):41–65, February 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218304272.

Borji:2022:PCG

[Bor22] Ali Borji. Pros and cons of GAN evaluation measures: New developments. Computer Vision and Image Understanding: CVIU, 215 (??):??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001685.

Boyd:2004:SOM

[Boy04] Jeffrey E. Boyd. Synchronization of oscillations for machine perception of gaits. Computer Vision and Image Understanding: CVIU, 96(1):35–59, October 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Benboudjema:2005:UIS

[BP05] Dalila Benboudjema and Wojciech Pieczynski. Unsupervised image segmentation using triplet Markov fields. Computer Vision and Image Understanding: CVIU, 99(3):476–498, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bugeau:2009:DSM

[BP09] Aurélie Bugeau and Patrick Pérez. Detection and segmentation of moving objects in complex scenes. Computer Vision and Image Understanding: CVIU, 113(4): 459–476, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bulo:2011:GBQ

[BPB11] Samuel Rota Bulò, Marcello Pelillo, and Immanuel M. Bomze. Graph-based quadratic optimization: a fast evolutionary approach. Computer Vision and Image Understanding: CVIU, 115 (7):984–995, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211000713.

Bogovic:2013:MOG

[BPB13] John A. Bogovic, Jerry L. Prince, and Pierre-Louis Bazin. A multiple object geometric deformable model for image segmentation. Computer Vision and Image Understanding: CVIU, 117(2):145–157, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001440.

Batenburg:2013:DAS

[BPBS13] K. Joost Batenburg, Willem Jan Palenstijn, Péter Balázs, and Jan Sijbers. Dynamic angle selection in binary tomography. Computer Vision and Image Understanding: CVIU, 117(4):306– 318, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001774.

Bhowmick:2017:DCH

[BPC^{+}17] Brojeshwar Bhowmick, Suvam Patra, Avishek Chatterjee, Venu Madhav Govindu, and Subhashis Banerjee. Divide and conquer: a hierarchical approach to large-scale structure-from-motion. Computer Vision and Image Understanding: CVIU, 157(??):190– 205, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300346.

Borlino:2022:SSM

[BPCT22] Francesco Cappio Borlino, Salvatore Polizzotto, Barbara Caputo, and Tatiana Tommasi. Self-supervision and meta-learning for oneshot unsupervised crossdomain detection. Computer Vision and Image Understanding: CVIU, 223 (??):??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001278.

Bratanic:2015:RTP

[BPLT15] Blaz Bratanic, Franjo Pernus, Bostjan Likar, and Dejan Tomazevic. Real-time pose estimation of rigid objects in heavily cluttered environments. Computer Vision and Image Understanding: CVIU, 141(??): 38–51, December 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001976.

Bhanu:2015:EIS

[BPQ15] Bir Bhanu, Andrea Prati, and Faisal Qureshi. Editorial introduction to the special issue on “Image Understanding for Real-World Distributed Video Networks” — Computer Vision and Image Understanding journal. Computer Vision and Image Understanding: CVIU, 134(??): 46–47, May 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000557.

Barreto:2010:SIO

[BPS10] Joao P. Barreto, Tomas Pajdla, and Akihiro Sugimoto. Special issue on omnidirectional vision, camera networks and non-conventional cameras. Computer Vision and Image Understanding: CVIU, 114(2):167, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Brun:2016:ARU

[BPSV16] Luc Brun, Gennaro Percannella, Alessia Saggese, and Mario Vento. Action recognition by using kernels on aclets sequences. Computer Vision and Image Understanding: CVIU, 144(??): 3–13, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001988.

Beveridge:1995:OGM

[BR95] J. Ross Beveridge and Edward M. Riseman. Optimal geometric model matching under full 3D perspective. Computer Vision and Image Understanding: CVIU, 61 (3):351–364, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1028/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1028/production/pdf.

Bhavsar:2012:RMS

[BR12] Arnav V. Bhavsar and Ambasamudram N. Rajagopalan. Range map superresolution-inpainting, and reconstruction from sparse data. Computer Vision and Image Understanding: CVIU, 116(4): 572–591, April 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002724.

Brand:1997:PBV

[Bra97] Matthew Brand. Physicsbased visual understanding. Computer Vision and Image Understanding: CVIU, 65(2):192–205, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0572/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0572/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0572/production/ref.

Bascon:2010:OPI

[BRA^{+}10] S. Maldonado Bascón, J. Acevedo Rodríguez, S. Lafuente Arroyo, A. Fernndez Caballero, and F. López-Ferreras. An optimization on pictogram identification for the road-sign recognition task using SVMs. Computer Vision and Image Understanding: CVIU, 114(3): 373–383, March 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bretto:2001:CGD

[Bre01] Alain Bretto. Comparability graphs and digital topology. Computer Vision and Image Understanding: CVIU, 82(1):33–41, April 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0901;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0901/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0901/ref.

Breuel:2003:ITG

[Bre03] Thomas M. Breuel. Implementation techniques for geometric branch-and-bound matching methods. Computer Vision and Image Understanding: CVIU, 90(3): 258–294, June 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bozorgtabar:2019:LSS

[BRET19] Behzad Bozorgtabar, Mohammad Saeed Rad, Hazim Kemal Ekenel, and Jean-Philippe Thiran. Learn to synthesize and synthesize to learn. Computer Vision and Image Understanding: CVIU, 185(??):1– 11, August 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219300657.

Brito:2017:ASM

[Bri17] José Henrique Brito. Autocalibration for structure from motion. Computer Vision and Image Understanding: CVIU, 157(??): 240–254, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216302053.

Bolle:2004:EAP

[BRP04] Ruud M. Bolle, Nalini K. Ratha, and Sharath Pankanti. Error analysis of pattern recognition systems—the subsets bootstrap. Computer Vision and Image Understanding: CVIU, 93(1): 1–33, January 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bui:2017:CDS

[BRPC17] T. Bui, L. Ribeiro, M. Ponti, and J. Collomosse. Compact descriptors for sketchbased image retrieval using a triplet loss convolutional neural network. Computer Vision and Image Understanding: CVIU, 164(??): 27–37, November 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301194.

Bartrina-Rapesta:2011:JRC

[BRSSAL11] Joan Bartrina-Rapesta, Joan Serra-Sagristà, and Francesc Aulí-Llinàs. JPEG2000 ROI coding through component priority for digital mammography. Computer Vision and Image Understanding: CVIU, 115(1):59–68, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barequet:1996:PLI

[BS96] Gill Barequet and Micha Sharir. Piecewise-linear interpolation between polygonal slices. Computer Vision and Image Understanding: CVIU, 63(2):251– 272, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0018/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0018/production/pdf.

Berman:1999:FID

[BS99a] Andrew P. Berman and Linda G. Shapiro. A flexible image database system for content-based retrieval. Computer Vision and Image Understanding: CVIU, 75(1–2):175– 195, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0772/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0772/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0772/production/ref.

Boyer:1999:GEI

[BS99b] Kim L. Boyer and Sudeep Sarkar. Guest Editors’ introduction: Perceptual organization in computer vision: Status, challenges, and potential. Computer Vision and Image Understanding: CVIU, 76(1):1– 5, October 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731429990797X.

Brejl:2000:DED

[BS00a] Marek Brejl and Milan Sonka. Directional 3D edge detection in anisotropic data: Detector design and performance assessment. Computer Vision and Image Understanding: CVIU, 77 (2):84–110, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0811;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0811/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0811/ref.

Bubna:2000:MST

[BS00b] Kishore Bubna and Charles V.

Stewart. Model selection techniques and merging rules for range data segmentation algorithms. Computer Vision and Image Understanding: CVIU, 80(2): 215–245, November 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0871;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0871/pdf;   http:
//www.idealibrary.com/

                                        links/doi/10.1006/cviu.
                                        2000.0871/ref.

Bartoli:2005:SMU

[BS05] Adrien Bartoli and Peter Sturm. Structure-from-motion using lines: Representation, triangulation, and bundle adjustment. Computer Vision and Image Understanding: CVIU, 100(3):416–441, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Baskurt:2019:VSS

[BS19] Kemal Batuhan Baskurt and Refik Samet. Video synopsis: a survey. Computer Vision and Image Understanding: CVIU, 181(??): 26–38, April 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219300190.

Ben-Shabat:2018:GBS

[BSALF18] Yizhak Ben-Shabat, Tamar Avraham, Michael Lindenbaum, and Anath Fischer. Graph based oversegmentation methods for 3D point clouds. Computer Vision and Image Understanding: CVIU, 174(??): 12–23, September 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421830078X.

Brunton:2014:RSS

[BSBW14] Alan Brunton, Augusto Salazar, Timo Bolkart, and Stefanie Wuhrer. Review of statistical shape spaces for 3D data with comparative analysis for human faces. Computer Vision and Image Understanding: CVIU, 128(??):1–17, November 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001131.

Boyer:2002:SSS

[BSF02] Kim L. Boyer, Ravi Srikantiah, and Patrick J. Flynn. Saliency sequential surface organization for free-form object recognition. Computer Vision and Image Understanding: CVIU, 88(3): 152–188, December 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bryner:2013:ESM

[BSH13] D. Bryner, A. Srivastava, and Q. Huynh. Elastic shapes models for improving segmentation of object boundaries in synthetic aperture sonar images. Computer Vision and Image Understanding: CVIU, 117(12):1695–1710, December 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213001318.

Bhowmick:2022:NPS

[BSH22] Alexy Bhowmick, Sarat Saharia, and Shyamanta M. Hazarika. Non-parametric scene parsing: Label transfer methods and datasets. Computer Vision and Image Understanding: CVIU, 219(??):??, June 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222000431.

Babu:2010:OAR

[BSM10] R. Venkatesh Babu, S. Suresh, and Anamitra Makur. Online adaptive radial basis function networks for robust object tracking. Computer Vision and Image Understanding: CVIU, 114(3): 297–310, March 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Binder:2013:ERM

[BSMK13] Alexander Binder, Wojciech Samek, Klaus-Robert Müller, and Motoaki Kawanabe. Enhanced representation and multi-task learning for image annotation. Computer Vision and Image Understanding: CVIU, 117(5): 466–478, May 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001713.

Bodis-Szomoru:2017:EEA

[BSRV17] András Bódis-Szomorú, Hayko Riemenschneider, and Luc Van Gool. Efficient edgeaware surface mesh reconstruction for urban scenes. Computer Vision and Image Understanding: CVIU, 157 (??):3–24, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300789.

Bisagno:2021:EGO

[BSZ^{+}21] Niccoló Bisagno, Cristiano Saltori, Bo Zhang, Francesco G. B. De Natale, and Nicola Conci. Embedding group and obstacle information in LSTM networks for human trajectory prediction in crowded scenes. Computer Vision and Image Understanding: CVIU, 203(??):Article 103126, February 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301454.

Bentoutou:2005:DST

[BT05] Y. Bentoutou and N. Taleb. A 3-D space–time motion detection for an invariant image registration approach in digital subtraction angiography. Computer Vision and Image Understanding: CVIU, 97(1):30–50, January 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Baig:2017:MHC

[BT17] Mohammad Haris Baig and Lorenzo Torresani. Multiple hypothesis colorization and its application to image compression. Computer Vision and Image Understanding: CVIU, 164(??):111– 123, November 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300267.

Buzuti:2023:FAD

[BT23] Lucas F. Buzuti and Carlos E. Thomaz. Fréchet AutoEncoder Distance: a new approach for evaluation of Generative Adversarial Networks. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001480.

Bleichrodt:2014:SAD

[BTB14] F. Bleichrodt, F. Tabak, and K. J. Batenburg. SDART: an algorithm for discrete tomography from noisy projections. Computer Vision and Image Understanding: CVIU, 129(??):63–74, December 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001295.

Badoual:2019:TDP

[BUD19] Anaïs Badoual, Michael Unser, and Adrien Depeursinge. Texture-driven parametric snakes for semiautomatic image segmentation. Computer Vision and Image Understanding: CVIU, 188(??):Article 102793, November 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314218303369.

Bucher:2021:HNT

[BVCP21] Maxime Bucher, Tuan-Hung Vu, Matthieu Cord, and Patrick Pérez. Handling new target classes in semantic segmentation with domain adaptation. Computer Vision and Image Understanding: CVIU, 212(??):

??, November 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001028.

Berendsen:2013:FFI

[BvdHL^{+}13] Floris F. Berendsen, Uulke A. van der Heide, Thomas R. Langerak, Alexis N. T. J. Kotte, and Josien P. W. Pluim. Free-form image registration regularized by a statistical shape model: application to organ segmentation in cervical MR. Computer Vision and Image Understanding: CVIU, 117(9):1119– 1127, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000660.

Berenguer-Vidal:2015:CAM

[BVVMMS15] R. Berenguer-Vidal, R. Verdú-Monedero, and J. Morales-Sánchez. Convergence analysis of multidimensional parametric deformable models. Computer Vision and Image Understanding: CVIU, 135(??):157– 177, June 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000247.

Banerjee:2021:RUC

[BVWS21] Sreya Banerjee, Rosaura G.

VidalMata, Zhangyang Wang, and Walter J. Scheirer. Report on UG^{2+} challenge Track 1: Assessing algorithms to improve video object detection and classification from unconstrained mobility platforms. Computer Vision and Image Understanding: CVIU, 213(??): ??, December 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001417.

Bai:2011:RTU

[BW11] Li Bai and Yan Wang. Road tracking using particle filters with partition sampling and auxiliary variables. Computer Vision and Image Understanding: CVIU, 115(10):1463–1471, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211001421.

Bolkart:2015:FMF

[BW15] Timo Bolkart and Stefanie Wuhrer. 3D faces in motion: Fully automatic registration and statistical analysis. Computer Vision and Image Understanding: CVIU, 131(??):100– 115, February 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214001404.

Brown:2017:GFS

[BWG17] Mark Brown, David Windridge, and Jean-Yves Guillemaut. A generalised framework for saliencybased point feature detection. Computer Vision and Image Understanding: CVIU, 157(??):117– 137, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216301424.

Bischof:2004:IIR

[BWL04] Horst Bischof, Horst Wildenauer, and Aleš Leonardis. Illumination insensitive recognition using eigenspaces. Computer Vision and Image Understanding: CVIU, 95 (1):86–104, July 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bolle:1998:CVV

[BY98] R. Bolle and B.-L. Yeo. Computer vision for visual computing: Techniques and applications. Computer Vision and Image Understanding: CVIU, 71(2):153–??, ???? 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bastanlar:2008:CVB

[BY08] Yalin Bastanlar and Yasemin Yardimci. Corner validation based on extracted corner properties. Computer Vision and Image Understanding: CVIU, 112(3):243–261, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bilir:2012:NRS

[BY12] S. C. Bilir and Y. Yemez.

Non-rigid 3D shape tracking from multiview video. Computer Vision and Image Understanding: CVIU, 116(11):1121–1134, November 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314212001026.

Bi:2023:LRE

[BYJG23] Qi Bi, Shaodi You, Wei Ji, and Theo Gevers. Learning rotation equivalent scene representation from instance-level semantics: a novel top-down perspective. Computer Vision and Image Understanding: CVIU, 229 (??):??, March 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000152.

Byeon:2018:UOF

[BYK^{+}18] Moonsub Byeon, Haanju Yoo, Kikyung Kim, Songhwai Oh, and Jin Young Choi. Unified optimization framework for localization and tracking of multiple targets with multiple cameras. Computer Vision and Image Understanding: CVIU, 166(??): 51–65, January 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301728.

Bing:2004:TSM

[BYN^{+}04] Cheng Bing, Wang Ying, Zheng Nanning, Bian Zhengzhong, and Zhang Yongping. A two-step method for preprocessing volume data. Computer Vision and Image Understanding: CVIU, 95(2): 150–164, August 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Babaee:2017:CSR

[BYR17] M. Babaee, Y. You, and G. Rigoll. Combined segmentation, reconstruction, and tracking of multiple targets in multi-view video sequences. Computer Vision and Image Understanding: CVIU, 154(??):166– 181, January 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421630114X.

Boshra:1999:CSA

[BZ99] Michael Boshra and Hong Zhang. A constraintsatisfaction approach for 3D object recognition by integrating 2D and 3D data. Computer Vision and Image Understanding: CVIU, 73(2):200–214, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0730/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0730/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0730/production/ref.

Bouwmans:2014:RPP

[BZ14] Thierry Bouwmans and El Hadi Zahzah. Robust PCA via Principal Component Pursuit: a review for a comparative evaluation in video surveillance. Computer Vision and Image Understanding: CVIU, 122 (??):22–34, May 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213002294.

Bartoccioni:2023:LMM

[BZP^{+}23] Florent Bartoccioni, Éloi Zablocki, Patrick Pérez, Matthieu Cord, and Karteek Alahari. LiDAR-Touch: Monocular metric depth estimation with a fewbeam LiDAR. Computer Vision and Image Understanding: CVIU, 227(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001795.

Basharat:2008:CBV

[BZS08] Arslan Basharat, Yun Zhai, and Mubarak Shah. Content based video matching using spatiotemporal volumes. Computer Vision and Image Understanding: CVIU, 110(3):360–377, June 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bagheri:2016:TMS

[BZS16] Saeid Bagheri, Jiang Yu Zheng, and Shivank Sinha. Temporal mapping of surveillance video for indexing and summarization. Computer Vision and Image Understanding: CVIU, 144(??): 237–257, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002581.

Chang:1997:LCC

[CA97] Yuh-Lin Chang and J. K.

Aggarwal. Line correspondences from cooperating spatial and temporal grouping processes for a sequence of images. Computer Vision and Image Understanding: CVIU, 67(2): 186–201, August 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0527/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0527/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0527/production/ref.

Cavallaro:2010:SIM

[CA10] Andrea Cavallaro and Hamid Aghajan. Special issue on multi-camera and multimodal sensor fusion. Computer Vision and Image Understanding: CVIU, 114(6): 609–610, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Comino:2017:EAC

[CACB17] Marc Comino, Carlos Andújar, Antonio Chica, and Pere Brunet. Error-aware construction and rendering of multi-scan panoramas from massive point clouds. Computer Vision and Image Understanding: CVIU, 157 (??):43–54, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301461.

Cherel:2024:PBS

[CAGN24] Nicolas Cherel, Andrés Almansa, Yann Gousseau, and Alasdair Newson. Patchbased stochastic attention for image editing. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223002461.

Chang:2020:GMB

[CALO20] Seunggyu Chang, Chanho Ahn, Minsik Lee, and Songhwai Oh. Graphmatching-based correspondence search for nonrigid point cloud registration. Computer Vision and Image Understanding: CVIU, 192(??):Article 102899, March 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219301766.

Cho:2023:ESU

[CAO^{+}23] Jae Won Cho, Dawit Mureja Argaw, Youngtaek Oh, Dong-Jin Kim, and In So Kweon. Empirical study on using adapters for debiased Visual Question Answering. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002229.

Carstensen:1996:ALM

[Car96] Jens Michael Carstensen.

An active lattice model in a Bayesian framework. Computer Vision and Image Understanding: CVIU, 63(2): 380–387, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0027/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0027/production/pdf.

Carlin:2001:MPS

[Car01] Mats Carlin. Measuring the performance of shape similarity retrieval methods. Computer Vision and Image Understanding: CVIU, 84(1):44–61, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/

cviu.2001.0935;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0935/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0935/ref.

Chen:1998:SID

[CB98] Francine R. Chen and Dan S. Bloomberg. Summarization of imaged documents without OCR. Computer Vision and Image Understanding: CVIU, 70(3): 307–320, June 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0688/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0688/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0688/production/ref.

Cooper:1995:CSU

[CBB95] Paul R. Cooper, Lawrence A. Birnbaum, and Matthew E. Brand. Causal scene understanding. Computer Vision and Image Understanding: CVIU, 62(2):215– 231, September 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1051/production;   http:
//www.idealibrary.com/

links/artid/cviu.1995.
1051/production/pdf.

Chaudhuri:2019:SGC

[CBB19] Ushasi Chaudhuri, Biplab Banerjee, and Avik Bhattacharya. Siamese graph convolutional network for content based remote sensing image retrieval. Computer Vision and Image Understanding: CVIU, 184 (??):22–30, July 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1077314219300578.

Cachier:2003:IFB

[CBD^{+}03] Pascal Cachier, Eric Bardinet, Didier Dormont, Xavier Pennec, and Nicholas Ayache. Iconic feature based nonrigid registration: the PASHA algorithm. Computer Vision and Image Understanding: CVIU, 89 (2–3):272–298, February/ March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2001:SED

[CBM01] Y. Chen, C. A. Z. Barcelos, and B. A. Mair. Smoothing and edge detection by time-varying coupled nonlinear diffusion equations. Computer Vision and Image Understanding: CVIU, 82(2):85–100, May 2001.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0903;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0903/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0903/ref.

Cakir:2017:OSH

[CBS17] Fatih Cakir, Sarah Adel Bargal, and Stan Sclaroff. Online supervised hashing. Computer Vision and Image Understanding: CVIU, 156(??):162– 173, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301606.

Chaudhuri:2004:ESE

[CBT^{+}04] A. Ray Chaudhuri, A. Basu, K. Tan, S. Bhandari, and B. B. Chaudhuri. An efficient set estimator in high dimensions: consistency and applications to fast data visualization. Computer Vision and Image Understanding: CVIU, 93(3):260– 287, March 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cornia:2023:FAI

[CBTC23] Marcella Cornia, Lorenzo Baraldi, Ayellet Tal, and Rita Cucchiara. Fullyattentive iterative networks for region-based controllable image and video captioning. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002370.

Cohen:1996:HHM

[CC96] Isaac Cohen and Laurent D.

Cohen. Hybrid hyperquadric model for 2-D and 3-D data fitting. Computer Vision and Image Understanding: CVIU, 63(3): 527–541, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0039/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0039/production/pdf.

Cretual:2000:DSP

[CC00] Armel Crétual and François Chaumette. Dynamic stabilization of a pan and tilt camera for submarine image visualization. Computer Vision and Image Understanding: CVIU, 79(1):

47–65, July 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0849;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0849/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0849/ref.

Chen:2001:ERA

[CC01] Teh-Chuan Chen and Kuo-Liang Chung. An efficient randomized algorithm for detecting circles. Computer Vision and Image Understanding: CVIU, 83 (2):172–191, August 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0923;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0923/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0923/ref.

Chowdhury:2003:FRM

[CC03] Amit K. Roy Chowdhury and Rama Chellappa. Face reconstruction from monocular video using uncertainty analysis and a generic model. Computer Vision and Image Understanding: CVIU, 91(1–2):188– 213, July/August 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cheng:2007:BSM

[CC07] Li Cheng and Terry Caelli.

Bayesian stereo matching. Computer Vision and Image Understanding: CVIU, 106(1):85–96, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2011:SSV

[CC11] Yu Chen and Roberto Cipolla. Single and sparse view 3D reconstruction by learning shape priors. Computer Vision and Image Understanding: CVIU, 115(5): 586–602, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cosar:2015:SDB

[CÇ15] Serhan Cosar and Müjdat Çetin. Sparsity-driven bandwidth-efficient decentralized tracking in visual sensor networks. Computer Vision and Image Understanding: CVIU, 139(??): 40–58, October 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421500123X.

Cristina:2016:MBH

[CC16] Stefania Cristina and Kenneth P. Camilleri. Modelbased head pose-free gaze estimation for assistive communication. Computer Vision and Image Understanding: CVIU, 149(??):157– 170, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000667.

Colombo:2011:SRT

[CCD11] Carlo Colombo, Dario Comanducci, and Alberto Del Bimbo. Shape reconstruction and texture sampling by active rectification and virtual view synthesis. Computer Vision and Image Understanding: CVIU, 115 (2):161–176, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2017:NFM

[CCF17] Junkai Chen, Zheru Chi, and Hong Fu. A new framework with multiple tasks for detecting and locating pain events in video. Computer Vision and Image Understanding: CVIU, 155(??):113– 123, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301813.

Chaquet:2013:SVD

[CCFC13] Jose M. Chaquet, Enrique J. Carmona, and Antonio Fernández-Caballero. A survey of video datasets for human action and activity recognition. Computer Vision and Image Understanding: CVIU, 117(6): 633–659, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000295.

Chang:2004:LTC

[CCL04] Fu Chang, Chun-Jen Chen, and Chi-Jen Lu. A lineartime component-labeling algorithm using contour tracing technique. Computer Vision and Image Understanding: CVIU, 93(2):206–220, February 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2017:ETS

[CCL^{+}17] Yisong Chen, Antoni B.

Chan, Zhouchen Lin, Kenji Suzuki, and Guoping Wang. Efficient tree-structured SfM by RANSAC generalized Procrustes analysis. Computer Vision and Image Understanding: CVIU, 157(??):179– 189, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300334.

Chaudhuri:1997:NAC

[CCP97] A. Ray Chaudhuri, B. B. Chaudhuri, and S. K. Parui. A novel approach to computation of the shape of a dot pattern and extraction of its perceptual border. Computer Vision and Image Understanding: CVIU, 68(3):257–275, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0550/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0550/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0550/production/ref.

Chang:2016:VBB

[CCPK16] Tommy Chang, Bharath Comandur, Johnny Park, and Avinash C. Kak. A variance-based Bayesian framework for improving land-cover classification through wide-area learning from large geographic regions. Computer Vision and Image Understanding: CVIU, 147(??): 3–22, June 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421630025X.

Cernadas:2005:AMR

[CCR^{+}05] E. Cernadas, P. Carrión, P. G. Rodriguez, E. Muriel, and T. Antequera. Analyzing magnetic resonance images of Iberian pork loin to predict its sensorial characteristics. Computer Vision and Image Understanding: CVIU, 98(2):344– 360, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Choy:1995:NSP

[CCS95] Steven S. O. Choy, Clifford Sze-Tsan Choy, and Wan-Chi Siu. New singlepass algorithm for parallel thinning. Computer Vision and Image Understanding: CVIU, 62(1): 69–77, July 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1042/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1042/production/pdf.

Choudhury:2001:RBR

[CCS01] Ragini Choudhury, Santanu Chaudhury, and J. B. Srivastava. Reconstruction based recognition of scenes with multiple repeated components. Computer Vision and Image Understanding: CVIU, 84(3):325–360, December 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ciocca:2014:USF

[CCSS14] Gianluigi Ciocca, Claudio Cusano, Simone Santini, and Raimondo Schettini. On the use of supervised features for unsupervised image categorization: an evaluation. Computer Vision and Image Understanding: CVIU, 122(??):155– 171, May 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421400023X.

Cardoso:2009:PDM

[CCTCR09] Jaime S. Cardoso, Pedro Carvalho, Luís F. Teixeira, and Luís Corte-Real. Partition-distance methods for assessing spatial segmentations of images and videos. Computer Vision and Image Understanding: CVIU, 113(7):811–823, July 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Choi:2012:RMO

[CCYC12] JinMin Choi, Hyung Jin Chang, Yung Jun Yoo, and Jin Young Choi. Robust moving object detection against fast illumination change. Computer Vision and Image Understanding: CVIU, 116(2):179–193, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002347.

Chen:2010:PPB

[CD10] Fan Chen and Christophe De Vleeschouwer. Personalized production of basketball videos from multisensored data under limited display resolution. Computer Vision and Image Understanding: CVIU, 114(6): 667–680, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chung:2013:RAM

[CD13] François Chung and Hervé Delingette. Regional appearance modeling based on the clustering of intensity profiles. Computer Vision and Image Understanding: CVIU, 117(6): 705–717, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000271.

Csurka:1999:FCB

[CDH99] Gabriella Csurka, David Demirdjian, and Radu Horaud. Finding the collineation between two projective reconstructions. Computer Vision and Image Understanding: CVIU, 75(3):260–268, September 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0782/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0782/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0782/production/ref.

Comic:2014:TMH

[CDIF14] Lidija Comić, Leila De Floriani, Federico Iuricich, and Ulderico Fugacci. Topological modifications and hierarchical representation of cell complexes in arbitrary dimensions. Computer Vision and Image Understanding: CVIU, 121(??): 2–12, April 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213002312.

Cerri:2014:CST

[CDJM14] Andrea Cerri, Barbara Di Fabio, Grzegorz Jabloński, and Filippo Medri. Comparing shapes through multiscale approximations of the matching distance. Computer Vision and Image Understanding: CVIU, 121 (??):43–56, April 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421300221X.

Chaperon:2011:RCP

[CDT11] Thomas Chaperon, Jacques Droulez, and Guillaume Thibault. Reliable camera pose and calibration from a small set of point and line correspondences: a probabilistic approach. Computer Vision and Image Understanding: CVIU, 115(5): 576–585, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2014:SAG

[CE14] Zezhi Chen and Tim Ellis. A self-adaptive Gaussian mixture model. Computer Vision and Image Understanding: CVIU, 122(??): 35–46, May 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000113.

Chi:2017:ETD

[CE17] Jianning Chi and Mark Eramian. Enhancing textural differences using waveletbased texture characteristics morphological component analysis: a preprocessing method for improving image segmentation. Computer Vision and Image Understanding: CVIU, 158 (??):49–61, May 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300140.

Canal:2016:RTH

[CEA16] Gerard Canal, Sergio Escalera, and Cecilio Angulo. A real-time Human– Robot Interaction system based on gestures for assistive scenarios. Computer Vision and Image Understanding: CVIU, 149(??): 65–77, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421600076X.

Cevikalp:2018:LSI

[CEO18] Hakan Cevikalp, Merve Elmas, and Savas Ozkan. Large-scale image retrieval using transductive support vector machines. Computer Vision and Image Understanding: CVIU, 173(??):

2–12, August 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301364.

Campbell:2001:SFF

[CF01] Richard J. Campbell and Patrick J. Flynn. A survey of free-form object representation and recognition techniques. Computer Vision and Image Understanding: CVIU, 81(2): 166–210, February 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0889;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0889/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0889/ref.

Cao:2007:CCL

[CF07] Xiaochun Cao and Hassan Foroosh. Camera calibration and light source orientation from solar shadows. Computer Vision and Image Understanding: CVIU, 105(1):60–72, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cheong:1998:EEV

[CFA98] LoongFah Cheong, Cornelia Fermüller, and Yiannis Aloimonos. Effects of errors in the viewing geometry on shape estimation. Computer Vision and Image Understanding: CVIU, 71(3):356– 372, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0649/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0649/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0649/production/ref.

Chen:2005:IVL

[CFB05] Xin Chen, Patrick J. Flynn, and Kevin W. Bowyer. IR and visible light face recognition. Computer Vision and Image Understanding: CVIU, 99(3):332–358, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Canton-Ferrer:2011:HMC

[CFCP11] Cristian Canton-Ferrer, Josep R. Casas, and Montse Pardàs. Human motion capture using scalable body models. Computer Vision and Image Understanding: CVIU, 115(10):1363–1374, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421100138X.

Castellani:2002:RMA

[CFM02] Umberto Castellani, Andrea Fusiello, and Vittorio Murino. Registration of multiple acoustic range views for underwater scene reconstruction. Computer Vision and Image Understanding: CVIU, 87(1–3): 78–89, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chou:2013:SIR

[CFM^{+}13] Chen-Rui Chou, Brandon Frederick, Gig Mageras, Sha Chang, and Stephen Pizer. 2D/3D image registration using regression learning. Computer Vision and Image Understanding: CVIU, 117(9):1095– 1106, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000647.

Chopin:2023:MBI

[CFM^{+}23] Jeremy Chopin, Jean-Baptiste Fasquel, Harold Mouchère, Rozenn Dahyot, and Isabelle Bloch. Model-based inexact graph matching on top of DNNs for semantic scene understanding. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001248.

Camps:1998:RPC

[CFS98] Octavia Camps, Patrick J. Flynn, and George C. Stockman. Recent progress in CAD-based computer vision: An introduction to the special issue. Computer Vision and Image Understanding: CVIU, 69(3): 251–252, March 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0662/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0662/production/pdf.

Cappabianco:2012:BTM

[CFYU12] Fábio A. M. Cappabianco, Alexandre X. Falcão, Clarissa L. Yasuda, and Jayaram K. Udupa. Brain tissue MR-image segmentation via optimum-path forest clustering. Computer Vision and Image Understanding: CVIU, 116(10):1047– 1059, October 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000999.

Cortelazzo:2004:MBI

[CG04] Guido M. Cortelazzo and Concettina Guerra. Modelbased and image-based 3D scene representation for interactive visualization. Computer Vision and Image Understanding: CVIU, 96(3):269–273, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Caglioti:2009:RBM

[CG09] Vincenzo Caglioti and Alessandro Giusti. Recovering ball motion from a single motion-blurred image. Computer Vision and Image Understanding: CVIU, 113(5):590–597, May 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cai:2023:GML

[CGD^{+}23] Zhixi Cai, Shreya Ghosh, Abhinav Dhall, Tom Gedeon, Kalin Stefanov, and Munawar Hayat. Glitch in the matrix: a large scale benchmark for content driven audio-visual forgery detection and localization. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001984.

Caillette:2008:RTD

[CGH08] Fabrice Caillette, Aphrodite Galata, and Toby Howard. Real-time 3-D human body tracking using learnt models of behaviour. Computer Vision and Image Understanding: CVIU, 109(2):112–125, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chang:2016:STH

[CGHTK16] Hyung Jin Chang, Guillermo Garcia-Hernando, Danhang Tang, and Tae-Kyun Kim. Spatio-Temporal Hough Forest for efficient detection-localisation-recognition of fingerwriting in egocentric camera. Computer Vision and Image Understanding: CVIU, 148 (??):87–96, July 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216000357.

Carpenter:1998:WWF

[CGL98] G. A. Carpenter, S. Grossberg, and G. W. Lesher. The what-and-where filter. A spatial mapping neural network for object recognition and image understanding. Computer Vision and Image Understanding: CVIU, 69(1):1–??, ???? 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chellappa:2021:GEA

[CGL^{+}21] Rama Chellappa, Diego Gragnaniello, Chang-Tsun Li, Francesco Marra, and Richa Singh. Guest editorial: Adversarial deep learning in biometrics and forensics. Computer Vision and Image Understanding: CVIU, 208–209(??): ??, July 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221000710.

Chakraborty:2013:LSC

[CGR13] Bhaskar Chakraborty, Jordi Gonzàlez, and F. Xavier Roca. Large scale continuous visual event recognition using max-margin Hough transformation framework. Computer Vision and Image Understanding: CVIU, 117(10):1356–1368, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314212001920.

Cakir:2011:NNB

[CGU11] Fatih Cakir, Ugur Güdükbay, and Özgür Ulusoy. Nearestneighbor based metric functions for indoor scene recognition. Computer Vision and Image Understanding: CVIU, 115(11):1483–1492, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211001780.

Castano:1996:PAP

[CH96] Rebecca L. Castaño and Seth Hutchinson. A probabilistic approach to perceptual grouping. Computer Vision and Image Understanding: CVIU, 64(3):399– 419, November 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0068/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0068/production/pdf.

Christy:1999:IPC

[CH99] Stéphane Christy and Radu Horaud. Iterative pose computation from line correspondences. Computer Vision and Image Understanding: CVIU, 73(1):137– 144, January 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.

0717/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0717/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0717/production/ref.

Castelan:2006:AHD

[CH06] Mario Castelán and Edwin R. Hancock. Acquiring height data from a single image of a face using local shape indicators. Computer Vision and Image Understanding: CVIU, 103(1): 64–79, July 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Corso:2009:IDF

[CH09] J. J. Corso and G. D. Hager.

Image description with features that summarize. Computer Vision and Image Understanding: CVIU, 113(4): 446–458, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chesi:2011:FMV

[CH11] G. Chesi and Y. S. Hung.

Fast multiple-view L_{2} triangulation with occlusion handling. Computer Vision and Image Understanding: CVIU, 115(2):211–223, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2017:SHG

[CH17] Liang Chen and Negar Hassanpour. Survey: How good are the current advances in image set based face identification? — Experiments on three popular benchmarks with a naïve approach. Computer Vision and Image Understanding: CVIU, 160 (??):1–23, July 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300577.

Chan:2021:CCE

[Cha21] Yi-Tung Chan. Comprehensive comparative evaluation of background subtraction algorithms in open sea environments. Computer Vision and Image Understanding: CVIU, 202(??):Article 103101, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301284.

Chan:2024:ELB

[Cha24] Yi-Tung Chan. Ensemble learning-based method for maritime background subtraction in open sea environments. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002394.

Cheng:2011:ELD

[CHC11] Xiangang Cheng, Yiqun Hu, and Liang-Tien Chia. Exploiting local dependencies with spatial-scale space (S-Cube) for near-duplicate retrieval. Computer Vision and Image Understanding: CVIU, 115(6):750– 758, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cheng:1996:AAI

[Che96] Yu Cheng. Analysis of affine invariants as approximate perspective invariants. Computer Vision and Image Understanding: CVIU, 63(2):197– 207, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0014/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0014/production/pdf.

Chen:1998:HDP

[Che98] Yung-Sheng Chen. Hidden deletable pixel detection using vector analysis in parallel thinning to obtain bias-reduced skeletons. Computer Vision and Image Understanding: CVIU, 71(3):294–311, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0647/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0647/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0647/production/ref.

Cheong:2000:SBS

[Che00] Loong-Fah Cheong. Scenebased shot change detection and comparative evaluation. Computer Vision and Image Understanding: CVIU, 79(2):224–235, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0858;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0858/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0858/ref.

Chen:2008:FLR

[Che08] Hui Chen. Focal length and registration correction for building panorama from photographs. Computer Vision and Image Understanding: CVIU, 112(2):225–230, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2024:TCT

[Che24] Jingqiang Chen. Transform, contrast and tell: Coherent entity-aware multiimage captioning. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002588.

Celik:2009:UST

[CHH09] Hasan Celik, Alan Hanjalic, and Emile A. Hendriks. Unsupervised and simultaneous training of multiple object detectors from unlabeled surveillance video. Computer Vision and Image Understanding: CVIU, 113(10):1076–1094, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cutzu:2005:DPP

[CHL05] Florin Cutzu, Riad Hammoud, and Alex Leykin. Distinguishing paintings from photographs. Computer Vision and Image Understanding: CVIU, 100(3): 249–273, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2021:LLF

[CHL21] Jiamin Chen, Jianguo Hu, and Shiren Li. Learning to locate for fine-grained image recognition. Computer Vision and Image Understanding: CVIU, 206(??):Article 103184, May 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422100028X.

Chakraborty:2012:SST

[CHMG12] Bhaskar Chakraborty, Michael B. Holte, Thomas B. Moeslund, and Jordi Gonzàlez. Selective spatio-temporal interest points. Computer Vision and Image Understanding: CVIU, 116(3): 396–410, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002128.

Calderara:2011:DAP

[CHP^{+}11] Simone Calderara, Uri Heinemann, Andrea Prati, Rita Cucchiara, and Naftali Tishby. Detecting anomalies in people’s trajectories using spectral graph analysis. Computer Vision and Image Understanding: CVIU, 115(8):1099– 1111, August 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000919.

Cox:1996:MLS

[CHRM96] Ingemar J. Cox, Sunita L.

Hingorani, Satish B. Rao, and Bruce M. Maggs. A maximum likelihood stereo algorithm. Computer Vision and Image Understanding: CVIU, 63(3): 542–567, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0040/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0040/production/pdf.

Caselles:2008:GVM

[CHSV08] V. Caselles, G. Haro, G. Sapiro, and J. Verdera. On geometric variational models for inpainting surface holes. Computer Vision and Image Understanding: CVIU, 111(3):351–373, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chung:2002:RVD

[Chu02] Ronald Chung. Relative viewing distance: a correspondence invariance under paraperspective projection.

Computer Vision and Image Understanding: CVIU, 86(1):1–31, April 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2013:RFB

[CHZ^{+}13] X. Chen, H. He, G. Zou, X. Zhang, X. Gu, and J. Hua. Ricci flowbased spherical parameterization and surface registration. Computer Vision and Image Understanding: CVIU, 117(9):1107– 1118, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213000659.

Caglayan:2022:WCM

[CICN22] Ali Caglayan, Nevrez Imamoglu, Ahmet Burak Can, and Ryosuke Nakamura. When CNNs meet random RNNs: Towards multi-level analysis for RGB-D object and scene recognition. Computer Vision and Image Understanding: CVIU, 217 (??):??, March 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222000133.

Collins:1998:ASA

[CJC^{+}98] Robert T. Collins, Christopher O. Jaynes, Yong-Qing Cheng, Xiaoguang Wang, Frank Stolle, Edward M. Riseman, and Allen R. Hanson. The ascender system: Automated site modeling from multiple aerial images. Computer Vision and Image Understanding: CVIU, 72(2):143– 162, November 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0729/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0729/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0729/production/ref.

Cord:2001:ABS

[CJC01] Matthieu Cord, Michel Jordan, and Jean-Pierre Cocquerez. Accurate building structure recovery from high resolution aerial imagery. Computer Vision and Image Understanding: CVIU, 82(2):138–173, May 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0905;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0905/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.

                                        2001.0905/ref.

Chin:2006:HSI

[CJL06] Chong Siew Chin, Andrew Teoh Beng Jin, and David Ngo Chek Ling. High security Iris verification system based on random secret integration. Computer Vision and Image Understanding: CVIU, 102(2):169– 177, May 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cui:2022:FCO

[CJWW22] Yutao Cui, Cheng Jiang, Limin Wang, and Gangshan Wu. Fully convolutional online tracking. Computer Vision and Image Understanding: CVIU, 224(??): ??, November 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001254.

Chen:2023:SCE

[CJYW23] Yadang Chen, Chuanjun Ji, Zhi-Xin Yang, and Enhua Wu. Spatial constraint for efficient semisupervised video object segmentation. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002230.

Choi:2009:RFP

[CK09] Ouk Choi and In So Kweon.

Robust feature point matching by preserving local geometric consistency. Computer Vision and Image Understanding: CVIU, 113(6): 726–742, June 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chang:2011:MSS

[CK11] Ming-Ching Chang and Benjamin B. Kimia. Measuring 3D shape similarity by graph-based matching of the medial scaffolds. Computer Vision and Image Understanding: CVIU, 115(5):707–720, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Califano:1996:DMD

[CKB96] Andrea Califano, Rick Kjeldsen, and Ruud M. Bolle. Data- and modeldriven multiresolution processing. Computer Vision and Image Understanding: CVIU, 63(1):27– 49, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0003/production;   http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0003/production/pdf.

Chai:2010:SIE

[CKB10] Sek Chai, Branislav Kisačanin, and Nikolaos Bellas. Special issue on embedded vision. Computer Vision and Image Understanding: CVIU, 114 (11):1115, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Choi:2014:CHW

[CKC14] Hong Seok Choi, In Su Kim, and Jin Young Choi. Combining histogram-wise and pixel-wise matchings for kernel tracking through constrained optimization. Computer Vision and Image Understanding: CVIU, 118 (??):61–70, January 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314213001720.

Chen:2018:RSA

[CKF18] Lei Chen, Wenyue Kuang, and Kun Fu. A resample strategy and artificial bee colony optimizationbased 3d range imaging registration. Computer Vision and Image Understanding: CVIU, 175(??):44–51, October 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218302601.

Chun:2012:ACM

[CKK^{+}12] Jinhee Chun, Natsuda Kaothanthong, Ryosei Kasai, Matias Korman, Martin Nöllenburg, and Takeshi Tokuyama. Algorithms for computing the maximum weight region decomposable into elementary shapes. Computer Vision and Image Understanding: CVIU, 116 (7):803–814, July 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000501.

Choi:2018:RTV

[CKL18] Janghoon Choi, Junseok Kwon, and Kyoung Mu Lee. Real-time visual tracking by deep reinforced decision making. Computer Vision and Image Understanding: CVIU, 171(??): 10–19, June 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300754.

Chiu:2009:LGN

[CKLP09] Han-Pang Chiu, Leslie Pack Kaelbling, and Tomás Lozano-Pérez. Learning to generate novel views of objects for class recognition. Computer Vision and Image Understanding: CVIU, 113(12): 1183–1197, December 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Castle:2011:WAA

[CKM11] Robert O. Castle, Georg Klein, and David W. Murray. Wide-area augmented reality using camera tracking and mapping in multiple regions. Computer Vision and Image Understanding: CVIU, 115(6):854– 867, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cho:2019:JPR

[CKP^{+}19] Yeong-Jun Cho, Su-A Kim, Jae-Han Park, Kyuewang Lee, and Kuk-Jin Yoon. Joint person re-identification and camera network topology inference in multiple cameras. Computer Vision and Image Understanding: CVIU, 180(??):34– 46, March 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219300037.

Chelali:2021:DSC

[CKPV21] Mohamed Chelali, Camille Kurtz, Anne Puissant, and Nicole Vincent. Deep-STaR:

Classification of image time series based on spatiotemporal representations. Computer Vision and Image Understanding: CVIU, 208– 209(??):??, July 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000655.

Criminisi:2005:ELA

[CKS^{+}05] Antonio Criminisi, Sing Bing Kang, Rahul Swaminathan, Richard Szeliski, and P. Anandan. Extracting layers and analyzing their specular properties using epipolarplane-image analysis. Computer Vision and Image Understanding: CVIU, 97(1): 51–85, January 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chung:1995:HTR

[CL95] Kuo-Liang Chung and Horn-Yi Lin. Hough transform on reconfigurable meshes. Computer Vision and Image Understanding: CVIU, 61(2):278– 284, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1020/production;   http:
//www.idealibrary.com/

links/artid/cviu.1995.
1020/production/pdf.

Crevier:1997:KBI

[CL97] Daniel Crevier and Richard Lepage. Knowledge-based image understanding systems: a survey. Computer Vision and Image Understanding: CVIU, 67(2):161– 185, August 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0520/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0520/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0520/production/ref.

Chen:2000:SEC

[CL00] Fei-Long Chen and Shiaur-Wehn Lin. Subpixel estimation of circle parameters using orthogonal circular detector. Computer Vision and Image Understanding: CVIU, 78(2): 206–221, May 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0836;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0836/pdf;   http:
//www.idealibrary.com/

links/doi/10.1006/cviu.
2000.0836/ref.

Chi:2008:GSC

[CL08] Yanling Chi and Maylor K. H. Leung. A general shape context framework for object identification. Computer Vision and Image Understanding: CVIU, 112(3):324–336, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chang:2015:LML

[CL15] Ju Yong Chang and Kyoung Mu Lee. Large margin learning of hierarchical semantic similarity for image classification. Computer Vision and Image Understanding: CVIU, 132(??): 3–11, March 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002264.

Chen:2017:WSC

[CL17] Jianhui Chen and James J.

Little. Where should cameras look at soccer games: Improving smoothness using the overlapped hidden Markov model. Computer Vision and Image Understanding: CVIU, 159(??): 59–73, June 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301709.

Chang:2018:PCB

[CL18] Ju Yong Chang and Kyoung Mu Lee. 2D–3D pose consistency-based conditional random fields for 3D human pose estimation. Computer Vision and Image Understanding: CVIU, 169 (??):52–61, April 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421830016X.

Christie:2017:RVL

[CLA^{+}17] Gordon Christie, Ankit Laddha, Aishwarya Agrawal, Stanisław Antol, Yash Goyal, Kevin Kochersberger, and Dhruv Batra. Resolving vision and language ambiguities together: Joint segmentation & prepositional attachment resolution in captioned scenes. Computer Vision and Image Understanding: CVIU, 163(??): 101–112, October 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301571.

Cho:2013:EGG

[CLCO13] Jungchan Cho, Minsik Lee, Chong-Ho Choi, and Songhwai Oh. EM–GPA: Generalized Procrustes analysis with hidden variables for 3D shape modeling. Computer Vision and Image Understanding: CVIU, 117(11):1549–1559, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001392.

Cha:2019:DPC

[CLCO19] Geonho Cha, Minsik Lee, Jungchan Cho, and Songhwai Oh. Deep pose consensus networks. Computer Vision and Image Understanding: CVIU, 182(??): 64–70, May 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218301462.

Cheng:2023:SMG

[CLCZ23] Bodong Cheng, Juncheng Li, Ying Chen, and Tieyong Zeng. Snow mask guided adaptive residual network for image snow removal. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001996.

Chen:1996:PVM

[CLD96] Yuan Chen, Noshir A. Langrana, and Atish K. Das. Perfecting vectorized mechanical drawings. Computer Vision and Image Understanding: CVIU, 63(2): 273–286, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0019/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0019/production/pdf.

Cheng:2023:PPC

[CLDP23] Xu Cheng, Haoyuan Li, Shuya Deng, and Yonghong Peng. POEM: a prototype cross and emphasis network for few-shot semantic segmentation. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001261.

Chen:2022:MVT

[CLFH22] Haoran Chen, Jianmin Li, Simone Frintrop, and Xiaolin Hu. The MSR-Video to Text dataset with clean annotations. Computer Vision and Image Understanding: CVIU, 225(??):

??, December 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422200159X.

Chang:2009:SRP

[CLK09] Ming-Ching Chang, Frederic Fol Leymarie, and Benjamin B. Kimia. Surface reconstruction from point clouds by transforming the medial scaffold. Computer Vision and Image Understanding: CVIU, 113(11): 1130–1146, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cheng:2014:SSM

[CLL^{+}14a] Jian Cheng, Cong Leng, Peng Li, Meng Wang, and Hanqing Lu. Semisupervised multi-graph hashing for scalable similarity search. Computer Vision and Image Understanding: CVIU, 124(??):12– 21, July 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214000794.

Coeurjolly:2014:MCP

[CLL14b] David Coeurjolly, Jacques-Olivier Lachaud, and Jérémy Levallois. Multigrid convergent principal curvature estimators in digital geometry. Computer Vision and Image Understanding: CVIU, 129(??):27–41, December 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001003.

Cheng:2017:NPD

[CLL17] Ziang Cheng, Yang Liu, and Guojun Liu. A new primal-dual algorithm for multilabel graph-cuts problems with approximate moves. Computer Vision and Image Understanding: CVIU, 165(??):75–84, December 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301340.

Cao:2021:FSA

[CLL^{+}21] Congqi Cao, Yajuan Li, Qinyi Lv, Peng Wang, and Yanning Zhang. Few-shot action recognition with implicit temporal alignment and pair similarity optimization. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000941.

Cho:2017:SIH

[CLO17] Jungchan Cho, Minsik Lee, and Songhwai Oh. Single image 3D human pose estimation using a procrustean normal distribution mixture model and model transformation. Computer Vision and Image Understanding: CVIU, 155(??):150– 161, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301734.

Chen:2023:SSC

[CLTW23] Chaofeng Chen, Wei Liu, Xiao Tan, and Kwan-Yee K. Wong. Semisupervised Cycle-GAN for face photo-sketch translation in the wild. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001558.

Canavan:2015:LLR

[CLZY15] Shaun Canavan, Peng Liu, Xing Zhang, and Lijun Yin. Landmark localization on 3D/4D range data using a shape index-based statistical shape model with global and local constraints. Computer Vision and Image Understanding: CVIU, 139(??):136– 148, October 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001344.

Chen:2013:OCI

[CLZZ13] Xiaowu Chen, Qing Li, Dongyue Zhao, and Qinping Zhao. Occlusion cues for image scene layering. Computer Vision and Image Understanding: CVIU, 117(1): 42–55, January 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001300.

Chen:2021:MBD

[CLZZ21] Xiuhong Chen, Yun Lu, Jun Zhang, and Xingyu Zhu. Margin-based discriminant embedding guided sparse matrix regression for image supervised feature selection. Computer Vision and Image Understanding: CVIU, 212(??): ??, November 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422100117X.

Chen:1995:DCO

[CM95] Yang Chen and Gérard Medioni. Description of complex objects from multiple range images using an inflating balloon model. Computer Vision and Image Understanding: CVIU, 61 (3):325–334, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1026/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1026/production/pdf.

Cho:1997:ISC

[CM97] Kyujin Cho and Peter Meer. Image segmentation from consensus information. Computer Vision and Image Understanding: CVIU, 68(1):72– 89, October 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0546/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0546/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0546/production/ref.

Casadei:1999:BUA

[CM99a] Stefano Casadei and Sanjoy Mitter. Beyond the uniqueness assumption: Ambiguity representation and redundancy elimination in the computation of a covering sample of salient contour cycles. Computer Vision and Image Understanding: CVIU, 76(1):19– 35, October 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0790/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0790/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0790/production/ref.

Cuisenaire:1999:FED

[CM99b] O. Cuisenaire and B. Macq. Fast Euclidean distance transformation by propagation using multiple neighborhoods. Computer Vision and Image Understanding: CVIU, 76(2):163– 172, November 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0783/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0783/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0783/production/ref.

Cao:2012:IFE

[CM12] Yanpeng Cao and John Mc-Donald. Improved feature extraction and matching in urban environments based on 3D viewpoint normalization. Computer Vision and Image Understanding: CVIU, 116(1):86–101, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211001974.

Chakraborty:2016:DSB

[CM16] Souradeep Chakraborty and Pabitra Mitra. A dense subgraph based algorithm for compact salient image region detection. Computer Vision and Image Understanding: CVIU, 145(??): 1–14, April 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002696.

Cheema:2021:SFD

[CM21] Usman Cheema and Seungbin Moon. Sejong face database: a multi-modal disguise face database. Computer Vision and Image Understanding: CVIU, 208– 209(??):??, July 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422100062X.

Cao:2009:HMR

[CMBP09] Dongwei Cao, Osama T. Masoud, Daniel Boley, and Nikolaos Papanikolopoulos. Human motion recognition using support vector machines. Computer Vision and Image Understanding: CVIU, 113(10):1064–1075, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Calderon:2004:MMA

[CMBV04] Felix Calderon, Jose L. Marroquin, Salvador Botello, and Baba C. Vemuri. The MPM–MAP algorithm for motion segmentation. Computer Vision and Image Understanding: CVIU, 95(2): 165–183, August 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cormier:2016:PVB

[CMCM16] Michael Cormier, Karyn Moffatt, Robin Cohen, and Richard Mann. Purely vision-based segmentation of web pages for assistive technology. Computer Vision and Image Understanding: CVIU, 148(??): 46–66, July 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216000527.

Christoudias:2006:NPL

[CMD06] C. Mario Christoudias, Louis-Philippe Morency, and Trevor Darrell. Nonparametric and light-field deformable models. Computer Vision and Image Understanding: CVIU, 104(1): 16–35, October 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cuevas:2016:DSF

[CMG16] Carlos Cuevas, Raquel Martínez, and Narciso García. Detection of stationary foreground objects: a survey. Computer Vision and Image Understanding: CVIU, 152(??):41–57, November 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216300972.

Cheriet:2013:LFO

[CMH13] Mohamed Cheriet, Reza Farrahi Moghaddam, and Rachid Hedjam. A learning framework for the optimization and automation of document binarization methods. Computer Vision and Image Understanding: CVIU, 117(3):269– 280, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001877.

Caracotte:2020:PSC

[CMM20] Jordan Caracotte, Fabio Morbidi, and El Mustapha Mouaddib. Photometric stereo with central panoramic cameras. Computer Vision and Image Understanding: CVIU, 201 (??):Article 103080, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314220301144.

Chlubna:2024:LAF

[CMZ24] Tomás Chlubna, Tomás Milet, and Pavel Zemcík. Lightweight all-focused light field rendering. Computer Vision and Image Understanding: CVIU, 244(??): ??, July 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314224001127.

Chung:1995:UMG

[CN95] Ronald Chung and Ramakant Nevatia. Use of monocular groupings and occlusion analysis in a hierarchical stereo system. Computer Vision and Image Understanding: CVIU, 62(3):245–268, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1053/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1053/production/pdf.

Cunado:2003:AED

[CNC03] David Cunado, Mark S.

Nixon, and John N. Carter. Automatic extraction and description of human gait models for recognition purposes. Computer Vision and Image Understanding: CVIU, 90(1):1–41, April 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cousty:2013:MFG

[CNDS13] Jean Cousty, Laurent Najman, Fabio Dias, and Jean Serra. Morphological filtering on graphs. Computer Vision and Image Understanding: CVIU, 117(4): 370–385, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421200183X.

Chessa:2016:IAV

[CNO^{+}16] Manuela Chessa, Nicoletta Noceti, Francesca Odone, Fabio Solari, Joan Sosa-García, and Luca Zini. An integrated artificial vision framework for assisting visually impaired users. Computer Vision and Image Understanding: CVIU, 149(??):209– 228, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002519.

Cheraghalikhani:2024:SAF

[CNO^{+}24] Milad Cheraghalikhani, Mehrdad Noori, David Osowiechi, Gustavo A. Vargas Hakim, Ismail Ben Ayed, and Christian Desrosiers. Structureaware feature stylization for domain generalization. Computer Vision and Image Understanding: CVIU, 244(??):??, July 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000973.

Ciocca:2018:CBF

[CNS18] Gianluigi Ciocca, Paolo Napoletano, and Raimondo Schettini. CNN-based features for retrieval and classification of food images. Computer Vision and Image Understanding: CVIU, 176–177(??):70–77, November/December 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218302467.

Choi:2016:SBA

[CO16] Kang-Sun Choi and Ki-Won Oh. Subsampling-based acceleration of simple linear iterative clustering for superpixel segmentation. Computer Vision and Image Understanding: CVIU, 146 (??):1–8, May 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000722.

Celik:2008:FFE

[ÇÖD08] Turgay Çelik, Hüseyin Özkaramanlı, and Hasan Demirel. Facial feature extraction using complex dual-tree wavelet transform. Computer Vision and Image Understanding: CVIU, 111 (2):229–246, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Coeurjolly:2012:FAA

[Coe12] David Coeurjolly. Fast and accurate approximation of digital shape thickness distribution in arbitrary dimension. Computer Vision and Image Understanding: CVIU, 116(12):1159– 1167, December 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001178.

Collins:1997:GVS

[Col97] Robert T. Collins. The geometry of visual space: About the incompatibility between science and mathematics — reply. Computer Vision and Image Understanding: CVIU, 65(3): 434–435, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0493/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0493/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0493/production/ref. See

[PRW97a].

Couprie:2013:TMR

[Cou13] Michel Couprie. Topological maps and robust hierarchical Euclidean skeletons in cubical complexes. Computer Vision and Image Understanding: CVIU, 117(4): 355–369, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001762.

Chu:2022:STV

[COV^{+}22] Chenhui Chu, Vinicius Oliveira, Felix Giovanni Virgo, Mayu Otani, Noa Garcia, and Yuta Nakashima. The semantic typology of visually grounded paraphrases. Computer Vision and Image Understanding: CVIU, 215(??): ??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001697.

Caelli:1998:SMI

[COW98] Terry Caelli, Erol Osman, and Geoff West. 3D shape matching and inspection using geometric features and relational learning. Computer Vision and Image Understanding: CVIU, 72(3):340–350, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0659/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0659/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0659/production/ref.

Cheong:2004:DDU

[CP04] Loong-Fah Cheong and Chin-Hwee Peh. Depth distortion under calibration uncertainty. Computer Vision and Image Understanding: CVIU, 93(3):221– 244, March 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cho:2009:FAD

[CP09] Minkook Cho and Hyeyoung Park. A feature analysis for dimension reduction based on a data generation model with class factors and environment factors. Computer Vision and Image Understanding: CVIU, 113(9): 1005–1016, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Combes:2020:EEI

[CP20] Benoit Combès and Sylvain Prima. An efficient EM-ICP algorithm for nonlinear registration of large 3D point sets. Computer Vision and Image Understanding: CVIU, 191(??):Article 102854, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219301481.

Chatterjee:2021:SSR

[CP21] Bodhiswatta Chatterjee and Charalambos Poullis. Semantic segmentation from remote sensor data and the exploitation of latent learning for classification of auxiliary tasks. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000953.

Chan:1999:BSR

[CPC99] Moses W. Chan, Zygmunt Pizlo, and David M. Chelberg. Binocular shape reconstruction: Psychological plausibility of the 8point algorithm. Computer Vision and Image Understanding: CVIU, 74(2): 121–137, May 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0748/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0748/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0748/production/ref.

Calderara:2008:HHE

[CPC08] Simone Calderara, Andrea Prati, and Rita Cucchiara. HECOL: Homography and epipolar-based consistent labeling for outdoor park surveillance. Computer Vision and Image Understanding: CVIU, 111(1):21– 42, July 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Crivelli:2016:VOT

[CPO16] Tomas Crivelli, Patrick Perez, and Lionel Oisel. Visual object trapping. Computer Vision and Image Understanding: CVIU, 153 (??):3–15, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314216301059.

Chang:2011:GFM

[CPP^{+}11] Ju Yong Chang, Haesol Park, In Kyu Park, Kyoung Mu Lee, and Sang Uk Lee. GPU-friendly multiview stereo reconstruction using surfel representation and graph cuts. Computer Vision and Image Understanding: CVIU, 115(5): 620–634, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2021:SAH

[CPPY21] Yun-Chun Chen, Marco Piccirilli, Robinson Piramuthu, and Ming-Hsuan Yang. Self-attentive 3D human pose and shape estimation from videos. Computer Vision and Image Understanding: CVIU, 213(??): ??, December 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001491.

Chum:2005:GEH

[CPS05] Ondřej Chum, Tomáš Pajdla, and Peter Sturm. The geometric error for homographies. Computer Vision and Image Understanding: CVIU, 97(1):86–102, January 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Craciun:2010:MVS

[CPS10] Daniela Craciun, Nicolas Paparoditis, and Francis Schmitt. Multi-view scans alignment for 3D spherical mosaicing in large-scale unstructured environments. Computer Vision and Image Understanding: CVIU, 114(11):1248–1263, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cheng:2007:MSM

[CPT07] Shinko Y. Cheng, Sangho Park, and Mohan M. Trivedi. Multi-spectral and multi-perspective video arrays for driver body tracking and activity analysis. Computer Vision and Image Understanding: CVIU, 106(2– 3):245–257, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chui:2003:NPM

[CR03] Haili Chui and Anand Rangarajan. A new point matching algorithm for nonrigid registration. Computer Vision and Image Understanding: CVIU, 89(2–3): 114–141, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Connor:2018:BRG

[CR18] Patrick Connor and Arun Ross. Biometric recognition by gait: A survey of modalities and features. Computer Vision and Image Understanding: CVIU, 167(??): 1–27, February 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300079.

Chatterjee:1997:NGS

[CRC97] Chanchal Chatterjee, Vwani P. Roychowdhury, and Edwin K. P. Chong. A nonlinear Gauss–Seidel algorithm for noncoplanar and coplanar camera calibration with convergence analysis. Computer Vision and Image Understanding: CVIU, 67(1): 58–80, July 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0516/production;   http:

//www.idealibrary.com/
links/artid/cviu.1997.
0516/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0516/production/ref.

Conze:2016:MRC

[CRCM16] Pierre-Henri Conze, Philippe Robert, Tomás Crivelli, and Luce Morin. Multi-reference combinatorial strategy towards longer long-term dense motion estimation. Computer Vision and Image Understanding: CVIU, 150 (??):66–80, September 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314216300340.

Crevier:1999:PME

[Cre99] Daniel Crevier. A probabilistic method for extracting chains of collinear segments. Computer Vision and Image Understanding: CVIU, 76(1):36– 53, October 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0785/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0785/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0785/production/ref;

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0787/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0787/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0787/production/ref.

Crevier:2008:ISA

[Cre08] Daniel Crevier. Image segmentation algorithm development using ground truth image data sets. Computer Vision and Image Understanding: CVIU, 112(2): 143–159, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:1998:FFO

[CS98] Jin-Long Chen and George C. Stockman. 3D free-form object recognition using indexing by contour features. Computer Vision and Image Understanding: CVIU, 71(3):334–355, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0648/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0648/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0648/production/ref.

Costa:2000:ORP

[CS00] Mauro S. Costa and Linda G.

Shapiro. 3D object recognition and pose with relational indexing. Computer Vision and Image Understanding: CVIU, 79(3): 364–407, September 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0865;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0865/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0865/ref.

Chau:2004:ADP

[CS04] Chun-Pong Chau and Wan-Chi Siu. Adaptive dualpoint Hough transform for object recognition. Computer Vision and Image Understanding: CVIU, 96(1): 1–16, October 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Coughlan:2007:DQB

[CS07] James Coughlan and Huiying Shen. Dynamic quantization for belief propagation in sparse spaces. Computer Vision and Image Understanding: CVIU, 106(1): 47–58, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2010:PEM

[CS10] Chong Chen and Dan Schonfeld. Pose estimation from multiple cameras based on Sylvester’s equation. Computer Vision and Image Understanding: CVIU, 114(6):652– 666, June 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Coeurjolly:2020:EDT

[CS20] David Coeurjolly and Isabelle Sivignon. Efficient distance transformation for path-based metrics. Computer Vision and Image Understanding: CVIU, 194(??):Article 102925, May 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220300175.

Chaudhuri:1996:REM

[CSC96] S. Chaudhuri, S. Sharma, and S. Chatterjee. Recursive estimation of motion parameters. Computer Vision and Image Understanding: CVIU, 64(3):434– 442, November 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/

links/artid/cviu.1996.
0070/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0070/production/pdf.

Cao:2024:ALA

[CSD^{+}24] Yiwen Cao, Yukun Su, Jingliang Deng, Yu Zhang, and Qingyao Wu. Adaptive locally-aligned transformer for low-light video enhancement. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002965.

Castrillon-Santana:2017:MTO

[CSDNR17] Modesto Castrillón-Santana, Maria De Marsico, Michele Nappi, and Daniel Riccio. MEG: Texture operators for multi-expert gender classification. Computer Vision and Image Understanding: CVIU, 156(??): 4–18, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301382.

Cohen:2003:FER

[CSG^{+}03] Ira Cohen, Nicu Sebe, Ashutosh Garg, Lawrence S. Chen, and Thomas S. Huang. Facial expression recognition from video sequences: temporal and static modeling. Computer Vision and Image Understanding: CVIU, 91(1–2): 160–187, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Casao:2024:DMT

[CSGM^{+}24] Sara Casao, Álvaro Serra-Gómez, Ana C. Murillo, Wendelin Böhmer, Javier Alonso-Mora, and Eduardo Montijano. Distributed multi-target tracking and active perception with mobile camera networks. Computer Vision and Image Understanding: CVIU, 238 (??):??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002564.

Csonde:2022:ORT

[CSK22] Gergely Csönde, Yoshihide Sekimoto, and Takehiro Kashiyama. Online realtime pedestrian tracking from medium altitude aerial footage with camera motion cancellation. Computer Vision and Image Understanding: CVIU, 217 (??):??, March 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000236.

Chen:2016:PBM

[CSLX16] Jiahui Chen, Hao Sheng, Chao Li, and Zhang Xiong. PSTG-based multi-label optimization for multi-target tracking. Computer Vision and Image Understanding: CVIU, 144(??):217– 227, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001307.

Ciesielski:2014:EAF

[CSMS14] Krzysztof Chris Ciesielski, Robin Strand, Filip Malmberg, and Punam K. Saha. Efficient algorithm for finding the exact minimum barrier distance. Computer Vision and Image Understanding: CVIU, 123(??): 53–64, June 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421400068X.

Chamveha:2013:HDE

[CSS^{+}13a] Isarun Chamveha, Yusuke Sugano, Daisuke Sugimura, Teera Siriteerakul, Takahiro Okabe, Yoichi Sato, and Akihiro Sugimoto. Head direction estimation from low resolution images with scene adaptation. Computer Vision and Image Understanding: CVIU, 117(10):1502– 1511, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001094.

Chessa:2013:ALM

[CSS13b] Manuela Chessa, Fabio Solari, and Silvio P. Sabatini. Adjustable linear models for optic flow based obstacle avoidance. Computer Vision and Image Understanding: CVIU, 117(6): 603–619, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000283.

Cassidy:2016:EVT

[CSV^{+}16] S. A. Cassidy, B. Stenger, L. Van Dongen, K. Yanagisawa, R. Anderson, V. Wan, S. Baron-Cohen, and R. Cipolla. Expressive visual text-to-speech as an assistive technology for individuals with autism spectrum conditions. Computer Vision and Image Understanding: CVIU, 148(??):193– 200, July 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215001861.

Chen:2024:NCC

[CSW^{+}24] Zhangfei Chen, Xuelong Si, Dan Wu, Fengnian Tian, Zhenxing Zheng, and Renfu Li. A novel camera calibration method based on known rotations and translations. Computer Vision and Image Understanding: CVIU, 243(??): ??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000778.

Cao:2008:RIC

[CSY08] Aize Cao, Qing Song, and Xulei Yang. Robust information clustering incorporating spatial information for breast mass detection in digitized mammograms. Computer Vision and Image Understanding: CVIU, 109(1):86–96, January 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cheng:2015:SDA

[CSZ^{+}15] Dong Seon Cheng, Francesco Setti, Nicola Zeni, Roberta Ferrario, and Marco Cristani. Semantically-driven automatic creation of training sets for object recognition. Computer Vision and Image Understanding: CVIU, 131(??):56–71, February 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421400157X.

Celik:2010:UCI

[CT10] Turgay Celik and Tardi Tjahjadi. Unsupervised colour image segmentation using dual-tree complex wavelet transform. Computer Vision and Image Understanding: CVIU, 114(7): 813–826, July 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Celik:2012:ACC

[CT12] Turgay Celik and Tardi Tjahjadi. Adaptive colour constancy algorithm using discrete wavelet transform. Computer Vision and Image Understanding: CVIU, 116(4):561– 571, April 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211002712.

Choudhury:2013:GRB

[CT13] Sruti Das Choudhury and Tardi Tjahjadi. Gait recognition based on shape and motion analysis of silhouette contours. Computer Vision and Image Understanding: CVIU, 117(12):1770– 1785, December 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001537.

Cootes:1995:ASM

[CTCG95] T. F. Cootes, C. J. Taylor, D. H. Cooper, and J. Graham. Active shape models — their training and application. Computer Vision and Image Understanding: CVIU, 61(1):38– 59, January 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1004/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1004/production/pdf.

Ching:1995:RVC

[CTE95] Wee-Soon Ching, Peng-Seng Toh, and Meng-Hwa Er. Robust vergence with concurrent identification of occlusion and specular highlights. Computer Vision and Image Understanding: CVIU, 62(3):298– 308, November 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1056/production;   http:
//www.idealibrary.com/

                                        links/artid/cviu.1995.
                                        1056/production/pdf.

Chen:1998:EAS

[CTF^{+}98] S. W. Chen, S. T. Tung, C. Y. Fang, Shen Cherng, and Anil K. Jain. Extended attributed string matching for shape recognition. Computer Vision and Image Understanding: CVIU, 70(1): 36–50, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0599/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0599/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0599/production/ref.

Chen:2020:MHP

[CTH20] Yucheng Chen, Yingli Tian, and Mingyi He. Monocular human pose estimation: a survey of deep learning-based methods. Computer Vision and Image Understanding: CVIU, 192(??):Article 102897, March 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219301778.

Cerutti:2013:ULN

[CTM^{+}13] Guillaume Cerutti, Laure Tougne, Julien Mille, Antoine Vacavant, and Didier Coquin. Understanding leaves in natural images — a model-based approach for tree species identification. Computer Vision and Image Understanding: CVIU, 117(10):1482– 1501, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001331.

Chen:2015:FPG

[CTWH15] Wei Chen, Yonghong Tian, Yaowei Wang, and Tiejun Huang. Fixed-point Gaussian Mixture Model for analysis-friendly surveillance video coding. Computer Vision and Image Understanding: CVIU, 142 (??):65–79, ???? 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002015.

Ciesielski:2010:AFFa

[CU10a] Krzysztof Chris Ciesielski and Jayaram K. Udupa. Affinity functions in fuzzy connectedness based image segmentation I: Equivalence of affinities. Computer Vision and Image Understanding: CVIU, 114(1):146–154, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ciesielski:2010:AFFb

[CU10b] Krzysztof Chris Ciesielski and Jayaram K. Udupa. Affinity functions in fuzzy connectedness based image segmentation II: Defining and recognizing truly novel affinities. Computer Vision and Image Understanding: CVIU, 114(1):155–166, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ciesielski:2011:FCD

[CU11] Krzysztof Chris Ciesielski and Jayaram K. Udupa. A framework for comparing different image segmentation methods and its use in studying equivalences between level set and fuzzy connectedness frameworks. Computer Vision and Image Understanding: CVIU, 115(6):721–734, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chudasama:2020:PEP

[CU20] Vishal Chudasama and Kishor Upla. E-ProSRNet: an enhanced progressive single image super-resolution approach. Computer Vision and Image Understanding: CVIU, 200(??):Article 103038, November 2020.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220300898.

Chen:2013:GAS

[CUAT13] Xinjian Chen, Jayaram K.

Udupa, Abass Alavi, and Drew A. Torigian. GC-ASM: Synergistic integration of graph-cut and active shape model strategies for medical image segmentation. Computer Vision and Image Understanding: CVIU, 117(5):513– 524, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001956.

Ciesielski:2007:IRF

[CUSZ07] Krzysztof Chris Ciesielski, Jayaram K. Udupa, Punam K. Saha, and Ying Zhuge. Iterative relative fuzzy connectedness for multiple objects with multiple seeds. Computer Vision and Image Understanding: CVIU, 107(3):160–182, September 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chellappa:2013:SSC

[CV13] Rama Chellappa and Baba Vemuri. Special section in celebration of Professor J. K. Aggarwal. Computer Vision and Image Understanding: CVIU, 117(10): 1203, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001410.

Casares:2010:LWS

[CVP10] Mauricio Casares, Senem Velipasalar, and Alvaro Pinto. Light-weight salient foreground detection for embedded smart cameras. Computer Vision and Image Understanding: CVIU, 114(11):1223–1237, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cui:2000:ABH

[CW00] Yuntao Cui and Juyang Weng. Appearance-based hand sign recognition from intensity image sequences. Computer Vision and Image Understanding: CVIU, 78(2):157–176, May 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0837;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0837/pdf;   http:
//www.idealibrary.com/

links/doi/10.1006/cviu.
2000.0837/ref.

Chang:2015:ODM

[CW15] Haw-Shiuan Chang and Yu-Chiang Frank Wang. Optimizing the decomposition for multiple foreground cosegmentation. Computer Vision and Image Understanding: CVIU, 141(??): 18–27, December 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001320.

Chen:2020:JSN

[CWC^{+}20] Beijing Chen, Yunqing Wu, Gouenou Coatrieux, Xiao Chen, and Yuhui Zheng. JSNet: a simulation network of JPEG lossy compression and restoration for robust image watermarking against JPEG attack. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 103015, August 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220300783.

Cao:2013:UDL

[CWH^{+}13] Xiaochun Cao, Xingxing Wei, Yahong Han, Yi Yang, Nicu Sebe, and Alexander Hauptmann. Unified dictionary learning and region tagging with hierarchical sparse representation. Computer Vision and Image Understanding: CVIU, 117(8): 934–946, August 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000623.

Cheng:2023:CCP

[CWL^{+}23] Cheng Cheng, Hang Wang, Xiang Liao, Gang Cheng, and Hongbin Sun. CPNet: Continuity Preservation Network for infrared video colorization. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223001960.

Choi:2013:STP

[CWLJ13] Jaesik Choi, Ziyu Wang, Sang-Chul Lee, and Won J. Jeon. A spatio-temporal pyramid matching for video retrieval. Computer Vision and Image Understanding: CVIU, 117(6):660– 669, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000325.

Chan:2022:PES

[CWLY22] Patrick P. K. Chan, Xiaotian Wang, Zhe Lin, and Daniel S. Yeung. Progressive editing with stacked generative adversarial network for multiple facial attribute editing. Computer Vision and Image Understanding: CVIU, 217(??): ??, March 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221001776.

Chen:2011:BPS

[CWO^{+}11] Shi Chen, Jinqiao Wang, Yi Ouyang, Bo Wang, Changsheng Xu, and Hanqing Lu. Boosting partsense multi-feature learners toward effective object detection. Computer Vision and Image Understanding: CVIU, 115(3):364– 374, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2024:FTF

[CWS^{+}24] Xu Chen, Lei Wu, Yongliang Su, Lei Meng, and Xiangxu Meng. Font transformer for few-shot font generation. Computer Vision and Image Understanding: CVIU, 245 (??):??, August 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224001243.

Cheng:2022:AFD

[CWW^{+}22] Rui Cheng, Yuzhe Wu, Jia Wang, Mingming Ma, Yi Niu, and Guangming Shi. Adaptive feature denoising based deep convolutional network for single image super-resolution. Computer Vision and Image Understanding: CVIU, 223 (??):??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001060.

Cheong:2011:BSA

[CX11] Loong-Fah Cheong and Xu Xiang. Behaviour of SFM algorithms with erroneous calibration. Computer Vision and Image Understanding: CVIU, 115(1):16– 30, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cao:2006:SCT

[CXFS06] Xiaochun Cao, Jiangjian Xiao, Hassan Foroosh, and Mubarak Shah. Selfcalibration from turn-table sequences in presence of zoom and focus. Computer Vision and Image Understanding: CVIU, 102(3): 227–237, June 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2024:MII

[CXYZ24] Yuantao Chen, Runlong Xia, Kai Yang, and Ke Zou. MFMAM: Image inpainting via multi-scale feature module with attention module. Computer Vision and Image Understanding: CVIU, 238 (??):??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002631.

Choi:2010:ASE

[CYC10] JinMin Choi, Yung Jun Yoo, and Jin Young Choi. Adaptive shadow estimator for removing shadow of moving object. Computer Vision and Image Understanding: CVIU, 114(9):1017– 1029, September 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2022:RDD

[CYD^{+}22] Jiyou Chen, Gaobo Yang, Xiangling Ding, Zhiqing Guo, and Shuai Wang. Robust detection of dehazed images via dualstream CNNs with adaptive feature fusion. Computer Vision and Image Understanding: CVIU, 217 (??):??, March 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000017.

Coughlan:2000:EDT

[CYES00] James Coughlan, Alan Yuille, Camper English, and Dan Snow. Efficient deformable template detection and localization without user initialization. Computer Vision and Image Understanding: CVIU, 78 (3):303–319, June 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0842;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0842/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0842/ref.

Cuevas:2016:LDI

[CYG16] Carlos Cuevas, Eva María Yáñez, and Narciso García. Labeled dataset for integral evaluation of moving object detection algorithms: LASIESTA. Computer Vision and Image Understanding: CVIU, 152(??):103– 117, November 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301138.

Chen:2011:HPR

[CYNO11] Cheng Chen, Yi Yang, Feiping Nie, and Jean-Marc Odobez. 3D human pose recovery from image by efficient visual feature selection. Computer Vision and Image Understanding: CVIU, 115(3):290– 299, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2010:CHP

[CYP^{+}10] Chung-Hao Chen, Yi Yao, David Page, Besma Abidi, Andreas Koschan, and Mongi Abidi. Camera handoff and placement for automated tracking systems with multiple omnidirectional cameras. Computer Vision and Image Understanding: CVIU, 114(2): 179–197, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2023:ARF

[CYY^{+}23] Jiawei Chen, Xiao Yang, Heng Yin, Mingzhi Ma, Bihui Chen, Jianteng Peng, Yandong Guo, Zhaoxia Yin, and Hang Su. AdvFAS: a robust face anti-spoofing framework against adversarial examples. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001595.

Chen:2014:FAS

[CZ14] Cheng Chen and Guoyan Zheng. Fully automatic segmentation of AP pelvis Xrays via random forest regression with efficient feature selection and hierarchical sparse shape composition. Computer Vision and Image Understanding: CVIU, 126(??):1–10, September 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001027.

Chung:2018:LLR

[CZ18] Joon Son Chung and Andrew Zisserman. Learning to lip read words by watching videos. Computer Vision and Image Understanding: CVIU, 173(??): 76–85, August 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218300134.

Cheng:2015:SSL

[CZHT15] Yanhua Cheng, Xin Zhao, Kaiqi Huang, and Tieniu Tan. Semi-supervised learning and feature evaluation for RGB-D object recognition. Computer Vision and Image Understanding: CVIU, 139(??):149– 160, October 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001083.

Cai:2024:DOT

[CZL^{+}24] Huayue Cai, Xiang Zhang, Long Lan, Changcheng Xiao, Chuanfu Xu, Jie Liu, and Zhigang Luo. Discriminative object tracking by domain contrast. Computer Vision and Image Understanding: CVIU, 239(??): ??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002710.

Chen:2020:NTA

[CZS^{+}20] Fangyi Chen, Chenchen Zhu, Zhiqiang Shen, Han Zhang, and Marios Savvides. NCMS: Towards accurate anchor free object detection through l_{2} norm calibration and multi-feature selection. Computer Vision and Image Understanding: CVIU, 200(??):Article 103050, November 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S1077314220300953.

Csurka:1997:CUF

[CZZF97] Gabriella Csurka, Cyril Zeller, Zhengyou Zhang, and Olivier D. Faugeras. Characterizing the uncertainty of the fundamental matrix. Computer Vision and Image Understanding: CVIU, 68(1):18– 36, October 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0531/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0531/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0531/production/ref.

Cui:2007:LBD

[CZZS07] Jinshi Cui, Hongbin Zha, Huijing Zhao, and Ryosuke Shibasaki. Laser-based detection and tracking of multiple people in crowds. Computer Vision and Image Understanding: CVIU, 106(2– 3):300–312, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dubourvieux:2022:FAG

[DAL^{+}22] Fabian Dubourvieux, Romaric Audigier, Angélique Loesch, Samia Ainouz, and Stéphane Canu. A formal approach to good practices in Pseudo-Labeling for Unsupervised Domain Adaptive Re-Identification. Computer Vision and Image Understanding: CVIU, 223 (??):??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001102.

Damiand:2008:TMI

[Dam08] Guillaume Damiand. Topological model for 3D image representation: Definition and incremental extraction algorithm. Computer Vision and Image Understanding: CVIU, 109(3):260– 289, March 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Diop:2012:INM

[DAM12] El Hadji S. Diop, Radjesvarane Alexandre, and Lionel Moisan. Intrinsic nonlinear multiscale image decomposition: a 2D empirical mode decompositionlike tool. Computer Vision and Image Understanding: CVIU, 116(1):102– 119, January 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001986.

Daniilidis:1997:FSM

[Dan97]  Konstantinos  Daniilidis.

Fixation simplifies 3D motion estimation. Computer Vision and Image Understanding: CVIU, 68(2):158– 169, November 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0535/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0535/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0535/production/ref.

Davatzikos:1997:STR

[Dav97] Christos Davatzikos. Spatial transformation and registration of brain images using elastically deformable models. Computer Vision and Image Understanding: CVIU, 66(2): 207–222, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0605/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0605/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0605/production/ref.

Das:2017:HAV

[DAZ^{+}17] Abhishek Das, Harsh Agrawal, Larry Zitnick, Devi Parikh, and Dhruv Batra. Human attention in visual question answering: Do humans and deep networks look at the same regions? Computer Vision and Image Understanding: CVIU, 163(??): 90–100, October 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301649.

DelaTorre:2003:RPC

[DB03] Fernando De la Torre and Michael J. Black. Robust parameterized component analysis: theory and applications to 2D facial appearance models. Computer Vision and Image Understanding: CVIU, 91(1–2):53–71, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ducournau:2014:RWD

[DB14] Aurélien Ducournau and Alain Bretto. Random walks in directed hypergraphs and application to semi-supervised image segmentation. Computer Vision and Image Understanding: CVIU, 120(??):91– 102, March 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213002038.

Dutt:2013:APO

[DBB13] Mousumi Dutt, Arindam Biswas, and Partha Bhowmick. Approximate partitioning of 2D objects into orthogonally convex components. Computer Vision and Image Understanding: CVIU, 117(4): 326–341, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001853.

Draper:2003:RFP

[DBBB03] Bruce A. Draper, Kyungim Baek, Marian Stewart Bartlett, and J. Ross Beveridge. Recognizing faces with PCA and ICA. Computer Vision and Image Understanding: CVIU, 91 (1–2):115–137, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dutt:2014:FSI

[DBBB14] Mousumi Dutt, Arindam Biswas, Partha Bhowmick, and Bhargab B. Bhattacharya. On the family of shortest isothetic paths in a digital object

— an algorithm with applications. Computer Vision and Image Understanding: CVIU, 129(??):75–88, December 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001520.

Damiand:2004:TMT

[DBF04] Guillaume Damiand, Yves Bertrand, and Christophe Fiorio. Topological model for two-dimensional image representation: definition and optimal extraction algorithm. Computer Vision and Image Understanding: CVIU, 93(2):111–154, February 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dimiccoli:2017:SCS

[DBT^{+}17] Mariella Dimiccoli, Marc Bolaños, Estefania Talavera, Maedeh Aghaei, Stavri G. Nikolov, and Petia Radeva. SR-clustering: Semantic regularized clustering for egocentric photo streams segmentation. Computer Vision and Image Understanding: CVIU, 155(??): 55–69, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301576.

Dankers:2007:MZS

[DBZ07] Andrew Dankers, Nick Barnes, and Alex Zelinsky.

MAP ZDF segmentation and tracking using active stereo vision: Hand tracking case study. Computer Vision and Image Understanding: CVIU, 108(1–2):74–86, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Deshpande:1998:REI

[DC98] S. G. Deshpande and S. Chaudhuri. Recursive estimation of illuminant motion from flow field and simultaneous recovery of shape. Computer Vision and Image Understanding: CVIU, 72(1):10– 20, October 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0657/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0657/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0657/production/ref.

Dornaika:2000:CSM

[DC00a] F. Dornaika and R. Chung. Cooperative stereo-motion: Matching and reconstruction. Computer Vision and Image Understanding: CVIU, 79(3):408–427, September 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0867;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0867/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0867/ref.

Drummond:2000:LTS

[DC00b] Tom Drummond and Terry Caelli. Learning taskspecific object recognition and scene understanding. Computer Vision and Image Understanding: CVIU, 80(3): 315–348, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0882;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0882/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0882/ref.

Dornaika:2001:AAC

[DC01] F. Dornaika and R. Chung.

An algebraic approach to camera self-calibration. Computer Vision and Image Understanding: CVIU, 83(3):195–215, September 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2001.0925;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0925/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0925/ref.

DelBimbo:1999:GEI

[DCCL99] Alberto Del Bimbo, Vittorio Castelli, Shih-Fu Chang, and Chung-Sheng Li. GUEST EDITORS’ IN-TRODUCTION: Contentbased access of image and video libraries. Computer Vision and Image Understanding: CVIU, 75(1–2):1– 2, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0775/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0775/production/pdf.

deCampos:2012:ISL

[dCCP12] Teófilo de Campos, Gabriela Csurka, and Florent Perronnin. Images as sets of locally weighted features. Computer Vision and Image Understanding: CVIU, 116(1):68–85, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211001822.

Denis:2007:SSQ

[DCFM07] Patrice Denis, Philippe Carre, and Christine Fernandez-Maloigne. Spatial and spectral quaternionic approaches for colour images. Computer Vision and Image Understanding: CVIU, 107 (1–2):74–87, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dee:2012:BSS

[DCH12] Hannah M. Dee, Anthony G. Cohn, and David C. Hogg. Building semantic scene models from unconstrained video. Computer Vision and Image Understanding: CVIU, 116(3): 446–456, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002025.

Drapaca:2005:STB

[DCS05] Corina S. Drapaca, Valerie Cardenas, and Colin Studholme. Segmentation of tissue boundary evolution from brain MR image sequences using multi-phase level sets. Computer Vision and Image Understanding: CVIU, 100(3):312–329, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dickinson:1997:AOR

[DCTO97] Sven J. Dickinson, Henrik I. Christensen, John K. Tsotsos, and Göran Olofsson. Active object recognition integrating attention and viewpoint control. Computer Vision and Image Understanding: CVIU, 67(3):239–260, September 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0532/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0532/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0532/production/ref.

DelBimbo:2011:PFB

[DD11a] Alberto Del Bimbo and Fabrizio Dini. Particle filterbased visual tracking with a first order dynamic model and uncertainty adaptation. Computer Vision and Image Understanding: CVIU, 115(6):771–786, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dong:2011:ALM

[DD11b] Y. Dong and G. N. DeSouza. Adaptive learning of multisubspace for foreground detection under illumination changes. Computer Vision and Image Understanding: CVIU, 115(1):31–49, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

DelBimbo:2010:EDV

[DDLP10] A. Del Bimbo, F. Dini, G. Lisanti, and F. Pernici. Exploiting distinctive visual landmark maps in pan– tilt–zoom camera networks. Computer Vision and Image Understanding: CVIU, 114(6):611–623, June 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Deng:2012:CTG

[DDWZ12] Yue Deng, Qionghai Dai, Ruiping Wang, and Zengke Zhang. Commute time guided transformation for feature extraction. Computer Vision and Image Understanding: CVIU, 116(4): 473–483, April 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002578.

Du:2023:DTS

[DDZ^{+}23] Yong Du, Junjie Deng, Yulong Zheng, Junyu Dong, and Shengfeng He. DSD-Net: Toward single image deraining with self-paced curricular dual stimulations. Computer Vision and Image Understanding: CVIU, 230(??):??, April 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000371.

Demi:1996:CTE

[Dem96] Marcello Demi. Contour tracking by enhancing corners and junctions. Computer Vision and Image Understanding: CVIU, 63(1): 118–134, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0008/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0008/production/pdf.

Demi:2005:GLC

[Dem05] M. Demi. On the gray-level central and absolute central moments and the mass center of the gray-level variability in low-level image processing. Computer Vision and Image Understanding: CVIU, 97(2):180–208, February 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dahi:2017:EBM

[DETE17] Ilias Dahi, Miloud Chikr El Mezouar, Nasreddine Taleb, and Mohamed Elbahri. An edge-based method for effective abandoned luggage detection in complex surveillance videos. Computer Vision and Image Understanding: CVIU, 158(??): 141–151, May 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300243.

Delamarre:2001:AMM

[DF01] Quentin Delamarre and Olivier Faugeras. 3D articulated models and multiview tracking with physical forces. Computer Vision and Image Understanding: CVIU, 81(3):328– 357, March 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0892;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0892/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0892/ref.

Dalley:2002:PWR

[DF02] Gerald Dalley and Patrick Flynn. Pair-wise range image registration: a study in outlier classification. Computer Vision and Image Understanding: CVIU, 87(1– 3):104–115, July 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dong:2022:LCD

[DFH^{+}22] Hanze Dong, Yanwei Fu, Sung Ju Hwang, Leonid Sigal, and Xiangyang Xue. Learning the compositional domains for generalized zero-shot learning. Computer Vision and Image Understanding: CVIU, 221 (??):??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000649.

Darrell:2015:ICS

[DFJL15] Trevor Darrell, Vittorio Ferrari, Frederic Jurie, and Vincent Lepetit. Introduction to the CVIU special issue on “Parts and Attributes: Mid-level representation for object recognition, scene classification and object detection”. Computer Vision and Image Understanding: CVIU, 138(??): 85, September 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001411.

Dong:2023:WSF

[DFLS23] Zihao Dong, Tiyu Fang, Jinping Li, and Xiuli Shao. Weakly supervised finegrained semantic segmentation via spatial correlationguided learning. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001959.

Lascio:2013:RTA

[DFP^{+}13] Rosario Di Lascio, Pasquale Foggia, Gennaro Percannella, Alessia Saggese, and Mario Vento. A real time algorithm for people tracking using contextual reasoning. Computer Vision and Image Understanding: CVIU, 117(8):892– 908, August 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000908.

Deng:2023:AMS

[DFP23] Bowen Deng, Andrew P.

French, and Michael P. Pound. Addressing multiple salient object detection via dual-space long-range dependencies. Computer Vision and Image Understanding: CVIU, 235(??):

??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422300156X.

Durou:2008:NMS

[DFS08] Jean-Denis Durou, Maurizio Falcone, and Manuela Sagona. Numerical methods for shape-from-shading: a new survey with benchmarks. Computer Vision and Image Understanding: CVIU, 109(1):22–43, January 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Damodaran:2020:EOT

[DFSC20] Bharath Bhushan Damodaran, Rémi Flamary, Vivien Seguy, and Nicolas Courty. An entropic optimal transport loss for learning deep neural networks under label noise in remote sensing images. Computer Vision and Image Understanding: CVIU, 191(??):Article 102863, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219301559.

Derrode:2001:REF

[DG01] Stéphane Derrode and Faouzi Ghorbel. Robust and efficient Fourier-Mellin transform approximations for gray-level image reconstruction and complete invariant description. Computer Vision and Image Understanding: CVIU, 83(1): 57–78, July 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0922;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0922/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0922/ref.

Dorini:2011:UFT

[DG11] Leyza Baldo Dorini and Siome Klein Goldenstein. Unscented feature tracking. Computer Vision and Image Understanding: CVIU, 115(1):8–15, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Daubney:2012:EPA

[DGC12] Ben Daubney, David Gibson, and Neill Campbell. Estimating pose of articulated objects using lowlevel motion. Computer Vision and Image Understanding: CVIU, 116(3): 330–346, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002049.

Delakis:2008:AIS

[DGG08] Manolis Delakis, Guillaume Gravier, and Patrick Gros. Audiovisual integration with Segment Models for tennis video parsing. Computer Vision and Image Understanding: CVIU, 111 (2):142–154, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Daul:1998:HTN

[DGH98] Christian Daul, Pierre Graebling, and Ernest Hirsch. From the Hough transform to a new approach for the detection and approximation of elliptical arcs. Computer Vision and Image Understanding: CVIU, 72(3):215–236, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0696/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0696/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0696/production/ref.

Dave:2022:TTC

[DGRS22] Ishan Dave, Rohit Gupta, Mamshad Nayeem Rizve, and Mubarak Shah. TCLR: Temporal contrastive learning for video representation.

Computer Vision and Image Understanding: CVIU, 219(??):??, June 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000376.

Demirdjian:2000:MED

[DH00] David Demirdjian and Radu Horaud. Motion-egomotion discrimination and motion segmentation from imagepair streams. Computer Vision and Image Understanding: CVIU, 78(1): 53–68, April 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0827;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0827/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0827/ref.

Du:2019:WPB

[DH19] Lingshuang Du and Haifeng Hu. Weighted patchbased manifold regularization dictionary pair learning model for facial expression recognition using iterative optimization classification strategy. Computer Vision and Image Understanding: CVIU, 186(??):13–24, September 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219300943.

Drouin:2008:IDO

[DHP08] Stéphane Drouin, Patrick Hébert, and Marc Parizeau. Incremental discovery of object parts in video sequences. Computer Vision and Image Understanding: CVIU, 110(1):60–74, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Drosou:2012:SAH

[DIMT12] Anastasios Drosou, Dimosthenis Ioannidis, Konstantinos Moustakas, and Dimitrios Tzovaras. Spatiotemporal analysis of human activities for biometric authentication. Computer Vision and Image Understanding: CVIU, 116(3): 411–421, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211002098.

Drew:2014:ZII

[DJF14] Mark S. Drew, Hamid Reza Vaezi Joze, and Graham D. Finlayson. The Zeta-image, illuminant estimation, and specularity manipulation. Computer Vision and Image Understanding: CVIU, 127(??):

1–13, October 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001532.

Dubuisson-Jolly:2001:TDT

[DJG01] Marie-Pierre Dubuisson-Jolly and Alok Gupta. Tracking deformable templates using a shortest path algorithm. Computer Vision and Image Understanding: CVIU, 81 (1):26–45, January 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0883;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0883/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0883/ref.

Dadashi:2013:AFN

[DK13] Roghayeh Dadashi and Hamidreza Rashidy Kanan. AVCD-FRA: a novel solution to automatic video cut detection using fuzzy-rule-based approach. Computer Vision and Image Understanding: CVIU, 117(7): 807–817, July 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000544.

Das:2022:IID

[DKG22] Partha Das, Sezer Karaoglu, and Theo Gevers. Intrinsic image decomposition using physics-based cues and CNNs. Computer Vision and Image Understanding: CVIU, 223(??): ??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001163.

Das:1997:RID

[DL97] Atish K. Das and Noshir A. Langrana. Recognition and integration of dimension sets in vectorized engineering drawings. Computer Vision and Image Understanding: CVIU, 68(1):90– 108, October 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0537/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0537/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0537/production/ref.

Draper:2005:ESA

[DL05] Bruce A. Draper and Albert Lionelle. Evaluation of selective attention under similarity transformations.

Computer Vision and Image Understanding: CVIU, 100(1–2):152–171, October/ November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Drbohlav:2010:TCI

[DL10] Ondřej Drbohlav and Aleš Leonardis. Towards correct and informative evaluation methodology for texture classification under varying viewpoint and illumination. Computer Vision and Image Understanding: CVIU, 114(4):439–449, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

deLavarene:2007:PIL

[dLAH07] Brice Chaix de Lavarène, David Alleysson, and Jeanny Hérault. Practical implementation of LMMSE demosaicing using luminance and chrominance spaces. Computer Vision and Image Understanding: CVIU, 107(1–2):3–13, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Daudt:2019:MLL

[DLBG19] Rodrigo Caye Daudt, Bertrand Le Saux, Alexandre Boulch, and Yann Gousseau. Multitask learning for largescale semantic change detection. Computer Vision and Image Understanding: CVIU, 187(??):Article 102783, October 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219300992.

Darby:2014:TOP

[DLC14] John Darby, Baihua Li, and Nicholas Costen. Tracking object poses in the context of robust body pose estimates. Computer Vision and Image Understanding: CVIU, 127(??):57–72, October 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001362.

Dai:2024:WBN

[DLC^{+}24] Tianhong Dai, Wei Li, Xilei Cao, Jianzhuang Liu, Xu Jia, Ales Leonardis, Youliang Yan, and Shanxin Yuan. Wavelet-based network for high dynamic range imaging. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223002618.

Dimitrijevic:2006:HBP

[DLF06] M. Dimitrijevic, V. Lepetit, and P. Fua. Human body pose detection using Bayesian spatio-temporal templates. Computer Vision and Image Understanding: CVIU, 104(2–3):127–139, November/December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Deng:1999:WBL

[DLHT99] Peter Shaohua Deng, Hong-Yuan Mark Liao, Chin Wen Ho, and Hsiao-Rong Tyan. Wavelet-based off-line handwritten signature verification. Computer Vision and Image Understanding: CVIU, 76(3):173–190, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0799/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0799/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0799/production/ref.

Damen:2016:YDL

[DLMC16] Dima Damen, Teesid Leelasawassuk, and Walterio Mayol-Cuevas. You-Do, I-Learn: Egocentric unsupervised discovery of objects and their modes of interaction towards videobased guidance. Computer Vision and Image Understanding: CVIU, 149(??): 98–112, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216000709.

DOrazio:2009:VSR

[DLS^{+}09] T. D’Orazio, M. Leo, P. Spagnolo, M. Nitti, N. Mosca, and A. Distante. A visual system for real time detection of goal events during soccer matches. Computer Vision and Image Understanding: CVIU, 113(5): 622–632, May 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Desai:2015:EFD

[DLV15] Alok Desai, Dah-Jye Lee, and Dan Ventura. An efficient feature descriptor based on synthetic basis functions and uniqueness matching strategy. Computer Vision and Image Understanding: CVIU, 142 (??):37–49, ???? 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002003.

Delingette:2001:STC

[DM01] H. Delingette and J. Montagnat. Shape and topology constraints on parametric active contours. Computer Vision and Image Understanding: CVIU, 83 (2):140–171, August 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0920;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0920/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0920/ref.

Datta:2012:TDS

[DM12] Madhura Datta and C. A.

Murthy. Two dimensional synthetic face generation and verification using set estimation technique. Computer Vision and Image Understanding: CVIU, 116(9):1022–1031, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212000793.

Dunnhofer:2024:VTC

[DM24] Matteo Dunnhofer and Christian Micheloni. Visual tracking in cameraswitching outdoor sport videos: Benchmark and baselines for skiing. Computer Vision and Image Understanding: CVIU, 243 (??):??, June 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000596.

Dafni:2017:DMR

[DMAD17] Adi Dafni, Yael Moses, Shai Avidan, and Tali Dekel. Detecting moving regions in CrowdCam images. Computer Vision and Image Understanding: CVIU, 160 (??):36–44, July 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300656.

deMiranda:2010:SAW

[dMFU10] P. A. V. de Miranda, A. X. Falcão, and J. K. Udupa. Synergistic arc-weight estimation for interactive image segmentation using graphs. Computer Vision and Image Understanding: CVIU, 114(1):85–99, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Desmarais:2021:RHP

[DMSM21] Yann Desmarais, Denis Mottet, Pierre Slangen, and Philippe Montesinos. A review of 3D human pose estimation algorithms for markerless motion capture. Computer Vision and Image Understanding: CVIU, 212(??):??, November 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221001193.

Dunne:2010:EGC

[DMW10] Aubrey K. Dunne, John Mallon, and Paul F. Whelan. Efficient generic calibration method for general cameras with single centre of projection. Computer Vision and Image Understanding: CVIU, 114(2):220–233, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Doan:2024:ADG

[DNG^{+}24] Anh-Dzung Doan, Bach Long Nguyen, Surabhi Gupta, Ian Reid, Markus Wagner, and Tat-Jun Chin. Assessing domain gap for continual domain adaptation in object detection. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002655.

Doermann:1998:IRD

[Doe98]  David Doermann.  The

indexing and retrieval of document images: a survey. Computer Vision and Image Understanding: CVIU, 70(3):287– 298, June 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0687/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0687/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0687/production/ref;
http://www.idealibrary.
com/links/artid/cviu.1998.
0692/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0692/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0692/production/ref.

Demirci:2011:EMM

[DOSD11] M. Fatih Demirci, Yusuf Osmanlioglu, Ali Shokoufandeh, and Sven Dickinson. Efficient many-to-many feature matching under the l_{1} norm. Computer Vision and Image Understanding: CVIU, 115(7): 976–983, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000804.

Junior:2012:RCH

[dOSJVBS12] Jurandir de Oliveira Santos Junior, Alexandre Vrubel, Olga R. P. Bellon, and Luciano Silva. 3D reconstruction of cultural heritages: Challenges and advances on precise mesh integration. Computer Vision and Image Understanding: CVIU, 116(12):1195– 1207, December 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001166.

deLaGorce:2010:VAM

[dP10] Martin de La Gorce and Nikos Paragios. A variational approach to monocular hand-pose estimation. Computer Vision and Image Understanding: CVIU, 114 (3):363–372, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Demi:2000:FAC

[DPB00] M. Demi, M. Paterni, and A. Benassi. The first absolute central moment in low-level image processing. Computer Vision and Image Understanding: CVIU, 80(1):57–87, October 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0861;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.

                                        2000.0861/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0861/ref.

Demirkus:2015:HTG

[DPCA15] Meltem Demirkus, Doina Precup, James J. Clark, and Tal Arbel. Hierarchical temporal graphical model for head pose estimation and subsequent attribute classification in realworld videos. Computer Vision and Image Understanding: CVIU, 136(??): 128–145, July 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421500051X.

Delibasis:2014:RHS

[DPM14] K. K. Delibasis, V. P. Plagianakos, and I. Maglogiannis. Refinement of human silhouette segmentation in omni-directional indoor videos. Computer Vision and Image Understanding: CVIU, 128(??): 65–83, November 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214001386.

Das:2017:CAM

[DPRC17] Abir Das, Rameswar Panda, and Amit K. Roy-Chowdhury. Continuous adaptation of multi-camera person identification models through sparse non-redundant representative selection. Computer Vision and Image Understanding: CVIU, 156 (??):66–78, March 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314216301643.

Dowdall:2007:CT

[DPT07] Jonathan Dowdall, Ioannis T. Pavlidis, and Panagiotis Tsiamyrtzis. Coalitional tracking. Computer Vision and Image Understanding: CVIU, 106(2–3): 205–219, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Das:2004:FSS

[DR04] M. Das and E. M. Riseman. FOCUS: a system for searching for multi-colored objects in a diverse image database. Computer Vision and Image Understanding: CVIU, 94(1–3): 168–192, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Diaz:2008:SHP

[DRAB08] Javier Díaz, Eduardo Ros, Rodrigo Agís, and Jose Luis Bernier. Superpipelined high-performance opticalflow computation architecture. Computer Vision and Image Understanding: CVIU, 112(3):262–273, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dehne:1995:HAP

[DRCF95] Frank Dehne, Andrew Rau-Chaplin, and Afonso G. Ferreira. Hypercube algorithms for parallel processing of pointer-based quadtrees. Computer Vision and Image Understanding: CVIU, 62(1):1–10, July 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1037/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1037/production/pdf.

Debled-Rennesson:2013:SID

[DRDKE13] Isabelle Debled-Rennesson, Eric Domenjoud, Bertrand Kerautret, and Philippe Even. Special issue on discrete geometry for computer imagery. Computer Vision and Image Understanding: CVIU, 117(4): 305, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000167.

Drew:1996:DSO

[Dre96] Mark S. Drew. Direct solution of orientation-from-color problem using a modification of Pentland’s light source direction estimator. Computer Vision and Image Understanding: CVIU, 64(2):286– 299, September 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0059/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0059/production/pdf.

Denman:2003:CBA

[DRK03] H. Denman, N. Rea, and A. Kokaram. Contentbased analysis for video from snooker broadcasts. Computer Vision and Image Understanding: CVIU, 92 (2–3):176–195, November/ December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dalsasso:2024:MSS

[DRTT24] Emanuele Dalsasso, Clément Rambour, Nicolas Trouvé, and Nicolas Thome. MERLIN-Seg: Self-supervised despeckling for label-efficient semantic segmentation. Computer Vision and Image Understanding: CVIU, 241 (??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000213.

Davis:2007:BSU

[DS07] James W. Davis and Vinay Sharma. Backgroundsubtraction using contourbased fusion of thermal and visible imagery. Computer Vision and Image Understanding: CVIU, 106(2–3): 162–182, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Damiand:2011:PAS

[DSdlH^{+}11] Guillaume Damiand, Christine Solnon, Colin de la Higuera, Jean-Christophe Janodet, and Émilie Samuel. Polynomial algorithms for subisomorphism of nD open combinatorial maps. Computer Vision and Image Understanding: CVIU, 115(7): 996–1010, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211000816.

daSilva:2012:IMD

[dSdSF^{+}12] André Tavares da Silva, Jefersson Alex dos Santos, Alexandre Xavier Falcão, Ricardo da S. Torres, and Léo Pini Magalhães. Incorporating multiple distance spaces in optimumpath forest classification to improve feedback-based learning. Computer Vision and Image Understanding: CVIU, 116(4):510– 523, April 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421100261X.

Evangelista:2024:SMT

[dSEdSPdMF24] Raphael dos S. Evangelista, Andre Luiz da S. Pereira, Rogério Ferreira de Moraes, and Leandro A. F. Fernandes. Semantic manipulation through the lens of Geometric Algebra. Computer Vision and Image Understanding: CVIU, 239(??): ??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002795.

Dufournaud:2004:IMS

[DSH04] Yves Dufournaud, Cordelia Schmid, and Radu Horaud. Image matching with scale adjustment. Computer Vision and Image Understanding: CVIU, 93(2):175–194, February 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Duong:2020:EMO

[DSK^{+}20] Nam-Duong Duong, Catherine Soladié, Amine Kacete, Pierre-Yves Richard, and Jérôme Royan. Efficient multi-output scene coordinate prediction for fast and accurate camera relocalization from a single RGB image. Computer Vision and Image Understanding: CVIU, 190(??):Article 102850, January 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219301468.

deSouza:2014:HHN

[dSM14] Gustavo Botelho de Souza and Aparecido Nilceu Marana. HTS and HTSn: New shape descriptors based on Hough transform statistics. Computer Vision and Image Understanding: CVIU, 127(??):43–56, October 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001374.

Denton:2008:CSI

[DSNN08] Trip Denton, Ali Shokoufandeh, John Novatnack, and Ko Nishino. Canonical subsets of image features. Computer Vision and Image Understanding: CVIU, 112(1): 55–66, October 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dutagaci:2010:SMR

[DSY10] Helin Dutağacı, Bülent Sankur, and Yücel Yemez. Subspace methods for retrieval of general 3D models. Computer Vision and Image Understanding: CVIU, 114 (8):865–886, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

DaVitoriaLobo:1996:CED

[DT96a] Niels Da Vitoria Lobo and John K. Tsotsos. Computing egomotion and detecting independent motion from image motion using collinear points. Computer Vision and Image Understanding: CVIU, 64(1): 21–52, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0044/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0044/production/pdf.

Dalmia:1996:HSE

[DT96b] Arun K. Dalmia and Mohan Trivedi. High-speed extraction of 3D structure of selectable quality using a translating camera. Computer Vision and Image Understanding: CVIU, 64(1): 97–110, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0047/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0047/production/pdf.

Dudek:1997:SRR

[DT97] Gregory Dudek and John K. Tsotsos. Shape representation and recognition from multiscale curvature. Computer Vision and Image Understanding: CVIU, 68(2):170– 189, November 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0533/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0533/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0533/production/ref.

Daliri:2009:CSU

[DT09] Mohammad Reza Daliri and Vincent Torre. Classification of silhouettes using contour fragments. Computer Vision and Image Understanding: CVIU, 113(9): 1017–1025, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Daliri:2010:SRB

[DT10] Mohammad Reza Daliri and Vincent Torre. Shape recognition based on Kernel-edit distance. Computer Vision and Image Understanding: CVIU, 114(10):1097–1103, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

DenHartog:1996:KBI

[DTG96] J. E. Den Hartog, T. K. Ten Kate, and J. J. Gerbrands. Knowledge-based interpretation of utility maps. Computer Vision and Image Understanding: CVIU, 63(1): 105–117, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0007/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0007/production/pdf.

Deng:2017:UOR

[DTL17] Zhuo Deng, Sinisa Todorovic, and Longin Jan Latecki. Unsupervised object region proposals for RGB-D indoor scenes. Computer Vision and Image Understanding: CVIU, 154(??): 127–136, January 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301011.

Delibasis:1997:DFD

[DUC97] K. Delibasis, P. E. Undrill, and G. G. Cameron. Designing Fourier descriptorbased geometric models for object interpretation in medical images using genetic algorithms. Computer Vision and Image Understanding: CVIU, 66(3): 286–300, June 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0505/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0505/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0505/production/ref.

Dori:1998:SRD

[DV98] Dov Dori and Yelena Velkovitch. Segmentation and recognition of dimensioning text from engineering drawings. Computer Vision and Image Understanding: CVIU, 69(2):196– 201, February 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0585/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0585/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0585/production/ref.

Demirci:2008:ITL

[DvLV08] M. Fatih Demirci, Reinier H.

van Leuken, and Remco C. Veltkamp. Indexing through Laplacian spectra. Computer Vision and Image Understanding: CVIU, 110(3): 312–325, June 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dang:2011:SIP

[DWB11] Trung Kien Dang, Marcel Worring, and The Duy Bui. A semi-interactive panorama based 3D reconstruction framework for indoor scenes. Computer Vision and Image Understanding: CVIU, 115(11):1516– 1524, November 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001664.

Dansereau:2016:SCD

[DWC16] Donald G. Dansereau, Stefan B. Williams, and Peter I. Corke. Simple change detection from mobile light field cameras. Computer Vision and Image Understanding: CVIU, 145(??): 160–171, April 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002726.

Dong:2024:TLC

[DWC^{+}24] Pengwei Dong, Bo Wang, Runmin Cong, Hai-Han Sun, and Chongyi Li. Transformer with large convolution kernel decoder network for salient object detection in optical remote sensing images. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002977.

Ding:2019:CDM

[DWL19] Mingyu Ding, Zhe Wang, and Zhiwu Lu. Crossdomain mapping learning for transductive zero-shot learning. Computer Vision and Image Understanding: CVIU, 187(??):Article 102784, October 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1077314219301006.

Dong:2024:TIS

[DWL^{+}24] Pei Dong, Lei Wu, Ruichen Li, Xiangxu Meng, and Lei Meng. Text to image synthesis with multigranularity feature aware enhancement Generative Adversarial Networks. Computer Vision and Image Understanding: CVIU, 245 (??):??, August 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224001231.

Duan:2023:CDM

[DWLW23] Yiqun Duan, Zhen Wang, Yi Li, and Jingya Wang. Cross-domain multi-style merge for image captioning. Computer Vision and Image Understanding: CVIU, 228 (??):??, February 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001953.

DeSmedt:2019:HHG

[DWV19] Quentin De Smedt, Hazem Wannous, and Jean-Philippe Vandeborre. Heterogeneous hand gesture recognition using 3D dynamic skeletal data. Computer Vision and Image Understanding: CVIU, 181(??): 60–72, April 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219300153.

Deng:2012:SCH

[DWW^{+}12] Xiaoming Deng, Fuchao Wu, Yihong Wu, Fuqing Duan, Liang Chang, and Hongan Wang. Selfcalibration of hybrid central catadioptric and perspective cameras. Computer Vision and Image Understanding: CVIU, 116(6): 715–729, June 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212000355.

Ding:1998:CSI

[DY98] Yuan Ding and Tzay Y. Young. Complete shape from imperfect contour: a rule-based approach. Computer Vision and Image Understanding: CVIU, 70(2): 197–211, May 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0635/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0635/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0635/production/ref.

Dinh:2014:CTG

[DYM14] Thang Ba Dinh, Qian Yu, and Gérard Medioni. Cotrained generative and discriminative trackers with cascade particle filter. Computer Vision and Image Understanding: CVIU, 119 (??):41–56, February 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314213002075.

Dong:2014:EKB

[DZJB14] Zilong Dong, Guofeng Zhang, Jiaya Jia, and Hujun Bao. Efficient keyframebased real-time camera tracking. Computer Vision and Image Understanding: CVIU, 118(??):97–110, January 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001562.

Dai:2007:PDT

[DZL07] Congxia Dai, Yunfei Zheng, and Xin Li. Pedestrian detection and tracking in infrared imagery using shape and appearance. Computer Vision and Image Understanding: CVIU, 106(2–3): 288–299, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Drory:2017:ADT

[DZLH17] Ami Drory, Gao Zhu, Hongdong Li, and Richard Hartley. Automated detection and tracking of slalom paddlers from broadcast image sequences using cascade classifiers and discriminative correlation filters. Computer Vision and Image Understanding: CVIU, 159(??):116– 127, June 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216301990.

Dang:2024:PUP

[DZQ24] Jiachen Dang, Yong Zhong, and Xiaolin Qin. PPformer: Using pixel-wise and patchwise cross-attention for lowlight image enhancement. Computer Vision and Image Understanding: CVIU, 241(??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000110.

Du:2023:PST

[DZZ^{+}23] Xiangcheng Du, Zhao Zhou, Yingbin Zheng, Xingjiao Wu, Tianlong Ma, and Cheng Jin. Progressive scene text erasing with self-supervision. Computer Vision and Image Understanding: CVIU, 233(??):

??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000929.

Eggers:1995:HSI

[EA95] Daniel D. Eggers and Eugene Ackerman. High speed image rotation in embedded systems. Computer Vision and Image Understanding: CVIU, 61(2):270– 277, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1019/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1019/production/pdf.

Elguebaly:2013:FAG

[EB13] Tarek Elguebaly and Nizar Bouguila. Finite asymmetric generalized Gaussian mixture models learning for infrared object detection. Computer Vision and Image Understanding: CVIU, 117(12):1659–1671, December 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001379.

Eskil:2014:FER

[EB14] M. Taner Eskil and Kristin S.

Benli. Facial expression recognition based on anatomy. Computer Vision and Image Understanding: CVIU, 119(??):1–14, February 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213002063.

Erol:2007:VBH

[EBN^{+}07] Ali Erol, George Bebis, Mircea Nicolescu, Richard D. Boyle, and Xander Twombly. Vision-based hand pose estimation: a review. Computer Vision and Image Understanding: CVIU, 108(1–2): 52–73, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Emberton:2018:UIV

[ECC18] Simon Emberton, Lars Chittka, and Andrea Cavallaro. Underwater image and video dehazing with pure haze region segmentation. Computer Vision and Image Understanding: CVIU, 168(??):145– 156, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301418.

Erichson:2016:RLR

[ED16] N. Benjamin Erichson and Carl Donovan. Randomized low-rank Dynamic Mode Decomposition for motion detection. Computer Vision and Image Understanding: CVIU, 146(??):40– 50, May 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216000503.

Erdem:2012:FBT

[EDB12] Erkut Erdem, Séverine Dubuisson, and Isabelle Bloch. Fragments based tracking with adaptive cue integration. Computer Vision and Image Understanding: CVIU, 116(7):827– 841, July 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212000525.

Escorcia:2020:GWA

[EDJ^{+}20] Victor Escorcia, Cuong D.

Dao, Mihir Jain, Bernard Ghanem, and Cees Snoek. Guess where? Actorsupervision for spatiotemporal action localization. Computer Vision and Image Understanding: CVIU, 192 (??):Article 102886, March 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S1077314219301687.

Edwards:2016:PAS

[EDX16] Michael Edwards, Jingjing Deng, and Xianghua Xie. From pose to activity: Surveying datasets and introducing CONVERSE. Computer Vision and Image Understanding: CVIU, 144 (??):73–105, March 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314215002209.

Ellis:2014:BIR

[EF14] Anna-Louise Ellis and James Ferryman. Biologicallyinspired robust motion segmentation using mutual information. Computer Vision and Image Understanding: CVIU, 122(??):47– 64, May 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214000228.

Eggert:1998:SRM

[EFF98] David W. Eggert, Andrew W. Fitzgibbon, and Robert B. Fisher. Simultaneous registration of multiple range views for use in reverse engineering of CAD models. Computer Vision and Image Understanding: CVIU, 69(3):253– 272, March 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0667/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0667/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0667/production/ref.

Eggers:1998:TFE

[Egg98] Hinnik Eggers. Two fast Euclidean distance transformations in Z^{2} based on sufficient propagation. Computer Vision and Image Understanding: CVIU, 69(1): 106–116, January 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0596/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0596/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0596/production/ref.

Guo:2007:PSI

[eGZW07] Cheng en Guo, Song-Chun Zhu, and Ying Nian Wu. Primal sketch: Integrating structure and texture. Computer Vision and Image Understanding: CVIU, 106(1): 5–19, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Elwarfalli:2021:FCN

[EH21] Hamed Elwarfalli and Russell C. Hardie. FIFNET: a convolutional neural network for motion-based multiframe super-resolution using fusion of interpolated frames. Computer Vision and Image Understanding: CVIU, 202(??):Article 103097, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220301260.

Escalante:2010:SAI

[EHG^{+}10] Hugo Jair Escalante, Carlos A. Hernández, Jesus A. Gonzalez, A. López-López, Manuel Montes, Eduardo F. Morales, L. Enrique Sucar, Luis Villaseñor, and Michael Grubinger. The segmented and annotated IAPR TC-12 benchmark. Computer Vision and Image Understanding: CVIU, 114(4):419– 428, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

EbroulIzquierdo:1998:IAM

[EK98] M. Ebroul Izquierdo and Silko Kruse. Image analysis for 3D modeling, rendering, and virtual view generation. Computer Vision and Image Understanding: CVIU, 71(2):231– 253, August 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0706/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0706/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0706/production/ref.

Escobar:2012:ARB

[EK12] María-José Escobar and Pierre Kornprobst. Action recognition via bio-inspired features: The richness of center-surround interaction. Computer Vision and Image Understanding: CVIU, 116 (5):593–605, May 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000185.

Ethier:2014:SMT

[EK14] Marc Ethier and Tomasz Kaczynski. Suspension models for testing shape similarity methods. Computer Vision and Image Understanding: CVIU, 121 (??):13–20, April 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213002385.

Ebrahimpour:2008:TDL

[EKY08] Reza Ebrahimpour, Ehsanollah Kabir, and Mohammad Reza Yousefi. Teacherdirected learning in viewindependent face recognition with mixture of experts using overlapping eigenspaces. Computer Vision and Image Understanding: CVIU, 111(2):195–206, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Eckhardt:2003:TDS

[EL03] Ulrich Eckhardt and Longin Jan Latecki. Topologies for the digital spaces and. Computer Vision and Image Understanding: CVIU, 90(3):295–312, June 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Elgammal:2007:NML

[EL07] Ahmed Elgammal and Chan-Su Lee. Nonlinear manifold learning for dynamic shape and dynamic appearance. Computer Vision and Image Understanding: CVIU, 106(1):31– 46, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ehrhardt:2019:TVM

[EMMV19] Sebastien Ehrhardt, Aron Monszpart, Niloy J. Mitra, and Andrea Vedaldi. Taking visual motion prediction to new heightfields. Computer Vision and Image Understanding: CVIU, 181(??): 14–25, April 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219300207.

Elich:2022:WSL

[EOPS22] Cathrin Elich, Martin R.

Oswald, Marc Pollefeys, and Joerg Stueckler. Weakly supervised learning of multiobject 3D scene decompositions using deep shape priors. Computer Vision and Image Understanding: CVIU, 220(??): ??, July 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000583.

ElJurdi:2021:HLP

[EPH^{+}21] Rosana El Jurdi, Caroline Petitjean, Paul Honeine, Veronika Cheplygina, and Fahed Abdallah. High-level prior-based loss functions for medical image segmentation: a survey. Computer Vision and Image Understanding: CVIU, 210(??):

??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000928.

Embrechts:1996:PED

[ER96] Hugo Embrechts and Dirk Roose. A parallel Euclidean distance transformation algorithm. Computer Vision and Image Understanding: CVIU, 63(1):15– 26, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0002/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0002/production/pdf.

Esteban:2004:SSF

[ES04] Carlos Hernández Esteban and Francis Schmitt. Silhouette and stereo fusion for 3D object modeling. Computer Vision and Image Understanding: CVIU, 96(3): 367–392, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Erdem:2006:ACL

[ES06] Ugur Murat Erdem and Stan Sclaroff. Automated camera layout to satisfy task-specific and floor planspecific coverage requirements. Computer Vision and Image Understanding: CVIU, 103(3):156–169, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ekenel:2010:VBD

[ESS10] Hazım Kemal Ekenel, Johannes Stallkamp, and Rainer Stiefelhagen. A video-based door monitoring system using local appearance-based face models. Computer Vision and Image Understanding: CVIU, 114(5):596–608, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Eskandari:2015:SOF

[ET15] Maryam Eskandari and Önsen Toygar. Selection of optimized features and weights on face-iris fusion using distance images. Computer Vision and Image Understanding: CVIU, 137(??):63– 75, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000454.

Evako:2006:TPC

[Eva06] Alexander V. Evako. Topological properties of closed digital spaces: One method of constructing digital models of closed continuous surfaces by using covers. Computer Vision and Image Understanding: CVIU, 102(2): 134–144, May 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Essa:2017:ASC

[EX17] Ehab Essa and Xianghua Xie. Automatic segmentation of cross-sectional coronary arterial images. Computer Vision and Image Understanding: CVIU, 165(??):97–110, December 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314217301881.

Edwards:2020:GCN

[EXP^{+}20] Michael Edwards, Xianghua Xie, Robert I. Palmer, Gary K. L. Tam, Rob Alcock, and Carl Roobottom. Graph convolutional neural network for multi-scale feature learning. Computer Vision and Image Understanding: CVIU, 194(??):Article 102881, May 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218304004.

Escalante:2011:EBM

[EyGS11] Hugo Jair Escalante, Manuel Montes y Goméz, and Luis Enrique Sucar. An energy-based model for region-labeling. Computer Vision and Image Understanding: CVIU, 115(6):787–803, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Feldmar:1997:PRF

[FAB97] Jacques Feldmar, Nicholas Ayache, and Fabienne Betting. 3D-2D projective registration of free-form curves and surfaces. Computer Vision and Image Understanding: CVIU, 65(3):403– 424, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0499/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0499/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0499/production/ref.

Fouquier:2012:SMB

[FAB12] Geoffroy Fouquier, Jamal Atif, and Isabelle Bloch. Sequential model-based segmentation and recognition of image structures driven by visual features and spatial relations. Computer Vision and Image Understanding: CVIU, 116(1): 146–165, January 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001998.

Farouk:2011:IRB

[Far11] R. M. Farouk. Iris recognition based on elastic graph matching and Gabor wavelets. Computer Vision and Image Understanding: CVIU, 115(8):1239– 1244, August 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001081.

Fakih:2014:AAS

[FAZ14] Adel Fakih, Daniel Asmar, and John Zelek. Augmenting analytic SFM filters with frame-to-frame features. Computer Vision and Image Understanding: CVIU, 129(??):1–14, December 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001829.

Fua:1997:IHC

[FB97] P. Fua and C. Brechbühler. Imposing hard constraints on deformable models through optimization in orthogonal subspaces. Computer Vision and Image Understanding: CVIU, 65(2):148– 162, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0568/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0568/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0568/production/ref.

Fiala:2005:PSR

[FB05] Mark Fiala and Anup Basu.

Panoramic stereo reconstruction using non-SVP optics. Computer Vision and Image Understanding: CVIU, 98(3):363–397, June 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ferraz:2012:SCB

[FB12] Luis Ferraz and Xavier Binefa. A sparse curvaturebased detector of affine invariant blobs. Computer Vision and Image Understanding: CVIU, 116(4): 524–537, April 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002621.

Fogelton:2016:EBD

[FB16] A. Fogelton and W. Benesova. Eye blink detection based on motion vectors analysis. Computer Vision and Image Understanding: CVIU, 148(??): 23–33, July 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300054.

Fogelton:2018:EBC

[FB18] Andrej Fogelton and Wanda Benesova. Eye blink completeness detection. Computer Vision and Image Understanding: CVIU, 176– 177(??):78–85, November/ December 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421830287X.

Faltemier:2008:UMI

[FBF08] Timothy C. Faltemier, Kevin W. Bowyer, and Patrick J. Flynn. Using multi-instance enrollment to improve performance of 3D face recognition. Computer Vision and Image Understanding: CVIU, 112(2): 114–125, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fortun:2015:OFM

[FBK15] Denis Fortun, Patrick Bouthemy, and Charles Kervrann. Optical flow modeling and computation: a survey. Computer Vision and Image Understanding: CVIU, 134 (??):1–21, May 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000429.

Fortun:2016:ALP

[FBK16] Denis Fortun, Patrick Bouthemy, and Charles Kervrann. Aggregation of local parametric candidates with exemplar-based occlusion handling for optical flow. Computer Vision and Image Understanding: CVIU, 145 (??):81–94, April 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002660.

Fond:2021:MIR

[FBS21] Antoine Fond, Marie-Odile Berger, and Gilles Simon. Model-image registration of a building’s facade based on dense semantic segmentation. Computer Vision and Image Understanding: CVIU, 206(??):Article 103185, May 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000291.

Fehr:2015:CBP

[FBZP15] Duc Fehr, William J. Beksi, Dimitris Zermas, and Nikolaos Papanikolopoulos. Covariance based point cloud descriptors for object detection and recognition. Computer Vision and Image Understanding: CVIU, 142 (??):80–93, ???? 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001368.

Furtado:2020:HVS

[FCM20] Pedro Furtado, Manuel Caldeira, and Pedro Martins. Human visual system vs convolution neural networks in food recognition task: an empirical comparison. Computer Vision and Image Understanding: CVIU, 191(??):Article 102878, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219301675.

Furuya:2024:SSL

[FCOK24] Takahiko Furuya, Zhoujie Chen, Ryutarou Ohbuchi, and Zhenzhong Kuang. Self-supervised learning of rotation-invariant 3D point set features using transformer and its self-distillation. Computer Vision and Image Understanding: CVIU, 244(??):??, July 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224001061.

Fejes:1999:DIM

[FD99] Sándor Fejes and Larry S.

Davis. Detection of independent motion using directional motion estimation. Computer Vision and Image Understanding: CVIU, 74 (2):101–120, May 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0751/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0751/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0751/production/ref.

Ferraris:2019:CDL

[FDC^{+}19] Vinicius Ferraris, Nicolas Dobigeon, Yanna Cavalcanti, Thomas Oberlin, and Marie Chabert. Coupled dictionary learning for unsupervised change detection between multimodal remote sensing images. Computer Vision and Image Understanding: CVIU, 189 (??):Article 102817, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219301274.

Feldmar:1997:EIA

[FDMA97] J. Feldmar, J. Declerck, G. Malandain, and N. Ayache. Extension of the ICP algorithm to nonrigid intensity-based registration of 3D volumes. Computer Vision and Image Understanding: CVIU, 66(2): 193–206, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0606/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0606/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0606/production/ref.

Frizza:2022:SAS

[FDSB22] Tristan Frizza, Donald G. Dansereau, Nagita Mehr Seresht, and Michael Bewley. Semantically accurate super-resolution Generative Adversarial Networks. Computer Vision and Image Understanding: CVIU, 221 (??):??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422200073X.

Fan:2021:RSS

[FDW21] Bin Fan, Yuchao Dai, and Ke Wang. Rolling-shutter-stereo-aware motion estimation and image correction. Computer Vision and Image Understanding: CVIU, 213(??):??, December 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221001405.

Farenzena:2009:SMG

[FF09] Michela Farenzena and Andrea Fusiello. Stabilizing 3D modeling with geometric constraints propagation. Computer Vision and Image Understanding: CVIU, 113(11):1147–1157, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Furnari:2023:SEA

[FF23] Antonino Furnari and Giovanni Maria Farinella. Streaming egocentric action anticipation: an evaluation scheme and approach. Computer Vision and Image Understanding: CVIU, 234 (??):??, September 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314223001431.

Fulgeri:2019:CAN

[FFA^{+}19] Federico Fulgeri, Matteo Fabbri, Stefano Alletto, Simone Calderara, and Rita Cucchiara. Can adversarial networks hallucinate occluded people with a plausible aspect? Computer Vision and Image Understanding: CVIU, 182(??): 71–80, May 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219300438.

Fei-Fei:2007:LGV

[FFFP07] Li Fei-Fei, Rob Fergus, and Pietro Perona. Learning generative visual models from few training examples: an incremental Bayesian approach tested on 101 object categories. Computer Vision and Image Understanding: CVIU, 106(1):59– 70, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ferri:2014:SSC

[FFL14] Massimo Ferri, Patrizio Frosini, and Claudia Landi. Special section on computational topology in image context. Computer Vision and Image Understanding: CVIU, 121(??): 1, April 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214000526.

Fasel:2005:GFR

[FFM05] Ian Fasel, Bret Fortenberry, and Javier Movellan. A generative framework for real time object detection and classification. Computer Vision and Image Understanding: CVIU, 98(1):182– 210, April 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fang:2004:ARS

[FFY^{+}04] C. Y. Fang, C. S. Fuh, P. S. Yen, S. Cherng, and S. W. Chen. An automatic road sign recognition system based on a computational model of human recognition processing. Computer Vision and Image Understanding: CVIU, 96(2):237–268, November 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Flynn:2001:SIE

[FHP01] Patrick J. Flynn, Adam Hoover, and P. Jonathon Phillips. Special issue on empirical evaluation of computer vision algorithms. Computer Vision and Image Understanding: CVIU, 84(1):1–4, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0948;   http:

                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0948/pdf.

Fan:2013:SSA

[FHSKP13] Guoliang Fan, Riad I. Hammoud, Firooz Sadjadi, and Behzad Kamgar-Parsi. Special section on Advances in Machine Vision Beyond the Visible Spectrum (BVS). Computer Vision and Image Understanding: CVIU, 117(12):1645–1646, December 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213001549.

Fang:2023:LDA

[FHZX23] Pengfei Fang, Yaojun Hu, Shipeng Zhu, and Hui Xue. On learning distribution alignment for video-based visible-infrared person reidentification. Computer Vision and Image Understanding: CVIU, 237(??): ??, December 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223002138.

Fielding:2000:CCO

[FK00] Gabriel Fielding and Moshe Kam. Computing the cost of occlusion. Computer Vision and Image Understanding: CVIU, 79 (2):324–329, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0853;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0853/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0853/ref.

Furukawa:2009:LRS

[FK09] Ryo Furukawa and Hiroshi Kawasaki. Laser range scanner based on self-calibration techniques using coplanarities and metric constraints. Computer Vision and Image Understanding: CVIU, 113(11):1118–1129, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fischer:1998:EBA

[FKL^{+}98] André Fischer, Thomas H.

Kolbe, Felicitas Lang, Armin B. Cremers, Wolfgang Förstner, Lutz Plümer, and Volker Steinhage. Extracting buildings from aerial images using hierarchical aggregation in 2D and 3D. Computer Vision and Image Understanding: CVIU, 72(2):185– 203, November 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/

                                        links/artid/cviu.1998.
                                        0721/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0721/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0721/production/ref.

Farinella:2016:SIAb

[FKL^{+}16a] Giovanni Maria Farinella, Takeo Kanade, Marco Leo, Gerard G. Medioni, and Mohan Trivedi. Special issue on assistive computer vision and robotics — “assistive solutions for mobility, communication and HMI”. Computer Vision and Image Understanding: CVIU, 149 (??):1–2, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300674.

Farinella:2016:SIAa

[FKL^{+}16b] Giovanni Maria Farinella, Takeo Kanade, Marco Leo, Gerard G. Medioni, and Mohan Trivedi. Special issue on assistive computer vision and robotics — Part I. Computer Vision and Image Understanding: CVIU, 148(??):1–2, July 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300571.

Fritz:2010:TBL

[FKS10] Mario Fritz, Geert-Jan M.

Kruijff, and Bernt Schiele. Tutor-based learning of visual categories using different levels of supervision. Computer Vision and Image Understanding: CVIU, 114(5):564–573, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ferrer:2011:GFM

[FKV^{+}11] M. Ferrer, D. Karatzas, E. Valveny, I. Bardaji, and H. Bunke. A generic framework for median graph computation based on a recursive embedding approach. Computer Vision and Image Understanding: CVIU, 115 (7):919–928, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000786.

Fieguth:1998:EMC

[FKW98] Paul W. Fieguth, William C.

Karl, and Alan S. Willsky. Efficient multiresolution counterparts to variational methods for surface reconstruction. Computer Vision and Image Understanding: CVIU, 70(2): 157–176, May 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/

                                        links/artid/cviu.1997.
                                        0630/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0630/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0630/production/ref.

Fua:1996:TAI

[FL96] P. Fua and Y. G. Leclerc. Taking advantage of imagebased and geometry-based constraints to recover 3-D surfaces. Computer Vision and Image Understanding: CVIU, 64(1): 111–127, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0048/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0048/production/pdf.

Fan:2009:FEM

[FL09] Shu-Kai S. Fan and Yen Lin. A fast estimation method for the generalized Gaussian mixture distribution on complex images. Computer Vision and Image Understanding: CVIU, 113(7): 839–853, July 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Figueroa:2006:TSP

[FLB06] Pascual J. Figueroa, Neucimar J. Leite, and Ricardo M. L. Barros. Tracking soccer players aiming their kinematical motion analysis. Computer Vision and Image Understanding: CVIU, 101 (2):122–135, February 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Felisberto:2006:ODR

[FLCdA06] Marcelo Kleber Felisberto, Heitor Silvério Lopes, Tania Mezzadri Centeno, and Lúcia Valéria Ramos de Arruda. An object detection and recognition system for weld bead extraction from digital radiographs. Computer Vision and Image Understanding: CVIU, 102(3):238– 249, June 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fu:2008:LAS

[FLHK08] Yun Fu, Zhu Li, Thomas S.

Huang, and Aggelos K. Katsaggelos. Locally adaptive subspace and similarity metric learning for visual data clustering and retrieval. Computer Vision and Image Understanding: CVIU, 110(3):390– 402, June 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fu:2023:SRB

[FLL^{+}23] Hongtao Fu, Wenze Liu, Yuliang Liu, Zhiguo Cao, and Hao Lu. SIERRA: a robust bilateral feature upsampler for dense prediction. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422300142X.

Fergani:2014:HST

[FLS^{+}14] K. Fergani, D. Lui, C. Scharfenberger, A. Wong, and D. A. Clausi. Hybrid structural and texture distinctiveness vector field convolution for region segmentation. Computer Vision and Image Understanding: CVIU, 125(??):85– 96, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214000770.

Fua:1999:AHO

[FM99] P. Fua and C. Miccio. Animated heads from ordinary images: a least-squares approach. Computer Vision and Image Understanding: CVIU, 75(3):247– 259, September 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0778/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0778/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0778/production/ref.

Fard:2022:FLP

[FM22] Ali Pourramezan Fard and Mohammad H. Mahoor. Facial landmark points detection using knowledge distillation-based neural networks. Computer Vision and Image Understanding: CVIU, 215(??): ??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001582.

Fu:2023:HIP

[FMG23] Yuanbin Fu, Jiayi Ma, and Xiaojie Guo. Hierarchical image peeling: a flexible scale-space filtering framework. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001492.

Falomir:2012:MQD

[FMGA^{+}12] Zoe Falomir, Lledó Museros, Luis Gonzalez-Abril, M. Teresa Escrig, and Juan A. Ortega. A model for the qualitative description of images based on visual and spatial features. Computer Vision and Image Understanding: CVIU, 116(6): 698–714, June 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421200032X.

Fradkin:2001:BDM

[FMR01] M. Fradkin, H. Maître, and M. Roux. Building detection from multiple aerial images in dense urban areas. Computer Vision and Image Understanding: CVIU, 82(3):181–207, June 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2001.0917;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0917/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0917/ref.

Fitschen:2017:RCE

[FMS17] Jan Henrik Fitschen, Jianwei Ma, and Sebastian Schuff. Removal of curtaining effects by a variational model with directional forward differences. Computer Vision and Image Understanding: CVIU, 155(??): 24–32, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216302065.

Firouzi:2014:ERM

[FN14] Hadi Firouzi and Homayoun Najjaran. Efficient and robust multi-template tracking using multi-start interactive hybrid search. Computer Vision and Image Understanding: CVIU, 120(??):70–80, March 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314213002403.

Furuya:2018:LPW

[FO18] Takahiko Furuya and Ryutarou Ohbuchi. Learning part-in-whole relation of 3D shapes for part-based 3D model retrieval. Computer Vision and Image Understanding: CVIU, 166(??): 102–114, January 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301984.

Freire-Obregon:2020:ARM

[FOCSB^{+}20] David Freire-Obregón, Modesto Castrillón-Santana, Paola Barra, Carmen Bisogni, and Michele Nappi. An attention recurrent model for human cooperation detection. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 102991, August 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422030062X.

Fablet:2008:IBR

[FPC^{+}08] R. Fablet, S. Pujolle, A. Chessel, A. Benzinou, and F. Cao. 2D imagebased reconstruction of shape deformation of biological structures using a levelset representation. Computer Vision and Image Understanding: CVIU, 111(3): 295–306, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Frahm:2012:SIV

[FPDK12] Jan-Michael Frahm, Marc Pollefeys, Frank Dellaert, and Jana Kosecka. Special issue on Virtual Representations and Modeling of Large-scale environments (VRML). Computer Vision and Image Understanding: CVIU, 116(1): 1, January 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002402.

Fu:2019:PAS

[FPMK19] Junsheng Fu, Said Pertuz, Jiri Matas, and Joni-Kristian Kämäräinen. Performance analysis of singlequery 6-DoF camera pose estimation in self-driving setups. Computer Vision and Image Understanding: CVIU, 186(??):58–73, September 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S1077314219300633.

Fabbrizzi:2022:SBV

[FPNK22] Simone Fabbrizzi, Symeon Papadopoulos, Eirini Ntoutsi, and Ioannis Kompatsiaris. A survey on bias in visual datasets. Computer Vision and Image Understanding: CVIU, 223(??): ??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001308.

Filipovych:2011:RSA

[FR11] Roman Filipovych and Eraldo Ribeiro. Robust sequence alignment for actor– object interaction recognition: Discovering actor– object states. Computer Vision and Image Understanding: CVIU, 115(2):177–193, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Faugeras:1998:DRU

[FRL^{+}98] Olivier Faugeras, Luc Robert, Stéphane Laveau, Gabriella Csurka, Cyril Zeller, Cyrille Gauclin, and Imad Zoghlami. 3-D reconstruction of urban scenes from image sequences. Computer Vision and Image Understanding: CVIU, 69(3):292– 309, March 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0665/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0665/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0665/production/ref.

Frintrop:2005:BLB

[FRNS05] Simone Frintrop, Erich Rome, Andreas Nüchter, and Hartmut Surmann. A Bimodal Laser-Based Attention System. Computer Vision and Image Understanding: CVIU, 100 (1–2):124–151, October/ November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Furukawa:2003:ARL

[FS03] Yasutaka Furukawa and Yoshihisa Shinagawa. Accurate and robust line segment extraction by analyzing distribution around peaks in Hough space. Computer Vision and Image Understanding: CVIU, 92(1):1– 25, October 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fermuller:2001:SOF

[FSA01] Cornelia Fermüller, David Shulman, and Yiannis Aloimonos. The statistics of optical flow. Computer Vision and Image Understanding: CVIU, 82(1): 1–32, April 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0900;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0900/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0900/ref.

Feigin:2022:CG

[FSG22] Yuri Feigin, Hedva Spitzer, and Raja Giryes. Cluster with GANs. Computer Vision and Image Understanding: CVIU, 225(??): ??, December 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001497.

Fujimura:2021:DCV

[FSI21] Yuki Fujimura, Motoharu Sonogashira, and Masaaki Iiyama. Dehazing cost volume for deep multi-view stereo in scattering media with airlight and scattering coefficient estimation. Computer Vision and Image Understanding: CVIU, 211 (??):??, October 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000977.

Falcon:2024:ISV

[FSL24] Alex Falcon, Giuseppe Serra, and Oswald Lanz. Improving semantic video retrieval models by training with a relevance-aware online mining strategy. Computer Vision and Image Understanding: CVIU, 245 (??):??, August 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224001164.

Fransens:2007:OFB

[FSV07] Rik Fransens, Christoph Strecha, and Luc Van Gool. Optical flow based superresolution: a probabilistic approach. Computer Vision and Image Understanding: CVIU, 106(1):106– 115, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fidrich:1998:SCP

[FT98] Márta Fidrich and Jean-Philippe Thirion. Stability of corner points in scale space: The effects of small nonrigid deformations. Computer Vision and Image Understanding: CVIU, 72(1):72– 83, October 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0661/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0661/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0661/production/ref.

Farhand:2023:FDS

[FT23] Sepehr Farhand and Gavriil Tsechpenakis. Foreground discovery in streaming videos with dynamic construction of content graphs. Computer Vision and Image Understanding: CVIU, 227 (??):??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001989.

Fernando:2015:LRL

[FTT15] Basura Fernando, Tatiana Tommasi, and Tinne Tuytelaars. Location recognition over large time lags. Computer Vision and Image Understanding: CVIU, 139(??):21–28, October 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421500137X.

Fan:1997:SCS

[FW97] Joel Fan and Lawrence B. Wolff. Surface curvature and shape reconstruction from unknown multiple illumination and integrability. Computer Vision and Image Understanding: CVIU, 65(2):347– 359, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0581/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0581/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0581/production/ref.

Fan:2018:HCS

[FWG18] Jingzhe Fan, Yan Wang, and Lei Guo. Hierarchical coherency sensitive hashing and interpolation with RANSAC for large displacement optical flow. Computer Vision and Image Understanding: CVIU, 175(??): 1–10, October 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218304223.

Fan:2020:VAD

[FWL^{+}20] Yaxiang Fan, Gongjian Wen, Deren Li, Shaohua Qiu, Martin D. Levine, and Fei Xiao. Video anomaly detection and localization via Gaussian mixture fully convolutional variational autoencoder. Computer Vision and Image Understanding: CVIU, 195(??):Article 102920, June 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218302674.

Fang:2023:IVI

[FWLQ23] Aiqing Fang, Junsheng Wu, Ying Li, and Ruimin Qiao. Infrared and visible image fusion via mutual information maximization. Computer Vision and Image Understanding: CVIU, 231 (??):??, June 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223000632.

Fang:2024:UMF

[FWZ^{+}24] Lixing Fang, Xiangxiang Wang, Junli Zhao, Zhenkuan Pan, Hui Li, and Yi Li. An unsupervised multi-focus image fusion method via dual-channel convolutional network and discriminator. Computer Vision and Image Understanding: CVIU, 244(??):??, July 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224001103.

Fang:2017:OHT

[FXWW17] Jianwu Fang, Hongke Xu, Qi Wang, and Tianjun Wu. Online hash tracking with spatio-temporal saliency auxiliary. Computer Vision and Image Understanding: CVIU, 160 (??):57–72, July 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300565.

Feldman:2006:HIP

[FY06] Thomas Feldman and Laurent Younes. Homeostatic image perception: an artificial system. Computer Vision and Image Understanding: CVIU, 102(1):70– 80, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fu:2011:SIF

[FYH11] Yun (Raymond) Fu, Shuicheng Yan, and Thomas S. Huang. Special issue on featureoriented image and video computing for extracting contexts and semantics. Computer Vision and Image Understanding: CVIU, 115 (3):289, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fang:2020:LDE

[FZ20] Yingying Fang and Tieyong Zeng. Learning deep edge prior for image denoising. Computer Vision and Image Understanding: CVIU, 200(??):Article 103044, November 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220300916.

Goncalves:2009:EPN

[GA09] Nuno Gonçalves and Helder Araújo. Estimating parameters of noncentral catadioptric systems using bundle adjustment. Computer Vision and Image Understanding: CVIU, 113(1):11– 28, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ghanem:2013:MDS

[GA13] Bernard Ghanem and Narendra Ahuja. Modeling dynamic swarms. Computer Vision and Image Understanding: CVIU, 117(1):1– 11, January 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001282.

Gutierrez:2001:SDV

[GAD01] J. Gutiérrez, G. Ayala, and M. E. Díaz. Set descriptors for visual evaluation of human corneal endothelia. Computer Vision and Image Understanding: CVIU, 84(2): 249–263, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0936;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0936/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0936/ref.

Gavrila:1999:VAH

[Gav99] D. M. Gavrila. The visual analysis of human movement: a survey. Computer Vision and Image Understanding: CVIU, 73(1):82– 98, January 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0716/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0716/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0716/production/ref.

Golland:1997:MC

[GB97] P. Golland and A. M.

Bruckstein. Motion from color. Computer Vision and Image Understanding: CVIU, 68(3):346–362, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0553/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0553/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0553/production/ref.

Gilbert:2008:IST

[GB08] Andrew Gilbert and Richard Bowden. Incremental, scalable tracking of objects inter camera. Computer Vision and Image Understanding: CVIU, 111(1):43– 58, July 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Galleguillos:2010:CBO

[GB10] Carolina Galleguillos and Serge Belongie. Context based object categorization: a critical survey. Computer Vision and Image Understanding: CVIU, 114(6): 712–722, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goncalves:2013:DTS

[GB13] Wesley Nunes Gonçalves and Odemir Martinez Bruno. Dynamic texture segmentation based on deterministic partially self-avoiding walks. Computer Vision and Image Understanding: CVIU, 117(9):1163–1174, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213000921.

Gilbert:2017:IVM

[GB17] Andrew Gilbert and Richard Bowden. Image and video mining through online learning. Computer Vision and Image Understanding: CVIU, 158(??):72– 84, May 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300292.

Ghassab:2022:PPV

[GB22] Vahid Khorasani Ghassab and Nizar Bouguila. Plug-and-play video superresolution using edge-preserving filtering. Computer Vision and Image Understanding: CVIU, 216(??): ??, February 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000029.

Gros:1998:ULP

[GBB98] Patrick Gros, Olivier Bournez, and Edmond Boyer. Using local planar geometric invariants to match and model images of line segments. Computer Vision and Image Understanding: CVIU, 69(2):135– 155, February 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0565/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0565/production/pdf;
http://www.idealibrary.

                                        com/links/artid/cviu.1997.
                                        0565/production/ref.

Ghorbel:2018:EKL

[GBB^{+}18] Enjie Ghorbel, Jacques Boonaert, Rémi Boutteau, Stéphane Lecoeuche, and Xavier Savatier. An extension of kernel learning methods using a modified Log-Euclidean distance for fast and accurate skeletonbased human action recognition. Computer Vision and Image Understanding: CVIU, 175(??):32–43, October 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218302649.

Gehrig:2012:ISP

[GBF12] Stefan K. Gehrig, Hernán Badino, and Uwe Franke. Improving sub-pixel accuracy for long range stereo. Computer Vision and Image Understanding: CVIU, 116(1):16–24, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211001792.

Gheissari:2006:PMB

[GBHS06] Niloofar Gheissari, Alireza Bab-Hadiashar, and David Suter. Parametric modelbased motion segmentation using surface selection criterion. Computer Vision and Image Understanding: CVIU, 102(2):214–226, May 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gouet-Brunet:2008:ORS

[GBL08] Valérie Gouet-Brunet and Bruno Lameyre. Object recognition and segmentation in videos by connecting heterogeneous visual features. Computer Vision and Image Understanding: CVIU, 111(1):86– 109, July 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gonzalez-Briones:2018:MAS

[GBVDC18] Alfonso González-Briones, Gabriel Villarrubia, Juan F. De Paz, and Juan M. Corchado. A multi-agent system for the classification of gender and age from images. Computer Vision and Image Understanding: CVIU, 172 (??):98–106, July 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300122.

Gard:2021:SMO

[GBY21] Nima A. Gard, Colin Bunker, and Alper Yilmaz. A spacetime model for one-shot active contour extraction scheme for human detection in image sequences. Computer Vision and Image Understanding: CVIU, 202(??):Article 103113, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422030134X.

Guclu:2019:SFR

[GC19] Oguzhan Guclu and Ahmet Burak Can. k-SLAM: a fast RGB-D SLAM approach for large indoor environments. Computer Vision and Image Understanding: CVIU, 184(??):31– 44, July 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S107731421930058X.

Grechka:2024:GGG

[GCC24] Asya Grechka, Guillaume Couairon, and Matthieu Cord. GradPaint: Gradientguided inpainting with diffusion models. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000092.

Giachetti:2018:NFH

[GCD^{+}18] Andrea Giachetti, Irina Mihaela Ciortan, Claudia Daffara, Giacomo Marchioro, Ruggero Pintus, and Enrico Gobbetti. A novel framework for highlight reflectance transformation imaging. Computer Vision and Image Understanding: CVIU, 168(??):118– 131, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301091.

Gomez:2007:PAU

[GCEC07] David Delgado Gomez, Line Harder Clemmensen, Bjarne K. Ersbøll, and Jens Michael Carstensen. Precise acquisition and unsupervised segmentation of multi-spectral images. Computer Vision and Image Understanding: CVIU, 106(2– 3):183–193, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gouiffes:2012:SLP

[GCFMT12] M. Gouiffès, C. Collewet, C. Fernandez-Maloigne, and A. Trémeau. A study on local photometric models and their application to robust tracking. Computer Vision and Image Understanding: CVIU, 116(8):

896–907, August 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421200063X.

Gosselin:2008:CVD

[GCPF08] Philippe Henri Gosselin, Matthieu Cord, and Sylvie Philipp-Foliguet. Combining visual dictionary, kernelbased similarity and learning strategy for image category retrieval. Computer Vision and Image Understanding: CVIU, 110(3): 403–417, June 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Garcia-Cobo:2023:HSC

[GCS23] Guillermo Garcia-Cobo and Juan C. SanMiguel. Human skeletons and change detection for efficient violence detection in surveillance videos. Computer Vision and Image Understanding: CVIU, 233(??): ??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001194.

Guo:2014:OTU

[GCT^{+}14] Yanwen Guo, Ye Chen, Feng Tang, Ang Li, Weitao Luo, and Mingming Liu. Object tracking using learned feature manifolds. Computer Vision and Image Understanding: CVIU, 118(??): 128–139, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001835.

Grogan:2019:DRC

[GD19] Mairéad Grogan and Rozenn Dahyot. L_{2} divergence for robust colour transfer. Computer Vision and Image Understanding: CVIU, 181 (??):39–49, April 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219300177.

Guislain:2017:FSI

[GDCM17] Maximilien Guislain, Julie Digne, Raphaëlle Chaine, and Gilles Monnier. Fine scale image registration in large-scale urban LIDAR point sets. Computer Vision and Image Understanding: CVIU, 157(??):90– 102, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216302016.

Gonzalez-Diaz:2011:IRC

[GDIIHK11] Rocio Gonzalez-Diaz, Adrian Ion, Mabel Iglesias-Ham, and Walter G. Kropatsch.

Invariant representative cocycles of cohomology generators using irregular graph pyramids. Computer Vision and Image Understanding: CVIU, 115(7):1011– 1022, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211000774.

Gritti:2014:SVS

[GDM14] Tommaso Gritti, Chris Damkat, and Gianluca Monaci. Semantic video scene segmentation and transfer. Computer Vision and Image Understanding: CVIU, 122(??):172– 181, May 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214000344.

Greenspan:2004:CDS

[GDR04] Hayit Greenspan, Guy Dvir, and Yossi Rubner. Context-dependent segmentation and matching in image databases. Computer Vision and Image Understanding: CVIU, 93(1):86– 109, January 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gimenez:2008:EAM

[GE08] David Gimenez and Adrian N.

Evans. An evaluation of area morphology scale-spaces for colour images. Computer Vision and Image Understanding: CVIU, 110(1):32– 42, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Green:1995:GRA

[GESB95] Kevin Green, David Eggert, Louise Stark, and Kevin Bowyer. Generic recognition of articulated objects through reasoning about potential function. Computer Vision and Image Understanding: CVIU, 62(2):177– 193, September 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1049/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1049/production/pdf.

Ghanei:2015:LST

[GF15] Shaho Ghanei and Karim Faez. Localizing scene texts by fuzzy inference systems and low rank matrix recovery model. Computer Vision and Image Understanding: CVIU, 142(??): 94–110, ???? 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421500212X.

Gao:2011:AIS

[GFL^{+}11] Xinbo Gao, Rong Fu, Xuelong Li, Dacheng Tao, Beichen Zhang, and Huigen Yang. Aurora image segmentation by combining patch and texture thresholding. Computer Vision and Image Understanding: CVIU, 115(3):390– 402, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Galteri:2019:DMM

[GFL^{+}19] Leonardo Galteri, Claudio Ferrari, Giuseppe Lisanti, Stefano Berretti, and Alberto Del Bimbo. Deep 3D morphable model refinement via progressive growing of conditional Generative Adversarial Networks. Computer Vision and Image Understanding: CVIU, 185(??):31–42, August 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219300773.

Gao:2024:CSL

[GFLW24] Liqing Gao, Wei Feng, Fan Lyu, and Liang Wan. Combinational sign language recognition. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000535.

Gevers:2004:CII

[GFS04] Theo Gevers, Graham Finlayson, and Raimondo Schettini. Color for image indexing and retrieval. Computer Vision and Image Understanding: CVIU, 94 (1–3):1–2, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gu:2013:ASU

[GFW13] Xiaodong Gu, Yu Fang, and Yuanyuan Wang. Attention selection using global topological properties based on pulse coupled neural network. Computer Vision and Image Understanding: CVIU, 117(10):1400– 1411, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001045.

Gong:2014:JVI

[GFY^{+}14] Jiulu Gong, Guoliang Fan, Liangjiang Yu, Joseph P. Havlicek, Derong Chen, and Ningjun Fan. Joint viewidentity manifold for infrared target tracking and recognition. Computer Vision and Image Understanding: CVIU, 118(??):211– 224, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213001859.

Gevrekci:2009:IRI

[GG09] Murat Gevrekci and Bahadir K. Gunturk. Illumination robust interest point detection. Computer Vision and Image Understanding: CVIU, 113(4):565– 571, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goldman:2020:CDC

[GG20] Eran Goldman and Jacob Goldberger. CRF with deep class embedding for large scale classification. Computer Vision and Image Understanding: CVIU, 191(??):Article 102865, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219301560.

Garcia-Garcia:2017:SEN

[GGGROE^{+}17] Alberto Garcia-Garcia, Jose Garcia-Rodriguez, Sergio Orts-Escolano, Sergiu Oprea, Francisco Gomez-Donoso, and Miguel Cazorla. A study of the effect of noise and occlusion on the accuracy of convolutional neural networks applied to 3D object recognition. Computer Vision and Image Understanding: CVIU, 164(??):124– 134, November 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301182.

Geng:2024:TRP

[GGH^{+}24] Xiaoyu Geng, Qiang Guo, Shuaixiong Hui, Ming Yang, and Caiming Zhang. Tensor robust PCA with nonconvex and nonlocal regularization. Computer Vision and Image Understanding: CVIU, 243(??):??, June 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000882.

Gomez-Garcia:2008:IRB

[GGMV08] Héctor Fernando Gómez-García, José L. Marroquín, and Johan Van Horebeek. Image registration based on kernel-predictability. Computer Vision and Image Understanding: CVIU, 112(2): 160–172, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Grossmann:2010:DCC

[GGO10] Etienne Grossmann, José António Gaspar, and Francesco Orabona. Discrete camera calibration from pixel streams. Computer Vision and Image Understanding: CVIU, 114(2):198–209, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gautam:2023:WEI

[GGP23] Sidharth Gautam, Tapan Kumar Gandhi, and B. K. Panigrahi. WMCP-EM: an integrated dehazing framework for visibility restoration in single image. Computer Vision and Image Understanding: CVIU, 229 (??):??, March 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223000280.

Greenspan:2001:CPF

[GGR01] Hayit Greenspan, Jacob Goldberger, and Lenny Ridel. A continuous probabilistic framework for image matching. Computer Vision and Image Understanding: CVIU, 84(3):384–406, December 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gondra:2008:CBI

[GH08] Iker Gondra and Douglas R.

Heisterkamp. Contentbased image retrieval with the normalized information distance. Computer Vision and Image Understanding: CVIU, 111(2):219–228, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gao:2010:HIC

[GHA10] Chunyu Gao, Hong Hua, and Narendra Ahuja. A hemispherical imaging camera. Computer Vision and Image Understanding: CVIU, 114(2):168–178, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gong:2004:MEM

[GHHX04] Yihong Gong, Mei Han, Wei Hua, and Wei Xu. Maximum entropy model-based baseball highlight detection and classification. Computer Vision and Image Understanding: CVIU, 96(2): 181–199, November 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gong:2017:NCS

[GHML17] Ming Gong, You Hao, Hanlin Mo, and Hua Li. Naturally combined shapecolor moment invariants under affine transformations. Computer Vision and Image Understanding: CVIU, 162(??):46–56, September 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301376.

Gros:1997:HUP

[GHMQ97] P. Gros, R. Hartley, R. Mohr, and L. Quan. How useful is projective geometry? Computer Vision and Image Understanding: CVIU, 65(3): 442–446, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0496/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0496/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0496/production/ref.

Godin:2009:SIN

[GHMT09] Guy Godin, Patrick Hébert, Takeshi Masuda, and Gabriel Taubin. Special issue on new advances in 3D imaging and modeling. Computer Vision and Image Understanding: CVIU, 113(11):1105–1106, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goutsias:1995:MOI

[GHS95] John Goutsias, Henk J.

A. M. Heijmans, and K. Sivakumar. Morphological operators for image sequences. Computer Vision and Image Understanding: CVIU, 62(3):326– 346, November 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1058/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1058/production/pdf.

Gao:2013:FEU

[GHZ^{+}13] Quanxue Gao, Xiujuan Hao, Qijun Zhao, Weiguo Shen, and Jingjie Ma. Feature extraction using twodimensional neighborhood margin and variation embedding. Computer Vision and Image Understanding: CVIU, 117(5):525– 531, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000027.

Gopalan:2010:CCL

[GJ10] Raghuraman Gopalan and David Jacobs. Comparing and combining lighting insensitive approaches for face recognition. Computer Vision and Image Understanding: CVIU, 114(1):135–145, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Galata:2001:LVL

[GJH01] Aphrodite Galata, Neil Johnson, and David Hogg. Learning variable-length Markov models of behavior. Computer Vision and Image Understanding: CVIU, 81(3):398–413, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0894;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0894/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0894/ref.

Gutierrez:2014:TEV

[GJMO14] Antonio Gutierrez, Maria Jose Jimenez, David Monaghan, and Noel E. O’Connor. Topological evaluation of volume reconstructions by voxel carving. Computer Vision and Image Understanding: CVIU, 121(??): 27–35, April 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213002233.

Gunsel:1996:RBD

[GJP96] Bilge Günsel, Anil K. Jain, and Erdal Panayirci. Reconstruction and boundary detection of range and intensity images using multiscale MRF representations. Computer Vision and Image Understanding: CVIU, 63(2): 353–366, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0025/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0025/production/pdf.

Grewe:1995:ILM

[GK95] Lynne Grewe and Avinash C.

Kak. Interactive learning of a multiple-attribute hash table classifier for fast object recognition. Computer Vision and Image Understanding: CVIU, 61(3): 387–416, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1030/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1030/production/pdf.

Ghosh:1998:SFR

[GK98] Pijush K. Ghosh and K. Vinod Kumar. Support function representation of convex bodies, its application in geometric computing, and some related representations. Computer Vision and Image Understanding: CVIU, 72(3):379– 403, December 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0674/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0674/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0674/production/ref.

Gaikwad:2023:RTD

[GK23] Bipin Gaikwad and Abhijit Karmakar. Real-time distributed video analytics for privacy-aware person search. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001297.

Gollmer:2014:UIS

[GKBW14] Sebastian T. Gollmer, Matthias Kirschner, Thorsten M. Buzug, and Stefan Wesarg. Using image segmentation for evaluating 3D statistical shape models built with groupwise correspondence optimization. Computer Vision and Image Understanding: CVIU, 125(??):283– 303, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001015.

Groenendijk:2020:BAT

[GKGM20] Rick Groenendijk, Sezer Karaoglu, Theo Gevers, and Thomas Mensink. On the benefit of adversarial training for monocular depth estimation. Computer Vision and Image Understanding: CVIU, 190(??):Article 102848, January 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219301456.

Guo:2021:ISD

[GKH^{+}21] Xuyang Guo, Meina Kan, Zhenliang He, Xingguang Song, and Shiguang Shan. Image style disentangling for instance-level facial attribute transfer. Computer Vision and Image Understanding: CVIU, 207(??): ??, June 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221000497.

Gao:2005:MKF

[GKK05] Jean Gao, Akio Kosaka, and Avinash C. Kak. A multi-Kalman filtering approach for video tracking of human-delineated objects in cluttered environments. Computer Vision and Image Understanding: CVIU, 99(1):1–57, July 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See erratum [Ano06h].

Girshick:2017:EDL

[GKL^{+}17] Ross Girshick, Iasonas Kokkinos, Ivan Laptev, Jitendra Malik, George Papandreou, Andrea Vedaldi, Xiaogang Wang, Shuicheng Yan, and Alan Yuille. Editorial- deep learning for computer vision. Computer Vision and Image Understanding: CVIU, 164(??): 1–2, November 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301972.

Giordano:2015:NLP

[GKPS15] D. Giordano, I. Kavasidis, S. Palazzo, and C. Spampinato. Nonparametric label propagation using mutual local similarity in nearest neighbors. Computer Vision and Image Understanding: CVIU, 131(??):116– 127, February 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001325.

Gross:1995:DPT

[GL95] Ari Gross and Longin Latecki. Digitizations preserving topological and differential geometric properties. Computer Vision and Image Understanding: CVIU, 62(3):370– 381, November 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1061/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1061/production/pdf.

Gross:1997:RDM

[GL97] Ari Gross and Longin Jan Latecki. A realistic digitization model of straight lines. Computer Vision and Image Understanding: CVIU, 67(2):131– 142, August 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0530/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0530/production/pdf;
http://www.idealibrary.

                                        com/links/artid/cviu.1997.
                                        0530/production/ref.

Guo:1998:DVS

[GL98] Baining Guo and Joseph Liu. Direct visible surface interpolation. Computer Vision and Image Understanding: CVIU, 72(3):328– 339, December 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0668/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0668/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0668/production/ref.

Gao:2019:CHI

[GL19] Qishuo Gao and Samsung Lim. Classification of hyperspectral images with convolutional neural networks and probabilistic relaxation. Computer Vision and Image Understanding: CVIU, 188(??):Article 102801, November 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S107731421930116X.

Guo:2024:DAT

[GL24] Kaiyu Guo and Brian C. Lovell. Domain-aware triplet loss in domain generalization. Computer Vision and Image Understanding: CVIU, 243(??): ??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000602.

Gonthier:2022:MIL

[GLG22] Nicolas Gonthier, Saïd Ladjal, and Yann Gousseau. Multiple instance learning on deep features for weakly supervised object detection with extreme domain shifts. Computer Vision and Image Understanding: CVIU, 214 (??):??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001430.

Gerke:2017:SPR

[GLM17] Sebastian Gerke, Antje Linnemann, and Karsten Müller. Soccer player recognition using spatial constellation features and jersey number recognition. Computer Vision and Image Understanding: CVIU, 159(??):105– 115, June 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300723.

Gori:2016:SVL

[GLMM16] Marco Gori, Marco Lippi, Marco Maggini, and Stefano Melacci. Semantic video labeling by developmental visual agents. Computer Vision and Image Understanding: CVIU, 146(??): 9–26, May 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216000655.

Germa:2010:VRD

[GLOC10] T. Germa, F. Lerasle, N. Ouadah, and V. Cadenat. Vision and RFID data fusion for tracking people in crowds by a mobile robot. Computer Vision and Image Understanding: CVIU, 114(6):641–651, June 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Glossop:1999:IHT

[GLR^{+}99] K. Glossop, P. J. G. Lisboa, P. C. Russell, A. Siddans, and G. R. Jones. An implementation of the Hough transformation for the identification and labelling of fixed period sinusoidal curves. Computer Vision and Image Understanding: CVIU, 74(1):96– 100, April 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0747/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0747/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0747/production/ref.

Guan:2023:PHP

[GLZF23] Shannan Guan, Haiyan Lu, Linchao Zhu, and Gengfa Fang. PoseGU: 3D human pose estimation with novel human pose generator and unbiased learning. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000954.

Galama:2019:IIG

[GM19] Ysbrand Galama and Thomas Mensink. IterGANs: Iterative GANs to learn and control 3D object transformation. Computer Vision and Image Understanding: CVIU, 189 (??):Article 102803, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219301171.

Gimenez:2014:UEM

[GMF14] Javier Gimenez, Jorge Martinez, and Ana Georgina Flesia. Unsupervised edge map scoring: a statistical complexity approach. Computer Vision and Image Understanding: CVIU, 122(??):131– 142, May 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214000319.

Gueziri:2016:GGR

[GML16] Houssem-Eddine Gueziri, Michael J. McGuffin, and Catherine Laporte. A generalized graph reduction framework for interactive segmentation of large images. Computer Vision and Image Understanding: CVIU, 150(??):44–57, September 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421630056X.

Guo:2021:MMP

[GML^{+}21] Yu Guo, Lichen Ma, Zhi Li, Xuan Wang, and Fei Wang. Monocular 3D multiperson pose estimation via predicting factorized correction factors. Computer Vision and Image Understanding: CVIU, 213(??):

??, December 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001223.

Garcia-Martin:2015:PPA

[GMM15] Álvaro García-Martín and José M. Martínez. Postprocessing approaches for improving people detection performance. Computer Vision and Image Understanding: CVIU, 133(??): 76–89, April 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001921.

Guimond:2000:ABM

[GMT00] Alexandre Guimond, Jean Meunier, and Jean-Philippe Thirion. Average brain models: a convergence study. Computer Vision and Image Understanding: CVIU, 77(2): 192–210, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0815;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0815/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0815/ref.

Gonzalez:2012:SUH

[GMW12] Jordi Gonzàlez, Thomas B. Moeslund, and Liang Wang. Semantic understanding of human behaviors in image sequences: From video-surveillance to videohermeneutics. Computer Vision and Image Understanding: CVIU, 116(3): 305–306, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212000100.

Gomes:2022:MHF

[GMZ^{+}22] Mouglas Eugênio Nasário Gomes, David Macêdo, Cleber Zanchettin, Paulo Salgado Gomes de Mattos-Neto, and Adriano Oliveira. Multi-human fall detection and localization in videos. Computer Vision and Image Understanding: CVIU, 220(??):??, July 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222000595.

Gruen:1998:ABE

[GN98] A. Gruen and R. Nevatia. Automatic building extraction from aerial images. Computer Vision and Image Understanding: CVIU, 72(2):99– 100, November 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0731/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0731/production/pdf.

Galego:2015:UAD

[GOF^{+}15] Ricardo Galego, Agustin Ortega, Ricardo Ferreira, Alexandre Bernardino, Juan Andrade-Cetto, and José Gaspar. Uncertainty analysis of the DLT-Lines calibration algorithm for cameras with radial distortion. Computer Vision and Image Understanding: CVIU, 140(??):115– 126, November 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001290.

Goh:2008:SSM

[Goh08] Wooi-Boon Goh. Strategies for shape matching using skeletons. Computer Vision and Image Understanding: CVIU, 110(3):326– 345, June 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goldberger:2005:RCP

[Gol05]  Jacob Goldberger.  Re-

constructing camera projection matrices from multiple pairwise overlapping views. Computer Vision and Image Understanding: CVIU, 97(3):283–296, March 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gong:2009:RTJ

[Gon09]  Minglun Gong.  Real-

time joint disparity and disparity flow estimation on programmable graphics hardware. Computer Vision and Image Understanding: CVIU, 113(1):90–100, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gottfried:2008:QSM

[Got08] Björn Gottfried. Qualitative similarity measures— The case of two-dimensional outlines. Computer Vision and Image Understanding: CVIU, 110(1):117– 133, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gumundsson:2010:IRS

[GPC^{+}10] Sigurjón Árni Guðmundsson, Montse Pardàs, Josep R. Casas, Jóhannes R. Sveinsson, Henrik Aanæs, and Rasmus Larsen. Improved 3D reconstruction in smartroom environments using ToF imaging. Computer Vision and Image Understanding: CVIU, 114(12): 1376–1384, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Grosgeorge:2013:GCS

[GPDR13] D. Grosgeorge, C. Petitjean, J.-N. Dacher, and S. Ruan. Graph cut segmentation with a statistical shape model in cardiac MRI. Computer Vision and Image Understanding: CVIU, 117(9):1027– 1035, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000738.

Gonfaus:2015:FAO

[GPG^{+}15] Josep M. Gonfaus, Marco Pedersoli, Jordi Gonzàlez, Andrea Vedaldi, and F. Xavier Roca. Factorized appearances for object detection. Computer Vision and Image Understanding: CVIU, 138(??):92–101, September 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000922.

Garcia:1999:UBM

[GPK99] P. García, M. Petrou, and S. Kamata. The use of Boolean model for texture analysis of grey images. Computer Vision and Image Understanding: CVIU, 74(3):227– 235, June 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0760/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0760/production/pdf.

Gong:2007:GOF

[GPY^{+}07] Haifeng Gong, Chunhong Pan, Qing Yang, Hanqing Lu, and Songde Ma. Generalized optical flow in the scale space. Computer Vision and Image Understanding: CVIU, 105(1):86–92, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gil:2005:EAO

[GR05] Debora Gil and Petia Radeva. Extending anisotropic operators to recover smooth shapes. Computer Vision and Image Understanding: CVIU, 99(1):110–125, July 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Godec:2013:HBT

[GRB13] M. Godec, P. M. Roth, and H. Bischof. Houghbased tracking of non-rigid objects. Computer Vision and Image Understanding: CVIU, 117(10):1245– 1256, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001890.

Gay:2018:FBS

[GRCD18] Paul Gay, Cosimo Rubino, Marco Crocco, and Alessio Del Bue. Factorization based structure from motion with object priors. Computer Vision and Image Understanding: CVIU, 172(??):124– 137, July 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301893.

Grevera:2004:DRS

[Gre04] George J. Grevera. The “dead reckoning” signed distance transform. Computer Vision and Image Understanding: CVIU, 95(3):317– 333, September 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Galbally:2013:IIR

[GRGB^{+}13] Javier Galbally, Arun Ross, Marta Gomez-Barrero, Julian Fierrez, and Javier Ortega-Garcia. Iris image reconstruction from binary templates: an efficient probabilistic approach based on genetic algorithms. Computer Vision and Image Understanding: CVIU, 117(10):1512–1525, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001070.

Granell:2019:ISC

[GRMH19] Emilio Granell, Verónica Romero, and Carlos-D. Martínez-Hinarejos. Imagespeech combination for interactive computer assisted transcription of handwritten documents. Computer Vision and Image Understanding: CVIU, 180(??): 74–83, March 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219300141.

Gong:1995:DRM

[GS95] Yihong Gong and Masao Sakauchi. Detection of regions matching specified chromatic features. Computer Vision and Image Understanding: CVIU, 61(2): 263–269, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1018/production;   http:

//www.idealibrary.com/
links/artid/cviu.1995.
1018/production/pdf.

Gauch:2006:FIU

[GS06] John M. Gauch and Abhishek Shivadas. Finding and identifying unknown commercials using repeated video sequence detection. Computer Vision and Image Understanding: CVIU, 103 (1):80–88, July 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goshtasby:2008:AWM

[GS08] Ardeshir Goshtasby and Martin Satter. An adaptive window mechanism for image smoothing. Computer Vision and Image Understanding: CVIU, 111(2): 155–169, August 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Garcia-Salguero:2022:CAT

[GSGJ22] Mercedes Garcia-Salguero and Javier Gonzalez-Jimenez. Certifiable algorithms for the two-view planar triangulation problem. Computer Vision and Image Understanding: CVIU, 225(??): ??, December 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001485.

Garcia-Salguero:2024:CPR

[GSGJ24] Mercedes Garcia-Salguero and Javier Gonzalez-Jimenez. Certifiable planar relative pose estimation with gravity prior. Computer Vision and Image Understanding: CVIU, 239(??): ??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002679.

Gupta:2002:CDG

[GSK02] Sumit Gupta, Kuntal Sengupta, and Ashraf A. Kassim. Compression of dynamic 3D geometry data using iterative closest point algorithm. Computer Vision and Image Understanding: CVIU, 87(1–3):116– 130, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Garcia-Sevilla:2001:AIS

[GSP01] Pedro García-Sevilla and Maria Petrou. Analysis of irregularly shaped texture regions. Computer Vision and Image Understanding: CVIU, 84(1):62– 76, October 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0941;   http:

//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0941/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0941/ref.

Ghebreab:2002:NIP

[GSP02] S. Ghebreab, A. W. M.

Smeulders, and P. R. Pfluger. Necklaces: Inhomogeneous and point-enhanced deformable models. Computer Vision and Image Understanding: CVIU, 86(2): 96–117, May 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Geronimo:2010:BBP

[GSPL10] David Gerónimo, Angel D.

Sappa, Daniel Ponsa, and Antonio M. López. 2D–3Dbased on-board pedestrian detection system. Computer Vision and Image Understanding: CVIU, 114(5): 583–595, May 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gavves:2012:VSL

[GSS12] Efstratios Gavves, Cees G. M. Snoek, and Arnold W. M. Smeulders. Visual synonyms for landmark image retrieval. Computer Vision and Image Understanding: CVIU, 116(2):238–249, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211002153.

Goshtasby:2003:NIR

[GSST03] Ardeshir Goshtasby, Lawrence Staib, Colin Studholme, and Demetri Terzopoulos. Nonrigid image registration: guest editors’ introduction. Computer Vision and Image Understanding: CVIU, 89(2–3):109–113, February/ March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goshtasby:2000:AVI

[GSU00] A. Ardeshir Goshtasby, Milan Sonka, and Jayaram Udupa. Analysis of volumetric images. Computer Vision and Image Understanding: CVIU, 77 (2):79–83, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.1999.0819;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        1999.0819/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        1999.0819/ref.

Gracias:2000:UVM

[GSV00] Nuno Gracias and José Santos-Victor. Underwater video mosaics as visual navigation maps. Computer Vision and Image Understanding: CVIU, 79(1): 66–91, July 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0848;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0848/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0848/ref.

Grossmann:2005:LSR

[GSV05] Etienne Grossmann and José Santos-Victor. Leastsquares 3D reconstruction from one or more views and geometric clues. Computer Vision and Image Understanding: CVIU, 99(2):151– 174, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Guidolin:2023:HRO

[GTMR23] Mattia Guidolin, Luca Tagliapietra, Emanuele Menegatti, and Monica Reggiani. Hi-ROS: Open-source multicamera sensor fusion for real-time people tracking. Computer Vision and Image Understanding: CVIU, 232(??):??, July 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223000747.

Giraud:2018:RSU

[GTP18] Rémi Giraud, Vinh-Thong Ta, and Nicolas Papadakis. Robust superpixels using color and contour features along linear path. Computer Vision and Image Understanding: CVIU, 170 (??):1–13, May 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218300067.

Guiducci:1998:RRS

[Gui98]  Antonio Guiducci.  3D

road reconstruction from a single view. Computer Vision and Image Understanding: CVIU, 70(2): 212–226, May 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0633/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0633/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0633/production/ref.

Guiducci:1999:PMP

[Gui99] Antonio Guiducci. Parametric model of the perspective projection of a road with applications to lane keeping and 3D road reconstruction. Computer Vision and Image Understanding: CVIU, 73(3):414– 427, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0737/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0737/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0737/production/ref.

Guiducci:2000:CCR

[Gui00] Antonio Guiducci. Camera calibration for road applications. Computer Vision and Image Understanding: CVIU, 79(2):250– 266, August 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0857;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0857/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0857/ref.

Gruber:2007:INM

[GW07] Amit Gruber and Yair Weiss. Incorporating nonmotion cues into 3D motion segmentation. Computer Vision and Image Understanding: CVIU, 108(3): 261–271, December 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gwak:2017:MOT

[Gwa17] Jeonghwan Gwak. Multiobject tracking through learning relational appearance features and motion patterns. Computer Vision and Image Understanding: CVIU, 162(??):103– 115, September 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300966.

Guttmann:2011:CAV

[GWCO11] Moshe Guttmann, Lior Wolf, and Danny Cohen-Or. Content aware video manipulation. Computer Vision and Image Understanding: CVIU, 115(12):1662– 1678, December 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001597.

Guo:2022:SSB

[GWFF22] Yun-Chih Guo, Tzu-Hsuan Weng, Robin Fischer, and Li-Chen Fu. 3D semantic segmentation based on spatial-aware convolution and shape completion for augmented reality applications. Computer Vision and Image Understanding: CVIU, 224(??): ??, November 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422200128X.

Gryn:2009:DMP

[GWT09] Jacob M. Gryn, Richard P.

Wildes, and John K. Tsotsos. Detecting motion patterns via direction maps with application to surveillance. Computer Vision and Image Understanding: CVIU, 113(2):291–307, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gao:2023:MVC

[GXC23] Chiwei Gao, Ziwei Xu, and Xiuhong Chen. Multiview clustering with Laplacian rank constraint based on symmetric and nonnegative low-rank representation. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002096.

Gao:2019:ICT

[GY19] Yiming Gao and Xiaoping Yang. Infimal convolution type regularization of TGV and shearlet transform for image restoration. Computer Vision and Image Understanding: CVIU, 182 (??):38–49, May 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219300311.

Guo:2021:FFD

[GYCS21] Zhiqing Guo, Gaobo Yang, Jiyou Chen, and Xingming Sun. Fake face detection via adaptive manipulation traces extraction network. Computer Vision and Image Understanding: CVIU, 204(??):Article 103170, March 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422100014X.

Guan:2018:MSR

[GYF18] Banglei Guan, Qifeng Yu, and Friedrich Fraundorfer. Minimal solutions for the rotational alignment of IMU-camera systems using homography constraints. Computer Vision and Image Understanding: CVIU, 170 (??):79–91, May 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300365.

Gao:2009:DOF

[GYTL09] Xinbo Gao, Yimin Yang, Dacheng Tao, and Xuelong Li. Discriminative optical flow tensor for video semantic analysis. Computer Vision and Image Understanding: CVIU, 113(3):372– 383, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Guan:2022:ADL

[GYW^{+}22] Bin Guan, Jinkun Yao, Shaoquan Wang, Guoshan Zhang, Yueming Zhang, Xinbo Wang, and Mengxuan Wang. Automatic detection and localization of thighbone fractures in Xray based on improved deep learning method. Computer Vision and Image Understanding: CVIU, 216(??): ??, February 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001764.

Guo:2023:DAF

[GYWZ23] Zhiqing Guo, Gaobo Yang, Dewang Wang, and Dengyong Zhang. A data augmentation framework by mining structured features for fake face image detection.

Computer Vision and Image Understanding: CVIU, 226 (??):??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001655.

Guo:2019:SDL

[GZ19] Guodong Guo and Na Zhang. A survey on deep learning based face recognition. Computer Vision and Image Understanding: CVIU, 189(??):Article 102805, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219301183.

Gu:2005:TOF

[GZJ05] Haisong Gu, Yongmian Zhang, and Qiang Ji. Task oriented facial behavior recognition with selective sensing. Computer Vision and Image Understanding: CVIU, 100(3):385–415, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gong:2023:SIF

[GZL^{+}23] Jiahao Gong, Sanyuan Zhao, Kin-Man Lam, Xin Gao, and Jianbing Shen. Spectrum-irrelevant finegrained representation for visible-infrared person reidentification. Computer Vision and Image Understanding: CVIU, 232(??): ??, July 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000838.

Gao:2005:IPG

[GZP05] Qigang Gao, Yun Zhang, and Alan Parslow. The influence of perceptual grouping on motion detection. Computer Vision and Image Understanding: CVIU, 100(3):442–457, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ge:2023:GGS

[GZX^{+}23] Yanliang Ge, Qiao Zhang, Tian-Zhu Xiang, Cong Zhang, Jing Zhang, and Hongbo Bi. GSNNet: Group semantic-guided neighbor interaction network for co-salient object detection. Computer Vision and Image Understanding: CVIU, 227 (??):??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001898.

Gou:2023:MNM

[GZY23] Chao Gou, Rui Zhong, and Yuezhao Yu. MAL-Net: Multiscale Attention Link Network for accurate eye center detection. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001303.

Hamker:2005:EAP

[Ham05]  Fred H. Hamker.  The

emergence of attention by population-based inference and its role in distributed processing and cognitive control of vision. Computer Vision and Image Understanding: CVIU, 100(1–2): 64–106, October/November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Horne:2016:SLP

[HAM^{+}16] Lachlan Horne, Jose Alvarez, Chris McCarthy, Mathieu Salzmann, and Nick Barnes. Semantic labeling for prosthetic vision. Computer Vision and Image Understanding: CVIU, 149(??):113– 125, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216000692.

Hansen:2010:FRP

[HASS10] Mark F. Hansen, Gary A.

Atkinson, Lyndon N. Smith, and Melvyn L. Smith. 3D face reconstructions from photometric stereo using near infrared and visible light. Computer Vision and Image Understanding: CVIU, 114(8):942–951, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hanbay:2015:CRI

[HAT^{+}15] Kazim Hanbay, Nuh Alpaslan, Muhammed Fatih Talu, Davut Hanbay, Ali Karci, and Adnan Fatih Kocamaz. Continuous rotation invariant features for gradient-based texture classification. Computer Vision and Image Understanding: CVIU, 132(??):87– 101, March 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002082.

Haala:1998:IUS

[HB98a] Norbert Haala and Claus Brenner. Interpretation of urban surface models using 2D building information. Computer Vision and Image Understanding: CVIU, 72(2):204– 214, November 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0720/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0720/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0720/production/ref.

Heisterkamp:1998:MPA

[HB98b] Douglas R. Heisterkamp and Prabir Bhattacharya. Matching 2D polygonal arcs by using a subgroup of the unit quaternions. Computer Vision and Image Understanding: CVIU, 69(2):246– 249, February 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0566/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0566/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0566/production/ref.

Ho:1998:PCC

[HB98c] Tin Kam Ho and Henry S. Baird. Pattern classification with compact distribution maps. Computer Vision and Image Understanding: CVIU, 70(1):101– 110, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0624/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0624/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0624/production/ref.

Hassan:2012:ASC

[HBB^{+}12] Waqas Hassan, Nagachetan Bangalore, Philip Birch, Rupert Young, and Chris Chatwin. An adaptive sample count particle filter. Computer Vision and Image Understanding: CVIU, 116(12):1208–1222, December 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001191.

Hollingsworth:2009:PDD

[HBF09] Karen Hollingsworth, Kevin W.

Bowyer, and Patrick J. Flynn. Pupil dilation degrades iris biometric performance. Computer Vision and Image Understanding: CVIU, 113(1):150–157, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hosni:2013:SAS

[HBG13] Asmaa Hosni, Michael Bleyer, and Margrit Gelautz. Secrets of adaptive support weight techniques for local stereo matching. Computer Vision and Image Understanding: CVIU, 117(6): 620–632, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000143.

Hesami:2010:RSL

[HBH10] Reyhaneh Hesami, Alireza BabHadiashar, and Reza HosseinNezhad. Range segmentation of large building exteriors: a hierarchical robust approach. Computer Vision and Image Understanding: CVIU, 114(4): 475–490, April 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hoseinnezhad:2011:EHB

[HBH11] Reza Hoseinnezhad and Alireza Bab-Hadiashar. An M-estimator for high breakdown robust estimation in computer vision. Computer Vision and Image Understanding: CVIU, 115(8): 1145–1156, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314211000956.

Hafner:2022:CMD

[HBKG22] Frank M. Hafner, Amran Bhuyian, Julian F. P.

Kooij, and Eric Granger. Cross-modal distillation for RGB-depth person reidentification. Computer Vision and Image Understanding: CVIU, 216(??): ??, February 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001806.

Hollingsworth:2011:GII

[HBL^{+}11] Karen Hollingsworth, Kevin W.

Bowyer, Stephen Lagree, Samuel P. Fenker, and Patrick J. Flynn. Genetically identical irises have texture similarity that is not detected by iris biometrics. Computer Vision and Image Understanding: CVIU, 115(11):1493–1502, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421100155X.

Holliday:2017:SDL

[HBL^{+}17] Andrew Holliday, Mohammadamin Barekatain, Johannes Laurmaa, Chetak Kandaswamy, and Helmut Prendinger. Speedup of deep learning ensembles for semantic segmentation using a model compression technique. Computer Vision and Image Understanding: CVIU, 164(??):16–26, November 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217300826.

Huang:2024:FIC

[HBZ^{+}24] Lan Huang, Xingyu Bai, Jia Zeng, Mengqiang Yu, Wei Pang, and Kangping Wang. FAM: Improving columnar vision transformer with feature attention mechanism. Computer Vision and Image Understanding: CVIU, 242(??):??, May 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000626.

Heidary:2013:PEA

[HC13a] Kaveh Heidary and H. John Caulfield. Presmoothing effects in artificial color image segmentation. Computer Vision and Image Understanding: CVIU, 117(3): 195–201, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001932.

Hu:2013:PEG

[HC13b] Rui Hu and John Collomosse. A performance evaluation of gradient field HOG descriptor for sketch based image retrieval. Computer Vision and Image Understanding: CVIU, 117(7): 790–806, July 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000349.

Hui:2013:DSM

[HC13c] Tak-Wai Hui and Ronald Chung. Determining shape and motion from nonoverlapping multi-camera rig: a direct approach using normal flows. Computer Vision and Image Understanding: CVIU, 117(8): 947–964, August 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000933.

Ho:2016:IPC

[HCC^{+}16] Edmond S. L. Ho, Jacky C. P. Chan, Donald C. K. Chan, Hubert P. H. Shum, Yiu ming Cheung, and Pong C. Yuen. Improving posture classification accuracy for depth sensor-based human activity monitoring in smart environments. Computer Vision and Image Understanding: CVIU, 148 (??):97–110, July 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000138.

Haritaoglu:2001:BDP

[HCHD01] Ismail Haritaoglu, Ross Cutler, David Harwood, and Larry S. Davis. Backpack: Detection of people carrying objects using silhouettes. Computer Vision and Image Understanding: CVIU, 81(3):385–397, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2000.0893;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0893/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0893/ref.

Hong:2021:SCG

[HCLZ21] Chaoqun Hong, Liang Chen, Yuxin Liang, and Zhiqiang Zeng. Stacked Capsule Graph Autoencoders for geometry-aware 3D head pose estimation. Computer Vision and Image Understanding: CVIU, 208–209 (??):??, July 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000680.

Hammoud:2007:AVA

[HD07] Riad Ibrahim Hammoud and James W. Davis. Advances in vision algorithms and systems beyond the visible spectrum. Computer Vision and Image Understanding: CVIU, 106(2–3): 145–147, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Han:2009:PFB

[HD09] Bohyung Han and Larry S.

Davis. Probabilistic fusionbased parameter estimation for visual tracking. Computer Vision and Image Understanding: CVIU, 113(4): 435–445, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hazra:2023:PPU

[HD23] Somnath Hazra and Pallab Dasgupta. Penalizing proposals using classifiers for semi-supervised object detection. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001522.

Huang:2012:PRF

[HDF12] Chen Huang, Xiaoqing Ding, and Chi Fang. Pose robust face tracking by combining view-based AAMs and temporal filters. Computer Vision and Image Understanding: CVIU, 116(7):

777–792, July 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000471.

Herold:2014:RHR

[HDG^{+}14] Catherine Herold, Vincent Despiegel, Stéphane Gentric, Séverine Dubuisson, and Isabelle Bloch. Recursive head reconstruction from multi-view video sequences. Computer Vision and Image Understanding: CVIU, 122(??):182– 201, May 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214000137.

Hoang:2020:SCQ

[HDL^{+}20] Tuan Hoang, Thanh-Toan Do, Huu Le, Dang-Khoa Le-Tan, and Ngai-Man Cheung. Simultaneous compression and quantization: a joint approach for efficient unsupervised hashing. Computer Vision and Image Understanding: CVIU, 191(??):Article 102852, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421930147X.

Hammouche:2008:MAT

[HDS08] Kamal Hammouche, Moussa Diaf, and Patrick Siarry. A multilevel automatic thresholding method based on a genetic algorithm for a fast image segmentation. Computer Vision and Image Understanding: CVIU, 109 (2):163–175, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Haering:1999:FCM

[HdVL99] Niels Haering and Niels da Vitoria Lobo. Features and classification methods to locate deciduous trees in images. Computer Vision and Image Understanding: CVIU, 75(1–2):133– 149, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0769/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0769/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0769/production/ref.

Han:2024:PAS

[HDZR24] Bing Han, Lixiang Deng, Yi Zheng, and Shuang Ren. S^{3}U-PVNet: Arbitraryscale point cloud upsampling via Point-Voxel Network based on Siamese selfsupervised learning. Computer Vision and Image Understanding: CVIU, 239 (??):??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002709.

Heijmans:1999:CMO

[Hei99] Henk J. A. M. Heijmans. Connected morphological operators for binary images. Computer Vision and Image Understanding: CVIU, 73(1):99–120, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0703/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0703/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0703/production/ref.

Heidemann:2004:CSC

[Hei04] Gunther Heidemann. Combining spatial and colour information for content based image retrieval. Computer Vision and Image Understanding: CVIU, 94(1–3): 234–270, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Henricsson:1998:RCA

[Hen98] Olof Henricsson. The role of color attributes and similarity grouping in 3D building reconstruction. Computer Vision and Image Understanding: CVIU, 72(2):163– 184, November 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0718/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0718/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0718/production/ref.

Hansard:2015:CCT

[HEPH15] Miles Hansard, Georgios Evangelidis, Quentin Pelorson, and Radu Horaud. Cross-calibration of timeof-flight and colour cameras. Computer Vision and Image Understanding: CVIU, 134(??):105– 115, May 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001830.

Hilton:2001:MPT

[HF01] Adrian Hilton and Pascal Fua. Modeling people toward vision-based understanding of a person’s shape, appearance, and movement.

Computer Vision and Image Understanding: CVIU, 81(3):227–230, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0907;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0907/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0907/ref.

Hillenbrand:2011:ESF

[HF11] Ulrich Hillenbrand and Alexander Fuchs. An experimental study of four variants of pose clustering from dense range data. Computer Vision and Image Understanding: CVIU, 115(10):1427–1448, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211001445.

Haag:1997:IEM

[HFKN97] Michael Haag, Thomas Frank, Henner Kollnig, and Hans-Hellmut Nagel. Influence of an explicitly modelled 3D scene on the tracking of partially occluded vehicles. Computer Vision and Image Understanding: CVIU, 65(2):206– 225, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0575/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0575/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0575/production/ref.

Hilton:2006:MPV

[HFR06] A. Hilton, P. Fua, and R. Ronfard. Modeling people: Vision-based understanding of a person’s shape, appearance, movement, and behaviour. Computer Vision and Image Understanding: CVIU, 104(2–3):87–89, November/December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hofmann:2011:HMA

[HG11] Michael Hofmann and Dariu M.

Gavrila. 3D Human model adaptation by frame selection and shape-texture optimization. Computer Vision and Image Understanding: CVIU, 115(11):1559– 1570, November 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001846.

Hoover:1998:SER

[HGB98] Adam Hoover, Dmitry Goldgof, and Kevin W. Bowyer. The space envelope: a representation for 3D scenes. Computer Vision and Image Understanding: CVIU, 69(3):310– 329, March 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0666/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0666/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0666/production/ref.

Hsu:2015:HDU

[HGP15] Fu-Chun Hsu, Jayavardhana Gubbi, and Marimuthu Palaniswami. Head detection using motion features and multi level pyramid architecture. Computer Vision and Image Understanding: CVIU, 137(??): 38–49, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000910.

Heber:2013:SBT

[HGR^{+}13] Markus Heber, Martin Godec, Matthias Rüther, Peter M. Roth, and Horst Bischof. Segmentationbased tracking by support fusion. Computer Vision and Image Understanding: CVIU, 117(6): 573–586, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000301.

Hurtut:2008:AIR

[HGS08] Thomas Hurtut, Yann Gousseau, and Francis Schmitt. Adaptive image retrieval based on the spatial organization of colors. Computer Vision and Image Understanding: CVIU, 112(2): 101–113, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hilton:2011:SII

[HGSM11] Adrian Hilton, Guy Godin, Chang Shu, and Takeshi Masuda. Special issue on 3D imaging and modelling. Computer Vision and Image Understanding: CVIU, 115(5):559–560, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Heinke:2005:SAI

[HH05] Dietmar Heinke and Glyn W.

Humphreys. Selective Attention for Identification Model: Simulating visual neglect. Computer Vision and Image Understanding: CVIU, 100(1–2):172–197, October/November 2005.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hansen:2007:ILM

[HH07] Dan Witzner Hansen and Riad I. Hammoud. An improved likelihood model for eye tracking. Computer Vision and Image Understanding: CVIU, 106(2–3):220– 230, May/June 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Huo:2012:MPT

[HH12] Feifei Huo and Emile A. Hendriks. Multiple people tracking and pose estimation with occlusion estimation. Computer Vision and Image Understanding: CVIU, 116(5):634– 647, May 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314212000033.

Hu:2019:DDF

[HH19] Weipeng Hu and Haifeng Hu. Discriminant deep feature learning based on joint supervision loss and multilayer feature fusion for heterogeneous face recognition. Computer Vision and Image Understanding: CVIU, 184 (??):9–21, July 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1077314219300566.

Hansard:2014:ADC

[HHAE14] Miles Hansard, Radu Horaud, Michel Amat, and Georgios Evangelidis. Automatic detection of calibration grids in time-of-flight images. Computer Vision and Image Understanding: CVIU, 121(??): 108–118, April 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000149.

Hassanein:2020:IMP

[HHG^{+}20] Allam S. Hassanein, Mohamed E. Hussein, Walid Gomaa, Yasushi Makihara, and Yasushi Yagi. Identifying motion pathways in highly crowded scenes: a non-parametric tracklet clustering approach. Computer Vision and Image Understanding: CVIU, 191 (??):Article 102710, February 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314218301887.

Hong:2016:EEI

[HHM^{+}16] Xin Hong, Yan Huang, Wenjun Ma, Sriram Varadarajan, Paul Miller, Weiru Liu, Maria Jose Santofimia Romero, Jesus Martinez del Rincon, and Huiyu Zhou. Evidential event inference in transport video surveillance. Computer Vision and Image Understanding: CVIU, 144(??):276– 297, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002477.

Heisele:2003:FRC

[HHWP03] Bernd Heisele, Purdy Ho, Jane Wu, and Tomaso Poggio. Face recognition: component-based versus global approaches. Computer Vision and Image Understanding: CVIU, 91(1– 2):6–21, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Harrison:2012:TPA

[HJ12] Adam P. Harrison and Dileepan Joseph. Translational photometric alignment of single-view image sequences. Computer Vision and Image Understanding: CVIU, 116(6): 765–776, June 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000227.

Huang:2016:MOT

[HJZ16] Shucheng Huang, Shuai Jiang, and Xia Zhu. Multiobject tracking via discriminative appearance modeling. Computer Vision and Image Understanding: CVIU, 153(??):77–87, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216300790.

Huq:2013:OFS

[HKA13] Shafik Huq, Andreas Koschan, and Mongi Abidi. Occlusion filling in stereo: Theory and experiments. Computer Vision and Image Understanding: CVIU, 117(6): 688–704, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000155.

Hamid:2014:VFT

[HKHE14] Raffay Hamid, Ramkrishan Kumar, Jessica Hodgins, and Irfan Essa. A visualization framework for team sports captured using multiple static cameras. Computer Vision and Image Understanding: CVIU, 118(??):171– 183, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213001768.

Hwang:2008:CDU

[HKK08] Youngbae Hwang, Jun-Sik Kim, and In-So Kweon. Change detection using a statistical model in an optimally selected color space. Computer Vision and Image Understanding: CVIU, 112(3):231–242, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Huo:2024:TEI

[HKL^{+}24] Jing Huo, Meihao Kong, Wenbin Li, Jing Wu, Yu-Kun Lai, and Yang Gao. Towards efficient image and video style transfer via distillation and learnable feature transformation. Computer Vision and Image Understanding: CVIU, 241 (??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000286.

Han:2022:QBD

[HKM22] Juan Han, Kit Ian Kou, and Jifei Miao. Quaternionbased dynamic mode decomposition for background modeling in color videos. Computer Vision and Image Understanding: CVIU, 224(??):??, November 2022.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314222001382.

Hasinoff:2006:BMV

[HKS06] Samuel W. Hasinoff, Sing Bing Kang, and Richard Szeliski. Boundary matting for view synthesis. Computer Vision and Image Understanding: CVIU, 103(1):22–32, July 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Huang:2014:NDR

[HKWC14] Zeyi Huang, Wenxiong Kang, Qiuxia Wu, and Xiaopeng Chen. A new descriptor resistant to affine transformation and monotonic intensity change. Computer Vision and Image Understanding: CVIU, 120 (??):117–125, March 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314213002014.

Huang:2016:MME

[HKZ^{+}16] Xiaohua Huang, Jukka Kortelainen, Guoying Zhao, Xiaobai Li, Antti Moilanen, Tapio Seppänen, and Matti Pietikäinen. Multi-modal emotion analysis from facial expressions and electroencephalogram. Computer Vision and Image Understanding: CVIU, 147(??): 114–124, June 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002106.

Hjelmaas:2001:FDS

[HL01] Erik Hjelmås and Boon Kee Low. Face detection: a survey. Computer Vision and Image Understanding: CVIU, 83(3): 236–274, September 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0921;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0921/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0921/ref.

Han:2013:BIT

[HL13] Yina Han and Guizhong Liu. Biologically inspired task oriented gist model for scene classification. Computer Vision and Image Understanding: CVIU, 117(1): 76–95, January 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421200135X.

Hussain:2023:EFM

[HL23] Sadia Hussain and Brejesh Lall. Extending function mixture network for improved spectral superresolution. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731422300214X.

Hadfield:2017:SRU

[HLB17] Simon Hadfield, Karel Lebeda, and Richard Bowden. Stereo reconstruction using top-down cues. Computer Vision and Image Understanding: CVIU, 157 (??):206–222, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314216301096.

Hsieh:1997:IRU

[HLF^{+}97] Jun-Wei Hsieh, Hong-Yuan Mark Liao, Kuo-Chin Fan, Ming-Tat Ko, and Yi-Ping Hung. Image registration using a new edgebased approach. Computer Vision and Image Understanding: CVIU, 67(2):112– 130, August 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0517/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0517/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0517/production/ref.

Hwang:2019:PML

[HLKK19] Youngbae Hwang, Joon-Young Lee, In So Kweon, and Seon Joo Kim. Probabilistic moving least squares with spatial constraints for nonlinear color transfer between images. Computer Vision and Image Understanding: CVIU, 180(??): 1–12, March 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218304491.

He:2023:DAM

[HLL^{+}23] Xuan He, Zhiyong Li, Jiacheng Lin, Ke Nai, Jin Yuan, Yifan Li, and Runmin Wang. Domain adaptive multigranularity proposal network for text detection under extreme traffic scenes. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000899.

Hartley:1997:RPR

[HM97] Richard I. Hartley and Roger Mohr. Reply to Pizlo, Rosenfeld, and Weiss. Computer Vision and Image Understanding: CVIU, 67(3):320– 323, September 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0644/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0644/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0644/production/ref. See

[PRW97a].

Huan:2010:IRB

[HMA10] Xiaoli Huan, Beddhu Murali, and Adel L. Ali. Image restoration based on the fast marching method and block based sampling. Computer Vision and Image Understanding: CVIU, 114 (8):847–856, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hofer:2017:ESA

[HMB17] Manuel Hofer, Michael Maurer, and Horst Bischof. Efficient 3D scene abstraction using line segments. Computer Vision and Image Understanding: CVIU, 157(??):167– 178, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216300236.

Hanbury:2010:SII

[HMC10] Allan Hanbury, Henning Müller, and Paul Clough. Special issue on image and video retrieval evaluation. Computer Vision and Image Understanding: CVIU, 114(4):409–410, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hemati:2022:NAG

[HMCT22] Sobhan Hemati, Mohammad Hadi Mehdizavareh, Shojaeddin Chenouri, and Hamid R. Tizhoosh. A non-alternating graph hashing algorithm for large-scale image search. Computer Vision and Image Understanding: CVIU, 219(??): ??, June 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314222000418.

Hinz:2007:TMS

[HMEB07] Stefan Hinz, Franz Meyer, Michael Eineder, and Richard Bamler. Traffic monitoring with spaceborne SAR-Theory, simulations, and experiments. Computer Vision and Image Understanding: CVIU, 106(2–3):231– 244, May/June 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Holte:2010:VIG

[HMF10] M. B. Holte, T. B. Moeslund, and P. Fihl. Viewinvariant gesture recognition using 3D optical flow and harmonic motion context. Computer Vision and Image Understanding: CVIU, 114(12):1353–1361, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hu:2015:DTD

[HML15] Tao Hu, Stefano Messelodi, and Oswald Lanz. Dynamic task decomposition for decentralized object tracking in complex scenes. Computer Vision and Image Understanding: CVIU, 134 (??):89–104, May 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000417.

Holt:1995:UAI

[HN95] Robert J. Holt and Arun N.

Netravali. Using affine invariants on perspective projections of plane curves. Computer Vision and Image Understanding: CVIU, 61 (1):112–121, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1995.
                                        1008/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1008/production/pdf.

Hongeng:2004:VBE

[HNB04] Somboon Hongeng, Ram Nevatia, and François Bremond. Video-based event recognition: activity representation and probabilistic recognition methods. Computer Vision and Image Understanding: CVIU, 96(2): 129–162, November 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hurley:2005:FFF

[HNC05] David J. Hurley, Mark S. Nixon, and John N. Carter. Force field feature extraction for ear biometrics. Computer Vision and Image Understanding: CVIU, 98(3):491–512, June 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hobby:2000:USL

[Hob00] John D. Hobby. Using shape and layout information to find signatures, text, and graphics. Computer Vision and Image Understanding: CVIU, 80 (1):88–110, October 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0868;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0868/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0868/ref.

Hodges:1995:FPO

[Hod95]  Jack Hodges.  Func-

tional and physical object characteristics and object recognition in improvisation. Computer Vision and Image Understanding: CVIU, 62(2):147– 163, September 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1046/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1046/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1995.
1047/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1047/production/pdf.

Hernandez:2007:VLT

[HOH^{+}07] Benjamín Hernández, Gustavo Olague, Riad Hammoud, Leonardo Trujillo, and Eva Romero. Visual learning of texture descriptors for facial expression recognition in thermal imagery. Computer Vision and Image Understanding: CVIU, 106(2–3):258–269, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hu:1996:HAE

[HP96] Jianying Hu and Theo Pavlidis. A hierarchical approach to efficient curvilinear object searching. Computer Vision and Image Understanding: CVIU, 63(2): 208–220, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0015/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0015/production/pdf.

Hansen:2005:ETW

[HP05] Dan Witzner Hansen and Arthur E. C. Pece. Eye tracking in the wild. Computer Vision and Image Understanding: CVIU, 98(1): 155–181, April 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hoey:2010:AHA

[HPvB^{+}10] Jesse Hoey, Pascal Poupart, Axel von Bertoldi, Tammy Craig, Craig Boutilier, and Alex Mihailidis. Automated handwashing assistance for persons with dementia using video and a partially observable Markov decision process. Computer Vision and Image Understanding: CVIU, 114(5):503– 519, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hu:2005:FCC

[HQN05] Qingmao Hu, Guoyu Qian, and Wiesław L. Nowinski. Fast connectedcomponent labelling in three-dimensional binary images based on iterative recursion. Computer Vision and Image Understanding: CVIU, 99(3):414–434, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hu:2012:SMU

[HQW^{+}12] Tingbo Hu, Baojun Qi, Tao Wu, Xin Xu, and Hangen He. Stereo matching using weighted dynamic programming on a single-direction four-connected tree. Computer Vision and Image Understanding: CVIU, 116(8): 908–921, August 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212000641.

Hausler:1999:FBO

[HR99] G. Häusler and D. Ritter. Feature-based object recognition and localization in 3D-space, using a single video image. Computer Vision and Image Understanding: CVIU, 73(1):64– 81, January 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0704/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0704/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0704/production/ref.

Hu:2009:AMF

[HRC09] Yiqun Hu, Deepu Rajan, and Liang-Tien Chia. Attention-from-motion: a factorization approach for detecting attention objects in motion. Computer Vision and Image Understanding: CVIU, 113(3):319– 331, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hasan:2016:ILH

[HRC16] Mahmudul Hasan and Amit K.

Roy-Chowdhury. Incremental learning of human activity models from videos. Computer Vision and Image Understanding: CVIU, 144(??):24– 35, March 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002489.

Han:2017:STR

[HRHZ17] Fei Han, Brian Reily, William Hoff, and Hao Zhang. Space-time representation of people based on 3D skeletal data: a review. Computer Vision and Image Understanding: CVIU, 158 (??):85–105, May 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300279.

Hartkens:2002:EOD

[HRS02] Thomas Hartkens, Karl Rohr, and H. Siegfried Stiehl. Evaluation of 3D operators for the detection of anatomical point landmarks in MR and CT images. Computer Vision and Image Understanding: CVIU, 86(2):118–136, May 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hartley:1997:T

[HS97] Richard I. Hartley and Peter Sturm. Triangulation. Computer Vision and Image Understanding: CVIU, 68(2):146– 157, November 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0547/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0547/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0547/production/ref.

Heo:2006:FRM

[HS06] Giseon Heo and Christopher G. Small. Form representations and means for landmarks: a survey and comparative study. Computer Vision and Image Understanding: CVIU, 102(2): 188–203, May 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Habibian:2014:RRV

[HS14] Amirhossein Habibian and Cees G. M. Snoek. Recommendations for recognizing video events by concept vocabularies. Computer Vision and Image Understanding: CVIU, 124(??): 110–122, July 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000290.

He:2016:HMD

[HSBS16] Sheng He, Petros Samara, Jan Burgers, and Lambert Schomaker. Historical manuscript dating based on temporal pattern codebook. Computer Vision and Image Understanding: CVIU, 152(??):167– 175, November 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301163.

Hannuksela:2007:VBM

[HSH07] Jari Hannuksela, Pekka Sangi, and Janne Heikkilä. Vision-based motion estimation for interaction with mobile devices. Computer Vision and Image Understanding: CVIU, 108 (1–2):188–195, October/ November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Huang:2020:HSM

[HSHA20] Ying Huang, Hubert P. H.

Shum, Edmond S. L. Ho, and Nauman Aslam. Highspeed multi-person pose estimation with deep feature transfer. Computer Vision and Image Understanding:

CVIU, 197–198(??):Article 103010, August 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218302261.

Hilton:1998:ISB

[HSIW98] A. Hilton, A. J. Stoddart, J. Illingworth, and T. Windeatt. Implicit surface-based geometric fusion. Computer Vision and Image Understanding: CVIU, 69(3):273– 291, March 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0664/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0664/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0664/production/ref.

Huhle:2010:FRC

[HSJS10] Benjamin Huhle, Timo Schairer, Philipp Jenke, and Wolfgang Straßer. Fusion of range and color images for denoising and resolution enhancement with a non-local filter. Computer Vision and Image Understanding: CVIU, 114(12):1336–1345, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Heo:2023:EDA

[HSK23] Jaehyuk Heo, Seungwan Seo, and Pilsung Kang. Exploring the differences in adversarial robustness between ViT- and CNN-based models using novel metrics. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001807.

Helferty:2007:CBS

[HSKH07] J. P. Helferty, A. J. Sherbondy, A. P. Kiraly, and W. E. Higgins. Computerbased system for the virtualendoscopic guidance of bronchoscopy. Computer Vision and Image Understanding: CVIU, 108 (1–2):171–187, October/ November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hershkovich:2016:PMI

[HSS^{+}16] Tsachi Hershkovich, Tamar Shalmon, Ohad Shitrit, Nir Halay, Bjoern H. Menze, Irit Dolgopyat, Itamar Kahn, Ilan Shelef, and Tammy Riklin Raviv. Probabilistic model for 3D interactive segmentation. Computer Vision and Image Understanding: CVIU, 151(??):47–60, October 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216300017.

Heath:1998:CED

[HSSB98] M. Heath, S. Sarkar, T. Sanocki, and K. Bowyer. Comparison of edge detectors. A methodology and initial study. Computer Vision and Image Understanding: CVIU, 69(1):38– ??, ???? 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hu:2024:ANB

[HSTL24] Chengyin Hu, Weiwen Shi, Ling Tian, and Wen Li. Adversarial Neon Beam: a light-based physical attack to DNNs. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223002576.

Hager:1998:XVP

[HT98] Gregory D. Hager and Kentaro Toyama. X vision: a portable substrate for real-time vision applications. Computer Vision and Image Understanding: CVIU, 69(1):023– 037, January 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0561/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0561/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0561/production/ref;
http://www.idealibrary.
com/links/artid/cviu.1997.
0586/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0586/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0586/production/ref;
http://www.idealibrary.
com/links/artid/cviu.1997.
0587/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0587/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0587/production/ref.

Hancock:2011:SIG

[HTEB11] Edwin R. Hancock, Andrea Torsello, Francisco Escolano, and Luc Brun. Special issue on graph-based representations in computer vision. Computer Vision and Image Understanding: CVIU, 115(7):903– 904, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211001093.

HajiRassouliha:2018:SPB

[HTNN18] Amir HajiRassouliha, Andrew J. Taberner, Martyn P. Nash, and Poul M. F. Nielsen. Subpixel phase-based image registration using Savitzky–Golay differentiators in gradientcorrelation. Computer Vision and Image Understanding: CVIU, 170(??):28– 39, May 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421730187X.

Hu:2008:IBN

[Hu08] Haifeng Hu. ICA-based neighborhood preserving analysis for face recognition. Computer Vision and Image Understanding: CVIU, 112(3):286–295, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hu:2011:MIN

[Hu11] Haifeng Hu. Multiscale illumination normalization for face recognition using dual-tree complex wavelet transform in logarithm domain. Computer Vision and Image Understanding: CVIU, 115(10):1384–1394, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421100141X.

Herda:2005:HIS

[HUF05] L. Herda, R. Urtasun, and P. Fua. Hierarchical implicit surface joint limits for human body tracking. Computer Vision and Image Understanding: CVIU, 99(2):189–209, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Haltakov:2016:GPN

[HUI16] Vladimir Haltakov, Christian Unger, and Slobodan Ilic. Geodesic pixel neighborhoods for 2D and 3D scene understanding. Computer Vision and Image Understanding: CVIU, 148 (??):164–180, July 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314215002520.

Hua:2006:SMF

[HW06] Gang Hua and Ying Wu.

Sequential mean field variational analysis of structured deformable shapes. Computer Vision and Image Understanding: CVIU, 101(2): 87–99, February 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hua:2007:DPA

[HW07] Gang Hua and Ying Wu. A decentralized probabilistic approach to articulated body tracking. Computer Vision and Image Understanding: CVIU, 108(3): 272–283, December 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hao:2021:WSI

[HWG21] Shengyu Hao, Gaoang Wang, and Renshu Gu. Weakly supervised instance segmentation using multiprior fusion. Computer Vision and Image Understanding: CVIU, 211(??): ??, October 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221001053.

Han:2021:PIA

[HWK^{+}21] Jian Han, Wei Wang, Sezer Karaoglu, Wei Zeng, and Theo Gevers. Pose invariant age estimation of face images in the wild. Computer Vision and Image Understanding: CVIU, 202 (??):Article 103123, January 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220301430.

Huang:2022:RRW

[HWL^{+}22] Xiaoshui Huang, Yangfu Wang, Sheng Li, Guofeng Mei, Zongyi Xu, Yucheng Wang, Jian Zhang, and Mohammed Bennamoun. Robust real-world point cloud registration by inlier detection. Computer Vision and Image Understanding: CVIU, 224(??): ??, November 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001345.

Huo:2024:GCR

[HWT^{+}24] Wei Huo, Ke Wang, Jun Tang, Nian Wang, and Dong Liang. GaitSCM: Causal representation learning for gait recognition. Computer Vision and Image Understanding: CVIU, 243(??): ??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000766.

Huang:2006:NCI

[HWW06] Kai-Qi Huang, Qiao Wang, and Zhen-Yang Wu. Natural color image enhancement and evaluation algorithm based on human visual system. Computer Vision and Image Understanding: CVIU, 103(1):52– 63, July 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Huang:2023:WSS

[HWW23] Zuxian Huang, Gangshan Wu, and Limin Wang. Webly-supervised semantic segmentation via curriculum learning. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422300190X.

He:2016:ISU

[HWZ16] Kun He, Dan Wang, and Xu Zhang. Image segmentation using the level set and improved-variation smoothing. Computer Vision and Image Understanding: CVIU, 152(??):29–40, November 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216300960.

Hao:2023:EMS

[HWZ^{+}23] Zeyu Hao, Hang Wang, Xuchong Zhang, Yuhai Li, and Hongbin Sun. Efficient multi-stage network with pixel-wise degradation prediction for real-time motion deblurring. Computer Vision and Image Understanding: CVIU, 233(??): ??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000735.

Hu:1998:MBS

[HY98] Jianming Hu and Hong Yan.

A model-based segmentation method for handwritten numeral strings. Computer Vision and Image Understanding: CVIU, 70(3): 383–403, June 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0689/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0689/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0689/production/ref.

Ho:2011:ARP

[HY11] Jeffrey Ho and Ming-Hsuan Yang. On affine registration of planar point sets using complex numbers. Computer Vision and Image Understanding: CVIU, 115(1): 50–58, January 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Han:2011:CFE

[HYJ11] Zhenjun Han, Qixiang Ye, and Jianbin Jiao. Combined feature evaluation for adaptive visual object tracking. Computer Vision and Image Understanding: CVIU, 115(1):69–80, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Huang:2024:FUI

[HYW^{+}24] Zefeng Huang, Shen Yang, Jin Wu, Lei Zhu, and Jin Liu. FusionDiff: a unified image fusion network based on diffusion probabilistic models. Computer Vision and Image Understanding: CVIU, 244(??): ??, July 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314224000924.

Hu:2019:EWS

[HZK19] Shell Xu Hu, Sergey Zagoruyko, and Nikos Komodakis. Exploring weight symmetry in deep neural networks. Computer Vision and Image Understanding: CVIU, 187(??):Article 102786, October 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S107731421930102X.

Huang:2011:CSA

[HZLM11] Junzhou Huang, Shaoting Zhang, Hongsheng Li, and Dimitris Metaxas. Composite splitting algorithms for convex optimization. Computer Vision and Image Understanding: CVIU, 115(12):1610–1622, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211001585.

Humenberger:2010:FSM

[HZW^{+}10] Martin Humenberger, Christian Zinner, Michael Weber, Wilfried Kubinger, and Markus Vincze. A fast stereo matching algorithm suitable for embedded realtime systems. Computer Vision and Image Understanding: CVIU, 114(11): 1180–1202, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ion:2011:MAS

[IAP^{+}11] Adrian Ion, Nicole M. Artner, Gabriel Peyré, Walter G. Kropatsch, and Laurent D. Cohen. Matching 2D and 3D articulated shapes using the eccentricity transform. Computer Vision and Image Understanding: CVIU, 115(6):817– 834, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Intille:2001:RPM

[IB01] Stephen S. Intille and Aaron F. Bobick. Recognizing planned, multiperson action. Computer Vision and Image Understanding: CVIU, 81(3):414– 445, March 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0896;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0896/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0896/ref.

Iqbal:2018:DSA

[IDY^{+}18] Umar Iqbal, Andreas Doering, Hashim Yasin, Björn Krüger, Andreas Weber, and Juergen Gall. A dualsource approach for 3D human pose estimation from single images. Computer Vision and Image Understanding: CVIU, 172(??): 37–49, July 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300511.

Imiya:1999:ECD

[IE99] Atsushi Imiya and Ulrich Eckhardt. The Euler characteristics of discrete objects and discrete quasiobjects. Computer Vision and Image Understanding: CVIU, 75(3):307– 318, September 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0791/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0791/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0791/production/ref.

Ikeuchi:1995:RPC

[IF95] Katsushi Ikeuchi and Patrick J.

Flynn. Recent progress in CAD-based vision. Computer Vision and Image Understanding: CVIU, 61(3): 293–294, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1023/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1023/production/pdf.

Imiya:1999:MAR

[IF99] Atsushi Imiya and Iris Fermin. Motion analysis by random sampling and voting process. Computer Vision and Image Understanding: CVIU, 73(3):309– 328, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0734/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0734/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0734/production/ref.

Imre:2015:CEM

[IH15] Evren Imre and Adrian Hilton. Covariance estimation for minimal geometry solvers via scaled unscented transformation. Computer Vision and Image Understanding: CVIU, 130(??): 18–34, January 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214001817.

Israel-Jost:2013:CSS

[IJDAB13] Vincent Israel-Jost, Jérôme Darbon, Elsa D. Angelini, and Isabelle Bloch. Conciliating syntactic and semantic constraints for multiphase and multi-channel region segmentation. Computer Vision and Image Understanding: CVIU, 117(8): 819–826, August 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000556.

Isler:2005:TTD

[IKST05] Volkan Isler, Sanjeev Khanna, John Spletzer, and Camillo J. Taylor. Target tracking with distributed sensors: The focus of attention problem. Computer Vision and Image Understanding: CVIU, 100(1–2):225–247, October/ November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Idrissi:2004:OIB

[ILRB04] Khalid Idrissi, Guillaume Lavoué, Julien Ricard, and Atilla Baskurt. Object of interest-based visual navigation, retrieval, and semantic content identification system. Computer Vision and Image Understanding: CVIU, 94(1–3):271– 294, April/June 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Immerkaer:1996:FNV

[Imm96] John Immerkær. Fast noise variance estimation. Computer Vision and Image Understanding: CVIU, 64(2): 300–302, September 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0060/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0060/production/pdf.

Ivins:1998:CAR

[IP98] Jim Ivins and John Porrill. Constrained active region models for fast tracking in color image sequences. Computer Vision and Image Understanding: CVIU, 72(1):54– 71, October 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0653/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0653/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0653/production/ref.

Impoco:2015:ILS

[IT15] G. Impoco and L. Tuminello. Incremental learning to segment micrographs. Computer Vision and Image Understanding: CVIU, 140(??):144– 152, November 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000533.

Iosifidis:2012:MVH

[ITNP12] Alexandros Iosifidis, Anastasios Tefas, Nikolaos Nikolaidis, and Ioannis Pitas. Multi-view human movement recognition based on fuzzy distances and linear discriminant analysis. Computer Vision and Image Understanding: CVIU, 116(3): 347–360, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002074.

Ip:1997:DPP

[IW97] Horace H. S. Ip and W. H.

Wong. Detecting perceptually parallel curves: Criteria and force-driven optimization. Computer Vision and Image Understanding: CVIU, 68(2):190– 208, November 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0552/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0552/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0552/production/ref.

Idrees:2017:TCA

[IZJ^{+}17] Haroon Idrees, Amir R. Zamir, Yu-Gang Jiang, Alex Gorban, Ivan Laptev, Rahul Sukthankar, and Mubarak Shah. The THUMOS challenge on action recognition for videos “in the wild”. Computer Vision and Image Understanding: CVIU, 155(??):1–23, February 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216301710.

Irschara:2012:LSD

[IZKB12] Arnold Irschara, Christopher Zach, Manfred Klopschitz, and Horst Bischof. Large-scale, dense city reconstruction from usercontributed photos. Computer Vision and Image Understanding: CVIU, 116(1): 2–15, January 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001810.

Jevnisek:2016:SGB

[JA16] Roy Josef Jevnisek and Shai Avidan. Semi global boundary detection. Computer Vision and Image Understanding: CVIU, 152(??): 21–28, November 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421630100X.

Jacobs:2001:LFM

[Jac01] David W. Jacobs. Linear fitting with missing data for structure-from-motion. Computer Vision and Image Understanding: CVIU, 82(1):57–81, April 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0906;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0906/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0906/ref.

Jiang:1999:EDR

[JB99] Xiaoyi Jiang and Horst Bunke. Edge detection in range images based on scan line approximation. Computer Vision and Image Understanding: CVIU, 73(2):183–199, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0715/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0715/production/pdf;
http://www.idealibrary.

                                        com/links/artid/cviu.1998.
                                        0715/production/ref.

Jin:2015:ASP

[JB15] Zhixing Jin and Bir Bhanu. Analysis-by-synthesis: Pedestrian tracking with crowd simulation models in a multi-camera video network. Computer Vision and Image Understanding: CVIU, 134(??):48– 63, May 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001933.

J:2023:ALH

[JB23] Preethi S. J. and Niranjana Krupa B. Analyzing lower half facial gestures for lip reading applications: Survey on vision techniques. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001182.

Jia:2008:VBD

[JBC08] Zhen Jia, Arjuna Balasuriya, and Subhash Challa. Vision based data fusion for autonomous vehicles target tracking using interacting multiple dynamic models. Computer Vision and Image Understanding: CVIU, 109(1):1–21, January 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jeyakar:2008:ROT

[JBR08] Jaideep Jeyakar, R. Venkatesh Babu, and K. R. Ramakrishnan. Robust object tracking with backgroundweighted local kernels. Computer Vision and Image Understanding: CVIU, 112(3):296–309, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jiang:2011:GBM

[JBWK11] Xiaoyi Jiang, Klaus Broelemann, Steffen Wachenfeld, and Antonio Krüger. Graph-based markerless registration of city maps using geometric hashing. Computer Vision and Image Understanding: CVIU, 115 (7):1032–1043, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314211000828.

Jang:1998:MSS

[JC98] Ben K. Jang and Roland T.

Chin. Morphological scale space for 2D shape smoothing. Computer Vision and Image Understanding: CVIU, 70(2):121– 141, May 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0626/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0626/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0626/production/ref.

Joshi:2006:SES

[JC06] Manjunath V. Joshi and Subhasis Chaudhuri. Simultaneous estimation of superresolved depth map and intensity field using photometric cue. Computer Vision and Image Understanding: CVIU, 101(1):31–44, January 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jiang:2021:SIR

[JCLZ21] Nanfeng Jiang, Weiling Chen, Liqun Lin, and Tiesong Zhao. Single image rain removal via multimodule deep grid network. Computer Vision and Image Understanding: CVIU, 202(??):Article 103106, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301302.

Jacot-Descombes:1997:APG

[JDP97] Alain Jacot-Descombes and Thierry Pun. Asynchronous perceptual grouping: From contours to relevant 2-D structures. Computer Vision and Image Understanding: CVIU, 66(1): 1–24, April 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0509/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0509/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0509/production/ref.

Jean:2011:OSP

[Jea11]  Yves D. Jean.  Opti-

cal signal processing with illumination-encoded filters. Computer Vision and Image Understanding: CVIU, 115(5):561–575, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jamieson:2012:DHO

[JEF^{+}12] Michael Jamieson, Yulia Eskin, Afsaneh Fazly, Suzanne Stevenson, and Sven J. Dickinson. Discovering hierarchical object models from captioned images. Computer Vision and Image Understanding: CVIU, 116(7):

842–853, July 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000495.

Jordan:1998:PCB

[JEK98] Corinne Le Buhan Jordan, Touradj Ebrahimi, and Murat Kunt. Progressive content-based shape compression for retrieval of binary images. Computer Vision and Image Understanding: CVIU, 71(2):198– 212, August 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0707/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0707/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0707/production/ref.

Junejo:2010:GCE

[JF10] Imran N. Junejo and Hassan Foroosh. GPS coordinates estimation and camera calibration from solar shadows. Computer Vision and Image Understanding: CVIU, 114(9):991–1003, September 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jiang:2020:VBE

[JGM20] Min Jiang, Guodong Guo, and Guowang Mu. Visual BMI estimation from face images using a label distribution based method. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 102985, August 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731422030059X.

Jang:2019:RFF

[JGP19] Youngkyoon Jang, Hatice Gunes, and Ioannis Patras. Registration-free Face-SSD: Single shot analysis of smiles, facial attributes, and affect in the wild. Computer Vision and Image Understanding: CVIU, 182(??): 17–29, May 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219300128.

Ju:2016:E

[JGSP16] Jingyi Ju, Bastian Goldluecke, Richard Szeliski, and Tomás Pajdla. Editorial. Computer Vision and Image Understanding: CVIU, 145 (??):139, April 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216000552.

Javaran:2017:NBI

[JHA17] Taiebeh Askari Javaran, Hamid Hassanpour, and Vahid Abolghasemi. Nonblind image deconvolution using a regularization based on re-blurring process. Computer Vision and Image Understanding: CVIU, 154(??):16–34, January 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216301473.

Jung:2024:RSU

[JHK24] Min Jae Jung, Seung Dae Han, and Joohee Kim. Re-scoring using imagelanguage similarity for fewshot object detection. Computer Vision and Image Understanding: CVIU, 241 (??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000377.

Jorquera:2019:PHD

[JHV19] Felipe Jorquera, Sergio Hernández, and Diego Vergara. Probability hypothesis density filter using determinantal point processes for multi object tracking. Computer Vision and Image Understanding: CVIU, 183 (??):33–41, June 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219300529.

Jain:2007:SMO

[JKM07] Vishal Jain, Benjamin B.

Kimia, and Joseph L. Mundy. Segregation of moving objects using elastic matching. Computer Vision and Image Understanding: CVIU, 108(3):230–242, December 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jam:2021:CRP

[JKW^{+}21] Jireh Jam, Connah Kendrick, Kevin Walker, Vincent Drouard, Jison Gee-Sern Hsu, and Moi Hoon Yap. A comprehensive review of past and present image inpainting methods. Computer Vision and Image Understanding: CVIU, 203 (??):Article 103147, February 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220301661.

Jiang:2012:CCM

[JLD12] Zhuolin Jiang, Zhe Lin, and Larry S. Davis. Class consistent k-means: Application to face and action recognition. Computer Vision and Image Understanding: CVIU, 116(6): 730–741, June 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212000367.

Jiang:2013:UTB

[JLD13] Zhuolin Jiang, Zhe Lin, and Larry S. Davis. A unified tree-based framework for joint action localization, recognition and segmentation. Computer Vision and Image Understanding: CVIU, 117(10):1345– 1355, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001749.

Jung:2013:WAP

[JLL13] Ho Yub Jung, Kyoung Mu Lee, and Sang Uk Lee. Window annealing for pixellabeling problems. Computer Vision and Image Understanding: CVIU, 117(3): 289–303, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212002056.

Jodelet:2022:BSC

[JLM22] Quentin Jodelet, Xin Liu, and Tsuyoshi Murata. Balanced softmax cross-entropy for incremental learning with and without memory. Computer Vision and Image Understanding: CVIU, 225(??):??, December 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222001606.

Jampour:2017:FIB

[JLY^{+}17] Mahdi Jampour, Chen Li, Lap-Fai Yu, Kun Zhou, Stephen Lin, and Horst Bischof. Face inpainting based on high-level facial attributes. Computer Vision and Image Understanding: CVIU, 161(??):29– 41, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300930.

Ji:2023:SSS

[JLZ23] Ruyi Ji, Jiaying Li, and Libo Zhang. Siamese selfsupervised learning for finegrained visual classification. Computer Vision and Image Understanding: CVIU, 229 (??):??, March 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000383.

Jodoin:2009:OFB

[JM09a] Pierre-Marc Jodoin and Max Mignotte. Optical-flow based on an edge-avoidance procedure. Computer Vision and Image Understanding: CVIU, 113(4):511– 531, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Just:2009:CST

[JM09b] Agnès Just and Sébastien Marcel. A comparative study of two state-of-the-art sequence processing techniques for hand gesture recognition. Computer Vision and Image Understanding: CVIU, 113(4):532– 543, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Julia:2011:SBI

[JMPG11] Carme Julià, Rodrigo Moreno, Domenec Puig, and Miguel Angel Garcia. Shape-based image segmentation through photometric stereo. Computer Vision and Image Understanding: CVIU, 115(1): 91–104, January 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jiang:2009:VWP

[JN09] Yu-Gang Jiang and Chong-Wah Ngo. Visual word proximity and linguistics for semantic video indexing and near-duplicate retrieval. Computer Vision and Image Understanding: CVIU, 113 (3):405–414, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jampani:2015:ISD

[JNLG15] Varun Jampani, Sebastian Nowozin, Matthew Loper, and Peter V. Gehler. The informed sampler: a discriminative approach to Bayesian inference in generative computer vision models. Computer Vision and Image Understanding: CVIU, 136 (??):32–44, July 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421500048X.

Jain:2011:GQ

[JO11] Brijnesh J. Jain and Klaus Obermayer. Graph quantization. Computer Vision and Image Understanding: CVIU, 115(7):946– 961, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211000920.

Jokinen:1998:ABM

[Jok98] Olli Jokinen. Area-based matching for simultaneous registration of multiple 3D profile maps. Computer Vision and Image Understanding: CVIU, 71(3):431– 447, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0639/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0639/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0639/production/ref;
http://www.idealibrary.
com/links/artid/cviu.1998.
0713/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0713/production/pdf.

Jones:1997:COH

[Jon97] Graeme A. Jones. Constraint, optimization, and hierarchy: Reviewing stereoscopic correspondence of complex features. Computer Vision and Image Understanding: CVIU, 65(1): 57–78, January 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0482/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0482/production/pdf;
http://www.idealibrary.

com/links/artid/cviu.1996.
0482/production/ref.

Jones:1999:CFS

[Jon99]  Ronald Jones.  Con-

nected filtering and segmentation using component trees. Computer Vision and Image Understanding: CVIU, 75(3):215– 228, September 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0777/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0777/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0777/production/ref.

Jones:2008:SII

[Jon08] Graeme A. Jones. Special issue on Intelligent Visual Surveillance. Computer Vision and Image Understanding: CVIU, 111(1): 1, July 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Joseph:1999:OPE

[Jos99] S. H. Joseph. Optimal pose estimation in two and three dimensions. Computer Vision and Image Understanding: CVIU, 73(2):215– 231, February 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0733/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0733/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0733/production/ref.

Jost:2005:ACC

[JOvW^{+}05] Timothée Jost, Nabil Ouerhani, Roman von Wartburg, René Müri, and Heinz Hügli. Assessing the contribution of color in visual attention. Computer Vision and Image Understanding: CVIU, 100(1–2):107–123, October/ November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ji:2022:DAA

[JPN^{+}22] Sangwoo Ji, Namgyu Park, Dongbin Na, Bin Zhu, and Jong Kim. Defending against attacks tailored to transfer learning via feature distancing. Computer Vision and Image Understanding: CVIU, 223(??): ??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422200114X.

Jung:2014:SRU

[JPP^{+}14] Ho Yub Jung, Haesol Park, In Kyu Park, Kyoung Mu Lee, and Sang Uk Lee. Stereo reconstruction using high-order likelihoods. Computer Vision and Image Understanding: CVIU, 125(??):223– 236, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000964.

Jia:2017:MML

[JRAJ17] Jieru Jia, Qiuqi Ruan, Gaoyun An, and Yi Jin. Multiple metric learning with query adaptive weights and multi-task re-weighting for person re-identification. Computer Vision and Image Understanding: CVIU, 160 (??):87–99, July 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300619.

Johnson-Roberson:2015:DSR

[JRBD^{+}15] Matthew Johnson-Roberson, Mitch Bryson, Bertrand Douillard, Oscar Pizarro, and Stefan B. Williams. Discovering salient regions on 3D photo-textured maps: Crowdsourcing interaction data from multitouch smartphones and tablets. Computer Vision and Image Understanding: CVIU, 131 (??):28–41, February 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314214001593.

Jaynes:2003:RRB

[JRH03] Christopher Jaynes, Edward Riseman, and Allen Hanson. Recognition and reconstruction of buildings from multiple aerial images. Computer Vision and Image Understanding: CVIU, 90(1): 68–98, April 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jiang:2024:RCF

[JRH24] Hai Jiang, Yang Ren, and Songchen Han. Revisiting coarse-to-fine strategy for low-light image enhancement with deep decomposition guided training. Computer Vision and Image Understanding: CVIU, 241 (??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422400033X.

Jamali-Rad:2021:LAL

[JRS21] Hadi Jamali-Rad and Attila Szabó. Lookahead adversarial learning for near real-time semantic segmentation. Computer Vision and Image Understanding: CVIU, 212(??): ??, November 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001156.

Jaimes:2007:MHC

[JS07] Alejandro Jaimes and Nicu Sebe. Multimodal human– computer interaction: a survey. Computer Vision and Image Understanding: CVIU, 108(1–2):116–134, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jung:2023:RKB

[JSC23] Seunghwan Jung, Yeong-Gil Shin, and Minyoung Chung. Robust kernel-based feature representation for 3D point cloud analysis via circular convolutional network. Computer Vision and Image Understanding: CVIU, 231(??):??, June 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000589.

Javed:2008:MIC

[JSRS08] Omar Javed, Khurram Shafique, Zeeshan Rasheed, and Mubarak Shah. Modeling inter-camera space– time and appearance relationships for tracking across non-overlapping views. Computer Vision and Image Understanding: CVIU, 109 (2):146–162, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Javanbakhat:2024:QMU

[JSWL24] Masoumeh Javanbakhat, Ludger Starke, Sonia Waiczies, and Christoph Lippert. Quantifying model uncertainty for semantic segmentation of fluorine-19 MRI using stochastic gradient MCMC. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000481.

Jiang:2017:IDU

[JSZY17] Yutong Jiang, Changming Sun, Yu Zhao, and Li Yang. Image dehazing using adaptive bi-channel priors on superpixels. Computer Vision and Image Understanding: CVIU, 165(??):17–32, December 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301777.

Jin:2017:FAW

[JT17] Xin Jin and Xiaoyang Tan.

Face alignment in-the-wild: a survey. Computer Vision and Image Understanding: CVIU, 162(??):1–22, September 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301455.

Jurie:1999:SSP

[Jur99] Frederic Jurie. Solution of the simultaneous pose and correspondence problem using Gaussian error model. Computer Vision and Image Understanding: CVIU, 73(3):357– 373, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0735/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0735/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0735/production/ref.

Janssen:1997:AVL

[JV97] Rik D. T. Janssen and Albert M. Vossepoel. Adaptive vectorization of line drawing images. Computer Vision and Image Understanding: CVIU, 65(1):38– 56, January 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0484/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0484/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0484/production/ref.

Jain:2020:LGE

[JVD^{+}20] Monika Jain, Subramanyam A. V., Simon Denman, Sridha Sridharan, and Clinton Fookes. LSTM guided ensemble correlation filter tracking with appearance model pool. Computer Vision and Image Understanding: CVIU, 195(??):Article 102935, June 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300229.

Jonk:1999:GID

[JvdBS99] Arnold Jonk, Rein van den Boomgaard, and Arnold Smeulders. Grammatical inference of dashed lines. Computer Vision and Image Understanding: CVIU, 74(3):212– 226, June 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/

links/artid/cviu.1999.
0753/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0753/production/pdf.

Ji:2005:SIE

[JWDF05] Qiang Ji, Harry Wechsler, Andrew Duchowski, and Myron Flickner. Special issue: eye detection and tracking. Computer Vision and Image Understanding: CVIU, 98(1):1– 3, April 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jeong:2004:IRU

[JWG04] Sangoh Jeong, Chee Sun Won, and Robert M. Gray. Image retrieval using color histograms generated by Gauss mixture vector quantization. Computer Vision and Image Understanding: CVIU, 94(1–3):44–66, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jiang:2014:PTS

[JY14] Jinwei Jiang and Alper Yilmaz. Persistent tracking of static scene features using geometry. Computer Vision and Image Understanding: CVIU, 120(??): 141–156, March 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213002002.

Jiang:2023:III

[JYL23] Jin Jiang, Xiaoyuan Yang, and Yixiao Li. IIANet: Information interactivity attention network with adversarial learning for infrared small object detection. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223002199.

Jiang:2011:AVE

[JYTK11] Fan Jiang, Junsong Yuan, Sotirios A. Tsaftaris, and Aggelos K. Katsaggelos. Anomalous video event detection using spatiotemporal context. Computer Vision and Image Understanding: CVIU, 115(3):323– 333, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jiang:2023:TGF

[JYX^{+}23] Weihao Jiang, Dongdong Yu, Zhaozhi Xie, Yaoyi Li, Zehuan Yuan, and Hongtao Lu. Trimap-guided feature mining and fusion network for natural image matting. Computer Vision and Image Understanding: CVIU, 230(??):??, April 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000255.

Jiang:2024:CTD

[JYY^{+}24] Yimin Jiang, Tingfei Yan, Mingze Yao, Huibing Wang, and Wenzhe Liu. Cascade transformers with dynamic attention for video question answering. Computer Vision and Image Understanding: CVIU, 242(??): ??, May 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731422400064X.

Jing:2023:SRE

[JZL^{+}23] Yingshan Jing, Ting Zhang, Zhaoying Liu, Yuewu Hou, and Changming Sun. Swin-ResUNet+: an edge enhancement module for road extraction from remote sensing images. Computer Vision and Image Understanding: CVIU, 237(??): ??, December 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422300187X.

Jiang:2016:ECW

[JZWD16] Hairong Jiang, Ting Zhang, Juan P. Wachs, and Bradley S. Duerstock. Enhanced control of a wheelchairmounted robotic manipulator using 3-D vision and multimodal interaction. Computer Vision and Image Understanding: CVIU, 149(??):21– 31, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300066.

Ji:2023:EED

[JZZ23] Fangting Ji, Xin Zhang, and Junlong Zhao. α-EGAN: α-energy distance GAN with an early stopping rule. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001285.

Jiang:2023:ISG

[JZZM23] Xingyu Jiang, Shihua Zhang, Xiao-Ping Zhang, and Jiayi Ma. Improving sparse graph attention for feature matching by informative keypoints exploration. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001832.

Keller:2008:GPI

[KA08] Y. Keller and A. Averbuch.

Global parametric image alignment via high-order approximation. Computer Vision and Image Understanding: CVIU, 109(3):244– 259, March 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Krishnamoorthi:2012:SBE

[KA12] R. Krishnamoorthi and G. Annapoorani. A simple boundary extraction technique for irregular pupil localization with orthogonal polynomials. Computer Vision and Image Understanding: CVIU, 116(2):262–273, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002013.

Knerr:1998:HMM

[KABP98] S. Knerr, E. Augustin, O. Baret, and D. Price. Hidden Markov model based word recognition and its application to legal amount reading on French checks. Computer Vision and Image Understanding: CVIU, 70(3):404–419, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0685/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0685/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0685/production/ref.

Kugelman:2023:EOC

[KACR^{+}23] Jason Kugelman, David Alonso-Caneiro, Scott A. Read, Stephen J. Vincent, and Michael J. Collins. Enhanced OCT chorioretinal segmentation in lowdata settings with semisupervised GAN augmentation using cross-localisation. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314223002321.

Kanth:1999:DRS

[KAES99] K. V. Ravi Kanth, Divyakant Agrawal, Amr El Abbadi, and Ambuj Singh. Dimensionality reduction for similarity searching in dynamic databases. Computer Vision and Image Understanding: CVIU, 75(1–2):59–72, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0762/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0762/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0762/production/ref;
http://www.idealibrary.
com/links/artid/cviu.1999.
0763/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0763/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0763/production/ref.

Kak:1995:E

[Kak95] Avi Kak. Editorial. Computer Vision and Image Understanding: CVIU, 61(2): 153, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1012/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1012/production/pdf.

Kak:1997:THF

[Kak97] Avi Kak. Tribute: Herb Freeman. Computer Vision and Image Understanding: CVIU, 68(3):255– ??, December 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0660/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0660/production/pdf.

Kimmel:1995:GSS

[KB95a] Ron Kimmel and Alfred M. Bruckstein. Global shape from shading. Computer Vision and Image Understanding: CVIU, 62(3):360– 369, November 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1060/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1060/production/pdf.

Kimmel:1995:TLS

[KB95b] Ron Kimmel and Alfred M. Bruckstein. Tracking level sets by level sets: a method for solving the shape from shading problem. Computer Vision and Image Understanding: CVIU, 62(1): 47–58, July 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1040/production;   http:
                                        //www.idealibrary.com/

links/artid/cviu.1995.
1040/production/pdf.

Kanai:1998:SID

[KB98] Junichi Kanai and Henry S.

Baird. Special issue on document image understanding and retrieval. Computer Vision and Image Understanding: CVIU, 70(3): 285–286, June 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0693/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0693/production/pdf.

Kiryati:2000:HHT

[KB00] Nahum Kiryati and Alfred M. Bruckstein. Heteroscedastic Hough transform (HtHT): An efficient method for robust line fitting in the ‘errors in the variables’ problem. Computer Vision and Image Understanding: CVIU, 78(1): 69–83, April 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0828;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0828/pdf;   http:
//www.idealibrary.com/

                                        links/doi/10.1006/cviu.
                                        1999.0828/ref.

Krausz:2012:LAV

[KB12] Barbara Krausz and Christian Bauckhage. Loveparade 2010: Automatic video analysis of a crowd disaster. Computer Vision and Image Understanding: CVIU, 116(3):307– 319, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211002037.

Khandelwal:2016:LGC

[KBAS16] Dinesh Khandelwal, Kush Bhatia, Chetan Arora, and Parag Singla. Lazy generic cuts. Computer Vision and Image Understanding: CVIU, 143(??):80–91, February 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215002362.

Kontschieder:2012:EHG

[KBD^{+}12] Peter Kontschieder, Samuel Rota Bulò, Michael Donoser, Marcello Pelillo, and Horst Bischof. Evolutionary Hough Games for coherent object detection. Computer Vision and Image Understanding: CVIU, 116(11): 1149–1158, November 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314212001142.

Kukelova:2010:FRN

[KBJ^{+}10] Zuzana Kukelova, Martin Byröd, Klas Josephson, Tomas Pajdla, and Kalle Åström. Fast and robust numerical solutions to minimal problems for cameras with radial distortion. Computer Vision and Image Understanding: CVIU, 114 (2):234–244, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kaltsa:2018:MHD

[KBKS18] Vagia Kaltsa, Alexia Briassouli, Ioannis Kompatsiaris, and Michael G. Strintzis. Multiple Hierarchical Dirichlet Processes for anomaly detection in traffic. Computer Vision and Image Understanding: CVIU, 169(??): 28–39, April 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300110.

Khan:2015:SHS

[KBMD15] Rahat Khan, Cécile Barat, Damien Muselet, and Christophe Ducottet. Spatial histograms of soft pairwise similar patches to improve the bag-of-visual-words model. Computer Vision and Image Understanding: CVIU, 132(??): 102–112, March 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001878.

Karavasilis:2012:NFM

[KBN12] Vasileios Karavasilis, Konstantinos Blekas, and Christophoros Nikou. A novel framework for motion segmentation and tracking by clustering incomplete trajectories. Computer Vision and Image Understanding: CVIU, 116(11):1135–1148, November 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001063.

Kennedy:2016:OAF

[KBWT16] Ryan Kennedy, Laura Balzano, Stephen J. Wright, and Camillo J. Taylor. Online algorithms for factorizationbased structure from motion. Computer Vision and Image Understanding: CVIU, 150(??):139– 152, September 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300364.

Kyatkin:1999:PMC

[KC99] Alexander B. Kyatkin and Gregory S. Chirikjian. Pattern matching as a correlation on the discrete motion group. Computer Vision and Image Understanding: CVIU, 74(1): 22–35, April 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0745/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0745/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0745/production/ref.

Kim:2022:LST

[KC22] Boeun Kim and Jin Young Choi. Learning spectral transform for 3D human motion prediction. Computer Vision and Image Understanding: CVIU, 223 (??):??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001266.

Khachan:2000:PRA

[KCD00] Mohammed Khachan, Patrick Chenin, and Hafsa Deddi. Polyhedral representation and adjacency graph in n-dimensional digital images. Computer Vision and Image Understanding: CVIU, 79(3):428–441, September 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0859;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0859/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0859/ref.

khadiri:2018:LDT

[kCE^{+}18] I. El khadiri, A. Chahi, Y. El Merabet, Y. Ruichek, and R. Touahni. Local directional ternary pattern: a new texture descriptor for texture classification. Computer Vision and Image Understanding: CVIU, 169 (??):14–27, April 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300043.

Kim:2017:MVF

[KCM^{+}17] Donghun Kim, Bharath Comandur, Henry Medeiros, Noha M. Elfiky, and Avinash C. Kak. Multi-view face recognition from single RGBD models of the faces. Computer Vision and Image Understanding: CVIU, 160(??):114– 132, July 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300693.

Kou:2018:PRF

[KCZ18] Wen Kou, Loong-Fah Cheong, and Zhiying Zhou. Proximal robust factorization for piecewise planar reconstruction. Computer Vision and Image Understanding: CVIU, 166(??):88–101, January 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301650.

Koplowitz:1996:HRC

[KD96] Jack Koplowitz and Joseph DeLeone. Hierarchical representation of chainencoded binary image contours. Computer Vision and Image Understanding: CVIU, 63(2):344– 352, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0024/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0024/production/pdf.

Kaess:2010:PSM

[KD10] Michael Kaess and Frank Dellaert. Probabilistic structure matching for visual SLAM with a multicamera rig. Computer Vision and Image Understanding: CVIU, 114(2):286–296, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kia:1998:SCP

[KDRC98] Omid E. Kia, David S. Doermann, Azriel Rosenfeld, and Rama Chellapa. Symbolic compression and processing of document images. Computer Vision and Image Understanding: CVIU, 70(3): 335–349, June 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0682/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0682/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0682/production/ref.

Kamenou:2023:LOB

[KdRM^{+}23] Eleni Kamenou, Jesus Martinez del Rincon, Paul Miller, Patricia Devlin-Hill, Samuel Budgett, Federico Angelini, and Charlotte Grinyer. LOFReg: an outlier-based regulariser for deep metric learning. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731422300200X.

Khatun:2020:JIV

[KDSF20] Amena Khatun, Simon Denman, Sridha Sridharan, and Clinton Fookes. Joint identification-verification for person re-identification: a four stream deep learning approach with improved quartet loss function. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 102989, August 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300618.

Kiforenko:2018:PEP

[KDT^{+}18] Lilita Kiforenko, Bertram Drost, Federico Tombari, Norbert Krüger, and Anders Glent Buch. A performance evaluation of point pair features. Computer Vision and Image Understanding: CVIU, 166(??): 66–80, January 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301601.

Kosmopoulos:2012:BFB

[KDV12] Dimitrios I. Kosmopoulos, Nikolaos D. Doulamis, and Athanasios S. Voulodimos. Bayesian filter based behavior recognition in workflows allowing for user feedback. Computer Vision and Image Understanding: CVIU, 116(3):422– 434, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211002050.

Kwon:2016:TSS

[KDV16] Junseok Kwon, Ralf Dragon, and Luc Van Gool. Tracking by switching state space models. Computer Vision and Image Understanding: CVIU, 153(??):29–36, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216000783.

Kwon:1999:ACF

[KdVL99] Young H. Kwon and Niels da Vitoria Lobo. Age classification from facial images. Computer Vision and Image Understanding: CVIU, 74(1):1–21, April 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0549/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0549/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0549/production/ref.

Kooij:2015:IMO

[KEG15] Julian F. P. Kooij, Gwenn Englebienne, and Dariu M. Gavrila. Identifying multiple objects from their appearance in inaccurate detections. Computer Vision and Image Understanding: CVIU, 136(??):103– 116, July 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000697.

Koller:2015:CSL

[KFN15] Oscar Koller, Jens Forster, and Hermann Ney. Continuous sign language recognition: Towards large vocabulary statistical recognition systems handling multiple signers. Computer Vision and Image Understanding: CVIU, 141(??):108– 125, December 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002088.

Kazantzidis:2018:VOG

[KFRD^{+}18] Ioannis Kazantzidis, Francisco Florez-Revuelta, Mickael Dequidt, Natasha Hill, and Jean-Christophe Nebel. Vide-omics: a genomicsinspired paradigm for video analysis. Computer Vision and Image Understanding: CVIU, 166(??):28–40, January 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217301662.

Kasiri:2017:FGA

[KFSM17] Soudeh Kasiri, Clinton Fookes, Sridha Sridharan, and Stuart Morgan. Finegrained action recognition of boxing punches from depth imagery. Computer Vision and Image Understanding: CVIU, 159(??): 143–153, June 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300668.

Khan:2014:ODS

[KG14] Zulfiqar Hasan Khan and Irene Yu-Hua Gu. Online domain-shift learning and object tracking based on nonlinear dynamic models and particle filters on Riemannian manifolds. Computer Vision and Image Understanding: CVIU, 125(??):97– 114, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000836.

Krejov:2017:GOT

[KGB17] Philip Krejov, Andrew Gilbert, and Richard Bowden. Guided optimisation through classification and regression for hand pose estimation. Computer Vision and Image Understanding: CVIU, 155(??):124– 138, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421630193X.

Kim:2005:RAD

[KGC05] Hae Yong Kim, Javier Giacomantone, and Zang Hee Cho. Robust anisotropic diffusion to produce enhanced statistical parametric map from noisy fMRI. Computer Vision and Image Understanding: CVIU, 99(3): 435–452, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2010:JRC

[KGFP10] Seon Joo Kim, David Gallup, Jan-Michael Frahm, and Marc Pollefeys. Joint radiometric calibration and feature tracking system with an application to stereo. Computer Vision and Image Understanding: CVIU, 114(5):574–582, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2010:ESC

[KGK10] Jun-Sik Kim, Pierre Gurdjos, and In So Kweon. Euclidean structure from confocal conics: Theory and application to camera calibration. Computer Vision and Image Understanding: CVIU, 114(7):803– 812, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kuznetsov:2019:IRP

[KGM19] A. O. Kuznetsov, A. V. Gorevoy, and A. S. Machikhin. Image rectification for prismbased stereoscopic optical systems. Computer Vision and Image Understanding: CVIU, 182(??):30– 37, May 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219300189.

Kucuktunc:2010:FCH

[KGU10] Onur Küçüktunç, Uğur Güdükbay, and Özgür Ulusoy. Fuzzy color histogrambased video segmentation. Computer Vision and Image Understanding: CVIU, 114 (1):125–134, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Khoubyari:1996:FFW

[KH96] Siamak Khoubyari and Jonathan J. Hull. Font and function word identification in document recognition. Computer Vision and Image Understanding: CVIU, 63(1):66– 74, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0005/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0005/production/pdf.

Kruger:2013:TOA

[KH13] Volker Krüger and Dennis Herzog. Tracking in object action space. Computer Vision and Image Understanding: CVIU, 117(7): 764–789, July 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000313.

Kim:2015:BWR

[KH15] Hansung Kim and Adrian Hilton. Block world reconstruction from spherical stereo image pairs. Computer Vision and Image Understanding: CVIU, 139 (??):104–121, October 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314215000831.

Kim:2023:IVI

[KH23] Jun-Hyung Kim and Youngbae Hwang. Infrared and visible image fusion using a guiding network to leverage perceptual similarity. Computer Vision and Image Understanding: CVIU, 227 (??):??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422200176X.

Kong:2005:RAV

[KHA^{+}05] Seong G. Kong, Jingu Heo, Besma R. Abidi, Joonki Paik, and Mongi A. Abidi. Recent advances in visual and infrared face recognition—a review. Computer Vision and Image Understanding: CVIU, 97 (1):103–135, January 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kita:2001:CBD

[KHB01] Yasuyo Kita, Ralph Highnam, and Michael Brady. Correspondence between different view breast X rays using curved epipolar lines. Computer Vision and Image Understanding: CVIU, 83(1):38–56, July 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0908;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0908/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0908/ref.

Ke:2022:WSF

[KHG22] Xiao Ke, Yanyan Huang, and WenZhong Guo. Weakly supervised fine-grained image classification via twolevel attention activation model. Computer Vision and Image Understanding: CVIU, 218(??): ??, April 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000364.

Kim:2012:PCM

[KHH^{+}12] Kunho Kim, Mohammad K.

Hasan, Jae-Pil Heo, Yu-Wing Tai, and Sung eui Yoon. Probabilistic cost model for nearest neighbor search in image retrieval. Computer Vision and Image Understanding: CVIU, 116(9):991–998, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212000768.

Kopuklu:2022:DCT

[KHH^{+}22] Okan Köpüklü, Stefan Hörmann, Fabian Herzog, Hakan Cevikalp, and Gerhard Rigoll. Dissected 3D CNNs: Temporal skip connections for efficient online video processing. Computer Vision and Image Understanding: CVIU, 215(??): ??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001594.

Kim:2010:SAM

[KHK10] Jun-Sik Kim, Myung Hwangbo, and Takeo Kanade. Spherical approximation for multiple cameras in motion estimation: Its applicability and advantages. Computer Vision and Image Understanding: CVIU, 114(10): 1068–1083, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kamal:2016:TLR

[KHR^{+}16] Mahdad Hosseini Kamal, Barmak Heshmat, Ramesh Raskar, Pierre Vandergheynst, and Gordon Wetzstein. Tensor low-rank and sparse light field photography. Computer Vision and Image Understanding: CVIU, 145 (??):172–181, April 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314215002465.

Kenmochi:1998:BED

[KII98] Yukiko Kenmochi, Atsushi Imiya, and Akira Ichikawa. Boundary extraction of discrete objects. Computer Vision and Image Understanding: CVIU, 71(3):281– 293, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0652/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0652/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0652/production/ref.

Kim:2004:EEU

[Kim04] Seung-Bum Kim. Eliminating extrapolation using point distribution criteria in scattered data interpolation.

Computer Vision and Image Understanding: CVIU, 95 (1):30–53, July 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2015:MCF

[Kim15] Minyoung Kim. Multipleconcept feature generative models for multi-label image classification. Computer Vision and Image Understanding: CVIU, 136(??): 69–78, July 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214002239.

Kim:2017:DSA

[Kim17]  Minyoung Kim.  Dual

soft assignment clustering algorithm for human action video clustering. Computer Vision and Image Understanding: CVIU, 155(??):106– 112, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216302041.

Kiselman:1996:RPD

[Kis96a] Christer O. Kiselman. Regularity properties of distance transformations in image analysis. Computer Vision and Image Understanding: CVIU, 64(3):390– 398, November 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0067/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0067/production/pdf.

Kishimoto:1996:CDC

[Kis96b] Kazuo Kishimoto. Characterizing digital convexity and straightness in terms of “Length” and “Total Absolute Curvature”. Computer Vision and Image Understanding: CVIU, 63(2): 326–333, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0022/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0022/production/pdf.

Kerola:2017:CVH

[KIS17] Tommi Kerola, Nakamasa Inoue, and Koichi Shinoda. Cross-view human action recognition from depth maps using spectral graph sequences. Computer Vision and Image Understanding: CVIU, 154(??):108– 126, January 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301588.

Kim:2007:RMB

[KK07] Sungho Kim and In So Kweon. Robust modelbased scene interpretation by multilayered context information. Computer Vision and Image Understanding: CVIU, 105(3):167– 187, March 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2009:CCB

[KK09] Jun-Sik Kim and In So Kweon. Camera calibration based on arbitrary parallelograms. Computer Vision and Image Understanding: CVIU, 113(1):1–10, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2011:MRP

[KK11] Jun-Sik Kim and In So Kweon. Metric reconstruction of planes utilizing offthe-plane features. Computer Vision and Image Understanding: CVIU, 115(1): 1–7, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kasaiezadeh:2013:MAS

[KK13] Alireza Kasaiezadeh and Amir Khajepour. Multiagent stochastic level set method in image segmentation. Computer Vision and Image Understanding: CVIU, 117(9):1147–1162, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000945.

Kane:2015:FLC

[KK15] Lalit Kane and Pritee Khanna. A framework for live and cross platform fingerspelling recognition using modified shape matrix variants on depth silhouettes. Computer Vision and Image Understanding: CVIU, 141(??):138– 151, December 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001630.

Kafle:2017:VQA

[KK17] Kushal Kafle and Christopher Kanan. Visual question answering: Datasets, algorithms, and future challenges. Computer Vision and Image Understanding: CVIU, 163(??):3–20, October 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301170.

Kim:2023:ECA

[KKCK23] Bum Jun Kim, Gyogwon Koo, Hyeyeon Choi, and Sang Woo Kim. Extending class activation mapping using Gaussian receptive field. Computer Vision and Image Understanding: CVIU, 231(??):??, June 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223000437.

Kini:2024:CVC

[KKKS24] Jyoti Kini, Fahad Shahbaz Khan, Salman Khan, and Mubarak Shah. CT-VOS: Cutout prediction and tagging for self-supervised video object segmentation. Computer Vision and Image Understanding: CVIU, 238 (??):??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223002400.

Kim:2024:ANA

[KKP24] Dogun Kim, Jin Kim, and Eunil Park. AFA-Net: Adaptive Feature Attention Network in image deblurring and super-resolution for improving license plate recognition. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422300259X.

Kajo:2023:TBC

[KKRK23] Ibrahim Kajo, Mohamed Kas, Yassine Ruichek, and Nidal Kamel. Tensor based completion meets adversarial learning: a win-win solution for change detection on unseen videos. Computer Vision and Image Understanding: CVIU, 226(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422200162X.

Kim:2023:ADI

[KKSC23] Jiwan Kim, Minchang Kim, Yeong-Gil Shin, and Minyoung Chung. Accurate depth image generation via overfit training of point cloud registration using local frame sets. Computer Vision and Image Understanding: CVIU, 226(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001667.

Kim:2023:RKV

[KKSH23] Hyeonwoo Kim, Hyungjoon Kim, Jonghwa Shim, and Eenjun Hwang. A robust kinship verification scheme using face age transformation. Computer Vision and Image Understanding: CVIU, 231(??): ??, June 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223000425.

Kadoury:2007:FDG

[KL07] Samuel Kadoury and Martin D. Levine. Face detection in gray scale images using locally linear embeddings. Computer Vision and Image Understanding: CVIU, 105(1):1–20, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kumar:2010:IMB

[KL10] Avin Kumar and Baoxin Li. On implementing motionbased Region of Interest detection on multi-core CELL. Computer Vision and Image Understanding: CVIU, 114(11):1139–1151, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2011:HAM

[KL11] Wonsik Kim and Kyoung Mu Lee. A hybrid approach for MRF optimization problems: Combination of stochastic sampling and deterministic algorithms. Computer Vision and Image Understanding: CVIU, 115(12):1623– 1637, December 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001640.

Komodakis:2011:SIO

[KLBP11] Nikos Komodakis, Georg Langs, Horst Bischof, and Nikos Paragios. Special issue on optimization for vision, graphics and medical imaging: theory and applications. Computer Vision and Image Understanding: CVIU, 115(12):1597, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002165.

Klette:2013:RCM

[Kle13] Gisela Klette. Recursive computation of minimumlength polygons. Computer Vision and Image Understanding: CVIU, 117(4): 386–392, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001865.

Kim:2014:RBP

[KLK14] Jungho Kim, Zhe Lin, and In So Kweon. Rao– Blackwellized particle filtering with Gaussian mixture models for robust visual tracking. Computer Vision and Image Understanding: CVIU, 125(??): 128–137, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000800.

Kooij:2016:MMH

[KLK^{+}16] J. F. P. Kooij, M. C. Liem, J. D. Krijnders, T. C. Andringa, and D. M. Gavrila. Multi-modal human aggression detection. Computer Vision and Image Understanding: CVIU, 144(??): 106–120, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001381.

Koch:2020:CMD

[KLKF20] Tobias Koch, Lukas Liebel, Marco Körner, and Friedrich Fraundorfer. Comparison of monocular depth estimation methods using geometrically relevant metrics on the IBims-1 dataset. Computer Vision and Image Understanding: CVIU, 191(??):Article 102877, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219301663.

Kim:2011:RBT

[KLL^{+}11] Kyung-Su Kim, Min-Jeong Lee, Ji-Won Lee, Tae-Woo Oh, and Hae-Yeoun Lee. Region-based tampering detection and recovery using homogeneity analysis in quality-sensitive imaging. Computer Vision and Image Understanding: CVIU, 115(9):1308– 1323, September 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001238.

Kim:2020:LIA

[KLO20] Nuri Kim, Donghoon Lee, and Songhwai Oh. Learning instance-aware object detection using determinantal point processes. Computer Vision and Image Understanding: CVIU, 201 (??):Article 103061, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220301041.

Koh:2021:SID

[KLY21] Jaihyun Koh, Jangho Lee, and Sungroh Yoon. Singleimage deblurring with neural networks: a comparative survey. Computer Vision and Image Understanding: CVIU, 203(??):Article 103134, February 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220301533.

Kowalczyk:2003:OCD

[KM03] M. Kowalczyk and W. S. Mokrzycki. Obtaining complete 2 D view representation of polyhedra using concept of seedling singleview area. Computer Vision and Image Understanding: CVIU, 91(3):280–301, September 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2017:IMS

[KM17] Chelhwon Kim and Roberto Manduchi. Indoor Manhattan spatial layout recovery from monocular videos via line matching. Computer Vision and Image Understanding: CVIU, 157(??): 223–239, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301448.

Krissian:2000:MBD

[KMA^{+}00] Karl Krissian, Grégoire Malandain, Nicholas Ayache, Régis Vaillant, and Yves Trousset. Model-based detection of tubular structures in 3D images. Computer Vision and Image Understanding: CVIU, 80(2): 130–171, November 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0866;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0866/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0866/ref.

Kakadiaris:1997:IOS

[KMB97] Ioannis Kakadiaris, Dimitris Metaxas, and Ruzena Bajcsy. Inferring 2D object structure from the deformation of apparent contours. Computer Vision and Image Understanding: CVIU, 65(2):129– 147, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0580/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0580/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0580/production/ref.

Kroon:2009:ELL

[KMBH09] Bart Kroon, Sander Maas, Sabri Boughorbel, and Alan Hanjalic. Eye localization in low and standard definition content with application to face matching. Computer Vision and Image Understanding: CVIU, 113(8): 921–933, August 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kuo:2011:IBT

[KMN11] Paul Kuo, Dimitrios Makris, and Jean-Christophe Nebel. Integration of bottomup/top-down approaches for 2D pose estimation using probabilistic Gaussian modelling. Computer Vision and Image Understanding: CVIU, 115(2):242–255, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Koutis:2011:CPM

[KMT11] Ioannis Koutis, Gary L. Miller, and David Tolliver. Combinatorial preconditioners and multilevel solvers for problems in computer vision and image processing. Computer Vision and Image Understanding: CVIU, 115(12):1638–1646, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211001627.

Kim:1999:URL

[KN99] ZuWhan Kim and Ramakant Nevatia. Uncertain reasoning and learning for feature grouping. Computer Vision and Image Understanding: CVIU, 76(3):278– 288, December 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0803/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0803/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0803/production/ref.

Khamene:2003:MSM

[KN03] Ali Khamene and Shahriar Negahdaripour. Motion and structure from multiple cues; image motion, shading flow, and stereo disparity. Computer Vision and Image Understanding: CVIU, 90(1):99–127, April 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2004:ADC

[KN04] ZuWhan Kim and Ramakant Nevatia. Automatic description of complex buildings from multiple images. Computer Vision and Image Understanding: CVIU, 96(1):60– 95, October 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kybic:2011:BOF

[KN11] Jan Kybic and Claudia Nieuwenhuis. Bootstrap optical flow confidence and uncertainty measure. Computer Vision and Image Understanding: CVIU, 115(10):1449–1462, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211001536.

Karavasilis:2015:VTU

[KNL15] Vasileios Karavasilis, Christophoros Nikou, and Aristidis Likas. Visual tracking using spatially weighted likelihood of Gaussian mixtures. Computer Vision and Image Understanding: CVIU, 140 (??):43–57, November 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314215001563.

Kurazume:2009:RFS

[KNO^{+}09] Ryo Kurazume, Kaori Nakamura, Toshiyuki Okada, Yoshinobu Sato, Nobuhiko Sugano, Tsuyoshi Koyama, Yumi Iwashita, and Tsutomu Hasegawa. 3D reconstruction of a femoral shape using a parametric model and two 2D fluoroscopic images. Computer Vision and Image Understanding: CVIU, 113(2):202–211, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kuzborskij:2017:SGA

[KOC17] Ilja Kuzborskij, Francesco Orabona, and Barbara Caputo. Scalable greedy algorithms for transfer learning. Computer Vision and Image Understanding: CVIU, 156(??):174– 185, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301370.

Kurazume:2017:ALS

[KON^{+}17] Ryo Kurazume, Souichiro Oshima, Shingo Nagakura, Yongjin Jeong, and Yumi Iwashita. Automatic largescale three dimensional modeling using cooperative multiple robots. Computer Vision and Image Understanding: CVIU, 157(??): 25–42, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300558.

Kumar:2010:ITC

[KORC10] S. Kumar, S. H. Ong, S. Ranganath, and F. T. Chew. Invariant texture classification for biomedical cell specimens via nonlinear polar map filtering. Computer Vision and Image Understanding: CVIU, 114(1):44–53, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kouzani:2003:LHF

[Kou03] A. Z. Kouzani. Locating human faces within images. Computer Vision and Image Understanding: CVIU, 91(3):247–279, September 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:1997:SSP

[KP97] Bang-Hwan Kim and Rae-Hong Park. Shape from shading and photometric stereo using surface approximation by Legendre polynomials. Computer Vision and Image Understanding: CVIU, 66(3): 255–270, June 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0515/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0515/production/pdf;

http://www.idealibrary.
com/links/artid/cviu.1997.
0515/production/ref.

Kruger:2000:OOR

[KP00] Norbert Krüger and Gabriele Peters. ORASSYLL: Object recognition with autonomously learned and sparse symbolic representations based on metrically organized local line detectors. Computer Vision and Image Understanding: CVIU, 77(1):48– 77, January 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0794/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0794/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0794/production/ref.

Krsek:2002:DIB

[KPH02] Pavel Krsek, Tomá Pajdla, and Václav Hlavá. Differential invariants as the base of triangulated surface registration. Computer Vision and Image Understanding: CVIU, 87(1–3): 27–38, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kolsch:2007:SIV

[KPKH07] Mathias Kölsch, Vladimir Pavlović, Branislav Kisačanin, and Thomas S. Huang. Special issue on vision for human–computer interaction. Computer Vision and Image Understanding: CVIU, 108(1–2):1–3, October/November 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kristan:2009:CWT

[KPPK09] Matej Kristan, Janez Perš, Matej Perše, and Stanislav Kovačič. Closed-world tracking of multiple interacting targets for indoorsports applications. Computer Vision and Image Understanding: CVIU, 113(5): 598–611, May 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kliot:1998:IBS

[KR98] Michael Kliot and Ehud Rivlin. Invariant-based shape retrieval in pictorial databases. Computer Vision and Image Understanding: CVIU, 71(2):182– 197, August 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0709/production;   http:
                                        //www.idealibrary.com/

links/artid/cviu.1998.
0709/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0709/production/ref.

Kundur:1999:NAV

[KR99] Sridhar R. Kundur and Daniel Raviv. Novel active vision-based visual threat cue for autonomous navigation tasks. Computer Vision and Image Understanding: CVIU, 73(2):169– 182, February 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0699/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0699/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0699/production/ref.

Kobyshev:2017:EAS

[KRBSV17] Nikolay Kobyshev, Hayko Riemenschneider, András Bódis-Szomorú, and Luc Van Gool. Efficient architectural structural element decomposition. Computer Vision and Image Understanding: CVIU, 157(??): 300–312, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300807.

Kuehne:2017:WSL

[KRG17] Hilde Kuehne, Alexander Richard, and Juergen Gall. Weakly supervised learning of actions from transcripts. Computer Vision and Image Understanding: CVIU, 163(??):78–89, October 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217301169.

Kilambi:2008:EPC

[KRJ^{+}08] Prahlad Kilambi, Evan Ribnick, Ajay J. Joshi, Osama Masoud, and Nikolaos Papanikolopoulos. Estimating pedestrian counts in groups. Computer Vision and Image Understanding: CVIU, 110(1):43–59, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kjellstrom:2011:VOA

[KRK11] Hedvig Kjellström, Javier Romero, and Danica Kragić. Visual object-action recognition: Inferring object affordances from human demonstration. Computer Vision and Image Understanding: CVIU, 115(1):81– 90, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kviatkovsky:2014:OAR

[KRS14] Igor Kviatkovsky, Ehud Rivlin, and Ilan Shimshoni. Online action recognition using covariance of shape and motion. Computer Vision and Image Understanding: CVIU, 129(??): 15–26, December 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001805.

Kehtarnavaz:1995:EAC

[KS95] N. D. Kehtarnavaz and W. Sohn. Error analysis of camera movements in stereo vehicle tracking systems. Computer Vision and Image Understanding: CVIU, 62(3):347– 359, November 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1059/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1059/production/pdf.

Kimia:1996:GHE

[KS96] Benjamin B. Kimia and Kaleem Siddiqi. Geometric heat equation and nonlinear diffusion of shapes and images. Computer Vision and Image Understanding: CVIU, 64(3):305– 322, November 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0062/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0062/production/pdf.

Karacal:2003:RDS

[KS03] Bilge Karaçal and Wesley Snyder. Reconstructing discontinuous surfaces from a given gradient field using partial integrability. Computer Vision and Image Understanding: CVIU, 92(1): 78–111, October 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Krivic:2004:PLO

[KS04] Jaka Krivic and Franc Solina. Part-level object recognition using superquadrics. Computer Vision and Image Understanding: CVIU, 95(1):105– 126, July 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kompella:2012:CRB

[KS12] Varun Raj Kompella and Peter Sturm. Collectivereward based approach for detection of semitransparent objects in single images. Computer Vision and Image Understanding: CVIU, 116(4):484– 499, April 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002591.

Kalinin:2015:GBA

[KS15] Pavel Kalinin and Aleksandr Sirota. A graph based approach to hierarchical image over-segmentation. Computer Vision and Image Understanding: CVIU, 130 (??):80–86, January 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314214001891.

Kong:2016:LHK

[KSF16] Yu Kong, Behnam Satarboroujeni, and Yun Fu. Learning hierarchical 3D kernel descriptors for RGB-D action recognition. Computer Vision and Image Understanding: CVIU, 144(??):14– 23, March 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002118.

Konyushkova:2019:GAL

[KSF19] Ksenia Konyushkova, Raphael Sznitman, and Pascal Fua. Geometry in active learning for binary and multiclass image segmentation. Computer Vision and Image Understanding: CVIU, 182 (??):1–16, May 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421930013X.

Kurtek:2013:SAM

[KSG^{+}13] Sebastian Kurtek, Jingyong Su, Cindy Grimm, Michelle Vaughan, Ross Sowell, and Anuj Srivastava. Statistical analysis of manual segmentations of structures in medical images. Computer Vision and Image Understanding: CVIU, 117(9):1036– 1050, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000714.

Keisler:2019:VSB

[KSG^{+}19] Ryan Keisler, Samuel W. Skillman, Sunny Gonnabathula, Justin Poehnelt, Xander Rudelis, and Michael S. Warren. Visual search over billions of aerial and satellite images. Computer Vision and Image Understanding: CVIU, 187(??):Article 102790, October 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1077314219301067.

Kruger:2020:RPC

[KSHE20] Julia Krüger, Sandra Schultz, Heinz Handels, and Jan Ehrhardt. Registration with probabilistic correspondences — accurate and robust registration for pathological and inhomogeneous medical data. Computer Vision and Image Understanding: CVIU, 190 (??):Article 102839, January 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218301899.

Kise:1998:SPI

[KSI98] Koichi Kise, Akinori Sato, and Motoi Iwata. Segmentation of page images using the area Voronoi diagram. Computer Vision and Image Understanding: CVIU, 70(3): 370–382, June 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0684/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0684/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0684/production/ref.

Kimmel:1995:SDM

[KSKB95] Ron Kimmel, Doron Shaked, Nahum Kiryati, and Alfred M. Bruckstein. Skeletonization via distance maps and level sets. Computer Vision and Image Understanding: CVIU, 62(3):382– 391, November 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1062/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1062/production/pdf.

Kim:2020:TCS

[KSL^{+}20] Eu Young Kim, Seung Yeon Shin, Soochahn Lee, Kyong Joon Lee, Kyoung Ho Lee, and Kyoung Mu Lee. Triplanar convolution with shared 2D kernels for 3D classification and shape retrieval. Computer Vision and Image Understanding: CVIU, 193(??):Article 102901, April 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219301791.

Khanloo:2012:LMF

[KSR^{+}12] Bahman Yari Saeed Khanloo, Ferdinand Stefanus, Mani Ranjbar, Ze-Nian Li, Nicolas Saunier, Tarek Sayed, and Greg Mori. A large margin framework for single camera offline tracking with hybrid cues. Computer Vision and Image Understanding: CVIU, 116(6): 676–689, June 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000215.

Kappes:2016:HOS

[KSRS16] Jörg Hendrik Kappes, Markus Speth, Gerhard Reinelt, and Christoph Schnörr. Higherorder segmentation via multicuts. Computer Vision and Image Understanding: CVIU, 143(??):104– 119, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002490.

Kang:1997:PFT

[KSS97] Sing Bing Kang, Richard Szeliski, and Heung-Yeung Shum. A parallel feature tracker for extended image sequences. Computer Vision and Image Understanding: CVIU, 67(3):296– 310, September 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0519/production;   http:
//www.idealibrary.com/

                                        links/artid/cviu.1996.
                                        0519/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0519/production/ref.

Kume:2015:BAU

[KSY15] Hideyuki Kume, Tomokazu Sato, and Naokazu Yokoya. Bundle adjustment using aerial images with twostage geometric verification. Computer Vision and Image Understanding: CVIU, 138(??):74–84, September 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421500096X.

Krotosky:2007:MIB

[KT07] Stephen J. Krotosky and Mohan M. Trivedi. Mutual information based registration of multimodal stereo videos for person tracking. Computer Vision and Image Understanding: CVIU, 106 (2–3):270–287, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kohli:2008:MUG

[KT08] Pushmeet Kohli and Philip H. S. Torr. Measuring uncertainty in graph cut solutions. Computer Vision and Image Understanding: CVIU, 112(1):30–38, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kalyoncu:2015:GLC

[KT15] Cem Kalyoncu and Önsen Toygar. Geometric leaf classification. Computer Vision and Image Understanding: CVIU, 133(??): 102–109, April 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002124.

Kakadiaris:2017:FRP

[KTE^{+}17] Ioannis A. Kakadiaris, George Toderici, Georgios Evangelopoulos, Georgios Passalis, Dat Chu, Xi Zhao, Shishir K. Shah, and Theoharis Theoharis. 3D–2D face recognition with pose and illumination normalization. Computer Vision and Image Understanding: CVIU, 154(??):137– 151, January 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300480.

Komodakis:2008:PVC

[KTP08] Nikos Komodakis, Georgios Tziritas, and Nikos Paragios. Performance vs computational efficiency for optimizing single and dynamic MRFs: Setting the state of the art with primal-dual strategies. Computer Vision and Image Understanding: CVIU, 112(1):14–29, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kwon:2017:LOU

[KTV17] Junseok Kwon, Radu Timofte, and Luc Van Gool. Leveraging observation uncertainty for robust visual tracking. Computer Vision and Image Understanding: CVIU, 158(??):62– 71, May 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217300310.

Khan:2019:SAV

[KU19] Sultan Daud Khan and Habib Ullah. A survey of advances in vision-based vehicle re-identification. Computer Vision and Image Understanding: CVIU, 182 (??):50–63, May 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421930030X.

Kawana:2018:ECN

[KUHY18] Yuki Kawana, Norimichi Ukita, Jia-Bin Huang, and Ming-Hsuan Yang. Ensemble convolutional neural networks for pose estimation. Computer Vision and Image Understanding: CVIU, 169 (??):62–74, April 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217302308.

Kuijper:2008:EES

[Kui08] Arjan Kuijper. Exploring and exploiting the structure of saddle points in Gaussian scale space. Computer Vision and Image Understanding: CVIU, 112(3):337–349, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kehl:2006:MTC

[KV06] Roland Kehl and Luc Van Gool. Markerless tracking of complex human motions from multiple views. Computer Vision and Image Understanding: CVIU, 104 (2–3):190–209, November/ December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Koster:1997:HLM

[KVdG^{+}97] André S. E. Koster, Koen L.

Vincken, Cornelis N. de Graaf, Olaf C. Zander, and Max A. Viergever. Heuristic linking models in multiscale image segmentation. Computer Vision and Image Understanding: CVIU, 65(3): 382–402, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0490/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0490/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0490/production/ref.

Kang:1999:CEC

[KW99] Sing Bing Kang and Richard Weiss. Characterization of errors in compositing panoramic images. Computer Vision and Image Understanding: CVIU, 73(2): 269–280, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0727/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0727/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0727/production/ref.

Keller:2000:FRB

[KW00] James M. Keller and Xiaomei Wang. A fuzzy rulebased approach to scene description involving spatial relationships. Computer Vision and Image Understanding: CVIU, 80 (1):21–41, October 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0872;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0872/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0872/ref.

Koppen:2012:BRM

[KW12] W. P. Koppen and M. Worring. Backtracking: Retrospective multi-target tracking. Computer Vision and Image Understanding: CVIU, 116(9):967–980, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212000653.

Koren:2006:ASE

[KY06] Raz Koren and Yitzhak Yitzhaky. Automatic selection of edge detector parameters based on spatial and statistical measures. Computer Vision and Image Understanding: CVIU, 102(2): 204–213, May 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kupcu:2019:DBI

[KY19] Emel Küpçü and Yücel Yemez. Diffusion-based isometric depth correspondence. Computer Vision and Image Understanding: CVIU, 189(??):Article 102808, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314219301195.

Koniusz:2013:CML

[KYM13] Piotr Koniusz, Fei Yan, and Krystian Mikolajczyk. Comparison of mid-level feature coding approaches and pooling strategies in visual concept detection. Computer Vision and Image Understanding: CVIU, 117(5): 479–492, May 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001725.

Kim:2014:STW

[KYYC14] Soo Wan Kim, Shimin Yin, Kimin Yun, and Jin Young Choi. Spatio-temporal weighting in local patches for direct estimation of camera motion in video stabilization. Computer Vision and Image Understanding: CVIU, 118(??):71–83, January 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001756.

Kosecka:2005:EMP

[KZ05] Jana Kosecká and Wei Zhang. Extraction, matching, and pose recovery based on dominant rectangular structures. Computer Vision and Image Understanding: CVIU, 100(3):274–293, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Klette:2012:ASD

[KŽ12] Reinhard Klette and Joviša Žunić. ADR shape descriptor — distance between shape centroids versus shape diameter. Computer Vision and Image Understanding: CVIU, 116(6): 690–697, June 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000331.

Lourakis:2005:ECC

[LA05] Manolis I. A. Lourakis and Antonis A. Argyros. Efficient, causal camera tracking in unprepared environments. Computer Vision and Image Understanding: CVIU, 99(2):259–290, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Law:2011:SVM

[LA11] Alvin J. Law and Daniel G. Aliaga. Single viewpoint model completion of symmetric objects for digital inspection. Computer Vision and Image Understanding: CVIU, 115(5):603– 610, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Leizea:2015:RTN

[LÁB15] Ibai Leizea, Hugo Álvarez, and Diego Borro. Real time non-rigid 3D surface tracking using particle filter. Computer Vision and Image Understanding: CVIU, 133 (??):51–65, April 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214002409.

Lesage:2016:APF

[LAFLB16] David Lesage, Elsa D. Angelini, Gareth Funka-Lea, and Isabelle Bloch. Adaptive particle filtering for coronary artery segmentation from 3D CT angiograms. Computer Vision and Image Understanding: CVIU, 151(??):29–46, October 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002532.

Lai:2000:RIM

[Lai00] Shang-Hong Lai. Robust image matching under partial occlusion and spatially varying illumination change. Computer Vision and Image Understanding: CVIU, 78(1):84–98, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0829;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0829/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0829/ref.

Lu:2010:CBO

[LAL^{+}10] ChengEn Lu, Nagesh Adluru, Haibin Ling, Guangxi Zhu, and Longin Jan Latecki. Contour based object detection using part bundles. Computer Vision and Image Understanding: CVIU, 114(7):827– 834, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Laurentini:1997:HMS

[Lau97] Aldo Laurentini. How many 2D silhouettes does it take to reconstruct a 3D object?

Computer Vision and Image Understanding: CVIU, 67 (1):81–87, July 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0508/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0508/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0508/production/ref.

Li:1998:MGP

[LB98] Fuxing Li and Michael Brady. Modeling the ground plane transformation for real-time obstacle detection. Computer Vision and Image Understanding: CVIU, 71 (1):137–152, July 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0645/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0645/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0645/production/ref.

Leonardis:2000:RRU

[LB00] Ale Leonardis and Horst Bischof. Robust recognition using eigenimages. Computer Vision and Image Understanding: CVIU, 78(1):99–118, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0830;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0830/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0830/ref.

Levine:2005:DRS

[LB05] Martin D. Levine and Jisnu Bhattacharyya. Detecting and removing specularities in facial images. Computer Vision and Image Understanding: CVIU, 100(3): 330–356, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:2008:TFT

[LB08] Sang-Chul Lee and Peter Bajcsy. Trajectory fusion for three-dimensional volume reconstruction. Computer Vision and Image Understanding: CVIU, 110(1): 19–31, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lim:2010:EEU

[LB10] John Lim and Nick Barnes.

Estimation of the epipole using optical flow at antipodal points. Computer Vision and Image Understanding: CVIU, 114(2):245–253, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Laurentini:2014:CAF

[LB14] Aldo Laurentini and Andrea Bottino. Computer analysis of face beauty: a survey. Computer Vision and Image Understanding: CVIU, 125(??):184– 199, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214000939.

Lekic:2019:ARC

[LB19] Vladimir Lekic and Zdenka Babic. Automotive radar and camera fusion using Generative Adversarial Networks. Computer Vision and Image Understanding: CVIU, 184(??): 1–8, July 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219300530.

Landi:2021:MAN

[LBC^{+}21] Federico Landi, Lorenzo Baraldi, Marcella Cornia, Massimiliano Corsini, and Rita Cucchiara. Multimodal attention networks for lowlevel vision-and-language navigation. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000990.

Loza:2010:NGM

[LBCA10] Artur Loza, David Bull, Nishan Canagarajah, and Alin Achim. Non-Gaussian model-based fusion of noisy images in the wavelet domain. Computer Vision and Image Understanding: CVIU, 114(1):54–65, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Larsen:2010:SIT

[LBK10] Rasmus Larsen, Erhardt Barth, and Andreas Kolb. Special issue on Time-of-Flight camera based computer vision. Computer Vision and Image Understanding: CVIU, 114(12):1317, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Loss:2009:IMS

[LBNS09] Leandro Loss, George Bebis, Mircea Nicolescu, and Alexei Skurikhin. An iterative multi-scale tensor voting scheme for perceptual grouping of natural shapes in cluttered backgrounds. Computer Vision and Image Understanding: CVIU, 113(1):126–149, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:2023:USH

[LBP23] Wonju Lee, Seok-Yong Byun, and Minje Park. Unsupervised soft-to-hard hashing with contrastive learning. Computer Vision and Image Understanding: CVIU, 233(??): ??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223000930.

Liu:2009:VRB

[LC09] David Liu and Tsuhan Chen. Video retrieval based on object discovery. Computer Vision and Image Understanding: CVIU, 113(3): 397–404, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lai:2011:FII

[LC11] Rongjie Lai and Tony F. Chan. A framework for intrinsic image processing on surfaces. Computer Vision and Image Understanding: CVIU, 115(12):1647– 1661, December 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001603.

Lee:2014:RTF

[LC14] Minsik Lee and Chong-Ho Choi. Real-time facial shape recovery from a single image under general, unknown lighting by rank relaxation. Computer Vision and Image Understanding: CVIU, 120(??): 59–69, March 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213002415.

Loh:2019:GKL

[LC19] Yuen Peng Loh and Chee Seng Chan. Getting to know low-light images with the Exclusively Dark dataset. Computer Vision and Image Understanding: CVIU, 178 (??):30–42, January 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314218304296.

Langrana:1997:FIV

[LCD97] Noshir A. Langrana, Yuan Chen, and Atish K. Das. Feature identification from vectorized mechanical drawings. Computer Vision and Image Understanding: CVIU, 68(2):127–145, November 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0548/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0548/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0548/production/ref.

Luo:2021:RSI

[LCG21] Minnan Luo, Xiaojun Chang, and Chen Gong. Reliable shot identification for complex event detection via visual-semantic embedding. Computer Vision and Image Understanding: CVIU, 213(??):??, December 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221001442.

Lu:2024:IRR

[LCG^{+}24] Jiale Lu, Lianggangxu Chen, Haoyue Guan, Shaohui Lin, Chunhua Gu, Changbo Wang, and Gaoqi He. Improving rare relation inferring for scene graph generation using bipartite graph network. Computer Vision and Image Understanding: CVIU, 239(??):

??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002813.

Liu:2014:SOV

[LCL^{+}14] Shaowei Liu, Peng Cui, Huanbo Luan, Wenwu Zhu, Shiqiang Yang, and Qi Tian. Social-oriented visual image search. Computer Vision and Image Understanding: CVIU, 118(??):30–39, January 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001604.

Luo:2017:SCC

[LCL^{+}17] Minnan Luo, Xiaojun Chang, Zhihui Li, Liqiang Nie, Alexander G. Hauptmann, and Qinghua Zheng. Simple to complex cross-modal learning to rank. Computer Vision and Image Understanding: CVIU, 163(??): 67–77, October 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301339.

Lu:2018:LCN

[LCLH18] Keyu Lu, Jianhui Chen, James J. Little, and Hangen He. Lightweight convolutional neural networks for player detection and classification. Computer Vision and Image Understanding: CVIU, 172(??):77– 87, July 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218300341.

Lee:2013:THR

[LCP13] Chan-Su Lee, SungYong Chun, and Shin Won Park. Tracking hand rotation and various grasping gestures from an IR camera using extended cylindrical manifold embedding. Computer Vision and Image Understanding: CVIU, 117(12):1711– 1723, December 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001616.

Li:2021:TTT

[LCS^{+}21] Yuezun Li, Ming-Ching Chang, Pu Sun, Honggang Qi, Junyu Dong, and Siwei Lyu. TransRPN: Towards the transferable adversarial perturbations using region proposal networks and beyond. Computer Vision and Image Understanding: CVIU, 213(??):??, December 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221001466.

Laptev:2007:LVA

[LCSL07] Ivan Laptev, Barbara Caputo, Christian Schüldt, and Tony Lindeberg. Local velocity-adapted motion events for spatio-temporal recognition. Computer Vision and Image Understanding: CVIU, 108(3):207–229, December 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lanman:2009:SSL

[LCT09] Douglas Lanman, Daniel Crispell, and Gabriel Taubin. Surround structured lighting: 3-D scanning with orthographic illumination. Computer Vision and Image Understanding: CVIU, 113(11):1107–1117, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2001:EEF

[LCZ^{+}01] B. Li, R. Chellappa, Q. Zheng, S. Der, N. Nasrabadi, L. Chan, and L. Wang. Experimental evaluation of FLIR ATR approaches — a comparative study. Computer Vision and Image Understanding: CVIU, 84 (1):5–24, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2001.0938;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0938/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0938/ref.

Li:2009:FCS

[LCZ09] Shu-Xiao Li, Hong-Xing Chang, and Cheng-Fei Zhu. Fast curvilinear structure extraction and delineation using density estimation. Computer Vision and Image Understanding: CVIU, 113(6):763–775, June 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2016:VCW

[LCZ^{+}16] Jingjing Liu, Chao Chen, Yan Zhu, Wei Liu, and Dimitris N. Metaxas. Video classification via weakly supervised sequence modeling. Computer Vision and Image Understanding: CVIU, 152(??):79–87, November 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215002325.

Liu:1998:GIL

[LD98] Wenyin Liu and Dov Dori. A generic integrated line detection algorithm and its object-process specification. Computer Vision and Image Understanding: CVIU, 70(3): 420–437, June 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0683/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0683/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0683/production/ref.

Li:2024:TTE

[LD24] Zhenyu Li and Zhenbiao Dong. TECD Attention: Texture-enhanced and crossdomain attention modeling for visual place recognition. Computer Vision and Image Understanding: CVIU, 240 (??):??, March 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000109.

Liu:2013:MRV

[LDC^{+}13] Ningning Liu, Emmanuel Dellandréa, Liming Chen, Chao Zhu, Yu Zhang, Charles-Edmond Bichot, Stéphane Bres, and Bruno Tellez. Multimodal recognition of visual concepts using histograms of textual concepts and selective weighted late fusion scheme. Computer Vision and Image Understanding: CVIU, 117(5): 493–512, May 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001695.

Liang:2024:SBS

[LDCX24] Shuang Liang, Weidong Dai, Yiyang Cai, and Chi Xie. Sketch-based 3D shape retrieval via teacher-student learning. Computer Vision and Image Understanding: CVIU, 239(??): ??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223002837.

Liang:2009:MCC

[LDD09] Jian Liang, Daniel DeMenthon, and David Doermann. Mosaicing of cameracaptured document images. Computer Vision and Image Understanding: CVIU, 113(4):572–579, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lefevre:2013:EEF

[LDGS^{+}13] Thierry Lefevre, Bernadette Dorizzi, Sonia Garcia-Salicetti, Nadege Lemperiere, and Stephane Belardi. Effective elliptic fitting for iris normalization. Computer Vision and Image Understanding: CVIU, 117(6):732– 745, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421300012X.

Lin:2014:ITC

[LDH^{+}14] Zijia Lin, Guiguang Ding, Mingqing Hu, Yunzhen Lin, and Shuzhi Sam Ge. Image tag completion via dualview linear sparse reconstructions. Computer Vision and Image Understanding: CVIU, 124(??):42– 60, July 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214000782.

Li:2015:EMF

[LDH^{+}15] Huibin Li, Huaxiong Ding, Di Huang, Yunhong Wang, Xi Zhao, Jean-Marie Morvan, and Liming Chen. An efficient multimodal 2D + 3D feature-based approach to automatic facial expression recognition. Computer Vision and Image Understanding: CVIU, 140(??): 83–92, November 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001587.

Lu:2019:ASE

[LDL^{+}19] Xuequan Lu, Zhigang Deng, Jun Luo, Wenzhi Chen, Sai-Kit Yeung, and Ying He. 3D articulated skeleton extraction using a single consumer-grade depth camera. Computer Vision and Image Understanding: CVIU, 188(??):Article 102792, November 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314218303606.

Lavest:1997:IRZ

[LDPD97] J. M. Lavest, C. Delherm, B. Peuchot, and N. Daucher. Implicit reconstruction by zooming. Computer Vision and Image Understanding: CVIU, 66(3): 301–315, June 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0511/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0511/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0511/production/ref.

Luo:2021:BBN

[LDT21] Yao Luo, Zhong-Hui Duan, and Jinhui Tang. Bi-branch network for dynamic scene deblurring. Computer Vision and Image Understanding: CVIU, 202(??):Article 103100, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301272.

Lee:2009:DSO

[LE09] Chan-Su Lee and Ahmed Elgammal. Dynamic shape outlier detection for human locomotion. Computer Vision and Image Understanding: CVIU, 113(3):332– 344, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lmaati:2010:DSE

[LEA^{+}10] Elmustapha Ait Lmaati, Ahmed El Oirrak, Driss Aboutajdine, Mohamed Daoudi, and Mohammed Najib Kaddioui. A 3-D Search engine based on Fourier series. Computer Vision and Image Understanding: CVIU, 114(1):1–7, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lezoray:2007:GRC

[LEB07] Olivier Lezoray, Abderrahim Elmoataz, and Sébastien Bougleux. Graph regularization for color image processing. Computer Vision and Image Understanding: CVIU, 107(1–2):38–55, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:2002:MVS

[Lee02] Joon Woong Lee. A machine vision system for lanedeparture detection. Computer Vision and Image Understanding: CVIU, 86(1): 52–78, April 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Laghrib:2018:SDD

[LEE^{+}18] Amine Laghrib, Mahmoud Ezzaki, Mohammed El Rhabi, Abdelilah Hakim, Pascal Monasse, and Said Raghay. Simultaneous deconvolution and denoising using a second order variational approach applied to image super resolution. Computer Vision and Image Understanding: CVIU, 168 (??):50–63, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421730142X.

Latecki:1995:WCS

[LER95] Longin Latecki, Ulrich Eckhardt, and Azriel Rosenfeld. Well-composed sets. Computer Vision and Image Understanding: CVIU, 61(1): 70–83, January 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1006/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1006/production/pdf.

Lejeune:1996:FPO

[LF96] André Lejeune and Frank P.

Ferrie. Finding the parts of objects in range images. Computer Vision and Image Understanding: CVIU, 64(2):230–247, September 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0056/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0056/production/pdf.

Luong:1998:DEU

[LF98] Q.-T. Luong and O. D.

Faugeras. On the determination of epipoles using cross-ratios. Computer Vision and Image Understanding: CVIU, 71 (1):1–18, July 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.

                                        0621/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0621/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0621/production/ref.

Lagger:2008:RML

[LF08] Pascal Lagger and Pascal Fua. Retrieving multiple light sources in the presence of specular reflections and texture. Computer Vision and Image Understanding: CVIU, 111(2):207–218, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2008:FBM

[LFL08] Zheng Liu, David S. Forsyth, and Robert Laganière. A feature-based metric for the quantitative evaluation of pixel-level image fusion. Computer Vision and Image Understanding: CVIU, 109(1):56–68, January 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2023:FRM

[LFLZ23] Yuewen Li, Wenquan Feng, Shuchang Lyu, and Qi Zhao. Feature reconstruction and metric based network for few-shot object detection. Computer Vision and Image Understanding: CVIU, 227 (??):??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001783.

Lee:2013:SAB

[LFMP13] Yooyoung Lee, James J. Filliben, Ross J. Micheals, and P. Jonathon Phillips. Sensitivity analysis for biometric systems: a methodology based on orthogonal experiment designs. Computer Vision and Image Understanding: CVIU, 117(5): 532–550, May 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000040.

Liem:2014:JMP

[LG14] Martijn C. Liem and Dariu M. Gavrila. Joint multi-person detection and tracking from overlapping cameras. Computer Vision and Image Understanding: CVIU, 128(??):36–50, November 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001301.

Li:2017:PRE

[LG17] Yingmao Li and Nicholas R.

Gans. Predictive RANSAC: Effective model fitting and tracking approach under heavy noise and outliers. Computer Vision and Image Understanding: CVIU, 161 (??):99–113, August 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S107731421730108X.

Lisanti:2024:CDM

[LG24] Giuseppe Lisanti and Nico Giambi. Conditioning diffusion models via attributes and semantic masks for face generation. Computer Vision and Image Understanding: CVIU, 244(??): ??, July 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224001073.

Liu:2016:LHO

[LGD16] Ni Liu, Georgy Gimel’farb, and Patrice Delmas. Learnable high-order MGRF models for contrast-invariant texture recognition. Computer Vision and Image Understanding: CVIU, 143(??):135–146, February 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314215001393.

Li:2018:VCA

[LGG^{+}18] Zhenyang Li, Kirill Gavrilyuk, Efstratios Gavves, Mihir Jain, and Cees G. M. Snoek. VideoL-STM convolves, attends and flows for action recognition. Computer Vision and Image Understanding: CVIU, 166(??):41–50, January 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301741.

Liu:2015:CCC

[LGL15] Yang Liu, Theo Gevers, and Xueqing Li. Color constancy by combining low–mid–high level image cues. Computer Vision and Image Understanding: CVIU, 140(??): 1–8, November 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001241.

Lavalle:1995:FCP

[LH95] Steven M. Lavalle and Seth A. Hutchinson. Framework for constructing probability distributions on the space of image segmentations. Computer Vision and Image Understanding: CVIU, 61(2):203– 230, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/

                                        links/artid/cviu.1995.
                                        1016/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1016/production/pdf.

Lee:1999:OCC

[LH99] Yu-Hua Lee and Shi-Jinn Horng. Optimal computing the chessboard distance transform on parallel processing systems. Computer Vision and Image Understanding: CVIU, 73(3):374– 390, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0741/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0741/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0741/production/ref.

Luo:2003:UFA

[LH03] Bin Luo and E. R. Hancock. A unified framework for alignment and correspondence. Computer Vision and Image Understanding: CVIU, 92(1):26–55, October 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2023:LGC

[LHG^{+}23] Haiyan Li, Yongqiang Han, Lei Guo, Mingchuan Tan, and Liping Zhou. LKDA-GAN: Cross-modality image synthesis via Generative Adversarial Network aggregating large kernel decomposable attention bottleneck block. Computer Vision and Image Understanding: CVIU, 237(??): ??, December 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002369.

Little:1997:DIR

[LHH97] J. A. Little, D. L. G.

Hill, and D. J. Hawkes. Deformations incorporating rigid structures. Computer Vision and Image Understanding: CVIU, 66(2): 223–232, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0608/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0608/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0608/production/ref.

Liu:1998:AVE

[LHH^{+}98] Hongche Liu, Tsai-Hong Hong, Martin Herman, Ted Camus, and Rama Chellappa. Accuracy vs efficiency trade-offs in optical flow algorithms. Computer Vision and Image Understanding: CVIU, 72(3):271– 286, December 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0675/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0675/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0675/production/ref.

Liu:1998:MMB

[LHHC98] Hongche Liu, Tsai-Hong Hong, Martin Herman, and Rama Chellappa. Motionmodel-based boundary extraction and a real-time implementation. Computer Vision and Image Understanding: CVIU, 70(1):87– 100, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0625/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0625/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0625/production/ref.

Liu:2009:FFS

[LHJ^{+}09] Chunxi Liu, Qingming Huang, Shuqiang Jiang, Liyuan Xing, Qixiang Ye, and Wen Gao. A framework for flexible summarization of racquet sports video using multiple modalities. Computer Vision and Image Understanding: CVIU, 113 (3):415–424, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:1997:PCE

[LHKC97] Yu-Hua Lee, Shi-Jinn Horng, Tzong-Wann Kao, and Yuung-Jih Chen. Parallel computation of the Euclidean distance transform on the mesh of trees and the hypercube computer. Computer Vision and Image Understanding: CVIU, 68(1): 109–119, October 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0539/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0539/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0539/production/ref.

Li:2021:LTL

[LHL^{+}21] Xinzhe Li, Jianqiang Huang, Yaoyao Liu, Qin Zhou, Shibao Zheng, Bernt Schiele, and Qianru Sun. Learning to teach and learn for semisupervised few-shot image classification. Computer Vision and Image Understanding: CVIU, 212(??): ??, November 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221001144.

Li:2023:ACP

[LHLZ23] Gang Li, Ping He, Huibin Li, and Fan Zhang. Adversarial composite prediction of normal video dynamics for anomaly detection. Computer Vision and Image Understanding: CVIU, 232(??):??, July 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223000668.

Li:2015:FHS

[LHSG15] Yi Li, Timothy M. Hospedales, Yi-Zhe Song, and Shaogang Gong. Free-hand sketch recognition by multi-kernel feature learning. Computer Vision and Image Understanding: CVIU, 137(??): 1–11, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215000375.

Lhuillier:2008:ASS

[Lhu08] Maxime Lhuillier. Automatic scene structure and camera motion using a catadioptric system. Computer Vision and Image Understanding: CVIU, 109(2): 186–203, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lhuillier:2018:SRS

[Lhu18]  Maxime Lhuillier.  Sur-

face reconstruction from a sparse point cloud by enforcing visibility consistency and topology constraints. Computer Vision and Image Understanding: CVIU, 175(??):52–71, October 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218302959.

Lhuillier:2023:EVD

[Lhu23] Maxime Lhuillier. Estimating the vertical direction in a photogrammetric 3D model, with application to visualization. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001947.

Lee:2014:QEB

[LHY14] Youngwoon Lee, Jae-Pil Heo, and Sung-Eui Yoon.

Quadra-embedding: Binary code embedding with low quantization error. Computer Vision and Image Understanding: CVIU, 125 (??):214–222, August 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314214000940.

Lee:2005:VTR

[LHYK05] Kuang-Chih Lee, Jeffrey Ho, Ming-Hsuan Yang, and David Kriegman. Visual tracking and recognition using probabilistic appearance manifolds. Computer Vision and Image Understanding: CVIU, 99(3):303–331, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liao:2019:RAU

[LHZY19] Zhongke Liao, Haifeng Hu, Junxuan Zhang, and Chang Yin. Residual attention unit for action recognition. Computer Vision and Image Understanding: CVIU, 189(??):Article 102821, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219301298.

Lillekjendlie:1997:CAF

[Lil97] Bjørn Lillekjendlie. Circular arcs fitted on a Riemann sphere. Computer Vision and Image Understanding: CVIU, 67(3):311– 317, September 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0529/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0529/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0529/production/ref.

Lingrand:2002:ESP

[Lin02] Diane Lingrand. An exhaustive study of particular cases leading to robust and accurate motion estimation. Computer Vision and Image Understanding: CVIU, 85(3):159–188, March 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2010:FFS

[Liu10] Yonghuai Liu. Free form shape registration using the barrier method. Computer Vision and Image Understanding: CVIU, 114(9): 1004–1016, September 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2023:MOT

[LJC^{+}23] Yi-Fan Li, Hong-Bing Ji, Xi Chen, Yu-Kun Lai, and Yong-Liang Yang. Multiobject tracking with robust object regression and association. Computer Vision and Image Understanding: CVIU, 227(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001643.

Li:2024:LKL

[LJC^{+}24] Yi-Fan Li, Hong-Bing Ji, Xi Chen, Yong-Liang Yang, and Yu-Kun Lai. Learning key lines for multiobject tracking. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000547.

Lee:2007:VAB

[LJHH07] Chang-Ock Lee, Kiwan Jeon, Youngsoo Ha, and Jooyoung Hahn. A variational approach to blending based on warping for nonoverlapped images. Computer Vision and Image Understanding: CVIU, 105 (2):112–120, February 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lin:2024:CLB

[LJJY24] Shaohui Lin, Bo Ji, Rongrong Ji, and Angela Yao. A closer look at branch classifiers of multi-exit architectures. Computer Vision and Image Understanding: CVIU, 239(??): ??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002801.

Li:2018:CBC

[LJZ18] Hui Li, Xixi Jia, and Lei Zhang. Clustering based content and color adaptive tone mapping. Computer Vision and Image Understanding: CVIU, 168(??): 37–49, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301789.

Lee:1997:SSG

[LK97] Kyoung Mu Lee and C.- C. Jay Kuo. Shape from shading with a generalized reflectance map model. Computer Vision and Image Understanding: CVIU, 67(2):143– 160, August 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.

0522/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0522/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0522/production/ref.

Lorenz:2000:GPB

[LK00] Cristian Lorenz and Nils Krahnstöver. Generation of point-based 3D statistical shape models for anatomical objects. Computer Vision and Image Understanding: CVIU, 77(2): 175–191, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0814;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0814/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0814/ref.

Liu:2003:FVS

[LK03] Yan Liu and John R. Kender. Fast video segment retrieval by Sort-Merge feature selection, boundary refinement, and lazy evaluation. Computer Vision and Image Understanding: CVIU, 92(2–3):147– 175, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lu:2000:SST

[LKK00] Rong Lu, Jan J. Koenderink, and Astrid M. L. Kappers. Specularities on surfaces with tangential hairs or grooves. Computer Vision and Image Understanding: CVIU, 78 (3):320–335, June 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0841;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0841/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0841/ref.

Li:2020:ECC

[LKZ20] He Li, Weihang Kong, and Shihui Zhang. Effective crowd counting using multi-resolution context and image quality assessmentguided training. Computer Vision and Image Understanding: CVIU, 201 (??):Article 103065, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220301065.

Lin:1997:DSU

[LL97a] Stephen Lin and Sang Wook Lee. Detection of specularity using stereo in color and polarization space. Computer Vision and Image Understanding: CVIU, 65(2): 336–346, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0577/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0577/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0577/production/ref.

Lindeberg:1997:SCE

[LL97b] Tony Lindeberg and Meng-Xiang Li. Segmentation and classification of edges using minimum description length approximation and complementary junction cues. Computer Vision and Image Understanding: CVIU, 67(1):88–98, July 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0510/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0510/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0510/production/ref.

Latecki:1999:CRS

[LL99] Longin Jan Latecki and Rolf Lakämper. Convexity rule for shape decomposition based on discrete contour evolution. Computer Vision and Image Understanding: CVIU, 73(3): 441–454, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0738/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0738/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0738/production/ref.

Leow:2004:AAA

[LL04] Wee Kheng Leow and Rui Li. The analysis and applications of adaptive-binning color histograms. Computer Vision and Image Understanding: CVIU, 94(1– 3):67–91, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2008:FCD

[LL08] Zhiming Liu and Chengjun Liu. Fusion of the complementary Discrete Cosine Features in the YIQ color space for face recognition. Computer Vision and Image Understanding: CVIU, 111(3):249–262, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Linde:2012:CCC

[LL12] Oskar Linde and Tony Lindeberg. Composed complexcue histograms: An investigation of the information content in receptive field based image descriptors for object recognition. Computer Vision and Image Understanding: CVIU, 116(4): 538–560, April 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211002633.

Presti:2017:BHM

[LL17] Liliana Lo Presti and Marco La Cascia. Boosting Hankel matrices for face emotion recognition and pain detection. Computer Vision and Image Understanding: CVIU, 156(??):19– 33, March 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421630162X.

Liu:2011:TBL

[LLC11] Yang Liu, Yan Liu, and Keith C. C. Chan. Tensorbased locally maximum margin classifier for image and video classification. Computer Vision and Image Understanding: CVIU, 115 (3):300–309, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2012:ORU

[LLC12] Ying-Ho Liu, Anthony J. T.

Lee, and Fu Chang. Object recognition using discriminative parts. Computer Vision and Image Understanding: CVIU, 116(7): 854–867, July 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000604.

Li:2013:HER

[LLC13] Yujian Li, Houjun Li, and Zhi Cai. Human eyebrow recognition in the matching-recognizing framework. Computer Vision and Image Understanding: CVIU, 117(2):170–181, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001452.

Li:2021:PRI

[LLCY21] Zhiyong Li, Jingyi Lv, Ying Chen, and Jin Yuan. Person re-identification with part prediction alignment. Computer Vision and Image Understanding: CVIU, 205 (??):Article 103172, April 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314221000163.

Liu:2009:ICV

[LLE^{+}09] Qingshan Liu, Xuelong Li, Ahmed Elgammal, Xian sheng Hua, Dong Xu, and Dacheng Tao. Introduction to Computer Vision and Image Understanding: the special issue on video analysis. Computer Vision and Image Understanding: CVIU, 113 (3):317–318, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2018:MNL

[LLF18] Yawei Li, Xiaofeng Li, and Zhizhong Fu. Modified non-local means for superresolution of hybrid videos. Computer Vision and Image Understanding: CVIU, 168 (??):64–78, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217302023.

Li:2014:CMS

[LLG^{+}14] Bo Li, Yijuan Lu, Afzal Godil, Tobias Schreck, Benjamin Bustos, Alfredo Ferreira, Takahiko Furuya, Manuel J. Fonseca, Henry Johan, Takahiro Matsuda, Ryutarou Ohbuchi, Pedro B. Pascoal, and Jose M. Saavedra. A comparison of methods for sketch-based 3D shape retrieval. Computer Vision and Image Understanding: CVIU, 119 (??):57–80, February 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314213002282.

Lu:2023:SGA

[LLG^{+}23] Junjie Lu, Shengyang Li, Weilong Guo, Manqi Zhao, Jian Yang, Yunfei Liu, and Zhuang Zhou. Siamese Graph Attention Networks for robust visual object tracking. Computer Vision and Image Understanding: CVIU, 229(??): ??, March 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000140.

Li:2024:SGC

[LLG^{+}24] Chuankun Li, Shuai Li, Yanbo Gao, Lijuan Zhou, and Wanqing Li. Static graph convolution with learned temporal and channelwise graph topology generation for skeleton-based action recognition. Computer Vision and Image Understanding: CVIU, 244 (??):??, July 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000936.

Liu:2023:PBB

[LLJ^{+}23] Hao Liu, Ce Li, Shangang Jin, Weizhe Gao, Fenghua Liu, Shaoyi Du, and Shihui Ying. PGF-BIQA: Blind image quality assessment via probability multi-grained cascade forest. Computer Vision and Image Understanding: CVIU, 232(??): ??, July 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223000759.

Li:2013:SPN

[LLL13] Zechao Li, Jing Liu, and Hanqing Lu. Structure preserving non-negative matrix factorization for dimensionality reduction. Computer Vision and Image Understanding: CVIU, 117(9): 1175–1189, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314213000891.

Li:2014:SEN

[LLL^{+}14] Yibao Li, Dongsun Lee, Chaeyoung Lee, Jihu Lee, Sanha Lee, Jisu Kim, Shinwoo Ahn, and Junseok Kim. Surface embedding narrow volume reconstruction from unorganized points. Computer Vision and Image Understanding: CVIU, 121(??): 100–107, April 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000289.

Li:2015:CSR

[LLL^{+}15a] Bo Li, Yijuan Lu, Chunyuan Li, Afzal Godil, Tobias Schreck, Masaki Aono, Martin Burtscher, Qiang Chen, Nihad Karim Chowdhury, Bin Fang, Hongbo Fu, Takahiko Furuya, Haisheng Li, Jianzhuang Liu, Henry Johan, Ryuichi Kosaka, Hitoshi Koyanagi, Ryutarou Ohbuchi, Atsushi Tatsuma, Yajuan Wan, Chaoli Zhang, et al. A comparison of 3D shape retrieval methods based on a large-scale benchmark supporting multimodal queries. Computer Vision and Image Understanding: CVIU, 131(??): 1–27, February 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002100.

Liu:2015:LRD

[LLL15b] Bingyuan Liu, Jing Liu, and Hanqing Lu. Learning representative and discriminative image representation by deep appearance and spatial coding. Computer Vision and Image Understanding: CVIU, 136(??): 23–31, July 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215000521.

Li:2020:MEP

[LLL^{+}20] Ru Li, Shuaicheng Liu, Guanghui Liu, Tiecheng Sun, and Jishun Guo. Multi-exposure photomontage with hand-held cameras. Computer Vision and Image Understanding: CVIU, 193(??):Article 102929, April 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220300199.

Li:2023:MGP

[LLLW23] Xiaobao Li, Qingyong Li, Fengjiao Liang, and Wen Wang. Multi-granularity Pseudo-label Collaboration for unsupervised person reidentification. Computer Vision and Image Understanding: CVIU, 227(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001941.

Le:2020:SNT

[LLNC20] Quyet-Tien Le, Patricia Ladret, Huu-Tuan Nguyen, and Alice Caplier. Study of naturalness in tone-mapped images. Computer Vision and Image Understanding: CVIU, 196(??):Article 102971, July 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300485.

Lai:2018:EAA

[LLNS18] Shang-Hong Lai, Vincent Lepetit, Ko Nishino, and Yoichi Sato. Editorial for ACCV’16 award papers. Computer Vision and Image Understanding: CVIU, 173 (??):46, August 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218304648.

Li:2022:ZSS

[LLNZ22] Jiangtong Li, Zhixin Ling, Li Niu, and Liqing Zhang. Zero-shot sketch-based image retrieval with structureaware asymmetric disentanglement. Computer Vision and Image Understanding: CVIU, 218(??): ??, April 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314222000406.

Li:2016:CNN

[LLP16] Hanxi Li, Yi Li, and Fatih Porikli. Convolutional neural net bagging for online visual tracking. Computer Vision and Image Understanding: CVIU, 153(??):120– 129, December 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300984.

Leichter:2010:MST

[LLR10] Ido Leichter, Michael Lindenbaum, and Ehud Rivlin. Mean Shift tracking with multiple reference color histograms. Computer Vision and Image Understanding: CVIU, 114(3):400– 408, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lazarou:2021:NSM

[LLS21a] Michalis Lazarou, Bo Li, and Tania Stathaki. A novel shape matching descriptor for real-time static hand gesture recognition. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000850.

Li:2021:DAM

[LLS21b] Bo Li, Sam Leroux, and Pieter Simoens. Decoupled appearance and motion learning for efficient anomaly detection in surveillance video. Computer Vision and Image Understanding: CVIU, 210 (??):??, September 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S107731422100093X.

Li:2023:RTS

[LLS^{+}23] Shijie Li, Junmin Liu, Weilin Shen, Jianyong Sun, and Chengli Tan. Robust teacher: Self-correcting pseudo-label-guided semisupervised learning for object detection. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001686.

Lopez:2000:MCB

[LLSV00] Antonio M. López, David Lloret, Joan Serrat, and Juan J. Villanueva. Multilocal creaseness based on the level-set extrinsic curvature.

Computer Vision and Image Understanding: CVIU, 77 (2):111–144, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0812;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0812/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0812/ref.

Li:2014:PMF

[LLTL14] Zechao Li, Jing Liu, Jinhui Tang, and Hanqing Lu. Projective Matrix Factorization with unified embedding for social image tagging. Computer Vision and Image Understanding: CVIU, 124 (??):71–78, July 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000253.

Liao:2024:ESC

[LLWX24] Yugang Liao, Junqing Li, Shuwei Wei, and Xiumei Xiao. Evolutionary Search via channel attention based parameter inheritance and stochastic uniform sampled training. Computer Vision and Image Understanding: CVIU, 243(??): ??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731422400081X.

Liu:2021:MCT

[LLWZ21] Rui Liu, Yi Liu, Rui Wang, and Yucong Zhou. Mutual calibration training: Training deep neural networks with noisy labels using dual-models. Computer Vision and Image Understanding: CVIU, 212(??): ??, November 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001211.

Li:2018:SRU

[LLY^{+}18] Huibin Li, Yibao Li, Ruixuan Yu, Jian Sun, and Junseok Kim. Surface reconstruction from unorganized points with l_{0} gradient minimization. Computer Vision and Image Understanding: CVIU, 169(??): 108–118, April 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300092.

Liu:2023:MMS

[LLZ23] Jianan Liu, Jian Liu, and Qiang Zhang. M2FINet: Modality-specific and modalityshared features interaction network for RGB-IR person re-identification. Computer Vision and Image Understanding: CVIU, 232 (??):??, July 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000887.

Latecki:1996:AST

[LM96] Longin Latecki and C. Min Ma. An algorithm for a 3D simplicity test. Computer Vision and Image Understanding: CVIU, 63(2): 388–393, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0028/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0028/production/pdf.

Lee:1999:GLR

[LM99a] Mi-Suen Lee and Gérard Medioni. Grouping … into regions, curves, and junctions. Computer Vision and Image Understanding: CVIU, 76(1):54– 69, October 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0787/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.

0787/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0787/production/ref.

Liao:1999:SSA

[LM99b] Chia-Wei Liao and Gérard Medioni. Simultaneous surface approximation and segmentation of complex objects. Computer Vision and Image Understanding: CVIU, 73(1):43– 63, January 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0694/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0694/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0694/production/ref.

Lee:2016:RDC

[LM16] Young Hoon Lee and Gérard Medioni. RGB-D camera based wearable navigation system for the visually impaired. Computer Vision and Image Understanding: CVIU, 149(??): 3–20, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300248.

Li:2009:AER

[LMC09] Shiying Li, Yoshitsugu Manabe, and Kunihiro Chihara. Accurately estimating reflectance parameters for color and gloss reproduction. Computer Vision and Image Understanding: CVIU, 113(2):308–316, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2016:LED

[LmCT16] Xin Liu, Yiu ming Cheung, and Yuan Yan Tang. Lip event detection using oriented histograms of regional optical flow and low rank affinity pursuit. Computer Vision and Image Understanding: CVIU, 148(??): 153–163, July 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002593.

Lopez-Molina:2011:GFE

[LMDB11] C. Lopez-Molina, B. De Baets, and H. Bustince. Generating fuzzy edge images from gradient magnitudes. Computer Vision and Image Understanding: CVIU, 115(11):1571–1580, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211001688.

Liu:2023:LTB

[LML^{+}23] Zhi Liu, Xingyu Mu, Yunhua Lu, Tingting Zhang, and Yingli Tian. Learning transformer-based attention region with multiple scales for occluded person re-identification. Computer Vision and Image Understanding: CVIU, 229 (??):??, March 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000322.

Liu:2022:RRO

[LMM22] Juncheng Liu, Steven Mills, and Brendan McCane. Roc-Net: Recursive octree network for efficient 3D processing. Computer Vision and Image Understanding: CVIU, 224(??): ??, November 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001333.

Labati:2019:NII

[LMP^{+}19] Ruggero Donida Labati, Enrique Muñoz, Vincenzo Piuri, Arun Ross, and Fabio Scotti. Non-ideal iris segmentation using Polar Spline RANSAC and illumination compensation. Computer Vision and Image Understanding: CVIU, 188(??):Article 102787, November 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314219301031.

Lazarevic-McManus:2008:OBC

[LMRMJ08] N. Lazarevic-McManus, J. R. Renno, D. Makris, and G. A. Jones. An objectbased comparative methodology for motion detection based on the F -measure. Computer Vision and Image Understanding: CVIU, 111 (1):74–85, July 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Leo:2017:CVA

[LMT^{+}17] M. Leo, G. Medioni, M. Trivedi, T. Kanade, and G. M. Farinella. Computer vision for assistive technologies. Computer Vision and Image Understanding: CVIU, 154(??):1–15, January 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216301357.

Lin:1998:BDD

[LN98] Chungan Lin and Ramakant Nevatia. Building detection and description from a single intensity image. Computer Vision and Image Understanding: CVIU, 72(2):101– 121, November 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0724/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0724/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0724/production/ref.

Lu:2010:LRC

[LN10] Xin Lu and Kiyoshi Nishiyama.

Low-resolution color-based visual tracking with statespace model identification. Computer Vision and Image Understanding: CVIU, 114(9):1045–1054, September 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Le:2021:AGD

[LNM^{+}21] Hoàng-Ân Lê, Tushar Nimbhorkar, Thomas Mensink, Anil S. Baslamisli, Sezer Karaoglu, and Theo Gevers. Automatic generation of dense non-rigid optical flow. Computer Vision and Image Understanding: CVIU, 212(??): ??, November 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221001181.

Le:2019:ANC

[LNN^{+}19] Trung-Nghia Le, Tam V. Nguyen, Zhongliang Nie, Minh-Triet Tran, and Akihiro Sugimoto. Anabranch network for camouflaged object segmentation. Computer Vision and Image Understanding: CVIU, 184 (??):45–56, July 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219300608.

Lopez-Nicolas:2014:UTM

[LNS14] G. López-Nicolás and C. Sagüés. Unitary torus model for conical mirror based catadioptric system. Computer Vision and Image Understanding: CVIU, 126(??): 67–79, September 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214001350.

Lohou:2010:DNT

[Loh10] Christophe Lohou. Detection of the non-topology preservation of Ma and Sonka’s algorithm, by the use of P -simple points. Computer Vision and Image Understanding: CVIU, 114 (3):384–399, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lu:2010:SAC

[LP10] Yan Lu and Shahram Payandeh. On the sensitivity analysis of camera calibration from images of spheres. Computer Vision and Image Understanding: CVIU, 114(1):8– 20, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Landabaso:2008:SIS

[LPC08] Jose-Luis Landabaso, Montse Pardàs, and Josep Ramon Casas. Shape from inconsistent silhouette. Computer Vision and Image Understanding: CVIU, 112(2): 210–224, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2020:PIR

[LPC^{+}20] Qing Li, Xiaojiang Peng, Liangliang Cao, Wenbin Du, Hao Xing, Yu Qiao, and Qiang Peng. Product image recognition with guidance learning and noisy supervision. Computer Vision and Image Understanding: CVIU, 196(??):Article 102963, July 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300436.

Liang:2001:PED

[LPH01] Jisheng Liang, Ihsin T. Phillips, and Robert M. Haralick. Performance evaluation of document structure extraction algorithms. Computer Vision and Image Understanding: CVIU, 84 (1):144–159, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0933;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0933/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0933/ref.

Lunn:2003:NBR

[LPR^{+}03] Karen E. Lunn, Keith D. Paulsen, David W. Roberts, Francis E. Kennedy, Alex Hartov, and Leah A. Platenik. Nonrigid brain registration: synthesizing full volume deformation fields from model basis solutions constrained by partial volume intraoperative data. Computer Vision and Image Understanding: CVIU, 89(2–3):299–317, February/ March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:2011:OLS

[LPS^{+}11] Soochahn Lee, Sang Hyun Park, Hackjoon Shim, Il Dong Yun, and Sang Uk Lee. Optimization of local shape and appearance probabilities for segmentation of knee cartilage in 3-D MR images. Computer Vision and Image Understanding: CVIU, 115(12):1710– 1720, December 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001639.

Lee:2023:SKD

[LPSK23] Hyoje Lee, Yeachan Park, Hyun Seo, and Myungjoo Kang. Self-knowledge distillation via dropout. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001005.

Lukac:2007:CIP

[LPV07] Rastislav Lukac, Konstantinos N. Plataniotis, and Anastasios N. Venetsanopoulos. Color image processing. Computer Vision and Image Understanding: CVIU, 107(1–2):1–2, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Losson:2013:CTA

[LPVM13] O. Losson, A. Porebski, N. Vandenbroucke, and L. Macaire. Color texture analysis using CFA chromatic co-occurrence matrices. Computer Vision and Image Understanding: CVIU, 117(7):747– 763, July 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213000532.

Liao:2008:FBR

[LPZ08] Iman Yi Liao, Maria Petrou, and Rongchun Zhao. A fractal-based relaxation algorithm for shape from terrain image. Computer Vision and Image Understanding: CVIU, 109(3):227– 243, March 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2021:NCD

[LQQS21] Zhiyang Li, Wenyu Qu, Heng Qi, and Milos Stojmenovic. Near-convex decomposition of 2D shape using visibility range. Computer Vision and Image Understanding: CVIU, 210 (??):??, September 2021.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314221000874.

Latecki:2002:RPN

[LR02] Longin Jan Latecki and Azriel Rosenfeld. Recovering a polygon from noisy data. Computer Vision and Image Understanding: CVIU, 86(1):32–51, April 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lerasle:1999:THL

[LRD99] F. Lerasle, G. Rives, and M. Dhome. Tracking of human limbs by multiocular vision. Computer Vision and Image Understanding: CVIU, 75(3):229– 246, September 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0759/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0759/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0759/production/ref.

Lonn:2019:SPO

[LRD19] Stefan Lonn, Petia Radeva, and Mariella Dimiccoli. Smartphone picture organization: a hierarchical approach. Computer Vision and Image Understanding: CVIU, 187(??):Article 102789, October 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1077314219301055.

Litany:2017:AAS

[LRF^{+}17] Or Litany, Tal Remez, Daniel Freedman, Lior Shapira, Alex Bronstein, and Ran Gal. ASIST: Automatic semantically invariant scene transformation. Computer Vision and Image Understanding: CVIU, 157(??):284– 299, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301102.

Leonardi:2024:EMS

[LRFF24] Rosario Leonardi, Francesco Ragusa, Antonino Furnari, and Giovanni Maria Farinella. Exploiting multimodal synthetic data for egocentric human-object interaction detection in an industrial scenario. Computer Vision and Image Understanding: CVIU, 242(??): ??, May 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000651.

Lopez-Rubio:2011:SAB

[LRLB11] Ezequiel López-Rubio and Rafael Marcos Luque-Baena. Stochastic approximation for background modelling. Computer Vision and Image Understanding: CVIU, 115(6):735–749, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lopez-Rubio:2015:FSA

[LRLR15] Francisco Javier López-Rubio and Ezequiel López-Rubio. Features for stochastic approximation based foreground detection. Computer Vision and Image Understanding: CVIU, 133 (??):30–50, April 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002458.

Lai:2008:RFR

[LRW08] Hung Lai, Venkatesh Ramanathan, and Harry Wechsler. Reliable face recognition using adaptive and robust correlation filters. Computer Vision and Image Understanding: CVIU, 111(3):329–350, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2019:SIR

[LRZ^{+}19] Siyuan Li, Wenqi Ren, Jiawan Zhang, Jinke Yu, and Xiaojie Guo. Single image rain removal via a deep decomposition-composition network. Computer Vision and Image Understanding: CVIU, 186(??):48–57, September 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219300785.

Li:2008:MSS

[LS08] Rui Li and Stan Sclaroff. Multi-scale 3D scene flow from binocular stereo sequences. Computer Vision and Image Understanding: CVIU, 110(1):75–90, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lo:2009:LFE

[LS09] Tsz-Wai Rachel Lo and J. Paul Siebert. Local feature extraction and matching on range images: 2.5D SIFT. Computer Vision and Image Understanding: CVIU, 113(12):1235–1250, December 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2012:SAI

[LS12] Lingyun Liu and Ioannis Stamos. A systematic approach for 2D-image to 3Drange registration in urban environments. Computer Vision and Image Understanding: CVIU, 116(1):25– 37, January 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001809.

Lelieveldt:2000:AMF

[LSB^{+}00] Boudewijn P. F. Lelieveldt, Milan Sonka, Lizann Bolinger, Thomas D. Scholz, Hein Kayser, Rob van der Geest, and Johan H. C. Reiber. Anatomical modeling with fuzzy implicit surface templates: Application to automated localization of the heart and lungs in thoracic MR volumes. Computer Vision and Image Understanding: CVIU, 80 (1):1–20, October 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0864;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0864/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0864/ref.

LeBesnerais:2008:DHM

[LSC08] G. Le Besnerais, M. Sanfourche, and F. Champagnat. Dense height map estimation from oblique aerial image sequences. Computer Vision and Image Understanding: CVIU, 109 (2):204–225, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2015:TDV

[LSCK15] Yibao Li, Jaemin Shin, Yongho Choi, and Junseok Kim. Three-dimensional volume reconstruction from slice data using phase-field models. Computer Vision and Image Understanding: CVIU, 137(??):115– 124, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215000351.

Liu:2003:FAQ

[LSCM03] Yanxi Liu, Karen L. Schmidt, Jeffrey F. Cohn, and Sinjini Mitra. Facial asymmetry quantification for expression invariant human identification. Computer Vision and Image Understanding: CVIU, 91(1–2): 138–159, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Larsen:2007:TEA

[LSD^{+}07] Rasmus Larsen, Mikkel B. Stegmann, Sune Darkner, Søren Forchhammer, Timothy F. Cootes, and Bjarne Kjær Ersbøll. Texture enhanced appearance models. Computer Vision and Image Understanding: CVIU, 106(1):20–30, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2024:MMG

[LSGY24] Zekun Li, Hock Soon Seah, Baolong Guo, and Muli Yang. MLGPnet: Multigranularity neural network for 3D shape recognition using pyramid data. Computer Vision and Image Understanding: CVIU, 239(??): ??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002849.

Li:2019:FGF

[LSH19] Minglei Li, Lei Sun, and Qiang Huo. Flow-guided feature propagation with occlusion aware detail enhancement for hand segmentation in egocentric videos. Computer Vision and Image Understanding: CVIU, 187(??):Article 102785, October 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S1077314219301018.

Lin:2002:NAA

[LSHT02] Chung-Yi Lin, Sheng-Wen Shih, Yi-Ping Hung, and Gregory Y. Tang. A new approach to automatic reconstruction of a 3-D world using active stereo vision. Computer Vision and Image Understanding: CVIU, 85 (2):117–143, February 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lindner:2010:TFS

[LSKK10] Marvin Lindner, Ingo Schiller, Andreas Kolb, and Reinhard Koch. Time-of-flight sensor calibration for accurate range sensing. Computer Vision and Image Understanding: CVIU, 114 (12):1318–1328, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2018:ATT

[LSL^{+}18] Xinzhao Li, Yuanqi Su, Yuehu Liu, Shaozhuo Zhai, and Ying Wu. Active target tracking: a simplified view aligning method for binocular camera model. Computer Vision and Image Understanding: CVIU, 175(??): 11–23, October 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218302820.

Li:2016:MEM

[LSP^{+}16] Rui Li, Pengcheng Shi, Jeff Pelz, Cecilia O. Alm, and Anne R. Haake. Modeling eye movement patterns to characterize perceptual skill in image-based diagnostic reasoning processes. Computer Vision and Image Understanding: CVIU, 151(??):138– 152, October 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000734.

Lukac:2004:SWV

[LSPV04] Rastislav Lukac, Bogdan Smolka, Konstantinos N. Plataniotis, and Anastasios N. Venetsanopoulos. Selection weighted vector directional filters. Computer Vision and Image Understanding: CVIU, 94(1–3): 140–167, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2024:LPM

[LSQL24] Yuezun Li, Pu Sun, Honggang Qi, and Siwei Lyu. LandmarkBreaker: a proactive method to obstruct DeepFakes via disrupting facial landmark extraction. Computer Vision and Image Understanding: CVIU, 240 (??):??, March 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422400016X.

Lee:2013:PFO

[LST13] Jehoon Lee, Romeil Sandhu, and Allen Tannenbaum. Particle filters and occlusion handling for rigid 2D– 3D pose tracking. Computer Vision and Image Understanding: CVIU, 117(8): 922–933, August 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421300088X.

Lopez-Sastre:2011:TMD

[LSTARMB11] R. J. López-Sastre, T. Tuytelaars, F. J. Acevedo-Rodríguez, and S. Maldonado-Bascón. Towards a more discriminative and semantic visual vocabulary. Computer Vision and Image Understanding: CVIU, 115(3): 415–425, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Livne:2012:HAP

[LSTF12] Micha Livne, Leonid Sigal, Nikolaus F. Troje, and David J. Fleet. Human attributes from 3D pose tracking. Computer Vision and Image Understanding: CVIU, 116(5):648– 660, May 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212000203.

Liang:2018:CHP

[LSW18] Shuang Liang, Xiao Sun, and Yichen Wei. Compositional human pose regression. Computer Vision and Image Understanding: CVIU, 176– 177(??):1–8, November/ December 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218304259.

Luettin:1997:SUP

[LT97] Juergen Luettin and Neil A.

Thacker. Speechreading using probabilistic models. Computer Vision and Image Understanding: CVIU, 65(2):163–178, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0570/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0570/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0570/production/ref.

Lachaud:2005:DMC

[LT05] J.-O. Lachaud and B. Taton. Deformable model with a complexity independent from image resolution. Computer Vision and Image Understanding: CVIU, 99(3):453–475, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Labussiere:2023:BAM

[LTAA23] Mathieu Labussière, Céline Teulière, and Omar Ait-Aider. Blur aware metric depth estimation with multi-focus plenoptic cameras. Computer Vision and Image Understanding: CVIU, 235(??):??, October 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223001820.

Liu:2014:MHD

[LTCT14] Weifeng Liu, Dacheng Tao, Jun Cheng, and Yuanyan Tang. Multiview Hessian discriminative sparse coding for image annotation. Computer Vision and Image Understanding: CVIU, 118(??):50–60, January 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213001550.

Lin:2023:USL

[LTL^{+}23] Yan-Bo Lin, Hung-Yu Tseng, Hsin-Ying Lee, Yen-Yu Lin, and Ming-Hsuan Yang. Unsupervised sound localization via iterative contrastive learning. Computer Vision and Image Understanding: CVIU, 227(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001801.

Li:2015:KRM

[LTY^{+}15] Yansheng Li, Yihua Tan, Jin-Gang Yu, Shengxiang Qi, and Jinwen Tian. Kernel regression in mixed feature spaces for spatiotemporal saliency detection. Computer Vision and Image Understanding: CVIU, 135(??):126– 140, June 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000260.

Lucchese:2001:FDT

[Luc01] Luca Lucchese. A frequency domain technique based on energy radial projections for robust estimation of global 2D affine transformations. Computer Vision and Image Understanding: CVIU, 81(1):72–116, January 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0885;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0885/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0885/ref.

Luong:1996:CRG

[LV96] Q.-T. Luong and T. Viéville. Canonical representations for the geometries of multiple projective views. Computer Vision and Image Understanding: CVIU, 64(2): 193–229, September 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0055/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0055/production/pdf.

LeGuyader:2011:CSR

[LV11] Carole Le Guyader and Luminita A. Vese. A combined segmentation and registration framework with a nonlinear elasticity smoother. Computer Vision and Image Understanding: CVIU, 115(12):1689–1709, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S1077314211001573.

Langerak:2015:ILF

[LvdHK^{+}15] T. R. Langerak, U. A.

van der Heide, A. N. T. J. Kotte, F. F. Berendsen, and J. P. W. Pluim. Improving label fusion in multiatlas based segmentation by locally combining atlas selection and performance estimation. Computer Vision and Image Understanding: CVIU, 130(??):71–79, January 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001866.

Lobel:2020:CCH

[LVS20] Hans Lobel, René Vidal, and Alvaro Soto. CompactNets: Compact hierarchical compositional networks for visual recognition. Computer Vision and Image Understanding: CVIU, 191(??):Article 102841, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218301905.

Lam:1997:HVU

[LVW97] C. P. Lam, S. Venkatesh, and G. A. W. West. Hypothesis verification using parametric models and active vision strategies. Computer Vision and Image Understanding: CVIU, 68(2):209–236, November 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0554/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0554/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0554/production/ref.

Limberger:2018:CBS

[LW18] Frederico A. Limberger and Richard C. Wilson. Curvature-based spectral signatures for non-rigid shape retrieval. Computer Vision and Image Understanding: CVIU, 172(??): 1–11, July 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218300535.

Li:2024:DPR

[LWFL24] Xiaofang Li, Weiwei Wang, Xiangchu Feng, and Min Li. Deep parametric Retinex decomposition model for low-light image enhancement. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000298.

Luo:2003:OBA

[LWH03] Ying Luo, Tzong-Der Wu, and Jenq-Neng Hwang. Object-based analysis and interpretation of human motion in sports video sequences by dynamic Bayesian networks. Computer Vision and Image Understanding: CVIU, 92(2–3):196– 216, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2023:RAR

[LWH^{+}23] Wen Li, Hengyou Wang, Lianzhi Huo, Qiang He, and Changlun Zhang. Robust attention ranking architecture with frequencydomain transform to defend against adversarial samples. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000978.

Li:2022:SMP

[LWLC22] Wenbo Li, Yi Wei, Siwei Lyu, and Ming-Ching Chang. Simultaneous multiperson tracking and activity recognition based on cohesive cluster search. Computer Vision and Image Understanding: CVIU, 214 (??):??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001454.

Liu:2023:GSK

[LWLP23] Fuchang Liu, Yu Wang, Zheng Li, and Zhigeng Pan. GEIKD: Self-knowledge distillation based on gated ensemble networks and influences-based label noise removal. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001510.

Liu:2012:SCW

[LWLS12] Junbin Liu, Tim Wark, Ruan Lakemond, and Sridha Sridharan. Self-calibration of wireless cameras with restricted degrees of freedom. Computer Vision and Image Understanding: CVIU, 116(10):1033–1046, October 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212000902.

Li:2017:STS

[LWLT17] Laquan Li, Jian Wang, Wei Lu, and Shan Tan. Simultaneous tumor segmentation, image restoration, and blur kernel estimation in PET using multiple regularizations. Computer Vision and Image Understanding: CVIU, 155(??):173– 194, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301552.

Lao:2016:HRD

[LWlZ16] Shihong Lao, Dong Wang, Fu li, and Haihong Zhang. Human running detection: Benchmark and baseline. Computer Vision and Image Understanding: CVIU, 153(??):143–150, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216000771.

Liu:2016:LPS

[LWSC16] Mengyi Liu, Ruiping Wang, Shiguang Shan, and Xilin Chen. Learning prototypes and similes on Grassmann manifold for spontaneous expression recognition. Computer Vision and Image Understanding: CVIU, 147(??):95– 101, June 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215001812.

Li:2021:LAW

[LWW^{+}21] Zheng Li, Chaofeng Wang, Jun Wang, Shihui Ying, and Jun Shi. Lightweight adaptive weighted network for single image superresolution. Computer Vision and Image Understanding: CVIU, 211(??): ??, October 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000989.

Lai:2017:EGH

[LWY^{+}17] Taotao Lai, Hanzi Wang, Yan Yan, Guobao Xiao, and David Suter. Efficient guided hypothesis generation for multi-structure epipolar geometry estimation. Computer Vision and Image Understanding: CVIU, 154(??):152– 165, January 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301564.

Li:2024:EIG

[LWYF24] Jingmeng Li, Hui Wei, Surun Yang, and Lukang Fu. Emerging image generation with flexible control of perceived difficulty. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002990.

Li:2014:VOT

[LWZC14] Shuxiao Li, Ou Wu, Chengfei Zhu, and Hongxing Chang. Visual object tracking using spatial context information and global tracking skills. Computer Vision and Image Understanding: CVIU, 125(??):1–15, August 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001847.

Li:2017:TLG

[LWZP17] Chao Li, Ping Wang, Xiangyu Zhu, and Huali Pi. Three-layer graph framework with the sumD feature for alpha matting. Computer Vision and Image Understanding: CVIU, 162(??):34–45, September 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301236.

Liang:2016:AIR

[LXFM16] Jinglian Liang, Chao Xu, Zhiyong Feng, and Xirong Ma. Affective interaction recognition using spatiotemporal features and context. Computer Vision and Image Understanding: CVIU, 144(??):155– 165, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002180.

Li:2023:CAV

[LXS^{+}23] Haoling Li, Mengqi Xue, Jie Song, Haofei Zhang, Wenqi Huang, Lingyu Liang, and Mingli Song. Constituent attention for vision transformers. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223002187.

Liu:2017:HMV

[LXW^{+}17] An-An Liu, Ning Xu, Yongkang Wong, Junnan Li, Yu-Ting Su, and Mohan Kankanhalli. Hierarchical & multimodal video captioning: Discovering and transferring multimodal knowledge for vision to language. Computer Vision and Image Understanding: CVIU, 163(??): 113–125, October 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300735.

Lv:2023:GAL

[LXW^{+}23] Ning Lv, Xuezhi Xiang, Xinyao Wang, Yulong Qiao, and Abdulmotaleb El Saddik. Global-aware and localaware enhancement network for person search. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001844.

Lv:2024:LFC

[LXW^{+}24] Ning Lv, Xuezhi Xiang, Xinyao Wang, Yulong Qiao, and Abdulmotaleb El Saddik. Learning feature contexts by transformer and CNN hybrid deep network for weakly supervised person search. Computer Vision and Image Understanding: CVIU, 239(??): ??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002862.

Lee:2005:LDI

[LY05] Joon Woong Lee and Un Kun Yi. A lanedeparture identification based on LBPE, Hough transform, and linear regression. Computer Vision and Image Understanding: CVIU, 99(3): 359–383, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Levine:2006:FRS

[LY06] Martin David Levine and Yingfeng Yu. Face recognition subject to variations in facial expression, illumination and pose using correlation filters. Computer Vision and Image Understanding: CVIU, 104(1):1– 15, October 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lhuillier:2013:MSR

[LY13] Maxime Lhuillier and Shuda Yu. Manifold surface reconstruction of an environment from sparse Structure-from-Motion data. Computer Vision and Image Understanding: CVIU, 117(11):1628– 1644, November 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213001525.

Luo:2013:CEH

[LYA13] Weilan Luo, Toshihiko Yamasaki, and Kiyoharu Aizawa. Cooperative estimation of human motion and surfaces using multiview videos. Computer Vision and Image Understanding: CVIU, 117(11):1560– 1574, November 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001367.

Li:2017:HVE

[LYBT17] Yu Li, Shaodi You, Michael S.

Brown, and Robby T. Tan. Haze visibility enhancement: a survey and quantitative benchmarking. Computer Vision and Image Understanding: CVIU, 165(??):1–16, December 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301595.

Liu:2024:DGF

[LYD24] Fangyi Liu, Mang Ye, and Bo Du. Domain generalized federated learning for person re-identification. Computer Vision and Image Understanding: CVIU, 241 (??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422400050X.

Liu:2007:SSO

[LYG07] Xin Liu, Hongxun Yaoand, and Wen Gao. Shape from silhouette outlines using an adaptive dandelion model. Computer Vision and Image Understanding: CVIU, 105 (2):121–130, February 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2019:TSN

[LYKY19] Jun Li, Can Yuce, Reinhard Klein, and Angela Yao. A two-streamed network for estimating finescaled depth maps from single RGB images. Computer Vision and Image Understanding: CVIU, 186(??): 25–36, September 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421930092X.

Lu:2017:ILM

[LYSK17] Guoyu Lu, Yan Yan, Nicu Sebe, and Chandra Kambhamettu. Indoor localization via multi-view images and videos. Computer Vision and Image Understanding: CVIU, 161(??):145– 160, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300802.

Liu:2012:LSF

[LYSS12] Jingen Liu, Yang Yang, Imran Saleemi, and Mubarak Shah. Learning semantic features for action recognition via diffusion maps. Computer Vision and Image Understanding: CVIU, 116(3):361– 377, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002116.

Li:2024:DGN

[LYW^{+}24] Chenglong Li, Xiaobin Yang, Guohao Wang, Aihua Zheng, Chang Tan, and Jin Tang. Disentangled generation network for enlarged license plate recognition and a unified dataset. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002606.

Li:2021:ILS

[LYX^{+}21] Guanbin Li, Pengxiang Yan, Yuan Xie, Guisheng Wang, Liang Lin, and Yizhou Yu. Instance-level salient object segmentation. Computer Vision and Image Understanding: CVIU, 207(??): ??, June 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221000515.

Liang:2024:SBC

[LYX^{+}24] Shuang Liang, Zhiqi Yan, Chi Xie, Hongming Zhu, and Jiewen Wang. Scribblebased complementary graph reasoning network for weakly supervised salient object detection. Computer Vision and Image Understanding: CVIU, 243(??): ??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000584.

Langer:1997:CLI

[LZ97a] Michael S. Langer and Steven W. Zucker. Casting light on illumination: a computational model and dimensional analysis of sources. Computer Vision and Image Understanding: CVIU, 65(2):322– 335, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0574/production;   http:
//www.idealibrary.com/

links/artid/cviu.1996.
0574/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0574/production/ref.

Lopresti:1997:UCS

[LZ97b] Daniel Lopresti and Jiangying Zhou. Using consensus sequence voting to correct OCR errors. Computer Vision and Image Understanding: CVIU, 67(1): 39–47, July 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0502/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0502/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0502/production/ref.

Li:2024:FLB

[LZ24] Guobin Li and Reyer Zwiggelaar. Feature learning based on connectivity estimation for unbiased mammography mass classification. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002643.

Li:2023:LFP

[LZB^{+}23] Zhaoyang Li, Xue Zhao, Chun Bao, Jie Cao, Dongxing Li, and Qun Hao. Lightweight feature point detection network with channel enhancement. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223000966.

Lin:2020:CSS

[LZC^{+}20] Songnan Lin, Jiawei Zhang, Jing Chen, Yongtian Wang, Yicun Liu, and Jimmy Ren. Cross-spectral stereo matching for facial disparity estimation in the dark. Computer Vision and Image Understanding: CVIU, 200(??):Article 103046, November 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S107731422030093X.

Lei:2014:OLD

[LZD^{+}14] Ying-Ke Lei, Ji-Wei Zou, Tianbao Dong, Zhu-Hong You, Yuan Yuan, and Yihua Hu. Orthogonal locally discriminant spline embedding for plant leaf recognition. Computer Vision and Image Understanding: CVIU, 119(??):116– 126, February 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213002324.

Liu:2017:RSC

[LZL^{+}17] Maofu Liu, Luming Zhang, Ya Liu, Huijun Hu, and Wei Fang. Recognizing semantic correlation in imagetext Weibo via feature space mapping. Computer Vision and Image Understanding: CVIU, 163(??):58–66, October 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300747.

Lv:2022:BDN

[LZL^{+}22] Xiaoqian Lv, Shengping Zhang, Qinglin Liu, Haozhe Xie, Bineng Zhong, and Huiyu Zhou. BacklitNet: a dataset and network for backlit image enhancement. Computer Vision and Image Understanding: CVIU, 218(??):??, April 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000340.

Lian:2010:QPB

[LZLP10] Wei Lian, Lei Zhang, Yan Liang, and Quan Pan.

A quadratic programming based cluster correspondence projection algorithm for fast point matching. Computer Vision and Image Understanding: CVIU, 114 (3):322–333, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2017:ESI

[LZmC^{+}17] Xin Liu, He Zhang, Yiu ming Cheung, Xinge You, and Yuan Yan Tang. Efficient single image dehazing and denoising: an efficient multi-scale correlated wavelet approach. Computer Vision and Image Understanding: CVIU, 162 (??):23–33, September 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314217301431.

Li:2024:AAN

[LZQL24] Yuezun Li, Cong Zhang, Honggang Qi, and Siwei Lyu. AdaNI: Adaptive Noise Injection to improve adversarial robustness. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002357.

Long:2016:ROM

[LZS16] Yang Long, Fan Zhu, and Ling Shao. Recognising occluded multi-view actions using local nearest neighbour embedding. Computer Vision and Image Understanding: CVIU, 144(??): 36–45, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001319.

Li:2024:LPL

[LZTX24] Tengjie Li, Sicong Zang, Shikui Tu, and Lei Xu. Lmser-pix2seq: Learning stable sketch representations for sketch healing. Computer Vision and Image Understanding: CVIU, 240 (??):??, March 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000122.

Li:2024:HCR

[LZW^{+}24] Changzhen Li, Jie Zhang, Shuzhe Wu, Xin Jin, and Shiguang Shan. Hierarchical compositional representations for few-shot action recognition. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223002916.

Liu:2003:MRM

[LZWP03] Feng Liu, Yueting Zhuang, Fei Wu, and Yunhe Pan. 3D motion retrieval with motion index tree. Computer Vision and Image Understanding: CVIU, 92 (2–3):265–284, November/ December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2021:SIS

[LZWX21] Yunan Liu, Shanshan Zhang, Chunpeng Wang, and Jie Xu. Single image superresolution via hybrid resolution NSST prediction. Computer Vision and Image Understanding: CVIU, 207(??):??, June 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000461.

Lin:2023:DAN

[LZYW23] Xin Lin, Li Zhu, Shuyu Yang, and Yaxiong Wang. Diff attention: a novel attention scheme for person re-identification. Computer Vision and Image Understanding: CVIU, 228(??): ??, February 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000036.

Li:2021:DTI

[LZZ^{+}21] Ailin Li, Lei Zhao, Zhiwen Zuo, Zhizhong Wang, Haibo Chen, Dongming Lu, and Wei Xing. Diversified text-to-image generation via deep mutual information estimation. Computer Vision and Image Understanding: CVIU, 211 (??):??, October 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422100103X.

Li:2022:SSS

[LZZ22] Xingye Li, Ling Zhang, and Zhigang Zhu. Snapshot-Net: Self-supervised feature learning for point cloud data segmentation using minimal labeled data. Computer Vision and Image Understanding: CVIU, 216(??): ??, February 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001740.

Li:2023:AGD

[LZZ^{+}23a] Zonglin Li, Shengping Zhang, Zhaoxin Zhang, Quanling Meng, Qinglin Liu, and Huiyu Zhou. Attention guided domain alignment for conditional face image generation. Computer Vision and Image Understanding: CVIU, 234 (??):??, September 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314223001200.

Liu:2023:RCA

[LZZ^{+}23b] Yuetong Liu, Rui Zhang, Yunfeng Zhang, Xiao Pan, Xunxiang Yao, Zhaorui Ni, and Huijian Han. Recurrent context-aware multistage network for single image deraining. Computer Vision and Image Understanding: CVIU, 227(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001904.

Mascetti:2016:RTL

[MAG^{+}16] Sergio Mascetti, Dragan Ahmetovic, Andrea Gerino, Cristian Bernareggi, Mario Busso, and Alessandro Rizzi. Robust traffic lights detection on mobile devices for pedestrians with visual impairment. Computer Vision and Image Understanding: CVIU, 148(??):123– 135, July 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002611.

Mahapatra:2016:CME

[Mah16] Dwarikanath Mahapatra.

Combining multiple expert annotations using semisupervised learning and graph cuts for medical image segmentation. Computer Vision and Image Understanding: CVIU, 151 (??):114–123, October 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314216000199.

Mishra:2016:EEM

[MAJ16] Anand Mishra, Karteek Alahari, and C. V. Jawahar. Enhancing energy minimization framework for scene text recognition with top-down cues. Computer Vision and Image Understanding: CVIU, 145(??): 30–42, April 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421600014X.

Milletari:2017:HCD

[MAK^{+}17] Fausto Milletari, Seyed-Ahmad Ahmadi, Christine Kroll, Annika Plate, Verena Rozanski, Juliana Maiostre, Johannes Levin, Olaf Dietrich, Birgit Ertl-Wagner, Kai Bötzel, and Nassir Navab. Hough-CNN: Deep learning for segmentation of deep brain regions in MRI and ultrasound. Computer Vision and Image Understanding: CVIU, 164(??):92–102, November 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300620.

Michel:2010:HML

[MAL10] Damien Michel, Antonis A. Argyros, and Manolis I. A. Lourakis. Horizon matching for localizing unordered panoramic images. Computer Vision and Image Understanding: CVIU, 114 (2):274–285, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Malti:2021:ERC

[Mal21] Abed Malti. On the exact recovery conditions of 3D human motion from 2D landmark motion with sparse articulated motion. Computer Vision and Image Understanding: CVIU, 202(??):Article 103072, January 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220301107.

Monga:1997:TNC

[MAM97] Olivier Monga, Nasser Armande, and Philippe Montesinos. Thin nets and crest lines: Application to satellite data and medical images. Computer Vision and Image Understanding: CVIU, 67(3):285– 295, September 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0507/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0507/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0507/production/ref.

Mandal:1999:FWH

[MAP99] M. K. Mandal, T. Aboulnasr, and S. Panchanathan. Fast wavelet histogram techniques for image indexing. Computer Vision and Image Understanding: CVIU, 75(1–2):99–110, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0766/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0766/production/pdf;
http://www.idealibrary.

                                        com/links/artid/cviu.1999.
                                        0766/production/ref.

Markussen:2007:LDD

[Mar07] Bo Markussen. Large deformation diffeomorphisms with application to optic flow. Computer Vision and Image Understanding: CVIU, 106(1):97–105, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Masuda:2002:RIM

[Mas02] Takeshi Masuda. Registration and integration of multiple range images by matching signed distance fields for object shape modeling. Computer Vision and Image Understanding: CVIU, 87(1–3):51–65, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Masuda:2009:LPH

[Mas09]  Takeshi Masuda.  Log-

polar height maps for multiple range image registration. Computer Vision and Image Understanding: CVIU, 113(11):1158–1169, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Maybank:1997:RPR

[May97]  S. J. Maybank.  Reply

to Pizlo, Rosenfeld, and Weiss. Computer Vision and Image Understanding: CVIU, 67(3):318– 319, September 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0643/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0643/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0643/production/ref. See

[PRW97a].

Mayer:1999:AOE

[May99] Helmut Mayer. Automatic object extraction from aerial imagery — a survey focusing on buildings. Computer Vision and Image Understanding: CVIU, 74(2): 138–149, May 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0750/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0750/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0750/production/ref.

Meng:2010:MML

[MAY^{+}10] Hongying Meng, Kofi Appiah, Shigang Yue, Andrew Hunter, Mervyn Hobden, Nigel Priestley, Peter Hobden, and Cy Pettit. A modified model for the Lobula Giant Movement Detector and its FPGA implementation. Computer Vision and Image Understanding: CVIU, 114(11):1238–1247, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Monga:1995:UPD

[MB95] Olivier Monga and Serge Benayoun. Using partial derivatives of 3D images to extract typical surface features. Computer Vision and Image Understanding: CVIU, 61(2):171– 189, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1014/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1014/production/pdf.

Min:2005:IRI

[MB05] Jaesik Min and Kevin W. Bowyer. Improved range image segmentation by analyzing surface fit patterns. Computer Vision and Image Understanding: CVIU, 97 (2):242–258, February 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Moore:2011:LBP

[MB11] S. Moore and R. Bowden.

Local binary patterns for multi-view facial expression recognition. Computer Vision and Image Understanding: CVIU, 115(4):541– 558, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Morimitsu:2017:ESL

[MBCJ17] Henrique Morimitsu, Isabelle Bloch, and Roberto M. Cesar-Jr. Exploring structure for long-term tracking of multiple objects in sports videos. Computer Vision and Image Understanding: CVIU, 159(??): 89–104, June 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216302004.

Mane:2022:SCH

[MBD^{+}22] Tejas Mane, Aylar Bayramova, Kostas Daniilidis, Philippos Mordohai, and Elena Bernardis. Singlecamera 3D head fitting for mixed reality clinical applications. Computer Vision and Image Understanding: CVIU, 218(??): ??, April 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314222000224.

Mahmud:2021:PDN

[MBHRC21] Tahmida Mahmud, Mohammad Billah, Mahmudul Hasan, and Amit K. Roy-Chowdhury. Prediction and description of near-future activities in video. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000746.

MacArthur:2002:ICB

[MBKB02] Sean D. MacArthur, Carla E. Brodley, Avinash C. Kak, and Lynn S. Broderick. Interactive content-based image retrieval using relevance feedback. Computer Vision and Image Understanding: CVIU, 88(2):55–75, November 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mille:2011:CNB

[MBMC11] Julien Mille, Romuald Boné, Pascal Makris, and Hubert Cardot. Corrigendum to “Narrow band region-based active contours and surfaces for 2D and 3D segmentation” [Comput. Vis. Image Understand. 113 (2009)

946–965]. Computer Vision and Image Understanding: CVIU, 115(2):286, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Maggio:2009:AAB

[MC09a] Emilio Maggio and Andrea Cavallaro. Accurate appearance-based Bayesian tracking for maneuvering targets. Computer Vision and Image Understanding: CVIU, 113(4):544– 555, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Marques:2009:ESD

[MC09b] Manuel Marques and João Costeira. Estimating 3D shape from degenerate sequences with missing data. Computer Vision and Image Understanding: CVIU, 113 (2):261–272, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

McGuigan:2020:ARA

[MC20] Matthew McGuigan and Jacqueline Christmas. Automating RTI: Automatic light direction detection and correcting non-uniform lighting for more accurate surface normals. Computer Vision and Image Understanding: CVIU, 192(??):

Article 102880, March 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314218302686.

Munusamy:2022:VCU

[MC22] Hemalatha Munusamy and Chandra Sekhar C. Video captioning using Semantically Contextual Generative Adversarial Network. Computer Vision and Image Understanding: CVIU, 221 (??):??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222000650.

Massoli:2021:DFR

[MCAF21] Fabio Valerio Massoli, Fabio Carrara, Giuseppe Amato, and Fabrizio Falchi. Detection of face recognition adversarial attacks. Computer Vision and Image Understanding: CVIU, 202 (??):Article 103103, January 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314220301296.

Martins:2013:GFA

[MCB13] Pedro Martins, Rui Caseiro, and Jorge Batista. Generative face alignment through 2.5D active appearance models. Computer Vision and Image Understanding: CVIU, 117(3): 250–268, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001944.

Martins:2020:TCF

[MCCRAC20] Isabel Martins, Pedro Carvalho, Luís Corte-Real, and José Luis Alba-Castro. Texture collinearity foreground segmentation for night videos. Computer Vision and Image Understanding: CVIU, 200(??):Article 103032, November 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220300862.

Molnar:2010:IRV

[MCF10] József Molnár, Dmitry Chetverikov, and Sándor Fazekas. Illuminationrobust variational optical flow using cross-correlation. Computer Vision and Image Understanding: CVIU, 114(10):1104–1114, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mrak:2009:FAS

[MĆK09] Marta Mrak, Janko Ćalić, and Ahmet Kondoz. Fast analysis of scalable video for adaptive browsing interfaces. Computer Vision and Image Understanding: CVIU, 113(3):425– 434, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Meng:2016:CMI

[MCL16] Fanman Meng, Jianfei Cai, and Hongliang Li. Cosegmentation of multiple image groups. Computer Vision and Image Understanding: CVIU, 146(??): 67–76, May 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000497.

Mhalla:2017:SFR

[MCM^{+}17] Ala Mhalla, Thierry Chateau, Houda Maâmatou, Sami Gazzah, and Najoua Essoukri Ben Amara. SMC faster R-CNN: Toward a scene-specialized multiobject detector. Computer Vision and Image Understanding: CVIU, 164(??):3– 15, November 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301200.

Mignotte:1999:TCM

[MCPB99] M. Mignotte, C. Collet, P. Pérez, and P. Bouthemy. Three-class Markovian segmentation of high-resolution sonar images. Computer Vision and Image Understanding: CVIU, 76(3):191– 204, December 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0804/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0804/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0804/production/ref.

Mignotte:2000:MRF

[MCPB00] M. Mignotte, C. Collet, P. Pérez, and P. Bouthemy. Markov random field and fuzzy logic modeling in sonar imagery: Application to the classification of underwater floor. Computer Vision and Image Understanding: CVIU, 79 (1):4–24, July 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0844;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0844/pdf;   http:
//www.idealibrary.com/

                                        links/doi/10.1006/cviu.
                                        2000.0844/ref.

Mavrinac:2010:ACM

[MCT10] Aaron Mavrinac, Xiang Chen, and Kemal Tepe. An automatic calibration method for stereo-based 3D distributed smart camera networks. Computer Vision and Image Understanding: CVIU, 114(8):952–962, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Maqueda:2015:HCI

[MdBJG15] Ana I. Maqueda, Carlos R. del Blanco, Fernando Jaureguizar, and Narciso García. Human-computer interaction based on visual hand-gesture recognition using volumetric spatiograms of local binary patterns. Computer Vision and Image Understanding: CVIU, 141(??):126– 137, December 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001629.

Maurin:2009:FAS

[MDdMG09] B. Maurin, C. Doignon, M. de Mathelin, and A. Gangi. A fast and automatic stereotactic registration with a single CT-slice. Computer Vision and Image Understanding: CVIU, 113 (8):878–890, August 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Macrini:2011:BGM

[MDFS11a] Diego Macrini, Sven Dickinson, David Fleet, and Kaleem Siddiqi. Bone graphs: Medial shape parsing and abstraction. Computer Vision and Image Understanding: CVIU, 115(7): 1044–1061, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000798.

Macrini:2011:OCU

[MDFS11b] Diego Macrini, Sven Dickinson, David Fleet, and Kaleem Siddiqi. Object categorization using bone graphs. Computer Vision and Image Understanding: CVIU, 115(8):1187– 1206, August 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000907.

Majhi:2024:HSN

[MDK^{+}24] Snehashis Majhi, Rui Dai, Quan Kong, Lorenzo Garattoni, Gianpiero Francesca, and François Brémond. Human-scene network: a novel baseline with selfrectifying loss for weakly supervised video anomaly detection. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000365.

Mondal:2023:FIF

[MDL^{+}23] Milton Mondal, Bishshoy Das, Brejesh Lall, Pushpendra Singh, Sumantra Dutta Roy, and Shiv Dutt Joshi. Feature independent filter pruning by successive layers analysis. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002084.

Mahpod:2021:FLL

[MDM^{+}21] Shahar Mahpod, Rig Das, Emanuele Maiorana, Yosi Keller, and Patrizio Campisi. Facial landmarks localization using cascaded neural networks. Computer Vision and Image Understanding: CVIU, 205(??):Article 103171, April 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000151.

Medeiros:2019:DTR

[MdOBA19] Heitor R. Medeiros, Felipe D. B. de Oliveira, Hansenclever F. Bassani, and Aluizio F. R. Araujo. Dynamic topology and relevance learning SOM-based algorithm for image clustering tasks. Computer Vision and Image Understanding: CVIU, 179(??): 19–30, February 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421830451X.

Moutzouris:2015:ETH

[MdRNM15] Alexandros Moutzouris, Jesus Martinez del Rincon, Jean-Christophe Nebel, and Dimitrios Makris. Efficient tracking of human poses using a manifold hierarchy. Computer Vision and Image Understanding: CVIU, 132(??):75– 86, March 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214002094.

Moganti:1998:SPC

[ME98a] Madhav Moganti and Fikret Ercal. Segmentation of printed circuit board images into basic patterns. Computer Vision and Image Understanding: CVIU, 70(1): 74–86, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0594/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0594/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0594/production/ref.

Moganti:1998:SLI

[ME98b] Madhav Moganti and Fikret Ercal. A subpattern level inspection system for printed circuit boards. Computer Vision and Image Understanding: CVIU, 70(1): 51–62, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0600/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0600/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0600/production/ref.

Molnar:2018:DGA

[ME18] József Molnár and Iván Eichhardt. A differential geometry approach to camera-independent image correspondence. Computer Vision and Image Understanding: CVIU, 169(??): 90–107, April 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300171.

Moganti:1996:API

[MEDT96] Madhav Moganti, Fikret Ercal, Cihan H. Dagli, and Shou Tsunekawa. Automatic PCB inspection algorithms: a survey. Computer Vision and Image Understanding: CVIU, 63(2): 287–313, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0020/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0020/production/pdf.

Manafifard:2017:SPT

[MEM17] M. Manafifard, H. Ebadi, and H. Abrishami Moghaddam. A survey on player tracking in soccer videos. Computer Vision and Image Understanding: CVIU, 159 (??):19–46, June 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300309.

Matov:2011:OFM

[MEYD11] Alexandre Matov, Marcus M. Edvall, Ge Yang, and Gaudenz Danuser. Optimalflow minimum-cost correspondence assignment in particle flow tracking. Computer Vision and Image Understanding: CVIU, 115(4): 531–540, April 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mackay:2011:MCA

[MFB11] Matthew D. Mackay, Robert G. Fenton, and Beno Benhabib. Multi-camera active surveillance of an articulated human form — an implementation strategy. Computer Vision and Image Understanding: CVIU, 115(10):1395–1413, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211001433.

Michalke:2010:BIV

[MFG10] Thomas Michalke, Jannik Fritsch, and Christian Goerick. A biologicallyinspired vision architecture for resource-constrained intelligent vehicles. Computer Vision and Image Understanding: CVIU, 114(5): 548–563, May 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mao:1995:IMF

[MFJ95] Jianchang Mao, Patrick J.

Flynn, and Anil K. Jain. Integration of multiple feature groups and multiple views into a 3D object recognition system. Computer Vision and Image Understanding: CVIU, 62(3):309– 325, November 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1057/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1057/production/pdf.

Mohammed:2020:PDP

[MFP^{+}20] Ahmed Mohammed, Ivar Farup, Marius Pedersen, Sule Yildirim, and Øistein Hovde. PS-DeVCEM: Pathology-sensitive deep learning model for video capsule endoscopy based on weakly labeled data. Computer Vision and Image Understanding: CVIU, 201(??):Article 103062, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220301053.

Medioni:2007:RRT

[MFS^{+}07] Gérard Medioni, Alexandre R. J. François, Matheen Siddiqui, Kwangsu Kim, and Hosub Yoon. Robust real-time vision for a personal service robot. Computer Vision and Image Understanding: CVIU, 108(1–2):196–203, October/ November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Marnissi:2023:IDA

[MFSB23a] Mohamed Amine Marnissi, Hajer Fradi, Anis Sahbani, and Najoua Essoukri Ben Amara. Improved domain adaptive object detector via adversarial feature learning. Computer Vision and Image Understanding: CVIU, 230(??):??, April 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223000401.

See corrigendum [MFSB23b].

Marnissi:2023:CID

[MFSB23b] MohamedAmine Marnissi, Hajer Fradi, Anis Sahbani, and Najoua Essoukri Ben Amara. Corrigendum to “Improved domain adaptive object detector via adversarial feature learning” [Comput. Vis. Image Underst. 230 (2023)

103660]. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000656.

See [MFSB23a].

Mason:1995:ASP

[MG95] Scott O. Mason and Armin Grün. Automatic sensor placement for accurate dimensional inspection. Computer Vision and Image Understanding: CVIU, 61(3): 454–467, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1034/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1034/production/pdf.

Moeslund:2001:SCV

[MG01] Thomas B. Moeslund and Erik Granum. A survey of computer vision-based human motion capture. Computer Vision and Image Understanding: CVIU, 81 (3):231–268, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0897;   http:

                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0897/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0897/ref.

Manfredi:2017:SMD

[MGCS17] Marco Manfredi, Costantino Grana, Rita Cucchiara, and Arnold W. M. Smeulders. Segmentation models diversity for object proposals. Computer Vision and Image Understanding: CVIU, 158 (??):40–48, May 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300819.

Matas:2000:RDL

[MGK00] J. Matas, C. Galambos, and J. Kittler. Robust detection of lines using the progressive probabilistic Hough transform. Computer Vision and Image Understanding: CVIU, 78 (1):119–137, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.1999.0831;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        1999.0831/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        1999.0831/ref.

Murino:2017:IVU

[MGLB17] Vittorio Murino, Shaogang Gong, Chen Change Loy, and Loris Bazzani. Image and video understanding in big data. Computer Vision and Image Understanding: CVIU, 156(??): 1–3, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300152.

Manzotti:2001:DEL

[MGMS01] R. Manzotti, A. Gasteratos, G. Metta, and G. Sandini. Disparity estimation on log-polar images and vergence control. Computer Vision and Image Understanding: CVIU, 83 (2):97–117, August 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0924;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0924/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0924/ref.

Morillas:2008:IIN

[MGPF08] Samuel Morillas, Valentín Gregori, and Guillermo Peris-Fajarnés. Isolating impulsive noise pixels in color images by peer group techniques. Computer Vision and Image Understanding: CVIU, 110(1):102– 116, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Moreno:2011:EPC

[MGPJ11] Rodrigo Moreno, Miguel Angel Garcia, Domenec Puig, and Carme Julià. Edgepreserving color image denoising through tensor voting. Computer Vision and Image Understanding: CVIU, 115(11):1536–1551, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211001706.

Melendez:2011:UTB

[MGPP11] Jaime Melendez, Miguel Angel Garcia, Domenec Puig, and Maria Petrou. Unsupervised texture-based image segmentation through pattern discovery. Computer Vision and Image Understanding: CVIU, 115(8): 1121–1133, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314211000968.

Moreno-Garcia:2015:CMC

[MGS15] Carlos Francisco Moreno-García and Francesc Serratosa. Consensus of multiple correspondences between sets of elements. Computer Vision and Image Understanding: CVIU, 142 (??):50–64, ???? 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001836.

Ma:2010:HSM

[MGW10] Yu Ma, Xiaodong Gu, and Yuanyuan Wang. Histogram similarity measure using variable bin size distance. Computer Vision and Image Understanding: CVIU, 114 (8):981–989, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mostafa:2013:FRL

[MHAF13] Eslam Mostafa, Riad Hammoud, Asem Ali, and Aly Farag. Face recognition in low resolution thermal images. Computer Vision and Image Understanding: CVIU, 117(12):1689–1694, December 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001409.

Moeslund:2006:SAV

[MHK06] Thomas B. Moeslund, Adrian Hilton, and Volker Krüger. A survey of advances in vision-based human motion capture and analysis. Computer Vision and Image Understanding: CVIU, 104 (2–3):90–126, November/ December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mikhnevich:2014:UVH

[MHL14] Maxim Mikhnevich, Patrick Hébert, and Denis Laurendeau. Unsupervised visual hull extraction in space, time and light domains. Computer Vision and Image Understanding: CVIU, 125(??):55– 71, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214000551.

Medrano:2009:MFA

[MHMO09] C. Medrano, J. E. Herrero, J. Martínez, and C. Orrite. Mean field approach for tracking similar objects. Computer Vision and Image Understanding: CVIU, 113 (8):907–920, August 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Medeiros:2010:PHB

[MHSP10] Henry Medeiros, Germán Holguín, Paul J. Shin, and Johnny Park. A parallel histogram-based particle filter for object tracking on SIMD-based smart cameras. Computer Vision and Image Understanding: CVIU, 114(11):1264–1272, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mai:2019:AMN

[MHX19] Sijie Mai, Haifeng Hu, and Jia Xu. Attentive matching network for few-shot learning. Computer Vision and Image Understanding: CVIU, 187(??):Article 102781, October 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1077314219300979.

Mignotte:2012:MBS

[Mig12]  Max Mignotte.  MDS-

based segmentation model for the fusion of contour and texture cues in natural images. Computer Vision and Image Understanding: CVIU, 116(9):981–990, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421200077X.

Miller:1999:ATI

[Mil99] Erik G. Miller. Alternative tilings for improved surface area estimates by local counting algorithms. Computer Vision and Image Understanding: CVIU, 74(3): 193–211, June 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0754/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0754/production/pdf.

Mille:2009:NBR

[Mil09] Julien Mille. Narrow band region-based active contours and surfaces for 2D and 3D segmentation. Computer Vision and Image Understanding: CVIU, 113(9): 946–965, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Meshgi:2016:OAP

[MiMO^{+}16] Kourosh Meshgi, Shin ichi Maeda, Shigeyuki Oba, Henrik Skibbe, Yu zhe Li, and Shin Ishii. An occlusion-aware particle filter tracker to handle complex and persistent occlusions. Computer Vision and Image Understanding: CVIU, 150(??):81–94, September 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216300649.

Moyou:2016:LSD

[MIP16] Mark Moyou, Koffi Eddy Ihou, and Adrian M. Peter. LBO-Shape densities: a unified framework for 2D and 3D shape classification on the hypersphere of wavelet densities. Computer Vision and Image Understanding: CVIU, 152(??):142– 154, November 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301060.

Mironica:2016:FKT

[MIUS16] Ionut Mironica, Bogdan Ionescu, Jasper Uijlings, and Nicu Sebe. Fisher kernel temporal variationbased relevance feedback for video retrieval. Computer Vision and Image Understanding: CVIU, 143(??): 38–51, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002155.

Martinez:2011:TMG

[MJ11] Antonio Martínez and Juan José Jiménez. Tracking by means of geodesic region models applied to multidimensional and complex medical images. Computer Vision and Image Understanding: CVIU, 115(8):1083– 1098, August 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000932.

Mozafari:2017:CBA

[MJ17] Azadeh Sadat Mozafari and Mansour Jamzad. Clusterbased adaptive SVM: a latent subdomains discovery method for domain adaptation problems. Computer Vision and Image Understanding: CVIU, 162(??):116– 134, September 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301145.

McKenna:2000:TGP

[MJD^{+}00] Stephen J. McKenna, Sumer Jabri, Zoran Duric, Azriel Rosenfeld, and Harry Wechsler. Tracking groups of people. Computer Vision and Image Understanding: CVIU, 80(1):42– 56, October 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0870;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0870/pdf;   http:
//www.idealibrary.com/

links/doi/10.1006/cviu.
2000.0870/ref.

Moon:2016:FUR

[MJPS16] Hyungil Moon, Geonhwan Ju, Seyoun Park, and Hayong Shin. 3D freehand ultrasound reconstruction using a piecewise smooth Markov random field. Computer Vision and Image Understanding: CVIU, 151(??): 101–113, October 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002738.

Mann:1997:CPS

[MJS97] Richard Mann, Allan Jepson, and Jeffrey Mark Siskind. The computational perception of scene dynamics. Computer Vision and Image Understanding: CVIU, 65(2):113– 128, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0576/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0576/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0576/production/ref.

Medasani:2001:CID

[MK01] Swarup Medasani and Raghu Krishnapuram. Categorization of image databases for efficient retrieval using robust mixture decomposition. Computer Vision and Image Understanding: CVIU, 83(3): 216–235, September 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0926;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0926/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0926/ref.

Mahpod:2018:KVU

[MK18] Shahar Mahpod and Yosi Keller. Kinship verification using multiview hybrid distance learning. Computer Vision and Image Understanding: CVIU, 167(??): 28–36, February 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421730228X.

Moreshet:2024:ABM

[MK24] Aviad Moreshet and Yosi Keller. Attention-based multimodal image matching. Computer Vision and Image Understanding: CVIU, 241(??):??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000304.

Morales:2015:IPR

[MKF15] Aythami Morales, Ajay Kumar, and Miguel A. Ferrer. Interdigital palm region for biometric identification. Computer Vision and Image Understanding: CVIU, 142(??):125– 133, ???? 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421500199X.

Matas:2002:MNS

[MKK02] J. Matas, D. Koubaroulis, and J. Kittler. The multimodal neighborhood signature for modeling object color appearance and applications in object recognition and image retrieval. Computer Vision and Image Understanding: CVIU, 88(1):1–23, October 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mokhtarian:2001:CCF

[MKY01] F. Mokhtarian, N. Khalili, and P. Yuen. Curvature computation on free-form 3-D meshes at multiple scales. Computer Vision and Image Understanding: CVIU, 83(2):118–139, August 2001.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0919;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0919/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0919/ref.

Ma:2013:HSI

[ML13] Jun Ma and Le Lu. Hierarchical segmentation and identification of thoracic vertebra using learningbased edge detection and coarse-to-fine deformable model. Computer Vision and Image Understanding: CVIU, 117(9):1072–1083, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000763.

Ma:2015:TDH

[ML15] Chao Ma and Chuancai Liu. Two dimensional hashing for visual tracking. Computer Vision and Image Understanding: CVIU, 135 (??):83–94, June 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421500017X.

Mader:2018:DLS

[MLB^{+}18] Alexander Oliver Mader, Cristian Lorenz, Martin Bergtholdt, Jens von Berg, Hauke Schramm, Jan Modersitzki, and Carsten Meyer. Detection and localization of spatially correlated point landmarks in medical images using an automatically learned conditional random field. Computer Vision and Image Understanding: CVIU, 176– 177(??):45–53, November/ December 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421830300X.

Mekonnen:2013:CPT

[MLH13] A. A. Mekonnen, F. Lerasle, and A. Herbulot. Cooperative passers-by tracking with a mobile robot and external cameras. Computer Vision and Image Understanding: CVIU, 117(10): 1229–1244, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314212002044.

Min:2020:CEP

[MLJC20] Lihua Min, Zhenhua Li, Zhengmeng Jin, and Qiang Cui. Color edge preserving image colorization with a coupled natural vectorial total variation. Computer Vision and Image Understanding: CVIU, 196(??):Article 102981, July 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300576.

Murrugarra-Llerena:2021:IRM

[MLK21] Nils Murrugarra-Llerena and Adriana Kovashka. Image retrieval with mixed initiative and multimodal feedback. Computer Vision and Image Understanding: CVIU, 207(??): ??, June 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221000485.

Marcelli:1997:SIC

[MLP97] Angelo Marcelli, Natasha Likhareva, and Theo Pavlidis. Structural indexing for character recognition. Computer Vision and Image Understanding: CVIU, 66(3): 330–346, June 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0518/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0518/production/pdf;

http://www.idealibrary.
com/links/artid/cviu.1996.
0518/production/ref.

Mashrur:2024:RVQ

[MLZRK24] Akib Mashrur, Wei Luo, Nayyar A. Zaidi, and Antonio Robles-Kelly. Robust visual question answering via semantic cross modal augmentation. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002424.

Morimoto:2005:EGT

[MM05] Carlos H. Morimoto and Marcio R. M. Mimica. Eye gaze tracking techniques for interactive applications. Computer Vision and Image Understanding: CVIU, 98(1):4–24, April 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mokhtarian:2006:PEC

[MM06] Farzin Mokhtarian and Farahnaz Mohanna. Performance evaluation of corner detectors using consistency and accuracy measures. Computer Vision and Image Understanding: CVIU, 102(1):81–94, April 2006. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Melanitis:2021:LMC

[MM21] Nikos Melanitis and Petros Maragos. A linear method for camera pair self-calibration. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000679.

Maschino:2006:JRA

[MMA06] Emeric Maschino, Yves Maurin, and Philippe Andrey. Joint registration and averaging of multiple 3D anatomical surface models. Computer Vision and Image Understanding: CVIU, 101(1):16–30, January 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Monteiro:2018:DRA

[MMBG18] Nuno Barroso Monteiro, Simão Marto, João Pedro Barreto, and José Gaspar. Depth range accuracy for plenoptic cameras. Computer Vision and Image Understanding: CVIU, 168(??):104– 117, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300109.

Martinez:2004:CGP

[MMK04] Aleix M. Martínez, Pradit Mittrapiyanuruk, and Avinash C. Kak. On combining graph-partitioning with non-parametric clustering for image segmentation. Computer Vision and Image Understanding: CVIU, 95(1):72–85, July 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ma:2016:LGC

[MML^{+}16a] Miao Ma, Naresh Marturi, Yibin Li, Rustam Stolkin, and Ales Leonardis. A local-global coupled-layer puppet model for robust online human pose tracking. Computer Vision and Image Understanding: CVIU, 153(??):163–178, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301187.

Mollaret:2016:MMP

[MML^{+}16b] C. Mollaret, A. A. Mekonnen, F. Lerasle, I. Ferrané, J. Pinquier, B. Boudet, and P. Rumeau. A multimodal perception based assistive robotic system for the elderly. Computer Vision and Image Understanding: CVIU, 149(??):78– 97, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216000758.

Malladi:2023:TED

[MMLC23] Sai Phani Kumar Malladi, Jayanta Mukherjee, Mohamed-Chaker Larabi, and Santanu Chaudhury. Towards explainable deep visual saliency models. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001625.

Mittal:2009:SMC

[MMP09] Anurag Mittal, Antoine Monnet, and Nikos Paragios. Scene modeling and change detection in dynamic scenes: a subspace approach. Computer Vision and Image Understanding: CVIU, 113(1):63–79, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mishkin:2015:MFR

[MMP15] Dmytro Mishkin, Jiri Matas, and Michal Perdoch. MODS: Fast and robust method for two-view matching. Computer Vision and Image Understanding: CVIU, 141 (??):81–93, December 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314215001800.

Marchand-Maillet:1997:DCS

[MMS97] Stéphane Marchand-Maillet and Yazid M. Sharaiha. Discrete convexity, straightness, and the 16-neighborhood. Computer Vision and Image Understanding: CVIU, 66(3):316–329, June 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0521/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0521/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0521/production/ref.

Marchand-Maillet:1999:EOC

[MMS99] Stéphane Marchand-Maillet and Yazid M. Sharaiha. Euclidean ordering via chamfer distance calculations. Computer Vision and Image Understanding: CVIU, 73(3): 404–413, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.

                                        0743/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0743/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0743/production/ref.

Mansouri:2006:MCL

[MMV06] Abdol-Reza Mansouri, Amar Mitiche, and Carlos Vázquez. Multiregion competition: a level set extension of region competition to multiple region image partitioning. Computer Vision and Image Understanding: CVIU, 101 (3):137–150, March 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Madjidi:2006:RLA

[MN06] Hossein Madjidi and Shahriar Negahdaripour. On robustness and localization accuracy of optical flow computation for underwater color images. Computer Vision and Image Understanding: CVIU, 104(1):61– 76, October 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

McCane:2001:BOF

[MNCG01] B. McCane, K. Novins, D. Crannitch, and B. Galvin. On benchmarking optical flow. Computer Vision and Image Understanding: CVIU, 84(1):126– 143, October 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0930;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0930/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0930/ref.

Maki:2000:ASS

[MNE00] Atsuto Maki, Peter Nordlund, and Jan-Olof Eklundh. Attentional scene segmentation: Integrating depth and motion. Computer Vision and Image Understanding: CVIU, 78 (3):351–373, June 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0840;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0840/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0840/ref.

Meribout:2000:HTA

[MNHO00] Mahmoud Meribout, Mamoru Nakanishi, Eiichi Hosoya, and Takeshi Ogura. Hough transform algorithm for three-dimensional segment extraction and its parallel hardware implementation.

Computer Vision and Image Understanding: CVIU, 78(2):177–205, May 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2000.0834;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0834/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0834/ref.

Mai:2017:ELS

[MNL^{+}17] Tien-Dung Mai, Thanh Duc Ngo, Duy-Dinh Le, Duc Anh Duong, Kiem Hoang, and Shin’ichi Satoh. Efficient large-scale multi-class image classification by learning balanced trees. Computer Vision and Image Understanding: CVIU, 156(??): 151–161, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301631.

Medathati:2016:BIC

[MNMK16] N. V. Kartheek Medathati, Heiko Neumann, Guillaume S. Masson, and Pierre Kornprobst. Bioinspired computer vision: Towards a synergistic approach of artificial and biological vision. Computer Vision and Image Understanding: CVIU, 150(??):1– 30, September 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300339.

T:2018:GHQ

[MNR18] Nimisha T. M., Rajagopalan A.

N., and Aravind R. Generating high quality pan-shots from motion blurred videos. Computer Vision and Image Understanding: CVIU, 171 (??):20–33, June 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300742.

Michel:1998:GAT

[MNSK98] J. D. Michel, N. Nandhakumar, Tushar Saxena, and Deepak Kapur. Geometric, algebraic, and thermophysical techniques for object recognition in IR imagery. Computer Vision and Image Understanding: CVIU, 72(1):84– 97, October 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0669/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0669/production/pdf;
http://www.idealibrary.

                                        com/links/artid/cviu.1997.
                                        0669/production/ref.

McGuinness:2011:TAE

[MO11] Kevin McGuinness and Noel E. O’Connor. Toward automated evaluation of interactive segmentation. Computer Vision and Image Understanding: CVIU, 115(6):868–884, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mateo:2014:BPR

[MOB14] X. Mateo, X. Orriols, and X. Binefa. Bayesian perspective for the registration of multiple 3D views. Computer Vision and Image Understanding: CVIU, 118(??):84–96, January 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213001732.

Mokhtarian:1997:TMT

[Mok97] Farzin Mokhtarian. A theory of multiscale, torsionbased shape representation for space curves. Computer Vision and Image Understanding: CVIU, 68(1): 1–17, October 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.

0544/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0544/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0544/production/ref.

Mahmoudabadi:2017:DSM

[MOT17] Hamid Mahmoudabadi, Michael J.

Olsen, and Sinisa Todorovic. Detecting sudden moving objects in a series of digital images with different exposure times. Computer Vision and Image Understanding: CVIU, 158(??): 17–30, May 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300048.

Magee:2009:CVB

[MP09a] Derek Magee and Janez Pers. Computer vision based analysis in sport environments. Computer Vision and Image Understanding: CVIU, 113(5):589, May 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Montoliu:2009:GLS

[MP09b] R. Montoliu and F. Pla.

Generalized least squaresbased parametric motion estimation. Computer Vision and Image Understanding: CVIU, 113(7):790– 801, July 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Maddalena:2014:AMO

[MP14] Lucia Maddalena and Alfredo Petrosino. The 3dSOBS+ algorithm for moving object detection. Computer Vision and Image Understanding: CVIU, 122 (??):65–73, May 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213002269.

Mukherjee:2020:SSS

[MP20] Souvick Mukherjee and Saurabh Prasad. A spatialspectral semisupervised deep learning framework using Siamese networks and angular loss. Computer Vision and Image Understanding: CVIU, 194(??):Article 102943, May 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218303011.

MacLean:2007:SIS

[MPF07] W. James MacLean, Nikos Paragios, and David Fleet. Special issue on spatial coherence for visual motion analysis. Computer Vision and Image Understanding: CVIU, 108(3):205–206, December 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Monnier:2024:SFD

[MPK24] Quentin Monnier, Tania Pouli, and Kidiyo Kpalma. Survey on fast dense video segmentation techniques. Computer Vision and Image Understanding: CVIU, 241(??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000407.

Martinel:2016:SEL

[MPM16] Niki Martinel, Claudio Piciarelli, and Christian Micheloni. A supervised extreme learning committee for food recognition. Computer Vision and Image Understanding: CVIU, 148 (??):67–86, July 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000436.

Morse:1998:ZIV

[MPPG98] B. S. Morse, S. M. Pizer, D. T. Puff, and C. Gu. Zoom-invariant vision of figural shape. effects on cores of image disturbances. Computer Vision and Image Understanding: CVIU, 69(1): 72–??, ???? 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Moreno:2014:FGC

[MPPP14] Juan C. Moreno, V. B. Surya Prasath, Hugo Proença, and K. Palaniappan. Fast and globally convex multiphase active contours for brain MRI segmentation. Computer Vision and Image Understanding: CVIU, 125(??):237– 250, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214000976.

Marcon:2008:FPA

[MPST08] Marco Marcon, Luca Piccarreta, Augusto Sarti, and Stefano Tubaro. Fast PDE approach to surface reconstruction from large cloud of points. Computer Vision and Image Understanding: CVIU, 112(3):274–285, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mayer:2021:AFD

[MPT21] Christoph Mayer, Matthieu Paul, and Radu Timofte. Adversarial feature distribution alignment for semisupervised learning. Computer Vision and Image Understanding: CVIU, 202 (??):Article 103109, January 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220301326.

Moctezuma:2023:VCC

[MRdRGC23] Daniela Moctezuma, Tania Ramírez-delReal, Guillermo Ruiz, and Othón González-Chávez. Video captioning: a comparative review of where we are and which could be the route. Computer Vision and Image Understanding: CVIU, 231 (??):??, June 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000516.

Murino:1996:GSP

[MRF96] Vittorio Murino, Carlo S.

Regazzoni, and Gian Luca Foresti. Grouping as a searching process for minimum-energy configurations of labelled random fields. Computer Vision and Image Understanding: CVIU, 64(1): 157–174, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0051/production;   http:
//www.idealibrary.com/

links/artid/cviu.1996.
0051/production/pdf.

Maas:2019:LNT

[MRH19] Alina E. Maas, Franz Rottensteiner, and Christian Heipke. A label noise tolerant random forest for the classification of remote sensing data based on outdated maps for training. Computer Vision and Image Understanding: CVIU, 188(??):Article 102782, November 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314219300980.

Mair:1997:SAD

[MRW^{+}97] Bernard A. Mair, Zoltán Réti, David C. Wilson, Edward A. Geiser, and Bryn David. A q-series approach to deblurring the discrete Gaussian. Computer Vision and Image Understanding: CVIU, 66(2): 247–254, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0609/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0609/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0609/production/ref.

Ma:1996:FPT

[MS96a] C. Min Ma and Milan Sonka. A fully parallel 3D thinning algorithm and its applications. Computer Vision and Image Understanding: CVIU, 64(3):420– 433, November 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0069/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0069/production/pdf.

Murray:1996:DUP

[MS96b] D. W. Murray and L. S.

Shapiro. Dynamic updating of planar structure and motion: The case of constant motion. Computer Vision and Image Understanding: CVIU, 63(1):169– 181, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0012/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0012/production/pdf.

Malassiotis:1997:MBJ

[MS97a] Sotiris Malassiotis and Michael G. Strintzis. Modelbased joint motion and structure estimation from stereo images. Computer Vision and Image Understanding: CVIU, 65(1):79– 94, January 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0481/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0481/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0481/production/ref.

Maxwell:1997:PBS

[MS97b] Bruce A. Maxwell and Steven A. Shafer. Physicsbased segmentation of complex objects using multiple hypotheses of image formation. Computer Vision and Image Understanding: CVIU, 65(2):269– 295, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0573/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0573/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0573/production/ref.

Maxwell:2000:SIM

[MS00] Bruce A. Maxwell and Steven A. Shafer. Segmentation and interpretation of multicolored objects with highlights. Computer Vision and Image Understanding: CVIU, 77(1):1– 24, January 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0801/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0801/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0801/production/ref.

Minetto:2012:IVS

[MSF^{+}12] R. Minetto, T. V. Spina, A. X. Falcão, N. J. Leite, J. P. Papa, and J. Stolfi. IFTrace: Video segmentation of deformable objects using the Image Foresting Transform. Computer Vision and Image Understanding: CVIU, 116(2):274–291, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002141.

Malmir:2017:DAO

[MSF^{+}17] Mohsen Malmir, Karan Sikka, Deborah Forster, Ian Fasel, Javier R. Movellan, and Garrison W. Cottrell. Deep active object recognition by joint label and action prediction. Computer Vision and Image Understanding: CVIU, 156(??): 128–137, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301655.

Mallik:2010:CBI

[MSG10] Joyita Mallik, Ashok Samal, and Scott L. Gardner. A content based image retrieval system for a biological specimen collection. Computer Vision and Image Understanding: CVIU, 114(7):745–757, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

MacLean:2010:LCM

[MSI10] W. James MacLean, Siraj Sabihuddin, and Jamin Islam. Leveraging cost matrix structure for hardware implementation of stereo disparity computation using dynamic programming. Computer Vision and Image Understanding: CVIU, 114(11):1126–1138, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mishkin:2017:SEC

[MSM17] Dmytro Mishkin, Nikolay Sergievskiy, and Jiri Matas. Systematic evaluation of convolution neural network advances on the Imagenet. Computer Vision and Image Understanding: CVIU, 161(??): 11–19, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300814.

Murabito:2018:TSD

[MSP^{+}18] Francesca Murabito, Concetto Spampinato, Simone Palazzo, Daniela Giordano, Konstantin Pogorelov, and Michael Riegler. Top-down saliency detection driven by visual classification. Computer Vision and Image Understanding: CVIU, 172 (??):67–76, July 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300407.

Magid:2007:CGM

[MSR07] Evgeni Magid, Octavian Soldea, and Ehud Rivlin. A comparison of Gaussian and mean curvature estimation methods on triangular meshes of range image data. Computer Vision and Image Understanding: CVIU, 107(3):139–159, September 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Miura:2009:BRE

[MSSS09] Jun Miura, Takumi Shimawaki, Takuro Sakiyama, and Yoshiaki Shirai. Ball route estimation under heavy occlusion in broadcast soccer video. Computer Vision and Image Understanding: CVIU, 113(5): 653–662, May 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Meer:2000:RCV

[MST00] Peter Meer, Charles V. Stewart, and David E. Tyler. Robust computer vision: An interdisciplinary challenge. Computer Vision and Image Understanding: CVIU, 78(1): 1–7, April 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.1999.0833;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        1999.0833/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        1999.0833/ref.

Marcon:2016:TMA

[MST16] Marco Marcon, Augusto Sarti, and Stefano Tubaro. Toothbrush motion analysis to help children learn proper tooth brushing. Computer Vision and Image Understanding: CVIU, 148(??): 34–45, July 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300030.

Machineni:2020:EED

[MSV^{+}20] Rakesh Chowdary Machineni, G. E. Spoorthi, Krishna Sumanth Vengala, Subrahmanyam Gorthi, and Rama Krishna Sai S. Gorthi. End-to-end deep learning-based fringe projection framework for 3D profiling of objects. Computer Vision and Image Understanding: CVIU, 199 (??):Article 103023, October 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220300825.

Mount:1996:AOT

[MSW96] David M. Mount, Ruth Silverman, and Angela Y. Wu. On the area of overlap of translated polygons. Computer Vision and Image Understanding: CVIU, 64(1): 53–61, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0045/production;   http:
//www.idealibrary.com/

                                        links/artid/cviu.1996.
                                        0045/production/pdf.

Medeiros:2015:ISM

[MSW15] R. S. Medeiros, J. Scharcanski, and A. Wong. Image segmentation via multi-scale stochastic regional texture appearance models. Computer Vision and Image Understanding: CVIU, 142 (??):23–36, ???? 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001253.

Metaxas:1997:GEI

[MT97] Dimitris Metaxas and Demetri Terzopoulos. GUEST EDITORS’ INTRODUC-TION. Computer Vision and Image Understanding: CVIU, 65(2):111– 112, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0593/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0593/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0593/production/ref.

Murino:2000:UCV

[MT00] Vittorio Murino and Andrea Trucco. Underwater computer vision and pattern recognition. Computer Vision and Image Understanding: CVIU, 79 (1):1–3, July 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0852;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0852/pdf.

M:2016:RHF

[MT16] José Oramas M. and Tinne Tuytelaars. Recovering hard-to-find object instances by sampling context-based object proposals. Computer Vision and Image Understanding: CVIU, 152(??): 118–130, November 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314216301151.

Morales:2011:RTA

[MTAA11] Néstor Morales, Jonay T.

Toledo, Leopoldo Acosta, and Rafael Arnay. Realtime adaptive obstacle detection based on an image database. Computer Vision and Image Understanding: CVIU, 115(9):1273– 1287, September 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001263.

Minetto:2014:STD

[MTC^{+}14] Rodrigo Minetto, Nicolas Thome, Matthieu Cord, Neucimar J. Leite, and Jorge Stolfi. SnooperText: a text detection system for automatic indexing of urban scenes. Computer Vision and Image Understanding: CVIU, 122(??): 92–104, May 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213001914.

Mancas-Thillou:2007:CTE

[MTG07] Céline Mancas-Thillou and Bernard Gosselin. Color text extraction with selective metric-based clustering. Computer Vision and Image Understanding: CVIU, 107 (1–2):97–107, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Moeslund:2017:CVS

[MTH^{+}17] Thomas B. Moeslund, Graham Thomas, Adrian Hilton, Peter Carr, and Irfan Essa. Computer vision in sports. Computer Vision and Image Understanding: CVIU, 159 (??):1–2, June 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300838.

Mayer:2021:TCG

[MTP21] Christoph Mayer, Radu Timofte, and Grégory Paul. Towards closing the gap in weakly supervised semantic segmentation with DC-NNs: Combining local and global models. Computer Vision and Image Understanding: CVIU, 208–209 (??):??, July 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000539.

Muthu:2023:GFI

[MTR^{+}23] Sundaram Muthu, Ruwan Tennakoon, Tharindu Rathnayake, Reza Hoseinnezhad, David Suter, and Alireza Bab-Hadiashar. Generalized framework for image and video object segmentation using affinity learning and message passing GNNS. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001923.

Mettes:2017:WDT

[MTV17] Pascal Mettes, Robby T.

Tan, and Remco C. Veltkamp.

Water detection through spatio-temporal invariant descriptors. Computer Vision and Image Understanding: CVIU, 154(??):182– 191, January 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300273.

Mindru:2004:MIR

[MTVM04] Florica Mindru, Tinne Tuytelaars, Luc Van Gool, and Theo Moons. Moment invariants for recognition under changing viewpoint and illumination. Computer Vision and Image Understanding: CVIU, 94(1– 3):3–27, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mikolajczyk:2011:ARA

[MU11] K. Mikolajczyk and H. Uemura. Action recognition with appearance– motion features and fast search trees. Computer Vision and Image Understanding: CVIU, 115(3):426– 438, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mukawa:1997:OMB

[Muk97] Naoki Mukawa. Opticalmodel-based analysis of consecutive images. Computer Vision and Image Understanding: CVIU, 66(1): 25–32, April 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0500/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0500/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0500/production/ref.

Munkelt:1995:ATG

[Mun95]  Olaf Munkelt.  Aspect-

trees: Generation and interpretation. Computer Vision and Image Understanding: CVIU, 61(3): 365–386, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1029/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1029/production/pdf.

Murray:1995:RRU

[Mur95] David W. Murray. Recovering range using virtual multicamera stereo. Computer Vision and Image Understanding: CVIU, 61(2): 285–291, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1021/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1021/production/pdf.

Madabhushi:2006:GST

[MUS06] Anant Madabhushi, Jayaram K. Udupa, and Andre Souza. Generalized scale: Theory, algorithms, and application to image inhomogeneity correction. Computer Vision and Image Understanding: CVIU, 101 (2):100–121, February 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mettes:2016:NSP

[MvGS16] Pascal Mettes, Jan C. van Gemert, and Cees G. M. Snoek. No spare parts: Sharing part detectors for image categorization. Computer Vision and Image Understanding: CVIU, 152(??):131– 141, November 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301035.

Macaire:2006:CIS

[MVP06] Ludovic Macaire, Nicolas Vandenbroucke, and Jack-Gérard Postaire. Color image segmentation by analysis of subset connectedness and color homogeneity properties. Computer Vision and Image Understanding: CVIU, 102(1):105– 116, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ma:2000:PTA

[MW00] Cherng-Min Ma and Shu-Yen Wan. Parallel thinning algorithms on 3D (18, 6) binary images. Computer Vision and Image Understanding: CVIU, 80(3): 364–378, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0879;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0879/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0879/ref.

Metternich:2013:TBR

[MW13] Michael J. Metternich and Marcel Worring. Track based relevance feedback for tracing persons in surveillance videos. Computer Vision and Image Understanding: CVIU, 117(3): 229–237, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001889.

Ma:2022:SOR

[MW22] Yunlong Ma and Chunyan Wang. SdcNet for object recognition. Computer Vision and Image Understanding: CVIU, 215(??): ??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221001703.

Ma:2007:FVP

[MWF07] Zhonghua Ma, Hong Ren Wu, and Dagan Feng. Fuzzy vector partition filtering technique for color image restoration. Computer Vision and Image Understanding: CVIU, 107(1–2):26–37, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Milun:1999:GRB

[MWL99] Elyse H. Milun, Deborah K. W. Walters, and Yiming Li. General ribbonbased thinning algorithms for stylus-generated images. Computer Vision and Image Understanding: CVIU, 76(3):267–277, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0807/production;   http:
                                        //www.idealibrary.com/

links/artid/cviu.1999.
0807/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0807/production/ref.

Meng:2024:DTE

[MWL^{+}24] Xiangyin Meng, Jie Wen, Yang Li, Chenlong Wang, and Jingzhen Zhang. DFNet-Trans: an end-to-end multibranching network for depth estimation for transparent objects. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002941.

Milun:1999:GRM

[MWLA99] Elyse H. Milun, Deborah K. W. Walters, Yiming Li, and Bemina Atanacio. General ribbons: a model for stylus-generated images. Computer Vision and Image Understanding: CVIU, 76(3):259– 266, December 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0806/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0806/production/pdf;
http://www.idealibrary.

                                        com/links/artid/cviu.1999.
                                        0806/production/ref.

Matsuyama:2004:RTS

[MWTN04] T. Matsuyama, X. Wu, T. Takai, and S. Nobuhara. Real-time 3D shape reconstruction, dynamic 3D mesh deformation, and high fidelity visualization for 3D video. Computer Vision and Image Understanding: CVIU, 96(3):393–434, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Masuda:1995:RMR

[MY95] Takeshi Masuda and Naokazu Yokoya. A robust method for registration and segmentation of multiple range images. Computer Vision and Image Understanding: CVIU, 61(3): 295–307, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1024/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1024/production/pdf.

Ma:2009:TTI

[MYC09] Yunqian Ma, Qian Yu, and Isaac Cohen. Target tracking with incomplete detection. Computer Vision and Image Understanding: CVIU, 113(4):580– 587, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mu:2014:GES

[MYC^{+}14] Yadong Mu, Yi Yang, Liangliang Cao, Shuicheng Yan, and Qi Tian. Guest editorial: Special issue on large scale multimedia semantic indexing. Computer Vision and Image Understanding: CVIU, 124(??): 1–2, July 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001052.

Martinez:2003:SIF

[MYK03] Aleix M. Martínez, Ming-Hsuan Yang, and David J. Kriegman. Special issue on face recognition. Computer Vision and Image Understanding: CVIU, 91(1– 2):1–5, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mirhosseini:1998:HFI

[MYLP98] Ali Reza Mirhosseini, Hong Yan, Kin-Man Lam, and Tuan Pham. Human face image recognition: An evidence aggregation approach. Computer Vision and Image Understanding: CVIU, 71(2):213– 230, August 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0710/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0710/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0710/production/ref.

Murtaza:2019:TTA

[MYV19] Fiza Murtaza, Muhammad Haroon Yousaf, and Sergio A. Velastin. TAB: Temporally aggregated bagof-discriminant-words for temporal action proposals. Computer Vision and Image Understanding: CVIU, 183 (??):42–52, June 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219300621.

Ma:2017:LNR

[MYYY17] Chao Ma, Chih-Yuan Yang, Xiaokang Yang, and Ming-Hsuan Yang. Learning a no-reference quality metric for single-image superresolution. Computer Vision and Image Understanding: CVIU, 158(??): 1–16, May 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421630203X.

Ma:2024:NSM

[MYZ^{+}24] Guoyuan Ma, Xiaofeng Yue, Juan Zhu, Zeyuan Liu, Zongheng Zhang, Yuan Zhou, and Chang Li. A novel slime mold algorithm for grayscale and color image contrast enhancement. Computer Vision and Image Understanding: CVIU, 240 (??):??, March 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000146.

Ma:2020:IVI

[MZ20] Jiayi Ma and Yi Zhou. Infrared and visible image fusion via gradientlet filter. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 103016, August 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220300771.

Michieli:2021:KDI

[MZ21] Umberto Michieli and Pietro Zanuttigh. Knowledge distillation for incremental learning in semantic segmentation. Computer Vision and Image Understanding: CVIU, 205(??):Article 103167, April 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000114.

Marzotto:2010:RTV

[MZB^{+}10] Roberto Marzotto, Paul Zoratti, Daniele Bagni, Andrea Colombari, and Vittorio Murino. A real-time versatile roadway path extraction and tracking on an FPGA platform. Computer Vision and Image Understanding: CVIU, 114(11): 1164–1179, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Martin:2005:HIO

[MZC^{+}05] José L. Martín, Aitzol Zuloaga, Carlos Cuadrado, Jesús Lázaro, and Unai Bidarte. Hardware implementation of optical flow constraint equation using FPGAs. Computer Vision and Image Understanding: CVIU, 98(3):462–490, June 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mei:2016:SIV

[MZL^{+}16] Xue Mei, Tianzhu Zhang, Huchuan Lu, Ming-Hsuan Yang, Kyoung Mu Lee, and Horst Bischof. Special issue on visual tracking. Computer Vision and Image Understanding:

CVIU, 153(??):1–2, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301746.

Naiel:2017:OMO

[NAS^{+}17] Mohamed A. Naiel, M. Omair Ahmad, M. N. S. Swamy, Jongwoo Lim, and Ming-Hsuan Yang. Online multiobject tracking via robust collaborative model and sample selection. Computer Vision and Image Understanding: CVIU, 154(??): 94–107, January 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300996.

Nichau:2010:PIN

[NB10] Marco Nichau and Volker Blanz. Pose-insensitive nose detection in TOFscans. Computer Vision and Image Understanding: CVIU, 114(12):1346–1352, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nabati:2020:VCU

[NB20] Masoomeh Nabati and Alireza Behrad. Video captioning using boosted and parallel long short-term memory networks. Computer Vision and Image Understanding: CVIU, 190 (??):Article 102840, January 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218301632.

Nunes:2004:MER

[NBDB04] J. C. Nunes, Y. Bouaoune, E. Delechelle, and Ph. Bunel. A multiscale elastic registration scheme for retinal angiograms. Computer Vision and Image Understanding: CVIU, 95(2):129– 149, August 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ngoc:2020:MBD

[NBFG20] Minh Ôn Vũ Ngo.c, Nicolas Boutry, Jonathan Fabrizio, and Thierry Géraud. A minimum barrier distance for multivariate images with applications. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 102993, August 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220300631.

Nezami:2021:POA

[NCDG21] Omid Mohamad Nezami, Akshay Chaturvedi, Mark Dras, and Utpal Garain.

Pick-Object-Attack: Typespecific adversarial attack for object detection. Computer Vision and Image Understanding: CVIU, 211 (??):??, October 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001016.

Nesi:1995:RAO

[NDBT95] P. Nesi, A. Del Bimbo, and D. Ben-Tzvi. A robust algorithm for optical flow estimation. Computer Vision and Image Understanding: CVIU, 62(1): 59–68, July 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1041/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1041/production/pdf.

Niessen:1997:MAI

[NDN^{+}97] W. J. Niessen, J. S. Duncan, M. Nielsen, L. M. J. Florack, B. M. ter Haar Romeny, and M. A. Viergever. A multiscale approach to image sequence analysis. Computer Vision and Image Understanding: CVIU, 65(2):259– 268, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0582/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0582/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0582/production/ref.

Noceti:2009:STC

[NDO09] Nicoletta Noceti, Elisabetta Delponte, and Francesca Odone. Spatio-temporal constraints for on-line 3D object recognition in videos. Computer Vision and Image Understanding: CVIU, 113(12):1198–1209, December 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Negahdaripour:1996:DCF

[Neg96] Shahriar Negahdaripour. Direct computation of the FOE with confidence measures. Computer Vision and Image Understanding: CVIU, 64(3):323–350, November 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0063/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0063/production/pdf.

Negahdaripour:2012:VMA

[Neg12] S. Negahdaripour. Visual motion ambiguities of a plane in 2-D FS sonar motion sequences. Computer Vision and Image Understanding: CVIU, 116(6): 754–764, June 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000379.

Nuchter:2010:SPR

[NESP10] Andreas Nüchter, Jan Elseberg, Peter Schneider, and Dietrich Paulus. Study of parameterizations for the rigid body transformations of the scan registration problem. Computer Vision and Image Understanding: CVIU, 114(8):963–980, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ng:2021:WSA

[NF21] Yan Bin Ng and Basura Fernando. Weakly supervised action segmentation with effective use of attention and self-attention. Computer Vision and Image Understanding: CVIU, 213(??): ??, December 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001429.

Neumann:2004:HCP

[NFA04] Jan Neumann, Cornelia Fermüller, and Yiannis Aloimonos. A hierarchy of cameras for 3D photography. Computer Vision and Image Understanding: CVIU, 96(3):274–293, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nascimento:2008:IIP

[NFM08] J. Nascimento, M. Figueiredo, and J. Marques. Independent increment processes for human motion recognition. Computer Vision and Image Understanding: CVIU, 109 (2):126–138, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nguyen:2013:FDS

[NFSD13] Kien Nguyen, Clinton Fookes, Sridha Sridharan, and Simon Denman. Featuredomain super-resolution for iris recognition. Computer Vision and Image Understanding: CVIU, 117(10): 1526–1535, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314213001306.

Neuenschwander:1997:VSF

[NFSK97] W. Neuenschwander, P. Fua, G. Székely, and O. Kübler. Velcro surfaces: Fast initialization of deformable models. Computer Vision and Image Understanding: CVIU, 65(2):237– 245, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0578/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0578/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0578/production/ref.

Nagao:1998:AMP

[NG98a] Kenji Nagao and W. E. L.

Grimson. Affine matching of planar sets. Computer Vision and Image Understanding: CVIU, 70(1): 1–22, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0623/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0623/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0623/production/ref.

Nagao:1998:UPI

[NG98b] Kenji Nagao and W. Eric. L. Grimson. Using photometric invariants for 3D object recognition. Computer Vision and Image Understanding: CVIU, 71(1): 74–93, July 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0603/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0603/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0603/production/ref.

Noori:2024:DID

[NGR24] Hossein Noori, Mohammad Hossein Gholizadeh, and Hossein Khodabakhshi Rafsanjani. Digital image defogging using joint Retinex theory and independent component analysis. Computer Vision and Image Understanding: CVIU, 245 (??):??, August 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224001140.

Nam:2014:MLS

[NHH14] Woonhyun Nam, Bohyung Han, and Joon Hee Han. Macrofeature layout selection for pedestrian localization and its acceleration using GPU. Computer Vision and Image Understanding: CVIU, 120(??): 46–58, March 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213002026.

Nam:2008:SBL

[NHK08] Yunyoung Nam, Eenjun Hwang, and Dongyoon Kim. A similarity-based leaf image retrieval scheme: Joining shape and venation features. Computer Vision and Image Understanding: CVIU, 110(2):245–259, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ni:2009:UMS

[NHSC09] Kangyu Ni, Byung-Woo Hong, Stefano Soatto, and Tony Chan. Unsupervised multiphase segmentation: a recursive approach. Computer Vision and Image Understanding: CVIU, 113(4): 502–510, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Niu:2015:VTN

[NHTG15] Zhenxing Niu, Gang Hua, Qi Tian, and Xinbo Gao. Visual Topic Network: Building better image representations for images in social media. Computer Vision and Image Understanding: CVIU, 136(??): 3–13, July 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215000259.

Nejhum:2010:OVT

[NHY10] S. M. Shahed Nejhum, Jeffrey Ho, and Ming-Hsuan Yang. Online visual tracking with histograms and articulating blocks. Computer Vision and Image Understanding: CVIU, 114(8):901–914, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Niu:2022:HUM

[NHZ^{+}22] Li Niu, Shengyuan Huang, Xing Zhao, Liwei Kang, Yiyi Zhang, and Liqing Zhang. Hallucinating uncertain motion and future for static image action recognition. Computer Vision and Image Understanding: CVIU, 215 (??):??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221001739.

Nicol:1995:SAR

[Nic95] C. J. Nicol. Systolic approach for real time connected component labeling. Computer Vision and Image Understanding: CVIU, 61(1):17–31, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1002/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1002/production/pdf.

Nishida:1995:SFE

[Nis95] Hirobumi Nishida. Structural feature extraction using multiple bases. Computer Vision and Image Understanding: CVIU, 62(1): 78–89, July 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1043/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1043/production/pdf.

Nishida:1996:SRI

[Nis96] Hirobumi Nishida. Shape recognition by integrating structural descriptions and geometrical/statistical transforms. Computer Vision and Image Understanding: CVIU, 64(2):248– 262, September 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0057/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0057/production/pdf.

Nishida:1997:ASD

[Nis97] Hirobumi Nishida. Analysis and synthesis of deformed patterns based on structural models. Computer Vision and Image Understanding: CVIU, 68(1): 59–71, October 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0541/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0541/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0541/production/ref.

Nishida:1999:SSI

[Nis99] Hirobumi Nishida. Structural shape indexing with feature generation models. Computer Vision and Image Understanding: CVIU, 73(1):121–136, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0712/production;   http:
                                        //www.idealibrary.com/

links/artid/cviu.1998.
0712/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0712/production/ref.

Newman:1995:SAV

[NJ95] Timothy S. Newman and Anil K. Jain. A survey of automated visual inspection. Computer Vision and Image Understanding: CVIU, 61(2):231– 262, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1017/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1017/production/pdf.

Negahdaripour:2000:MBC

[NK00] S. Negahdaripour and A. Khamene.

Motion-based compression of underwater video imagery for the operations of unmanned submersible vehicles. Computer Vision and Image Understanding: CVIU, 79(1): 162–183, July 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0845;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0845/pdf;   http:

                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0845/ref.

Noris:2011:WGT

[NKB11] Basilio Noris, Jean-Baptiste Keller, and Aude Billard. A wearable gaze tracking system for children in unconstrained environments. Computer Vision and Image Understanding: CVIU, 115(4):476–486, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nguyen:2024:STR

[NKL24] Hung Nguyen, Chanho Kim, and Fuxin Li. Space-time recurrent memory network. Computer Vision and Image Understanding: CVIU, 241(??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000249.

Ngo:2013:CSR

[NKPT13] Phuc Ngo, Yukiko Kenmochi, Nicolas Passat, and Hugues Talbot. Combinatorial structure of rigid transformations in 2D digital images. Computer Vision and Image Understanding: CVIU, 117(4): 393–408, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001786.

Nguyen:1996:RDO

[NL96] Quang-Loc Nguyen and Martin D. Levine. Representing 3-D objects in range images using geons. Computer Vision and Image Understanding: CVIU, 63(1): 158–168, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0011/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0011/production/pdf.

Nguyen:2017:SCO

[NL17] Thanh-Tin Nguyen and Maxime Lhuillier. Selfcalibration of omnidirectional multi-cameras including synchronization and rolling shutter. Computer Vision and Image Understanding: CVIU, 162(??): 166–184, September 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314217301492.

Nam:2023:RIF

[NL23] Ju-Hyeon Nam and Sang-Chul Lee. Random image frequency aggregation dropout in image classification for deep convolutional neural networks. Computer Vision and Image Understanding: CVIU, 232(??): ??, July 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223000644.

Noureddin:2005:NCD

[NLM05] B. Noureddin, P. D. Lawrence, and C. F. Man. A noncontact device for tracking gaze in a human computer interface. Computer Vision and Image Understanding: CVIU, 98(1):52– 82, April 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Niessen:2013:SIS

[NLW13] Wiro Niessen, Shuo Li, and Song Wang. Special issue on Shape Modeling in Medical Image Analysis. Computer Vision and Image Understanding: CVIU, 117(9): 965, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213001227.

Nian:2017:LEV

[NLW^{+}17] Fudong Nian, Teng Li, Yan Wang, Xinyu Wu, Bingbing Ni, and Changsheng Xu. Learning explicit video attributes from mid-level representation for video captioning. Computer Vision and Image Understanding: CVIU, 163(??):126– 138, October 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301327.

Ni:2023:HOI

[NMAL23] Zhifan Ni, Esteve Valls Mascaró, Hyemin Ahn, and Dongheui Lee. Human– object interaction prediction in videos through gaze following. Computer Vision and Image Understanding: CVIU, 233(??): ??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001212.

Nastar:1997:FIM

[NMP97] Chahab Nastar, Baback Moghaddam, and Alex Pentland. Flexible images: Matching and recognition using learned deformations. Computer Vision and Image Understanding: CVIU, 65(2):179– 191, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0583/production;   http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0583/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0583/production/ref.

Nikolaev:2004:LCS

[NN04] Dmitry P. Nikolaev and Petr P. Nikolayev. Linear color segmentation and its implementation. Computer Vision and Image Understanding: CVIU, 94(1–3): 115–139, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Natarajan:2013:HMC

[NN13] Pradeep Natarajan and Ramakant Nevatia. Hierarchical multi-channel hidden semi Markov graphical models for activity recognition. Computer Vision and Image Understanding: CVIU, 117(10):1329–1344, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421200166X.

Nguyen:2018:SMU

[NN18] Thanh Phuong Nguyen and Xuan Son Nguyen. Shape measurement using LIPsignature. Computer Vision and Image Understanding: CVIU, 171(??):83– 94, June 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218300699.

Nguyen:2020:MDP

[NNBN20] Thanh Tuan Nguyen, Thanh Phuong Nguyen, Frédéric Bouchara, and Xuan Son Nguyen. Momental directional patterns for dynamic texture recognition. Computer Vision and Image Understanding: CVIU, 194(??):Article 102882, May 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218302728.

Nguyen:2022:DLD

[NNN^{+}22] Thanh Thi Nguyen, Quoc Viet Hung Nguyen, Dung Tien Nguyen, Duc Thanh Nguyen, Thien Huynh-The, Saeid Nahavandi, Thanh Tam Nguyen, Quoc-Viet Pham, and Cuong M. Nguyen. Deep learning for deepfakes creation and detection: a survey. Computer Vision and Image Understanding: CVIU, 223(??): ??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001114.

Nguyen:2018:DST

[NNS^{+}18] Dung Nguyen, Kien Nguyen, Sridha Sridharan, David Dean, and Clinton Fookes. Deep spatio-temporal feature fusion with compact bilinear pooling for multimodal emotion recognition. Computer Vision and Image Understanding: CVIU, 174(??):33–42, September 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218300808.

Nitschke:2011:DCC

[NNT11] Christian Nitschke, Atsushi Nakazawa, and Haruo Takemura. Display-camera calibration using eye reflections and geometry constraints. Computer Vision and Image Understanding: CVIU, 115(6):835–853, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nordberg:2009:TT

[Nor09] Klas Nordberg. The triangulation tensor. Computer Vision and Image Understanding: CVIU, 113(9): 935–945, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nappi:2022:GEI

[NPBM22] Michele Nappi, Hugo Proença, Sambit Bakshi, and Vittorio Murino. Guest editorial introduction to the special issue on “Biometrics Based Methods for Healthcare Applications”. Computer Vision and Image Understanding: CVIU, 224(??): ??, November 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001370.

Narayanan:2016:VBA

[NPM^{+}16] Vishnu K. Narayanan, François Pasteau, Maud Marchal, Alexandre Krupa, and Marie Babel. Visionbased adaptive assistance and haptic guidance for safe wheelchair corridor following. Computer Vision and Image Understanding: CVIU, 149(??):171– 185, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000539.

Narappanawar:2011:GTB

[NRJ11] Nitin Narappanawar, B. Madhusudan Rao, and Maduri Joshi. Graph theory based segmentation of traced boundary into open and closed sub-sections. Computer Vision and Image Understanding: CVIU, 115 (11):1552–1558, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314211001676.

Nogly:1996:DTG

[NS96] Daniel Nogly and Markus Schladt. Digital topology on graphs. Computer Vision and Image Understanding: CVIU, 63(2):394– 396, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0029/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0029/production/pdf.

Noordmans:1998:DCI

[NS98] H. J. Noordmans and A. W. M. Smeulders. Detection and characterization of isolated and overlapping spots. Computer Vision and Image Understanding: CVIU, 70(1): 23–35, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0604/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0604/production/pdf;

http://www.idealibrary.
com/links/artid/cviu.1998.
0604/production/ref.

Nazare:2016:SFF

[NS16] Antonio C. Nazare, Jr. and William Robson Schwartz. A scalable and flexible framework for smart video surveillance. Computer Vision and Image Understanding: CVIU, 144(??):258– 275, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002349.

Normand:2013:MDD

[NSEA13] Nicolas Normand, Robin Strand, Pierre Evenou, and Aurore Arlicot. Minimaldelay distance transform for neighborhood-sequence distances in 2D and 3D. Computer Vision and Image Understanding: CVIU, 117(4): 409–417, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001798.

Naf:1997:VST

[NSK^{+}97] M. Näf, G. Székely, R. Kikinis, M. E. Shenton, and O. Kübler. 3D Voronoi skeletons and their usage for the characterization and recognition of 3D organ shape. Computer Vision and Image Understanding: CVIU, 66(2): 147–161, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0610/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0610/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0610/production/ref.

Negahdaripour:2010:DME

[NT10] S. Negahdaripour and Ali Taatian. 3-D motion estimation by integrating visual cues in 2-D multimodal opti-acoustic stereo sequences. Computer Vision and Image Understanding: CVIU, 114(8):928–941, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Niessen:1997:NMR

[NVWV97] Wiro J. Niessen, Koen L. Vincken, Joachim A. Weickert, and Max A. Viergever. Nonlinear multiscale representations for image segmentation. Computer Vision and Image Understanding: CVIU, 66(2): 233–245, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0614/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0614/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0614/production/ref.

Nie:2015:GRB

[NWJ15] Siqi Nie, Ziheng Wang, and Qiang Ji. A generative restricted Boltzmann machine based method for highdimensional motion data modeling. Computer Vision and Image Understanding: CVIU, 136(??):14– 22, July 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214002434.

Neverova:2017:HPE

[NWNT17] Natalia Neverova, Christian Wolf, Florian Nebout, and Graham W. Taylor. Hand pose estimation through semi-supervised and weaklysupervised learning. Computer Vision and Image Understanding: CVIU, 164 (??):56–67, November 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314217301686.

Noble:1997:CAG

[NWP97] Alison Noble, Dale Wilson, and Jean Ponce. On computing aspect graphs of smooth shapes from volumetric data. Computer Vision and Image Understanding: CVIU, 66(2): 179–192, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0615/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0615/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0615/production/ref.

Nga:2014:AER

[NY14] Do Hang Nga and Keiji Yanai. Automatic extraction of relevant video shots of specific actions exploiting Web data. Computer Vision and Image Understanding: CVIU, 118(??): 2–15, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213001598.

Nguyen:2023:DLB

[NZH^{+}23] Kien X. Nguyen, Liying Zheng, Ashley L. Hawke, Robert E. Carey, Scott P. Breloff, Kang Li, and Xi Peng. Deep learningbased estimation of wholebody kinematics from multiview images. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001601.

Oliveira:2018:BDE

[OAGN18] Saulo A. F. Oliveira, Shara S. A. Alves, João P. P. Gomes, and Ajalmar R. Rocha Neto. A bi-directional evaluation-based approach for image retargeting quality assessment. Computer Vision and Image Understanding: CVIU, 168(??): 172–181, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217302035.

Ortiz:2014:FRW

[OB14] Enrique G. Ortiz and Brian C. Becker. Face recognition for Web-scale datasets. Computer Vision and Image Understanding: CVIU, 118(??):153– 170, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001744.

Orrite:2004:ESM

[OBH04] Carlos Orrite, Sergio Blecua, and J. Elías Herrero. Erratum to “Shape matching of partially occluded curves invariant under projective transformation” [Comp. Vision Image Understanding 93 (2004) 34–64]. Computer Vision and Image Understanding: CVIU, 95(1):127, July 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See [OH04].

Ohn-Bar:2015:SSC

[OBTMT15] Eshed Ohn-Bar, Ashish Tawari, Sujitha Martin, and Mohan M. Trivedi. On surveillance for safety critical events: In-vehicle video networks for predictive driver assistance systems. Computer Vision and Image Understanding: CVIU, 134(??):130– 140, May 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214002070.

Ozden:2004:RTI

[OCVV04] Kemal Egemen Ozden, Kurt Cornelis, Luc Van Eycken, and Luc Van Gool. Reconstructing 3D trajectories of independently moving objects using generic constraints. Computer Vision and Image Understanding: CVIU, 96(3):453–471, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Okatani:1997:SRE

[OD97] Takayuki Okatani and Koichiro Deguchi. Shape reconstruction from an endoscope image by shape from shading technique for a point light source at the projection center. Computer Vision and Image Understanding: CVIU, 66(2): 119–131, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0613/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0613/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0613/production/ref.

Okatani:1999:CSG

[OD99] Takayuki Okatani and Koichiro Deguchi. Computation of the sign of the Gaussian curvature of a surface from multiple unknown illumination images without knowledge of the reflectance property. Computer Vision and Image Understanding: CVIU, 76(2):125– 134, November 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0792/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0792/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0792/production/ref.

Okatani:2001:UST

[OD01] Takayuki Okatani and Koichiro Deguchi. On uniqueness of solutions of the three-light-source photometric stereo: Conditions on illumination configuration and surface reflectance. Computer Vision and Image Understanding: CVIU, 81(2): 211–226, February 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0887;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0887/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0887/ref.

Okatani:2002:MFR

[OD02] Ikuko Shimizu Okatani and Koichiro Deguchi. A method for fine registration of multiple view range images considering the measurement error properties. Computer Vision and Image Understanding: CVIU, 87(1–3):66–77, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Oberkampf:1996:IPE

[ODD96] Denis Oberkampf, Daniel F.

DeMenthon, and Larry S. Davis. Iterative pose estimation using coplanar feature points. Computer Vision and Image Understanding: CVIU, 63(3): 495–511, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0037/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0037/production/pdf.

OramasM:2017:CBO

[ODT17] José Oramas M., Luc De Raedt, and Tinne Tuytelaars. Context-based object viewpoint estimation: a 2D relational approach. Computer Vision and Image Understanding: CVIU, 160(??):100– 113, July 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421730067X.

Olsson:2008:ISR

[OEK08] Carl Olsson, Anders P. Eriksson, and Fredrik Kahl. Improved spectral relaxation methods for binary quadratic optimization problems. Computer Vision and Image Understanding: CVIU, 112(1):3– 13, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Oblonsek:1998:FSB

[OG98] C. Oblonsek and N. Guid. A fast surface-based procedure for object reconstruction from 3D scattered points. Computer Vision and Image Understanding: CVIU, 69(2):185– 195, February 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0584/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0584/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0584/production/ref.

Oshin:2014:CRM

[OGB14] Olusegun Oshin, Andrew Gilbert, and Richard Bowden. Capturing relative motion and finding modes for action recognition in the wild. Computer Vision and Image Understanding: CVIU, 125(??):155– 171, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000848.

Oliver:2004:LRL

[OGH04] Nuria Oliver, Ashutosh Garg, and Eric Horvitz. Layered representations for learning and inferring office activity from multiple sensory channels. Computer Vision and Image Understanding: CVIU, 96(2):163– 180, November 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Orrite:2004:SMP

[OH04] Carlos Orrite and J. Elias Herrero. Shape matching of partially occluded curves invariant under projective transformation. Computer Vision and Image Understanding: CVIU, 93(1):34– 64, January 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See erratum [OBH04].

Oliver:2005:SPP

[OH05] Nuria Oliver and Eric Horvitz. Selective perception policies for guiding sensing and computation in multimodal systems: a comparative analysis. Computer Vision and Image Understanding: CVIU, 100 (1–2):198–224, October/ November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Osadchy:2008:USC

[OJRT08] Margarita Osadchy, David Jacobs, Ravi Ramamoorthi, and David Tucker. Using specularities in comparing 3D models and 2D images. Computer Vision and Image Understanding: CVIU, 111(3):275–294, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Osadchy:2004:EDU

[OK04] Margarita Osadchy and Daniel Keren. Efficient detection under varying illumination conditions and image plane rotations. Computer Vision and Image Understanding: CVIU, 93(3):245– 259, March 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Oliensis:2000:CSM

[Oli00] John Oliensis. A critique of structure-from-motion algorithms. Computer Vision and Image Understanding: CVIU, 80(2):

172–214, November 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0869;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0869/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0869/ref. See corri-

gendum [Oli01].

Oliensis:2001:CSM

[Oli01] John Oliensis. A critique of structure-from-motion algorithms: Volume 80, Number 2 (2000), pages 172– 214. Computer Vision and Image Understanding: CVIU, 84(3):407–408, December 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See [Oli00].

Olson:1999:CHT

[Ols99]  Clark F. Olson.  Con-

strained Hough transforms for curve detection. Computer Vision and Image Understanding: CVIU, 73(3): 329–345, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0728/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.

                                        0728/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0728/production/ref.

Okorie:2019:RBI

[OM19] Azubuike Okorie and Sokratis Makrogiannis. Region-based image registration for remote sensing imagery. Computer Vision and Image Understanding: CVIU, 189 (??):Article 102825, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219301304.

Ong:2006:VIE

[OMBH06] Eng-Jon Ong, Antonio S. Micilotta, Richard Bowden, and Adrian Hilton. Viewpoint invariant exemplarbased 3D human tracking. Computer Vision and Image Understanding: CVIU, 104 (2–3):178–189, November/ December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ogier:1998:MAD

[OMLL98] Jean Marc Ogier, Rémy Mullot, Jacques Labiche, and Yves Lecourtier. Multilevel approach and distributed consistency for technical map interpretation: Application to cadastral maps. Computer Vision and Image Understanding: CVIU, 70(3): 438–451, June 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0690/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0690/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0690/production/ref;
http://www.idealibrary.
com/links/artid/cviu.1998.
0708/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0708/production/pdf.

Olson:2007:VTM

[OMW^{+}07] Clark F. Olson, Larry H.

Matthies, John R. Wright, Rongxing Li, and Kaichang Di. Visual terrain mapping for Mars exploration. Computer Vision and Image Understanding: CVIU, 105(1):73–85, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Osman:2019:DDR

[OS19] Ahmed Osman and Wojciech Samek. DRAU: Dual Recurrent Attention Units for visual question answering. Computer Vision and Image Understanding: CVIU, 185(??):24–30, August 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219300761.

Ortego:2016:RBM

[OSM16] Diego Ortego, Juan C. SanMiguel, and José M. Martínez. Rejection based multipath reconstruction for background estimation in video sequences with stationary objects. Computer Vision and Image Understanding: CVIU, 147(??): 23–37, June 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421630008X.

Ortego:2017:SAQ

[OSM17] Diego Ortego, Juan C. SanMiguel, and José M. Martínez. Stand-alone quality estimation of background subtraction algorithms. Computer Vision and Image Understanding: CVIU, 162(??):87–102, September 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217301443.

Ozawa:2018:SCI

[OSY18] Keisuke Ozawa, Imari Sato, and Masahiro Yamaguchi.

Single color image photometric stereo for multicolored surfaces. Computer Vision and Image Understanding: CVIU, 171(??): 140–149, June 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300547.

Othmani:2020:AEF

[OTAH20] Alice Othmani, Abdul Rahman Taleb, Hazem Abdelkawy, and Abdenour Hadid. Age estimation from faces using deep learning: a comparative analysis. Computer Vision and Image Understanding: CVIU, 196(??):Article 102961, July 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300424.

Olson:1996:PPI

[OTL96] Thomas J. Olson, John R.

Taylor, and Robert J. Lockwood. Programming a pipelined image processor. Computer Vision and Image Understanding: CVIU, 64(3):351– 367, November 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.

                                        0064/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0064/production/pdf.

Ogiela:2006:GIL

[OTO06] Marek R. Ogiela, Ryszard Tadeusiewicz, and Lidia Ogiela. Graph image language techniques supporting radiological, hand image interpretations. Computer Vision and Image Understanding: CVIU, 103(2):112–120, August 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ozcan:2020:POD

[ÖÜ20] Abdullah H. Özcan and Cem Ünsalan. Probabilistic object detection and shape extraction in remote sensing data. Computer Vision and Image Understanding: CVIU, 195(??):Article 102953, June 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220300357.

Ozbulak:2021:ISA

[OVJ^{+}21] Utku Ozbulak, Baptist Vandersmissen, Azarakhsh Jalalvand, Ivo Couckuyt, Arnout Van Messem, and Wesley De Neve. Investigating the significance of adversarial attacks and their relation to interpretability for radar-based human activity recognition systems. Computer Vision and Image Understanding: CVIU, 202(??):Article 103111, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301338.

Onoe:1998:TRT

[OYTY98] Yoshio Onoe, Kazumasa Yamazawa, Haruo Takemura, and Naokazu Yokoya. Telepresence by real-time viewdependent image generation from omnidirectional video streams. Computer Vision and Image Understanding: CVIU, 71(2):154– 165, August 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0705/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0705/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0705/production/ref.

Olmschenk:2019:GSS

[OZT19] Greg Olmschenk, Zhigang Zhu, and Hao Tang. Generalizing semi-supervised generative adversarial networks to regression using feature contrasting. Computer Vision and Image Understanding: CVIU, 186(??):1–12, September 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219300955.

Papamarkos:2000:GLR

[PA00] Nikos Papamarkos and Antonios Atsalakis. Graylevel reduction using local spatial features. Computer Vision and Image Understanding: CVIU, 78 (3):336–350, June 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2000.0838;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0838/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0838/ref.

Park:2006:STM

[PA06] Sangho Park and J. K. Aggarwal. Simultaneous tracking of multiple body parts of interacting persons. Computer Vision and Image Understanding: CVIU, 102(1): 1–21, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Papadourakis:2010:MOT

[PA10a] Vasilis Papadourakis and Antonis Argyros. Multiple objects tracking in the presence of long-term occlusions. Computer Vision and Image Understanding: CVIU, 114(7):835– 846, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Phan:2010:CBR

[PA10b] Raymond Phan and Dimitrios Androutsos. Contentbased retrieval of logo and trademarks in unconstrained color image databases using Color Edge Gradient Co-occurrence Histograms. Computer Vision and Image Understanding: CVIU, 114(1):66–84, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Perdigoto:2013:CMP

[PA13] Luis Perdigoto and Helder Araujo. Calibration of mirror position and extrinsic parameters in axial non-central catadioptric systems. Computer Vision and Image Understanding: CVIU, 117(8):909– 921, August 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213000635.

Papanelopoulos:2019:RMA

[PAK19] Nikos Papanelopoulos, Yannis Avrithis, and Stefanos Kollias. Revisiting the medial axis for planar shape decomposition. Computer Vision and Image Understanding: CVIU, 179(??): 66–78, February 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218304284.

Paragios:2016:SJ

[Par16] Nikos Paragios. State of the journal. Computer Vision and Image Understanding: CVIU, 147(??): 1–2, June 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421600045X.

Patane:2013:MRS

[Pat13] Giuseppe Patanè. Multiresolutive sparse approximations of d-dimensional data. Computer Vision and Image Understanding: CVIU, 117(4):418– 428, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314212001816.

Puzicha:1999:MAG

[PB99] Jan Puzicha and Joachim M.

Buhmann. Multiscale annealing for grouping and unsupervised texture segmentation. Computer Vision and Image Understanding: CVIU, 76(3):213– 230, December 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0805/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0805/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0805/production/ref.

Papazov:2011:SGO

[PB11] Chavdar Papazov and Darius Burschka. Stochastic global optimization for robust point set registration. Computer Vision and Image Understanding: CVIU, 115(12):1598–1609, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211001561.

Pernici:2020:SSL

[PBD20] Federico Pernici, Matteo Bruni, and Alberto Del Bimbo. Self-supervised on-line cumulative learning from video streams.

Computer Vision and Image Understanding: CVIU, 197–198(??):Article 102983, August 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314220300588.

Prasad:2004:RBI

[PBG04] B. G. Prasad, K. K. Biswas, and S. K. Gupta. Regionbased image retrieval using integrated color, shape, and location index. Computer Vision and Image Understanding: CVIU, 94(1–3): 193–233, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pennisi:2016:ORT

[PBI16] Andrea Pennisi, Domenico D. Bloisi, and Luca Iocchi. Online real-time crowd behavior detection in video sequences. Computer Vision and Image Understanding: CVIU, 144(??):166– 176, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002052.

Roman:2017:SDO

[PBPD^{+}17] Philippe Pérez de San Roman, Jenny Benois-Pineau, Jean-Philippe Domenger, Florent Paclet, Daniel Cataert, and Aymar de Rugy. Saliency driven object recognition in egocentric videos with deep CNN: toward application in assistance to neuroprostheses. Computer Vision and Image Understanding: CVIU, 164(??): 82–91, November 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300462.

Peng:1999:PFR

[PBQ99] Jing Peng, Bir Bhanu, and Shan Qing. Probabilistic feature relevance learning for content-based image retrieval. Computer Vision and Image Understanding: CVIU, 75(1–2):150– 164, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0770/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0770/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0770/production/ref.

Puig:2012:COC

[PBSG12] Luis Puig, J. Bermúdez, Peter Sturm, and J. J. Guerrero. Calibration of omnidirectional cameras in practice: a comparison of methods. Computer Vision and Image Understanding: CVIU, 116(1):120– 137, January 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001858.

Pateraki:2014:VEP

[PBT14] Maria Pateraki, Haris Baltzakis, and Panos Trahanias. Visual estimation of pointed targets for robot guidance via fusion of face pose and hand orientation. Computer Vision and Image Understanding: CVIU, 120(??): 1–13, March 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213002373.

Popham:2014:ESF

[PBW14] Thomas Popham, Abhir Bhalerao, and Roland Wilson. Estimating scene flow using an interconnected patch surface model with belief-propagation inference. Computer Vision and Image Understanding: CVIU, 121(??):74– 85, April 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214000083.

Pearce:1999:IMH

[PC99] Adrian R. Pearce and Terry Caelli. Interactively matching hand-drawings using induction. Computer Vision and Image Understanding: CVIU, 73(3):391– 403, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0742/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0742/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0742/production/ref.

Parameswaran:2005:HAR

[PC05] Vasu Parameswaran and Rama Chellappa. Human action-recognition using mutual invariants. Computer Vision and Image Understanding: CVIU, 98(2): 294–324, May 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Perret:2015:CIP

[PC15] B. Perret and Ch. Collet.

Connected image processing with multivariate attributes: an unsupervised Markovian classification approach. Computer Vision and Image Understanding: CVIU, 133 (??):1–14, April 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214001908.

Perri:2013:ACT

[PCC13] Stefania Perri, Pasquale Corsonello, and Giuseppe Cocorullo. Adaptive Census Transform: a novel hardware-oriented stereovision algorithm. Computer Vision and Image Understanding: CVIU, 117(1):29– 41, January 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001336.

Postolski:2014:SFE

[PCJ14] Michal Postolski, Michel Couprie, and Marcin Janaszewski. Scale filtered Euclidean medial axis and its hierarchy. Computer Vision and Image Understanding: CVIU, 129(??):89–102, December 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001556.

Paparoditis:1998:BDR

[PCJC98] N. Paparoditis, M. Cord, M. Jordan, and J.-P. Cocquerez. Building detection and reconstruction from mid- and high-resolution aerial imagery. Computer Vision and Image Understanding: CVIU, 72(2):122– 142, November 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0722/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0722/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0722/production/ref.

Plizzari:2021:SBA

[PCM21] Chiara Plizzari, Marco Cannici, and Matteo Matteucci. Skeleton-based action recognition via spatial and temporal transformer networks. Computer Vision and Image Understanding: CVIU, 208– 209(??):??, July 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000631.

Provost:2004:HMS

[PCR^{+}04] J.-N. Provost, C. Collet, P. Rostaing, P. Pérez, and P. Bouthemy. Hierarchical Markovian segmentation of multispectral images for the reconstruction of water depth maps. Computer Vision and Image Understanding: CVIU, 93(2):155–174, February 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Paragios:2005:GAR

[PD05] Nikos Paragios and Rachid Deriche. Geodesic active regions and level set methods for motion estimation and tracking. Computer Vision and Image Understanding: CVIU, 97(3):259– 282, March 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pehlivan:2011:NPB

[PD11] Selen Pehlivan and Pınar Duygulu. A new posebased representation for recognizing actions from multiple cameras. Computer Vision and Image Understanding: CVIU, 115(2):140–151, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Papachristou:2014:MEP

[PD14] Christos Papachristou and Anastasios N. Delopoulos. A method for the evaluation of projective geometric consistency in weakly calibrated stereo with application to point matching. Computer Vision and Image Understanding: CVIU, 119(??): 81–101, February 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421300235X.

Parisot:2017:SSC

[PD17] Pascaline Parisot and Christophe De Vleeschouwer. Scenespecific classifier for effective and efficient team sport players detection from a single calibrated camera. Computer Vision and Image Understanding: CVIU, 159 (??):74–88, June 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300036.

Piva:2023:EIT

[PD23] Fabrizio J. Piva and Gijs Dubbelman. Exploiting image translations via ensemble self-supervised learning for Unsupervised Domain Adaptation. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422300125X.

Palmer:1996:PMB

[PDK96] P. L. Palmer, H. Dabis, and J. Kittler. Performance measure for boundary detection algorithms. Computer Vision and Image Understanding: CVIU, 63(3): 476–494, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0036/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0036/production/pdf.

Pavlidis:2007:IHP

[PDS^{+}07] I. Pavlidis, J. Dowdall, N. Sun, C. Puri, J. Fei, and M. Garbey. Interacting with human physiology. Computer Vision and Image Understanding: CVIU, 108(1–2):150–170, October/ November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Prasad:2006:EAP

[PDTE06] V. S. N. Prasad, Larry S. Davis, Son Dinh Tran, and Ahmed Elgammal. Edge affinity for pose-contour matching. Computer Vision and Image Understanding: CVIU, 104(1):36–47, October 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Paglieroni:2009:RAG

[PE09] David W. Paglieroni and Walter G. Eppler. Resolution analysis for Gradient Direction Matching of object model edges to overhead images. Computer Vision and Image Understanding: CVIU, 113(2):235–248, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pece:2007:CRG

[Pec07] Arthur E. C. Pece. On the computational rationale for generative models. Computer Vision and Image Understanding: CVIU, 106(1): 130–143, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pizer:1998:ZIV

[PEFM98] Stephen M. Pizer, David Eberly, Daniel S. Fritsch, and Bryan S. Morse. Zoominvariant vision of figural shape: The mathematics of cores. Computer Vision and Image Understanding: CVIU, 69(1):055– 071, January 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0563/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0563/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0563/production/ref;
http://www.idealibrary.
com/links/artid/cviu.1997.
0564/production;   http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0564/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0564/production/ref.

Penna:1999:MAN

[Pen99] Michael A. Penna. The motion analysis of nonrigid membranes. Computer Vision and Image Understanding: CVIU, 75(3):281– 306, September 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0780/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0780/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0780/production/ref.

Peng:2003:MCR

[Pen03]  Jing Peng.  Multi-class

relevance feedback contentbased image retrieval. Computer Vision and Image Understanding: CVIU, 90(1): 42–67, April 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Peng:2015:CCS

[Pen15] Xiaoming Peng. Combine color and shape in realtime detection of textureless objects. Computer Vision and Image Understanding: CVIU, 135(??): 31–48, June 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000442.

Peterfreund:1999:VSD

[Pet99] Natan Peterfreund. The velocity snake: Deformable contour for tracking in spatio-velocity space. Computer Vision and Image Understanding: CVIU, 73(3): 346–356, March 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0732/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0732/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0732/production/ref.

Peyre:2009:MMS

[Pey09]  Gabriel Peyré.  Mani-

fold models for signals and images. Computer Vision and Image Understanding: CVIU, 113(2):249–260, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pauwels:1999:FSR

[PF99] E. J. Pauwels and G. Frederix. Finding salient regions in images: Nonparametric clustering for image segmentation and grouping. Computer Vision and Image Understanding: CVIU, 75(1– 2):73–85, July 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314299907634.

Plankers:2001:TMP

[PF01] Ralf Plänkers and Pascal Fua. Tracking and modeling people in video sequences. Computer Vision and Image Understanding: CVIU, 81(3):285–302, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0891;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0891/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0891/ref.

Philipp-Foliguet:2009:FIR

[PFGG09] Sylvie Philipp-Foliguet, Julien Gony, and Philippe-Henri Gosselin. FReBIR: an image retrieval system based on fuzzy region matching. Computer Vision and Image Understanding: CVIU, 113(6):693–707, June 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Picard:2013:EIS

[PG13] David Picard and Philippe-Henri Gosselin. Efficient image signatures and similarities using tensor products of local descriptors. Computer Vision and Image Understanding: CVIU, 117(6): 680–687, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000337.

Pua:2004:RTR

[PGGM04] Kok Meng Pua, John M.

Gauch, Susan E. Gauch, and Jedrzej Z. Miadowicz. Real time repeated video sequence identification. Computer Vision and Image Understanding: CVIU, 93(3): 310–327, March 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pertuz:2015:FAS

[PGP15] Said Pertuz, Miguel Angel Garcia, and Domenec Puig. Focus-aided scene segmentation. Computer Vision and Image Understanding: CVIU, 133(??): 66–75, April 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421400191X.

Pham:2001:SMF

[Pha01]  Dzung L. Pham.  Spa-

tial models for fuzzy clustering. Computer Vision and Image Understanding: CVIU, 84(2): 285–297, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0951;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0951/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0951/ref.

Pham:2017:PWO

[Pha17] The-Anh Pham. Pair-wisely optimized clustering tree for feature indexing. Computer Vision and Image Understanding: CVIU, 154(??): 35–47, January 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301084.

Peng:2015:CSH

[PHH^{+}15] Xi Peng, Junzhou Huang, Qiong Hu, Shaoting Zhang, Ahmed Elgammal, and Dimitris Metaxas. From circle to 3-sphere: Head pose estimation by instance parameterization. Computer Vision and Image Understanding: CVIU, 136(??): 92–102, July 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000545.

Pang:2023:FFT

[PHHL23] Shanmin Pang, Xin He, Wenyu Hao, and Yang Long. Feature fine-tuning and attribute representation transformation for zeroshot learning. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001911.

Pang:2011:STD

[PHY^{+}11] Yanwei Pang, Qiang Hao, Yuan Yuan, Tanji Hu, Rui Cai, and Lei Zhang. Summarizing tourist destinations by mining usergenerated travelogues and photos. Computer Vision and Image Understanding: CVIU, 115(3):352– 363, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pan:2011:MSP

[PJW11] Yongsheng Pan, Won-Ki Jeong, and Ross Whitaker. Markov surfaces: a probabilistic framework for userassisted three-dimensional image segmentation. Computer Vision and Image Understanding: CVIU, 115 (10):1375–1383, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314211001408.

Pont:2005:RLG

[PK05] Sylvia C. Pont and Jan J. Koenderink. Reflectance from locally glossy thoroughly pitted surfaces. Computer Vision and Image Understanding: CVIU, 98(2):211–222, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Prakash:2018:ALD

[PK18] Tanmay Prakash and Avinash C. Kak. Active learning for designing detectors for infrequently occurring objects in wide-area satellite imagery. Computer Vision and Image Understanding: CVIU, 170 (??):92–108, May 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300390.

Pham:2018:EDR

[PKC^{+}18] Huy-Hieu Pham, Louahdi Khoudour, Alain Crouzil, Pablo Zegers, and Sergio A. Velastin. Exploiting deep residual networks for human action recognition from skeletal data. Computer Vision and Image Understanding: CVIU, 170(??): 51–66, May 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300389.

Pitie:2007:ACG

[PKD07] François Pitié, Anil C.

Kokaram, and Rozenn Dahyot. Automated colour grading using colour distribution transfer. Computer Vision and Image Understanding: CVIU, 107(1–2): 123–137, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Park:2023:EEL

[PKH23] Jaeyoo Park, Junha Kim, and Bohyung Han. Endto-end learning for weakly supervised video anomaly detection using Absorbing Markov Chain. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001789.

Perse:2009:TBA

[PKK^{+}09] Matej Perše, Matej Kristan, Stanislav Kovačič, Goran Vučkovič, and Janez Perš. A trajectory-based analysis of coordinated team activity in a basketball game. Computer Vision and Image Understanding: CVIU, 113(5):612–621, May 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Palmer:1997:OLF

[PKP97] P. L. Palmer, J. Kittler, and M. Petrou. An optimizing line finder using a Hough transform algorithm. Computer Vision and Image Understanding: CVIU, 67(1):1–23, July 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0491/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0491/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0491/production/ref.

Pintea:2016:LSG

[PKvGS16] S. L. Pintea, S. Karaoglu, J. C. van Gemert, and A. W. M. Smeulders. Large scale Gaussian process for overlap-based object proposal scoring. Computer Vision and Image Understanding: CVIU, 150(??):95–108, September 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216300492.

Pizzi:2024:VSD

[PKZP^{+}24] Ed Pizzi, Giorgos Kordopatis-Zilos, Hiral Patel, Gheorghe Postelnicu, Sugosh Nagavara Ravindra, Akshay Gupta, Symeon Papadopoulos, Giorgos Tolias, and Matthijs Douze. The 2023 video similarity dataset and challenge. Computer Vision and Image Understanding: CVIU, 243(??): ??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731422400078X.

Pece:2007:GMB

[PL07] Arthur E. C. Pece and Rasmus Larsen. Generative model based vision. Computer Vision and Image Understanding: CVIU, 106(1): 3–4, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Potetz:2008:EBP

[PL08] Brian Potetz and Tai Sing Lee. Efficient belief propagation for higher-order cliques using linear constraint nodes. Computer Vision and Image Understanding: CVIU, 112(1):39– 54, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Perko:2010:FVC

[PL10] Roland Perko and Aleš Leonardis. A framework for visual-context-aware object detection in still images. Computer Vision and Image Understanding: CVIU, 114(6):700–711, June 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pla:1996:RPC

[Pla96] Filiberto Pla. Recognition of partial circular shapes from segmented contours. Computer Vision and Image Understanding: CVIU, 63(2): 334–343, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0023/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0023/production/pdf.

Phung:2016:PLD

[PLB16] Son Lam Phung, Manh Cuong Le, and Abdesselam Bouzerdoum. Pedestrian lane detection in unstructured scenes for assistive navigation. Computer Vision and Image Understanding: CVIU, 149(??):186– 196, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000369.

Pottmann:2004:RI

[PLH04] Helmut Pottmann, Stefan Leopoldseder, and Michael Hofer. Registration without ICP. Computer Vision and Image Understanding: CVIU, 95(1):54– 71, July 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pei:2014:ESI

[PLJS14] Deli Pei, Zhenguo Li, Rongrong Ji, and Fuchun Sun. Efficient semantic image segmentation with multi-class ranking prior. Computer Vision and Image Understanding: CVIU, 120(??):81– 90, March 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001926.

Pintelas:2023:MVC

[PLKP23] Emmanuel Pintelas, Ioannis E. Livieris, Sotiris Kotsiantis, and Panagiotis Pintelas. A multi-view-CNN framework for deep representation learning in image classification. Computer Vision and Image Understanding: CVIU, 232 (??):??, July 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422300067X.

Park:2000:LFM

[PLL00] Sang Ho Park, Kyoung Mu Lee, and Sang Uk Lee. A line feature matching technique based on an eigenvector approach. Computer Vision and Image Understanding: CVIU, 77 (3):263–283, March 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0808;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0808/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0808/ref.

Park:2003:RPO

[PLLL03] Bo Gun Park, Kyoung Mu Lee, Sang Uk Lee, and Jin Hak Lee. Recognition of partially occluded objects using probabilistic ARG (attributed relational graph)- based matching. Computer Vision and Image Understanding: CVIU, 90(3):217– 241, June 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Peng:2021:MFI

[PLYW21] Hong Peng, Bo Li, Qian Yang, and Jun Wang. Multifocus image fusion approach based on CNP systems in NSCT domain. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000722.

Pla:1997:MFP

[PM97] Filiberto Pla and John A.

Marchant. Matching feature points in image sequences through a regionbased method. Computer Vision and Image Understanding: CVIU, 66(3): 271–285, June 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0512/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0512/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0512/production/ref.

Poiesi:2013:MTT

[PMC13] Fabio Poiesi, Riccardo Mazzon, and Andrea Cavallaro. Multi-target tracking on confidence maps: an application to people tracking. Computer Vision and Image Understanding: CVIU, 117(10):1257– 1272, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001634.

Potje:2022:LGA

[PMCN22] Guilherme Potje, Renato Martins, Felipe Cadar, and Erickson R. Nascimento. Learning geodesic-aware local features from RGB-D images. Computer Vision and Image Understanding: CVIU, 219(??): ??, June 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731422200039X.

Peruzzo:2023:INP

[PMG^{+}23] Elia Peruzzo, Willi Menapace, Vidit Goel, Federica Arrigoni, Hao Tang, Xingqian Xu, Arman Chopikyan, Nikita Orlov, Yuxiao Hu, Humphrey Shi, Nicu Sebe, and Elisa Ricci. Interactive neural painting. Computer Vision and Image Understanding: CVIU, 235(??):

??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001583.

Patel:2017:EPF

[PMR17] Bhavik Patel, R. P. Maheshwari, and Balasubramanian Raman. Evaluation of periocular features for kinship verification in the wild. Computer Vision and Image Understanding: CVIU, 160 (??):24–35, July 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421730070X.

Pluim:2000:IAM

[PMV00] Josien P. W. Pluim, J. B. Antoine Maintz, and Max A. Viergever. Interpolation artefacts in mutual information-based image registration. Computer Vision and Image Understanding: CVIU, 77(2): 211–232, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0816;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0816/pdf;   http:
//www.idealibrary.com/

                                        links/doi/10.1006/cviu.
                                        1999.0816/ref.

Planitz:2005:CFS

[PMW05] B. M. Planitz, A. J. Maeder, and J. A. Williams. The correspondence framework for 3D surface matching algorithms. Computer Vision and Image Understanding: CVIU, 97(3):347– 383, March 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pemasiri:2021:MMS

[PNSF21] Akila Pemasiri, Kien Nguyen, Sridha Sridharan, and Clinton Fookes. Multi-modal semantic image segmentation. Computer Vision and Image Understanding: CVIU, 202(??):Article 103085, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422030117X.

Poppe:2007:VBH

[Pop07] Ronald Poppe. Vision-based human motion analysis: an overview. Computer Vision and Image Understanding: CVIU, 108(1–2):4–18, October/November 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Portegys:2000:RHP

[Por00] Thomas E. Portegys. Recognizing hand-printed digits with a distance quasimetric. Computer Vision and Image Understanding: CVIU, 80(3): 289–294, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0876;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0876/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0876/ref.

Petrou:2006:TRS

[PPT06] M. Petrou, R. Piroddi, and A. Talebpour. Texture recognition from sparsely and irregularly sampled data. Computer Vision and Image Understanding: CVIU, 102(1):95–104, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Peng:2011:MFE

[PQML11] Bo Peng, Gang Qian, Yunqian Ma, and Baoxin Li. Multifactor feature extraction for human movement recognition. Computer Vision and Image Understanding: CVIU, 115(3):375– 389, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pickering:2003:EKF

[PR03] Marcus J. Pickering and Stefan Rüger. Evaluation of key frame-based retrieval techniques for video. Computer Vision and Image Understanding: CVIU, 92 (2–3):217–235, November/ December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Perez-Rua:2016:OGM

[PRCP16] Juan-Manuel Pérez-Rúa, Tomas Crivelli, and Patrick Pérez. Object-guided motion estimation. Computer Vision and Image Understanding: CVIU, 153(??): 88–99, December 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300546.

Papaodysseus:2014:IWA

[PRG^{+}14] C. Papaodysseus, P. Rousopoulos, F. Giannopoulos, S. Zannos, D. Arabadjis, M. Panagopoulos, E. Kalfa, C. Blackwell, and S. Tracy. Identifying the writer of ancient inscriptions and Byzantine codices. A novel approach. Computer Vision and Image Understanding: CVIU, 121 (??):57–73, April 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000101.

Pansari:2019:LPB

[PRK19] Pankaj Pansari, Chris Russell, and M. Pawan Kumar. Linear programming-based submodular extensions for marginal estimation. Computer Vision and Image Understanding: CVIU, 189 (??):Article 102824, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219301316.

Paragios:2003:NRR

[PRR03] Nikos Paragios, Mikael Rousson, and Visvanathan Ramesh. Non-rigid registration using distance functions. Computer Vision and Image Understanding: CVIU, 89(2–3):142– 165, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pizlo:1997:GVS

[PRW97a] Zygmunt Pizlo, Azriel Rosenfeld, and Isaac Weiss. The geometry of visual space: About the incompatibility between science and mathematics. Computer Vision and Image Understanding: CVIU, 65(3):

425–433, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0492/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0492/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0492/production/ref. See

comments [Col97, Åst97, Ver97, HM97].

Pizlo:1997:VSM

[PRW97b] Zygmunt Pizlo, Azriel Rosenfeld, and Isaac Weiss. Visual space: Mathematics, engineering, and science. Computer Vision and Image Understanding: CVIU, 65(3):450– 454, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0498/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0498/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0498/production/ref. See

replies [May97, HM97].

Pham:2005:ORU

[PS05] Thang V. Pham and Arnold W. M. Smeulders. Object recognition with uncertain geometry and uncertain part detection. Computer Vision and Image Understanding: CVIU, 99(2):241– 258, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Proenca:2012:FCS

[PS12] Hugo Proença and Gil Santos. Fusing color and shape descriptors in the recognition of degraded iris images acquired at visible wavelengths. Computer Vision and Image Understanding: CVIU, 116(2):167–178, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002359.

Pisharady:2015:RMD

[PS15] Pramod Kumar Pisharady and Martin Saerbeck. Recent methods and databases in vision-based hand gesture recognition: a review. Computer Vision and Image Understanding: CVIU, 141(??):152–165, December 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215001794.

Parida:2022:DST

[PS22] Kranti Kumar Parida and Gaurav Sharma. Discriminative semantic transitive consistency for cross-modal learning. Computer Vision and Image Understanding: CVIU, 219(??): ??, June 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000339.

Pennisi:2024:FFL

[PSB^{+}24] Matteo Pennisi, Federica Proietto Salanitri, Giovanni Bellitto, Bruno Casella, Marco Aldinucci, Simone Palazzo, and Concetto Spampinato. FedER: Federated learning through Experience Replay and privacypreserving data synthesis. Computer Vision and Image Understanding: CVIU, 238 (??):??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422300262X.

Papadopoulos:2011:CSO

[PSE^{+}11] G. Th. Papadopoulos, C. Saathoff, H. J. Escalante, V. Mezaris, I. Kompatsiaris, and M. G. Strintzis. A comparative study of object-level spatial context techniques for semantic image analysis. Computer Vision and Image Understanding: CVIU, 115(9):1288–1307, September 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211001275.

Pechuk:2008:LFB

[PSR08] Michael Pechuk, Octavian Soldea, and Ehud Rivlin. Learning function-based object classification from 3D imagery. Computer Vision and Image Understanding: CVIU, 110(2):173– 191, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pan:2021:ECI

[PSY^{+}21] Jinshan Pan, Deqing Sun, Jian Yang, Wangmeng Zuo, Paolo Favaro, Yasuyuki Matsushita, and Ming-Hsuan Yang. Editorial for CVIU DL for image restoration. Computer Vision and Image Understanding: CVIU, 208–209(??): ??, July 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221000667.

Pei:2013:LPV

[PSYZ13] Mingtao Pei, Zhangzhang Si, Benjamin Z. Yao, and Song-Chun Zhu. Learning and parsing video events with goal and intent prediction. Computer Vision and Image Understanding: CVIU, 117(10):1369–1383, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421200197X.

Park:2008:UHI

[PT08] Sangho Park and Mohan M. Trivedi. Understanding human interactions with track and body synergies (TBS) captured from multiple views. Computer Vision and Image Understanding: CVIU, 111(1):2– 20, July 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Park:2015:SBM

[PT15] Jae-Hyuck Park and Yu-Wing Tai. A simulation based method for vehicle motion prediction. Computer Vision and Image Understanding: CVIU, 136 (??):79–91, July 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215000508.

Pothos:2012:LSD

[PTE12] Vasileios K. Pothos, Christos Theoharatos, and George Economou. A local spectral distribution approach to face recognition. Computer Vision and Image Understanding: CVIU, 116(6): 663–675, June 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000318.

Psalta:2024:TBA

[PTK24] Athena Psalta, Vasileios Tsironis, and Konstantinos Karantzalos. Transformerbased assignment decision network for multiple object tracking. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000389.

Poggi:2020:LCM

[PTM20] Matteo Poggi, Fabio Tosi, and Stefano Mattoccia. Learning a confidence measure in the disparity domain from O(1) features. Computer Vision and Image Understanding: CVIU, 193 (??):Article 102905, April 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220300023.

Pudney:1998:DOH

[Pud98] Chris Pudney. Distanceordered homotopic thinning: a skeletonization algorithm for 3D digital images. Computer Vision and Image Understanding: CVIU, 72(3):404–413, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0680/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0680/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0680/production/ref.

Pun:2003:RIT

[Pun03] Chi-Man Pun. Rotationinvariant texture feature for image retrieval. Computer Vision and Image Understanding: CVIU, 89(1):24– 43, January 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Poli:1997:SRD

[PV97] Riccardo Poli and Guido Valli. Shape from radiological density. Computer Vision and Image Understanding: CVIU, 65(3):361– 381, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.

0489/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0489/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0489/production/ref.

Pauwels:2006:OIR

[PV06] Karl Pauwels and Marc M.

Van Hulle. Optimal instantaneous rigid motion estimation insensitive to local minima. Computer Vision and Image Understanding: CVIU, 104(1):77–86, October 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pellegrini:2013:TMC

[PV13] Stefano Pellegrini and Luc Van Gool. Tracking with a mixed continuousdiscrete Conditional Random Field. Computer Vision and Image Understanding: CVIU, 117(10):1215– 1228, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001683.

Pereira:2014:CMD

[PV14] Jose Costa Pereira and Nuno Vasconcelos. Crossmodal domain adaptation for text-based regularization of image semantics in image retrieval systems. Computer Vision and Image Understanding: CVIU, 124(??): 123–135, July 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421400054X.

Peng:2015:PLI

[PV15] Renbin Peng and Pramod K. Varshney. On performance limits of image segmentation algorithms. Computer Vision and Image Understanding: CVIU, 132(??): 24–38, March 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214002240.

Park:2006:SMC

[PW06] Eunkwang Park and Kwangyun Wohn. Stereo and motion correspondences using nonlinear optimization method. Computer Vision and Image Understanding: CVIU, 101 (3):194–203, March 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pan:2023:GGM

[PW23] Sicong Pan and Hui Wei. A global generalized maximum coverage-based solution to the non-model-based view planning problem for object reconstruction. Computer Vision and Image Understanding: CVIU, 226(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001631.

Park:2023:CMA

[PWL^{+}23] Yeonju Park, Sangmin Woo, Sumin Lee, Muhammad Adi Nugroho, and Changick Kim. Cross-modal alignment and translation for missing modality action recognition. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001856.

Paisitkriangkrai:2017:SLM

[PWSvdH17] Sakrapee Paisitkriangkrai, Lin Wu, Chunhua Shen, and Anton van den Hengel. Structured learning of metric ensembles with application to person reidentification. Computer Vision and Image Understanding: CVIU, 156(??): 51–65, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301667.

Peng:2016:BVW

[PWWQ16] Xiaojiang Peng, Limin Wang, Xingxing Wang, and Yu Qiao. Bag of visual words and fusion methods for action recognition: Comprehensive study and good practice. Computer Vision and Image Understanding: CVIU, 150(??):109– 125, September 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300091.

Pang:2014:ELL

[PXTZ14] Shanmin Pang, Jianru Xue, Qi Tian, and Nanning Zheng. Exploiting local linear geometric structure for identifying correct matches. Computer Vision and Image Understanding: CVIU, 128(??):51–64, November 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001337.

Palanivel:2008:MPA

[PY08] S. Palanivel and B. Yegnanarayana. Multimodal person authentication using speech, face and visual speech. Computer Vision and Image Understanding: CVIU, 109(1):44–55, January 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Park:2019:PPD

[PY19] Min-Gyu Park and Kuk-Jin Yoon. As-planar-as-possible depth map estimation. Computer Vision and Image Understanding: CVIU, 181(??):50– 59, April 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219300165.

Perez-Yus:2017:SDO

[PYGGLNG17] A. Perez-Yus, D. Gutierrez-Gomez, G. Lopez-Nicolas, and J. J. Guerrero. Stairs detection with odometryaided traversal from a wearable RGB-D camera. Computer Vision and Image Understanding: CVIU, 154(??):192– 205, January 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300315.

Polat:2003:RTH

[PYS03] Ediz Polat, Mohammed Yeasin, and Rajeev Sharma. Robust tracking of human body parts for collaborative human computer interaction. Computer Vision and Image Understanding: CVIU, 89(1):44–69, January 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pandey:2017:AVS

[PYWZ17] Dinesh Pandey, Xiaoxia Yin, Hua Wang, and Yanchun Zhang. Accurate vessel segmentation using maximum entropy incorporating line detection and phase-preserving denoising. Computer Vision and Image Understanding: CVIU, 155(??):162– 172, February 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216302028.

Paragios:2008:DOC

[PZ08] Nikos Paragios and Ramin Zabih. Discrete optimization in computer vision. Computer Vision and Image Understanding: CVIU, 112(1):1–2, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Paragios:2009:CDO

[PZ09] Nikos Paragios and Ramin Zabih. Corrigendum to “Discrete optimization in computer vision” [Comput. Vis. Image Understanding 112 (2008) 1–2]. Computer Vision and Image Understanding: CVIU, 113(4):

588, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See [PZ08].

Pribyl:2017:APE

[PZC17] Bronislav Pribyl, Pavel Zemcík, and Martin Cadík. Absolute pose estimation from line correspondences using direct linear transformation. Computer Vision and Image Understanding: CVIU, 161(??):130– 144, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300784.

Peng:2021:SIL

[PZM^{+}21] Can Peng, Kun Zhao, Sam Maksoud, Meng Li, and Brian C. Lovell. SID: Incremental learning for anchorfree object detection via selective and inter-related distillation. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000734.

Prankl:2013:IOM

[PZV13] Johann Prankl, Michael Zillich, and Markus Vincze. Interactive object modelling based on piecewise planar surface patches. Computer Vision and Image Understanding: CVIU, 117(6): 718–731, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421300026X.

Pan:2013:EAF

[PZX13] Hong Pan, Yaping Zhu, and Liangzheng Xia. Efficient and accurate face detection using heterogeneous feature descriptors and feature selection. Computer Vision and Image Understanding: CVIU, 117(1):12– 28, January 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001294.

Qazi:2011:PML

[QAB^{+}11] Imtnan-Ul-Haque Qazi, Olivier Alata, Jean-Christophe Burie, Mohamed Abadi, Ahmed Moussa, and Christine Fernandez-Maloigne. Parametric models of linear prediction error distribution for color texture and satellite image segmentation. Computer Vision and Image Understanding: CVIU, 115(8):1245–1262, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211000944.

Qin:2021:MSA

[QBZ21] Jia Qin, Huihui Bai, and Yao Zhao. Multi-scale attention network for image inpainting. Computer Vision and Image Understanding: CVIU, 204(??):Article 103155, March 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301740.

Qian:2004:BSC

[QC04] Gang Qian and Rama Chellappa. Bayesian selfcalibration of a moving camera. Computer Vision and Image Understanding: CVIU, 95(3):287–316, September 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Qiu:2020:PDA

[QCH20] Wenjie Qiu, Wendong Chen, and Haifeng Hu. Partial domain adaptation based on shared class oriented adversarial network. Computer Vision and Image Understanding: CVIU, 199 (??):Article 103018, October 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220300795.

Quan:2023:AKD

[QCL^{+}23] Zhenzhen Quan, Qingshan Chen, Yujun Li, Zhi Liu, and Yan Cui. ARCTIC: a knowledge distillation approach via attention-based relation matching and activation region constraint for RGB-to-infrared videos action recognition. Computer Vision and Image Understanding: CVIU, 237(??): ??, December 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002333.

Quan:2019:ASR

[QCXJ19] Yuhui Quan, Yixin Chen, Ruotao Xu, and Hui Ji. Attention with structure regularization for action recognition. Computer Vision and Image Understanding: CVIU, 187(??):Article 102794, October 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1077314218301954.

Quach:2017:NCO

[QDLB17] Kha Gia Quach, Chi Nhan Duong, Khoa Luu, and Tien D. Bui. Non-convex online robust PCA: Enhance sparsity via l_{p}-norm minimization. Computer Vision and Image Understanding: CVIU, 158(??): 126–140, May 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300450.

Qiao:2023:DSR

[QGZ^{+}23] Xin Qiao, Chenyang Ge, Youmin Zhang, Yanhui Zhou, Fabio Tosi, Matteo Poggi, and Stefano Mattoccia. Depth superresolution from explicit and implicit high-frequency features. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002217.

Qureshi:2012:CAA

[QKH^{+}12] Rashid Jalal Qureshi, Laszlo Kovacs, Balazs Harangi, Brigitta Nagy, Tunde Peto, and Andras Hajdu. Combining algorithms for automatic detection of optic disc and macula in fundus images. Computer Vision and Image Understanding: CVIU, 116(1):138– 145, January 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001883.

Qu:1996:STA

[QL96] Xiaoqing Qu and Xiaobo Li. A 3D surface tracking algorithm. Computer Vision and Image Understanding: CVIU, 64(1): 147–156, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0050/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0050/production/pdf.

Qin:2017:FAR

[QLY^{+}17] Jie Qin, Li Liu, Mengyang Yu, Yunhong Wang, and Ling Shao. Fast action retrieval from videos via feature disaggregation. Computer Vision and Image Understanding: CVIU, 156(??):104– 116, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301436.

Quan:2017:SLS

[QSX17] Yuhui Quan, Yuping Sun, and Yong Xu. Spatiotemporal lacunarity spectrum for dynamic texture classification. Computer Vision and Image Understanding: CVIU, 165(??):85–96, December 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301704.

Qiao:2023:MPM

[QSXS23] Nianzu Qiao, Jia Sun, Duo Xu, and Changyin Sun. Multi-patch multiscale model for motion deblurring with high-frequency information. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001509.

Qian:2010:AIT

[QT10] Xiaoning Qian and Hemant D. Tagare. Adapting indexing trees to data distribution in feature spaces. Computer Vision and Image Understanding: CVIU, 114 (1):111–124, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Qu:2022:IIS

[QTLP22] Shaojun Qu, Huang Tan, Qiaoliang Li, and Zili Peng. Interactive image segmentation based on the appearance model and orientation energy. Computer Vision and Image Understanding: CVIU, 217(??): ??, March 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000121.

Quan:1998:JIT

[QV98] Long Quan and Françoise Veillon. Joint invariants of a triplet of coplanar conics: Stability and discriminating power for object recognition. Computer Vision and Image Understanding: CVIU, 70(1): 111–119, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0617/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0617/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0617/production/ref.

Qiu:2020:TMM

[QWHW20] Ji Qiu, Lide Wang, Yu Hen Hu, and Yin Wang. Two motion models for improving video object tracking performance. Computer Vision and Image Understanding: CVIU, 195(??):Article 102951, June 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300345.

Qian:2024:CCT

[QZY^{+}24] Huifang Qian, Jialun Zhang, Jianping Yi, Zhenyu Shi, and Yimin Zhang. CTM: Cross-time temporal module for fine-grained action recognition. Computer Vision and Image Understanding: CVIU, 244(??): ??, July 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000948.

Rzeszutek:2015:FER

[RA15] Richard Rzeszutek and Dimitrios Androutsos. A framework for estimating relative depth in video. Computer Vision and Image Understanding: CVIU, 133 (??):15–29, April 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421500003X.

Rekik:2013:TGP

[RAC^{+}13] Islem Rekik, Stéphanie Allassonnière, Olivier Clatz, Ezequiel Geremia, Erin Stretton, Hervé Delingette, and Nicholas Ayache. Tumor growth parameters estimation and source localization from a unique time point: Application to lowgrade gliomas. Computer Vision and Image Understanding: CVIU, 117(3): 238–249, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001476.

Radeva:1997:DBS

[RAH97] Petia Radeva, Amir A. Amini, and Jiantao Huang. Deformable B-solids and implicit snakes for 3D localization and tracking of SPAMM MRI data. Computer Vision and Image Understanding: CVIU, 66(2): 163–178, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0611/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0611/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0611/production/ref.

Raveaux:2011:LGP

[RAHT11] Romain Raveaux, Sébastien Adam, Pierre Héroux, and Éric Trupin. Learning graph prototypes for shape recognition. Computer Vision and Image Understanding: CVIU, 115(7):905– 918, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421100083X.

Ribeiro:2016:RFD

[RAP16] Pedro Canotilho Ribeiro, Romaric Audigier, and Quoc Cuong Pham. RI-MOC, a feature to discriminate unstructured motions: Application to violence detection for videosurveillance. Computer Vision and Image Understanding: CVIU, 144(??):121– 143, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002374.

Riahi:2016:OMO

[RB16] Dorra Riahi and Guillaume-Alexandre Bilodeau. Online multi-object tracking by detection based on generative appearance models. Computer Vision and Image Understanding: CVIU, 152(??):88–102, November 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301023.

Rantoson:2018:DMB

[RB18] Rindra Rantoson and Adrien Bartoli. A 3D deformable model-based framework for the retrieval of near-isometric flattenable objects using Bag-of-Visual-Words. Computer Vision and Image Understanding: CVIU, 167 (??):89–108, February 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S107731421730139X.

Regmi:2019:CVI

[RB19] Krishna Regmi and Ali Borji. Cross-view image synthesis using geometryguided conditional GANs. Computer Vision and Image Understanding: CVIU, 187(??):Article 102788, October 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S1077314219301043.

Rodrigues:2020:PCC

[RBA20] Pedro M. C. Rodrigues, João P. Barreto, and Michel Antunes. Photometric camera characterization from a single image with invariance to light intensity and vignetting. Computer Vision and Image Understanding: CVIU, 192(??):Article 102887, March 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219301699.

Ross:2022:DHC

[RBC22] Arun Ross, Sudipta Banerjee, and Anurag Chowdhury. Deducing health cues from biometric data. Computer Vision and Image Understanding: CVIU, 221(??): ??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000522.

Roelandts:2014:RRE

[RBdDS14] Tom Roelandts, K. Joost Batenburg, Arnold J. den Dekker, and Jan Sijbers. The reconstructed residual error: a novel segmentation evaluation measure for reconstructed images in tomography. Computer Vision and Image Understanding: CVIU, 126(??):28–37, September 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421400126X.

Rajagopalan:1997:SVA

[RC97] A. N. Rajagopalan and S. Chaudhuri. Spacevariant approaches to recovery of depth from defocused images. Computer Vision and Image Understanding: CVIU, 68(3):309– 329, December 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0534/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0534/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0534/production/ref.

Richard:2003:NIR

[RC03] Frédéric J. P. Richard and Laurent D. Cohen. A new image registration technique with free boundary constraints: application to mammography. Computer Vision and Image Understanding: CVIU, 89(2–3): 166–196, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rouchdy:2013:GVA

[RC13] Youssef Rouchdy and Laurent D. Cohen. Geodesic voting for the automatic extraction of tree structures. Methods and applications. Computer Vision and Image Understanding: CVIU, 117(10):1453–1467, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001057.

Ryoo:2013:PDD

[RCJ^{+}13] M. S. Ryoo, Sunglok Choi, Ji Hoon Joung, Jae-Yeong Lee, and Wonpil Yu. Personal driving diary: Automated recognition of driving events from first-person videos. Computer Vision and Image Understanding: CVIU, 117(10):1299–1312, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000118.

Redondo-Cabrera:2019:ULV

[RCLS19] Carolina Redondo-Cabrera and Roberto Lopez-Sastre. Unsupervised learning from videos using temporal coherency deep networks. Computer Vision and Image Understanding: CVIU, 179 (??):79–89, February 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314218301772.

Rios-Cabrera:2014:BMD

[RCT14] Reyes Rios-Cabrera and Tinne Tuytelaars. Boosting masked dominant orientation templates for efficient object detection. Computer Vision and Image Understanding: CVIU, 120(??): 103–116, March 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213002397.

Rios-Cabrera:2012:EMC

[RCTV12] Reyes Rios-Cabrera, Tinne Tuytelaars, and Luc Van Gool. Efficient multi-camera vehicle detection, tracking, and identification in a tunnel surveillance application. Computer Vision and Image Understanding: CVIU, 116(6):742– 753, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314212000380.

Rahmati:2015:WSM

[RDA^{+}15] Hodjat Rahmati, Ralf Dragon, Ole Morten Aamo, Lars Adde, Øyvind Stavdahl, and Luc Van Gool. Weakly supervised motion segmentation with particle matching. Computer Vision and Image Understanding: CVIU, 140(??):30–42, November 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215001575.

Radgui:2011:OFE

[RDM^{+}11] A. Radgui, C. Demonceaux, E. Mouaddib, M. Rziza, and D. Aboutajdine. Optical flow estimation from multichannel spherical image decomposition. Computer Vision and Image Understanding: CVIU, 115(9):1263– 1272, September 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421100124X.

Rivlin:1995:RFP

[RDR95] Ehud Rivlin, Sven J. Dickinson, and Azriel Rosenfeld. Recognition by functional parts. Computer Vision and Image Understanding: CVIU, 62(2):164– 176, September 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1048/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1048/production/pdf.

Rastegar:2024:BNM

[RDS24] Sarah Rastegar, Hazel Doughty, and Cees G. M. Snoek. Background no more: Action recognition across domains by causal interventions. Computer Vision and Image Understanding: CVIU, 242(??): ??, May 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314224000560.

Ryan:2015:ECC

[RDSF15] David Ryan, Simon Denman, Sridha Sridharan, and Clinton Fookes. An evaluation of crowd counting methods, features and regression models. Computer Vision and Image Understanding: CVIU, 130(??): 1–17, January 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214001611.

Rambour:2019:USR

[RDT^{+}19] Clément Rambour, Loïc Denis, Florence Tupin, Hélène Oriot, Yue Huang, and Laurent Ferro-Famil. Urban surface reconstruction in SAR tomography by graph-cuts. Computer Vision and Image Understanding: CVIU, 188(??):Article 102791, November 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314219301079.

Rigas:2015:EMV

[REF15] Ioannis Rigas, George Economou, and Spiros Fotopoulos. Efficient modeling of visual saliency based on local sparse representation and the use of Hamming distance. Computer Vision and Image Understanding: CVIU, 134(??):33– 45, May 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000223.

Reid:2016:ACC

[Rei16] Ian Reid. 12th Asian conference on computer vision. Computer Vision and Image Understanding: CVIU, 146(??):51, May 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300108.

Remondino:2004:DRS

[Rem04] Fabio Remondino. 3-D reconstruction of static human body shape from image sequence. Computer Vision and Image Understanding: CVIU, 93(1):65– 85, January 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Robertson:2002:PER

[RF02] Craig Robertson and Robert B.

Fisher. Parallel evolutionary registration of range data. Computer Vision and Image Understanding: CVIU, 87(1–3):39–50, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ripas:2023:IPS

[RF23] Roger Ripas and Leandro A. F. Fernandes. Improving the planarity and sharpness of monocularly estimated depth images using the Phong reflection model. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001066.

Rothwell:1997:CPR

[RFC97] Charlie Rothwell, Olivier Faugeras, and Gabriella Csurka. A comparison of projective reconstruction methods for pairs of views. Computer Vision and Image Understanding: CVIU, 68(1):37– 58, October 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0525/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0525/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0525/production/ref.

Ragusa:2023:MME

[RFF23] Francesco Ragusa, Antonino Furnari, and Giovanni Maria Farinella. MEC-CANO: a multimodal egocentric dataset for humans behavior understanding in the industrial-like domain. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001443.

Rodrigues:2002:SIR

[RFL02] Marcos Rodrigues, Robert Fisher, and Yonghuai Liu. Special issue on registration and fusion of range images. Computer Vision and Image Understanding: CVIU, 87 (1–3):1–7, July 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rodin:2021:PFF

[RFMF21] Ivan Rodin, Antonino Furnari, Dimitrios Mavroeidis, and Giovanni Maria Farinella. Predicting the future from first person (egocentric) vision: a survey. Computer Vision and Image Understanding: CVIU, 211(??): ??, October 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000965.

Rohr:2003:SBE

[RFS03] K. Rohr, M. Fornefett, and H. S. Stiehl. Splinebased elastic image registration: integration of landmark errors and orientation attributes. Computer Vision and Image Understanding: CVIU, 90(2):153– 168, May 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rocha:2010:PRS

[RG10] Anderson Rocha and Siome Goldenstein. Progressive randomization: Seeing the unseen. Computer Vision and Image Understanding: CVIU, 114(3):349– 362, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rathee:2016:MDM

[RG16] Neeru Rathee and Dinesh Ganotra. Multiview Distance Metric Learning on facial feature descriptors for automatic pain intensity detection. Computer Vision and Image Understanding: CVIU, 147(??): 77–86, June 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002684.

Richard:2017:BWE

[RG17] Alexander Richard and Juergen Gall. A bagof-words equivalent recurrent neural network for action recognition. Computer Vision and Image Understanding: CVIU, 156(??): 79–91, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301680.

Ryoo:2010:TDI

[RGA10] M. S. Ryoo, Kristen Grauman, and J. K. Aggarwal. A task-driven intelligent workspace system to provide guidance feedback. Computer Vision and Image Understanding: CVIU, 114(5):520–534, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rigoutsos:1995:BAM

[RH95] Isidore Rigoutsos and Robert Hummel. A Bayesian approach to model matching with geometric hashing. Computer Vision and Image Understanding: CVIU, 62 (1):11–26, July 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/

                                        links/artid/cviu.1995.
                                        1038/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1038/production/pdf.

Ragheb:2006:TNV

[RH06] Hossein Ragheb and Edwin R. Hancock. Testing new variants of the Beckmann–Kirchhoff model against radiance data. Computer Vision and Image Understanding: CVIU, 102(2): 145–168, May 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Raviv:2000:VLN

[RJ00] Daniel Raviv and Kunal Joarder. The visual looming navigation cue: a unified approach. Computer Vision and Image Understanding: CVIU, 79(3): 331–363, September 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2000.0862;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0862/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0862/ref.

Raftopoulos:2011:GLT

[RK11] Konstantinos A. Raftopoulos and Stefanos D. Kollias. The Global–Local transformation for noise resistant shape representation. Computer Vision and Image Understanding: CVIU, 115(8):1170–1186, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421100097X.

Rogelj:2003:PSM

[RKG03] Peter Rogelj, Stanislav Kovaçiç, and James C. Gee. Point similarity measures for non-rigid registration of multi-modal data. Computer Vision and Image Understanding: CVIU, 92(1): 112–140, October 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ryu:2022:MIC

[RKKK22] Junhwan Ryu, Jongchan Kim, Heegon Kim, and Sungho Kim. Multispectral interaction convolutional neural network for pedestrian detection. Computer Vision and Image Understanding: CVIU, 223 (??):??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001321.

Rupprecht:2018:LPA

[RKL^{+}18] Christian Rupprecht, Ansh Kapil, Nan Liu, Lamberto Ballan, and Federico Tombari. Learning without prejudice: Avoiding bias in webly-supervised action recognition. Computer Vision and Image Understanding: CVIU, 173(??): 24–32, August 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301406.

Roshtkhari:2013:LRT

[RL13] Mehrsan Javan Roshtkhari and Martin D. Levine. An on-line, real-time learning method for detecting anomalies in videos using spatio-temporal compositions. Computer Vision and Image Understanding: CVIU, 117(10):1436– 1452, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213001239.

Royer:2017:CRC

[RLB17] Emilien Royer, Thibault Lelore, and Frédéric Bouchara. COnfusion REduction (CORE) algorithm for local descriptors, floating-point and binary cases. Computer Vision and Image Understanding: CVIU, 158(??):115– 125, May 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421730005X.

Ryberg:2011:AEG

[RLC^{+}11] A. Ryberg, B. Lennartson, A.-K. Christiansson, M. Ericsson, and L. Asplund. Analysis and evaluation of a general camera model. Computer Vision and Image Understanding: CVIU, 115(11):1503– 1515, November 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001548.

Rozantsev:2015:RSI

[RLF15] Artem Rozantsev, Vincent Lepetit, and Pascal Fua. On rendering synthetic images for training an object detector. Computer Vision and Image Understanding: CVIU, 137(??): 24–37, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002446.

Rigamonti:2014:RSI

[RLG^{+}14] Roberto Rigamonti, Vincent Lepetit, Germán González, Engin Türetken, Fethallah Benmansour, Matthew Brown, and Pascal Fua. On the relevance of sparsity for image classification. Computer Vision and Image Understanding: CVIU, 125(??):115– 127, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000757.

Ruffieux:2015:GRC

[RLMK15] Simon Ruffieux, Denis Lalanne, Elena Mugellini, and Omar Abou Khaled. Gesture recognition corpora and tools: a scripted ground truthing method. Computer Vision and Image Understanding: CVIU, 131(??): 72–87, February 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001568.

Ramalingam:2006:GSM

[RLS06] Srikumar Ramalingam, Suresh K. Lodha, and Peter Sturm. A generic structure-from-motion framework. Computer Vision and Image Understanding: CVIU, 103(3): 218–228, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Robert:1998:FBI

[RM98] L. Robert and G. Malandain. Fast binary image processing using binary decision diagrams. Computer Vision and Image Understanding: CVIU, 72(1): 1–9, October 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0655/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0655/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0655/production/ref.

Rivera:2002:ARC

[RM02] Mariano Rivera and Jose L.

Marroquin. Adaptive rest condition potentials: First and second order edgepreserving regularization. Computer Vision and Image Understanding: CVIU, 88 (2):76–93, November 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Raytchev:2003:URM

[RM03] Bisser Raytchev and Hiroshi Murase. Unsupervised recognition of multiview face sequences based on pairwise clustering with attraction and repulsion. Computer Vision and Image Understanding: CVIU, 91(1–2):22–52, July/August 2003. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Rosin:2006:SCM

[RM06] Paul L. Rosin and Christine L. Mumford. A symmetric convexity measure. Computer Vision and Image Understanding: CVIU, 103 (2):101–111, August 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rakshit:2022:FGF

[RMC^{+}22] Sayan Rakshit, Anwesh Mohanty, Ruchika Chavhan, Biplab Banerjee, Gemma Roig, and Subhasis Chaudhuri. FRIDA — generative feature replay for incremental domain adaptation. Computer Vision and Image Understanding: CVIU, 217 (??):??, March 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222000091.

Rahimi:2008:RDD

[RMD08] Ali Rahimi, Louis-Philippe Morency, and Trevor Darrell. Reducing drift in differential tracking. Computer Vision and Image Understanding: CVIU, 109(2):97– 111, February 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ruan:2002:FMS

[RMFB02] Su Ruan, Bruno Moretti, Jalal Fadili, and Daniel Bloyet. Fuzzy Markovian segmentation in application of magnetic resonance images. Computer Vision and Image Understanding: CVIU, 85(1):54– 69, January 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Reno:2017:TPA

[RMN^{+}17] Vito Renò, Nicola Mosca, Massimiliano Nitti, Tiziana D’Orazio, Cataldo Guaragnella, Donato Campagnoli, Andrea Prati, and Ettore Stella. A technology platform for automatic high-level tennis game analysis. Computer Vision and Image Understanding: CVIU, 159(??):164– 175, June 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300012.

Ruwa:2019:TAN

[RMS^{+}19] Nelson Ruwa, Qirong Mao, Heping Song, Hongjie Jia, and Ming Dong. Triple attention network for sentimental visual question answering. Computer Vision and Image Understanding: CVIU, 189 (??):Article 102829, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421930133X.

Raudies:2012:REM

[RN12] Florian Raudies and Heiko Neumann. A review and evaluation of methods estimating ego-motion. Computer Vision and Image Understanding: CVIU, 116(5): 606–633, May 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212000021.

Robert:1996:CCF

[Rob96a] Luc Robert. Camera calibration without feature extraction. Computer Vision and Image Understanding: CVIU, 63(2):314– 325, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0021/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0021/production/pdf.

Robinson:1996:LSC

[Rob96b] Julia Jean Robinson. Line symmetry of convex digital regions. Computer Vision and Image Understanding: CVIU, 64(2):263– 285, September 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0058/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0058/production/pdf.

Rogez:2014:EPG

[ROGT14] Grégory Rogez, Carlos Orrite, J. J. Guerrero, and Philip H. S. Torr. Exploiting projective geometry for view-invariant monocular human motion analysis in man-made environments. Computer Vision and Image Understanding: CVIU, 120(??):126– 140, March 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213002439.

Ren:2009:TSB

[ROJX09] Jinchang Ren, James Orwell, Graeme A. Jones, and Ming Xu. Tracking the soccer ball using multiple fixed cameras. Computer Vision and Image Understanding: CVIU, 113(5):633– 642, May 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rosenfeld:1995:IAC

[Ros95] Azriel Rosenfeld. Image analysis and computer vision: 1994. Computer Vision and Image Understanding: CVIU, 62(1): 90–143, July 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1044/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1044/production/pdf.

Rosenfeld:1996:IAC

[Ros96] Azriel Rosenfeld. Image analysis and computer vision: 1995. Computer Vision and Image Understanding: CVIU, 63(3): 568–612, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0041/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0041/production/pdf.

Rosenfeld:1997:IAC

[Ros97] Azriel Rosenfeld. Image analysis and computer vision: 1996. Computer Vision and Image Understanding: CVIU, 66(1): 33–93, April 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0602/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0602/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0602/production/ref.

Rosenfeld:1998:IAC

[Ros98] Azriel Rosenfeld. Image analysis and computer vision: 1997. Computer Vision and Image Understanding: CVIU, 70(2): 239–284, May 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0697/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0697/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0697/production/ref.

Rosenfeld:1999:IAC

[Ros99a] Azriel Rosenfeld. Image analysis and computer vision: 1998. Computer Vision and Image Understanding: CVIU, 74(1): 36–95, April 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.

                                        0746/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0746/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0746/production/ref.

Rosin:1999:MCP

[Ros99b] Paul L. Rosin. Measuring corner properties. Computer Vision and Image Understanding: CVIU, 73(2):291– 307, February 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0719/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0719/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0719/production/ref.

Rosenfeld:2000:CLR

[Ros00a]  Azriel Rosenfeld.  Clas-

sifying the literature related to computer vision and image analysis. Computer Vision and Image Understanding: CVIU, 79 (2):308–323, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2000.0851;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.

2000.0851/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0851/ref.

Rosenfeld:2000:IAC

[Ros00b] Azriel Rosenfeld. Image Analysis and Computer Vision: 1999. Computer Vision and Image Understanding: CVIU, 78(2): 222–302, May 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0835;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0835/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0835/ref.

Rosenfeld:2001:IAC

[Ros01] Azriel Rosenfeld. From image analysis to computer vision: An annotated bibliography, 1955–1979. Computer Vision and Image Understanding: CVIU, 84(2): 298–324, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0953;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0953/pdf;   http:
//www.idealibrary.com/

                                        links/doi/10.1006/cviu.
                                        2001.0953/ref.

Rosin:2002:TCD

[Ros02] Paul L. Rosin. Thresholding for change detection. Computer Vision and Image Understanding: CVIU, 86(2): 79–95, May 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rosin:2008:TCR

[Ros08] Paul L. Rosin. A twocomponent rectilinearity measure. Computer Vision and Image Understanding: CVIU, 109(2):176–185, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rosin:2010:IPU

[Ros10] Paul L. Rosin. Image processing using 3-state cellular automata. Computer Vision and Image Understanding: CVIU, 114(7):790– 802, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rameau:2022:MCG

[RPBK22] François Rameau, Jinsun Park, Oleksandr Bailo, and In So Kweon. MC-Calib: a generic and robust calibration toolbox for multicamera systems. Computer Vision and Image Understanding: CVIU, 217(??): ??, March 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221001818.

Rashwan:2012:IRV

[RPG12] Hatem A. Rashwan, Domenec Puig, and Miguel Angel Garcia. Improving the robustness of variational optical flow through tensor voting. Computer Vision and Image Understanding: CVIU, 116(9):953–966, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212000756.

Rubner:2001:EED

[RPTB01] Yossi Rubner, Jan Puzicha, Carlo Tomasi, and Joachim M. Buhmann. Empirical evaluation of dissimilarity measures for color and texture. Computer Vision and Image Understanding: CVIU, 84(1):25–43, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0934;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0934/pdf;   http:

                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0934/ref.

Rivlin:1995:NF

[RR95] Ehud Rivlin and Azriel Rosenfeld. Navigational functionalities. Computer Vision and Image Understanding: CVIU, 62(2):232– 244, September 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1052/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1052/production/pdf.

Robertson:2006:GMH

[RR06] Neil Robertson and Ian Reid. A general method for human activity recognition in video. Computer Vision and Image Understanding: CVIU, 104 (2–3):232–248, November/ December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rivera-Rubio:2016:AHI

[RRAR^{+}16] Jose Rivera-Rubio, Kai Arulkumaran, Hemang Rishi, Ioannis Alexiou, and Anil A. Bharath. An assistive haptic interface for appearancebased indoor navigation. Computer Vision and Image Understanding: CVIU, 149 (??):126–145, August 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314216000680.

Rahman:2013:OAR

[RRK13] Sejuti Rahman and Antonio Robles-Kelly. An optimisation approach to the recovery of reflection parameters from a single hyperspectral image. Computer Vision and Image Understanding: CVIU, 117(12):1672– 1688, December 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001288.

Rad:2020:AAL

[RRL20] Mahdi Rad, Peter M.

Roth, and Vincent Lepetit. ALCN: Adaptive local contrast normalization. Computer Vision and Image Understanding: CVIU, 194(??):Article 102947, May 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300308.

Raskin:2011:DRU

[RRR11] Leonid Raskin, Michael Rudzsky, and Ehud Rivlin. Dimensionality reduction using a Gaussian Process Annealed Particle Filter for tracking and classification of articulated body motions. Computer Vision and Image Understanding: CVIU, 115(4):503–519, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rosales:2003:FHG

[RS03] Rómer Rosales and Stan Sclaroff. A framework for heading-guided recognition of human activity. Computer Vision and Image Understanding: CVIU, 91(3): 335–367, September 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rezaei:2023:DBD

[RSBA23] Mina Rezaei, Farzin Soleymani, Bernd Bischl, and Shekoofeh Azizi. Deep Bregman divergence for selfsupervised representations learning. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001819.

Ramalingam:2010:GSC

[RSL10] Srikumar Ramalingam, Peter Sturm, and Suresh K. Lodha. Generic selfcalibration of central cameras. Computer Vision and Image Understanding: CVIU, 114(2):210–219, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ribnick:2012:RAP

[RSPD12] Evan Ribnick, Ravishankar Sivalingam, Nikolaos Papanikolopoulos, and Kostas Daniilidis. Reconstructing and analyzing periodic human motion from stationary monocular views. Computer Vision and Image Understanding: CVIU, 116(7): 815–826, July 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000513.

Rivlin:2007:ESV

[RSS07] Guy Froimovich Ehud Rivlin, Ilan Shimshoni, and Octavian Soldea. Efficient search and verification for function based classification from real range images. Computer Vision and Image Understanding: CVIU, 105 (3):200–217, March 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rai:2022:TRC

[RSY22] Abhinav Rai, Fadime Sener, and Angela Yao. Transformed ROIs for capturing visual transformations in videos. Computer Vision and Image Understanding: CVIU, 224(??): ??, November 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001369.

Rossi:2014:CFS

[RT14] Luca Rossi and Andrea Torsello. Coarse-to-fine skeleton extraction for high resolution 3D meshes. Computer Vision and Image Understanding: CVIU, 118 (??):140–152, January 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314213001938.

Rumpler:2017:EMS

[RTM^{+}17] Markus Rumpler, Alexander Tscharf, Christian Mostegel, Shreyansh Daftry, Christof Hoppe, Rudolf Prettenthaler, Friedrich Fraundorfer, Gerhard Mayer, and Horst Bischof. Evaluations on multi-scale camera networks for precise and geo-accurate reconstructions from aerial and terrestrial images with user guidance. Computer Vision and Image Understanding: CVIU, 157(??):255– 273, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216300327.

Rivlin:1997:DIO

[RW97] Ehud Rivlin and Isaac Weiss. Deformation invariants in object recognition. Computer Vision and Image Understanding: CVIU, 65(1):95– 108, January 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0478/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0478/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0478/production/ref.

Robey:1995:IUP

[RWV95] M. Robey, G. West, and S. Venkatesh. An investigation into the use of physical modeling for the prediction of various feature types visible from different viewpoints. Computer Vision and Image Understanding: CVIU, 61(3): 417–429, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.

                                        1031/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1031/production/pdf.

Reinhardt:2000:CBS

[RWWH00] Joseph M. Reinhardt, Andrien J. Wang, Thomas P. Weldon, and William E. Higgins. Cue-based segmentation of 4D cardiac image sequences. Computer Vision and Image Understanding: CVIU, 77(2): 251–262, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.1999.0818;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        1999.0818/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        1999.0818/ref.

Ren:2022:PMS

[RXDS22] Guangyu Ren, Yanchun Xie, Tianhong Dai, and Tania Stathaki. Progressive multi-scale fusion network for RGB-D salient object detection. Computer Vision and Image Understanding: CVIU, 223(??): ??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001126.

Reisfeld:1998:PFI

[RY98] Daniel Reisfeld and Yehezkel Yeshurun. Preprocessing of face images: Detection of features and pose normalization. Computer Vision and Image Understanding: CVIU, 71(3):413– 430, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0640/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0640/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0640/production/ref.

Rao:2024:CCM

[RYLZ24] Qi Rao, Xin Yu, Guang Li, and Linchao Zhu. CMGNet: Collaborative multi-modal graph network for video captioning. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002448.

Rosin:2005:MR

[RŽ05] Paul L. Rosin and Joviša Žunić. Measuring rectilinearity. Computer Vision and Image Understanding: CVIU, 99(2):175–188, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Reily:2017:RTG

[RZH17] Brian Reily, Hao Zhang, and William Hoff. Realtime gymnast detection and performance analysis with a portable 3D camera. Computer Vision and Image Understanding: CVIU, 159(??):154– 163, June 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216301953.

Ren:2023:EEW

[RZZ23] Jinwei Ren, Jianke Zhu, and Jialiang Zhang. End-to-end weakly-supervised singlestage multiple 3D hand mesh reconstruction from a single RGB image. Computer Vision and Image Understanding: CVIU, 232 (??):??, July 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223000863.

Sull:1995:IMS

[SA95] Sanghoon Sull and Narendra Ahuja. Integrated matching and segmentation of multiple features in two views. Computer Vision and Image Understanding: CVIU, 62(3):279– 297, November 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1055/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1055/production/pdf.

Sabata:1996:SCM

[SA96] Bikash Sabata and J. K.

Aggarwal. Surface correspondence and motion computation from a pair of range images. Computer Vision and Image Understanding: CVIU, 63(2):232– 250, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0017/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0017/production/pdf.

Stamos:2002:GTR

[SA02] Ioannis Stamos and Peter K.

Allen. Geometry and texture recovery of scenes of large scale. Computer Vision and Image Understanding: CVIU, 88(2):94–118, November 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sato:2004:TSV

[SA04] Koichi Sato and J. K. Aggarwal. Temporal spatiovelocity transform and its application to tracking and interaction. Computer Vision and Image Understanding: CVIU, 96(2):100–128, November 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Singh:2015:LRT

[SA15] Abhishek Singh and Narendra Ahuja. Learning ramp transformation for single image super-resolution. Computer Vision and Image Understanding: CVIU, 135(??):109–125, June 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314215000181.

Skaff:2009:SBA

[SAC09] Sandra Skaff, Tal Arbel, and James J. Clark. A sequential Bayesian approach to color constancy using nonuniform filters. Computer Vision and Image Understanding: CVIU, 113(9): 993–1004, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sharma:2012:RPI

[SAC^{+}12] Abhishek Sharma, Murad Al Haj, Jonghyun Choi, Larry S. Davis, and David W. Jacobs. Robust pose invariant face recognition using coupled latent space discriminant analysis. Computer Vision and Image Understanding: CVIU, 116(11):1095– 1110, November 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001051.

Serino:2014:SBO

[SAdB14] Luca Serino, Carlo Arcelli, and Gabriella Sanniti di Baja. From skeleton branches to object parts. Computer Vision and Image Understanding: CVIU, 129(??):42–51, December 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001040.

Saha:2005:TSL

[Sah05] Punam Kumar Saha. Tensor scale: a local morphometric parameter with applications to computer vision and image processing. Computer Vision and Image Understanding: CVIU, 99(3): 384–413, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sela:2015:CCS

[SAK15] Matan Sela, Yonathan Aflalo, and Ron Kimmel. Computational caricaturization of surfaces. Computer Vision and Image Understanding: CVIU, 141(??):1– 17, December 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001277.

Schwartz:2024:MMF

[SAK^{+}24] Eli Schwartz, Assaf Arbelle, Leonid Karlinsky, Sivan Harary, Florian Scheidegger, Sivan Doveh, and Raja Giryes. MAEDAY: MAE for few- and zero-shot AnomalY-Detection. Computer Vision and Image Understanding: CVIU, 241 (??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000390.

Sun:2016:LAG

[SAL16] Lei Sun, Haizhou Ai, and Shihong Lao. Localizing activity groups in videos. Computer Vision and Image Understanding: CVIU, 144(??):144– 154, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002192.

Sapiro:1997:CS

[Sap97] Guillermo Sapiro. Color snakes. Computer Vision and Image Understanding: CVIU, 68(2):247– 253, November 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0562/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0562/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0562/production/ref.

Sanroma:2012:NGM

[SAS12] Gerard Sanromà, René Alquézar, and Francesc Serratosa. A new graph matching method for point-set correspondence using the EM algorithm and softassign. Computer Vision and Image Understanding: CVIU, 116(2):292–304, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002360.

Solis-Arrazola:2024:EIB

[SASYGCRG24] Manuel A. Solis-Arrazola, Raul E. Sanchez-Yañez, Carlos H. Garcia-Capulin, and Horacio Rostro-Gonzalez. Enhancing image-based facial expression recognition through muscle activationbased facial feature extraction. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000080.

Saund:1999:POO

[Sau99] Eric Saund. Perceptual organization of occluding contours of opaque surfaces. Computer Vision and Image Understanding: CVIU, 76(1):70–82, October 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0789/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0789/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0789/production/ref.

Savran:2023:MTB

[Sav23]  Arman Savran.  Multi-

timescale boosting for efficient and improved event camera face pose alignment. Computer Vision and Image Understanding: CVIU, 236(??):??, November 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223001972.

Shumitskaya:2024:TAR

[SAV24] Ekaterina Shumitskaya, Anastasia Antsiferova, and Dmitriy Vatolin. Towards adversarial robustness verification of no-reference imageand video-quality metrics. Computer Vision and Image Understanding: CVIU, 240 (??):??, March 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422300293X.

Sarkar:1995:UPI

[SB95] Sudeep Sarkar and Kim L.

Boyer. Using perceptual inference networks to manage vision processes. Computer Vision and Image Understanding: CVIU, 62(1): 27–46, July 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1039/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1039/production/pdf.

Sahabi:1996:AED

[SB96a] Hossein Sahabi and Anup Basu. Analysis of error in depth perception with vergence and spatially varying sensing. Computer Vision and Image Understanding: CVIU, 63(3): 447–461, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0034/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0034/production/pdf.

Shaked:1996:CA

[SB96b] Doron Shaked and Alfred M. Bruckstein. The curve axis. Computer Vision and Image Understanding: CVIU, 63(2):367– 379, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0026/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0026/production/pdf.

Sarkar:1998:QMC

[SB98a] Sudeep Sarkar and Kim L. Boyer. Quantitative measures of change based on feature organization: Eigenvalues and eigenvectors. Computer Vision and Image Understanding: CVIU, 71(1): 110–136, July 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0637/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0637/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0637/production/ref.

Sengupta:1998:MPU

[SB98b] Kuntal Sengupta and Kim L.

Boyer. Modelbase partitioning using property matrix spectra. Computer Vision and Image Understanding: CVIU, 70(2): 177–196, May 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0631/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0631/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0631/production/ref.

Shaked:1998:PMA

[SB98c] Doron Shaked and Alfred M. Bruckstein. Pruning medial axes. Computer Vision and Image Understanding: CVIU, 69(2):156– 169, February 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0598/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0598/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0598/production/ref.

Stevens:2000:LSI

[SB00] Mark R. Stevens and J. Ross Beveridge. Localized scene interpretation from 3D models, range, and optical data. Computer Vision and Image Understanding: CVIU, 80(2): 111–129, November 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0821;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0821/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0821/ref.

Svensson:2002:DDT

[SB02] Stina Svensson and Gunilla Borgefors. Digital distance transforms in 3D images using information from neighbourhoods up to 5 × 5 × 5. Computer Vision and Image Understanding: CVIU, 88(1):24–53, October 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Strand:2005:DTT

[SB05] Robin Strand and Gunilla Borgefors. Distance transforms for three-dimensional grids with non-cubic voxels. Computer Vision and Image Understanding: CVIU, 100(3):294–311, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shang:2013:FRF

[SB13] Changjing Shang and Dave Barnes. Fuzzy-rough feature selection aided support vector machines for Mars image classification. Computer Vision and Image Understanding: CVIU, 117(3): 202–213, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001968.

Sankar:2018:MBA

[SB18] Shrinivasan Sankar and Adrien Bartoli. Modelbased active learning to detect an isometric deformable object in the wild with a deep architecture. Computer Vision and Image Understanding: CVIU, 171 (??):69–82, June 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218300705.

Shao:2022:WBM

[SB22] Jun Shao and Tien D. Bui. Wavelet-based multi-level generative adversarial networks for face aging. Computer Vision and Image Understanding: CVIU, 223 (??):??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001084.

Saha:2010:DIE

[SBB10] Sankalita Saha, Neal K. Bambha, and Shuvra S. Bhattacharyya. Design and implementation of embedded computer vision systems based on particle filters. Computer Vision and Image Understanding: CVIU, 114(11):1203–1214, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Schedl:2018:OSV

[SBB18] David C. Schedl, Clemens Birklbauer, and Oliver Bimber. Optimized sampling for view interpolation in light fields using local dictionaries. Computer Vision and Image Understanding: CVIU, 168(??):93– 103, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301212.

Shipitsin:2022:GGA

[SBD22] Viktor Shipitsin, Iaroslav Bespalov, and Dmitry V. Dylov. GAFL: Global adaptive filtering layer for computer vision. Computer Vision and Image Understanding: CVIU, 223(??): ??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001072.

Song:2017:SFI

[SBH^{+}17] Yibing Song, Linchao Bao, Shengfeng He, Qingxiong Yang, and Ming-Hsuan Yang. Stylizing face images via multiple exemplars. Computer Vision and Image Understanding: CVIU, 162(??):135– 145, September 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301467.

Sarafianos:2016:HPE

[SBIK16] Nikolaos Sarafianos, Bogdan Boteanu, Bogdan Ionescu, and Ioannis A. Kakadiaris. 3D human pose estimation: a review of the literature and analysis of covariates. Computer Vision and Image Understanding: CVIU, 152(??):1–20, November 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301369.

Shyu:1999:APL

[SBK^{+}99] Chi-Ren Shyu, Carla E. Brodley, Avinash C. Kak, Akio Kosaka, Alex M. Aisen, and Lynn S. Broderick. ASSERT: a physicianin-the-loop content-based retrieval system for HRCT image databases. Computer Vision and Image Understanding: CVIU, 75(1–2): 111–132, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0768/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0768/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0768/production/ref.

Shokoufandeh:2006:RMC

[SBM^{+}06] Ali Shokoufandeh, Lars Bretzner, Diego Macrini, M. Fatih Demirci, Clas Jönsson, and Sven Dickinson. The representation and matching of categorical shape. Computer Vision and Image Understanding: CVIU, 103(2):139–154, August 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Staranowicz:2015:PAC

[SBMM15] Aaron N. Staranowicz, Garrett R. Brown, Fabio Morbidi, and Gian-Luca Mariottini. Practical and accurate calibration of RGB-D cameras using spheres. Computer Vision and Image Understanding: CVIU, 137(??):102– 114, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000703.

Shivakumara:2024:NDC

[SBN^{+}24] Palaiahnakote Shivakumara, Ayan Banerjee, Lokesh Nandanwar, Umapada Pal, Apostolos Antonacopoulos, Tong Lu, and Michael Blumenstein. A new deep CNN for 3D text localization in the wild through shadow removal. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223002436.

Svensson:2017:SCT

[SBPF17] Carl-Magnus Svensson, Karen Grace Bondoc, Georg Pohnert, and Marc Thilo Figge. Segmentation of clusters by template rotation expectation maximization. Computer Vision and Image Understanding: CVIU, 154 (??):64–72, January 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314216301114.

Sohn:2004:VVC

[SBS04] Bong-Soo Sohn, Chandrajit Bajaj, and Vinay Siddavanahalli. Volumetric video compression for interactive playback. Computer Vision and Image Understanding: CVIU, 96(3):435–452, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sudhir:1997:FPC

[SBZ97] G. Sudhir, Subhashis Banerjee, and Andrew Zisserman. Finding point correspondences in motion sequences preserving affine structure. Computer Vision and Image Understanding: CVIU, 68(2):237– 246, November 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0545/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0545/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0545/production/ref.

Saha:1996:DTU

[SC96] P. K. Saha and B. B. Chaudhuri. 3D digital topology under binary transformation with applications. Computer Vision and Image Understanding: CVIU, 63(3): 418–429, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0032/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0032/production/pdf.

Sziranyi:1998:TCS

[SC98] Tamás Szirányi and Márton Csapodi. Texture classification and segmentation by cellular neural networks using genetic learning. Computer Vision and Image Understanding: CVIU, 71(3):255–270, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0646/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0646/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0646/production/ref.

Sato:1999:EGT

[SC99] Jun Sato and Roberto Cipolla. Extracting group transformations from image moments. Computer Vision and Image Understanding: CVIU, 73(1):29– 42, January 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0702/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0702/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0702/production/ref.

Sarkar:2000:MPS

[SC00a] Sudeep Sarkar and Srikanth Chavali. Modeling parameter space behavior of vision systems using Bayesian networks. Computer Vision and Image Understanding: CVIU, 79(2):185– 223, August 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0854;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0854/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0854/ref.

Squire:2000:ISC

[SC00b] David McG. Squire and Terry M. Caelli. Invariance signatures: Characterizing contours by their departures from invariance. Computer Vision and Image Understanding: CVIU, 77(3):284–316, March 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0809;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0809/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0809/ref.

SanMiguel:2015:TVP

[SC15] Juan C. SanMiguel and Andrea Cavallaro. Temporal validation of Particle Filters for video tracking. Computer Vision and Image Understanding: CVIU, 131(??):42–55, February 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001581.

Saval-Calvo:2018:NRR

[SCALFG^{+}18] Marcelo Saval-Calvo, Jorge Azorin-Lopez, Andres Fuster-Guillo, Victor Villena-Martinez, and Robert B. Fisher. 3D non-rigid registration using color: Color Coherent Point Drift. Computer Vision and Image Understanding: CVIU, 169 (??):119–135, April 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314218300080.

Segalin:2017:SPT

[SCC17] Cristina Segalin, Dong Seon Cheng, and Marco Cristani. Social profiling through image understanding: Personality inference using convolutional neural networks. Computer Vision and Image Understanding: CVIU, 156(??):34– 50, March 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216301679.

Su:2022:VRD

[SCC^{+}22] Yu-Chuan Su, Soravit Changpinyo, Xiangning Chen, Sathish Thoppay, Cho-Jui Hsieh, Lior Shapira, Radu Soricut, Hartwig Adam, Matthew Brown, Ming-Hsuan Yang, and Boqing Gong. 2.5D visual relationship detection. Computer Vision and Image Understanding: CVIU, 224(??): ??, November 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001357.

Sturm:2005:FLC

[SCCP05] P. Sturm, Z. L. Cheng, P. C. Y. Chen, and A. N. Poo. Focal length calibration from two views: method and analysis of singular cases. Computer Vision and Image Understanding: CVIU, 99(1):58– 95, July 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Santamaria:2011:CSS

[SCD11] J. Santamaría, O. Cordón, and S. Damas. A comparative study of state-of-the-art evolutionary image registration methods for 3D modeling. Computer Vision and Image Understanding: CVIU, 115(9):1340– 1354, September 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001366.

Salvador:2004:CSS

[SCE04] Elena Salvador, Andrea Cavallaro, and Touradj Ebrahimi. Cast shadow segmentation using invariant color features. Computer Vision and Image Understanding: CVIU, 95(2):238– 259, August 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Szpak:2014:SDB

[SCEvdH14] Zygmunt L. Szpak, Wojciech Chojnacki, Anders Eriksson, and Anton van den Hengel. Sampson distance based joint estimation of multiple homographies with uncalibrated cameras. Computer Vision and Image Understanding: CVIU, 125(??):200– 213, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214000952.

Santana-Cedres:2017:ACP

[SCGAF^{+}17] Daniel Santana-Cedrés, Luis Gomez, Miguel Alemán-Flores, Agustín Salgado, Julio Esclarín, Luis Mazorra, and Luis Alvarez. Automatic correction of perspective and optical distortions. Computer Vision and Image Understanding: CVIU, 161(??):1–10, August 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421730111X.

Schindler:2006:GCS

[Sch06] Konrad Schindler. Geometry and construction of straight lines in log-polar images. Computer Vision and Image Understanding: CVIU, 103(3):196–207, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shin:2013:MOR

[SCL13] Young Min Shin, Minsu Cho, and Kyoung Mu Lee. Multi-object reconstruction from dynamic scenes: an object-centered approach. Computer Vision and Image Understanding: CVIU, 117(11):1575– 1588, November 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001276.

Somasundaram:2014:ARU

[SCMP14] Guruprasad Somasundaram, Anoop Cherian, Vassilios Morellas, and Nikolaos Papanikolopoulos. Action recognition using global spatio-temporal features derived from sparse representations. Computer Vision and Image Understanding: CVIU, 123(??): 1–13, June 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000095.

Strand:2013:MBD

[SCMS13] Robin Strand, Krzysztof Chris Ciesielski, Filip Malmberg, and Punam K. Saha. The minimum barrier distance. Computer Vision and Image Understanding: CVIU, 117(4):429– 437, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001750.

Setti:2017:HDN

[SCR^{+}17] Francesco Setti, Davide Conigliaro, Paolo Rota, Chiara Bassetti, Nicola Conci, Nicu Sebe, and Marco Cristani. The S-Hock dataset: a new benchmark for spectator crowd analysis. Computer Vision and Image Understanding: CVIU, 159 (??):47–58, June 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300024.

Ser:1999:GAE

[SCS99] P. K. Ser, Clifford S. T.

Choy, and W. C. Siu. Genetic algorithm for the extraction of nonanalytic objects from multiple dimensional parameter space. Computer Vision and Image Understanding: CVIU, 73 (1):1–13, January 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0695/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0695/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0695/production/ref.

Solari:2014:INA

[SCS14] Fabio Solari, Manuela Chessa, and Silvio P. Sabatini. An integrated neuromimetic architecture for direct motion interpretation in the logpolar domain. Computer Vision and Image Understanding: CVIU, 125(??): 37–54, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000381.

Sobieranski:2011:LND

[SCvW11] Antonio Carlos Sobieranski, Eros Comunello, and Aldo von Wangenheim.

Learning a nonlinear distance metric for supervised region-merging image segmentation. Computer Vision and Image Understanding: CVIU, 115(2):127–139, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Svensson:2003:SCS

[SdB03] Stina Svensson and Gabriella Sanniti di Baja. Simplifying curve skeletons in volume images. Computer Vision and Image Understanding: CVIU, 90(3):242– 257, June 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sharma:2022:FAF

[SDK22] Shailza Sharma, Abhinav Dhall, and Vinay Kumar. Frequency aware face hallucination generative adversarial network with semantic structural constraint. Computer Vision and Image Understanding: CVIU, 223 (??):??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422200131X.

Schieber:2024:ISD

[SDK^{+}24] Hannah Schieber, Kubilay Can Demir, Constantin Kleinbeck, Seung Hee Yang, and Daniel Roth. Indoor synthetic data generation: a systematic review. Computer Vision and Image Understanding: CVIU, 240 (??):??, March 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002874.

Saez:2011:EMS

[SE11] Juan M. Sáez and Francisco Escolano. 6DOF entropy minimization SLAM for stereo-based wearable devices. Computer Vision and Image Understanding: CVIU, 115(2):270–285, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sanchez-Escobedo:2015:SFS

[SECS15] Dalila Sánchez-Escobedo, Mario Castelán, and William A. P. Smith. Statistical 3D face shape estimation from occluding contours. Computer Vision and Image Understanding: CVIU, 142(??):111– 124, ???? 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001885.

Schonborn:2015:BMG

[SEFV15] Sandro Schönborn, Bernhard Egger, Andreas Forster, and Thomas Vetter. Background modeling for generative image models. Computer Vision and Image Understanding: CVIU, 136 (??):117–127, July 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314215000235.

Seales:1995:BTD

[SF95] W. Brent Seales and Olivier D. Faugeras. Building threedimensional object models from image sequences. Computer Vision and Image Understanding: CVIU, 61(3): 308–324, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1025/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1025/production/pdf.

Soucy:1997:SRR

[SF97] G. Soucy and F. P. Ferrie. Surface recovery from range images using curvature and motion consistency. Computer Vision and Image Understanding: CVIU, 65(1): 1–18, January 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/

links/artid/cviu.1996.
0485/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0485/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0485/production/ref.

Shi:2007:QCT

[SF07] Lilong Shi and Brian Funt.

Quaternion color texture segmentation. Computer Vision and Image Understanding: CVIU, 107(1–2): 88–96, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Smith:2016:HPR

[SF16] William Smith and Fufu Fang. Height from photometric ratio with modelbased light source selection. Computer Vision and Image Understanding: CVIU, 145(??):128– 138, April 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002635.

Sabokrou:2018:DAF

[SFF^{+}18] Mohammad Sabokrou, Mohsen Fayyaz, Mahmood Fathy, Zahra. Moayed, and Reinhard Klette. Deep-anomaly: Fully convolutional neural network for fast anomaly detection in crowded scenes.

Computer Vision and Image Understanding: CVIU, 172 (??):88–97, July 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218300249.

Santini:2018:PRP

[SFK18] Thiago Santini, Wolfgang Fuhl, and Enkelejda Kasneci. PuRe: Robust pupil detection for real-time pervasive eye tracking. Computer Vision and Image Understanding: CVIU, 170 (??):40–50, May 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218300146.

Shavit:2024:LSM

[SFK24] Yoli Shavit, Ron Ferens, and Yosi Keller. Learning single and multi-scene camera pose regression with transformer encoders. Computer Vision and Image Understanding: CVIU, 243(??): ??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314224000638.

Sun:2008:CVM

[SFWG08] Yaoru Sun, Robert Fisher, Fang Wang, and Herman Martins Gomes. A computer vision model for visual-object-based attention and eye movements. Computer Vision and Image Understanding: CVIU, 112(2):126–142, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sengur:2011:CTI

[SG11] Abdulkadir Sengur and Yanhui Guo. Color texture image segmentation based on neutrosophic set and wavelet transformation. Computer Vision and Image Understanding: CVIU, 115(8):1134–1144, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211000993.

Srikantha:2017:WSD

[SG17] Abhilash Srikantha and Juergen Gall. Weak supervision for detecting object classes from activities. Computer Vision and Image Understanding: CVIU, 156(??):138– 150, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301400.

Shu:2012:AAM

[SGA12] Xianbiao Shu, Chunyu Gao, and Narendra Ahuja. Aperture access and manipulation for computational imaging. Computer Vision and Image Understanding: CVIU, 116(2):222–237, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002384.

Shin:2001:CED

[SGB01] Min C. Shin, Dmitry B. Goldgof, and Kevin W. Bowyer. Comparison of edge detector performance through use in an object recognition task. Computer Vision and Image Understanding: CVIU, 84 (1):160–178, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0932;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0932/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0932/ref.

Song:2001:MPB

[SGDP01] Yang Song, Luis Goncalves, Enrico Di Bernardo, and Pietro Perona. Monocular perception of biological motion in Johansson displays. Computer Vision and Image Understanding: CVIU, 81(3):303–327, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0890;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0890/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0890/ref.

Stefanov:2007:RTH

[SGH07] Nikolay Stefanov, Aphrodite Galata, and Roger Hubbold. A real-time hand tracker using variable-length Markov models of behaviour. Computer Vision and Image Understanding: CVIU, 108(1–2):98–115, October/ November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Stokman:2000:CMI

[SGK00] H. M. G. Stokman, Th.

Gevers, and J. J. Koenderink. Color measurement by imaging spectrometry. Computer Vision and Image Understanding: CVIU, 79(2):236–249, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0860;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.

                                        2000.0860/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2000.0860/ref.

Serra:2015:GGL

[SGMC15] Giuseppe Serra, Costantino Grana, Marco Manfredi, and Rita Cucchiara. GOLD: Gaussians of Local Descriptors for image representation. Computer Vision and Image Understanding: CVIU, 134(??):22– 32, May 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215000193.

Sabatini:2010:CHC

[SGS^{+}10] Silvio P. Sabatini, Giulia Gastaldi, Fabio Solari, Karl Pauwels, Marc M. Van Hulle, Javier Diaz, Eduardo Ros, Nicolas Pugeault, and Norbert Krüger. A compact harmonic code for early vision based on anisotropic frequency channels. Computer Vision and Image Understanding: CVIU, 114(6): 681–699, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Su:2021:MPC

[SGZ21] Hang Su, Shaogang Gong, and Xiatian Zhu. Multiperspective cross-class domain adaptation for open logo detection. Computer Vision and Image Understanding: CVIU, 204(??): Article 103156, March 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220301752.

Starck:2008:MBH

[SH08] Jonathan Starck and Adrian Hilton. Model-based human shape reconstruction from multiple views. Computer Vision and Image Understanding: CVIU, 111(2): 179–194, August 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Schmitt:2009:MMD

[SH09] Oliver Schmitt and Maria Hasse. Morphological multiscale decomposition of connected regions with emphasis on cell clusters. Computer Vision and Image Understanding: CVIU, 113 (2):188–201, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shah:2005:GSS

[Sha05] Jayant Shah. Gray skeletons and segmentation of shapes. Computer Vision and Image Understanding: CVIU, 99 (1):96–109, July 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shapiro:2006:ASL

[Sha06] Vladimir Shapiro. Accuracy of the straight line Hough Transform: The nonvoting approach. Computer Vision and Image Understanding: CVIU, 103(1):1– 21, July 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shah:2011:CTS

[Sha11]  Mili Shah.  Comparing

two sets of corresponding six degree of freedom data. Computer Vision and Image Understanding: CVIU, 115(10):1355–1362, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211001378.

Shekhovtsov:2016:HOM

[She16]  Alexander  Shekhovtsov.

Higher order maximum persistency and comparison theorems. Computer Vision and Image Understanding: CVIU, 143(??):54–79, February 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215000958.

Seibold:2017:MBM

[SHE17] Clemens Seibold, Anna Hilsmann, and Peter Eisert. Model-based motion blur estimation for the improvement of motion tracking. Computer Vision and Image Understanding: CVIU, 160 (??):45–56, July 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300590.

Shimshoni:1999:EUL

[Shi99] Ilan Shimshoni. On estimating the uncertainty in the location of image points in 3D recognition from match sets of different sizes. Computer Vision and Image Understanding: CVIU, 74(3): 163–173, June 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0755/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0755/production/pdf.

Sibbing:2011:MRS

[SHK11] Dominik Sibbing, Martin Habbecke, and Leif Kobbelt. Markerless reconstruction and synthesis of dynamic facial expressions. Computer Vision and Image Understanding: CVIU, 115(5):

668–680, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sim:1998:AMK

[SHKP98] Dong-Gyu Sim, Young Kug Ham, In Kwon Kim, and Rae-Hong Park. Analysis of mixed Korean documents using the branch and bound algorithm based on DP matching. Computer Vision and Image Understanding: CVIU, 71(3):373– 384, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0651/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0651/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0651/production/ref.

Stutz:2018:SES

[SHL18] David Stutz, Alexander Hermans, and Bastian Leibe. Superpixels: an evaluation of the state-of-the-art. Computer Vision and Image Understanding: CVIU, 166(??):1–27, January 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300589.

Suzuki:2003:LTC

[SHS03] Kenji Suzuki, Isao Horiba, and Noboru Sugie. Lineartime connected-component labeling based on sequential local operations. Computer Vision and Image Understanding: CVIU, 89(1):1– 23, January 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Song:2023:DCH

[SHS^{+}23] Ge Song, Kai Huang, Hanwen Su, Fengyi Song, and Ming Yang. Deep continual hashing for realworld multi-label image retrieval. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001224.

Siry:2023:IBD

[SHSJ23] Rodrigue Siry, Louis Hémadou, Loïc Simon, and Frédéric Jurie. On the inductive biases of deep domain adaptation. Computer Vision and Image Understanding: CVIU, 233(??): ??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000942.

Shi:2024:EED

[SHW24] Fengyuan Shi, Weilin Huang, and Limin Wang. Endto-end dense video grounding via parallel regression. Computer Vision and Image Understanding: CVIU, 242(??):??, May 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000614.

Sclaroff:2003:ABR

[SI03] Stan Sclaroff and John Isidoro. Active blobs: region-based, deformable appearance models. Computer Vision and Image Understanding: CVIU, 89 (2–3):197–225, February/ March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Soviany:2021:CSP

[SIRS21] Petru Soviany, Radu Tudor Ionescu, Paolo Rota, and Nicu Sebe. Curriculum self-paced learning for cross-domain object detection. Computer Vision and Image Understanding: CVIU, 204(??):Article 103166, March 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000102.

Sumi:2007:RVA

[SIT07] Yasushi Sumi, Yutaka Ishiyama, and Fumiaki Tomita. Robot-vision architecture for real-time 6-DOF object localization. Computer Vision and Image Understanding: CVIU, 105(3): 218–230, March 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Satherley:2001:VCV

[SJ01] Richard Satherley and Mark W. Jones. Vector-city vector distance transform. Computer Vision and Image Understanding: CVIU, 82(3):238–254, June 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0915;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0915/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0915/ref.

Sharma:2015:LHO

[SJ15a] Gaurav Sharma and Frédéric Jurie. Local Higher-Order Statistics (LHS) describing images with statistics of local non-binarized pixel patterns. Computer Vision and Image Understanding: CVIU, 142(??):13– 22, ???? 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215002027.

Sicre:2015:DPM

[SJ15b] Ronan Sicre and Frédéric Jurie. Discriminative part model for visual recognition. Computer Vision and Image Understanding: CVIU, 141(??):28–37, December 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215001642.

Spies:2002:RFE

[SJB02] Hagen Spies, Bernd Jähne, and John L. Barron. Range flow estimation. Computer Vision and Image Understanding: CVIU, 85(3):209– 231, March 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Saraee:2020:VCA

[SJB20] Elham Saraee, Mona Jalal, and Margrit Betke. Visual complexity analysis using deep intermediate-layer features. Computer Vision and Image Understanding: CVIU, 195(??):Article 102949, June 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300333.

Sochor:2017:TSC

[SJH17] Jakub Sochor, Roman Juránek, and Adam Herout. Traffic surveillance camera calibration by 3D model bounding box alignment for accurate vehicle speed measurement. Computer Vision and Image Understanding: CVIU, 161(??): 87–98, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301108.

Si:2021:SAH

[SJSL21] Jiaxin Si, Fei Jiang, Ruimin Shen, and Hongtao Lu. Small and accurate heatmapbased face alignment via distillation strategy and cascaded architecture. Computer Vision and Image Understanding: CVIU, 203 (??):Article 103125, February 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220301442.

Schmugge:2007:OEA

[SJST07] Stephen J. Schmugge, Sriram Jayaram, Min C. Shin, and Leonid V. Tsap. Objective evaluation of approaches of skin detection using ROC analysis.

Computer Vision and Image Understanding: CVIU, 108(1–2):41–51, October/ November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Soffer:1998:GCH

[SK98] Menashe Soffer and Nahum Kiryati. Guaranteed convergence of the Hough transform. Computer Vision and Image Understanding: CVIU, 69(2):119– 134, February 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0557/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0557/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0557/production/ref.

Sablatnig:2002:MBR

[SK02] Robert Sablatnig and Martin Kampel. Model-based registration of front- and backviews of rotationally symmetric objects. Computer Vision and Image Understanding: CVIU, 87(1– 3):90–103, July 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shtern:2015:SGF

[SK15] Alon Shtern and Ron Kimmel. Spectral gradient fields embedding for nonrigid shape matching. Computer Vision and Image Understanding: CVIU, 140(??): 21–29, November 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000387.

Saeki:2023:MPA

[SKA23] Shozo Saeki, Minoru Kawahara, and Hirohisa Aman. Multi proxy anchor family loss for several types of gradients. Computer Vision and Image Understanding: CVIU, 229(??): ??, March 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000346.

Shimshoni:1996:GSS

[SKB96] Ilan Shimshoni, Ron Kimmel, and Alfred M. Bruckstein. Global shape from shading. Computer Vision and Image Understanding: CVIU, 64(1): 188–189, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0053/production;   http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0053/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0054/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0054/production/pdf.

Sun:2013:ODS

[SKBS13] Min Sun, Shyam Sunder Kumar, Gary Bradski, and Silvio Savarese. Object detection, shape recovery, and 3D modelling by depth-encoded Hough voting. Computer Vision and Image Understanding: CVIU, 117(9):1190–1202, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213000969.

Shekhovtsov:2008:EMD

[SKH08]  Alexander  Shekhovtsov,

Ivan Kovtun, and Václav Hlaváč. Efficient MRF deformation model for nonrigid image matching. Computer Vision and Image Understanding: CVIU, 112(1): 91–99, October 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Saada:2022:MOT

[SKLM22] Mohamad Saada, Christos Kouppas, Baihua Li, and Qinggang Meng. A multiobject tracker using dynamic Bayesian networks and a residual neural network based similarity estimator. Computer Vision and Image Understanding: CVIU, 225(??):??, December 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222001473.

Sminchisescu:2006:CMC

[SKM06] Cristian Sminchisescu, Atul Kanaujia, and Dimitris Metaxas. Conditional models for contextual human motion recognition. Computer Vision and Image Understanding: CVIU, 104 (2–3):210–220, November/ December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sakane:1995:PFA

[SKOS95] Shigeyuki Sakane, Toshiji Kuruma, Toru Omata, and Tomomasa Sato. Planning focus of attention for multifingered hand with consideration of time-varying aspects. Computer Vision and Image Understanding: CVIU, 61(3): 445–453, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1033/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1033/production/pdf.

Strandmark:2011:PDV

[SKS11] Petter Strandmark, Fredrik Kahl, and Thomas Schoenemann. Parallel and distributed vision algorithms using dual decomposition. Computer Vision and Image Understanding: CVIU, 115(12):1721– 1732, December 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001652.

Struski:2022:LLC

[SKS^{+}22] Lukasz Struski, Szymon Knop, Przemysław Spurek, Wiktor Daniec, and Jacek Tabor. LocoGAN — locally convolutional GAN. Computer Vision and Image Understanding: CVIU, 221 (??):??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222000728.

Smith:2008:RCR

[SKSR08] Eric R. Smith, Bradford J. King, Charles V. Stewart, and Richard J. Radke. Registration of combined range– intensity scans: Initialization through verification. Computer Vision and Image Understanding: CVIU, 110(2):226–244, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Stafylakis:2018:PBA

[SKT18] Themos Stafylakis, Muhammad Haris Khan, and Georgios Tzimiropoulos. Pushing the boundaries of audiovisual word recognition using residual networks and LSTMs. Computer Vision and Image Understanding: CVIU, 176– 177(??):22–32, November/ December 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218303904.

Shinozaki:2009:CCI

[SKU^{+}09] Megumi Shinozaki, Masato Kusanagi, Kazunori Umeda, Guy Godin, and Marc Rioux. Correction of color information of a 3D model using a range intensity image. Computer Vision and Image Understanding: CVIU, 113(11):1170–1179, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Smeets:2013:MLS

[SKVS13] Dirk Smeets, Johannes Keustermans, Dirk Vandermeulen, and Paul Suetens. meshSIFT: Local surface features for 3D face recognition under expression variations and partial data. Computer Vision and Image Understanding: CVIU, 117 (2):158–169, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314212001324.

Soucy:1996:MSM

[SL96] Marc Soucy and Denis Laurendeau. Multiresolution surface modeling based on hierarchical triangulation. Computer Vision and Image Understanding: CVIU, 63 (1):1–14, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0001/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0001/production/pdf.

Smith:1999:ICQ

[SL99] John R. Smith and Chung-Sheng Li. Image classification and querying using composite region templates. Computer Vision and Image Understanding: CVIU, 75(1–2):165– 174, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0771/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0771/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0771/production/ref.

Strobl:2016:SCF

[SL16a] Klaus H. Strobl and Martin Lingenauber. Stepwise calibration of focused plenoptic cameras. Computer Vision and Image Understanding: CVIU, 145(??): 140–147, April 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421500274X.

Su:2016:DLM

[SL16b] Lin Su and Martin Levine.

Does “lie to me” lie to you? An evaluation of facial clues to high-stakes deception. Computer Vision and Image Understanding: CVIU, 147 (??):52–68, June 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000345.

Sarbolandi:2015:KRS

[SLK15] Hamed Sarbolandi, Damien Lefloch, and Andreas Kolb. Kinect range sensing: Structuredlight versus time-of-flight Kinect. Computer Vision and Image Understanding: CVIU, 139(??):1–20, October 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215001071.

Seo:2023:CFA

[SLK23] Seungwan Seo, Yunseung Lee, and Pilsung Kang. Cost-free adversarial defense: Distance-based optimization for model robustness without adversarial training. Computer Vision and Image Understanding: CVIU, 227(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001771.

Song:2001:MBO

[SLL01] Bong Seop Song, Kyoung Mu Lee, and Sang Uk Lee. Model-based object recognition using geometric invariants of points and lines. Computer Vision and Image Understanding: CVIU, 84(3):361–383, December 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sole:2001:CED

[SLS01] Andres Fco. Solé, Antonio López, and Guillermo Sapiro. Crease enhancement diffusion. Computer Vision and Image Understanding: CVIU, 84(2): 241–248, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0945;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0945/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0945/ref.

Sebe:2003:VRS

[SLS03] Nicu Sebe, Michael S.

Lew, and Arnold W. M. Smeulders. Video retrieval and summarization. Computer Vision and Image Understanding: CVIU, 92 (2–3):141–146, November/ December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sclaroff:1999:UTV

[SLST99] Stan Sclaroff, Marco La Cascia, Saratendu Sethi, and Leonid Taycher. Unifying textual and visual cues for content-based image retrieval on the world wide web. Computer Vision and Image Understanding: CVIU, 75(1–2):86–98, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0765/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0765/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0765/production/ref.

Syeda-Mahmood:1997:DMD

[SM97] Tanveer Fathima Syeda-Mahmood. Data- and model-driven selection using parallel line groups. Computer Vision and Image Understanding: CVIU, 67(3): 205–222, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0542/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0542/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0542/production/ref.

Syeda-Mahmood:1999:DPS

[SM99] T. F. Syeda-Mahmood. Detecting perceptually salient texture regions in images. Computer Vision and Image Understanding: CVIU, 76(1):93–108, October 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0784/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0784/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0784/production/ref.

Sekkati:2006:JOF

[SM06] H. Sekkati and A. Mitiche.

Joint optical flow estimation, segmentation, and 3D interpretation with level sets. Computer Vision and Image Understanding: CVIU, 103(2):89–100, August 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sugano:2010:HOI

[SM10] Hiroki Sugano and Ryusuke Miyamoto. Highly optimized implementation of OpenCV for the Cell Broadband Engine. Computer Vision and Image Understanding: CVIU, 114(11): 1273–1281, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

SanMiguel:2012:SBP

[SM12] Juan C. SanMiguel and José M. Martínez. A semantic-based probabilistic approach for realtime video event recognition. Computer Vision and Image Understanding: CVIU, 116(9):937–952, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314212000665.

Sparks:2013:SSM

[SM13a] Rachel Sparks and Anant Madabhushi. Statistical shape model for manifold regularization: Gleason grading of prostate histology. Computer Vision and Image Understanding: CVIU, 117(9):1138– 1146, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213000672.

Subburaman:2013:AST

[SM13b] Venkatesh Bala Subburaman and Sébastien Marcel. Alternative search techniques for face detection using location estimation and binary features. Computer Vision and Image Understanding: CVIU, 117(5): 551–570, May 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000039.

Stein:2017:RCA

[SM17] Sebastian Stein and Stephen J.

McKenna. Recognising complex activities with histograms of relative tracklets. Computer Vision and Image Understanding: CVIU, 154(??):82–93, January 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301345.

Shedligeri:2021:HFR

[SM21] Prasan Shedligeri and Kaushik Mitra. High frame rate optical flow estimation from event sensors via intensity estimation. Computer Vision and Image Understanding: CVIU, 208–209(??): ??, July 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221000527.

Shin:2022:LCM

[SM22] Jungkyoo Shin and Jinyoung Moon. Learning to combine the modalities of language and video for temporal moment localization. Computer Vision and Image Understanding: CVIU, 217 (??):??, March 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222000145.

Sousa:2023:SRI

[SM23] Eduardo V. Sousa and Douglas G. Macharet. Structural reasoning for imagebased social relation recognition. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001650.

Sokhandan:2024:VTV

[SM24] Alireza Sokhandan and Amirhassan Monadjemi. Visual tracking in video sequences based on biologically inspired mechanisms. Computer Vision and Image Understanding: CVIU, 239 (??):??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218303898.

Singh:2008:LVC

[SMD^{+}08] Vikas Singh, Lopamudra Mukherjee, Petru M. Dinu, Jinhui Xu, and Kenneth R. Hoffmann. Limited view CT reconstruction and segmentation via constrained metric labeling. Computer Vision and Image Understanding: CVIU, 112(1):67– 80, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Syeda-Mahmood:2004:CSI

[SMHH04] Tanveer Syeda-Mahmood, Ismail Haritaoglu, and Thomas Huang. CVIU special issue on event detection in video. Computer Vision and Image Understanding: CVIU, 96(2):97–99, November 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sarkar:2002:POB

[SMK02] Sudeep Sarkar, Daniel Majchrzak, and Kishore Korimilli. Perceptual organization based computational model for robust segmentation of moving objects. Computer Vision and Image Understanding: CVIU, 86(3):141– 170, June 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Selinger:1999:PGH

[SN99] Andrea Selinger and Randal C. Nelson. A perceptual grouping hierarchy for appearance-based 3D object recognition. Computer Vision and Image Understanding: CVIU, 76(1):83– 92, October 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0788/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0788/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0788/production/ref.

Shahid:2007:ILS

[SO07] Kamran Shahid and Galina Okouneva. Intelligent LI-DAR scanning region selection for satellite pose estimation. Computer Vision and Image Understanding: CVIU, 107(3):203–209, September 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Smeaton:2010:VSB

[SOD10] Alan F. Smeaton, Paul Over, and Aiden R. Doherty. Video shot boundary detection: Seven years of TRECVid activity. Computer Vision and Image Understanding: CVIU, 114(4): 411–418, April 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sobh:1995:IIR

[SOJ^{+}95] Tarek M. Sobh, J. Owen, C. Jaynes, M. Dekhil, and T. C. Henderson. Industrial inspection and reverse engineering. Computer Vision and Image Understanding: CVIU, 61(3): 468–474, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1035/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1035/production/pdf.

Shukla:2017:NCM

[SOJ17] Deepika Shukla, Aparajita Ojha, and Rajib Kumar Jha. A new composite multi-constrained differential-radon warping approach for digital video affine motion stabilization. Computer Vision and Image Understanding: CVIU, 155(??):83–105, February 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301977.

Sakurada:2016:HMM

[SOK16] Ken Sakurada, Takayuki Okatani, and Kris M. Kitani. Hybrid macro-micro visual analysis for city-scale state estimation. Computer Vision and Image Understanding: CVIU, 146(??):

86–98, May 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216000710.

Sashida:2014:CML

[SOL14] Satoshi Sashida, Yutaka Okabe, and Hwee Kuan Lee. Comparison of multilabel graph cuts method and Monte Carlo simulation with block-spin transformation for the piecewise constant Mumford–Shah segmentation model. Computer Vision and Image Understanding: CVIU, 119(??): 15–26, February 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213002051.

Sashida:2016:AMC

[SOL16] Satoshi Sashida, Yutaka Okabe, and Hwee Kuan Lee. Application of Monte Carlo simulation with blockspin transformation based on the Mumford–Shah segmentation model to threedimensional biomedical images. Computer Vision and Image Understanding: CVIU, 152(??):176– 189, November 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300777.

Shimshoni:1997:RSP

[SP97a] Ilan Shimshoni and Jean Ponce. Recovering the shape of polyhedra using line-drawing analysis and complex reflectance models. Computer Vision and Image Understanding: CVIU, 65(2):296– 310, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0569/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0569/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0569/production/ref.

Sim:1997:TSA

[SP97b] Dong-Gyu Sim and Rae-Hong Park. A twostage algorithm for motion discontinuity-preserving optical flow estimation. Computer Vision and Image Understanding: CVIU, 65(1): 19–37, January 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0483/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.

                                        0483/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0483/production/ref.

Sinha:2006:PTZ

[SP06] Sudipta N. Sinha and Marc Pollefeys. Pan-tilt-zoom camera calibration and high-resolution mosaic generation. Computer Vision and Image Understanding: CVIU, 103(3):170–183, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Stergiou:2019:AHH

[SP19] Alexandros Stergiou and Ronald Poppe. Analyzing human-human interactions: a survey. Computer Vision and Image Understanding: CVIU, 188(??):Article 102799, November 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314219301158.

S:2023:CDF

[SP23a] Saranya M. S. and Geetha P. Cross-domain fashion cloth retrieval via novel attentionguided cascade neural network and clothing parsing. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001571.

Sarang:2023:TLS

[SP23b] Nima Sarang and Charalambos Poullis. Tractable large-scale deep reinforcement learning. Computer Vision and Image Understanding: CVIU, 232(??): ??, July 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000693.

Salvagnini:2015:NMI

[SPC^{+}15] Pietro Salvagnini, Federico Pernici, Marco Cristani, Giuseppe Lisanti, Alberto Del Bimbo, and Vittorio Murino. Non-myopic information theoretic sensor management of a single pan-tilt-zoom camera for multiple object detection and tracking. Computer Vision and Image Understanding: CVIU, 134(??): 74–88, May 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002392.

Spetsakis:1997:OFE

[Spe97] Minas E. Spetsakis. Optical flow estimation using discontinuity conforming filters. Computer Vision and Image Understanding: CVIU, 68(3):276– 289, December 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0555/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0555/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0555/production/ref.

Spitz:1998:ACD

[Spi98] A. Lawrence Spitz. Analysis of compressed document images for dominant skew, multiple skew, and logotype detection. Computer Vision and Image Understanding: CVIU, 70(3): 321–334, June 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0686/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0686/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0686/production/ref.

Shyu:2002:UHP

[SPK^{+}02] Chi-Ren Shyu, Christina Pavlopoulou, Avinash C. Kak, Carla E. Brodley, and Lynn S. Broderick. Using human perceptual categories for content-based retrieval from a medical image database. Computer Vision and Image Understanding: CVIU, 88(3):119–151, December 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Spampinato:2014:TBK

[SPK14] C. Spampinato, S. Palazzo, and I. Kavasidis. A textonbased kernel density estimation approach for background modeling under extreme conditions. Computer Vision and Image Understanding: CVIU, 122(??): 74–83, May 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213002348.

Singha:2023:IST

[SPK23] Tanmay Singha, Duc-Son Pham, and Aneesh Krishna. Improved Short-term Dense Bottleneck network for efficient scene analysis. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001753.

Stamos:2017:SIL

[SPQ^{+}17] Ioannis Stamos, Marc Pollefeys, Long Quan, Philippos Mordohai, and Yasutaka Furukawa. Special issue on large-scale 3D modeling of urban indoor or outdoor scenes from images and range scans. Computer Vision and Image Understanding: CVIU, 157 (??):1–2, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300413.

Sortino:2023:TBI

[SPRS23] Renato Sortino, Simone Palazzo, Francesco Rundo, and Concetto Spampinato. Transformer-based image generation from scene graphs. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001017.

Savelonas:2018:SSS

[SPT^{+}18] Michalis A. Savelonas, Ioannis Pratikakis, Theoharis Theoharis, Georgios Thanellas, Frédéric Abad, and Rémy Bendahan. Spatially sensitive statistical shape analysis for pedestrian recognition from LI-DAR data. Computer Vision and Image Understanding: CVIU, 171(??): 1–9, June 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218300766.

Steinberg:2015:HBM

[SPW15] Daniel M. Steinberg, Oscar Pizarro, and Stefan B. Williams. Hierarchical Bayesian models for unsupervised scene understanding. Computer Vision and Image Understanding: CVIU, 131(??):128– 144, February 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001313.

Sharma:2023:PBR

[SR23] Renu Sharma and Arun Ross. Periocular biometrics and its relevance to partially masked faces: a survey. Computer Vision and Image Understanding: CVIU, 226 (??):??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001618.

Schlett:2021:DLB

[SRB21] Torsten Schlett, Christian Rathgeb, and Christoph Busch. Deep learning-based single image face depth data enhancement. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000916.

Segvic:2009:MLF

[ŠRDC09] Siniša Šegvić, Anthony Remazeilles, Albert Diosi, and François Chaumette. A mapping and localization framework for scalable appearance-based navigation. Computer Vision and Image Understanding: CVIU, 113(2):172–187, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Suau:2013:DEE

[SRHC13] Xavier Suau, Javier Ruiz-Hidalgo, and Josep R. Casas. Detecting endeffectors on 2.5D data using geometric deformable models: Application to human pose estimation. Computer Vision and Image Understanding: CVIU, 117(3): 281–288, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001907.

Singh:2020:RHF

[SRM20] Vikram Singh, Keerthan Ramnath, and Anurag Mittal. Refining highfrequencies for sharper super-resolution and deblurring. Computer Vision and Image Understanding: CVIU, 199(??):Article 103034, October 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300874.

Seewald:2019:TAM

[SRO^{+}19] Lucas Adams Seewald, Vinicius Facco Rodrigues, Malte Ollenschläger, Rodolfo Stoffel Antunes, Cristiano André da Costa, Rodrigo da Rosa Righi, Luiz Gonzaga da Silveira, Andreas Maier, Björn Eskofier, and Rebecca Fahrig. Toward analyzing mutual interference on infrared-enabled depth cameras. Computer Vision and Image Understanding: CVIU, 178(??):1–15, January 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218303400.

Subashini:2010:AAB

[SRP10] T. S. Subashini, V. Ramalingam, and S. Palanivel. Automated assessment of breast tissue density in digital mammograms. Computer Vision and Image Understanding: CVIU, 114(1): 33–43, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sole-Ribalta:2011:MAC

[SRS11] Albert Solé-Ribalta and Francesc Serratosa. Models and algorithms for computing the common labelling of a set of attributed graphs. Computer Vision and Image Understanding: CVIU, 115 (7):929–945, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000750.

Sun:2001:ECF

[SRT01] Zhaohui Sun, Visvanathan Ramesh, and A. Murat Tekalp. Error characterization of the factorization method. Computer Vision and Image Understanding: CVIU, 82(2): 110–137, May 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0910;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0910/pdf;   http:
//www.idealibrary.com/

links/doi/10.1006/cviu.
2001.0910/ref.

Shirazi:2009:OED

[SS09] Sayed Kamaledin Ghiasi Shirazi and Reza Safabakhsh. Omnidirectional edge detection. Computer Vision and Image Understanding: CVIU, 113(4):556– 564, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sundareswara:2011:BDC

[SS11] Rashmi Sundareswara and Paul R. Schrater. Bayesian discounting of camera parameter uncertainty for optimal 3D reconstruction from images. Computer Vision and Image Understanding: CVIU, 115(1):117–126, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Savran:2017:NRR

[SS17a] Arman Savran and Bülent Sankur. Non-rigid registration based model-free 3D facial expression recognition. Computer Vision and Image Understanding: CVIU, 162(??):146– 165, September 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301352.

Sultani:2017:AAA

[SS17b] Waqas Sultani and Mubarak Shah. Automatic action annotation in weakly labeled videos. Computer Vision and Image Understanding: CVIU, 161(??): 77–86, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300796.

Sultani:2021:HAR

[SS21] Waqas Sultani and Mubarak Shah. Human action recognition in drone videos using a few aerial training examples. Computer Vision and Image Understanding: CVIU, 206(??):Article 103186, May 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000308.

Smith:2006:IML

[SSdVL06] Paul Smith, Mubarak Shah, and Niels da Vitoria Lobo. Integrating multiple levels of zoom to enable activity analysis. Computer Vision and Image Understanding: CVIU, 103(1):33– 51, July 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Schops:2017:LSO

[SSHP17] Thomas Schöps, Torsten Sattler, Christian Häne, and Marc Pollefeys. Largescale outdoor 3D reconstruction on a mobile device. Computer Vision and Image Understanding: CVIU, 157(??):151– 166, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301412.

Spanhel:2020:LFA

[ŠSJ^{+}20] Jakub Špaňhel, Jakub Sochor, Roman Juránek, Petr Dobeš, Vojtěch Bartl, and Adam Herout. Learning feature aggregation in temporal domain for reidentification. Computer Vision and Image Understanding: CVIU, 192(??):Article 102883, March 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421830393X.

Saracchini:2012:RMS

[SSL^{+}12] Rafael F. V. Saracchini, Jorge Stolfi, Helena C. G. Leitão, Gary A. Atkinson, and Melvyn L. Smith. A robust multi-scale integration method to obtain the depth from gradient maps. Computer Vision and Image Understanding: CVIU, 116(8):

882–895, August 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212000537.

Shi:2006:HEL

[SSM06] J. Shi, A. Samal, and D. Marx. How effective are landmarks and their geometry for face recognition? Computer Vision and Image Understanding: CVIU, 102(2):117–133, May 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Socolinsky:2003:FRV

[SSN03] Diego A. Socolinsky, Andrea Selinger, and Joshua D. Neuheisel. Face recognition with visible and thermal infrared imagery. Computer Vision and Image Understanding: CVIU, 91(1–2): 72–114, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Soladie:2013:IRF

[SSS13] Catherine Soladié, Nicolas Stoiber, and Renaud Séguier. Invariant representation of facial expressions for blended expression recognition on unknown subjects. Computer Vision and Image Understanding: CVIU, 117(11):1598– 1609, November 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001355.

Sturm:2006:SIO

[SST06] Peter Sturm, Tomas Svoboda, and Seth Teller. Special issue: Omnidirectional vision and camera networks. Computer Vision and Image Understanding: CVIU, 103(3):155, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sato:1996:DPC

[ST96] Yoshinobu Sato and Shinichi Tamura. Detecting planar and curved symmetries of 3D shapes from a range image. Computer Vision and Image Understanding: CVIU, 64(1): 175–187, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0052/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0052/production/pdf.

Shubina:2010:VSO

[ST10] Ksenia Shubina and John K.

Tsotsos. Visual search for an object in a 3D environment using a mobile robot. Computer Vision and Image Understanding: CVIU, 114(5):535–547, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Song:2020:DCO

[ST20] Ge Song and Xiaoyang Tan. Deep code operation network for multi-label image retrieval. Computer Vision and Image Understanding: CVIU, 193(??):Article 102916, April 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220300114.

Stark:1995:FOR

[Sta95] L. Stark. Functionality in object recognition. Computer Vision and Image Understanding: CVIU, 62(2): 145–??, ???? 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Song:2014:ERB

[STC14] Fengyi Song, Xiaoyang Tan, and Songcan Chen. Exploiting relationship between attributes for improved face verification. Computer Vision and Image Understanding: CVIU, 122(??):143– 154, May 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214000368.

Sanchez:2016:LTK

[STC^{+}16] Carlos Sánchez, Pierluigi Taddei, Simone Ceriani, Erik Wolfart, and Vítor Sequeira. Localization and tracking in known large environments using portable real-time 3D sensors. Computer Vision and Image Understanding: CVIU, 149(??):197– 208, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002568.

Salti:2014:SUS

[STD14] Samuele Salti, Federico Tombari, and Luigi Di Stefano. SHOT: Unique signatures of histograms for surface and texture description. Computer Vision and Image Understanding: CVIU, 125(??):251– 264, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000988.

Stevens:2001:EIC

[Ste01] Mark R. Stevens. Evaluating 2D image comparison metrics for 3D scene interpretation. Computer Vision and Image Understanding: CVIU, 84(1):179– 197, October 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0940;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0940/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0940/ref.

Steger:2013:UEL

[Ste13]  Carsten Steger.  Unbi-

ased extraction of lines with parabolic and Gaussian profiles. Computer Vision and Image Understanding: CVIU, 117(2):97– 112, February 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421200118X.

Sadri:2018:RVD

[STHBH18] Alireza Sadri, Ruwan Tennakoon, Reza Hosseinnezhad, and Alireza Bab-Hadiashar. Robust visual data segmentation: Sampling from distribution of model parameters. Computer Vision and Image Understanding: CVIU, 174(??):82–94, September 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218301140.

Sebe:2008:SMC

[STLH08] Nicu Sebe, Qi Tian, Michael S.

Lew, and Thomas S. Huang. Similarity Matching in Computer Vision and Multimedia. Computer Vision and Image Understanding: CVIU, 110(3):309– 311, June 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sakurada:2017:TCM

[STO17] Ken Sakurada, Daiki Tetsuka, and Takayuki Okatani. Temporal city modeling using street level imagery. Computer Vision and Image Understanding: CVIU, 157 (??):55–71, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300280.

Savarese:2009:SIR

[STV09] Silvio Savarese, Tinne Tuytelaars, and Luc Van Gool. Special issue on 3D representation for object and scene recognition. Computer Vision and Image Understanding: CVIU, 113(12): 1181–1182, December 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Saha:2001:FCO

[SU01a] Punam K. Saha and Jayaram K. Udupa. Fuzzy connected object delineation: Axiomatic path strength definition and the case of multiple seeds. Computer Vision and Image Understanding: CVIU, 83(3): 275–295, September 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0927;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0927/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0927/ref.

Saha:2001:RFC

[SU01b] Punam K. Saha and Jayaram K. Udupa. Relative fuzzy connectedness among multiple objects: Theory, algorithms, and applications in image segmentation. Computer Vision and Image Understanding: CVIU, 82(1):42–56, April 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0902;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0902/pdf;   http:
//www.idealibrary.com/

links/doi/10.1006/cviu.
2000.0902/ref.

Saha:2000:SBF

[SUO00] Punam K. Saha, Jayaram K.

Udupa, and Dewey Odhner. Scale-based fuzzy connected image segmentation: Theory, algorithms, and validation. Computer Vision and Image Understanding: CVIU, 77(2): 145–174, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.1999.0813;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0813/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0813/ref.

Super:2002:FRI

[Sup02] Boaz J. Super. Fast retrieval of isolated visual shapes. Computer Vision and Image Understanding: CVIU, 85(1):1–21, January 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Stenger:2015:E

[SUS^{+}15] Bjorn Stenger, Norimichi Ukita, Yoichi Sato, Pascal Fua, and David Fleet. Editorial. Computer Vision and Image Understanding: CVIU, 141(??):94, December 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215002210.

Sobral:2014:CRB

[SV14] Andrews Sobral and Antoine Vacavant. A comprehensive review of background subtraction algorithms evaluated with synthetic and real videos. Computer Vision and Image Understanding: CVIU, 122 (??):4–21, May 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213002361.

Subramaniam:2022:CSI

[SVA^{+}22] Arulkumar Subramaniam, Jayesh Vaidya, Muhammed Abdul Majeed Ameen, Athira Nambiar, and Anurag Mittal. Co-segmentation inspired attention module for video-based computer vision tasks. Computer Vision and Image Understanding: CVIU, 223(??): ??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001138.

Saygili:2015:ASS

[SvdMH15] Gorkem Saygili, Laurens van der Maaten, and Emile A. Hendriks. Adaptive stereo similarity fusion using confidence measures. Computer Vision and Image Understanding: CVIU, 135 (??):95–108, June 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000399.

Speth:2021:UFR

[SVF^{+}21] Jeremy Speth, Nathan Vance, Patrick Flynn, Kevin Bowyer, and Adam Czajka. Unifying frame rate and temporal dilations for improved remote pulse detection. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000904.

Strezoski:2023:MMT

[SvNW23] Gjorgji Strezoski, Nanne van Noord, and Marcel Worring. MATTE: Multi-task multi-scale attention. Computer Vision and Image Understanding: CVIU, 228 (??):??, February 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000024.

Santos-Victor:1997:VBD

[SVS97] José Santos-Victor and Giulio Sandini. Visual behaviors for docking. Computer Vision and Image Understanding: CVIU, 67(3): 223–238, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0528/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0528/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0528/production/ref.

Sefidgar:2015:DKC

[SVSM15] Yasaman S. Sefidgar, Arash Vahdat, Stephen Se, and Greg Mori. Discriminative key-component models for interaction detection and recognition. Computer Vision and Image Understanding: CVIU, 135(??): 16–30, June 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215000466.

Shih:2004:FED

[SW04] Frank Y. Shih and Yi-Ta Wu. Fast Euclidean distance transformation in two scans using a 3 × 3 neighborhood. Computer Vision and Image Understanding: CVIU, 93(2):195–205, February 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shih:2005:DBM

[SW05] Frank Y. Shih and Yi-Ta Wu. Decomposition of binary morphological structuring elements based on genetic algorithms. Computer Vision and Image Understanding: CVIU, 99(2):291– 302, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Schnieders:2013:CLC

[SW13] Dirk Schnieders and Kwan-Yee K. Wong. Camera and light calibration from reflections on a sphere. Computer Vision and Image Understanding: CVIU, 117(10):1536–1547, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001082.

Shi:2017:LFM

[SW17] Jiacha Shi and Xuanyin Wang. A local feature with multiple line descriptors and its speeded-up matching algorithm. Computer Vision and Image Understanding: CVIU, 162(??):57–70, September 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217301479.

Saha:2002:FDT

[SWG02] Punam K. Saha, Felix W. Wehrli, and Bryon R. Gomberg. Fuzzy distance transform: Theory, algorithms, and applications. Computer Vision and Image Understanding: CVIU, 86(3):171–190, June 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sun:2022:EBM

[SWMM22] Wenzhang Sun, Lu Wang, Shaopeng Ma, and Qinwen Ma. Estimating 3D body mesh without SMPL annotations via alternating successive convex approximation. Computer Vision and Image Understanding: CVIU, 224(??): ??, November 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001175.

Stolpner:2011:SML

[SWS11] Svetlana Stolpner, Sue Whitesides, and Kaleem Siddiqi. Sampled medial loci for 3D shape representation. Computer Vision and Image Understanding: CVIU, 115(5):695–706, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Singh:2000:MMU

[SWYP00] Hanumant Singh, Louis Whitcomb, Dana Yoerger, and Oscar Pizarro. Microbathymetric mapping from underwater vehicles in the deep ocean. Computer Vision and Image Understanding: CVIU, 79(1): 143–161, July 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0850;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0850/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0850/ref.

Shi:2023:CPL

[SXY^{+}23] Mengchen Shi, Fei Xie, Jiquan Yang, Jing Zhao, Xixiang Liu, and Fan Wang. Cutout with patchloss augmentation for improving generative adversarial networks against instability. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001418.

Song:2024:BRP

[SXZZ24] Yukai Song, Guangxin Xu, Xiaoyan Zhang, and Zhijun Zhang. BiPR-RL: Portrait relighting via bi-directional consistent deep reinforcement learning. Computer Vision and Image Understanding: CVIU, 239(??): ??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223002692.

Sahillioglu:2010:CFS

[SY10] Y. Sahillioğlu and Y. Yemez. Coarse-to-fine surface reconstruction from silhouettes and range data using mesh deformation. Computer Vision and Image Understanding: CVIU, 114(3):334– 348, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Srestasathiern:2011:PSR

[SY11] P. Srestasathiern and A. Yilmaz. Planar shape representation and matching under projective transformation. Computer Vision and Image Understanding: CVIU, 115(11):1525–1535, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421100169X.

Shimoda:2020:WSS

[SY20] Wataru Shimoda and Keiji Yanai. Weakly supervised semantic segmentation using distinct class specific saliency maps. Computer Vision and Image Understanding: CVIU, 191(??): Article 102712, February 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314218301863.

Sun:2023:UVA

[SY23] Qiyue Sun and Yang Yang.

Unsupervised video anomaly detection based on multitimescale trajectory prediction. Computer Vision and Image Understanding: CVIU, 227(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422200193X.

Sakalli:1999:RBS

[SYF99] Mustafa Sakalli, Hong Yan, and Alan Fu. A regionbased scheme using RKLT and predictive classified vector quantization. Computer Vision and Image Understanding: CVIU, 75(3):269– 280, September 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0776/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0776/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0776/production/ref.

Shaked:1996:DSR

[SYK96] D. Shaked, O. Yaron, and N. Kiryati. Deriving stopping rules for the probabilistic Hough transform by sequential analysis. Computer Vision and Image Understanding: CVIU, 63(3): 512–526, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0038/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0038/production/pdf.

Song:2024:AST

[SYL^{+}24] Dan Song, Yuanxiang Yang, Wenhui Li, Zhuang Shao, Weizhi Nie, Xuanya Li, and An-An Liu. Adaptive semantic transfer network for unsupervised 2D imagebased 3D model retrieval. Computer Vision and Image Understanding: CVIU, 238 (??):??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002382.

Srivastava:2013:UOG

[SYPK13] Gaurav Srivastava, Josiah A.

Yoder, Johnny Park, and Avinash C. Kak. Using objective ground-truth labels created by multiple annotators for improved video classification: a comparative study. Computer Vision and Image Understanding: CVIU, 117(10):1384– 1399, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421300129X.

Shivakumara:2015:NGS

[SYZ^{+}15] Palaiahnakote Shivakumara, Zehuan Yuan, Danni Zhao, Tong Lu, and Chew Lim Tan. New Gradient-Spatial-Structural Features for video script identification. Computer Vision and Image Understanding: CVIU, 130(??):35–53, January 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001854.

Schaffalitzky:2003:ALM

[SZ03] F. Schaffalitzky and A. Zisserman. Automated location matching in movies. Computer Vision and Image Understanding: CVIU, 92 (2–3):236–264, November/ December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shih:2007:LOC

[SZ07] Frank Y. Shih and Kai Zhang. Locating object contours in complex background using improved snakes. Computer Vision and Image Understanding: CVIU, 105(2):93–98, February 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shakeri:2016:CSS

[SZ16] Moein Shakeri and Hong Zhang. COROLA: a sequential solution to moving object detection using lowrank approximation. Computer Vision and Image Understanding: CVIU, 146 (??):27–39, May 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216000540.

Sun:2021:ISI

[SZB^{+}21] Ziyi Sun, Yunfeng Zhang, Fangxun Bao, Kai Shao, Xinxin Liu, and Caiming Zhang. ICycleGAN: Single image dehazing based on iterative dehazing model and CycleGAN. Computer Vision and Image Understanding: CVIU, 203(??):Article 103133, February 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301521.

Sushko:2024:GNS

[SZGK24] Vadim Sushko, Dan Zhang, Juergen Gall, and Anna Khoreva. Generating novel scene compositions from single images and videos. Computer Vision and Image Understanding: CVIU, 239 (??):??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002680.

Shen:2023:LNL

[SZL^{+}23] Lili Shen, Bo Zhao, Qunxia Li, Chuhe Zhang, Xichun Sun, and Bo Peng. Local to non-local: Multi-scale progressive attention network for image restoration. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001054.

Sultani:2017:UAP

[SZS17] Waqas Sultani, Dong Zhang, and Mubarak Shah. Unsupervised action proposal ranking through proposal recombination. Computer Vision and Image Understanding: CVIU, 161(??): 42–50, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301133.

Shi:2021:DPI

[SZW^{+}21] Miaowen Shi, Fan Zhang, Suwei Wang, Caiming Zhang, and Xuemei Li. Detail preserving image denoising with patch-based structure similarity via sparse representation and SVD. Computer Vision and Image Understanding: CVIU, 206(??):Article 103173, May 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000175.

Tzevanidis:2011:ULB

[TA11] Konstantinos Tzevanidis and Antonis Argyros. Unsupervised learning of background modeling parameters in multicamera systems. Computer Vision and Image Understanding: CVIU, 115 (1):105–116, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tavakoli:2013:SSB

[TA13] Vahid Tavakoli and Amir A.

Amini. A survey of shaped-based registration and segmentation techniques for cardiac images. Computer Vision and Image Understanding: CVIU, 117(9):966–989, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000775.

Tachos:2017:MDD

[TABK17] Stavros Tachos, Konstantinos Avgerinakis, Alexia Briassouli, and Ioannis Kompatsiaris. Mining discriminative descriptors for goalbased activity detection. Computer Vision and Image Understanding: CVIU, 160 (??):73–86, July 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300607.

Tian:2021:TDD

[TAC21] Qing Tian, Tal Arbel, and James J. Clark. Task dependent deep LDA pruning of neural networks. Computer Vision and Image Understanding: CVIU, 203(??): Article 103154, February 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314220301739.

Tian:2023:GPP

[TAC23] Qing Tian, Tal Arbel, and James J. Clark. Grow– push–prune: Aligning deep discriminants for effective structural network compression. Computer Vision and Image Understanding: CVIU, 231(??): ??, June 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223000620.

Tamaki:2009:RIR

[TAK09] Toru Tamaki, Toshiyuki Amano, and Kazufumi Kaneda. Representing images of a rotating object with cyclic permutation for view-based pose estimation. Computer Vision and Image Understanding: CVIU, 113(12):1210–1221, December 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tripathi:2022:MCD

[TAK^{+}22] Pavani Tripathi, Yasmeena Akhter, Mahapara Khurshid, Aditya Lakra, Rohit Keshari, Mayank Vatsa, and Richa Singh. MTCD: Cataract detection via near infrared eye images. Computer Vision and Image Understanding: CVIU, 214 (??):??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001478.

Tanaka:1995:ABS

[Tan95] Hiromi T. Tanaka. Accuracybased sampling and reconstruction with adaptive meshes for parallel hierarchical triangulation. Computer Vision and Image Understanding: CVIU, 61(3): 335–350, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1027/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1027/production/pdf.

Tang:2011:ARF

[Tan11] Min Tang. Automatic registration and fast volume reconstruction from serial histology sections. Computer Vision and Image Understanding: CVIU, 115(8): 1112–1120, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314211000865.

Taylor:2000:RAO

[Tay00] Camillo J. Taylor. Reconstruction of articulated objects from point correspondences in a single uncalibrated image. Computer Vision and Image Understanding: CVIU, 80(3): 349–363, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0878;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0878/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0878/ref.

Tarel:1999:CFR

[TB99] Jean-Philippe Tarel and Nozha Boujemaa. A coarse to fine 3D registration method based on robust fuzzy clustering. Computer Vision and Image Understanding: CVIU, 73(1):14– 28, January 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0673/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.

0673/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0673/production/ref.

Torabi:2013:LBR

[TB13] Atousa Torabi and Guillaume-Alexandre Bilodeau. A LSSbased registration of stereo thermal-visible videos of multiple people using belief propagation. Computer Vision and Image Understanding: CVIU, 117(12):1736– 1747, December 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001264.

Tsaregorodtsev:2023:PSB

[TB23] Alexander Tsaregorodtsev and Vasileios Belagiannis. ParticleAugment: Sampling-based data augmentation. Computer Vision and Image Understanding: CVIU, 228(??): ??, February 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000139.

Tomei:2021:VAD

[TBC^{+}21] Matteo Tomei, Lorenzo Baraldi, Simone Calderara, Simone Bronzin, and Rita Cucchiara. Video action detection by learning graphbased spatio-temporal interactions. Computer Vision and Image Understanding: CVIU, 206(??):Article 103187, May 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422100031X.

Tolias:2015:RTC

[TBFJ15] Giorgos Tolias, Andrei Bursuc, Teddy Furon, and Hervé Jégou. Rotation and translation covariant match kernels for image retrieval. Computer Vision and Image Understanding: CVIU, 140(??):9–20, November 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215001356.

Tong:2024:LGP

[TBZ^{+}24] Jinghui Tong, Yaqiu Bi, Cong Zhang, Hongbo Bi, and Ye Yuan. Local to global purification strategy to realize collaborative camouflaged object detection. Computer Vision and Image Understanding: CVIU, 241(??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000134.

Tsechpenakis:2011:DPM

[TC11] Gavriil Tsechpenakis and Sotirios P. Chatzis. Deformable probability maps: Probabilistic shape and appearance-based object segmentation. Computer Vision and Image Understanding: CVIU, 115(8): 1157–1169, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314211000853.

Thacker:2008:PCC

[TCB^{+}08] Neil A. Thacker, Adrian F.

Clark, John L. Barron, J. Ross Beveridge, Patrick Courtney, William R. Crum, Visvanathan Ramesh, and Christine Clark. Performance characterization in computer vision: a guide to best practices. Computer Vision and Image Understanding: CVIU, 109(3): 305–334, March 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tian:2018:OSP

[TCM18] Qing Tian, Songcan Chen, and Tinghuai Ma. Ordinal space projection learning via neighbor classes representation. Computer Vision and Image Understanding: CVIU, 174(??):24–32, September 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218300791.

Thi:2012:ILA

[TCZ^{+}12] Tuan Hue Thi, Li Cheng, Jian Zhang, Li Wang, and Shinichi Satoh. Integrating local action elements for action analysis. Computer Vision and Image Understanding: CVIU, 116(3): 378–395, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211002062.

Tan:2004:DVC

[TD04] Robin Tan and James W. Davis. Differential video coding of face and gesture events in presentation videos. Computer Vision and Image Understanding: CVIU, 96(2):200–215, November 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Trinh:2019:IIV

[TD19] Dinh-Hoan Trinh and Christian Daul. On illuminationinvariant variational optical flow for weakly textured scenes. Computer Vision and Image Understanding: CVIU, 179(??): 1–18, February 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421830256X.

Tek:2010:PDI

[TDK10] F. Boray Tek, Andrew G.

Dempster, and İzzet Kale. Parasite detection and identification for automated thin blood film malaria diagnosis. Computer Vision and Image Understanding: CVIU, 114(1):21–32, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tran:2012:MPD

[TDT12] Cuong Tran, Anup Doshi, and Mohan Manubhai Trivedi. Modeling and prediction of driver behavior by foot gesture analysis. Computer Vision and Image Understanding: CVIU, 116(3): 435–445, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002086.

Timofte:2015:SSR

[TDV15] Radu Timofte, Vincent De Smet, and Luc Van Gool. Semantic super-resolution: When and where is it useful? Computer Vision and Image Understanding: CVIU, 142 (??):1–12, ???? 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002039.

Tisse:2007:ONS

[TDWH07] Christel-Loic Tisse, Hugh Durrant-Whyte, and R. Andrew Hicks. An optical navigation sensor for micro aerial vehicles. Computer Vision and Image Understanding: CVIU, 105(1):21– 29, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tursun:2020:MOS

[TDZ^{+}20] Osman Tursun, Simon Denman, Rui Zeng, Sabesan Sivapalan, Sridha Sridharan, and Clinton Fookes. MTRNet++: One-stage mask-based scene text eraser. Computer Vision and Image Understanding: CVIU, 201(??):Article 103066, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314220301077.

Thorstensen:2011:DMF

[TÉSK11] Nicolas Thorstensen, Patrick Étyngier, Florent Ségonne, and Renaud Keriven. Diffusion maps as a framework for shape modeling. Computer Vision and Image Understanding: CVIU, 115(4):520–530, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tirkaz:2015:IVA

[TESY15] Caglar Tirkaz, Jacob Eisenstein, T. Metin Sezgin, and Berrin Yanikoglu. Identifying visual attributes for object recognition from text and taxonomy. Computer Vision and Image Understanding: CVIU, 137(??): 12–23, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000405.

Taycher:2007:COF

[TFD07] Leonid Taycher, John W.

Fisher III, and Trevor Darrell. Combining object and feature dynamics in probabilistic tracking. Computer Vision and Image Understanding: CVIU, 108(3): 243–260, December 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Thomas:2009:SRR

[TFL^{+}09] Alexander Thomas, Vittorio Ferrari, Bastian Leibe, Tinne Tuytelaars, and Luc Van Gool. Shape-from-recognition: Recognition enables meta-data transfer. Computer Vision and Image Understanding: CVIU, 113(12):1222–1234, December 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tarbox:1995:IIV

[TG95a] G. H. Tarbox and S. N. Gottschlich. IVIS: An Integrated Volumetric Inspection System. Computer Vision and Image Understanding: CVIU, 61(3): 430–444, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1032/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1032/production/pdf.

Tarbox:1995:PCS

[TG95b] G. H. Tarbox and S. N. Gottschlich. Planning for complete sensor coverage in inspection. Computer Vision and Image Understanding: CVIU, 61(1):84– 111, January 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1007/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1007/production/pdf.

Thirion:1995:CDC

[TG95c] Jean-Philippe Thirion and Alexis Gourdon. Computing the differential characteristics of isointensity surfaces. Computer Vision and Image Understanding: CVIU, 61(2):190– 202, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1015/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1015/production/pdf.

Taati:2011:LSD

[TG11] Babak Taati and Michael Greenspan. Local shape descriptor selection for object recognition in range data. Computer Vision and Image Understanding: CVIU, 115(5):681–694, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Toldo:2015:HSM

[TGFF15] Roberto Toldo, Riccardo Gherardi, Michela Farenzena, and Andrea Fusiello. Hierarchical structure-and-motion recovery from uncalibrated images. Computer Vision and Image Understanding: CVIU, 140(??): 127–143, November 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314215001228.

Thomas:2017:CVS

[TGM^{+}17] Graham Thomas, Rikke Gade, Thomas B. Moeslund, Peter Carr, and Adrian Hilton. Computer vision for sports: Current applications and research topics. Computer Vision and Image Understanding: CVIU, 159(??): 3–18, June 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300711.

Tang:2023:FPM

[TGQ23] Wenming Tang, Yuanhao Gong, and Guoping Qiu. Feature preserving 3D mesh denoising with a dense local graph neural network. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000905.

Tsap:1998:ENF

[TGSH98] Leonid V. Tsap, Dmitry B. Goldgof, Sudeep Sarkar, and Wen-Chen Huang. Efficient nonlinear finite element modeling of nonrigid objects via optimization of mesh models. Computer Vision and Image Understanding: CVIU, 69(3):330– 350, March 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0663/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0663/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0663/production/ref;
http://www.idealibrary.
com/links/artid/cviu.1998.
0698/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0698/production/pdf.

Torsello:2004:SMS

[TH04] Andrea Torsello and Edwin R. Hancock. A skeletal measure of 2D shape similarity. Computer Vision and Image Understanding: CVIU, 95(1):1–29, July 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tang:2006:CSA

[TH06] W. K. Tang and Y. S. Hung.

A column-space approach to projective reconstruction. Computer Vision and Image Understanding: CVIU, 101 (3):166–176, March 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tang:2023:SNS

[THH^{+}23] Xikai Tang, Karim Habashy, Fangzheng Huang, Chao Li, and Dayan Ban. SCA-Net: Spatial and channel attention-based network for 3D point clouds. Computer Vision and Image Understanding: CVIU, 232 (??):??, July 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422300070X.

Tubic:2003:VAI

[THL03] Dragan Tubic, Patrick Hébert, and Denis Laurendeau. A volumetric approach for interactive 3D modeling. Computer Vision and Image Understanding: CVIU, 92(1):56–77, October 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Thomee:2013:SIV

[THL13] Bart Thomee, Mark Huiskes, and Michael S. Lew. Special issue on visual concept detection in the MIR-FLICKR/ImageCLEF benchmark. Computer Vision and Image Understanding: CVIU, 117(5):451– 452, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213000350.

Thonnat:2010:SII

[Tho10] Monique Thonnat. Special issue on intelligent vision systems. Computer Vision and Image Understanding: CVIU, 114(5):501– 502, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tsai:1998:OSP

[THT^{+}98] Horng-Ren Tsai, Shi-Jinn Horng, Shun-Shan Tsai, Shung-Shing Lee, Tzong-Wann Kao, and Chia-Ho Chen. Optimal speedup parallel image template matching algorithms on processor arrays with a reconfigurable bus system. Computer Vision and Image Understanding: CVIU, 71(3):393–412, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0638/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0638/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0638/production/ref.

Tan:2001:DSR

[TI01] Joo Kooi Tan and Seiji Ishikawa. Deformable shape recovery by factorization based on a spatiotemporal measurement matrix. Computer Vision and Image Understanding: CVIU, 82(2):101–109, May 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0904;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0904/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0904/ref.

Tek:1997:VSM

[TK97] Hüseyin Tek and Benjamin B. Kimia. Volumetric segmentation of medical images by three-dimensional bubbles. Computer Vision and Image Understanding: CVIU, 65(2):246– 258, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0579/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0579/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0579/production/ref.

Tolias:2014:TLS

[TKAK14] Giorgos Tolias, Yannis Kalantidis, Yannis Avrithis, and Stefanos Kollias. Towards large-scale geometry indexing by feature selection. Computer Vision and Image Understanding: CVIU, 120(??):31– 45, March 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213002336.

Tsiotsios:2016:MEP

[TKDN16] C. Tsiotsios, T. K. Kim, A. J. Davison, and S. G. Narasimhan. Model effectiveness prediction and system adaptation for photometric stereo in murky water. Computer Vision and Image Understanding: CVIU, 150(??):126– 138, September 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000746.

Tanaka:2024:FAR

[TKK24] Nariki Tanaka, Hiroshi Kera, and Kazuhiko Kawamoto. Fourier analysis on robustness of graph convolutional neural networks for skeleton-based action recognition. Computer Vision and Image Understanding: CVIU, 240(??):

??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314224000171.

Truc:2009:VEF

[TKL^{+}09] Phan T. H. Truc, Md. A. U. Khan, Young-Koo Lee, Sungyoung Lee, and Tae-Seong Kim. Vessel enhancement filter using directional filter bank. Computer Vision and Image Understanding: CVIU, 113(1): 101–112, January 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tapaswi:2021:LTS

[TKL21] Makarand Tapaswi, Vijay Kumar, and Ivan Laptev. Long term spatio-temporal modeling for action detection. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000862.

Timofte:2016:PPC

[TKV16] Radu Timofte, Junseok Kwon, and Luc Van Gool. PICASO: PIxel correspondences and SOft match selection for real-time tracking. Computer Vision and Image Understanding: CVIU, 153(??):151–162, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216000473.

Tung:2015:ISA

[TL15] Frederick Tung and James J.

Little. Improving scene attribute recognition using web-scale object detectors. Computer Vision and Image Understanding: CVIU, 138(??):86–91, September 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215001265.

Tung:2016:SPN

[TL16] Frederick Tung and James J.

Little. Scene parsing by nonparametric label transfer of content-adaptive windows. Computer Vision and Image Understanding: CVIU, 143(??):191– 200, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001848.

Tabernik:2015:ADP

[TLB^{+}15] Domen Tabernik, Ales Leonardis, Marko Boben, Danijel Skocaj, and Matej Kristan. Adding discriminative power to a generative hierarchical compositional model using histograms of compositions. Computer Vision and Image Understanding: CVIU, 138(??):102– 113, September 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000909.

Teng:2005:AOF

[TLCH05] Chin-Hung Teng, Shang-Hong Lai, Yung-Sheng Chen, and Wen-Hsing Hsu. Accurate optical flow computation under non-uniform brightness variations. Computer Vision and Image Understanding: CVIU, 97(3): 315–346, March 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Theoharatos:2006:CSO

[TLEF06] Christos Theoharatos, Nikolaos Laskaris, George Economou, and Spiros Fotopoulos. Combining self-organizing neural nets with multivariate statistics for efficient color image retrieval. Computer Vision and Image Understanding: CVIU, 102(3): 250–258, June 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tan:2023:DTS

[TLFM23] Thon Tan, Joanne Mun-Yee Lim, Ji Jinn Foo, and Ramachandran Muniandy. 3D detection transformer: Set prediction of objects using point clouds. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001881.

Tang:2022:TAS

[TLH22] Yiming Tang, Yufei Liu, and Hong Huang. Target-aware and spatial-spectral discriminant feature joint correlation filters for hyperspectral video object tracking. Computer Vision and Image Understanding: CVIU, 223 (??):??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001151.

Tsotsos:2005:AVM

[TLMT^{+}05] John K. Tsotsos, Yueju Liu, Julio C. Martinez-Trujillo, Marc Pomplun, Evgueni Simine, and Kunhao Zhou. Attending to visual motion. Computer Vision and Image Understanding: CVIU, 100(1–2):3–40, October/November 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tamaazousti:2017:VLI

[TLP^{+}17] Youssef Tamaazousti, Hervé Le Borgne, Adrian Popescu, Etienne Gadeski, Alexandru Ginsca, and Céline Hudelot. Vision-language integration using constrained local semantic features. Computer Vision and Image Understanding: CVIU, 163(??): 41–57, October 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301121.

Tong:2012:SSF

[TLWT12] Yan Tong, Xiaoming Liu, Frederick W. Wheeler, and Peter H. Tu. Semisupervised facial landmark annotation. Computer Vision and Image Understanding: CVIU, 116(8):922– 935, August 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212000616.

Tan:2016:DDS

[TLY^{+}16] Chaowei Tan, Kang Li, Zhennan Yan, Dong Yang, Shaoting Zhang, Hui Jing Yu, Klaus Engelke, Colin Miller, and Dimitris Metaxas. A detection-driven and sparsity-constrained deformable model for fascia lata labeling and thigh intermuscular adipose quantification. Computer Vision and Image Understanding: CVIU, 151(??):80–89, October 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216300029.

Tordoff:2004:IRD

[TM04] Ben Tordoff and David W.

Murray. The impact of radial distortion on the self-calibration of rotating cameras. Computer Vision and Image Understanding: CVIU, 96(1):17–34, October 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tordoff:2007:MRZ

[TM07] B. J. Tordoff and D. W.

Murray. A method of reactive zoom control from uncertainty in tracking. Computer Vision and Image Understanding: CVIU, 105 (2):131–144, February 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Torabi:2012:IIF

[TMB12] Atousa Torabi, Guillaume Massé, and Guillaume-Alexandre Bilodeau. An iterative integrated framework for thermal-visible image registration, sensor fusion, and people tracking for video surveillance applications. Computer Vision and Image Understanding: CVIU, 116(2):210–221, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002335.

Tillett:2000:EDF

[TML00] Robin Tillett, Nigel McFarlane, and Jeff Lines. Estimating dimensions of freeswimming fish using 3D point distribution models. Computer Vision and Image Understanding: CVIU, 79(1):123–141, July 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0847;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0847/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0847/ref.

Tavares:2016:CFE

[TMM16] Gonçalo Tavares, André Mourão, and João Magalhães. Crowdsourcing facial expressions for affectiveinteraction. Computer Vision and Image Understanding: CVIU, 147(??):102– 113, June 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216000461.

Tsechpenakis:2006:LBD

[TMN06] Gabriel Tsechpenakis, Dimitris Metaxas, and Carol Neidle. Learning-based dynamic coupling of discrete and continuous trackers. Computer Vision and Image Understanding: CVIU, 104 (2–3):140–156, November/ December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Takai:2009:DSA

[TMNM09] Takeshi Takai, Atsuto Maki, Koichiro Niinuma, and Takashi Matsuyama. Difference sphere: an approach to near light source estimation. Computer Vision and Image Understanding: CVIU, 113(9):966–978, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tao:2013:RNR

[TMQM13] Lili Tao, Stephen J. Mein, Wei Quan, and Bogdan J. Matuszewski. Recursive non-rigid structure from motion with online learned shape prior. Computer Vision and Image Understanding: CVIU, 117(10):1287– 1298, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000726.

Thong:2020:OCD

[TMS20] William Thong, Pascal Mettes, and Cees G. M. Snoek. Open cross-domain visual search. Computer Vision and Image Understanding: CVIU, 200(??):Article 103045, November 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220300928.

Thurley:2005:IVC

[TN05] Matthew J. Thurley and Kim C. Ng. Identifying, visualizing, and comparing regions in irregularly spaced 3D surface data. Computer Vision and Image Understanding: CVIU, 98(2):239– 270, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Todorovic:2007:ICS

[TN07] Sinisa Todorovic and Michael C. Nechyba. Interpretation of complex scenes using dynamic tree-structure Bayesian networks. Computer Vision and Image Understanding: CVIU, 106(1):

71–84, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Thurley:2008:ISE

[TN08] Matthew J. Thurley and Kim C. Ng. Identification and sizing of the entirely visible rocks from a 3D surface data segmentation of laboratory rock piles. Computer Vision and Image Understanding: CVIU, 111(2): 170–178, August 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tomenotti:2024:HPE

[TNO24] Federico Figari Tomenotti, Nicoletta Noceti, and Francesca Odone. Head pose estimation with uncertainty and an application to dyadic interaction detection. Computer Vision and Image Understanding: CVIU, 243 (??):??, June 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000808.

Thomas:1999:DNM

[TO99] J. Inigo Thomas and John Oliensis. Dealing with noise in multiframe structure from motion. Computer Vision and Image Understanding: CVIU, 76(2):109– 124, November 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0779/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0779/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0779/production/ref.

Traver:2005:SME

[TP05] V. Javier Traver and Filiberto Pla. Similarity motion estimation and active tracking through spatialdomain projections on logpolar images. Computer Vision and Image Understanding: CVIU, 97(2):209–241, February 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Teney:2014:MFD

[TP14] Damien Teney and Justus Piater. Multiview feature distributions for object detection and continuous pose estimation. Computer Vision and Image Understanding: CVIU, 125(??):265– 282, August 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421400099X.

Tao:2016:CSP

[TPD^{+}16] Lili Tao, Adeline Paiement, Dima Damen, Majid Mirmehdi, Sion Hannuna, Massimo Camplani, Tilo Burghardt, and Ian Craddock. A comparative study of pose representation and dynamics modelling for online motion quality assessment. Computer Vision and Image Understanding: CVIU, 148(??):136– 152, July 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421500260X.

Thermos:2020:DSL

[TPDP20] Spyridon Thermos, Georgios Th. Papadopoulos, Petros Daras, and Gerasimos Potamianos. Deep sensorimotor learning for RGB-D object recognition. Computer Vision and Image Understanding: CVIU, 190 (??):Article 102844, January 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219301432.

Theodoridis:2015:OMA

[TPNP15] Theodoris Theodoridis, Konstantinos Papachristou, Nikos Nikolaidis, and Ioannis Pitas. Object motion analysis description in stereo video content. Computer Vision and Image Understanding: CVIU, 141(??): 52–66, December 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215001551.

Trucco:2000:FTV

[TPR^{+}00] E. Trucco, Y. R. Petillot, I. Tena Ruiz, K. Plakas, and D. M. Lane. Feature tracking in video and sonar subsea sequences with applications. Computer Vision and Image Understanding: CVIU, 79(1): 92–122, July 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0846;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0846/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0846/ref.

Theologou:2015:COM

[TPT15] Panagiotis Theologou, Ioannis Pratikakis, and Theoharis Theoharis. A comprehensive overview of methodologies and performance evaluation frameworks in 3D mesh segmentation. Computer Vision and Image Understanding: CVIU, 135 (??):49–82, June 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000028.

Theologou:2017:PBO

[TPT17] Panagiotis Theologou, Ioannis Pratikakis, and Theoharis Theoharis. Partbased 3D object retrieval via multi-label optimization. Computer Vision and Image Understanding: CVIU, 160(??):148– 157, July 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300681.

Tresadern:2009:VSH

[TR09] Philip A. Tresadern and Ian D. Reid. Video synchronization from human motion using rank constraints. Computer Vision and Image Understanding: CVIU, 113 (8):891–906, August 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Toth:2013:SSP

[TRG^{+}13] Robert Toth, Justin Ribault, John Gentile, Dan Sperling, and Anant Madabhushi. Simultaneous segmentation of prostatic zones using Active Appearance Models with multiple coupled levelsets. Computer Vision and Image Understanding: CVIU, 117(9):1051– 1060, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000702.

Tim:2020:DEB

[TRPD20] Stefen Chan Wai Tim, Michele Rombaut, Denis Pellerin, and Anuvabh Dutt. Descriptor extraction based on a multilayer dictionary architecture for classification of natural images. Computer Vision and Image Understanding: CVIU, 191(??):Article 102708, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218301784.

Tang:2000:OSI

[TS00a] Xiaoou Tang and W. Kenneth Stewart. Optical and sonar image classification: Wavelet packet transform vs Fourier transform. Computer Vision and Image Understanding: CVIU, 79(1): 25–46, July 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0843;   http:

//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0843/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0843/ref.

Tari:2000:NLS

[TS00b] Sibel Tari and Jayant Shah.

Nested local symmetry set. Computer Vision and Image Understanding: CVIU, 79(2):267–280, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0856;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0856/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0856/ref.

Tissainayagam:2001:PPA

[TS01] P. Tissainayagam and D. Suter.

Performance prediction analysis of a point feature tracker based on different motion models. Computer Vision and Image Understanding: CVIU, 84 (1):104–125, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0939;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0939/pdf;   http:

                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0939/ref.

Thomas:2011:RRR

[TS11] Diego Thomas and Akihiro Sugimoto. Robustly registering range images using local distribution of albedo. Computer Vision and Image Understanding: CVIU, 115(5):649–667, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tichy:2016:NPB

[TS16] Ondrej Tichý and Václav Smídl. Non-parametric Bayesian models of response function in dynamic image sequences. Computer Vision and Image Understanding: CVIU, 151(??): 90–100, October 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002544.

Thomas:2017:MLS

[TS17] Diego Thomas and Akihiro Sugimoto. Modeling large-scale indoor scenes with rigid fragments using RGB-D cameras. Computer Vision and Image Understanding: CVIU, 157(??): 103–116, April 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301941.

Tonioni:2019:DIH

[TS19] Alessio Tonioni and Luigi Di Stefano. Domain invariant hierarchical embedding for grocery products recognition. Computer Vision and Image Understanding: CVIU, 182(??):81– 92, May 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219300414.

Takeda:2024:UDL

[TS24] Keita Takeda and Tomoya Sakai. Unsupervised deep learning of foreground objects from low-rank and sparse dataset. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000201.

Tsai:1996:PAG

[Tsa96] Frank C. D. Tsai. A probabilistic approach to geometric hashing using line features. Computer Vision and Image Understanding: CVIU, 63(1):182– 195, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0013/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0013/production/pdf.

Turchini:2017:ULA

[TSD17] Francesco Turchini, Lorenzo Seidenari, and Alberto Del Bimbo. Understanding and localizing activities from correspondences of clustered trajectories. Computer Vision and Image Understanding: CVIU, 159(??):128– 142, June 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216301965.

Tang:2014:EDL

[TSL14] Jun Tang, Ling Shao, and Xuelong Li. Efficient dictionary learning for visual categorization. Computer Vision and Image Understanding: CVIU, 124(??): 91–98, July 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214000332.

Tari:1997:ESS

[TSP97] Z. Sibel Göktepe Tari, Jayant Shah, and Homer Pien. Extraction of shape skeletons from grayscale images. Computer Vision and Image Understanding: CVIU, 66(2): 133–146, May 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0612/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0612/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0612/production/ref.

Tang:2014:SRR

[TST14] Yunqi Tang, Zhenan Sun, and Tieniu Tan. Slice representation of range data for head pose estimation. Computer Vision and Image Understanding: CVIU, 128(??):18–35, November 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001271.

Tiwari:2016:NSB

[TT16] Deepshikha Tiwari and Vipin Tyagi. A novel scheme based on local binary pattern for dynamic texture recognition. Computer Vision and Image Understanding: CVIU, 150(??):58–65, September 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216300352.

Tu:2007:FMT

[TTH07] Jilin Tu, Hai Tao, and Thomas Huang. Face as mouse through visual face tracking. Computer Vision and Image Understanding: CVIU, 108(1–2):35–40, October/November 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tateno:2017:LSL

[TTN17] Keisuke Tateno, Federico Tombari, and Nassir Navab. Large scale and long standing simultaneous reconstruction and segmentation. Computer Vision and Image Understanding: CVIU, 157(??):138– 150, April 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216300662.

Tai:2021:SRB

[TTXT21] Yuan Tai, Yihua Tan, Shengzhou Xiong, and Jinwen Tian. Subspace reconstruction based correlation filter for object tracking. Computer Vision and Image Understanding: CVIU, 212(??): ??, November 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001168.

Takala:1999:DTA

[TV99] Jarmo H. Takala and Jouko O. Viitanen. Distance transform algorithm for bitserial SIMD architectures. Computer Vision and Image Understanding: CVIU, 74 (2):150–161, May 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0756/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0756/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0756/production/ref;
http://www.idealibrary.
com/links/artid/cviu.1999.
0757/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0757/production/pdf.

Turaga:2009:UVR

[TVC09] Pavan Turaga, Ashok Veeraraghavan, and Rama Chellappa. Unsupervised view and rate invariant clustering of video sequences. Computer Vision and Image Understanding: CVIU, 113(3): 353–371, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tu:2016:EBC

[TVE^{+}16] Liyun Tu, Jared Vicory, Shireen Elhabian, Beatriz Paniagua, Juan Carlos Prieto, James N. Damon, Ross Whitaker, Martin Styner, and Stephen M. Pizer. Entropy-based correspondence improvement of interpolated skeletal models. Computer Vision and Image Understanding: CVIU, 151(??):72–79, October 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215002441.

Tahoces:2008:ICM

[TVLS08] Pablo G. Tahoces, J. Ramón Varela, María J. Lado, and Miguel Souto. Image compression: Maxshift ROI encoding options in JPEG2000. Computer Vision and Image Understanding: CVIU, 109(2):139–145, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Timofte:2018:GEV

[TVY^{+}18] Radu Timofte, Luc Van Gool, Ming-Hsuan Yang, Shai Avidan, Yasuyuki Matsushita, and Qingxiong Yang. Guest editorial:

Vision and computational photography and graphics. Computer Vision and Image Understanding: CVIU, 168 (??):1–2, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300274.

Takacs:1998:DMM

[TW98] Barnabás Takács and Harry Wechsler. A dynamic and multiresolution model of visual attention and its application to facial landmark detection. Computer Vision and Image Understanding: CVIU, 70(1): 63–73, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0619/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0619/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0619/production/ref.

Tang:2023:EDC

[TWQW23] Fulin Tang, Shaohuan Wu, Zhengda Qian, and Yihong Wu. Efficient 6-DoF camera pose tracking with circular edges. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001479.

Tabbone:2006:NSD

[TWS06] S. Tabbone, L. Wendling, and J.-P. Salmon. A new shape descriptor defined on the Radon transform. Computer Vision and Image Understanding: CVIU, 102(1): 42–51, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tribou:2014:SRM

[TWW14] Michael J. Tribou, Steven L. Waslander, and David W. L. Wang. Scale recovery in multicamera cluster SLAM with non-overlapping fields of view. Computer Vision and Image Understanding: CVIU, 126(??):53–66, September 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001283.

Tankus:2001:CBV

[TY01] Ariel Tankus and Yehezkel Yeshurun. Convexity-based visual camouflage breaking. Computer Vision and Image Understanding: CVIU, 82(3):208–237, June 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0912;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0912/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0912/ref.

Tankus:2005:SCD

[TY05] Ariel Tankus and Yehezkel Yeshurun. Scene-consistent detection of feature points in video sequences. Computer Vision and Image Understanding: CVIU, 97(1):1– 29, January 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tran:2022:VFI

[TY22] Quang Nhat Tran and Shih-Hsuan Yang. Video frame interpolation via down-up scale generative adversarial networks. Computer Vision and Image Understanding: CVIU, 220(??): ??, July 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000546.

Tian:2018:PSP

[TYDH18] Yonghong Tian, Jiaying Yan, Siwei Dong, and Tiejun Huang. PA-Search: Predicting units adaptive motion search for surveillance video coding. Computer Vision and Image Understanding: CVIU, 170 (??):14–27, May 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300353.

Tang:2021:NID

[TYH^{+}21] Qunfang Tang, Jie Yang, Xiangjian He, Wenjing Jia, Qingnian Zhang, and Haibo Liu. Nighttime image dehazing based on Retinex and dark channel prior using Taylor series expansion. Computer Vision and Image Understanding: CVIU, 202(??):Article 103086, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301168.

Torr:2000:MNR

[TZ00] P. H. S. Torr and A. Zisserman. MLESAC: a new robust estimator with application to estimating image geometry. Computer Vision and Image Understanding: CVIU, 78(1):138– 156, April 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/

cviu.1999.0832;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0832/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
1999.0832/ref.

Tao:2022:ACN

[TZL^{+}22] Jianwei Tao, Xiankun Zhang, Xuexiong Luo, Yuan Wang, Chen Song, and Yue Sun. Adaptive capsule network. Computer Vision and Image Understanding: CVIU, 218(??):??, April 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000327.

Tan:2021:CSP

[TZLT21] Hui Li Tan, Hongyuan Zhu, Joo-Hwee Lim, and Cheston Tan. A comprehensive survey of procedural video datasets. Computer Vision and Image Understanding: CVIU, 202(??):Article 103107, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301314.

Torr:1998:RDD

[TZM98] P. H. S. Torr, A. Zisserman, and S. J. Maybank. Robust detection of degenerate configurations while estimating the fundamental matrix. Computer Vision and Image Understanding: CVIU, 71(3):312– 333, September 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0559/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0559/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0559/production/ref.

Tu:2008:SMR

[TZY08] Zhuowen Tu, Songfeng Zheng, and Alan Yuille. Shape matching and registration by data-driven EM. Computer Vision and Image Understanding: CVIU, 109 (3):290–304, March 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Unsalan:2005:SDH

[ÜB05] Cem Ünsalan and Kim L. Boyer. A system to detect houses and residential street networks in multispectral satellite images. Computer Vision and Image Understanding: CVIU, 98(3):423– 461, June 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Uz:2009:MBT

[UBEP09] Tamer Uz, George Bebis, Ali Erol, and Salil Prabhakar. Minutiae-based template synthesis and matching for fingerprint authentication. Computer Vision and Image Understanding: CVIU, 113(9):979–992, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Unsalan:2001:CBP

[ÜE01] Cem Ünsalan and Aytül Erçil. Conversions between parametric and implicit forms using polar/spherical coordinate representations. Computer Vision and Image Understanding: CVIU, 81(1):1–25, January 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0881;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0881/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0881/ref.

Urtasun:2006:TMM

[UFF06] Raquel Urtasun, David J.

Fleet, and Pascal Fua. Temporal motion models for monocular and multiview 3D human body tracking. Computer Vision and Image Understanding: CVIU, 104 (2–3):157–177, November/ December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

UlHaq:2020:DMD

[UFK20] Israr Ul Haq, Keisuke Fujii, and Yoshinobu Kawahara. Dynamic mode decomposition via dictionary learning for foreground modeling in videos. Computer Vision and Image Understanding: CVIU, 199(??):Article 103022, October 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220300813.

UlHaq:2022:DMD

[UIK22] Israr Ul Haq, Tomoharu Iwata, and Yoshinobu Kawahara. Dynamic mode decomposition via convolutional autoencoders for dynamics modeling in videos. Computer Vision and Image Understanding: CVIU, 216 (??):??, February 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422100182X.

Uddin:2022:SFF

[UJ22] S. M. Nadim Uddin and Yong Ju Jung. SIFNet: Free-form image inpainting using color split-inpaint-fuse approach. Computer Vision and Image Understanding: CVIU, 221(??): ??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000613.

Ukita:2012:GPM

[UK12a] Norimichi Ukita and Takeo Kanade. Gaussian process motion graph models for smooth transitions among multiple actions. Computer Vision and Image Understanding: CVIU, 116(4): 500–509, April 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211002608.

Ukita:2012:RCR

[UK12b] Norimichi Ukita and Takeo Kanade. Reference consistent reconstruction of 3D cloth surface. Computer Vision and Image Understanding: CVIU, 116(8): 869–881, August 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000628.

Ukita:2005:RTC

[UM05] Norimichi Ukita and Takashi Matsuyama. Real-time cooperative multi-target tracking by communicating active vision agents. Computer Vision and Image Understanding: CVIU, 97(2): 137–179, February 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ukita:2016:PRI

[UMH16] Norimichi Ukita, Yusuke Moriguchi, and Norihiro Hagita. People re-identification across non-overlapping cameras using group features. Computer Vision and Image Understanding: CVIU, 144 (??):228–236, March 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S107731421500140X.

Ukita:2016:HOF

[UO16] Norimichi Ukita and Asami Okada. High-order framewise smoothness-constrained globally-optimal tracking. Computer Vision and Image Understanding: CVIU, 153(??):130–142, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314216300650.

Ulges:2010:LAC

[USKB10] Adrian Ulges, Christian Schulze, Markus Koch, and Thomas M. Breuel. Learning automatic concept detectors from online video. Computer Vision and Image Understanding: CVIU, 114(4):429–438, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Urfalioglu:2011:AEA

[UTB^{+}11] Onay Urfalioglu, Thorsten Thormählen, Hellward Broszio, Patrick Mikulastik, and A. Enis Cetin. Algebraic error analysis of collinear feature points for camera parameter estimation. Computer Vision and Image Understanding: CVIU, 115(4): 467–475, April 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ukita:2018:SWS

[UU18] Norimichi Ukita and Yusuke Uematsu. Semi- and weaklysupervised human pose estimation. Computer Vision and Image Understanding: CVIU, 170(??):67– 78, May 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218300158.

Urban:2017:MFO

[UWH17] Steffen Urban, Martin Weinmann, and Stefan Hinz. mdBRIEF — a fast online-adaptable, distorted binary descriptor for realtime applications using calibrated wide-angle or fisheye cameras. Computer Vision and Image Understanding: CVIU, 162(??): 71–86, September 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301480.

Udomkesmalee:1997:IPP

[UZC97] Suraphol Udomkesmalee, David Q. Zhu, and Cheng-Chih Chu. Image processing for planetary limb/ terminator extraction. Computer Vision and Image Understanding: CVIU, 67(3): 274–284, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0543/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0543/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0543/production/ref.

Vemulapalli:2016:RFR

[VAC16] Raviteja Vemulapalli, Felipe Arrate, and Rama Chellappa. R3DG features: Relative 3D geometrybased skeletal representations for human action recognition. Computer Vision and Image Understanding: CVIU, 152(??):155– 166, November 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300297.

Valentine:2010:ECC

[VAWW10] Brian Valentine, Senyo Apewokin, Linda Wills, and Scott Wills. An efficient, chromatic clustering-based background model for embedded vision platforms. Computer Vision and Image Understanding: CVIU, 114(11):1152–1163, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Vaidya:1998:DPS

[VB98] Nitin M. Vaidya and Kim L.

Boyer. Discontinuitypreserving surface reconstruction using stochastic differential equations. Computer Vision and Image Understanding: CVIU, 72(3): 257–270, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0700/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0700/production/pdf;
http://www.idealibrary.

                                        com/links/artid/cviu.1998.
                                        0700/production/ref.

Vo:2016:SIN

[VB16] Nam N. Vo and Aaron F. Bobick. Sequential Interval Network for parsing complex structured activity. Computer Vision and Image Understanding: CVIU, 143(??):147– 158, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001599.

V:2019:SSF

[VBA19] Arun P. V., Krishna Mohan B., and Porwal A. Spatialspectral feature based approach towards convolutional sparse coding of hyperspectral images. Computer Vision and Image Understanding: CVIU, 188 (??):Article 102797, November 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314219301146.

Vanrell:2004:IOC

[VBS^{+}04] M. Vanrell, R. Baldrich, A. Salvatella, R. Benavente, and F. Tous. Induction operators for a computational colour–texture representation. Computer Vision and Image Understanding: CVIU, 94(1–3):92–114, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Vasileiadis:2019:MPP

[VBT19] Manolis Vasileiadis, Christos-Savvas Bouganis, and Dimitrios Tzovaras. Multiperson 3D pose estimation from 3D cloud data using 3D convolutional neural networks. Computer Vision and Image Understanding: CVIU, 185(??): 12–23, August 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1077314219300669.

Valle:2019:FAU

[VBVB19] Roberto Valle, José M. Buenaposada, Antonio Valdés, and Luis Baumela. Face alignment using a 3D deeply-initialized ensemble of regression trees. Computer Vision and Image Understanding: CVIU, 189 (??):Article 102846, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314219301444.

Vaca-Castano:2017:ISI

[VCDS^{+}17] Gonzalo Vaca-Castano, Samarjit Das, Joao P. Sousa, Niels D. Lobo, and Mubarak Shah. Improved scene identification and object detection on egocentric vision of daily activities. Computer Vision and Image Understanding: CVIU, 156(??): 92–103, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216301692.

Vaca-Castano:2019:HOD

[VCLS19] Gonzalo Vaca-Castano, Niels DaVitoria Lobo, and Mubarak Shah. Holistic object detection and image understanding. Computer Vision and Image Understanding: CVIU, 181(??): 1–13, April 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219300219.

Venetianer:2010:PEI

[VD10] Péter L. Venetianer and Hongli Deng. Performance evaluation of an intelligent video surveillance system — a case study. Computer Vision and Image Understanding: CVIU, 114(11): 1292–1302, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Verri:1997:TMV

[Ver97] Alessandro Verri. There is more to vision than geometry — reply to Pizlo, Rosenfeld, and Weiss. Computer Vision and Image Understanding: CVIU, 65(3): 439–441, March 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0495/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0495/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0495/production/ref. See

[PRW97a].

Vieville:1996:FOE

[VF96] T. Viéville and O. D.

Faugeras. The first order expansion of motion equations in the uncalibrated case. Computer Vision and Image Understanding: CVIU, 64 (1):128–146, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0049/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0049/production/pdf.

Vo:2017:HNW

[VGLP17] Phong D. Vo, Alexandru Ginsca, Hervé Le Borgne, and Adrian Popescu. Harnessing noisy Web images for deep representation. Computer Vision and Image Understanding: CVIU, 164(??):68–81, November 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217300255.

Versteegen:2016:TMN

[VGR16] Ralph Versteegen, Georgy Gimel’farb, and Patricia Riddle. Texture modelling with nested highorder Markov–Gibbs random fields. Computer Vision and Image Understanding: CVIU, 143(??):120– 134, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002453.

Villamizar:2016:IMO

[VGSMN16] Michael Villamizar, Anaís Garrell, Alberto Sanfeliu, and Francesc Moreno-Noguer. Interactive multiple object learning with scanty human supervision. Computer Vision and Image Understanding: CVIU, 149(??):51– 64, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300042.

vanGemert:2010:CCC

[vGSV^{+}10] Jan C. van Gemert, Cees G. M. Snoek, Cor J. Veenman, Arnold W. M. Smeulders, and Jan-Mark Geusebroek. Comparing compact codebooks for visual categorization. Computer Vision and Image Understanding: CVIU, 114(4):450– 462, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Verma:2017:SVA

[VJ17] Yashaswi Verma and C. V.

Jawahar. A support vector approach for crossmodal search of images and texts. Computer Vision and Image Understanding: CVIU, 154(??):48–63, January 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301540.

Vakhitov:2018:SNL

[VKL18] Alexander Vakhitov, Andrey Kuzmin, and Victor Lempitsky. Set2Model networks: Learning discriminatively to learn generative models. Computer Vision and Image Understanding: CVIU, 173(??): 13–23, August 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301388.

Vrigkas:2014:MMC

[VKNK14] Michalis Vrigkas, Vasileios Karavasilis, Christophoros Nikou, and Ioannis A. Kakadiaris. Matching mixtures of curves for human action recognition. Computer Vision and Image Understanding: CVIU, 119(??): 27–40, February 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213002270.

Vijayakumar:1998:IBR

[VKP98] B. Vijayakumar, David Kriegman, and Jean Ponce. Invariant-based recognition of complex curved 3D objects from image contours. Computer Vision and Image Understanding: CVIU, 72(3):287–303, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0701/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0701/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0701/production/ref.

Vogler:2001:FRS

[VM01] Christian Vogler and Dimitris Metaxas. A framework for recognizing the simultaneous aspects of American Sign Language. Computer Vision and Image Understanding: CVIU, 81 (3):358–384, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0895;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0895/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0895/ref.

Vascon:2016:DCG

[VMC^{+}16] Sebastiano Vascon, Eyasu Z.

Mequanint, Marco Cristani, Hayley Hung, Marcello Pelillo, and Vittorio Murino. Detecting conversational groups in images and sequences: a robust gametheoretic approach. Computer Vision and Image Understanding: CVIU, 143 (??):11–24, February 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314215002076.

Vojir:2016:OAH

[VMN16] Tomas Vojir, Jiri Matas, and Jana Noskova. Online adaptive hidden Markov model for multi-tracker fusion. Computer Vision and Image Understanding: CVIU, 153(??):109–119, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216300509.

Vandenbroucke:2003:CIS

[VMP03] Nicolas Vandenbroucke, Ludovic Macaire, and Jack-Gérard Postaire. Color image segmentation by pixel classification in an adapted hybrid color space. application to soccer image analysis. Computer Vision and Image Understanding: CVIU, 90(2):190–216, May 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

VanGool:1995:CDS

[VMUO95] Luc Van Gool, Theo Moons, Dorin Ungureanu, and Andre Oosterlinck. The characterization and detection of skewed symmetry. Computer Vision and Image Understanding: CVIU, 61(1): 138–150, January 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1010/production;   http:

//www.idealibrary.com/
links/artid/cviu.1995.
1010/production/pdf.

Varga:2014:LGU

[VNNB14] László G. Varga, László G.

Nyúl, Antal Nagy, and Péter Balázs. Local and global uncertainty in binary tomographic reconstruction. Computer Vision and Image Understanding: CVIU, 129(??):52–62, December 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214001179.

Vandersmissen:2024:CQE

[VO24] Benjamin Vandersmissen and José Oramas. On the coherency of quantitative evaluation of visual explanations. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000158.

Valem:2023:GCN

[VPL23] Lucas Pascotti Valem, Daniel Carlos Guimarães Pedronette, and Longin Jan Latecki. Graph convolutional networks based on manifold learning for semi-supervised image classification. Computer Vision and Image Understanding: CVIU, 227 (??):??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001965.

Vecchio:2023:MGT

[VPP^{+}23] Giuseppe Vecchio, Luca Prezzavento, Carmelo Pino, Francesco Rundo, Simone Palazzo, and Concetto Spampinato. MeT: a graph transformer for semantic segmentation of 3D meshes. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001534.

Vacavant:2013:CMS

[VRKL13] Antoine Vacavant, Tristan Roussillon, Bertrand Kerautret, and Jacques-Olivier Lachaud. A combined multiscale/irregular algorithm for the vectorization of noisy digital contours. Computer Vision and Image Understanding: CVIU, 117(4): 438–450, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001841.

Veeraraghavan:2006:RTD

[VSP06] Harini Veeraraghavan, Paul Schrater, and Nikos Papanikolopoulos. Robust target detection and tracking through integration of motion, color, and geometry. Computer Vision and Image Understanding: CVIU, 103 (2):121–138, August 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Vacavant:2014:SSB

[VTRC14] Antoine Vacavant, Laure Tougne, Lionel Robinault, and Thierry Chateau. Special section on background models comparison. Computer Vision and Image Understanding: CVIU, 122 (??):1–3, May 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000587.

VandenWyngaerd:2002:ACP

[VV02] Joris Vanden Wyngaerd and Luc Van Gool. Automatic crude patch registration: Toward automatic 3D model building. Computer Vision and Image Understanding: CVIU, 87(1–3): 8–26, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Varadarajan:2015:FCR

[VWMZ15] Sriram Varadarajan, Hongbin Wang, Paul Miller, and Huiyu Zhou. Fast convergence of regularised regionbased mixture of Gaussians for dynamic background modelling. Computer Vision and Image Understanding: CVIU, 136 (??):45–58, July 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002422.

Vincze:2009:IVS

[VZP^{+}09] Markus Vincze, Michael Zillich, Wolfgang Ponweiser, Vaclav Hlavac, Jiri Matas, Stepan Obdrzalek, Hilary Buxton, Jonathan Howell, Kingsley Sage, Antonis Argyros, Christoph Eberst, and Gerald Umgeher. Integrated vision system for the semantic interpretation of activities where a person handles objects. Computer Vision and Image Understanding: CVIU, 113(6): 682–692, June 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wildes:2000:REF

[WALL00] Richard P. Wildes, Michael J. Amabile, Ann-Marie Lanzillotto, and Tzong-Shyng Leu. Recovering estimates of fluid flow from image sequence data. Computer Vision and Image Understanding: CVIU, 80(2): 246–266, November 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0874;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0874/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0874/ref.

Wang:2017:SRG

[WAPB17] Fan Wang, Samia Ainouz, Caroline Petitjean, and Abdelaziz Bensrhair. Specularity removal: a global energy minimization approach based on polarization imaging. Computer Vision and Image Understanding: CVIU, 158(??): 31–39, May 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217300474.

Wang:2014:TYE

[WASF14] Xinchao Wang, Vitaly Ablavsky, Horesh Ben Shitrit, and Pascal Fua. Take your eyes off the ball: Improving ball-tracking by focusing on team play. Computer Vision and Image Understanding: CVIU, 119(??):102–115, February 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314213002300.

Wang:1997:DCE

[WB97] Yang Wang and Prabir Bhattacharya. Digital connectivity and extended wellcomposed sets for gray images. Computer Vision and Image Understanding: CVIU, 68(3):330–345, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0551/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0551/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0551/production/ref.

Williams:2001:SRM

[WB01] John Williams and Mohammed Bennamoun. Simultaneous registration of multiple corresponding point sets. Computer Vision and Image Understanding: CVIU, 81(1):117– 142, January 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2000.0884;   http:

//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0884/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0884/ref.

Wu:2011:MRF

[WB11] Dijia Wu and Kim L. Boyer.

Markov random field based phase demodulation of interferometric images. Computer Vision and Image Understanding: CVIU, 115(6): 759–770, June 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2012:AGS

[WB12] Quan Wang and Kim L.

Boyer. The active geometric shape model: a new robust deformable shape model and its applications. Computer Vision and Image Understanding: CVIU, 116(12):1178–1194, December 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001154.

Whitelam:2015:AEL

[WB15] Cameron Whitelam and Thirimachos Bourlai. Accurate eye localization in the Short Waved Infrared Spectrum through summation range filters. Computer Vision and Image Understanding: CVIU, 139(??): 59–72, October 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215000946.

Wu:2016:GOC

[WB16] Zheng Wu and Margrit Betke. Global optimization for coupled detection and data association in multiple object tracking. Computer Vision and Image Understanding: CVIU, 143(??): 25–37, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002167.

Williams:2014:PES

[WBS14] Ian Williams, Nicholas Bowring, and David Svoboda. A performance evaluation of statistical tests for edge detection in textured images. Computer Vision and Image Understanding: CVIU, 122(??): 115–130, May 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214000356.

Wang:1999:PMR

[WC99] Ranxiao Frances Wang and James E. Cutting. A probabilistic model for recovering camera translation. Computer Vision and Image Understanding: CVIU, 76(3):205–212, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0798/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0798/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0798/production/ref.

Wu:2010:CCG

[WCF10] Lin Wu, Xiaochun Cao, and Hassan Foroosh. Camera calibration and geo-location estimation from two shadow trajectories. Computer Vision and Image Understanding: CVIU, 114(8):915–927, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wilson:1998:SMA

[WCH98] Richard C. Wilson, Andrew D. J. Cross, and Edwin R. Hancock. Structural matching with active triangulations. Computer Vision and Image Understanding: CVIU, 72(1):21– 38, October 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0656/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1997.
                                        0656/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0656/production/ref.

Wong:2013:SSF

[WCYS13] Hoi Sim Wong, Tat-Jun Chin, Jin Yu, and David Suter. A simultaneous sample-and-filter strategy for robust multi-structure model fitting. Computer Vision and Image Understanding: CVIU, 117(12):1755– 1769, December 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213001719.

Wang:2002:ITM

[WCZ02] Qing Wang, Zheru Chi, and Rongchun Zhao. Image thresholding by maximizing the index of nonfuzziness of the 2-D grayscale histogram. Computer Vision and Image Understanding: CVIU, 85(2):100–116, February 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2007:VIS

[WCZ^{+}07] Qi Wang, Xilin Chen, Liang-Guo Zhang, Chunli Wang, and Wen Gao. Viewpoint invariant sign language recognition. Computer Vision and Image Understanding: CVIU, 108(1–2):87–97, October/ November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2020:TDC

[WCZ^{+}20] Bisheng Wang, Guo Cao, Licun Zhou, Youqiang Zhang, and Yanfeng Shang. Task differentiation: Constructing robust branches for precise object detection. Computer Vision and Image Understanding: CVIU, 199 (??):Article 103030, October 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220300849.

Wan:2023:VCA

[WCZ23] Xiaoyue Wan, Zhuo Chen, and Xu Zhao. View consistency aware holistic triangulation for 3D human pose estimation. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002102.

Wang:1996:RTD

[WD96] Wendong Wang and James H. Duncan. Recovering the three-dimensional motion and structure of multiple moving objects from binocular image flows. Computer Vision and Image Understanding: CVIU, 63(3): 430–446, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0033/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0033/production/pdf.

Wagner:2014:TTA

[WD14] Hubert Wagner and Pawel Dlotko. Towards topological analysis of high-dimensional feature spaces. Computer Vision and Image Understanding: CVIU, 121(??): 21–26, April 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214000125.

Widynski:2012:FSC

[WDB12] Nicolas Widynski, Séverine Dubuisson, and Isabelle Bloch. Fuzzy spatial constraints and ranked partitioned sampling approach for multiple object tracking. Computer Vision and Image Understanding:

CVIU, 116(10):1076–1094, October 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001038.

Wen:2020:UDN

[WDC^{+}20] Longyin Wen, Dawei Du, Zhaowei Cai, Zhen Lei, Ming-Ching Chang, Honggang Qi, Jongwoo Lim, Ming-Hsuan Yang, and Siwei Lyu. UA-DETRAC: a new benchmark and protocol for multi-object detection and tracking. Computer Vision and Image Understanding: CVIU, 193(??): Article 102907, April 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220300035.

Wang:2024:MNM

[WDC^{+}24] Hang Wang, Zhenyu Ding, Cheng Cheng, Yuhai Li, and Hongbin Sun. MDC-Net: Multi-domain constrained kernel estimation network for blind image super resolution. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422300245X.

Wang:2012:DMB

[WDN^{+}12] Tao Wang, Guojun Dai, Bingbing Ni, De Xu, and François Siewe. A distance measure between labeled combinatorial maps. Computer Vision and Image Understanding: CVIU, 116(12):1168–1177, December 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314212001130.

Wanderley:2002:SFP

[WF02] Juliana F. Camapum Wanderley and Mark H. Fisher. Spatial-feature parametric clustering applied to motion-based segmentation in camouflage. Computer Vision and Image Understanding: CVIU, 85(2):144– 157, February 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Woodard:2005:FSB

[WF05] Damon L. Woodard and Patrick J. Flynn. Finger surface as a biometric identifier. Computer Vision and Image Understanding: CVIU, 100(3):357–384, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2023:OOT

[WG23] Hongmei Wang and Fan Guo. Online object tracking based interactive attention. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001893.

Wang:2014:ROD

[WGAD14] Yifei Wang, Yuan Gao, Alin Achim, and Naim Dahnoun. Robust obstacle detection based on a novel disparity calculation method and G-disparity. Computer Vision and Image Understanding: CVIU, 123(??): 23–40, June 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421400040X.

Wang:2024:EVA

[WGGH24] Zhiqiang Wang, Xiaojing Gu, Xingsheng Gu, and Jingyu Hu. Enhancing video anomaly detection with learnable memory network: a new approach to memorybased auto-encoders. Computer Vision and Image Understanding: CVIU, 241 (??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000274.

Wang:2021:CBD

[WGGHvdW21] Yaxing Wang, Abel Gonzalez-Garcia, Luis Herranz, and Joost van de Weijer. Controlling biases and diversity in diverse image-to-image translation. Computer Vision and Image Understanding: CVIU, 202(??):Article 103082, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220301156.

Wen:2020:RND

[WGZL20] Zhijie Wen, Jiawei Guan, Tieyong Zeng, and Ying Li. Residual network with detail perception loss for single image super-resolution. Computer Vision and Image Understanding: CVIU, 199(??):Article 103007, October 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314220300734.

Williams:1996:PCO

[WH96] Lance R. Williams and Allen R. Hanson. Perceptual completion of occluded surfaces. Computer Vision and Image Understanding: CVIU, 64 (1):1–20, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0043/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0043/production/pdf.

Wilson:2000:BVA

[WH00] Richard C. Wilson and Edwin R. Hancock. Biasvariance analysis for controlling adaptive surface meshes. Computer Vision and Image Understanding: CVIU, 77(1):25– 47, January 2000. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0810/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0810/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0810/production/ref.

Wang:2001:STM

[WH01] Xiaoguang Wang and Allen R.

Hanson. Surface texture and microstructure extraction from multiple aerial images. Computer Vision and Image Understanding: CVIU, 83(1):1–37, July 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2001.0916;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0916/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0916/ref.

Wang:2018:BBS

[WH18] Boyu Wang and Minh Hoai. Back to the beginning: Starting point detection for early recognition of ongoing human actions. Computer Vision and Image Understanding: CVIU, 175(??): 24–31, October 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218303886.

Wang:2014:TFI

[WHC14] Tinghuai Wang, Bo Han, and John Collomosse. Touch-Cut: Fast image and video segmentation using singletouch interaction. Computer Vision and Image Understanding: CVIU, 120(??): 14–30, March 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421300204X.

Wu:2020:RMM

[WHGZ20] Min Wu, Congying Han, Tiande Guo, and Tong Zhao. Registration and matching method for directed point set with orientation attributes and local information. Computer Vision and Image Understanding: CVIU, 191(??):Article 102866, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219301572.

Wang:2023:OFE

[WHJK23] Jian Wang, Ziwei Han, Wenjing Jiang, and Junseok Kim. 3D object feature extraction and classification using 3D MF-DFA. Computer Vision and Image Understanding: CVIU, 232 (??):??, July 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000681.

Wu:2014:RDI

[WHL14] Yihong Wu, Zhanyi Hu, and Youfu Li. Radial distortion invariants and lens evaluation under a singleoptical-axis omnidirectional camera. Computer Vision and Image Understanding: CVIU, 126(??):11–27, September 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001039.

Wang:2020:TSD

[WHL^{+}20] Shuhui Wang, Ling Hu, Liang Li, Weigang Zhang, and Qingming Huang. Twostream deep sparse network for accurate and efficient image restoration. Computer Vision and Image Understanding: CVIU, 200(??):Article 103029, November 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314220300850.

Wang:2021:RTA

[WHL^{+}21] Xinggang Wang, Zhaojin Huang, Bencheng Liao, Lichao Huang, Yongchao Gong, and Chang Huang. Real-time and accurate object detection in compressed video by long short-term feature aggregation. Computer Vision and Image Understanding: CVIU, 206(??):Article 103188, May 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000321.

Wang:2009:SSK

[WHM^{+}09] Meng Wang, Xian-Sheng Hua, Tao Mei, Richang Hong, Guojun Qi, Yan Song, and Li-Rong Dai. Semi-supervised kernel density estimation for video annotation. Computer Vision and Image Understanding: CVIU, 113(3):384– 396, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wu:2008:MNR

[WHN08] Xiao Wu, Alexander G.

Hauptmann, and Chong-Wah Ngo. Measuring novelty and redundancy with multiple modalities in crosslingual broadcast news. Computer Vision and Image Understanding: CVIU, 110(3):418–431, June 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wu:2024:CCP

[WHS^{+}24] Pingfan Wu, Hengyi Huang, Han Sun, Dong Liang, and Ningzhong Liu. CPRNC: Channels pruning via reverse neuron crowding for model compression. Computer Vision and Image Understanding: CVIU, 240 (??):??, March 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000237.

Weng:2023:FUC

[WHY^{+}23] Jianfeng Weng, Kun Hu, Tingting Yao, Jingya Wang, and Zhiyong Wang. Federated Unsupervised Cluster-Contrastive learning for person re-identification: a coarse-to-fine approach. Computer Vision and Image Understanding: CVIU, 237 (??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314223002114.

Wang:2007:MVF

[WJ07] Peng Wang and Qiang Ji. Multi-view face and eye detection using discriminant features. Computer Vision and Image Understanding: CVIU, 105(2):99–111, February 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wheeler:2021:SDS

[WK21] Bradley J. Wheeler and Hassan A. Karimi. A semantically driven selfsupervised algorithm for detecting anomalies in image sets. Computer Vision and Image Understanding: CVIU, 213(??):??, December 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314221001235.

Wang:2016:ILG

[WKI^{+}16] Chaohui Wang, Nikos Komodakis, Hiroshi Ishikawa, Olga Veksler, and Endre Boros. Inference and learning of graphical models: Theory and applications in computer vision and image analysis. Computer Vision and Image Understanding: CVIU, 143(??): 52–53, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000023.

Wang:2013:MRF

[WKP13] Chaohui Wang, Nikos Komodakis, and Nikos Paragios. Markov Random Field modeling, inference & learning in computer vision & image understanding: a survey. Computer Vision and Image Understanding: CVIU, 117(11):1610–1627, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001343.

Wu:2022:SAN

[WKT22] Shuang Wu, Mohan Kankanhalli, and Anthony K. H. Tung. Superclass-aware network for few-shot learning.

Computer Vision and Image Understanding: CVIU, 216 (??):??, February 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001788.

Weyand:2015:VLR

[WL15] Tobias Weyand and Bastian Leibe. Visual landmark recognition from Internet photo collections: a large-scale evaluation. Computer Vision and Image Understanding: CVIU, 135 (??):1–15, June 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000363.

Wagner:1999:RME

[WLD99] Robert Wagner, Feiyu Liu, and Klaus Donner. Robust motion estimation for calibrated cameras from monocular image sequences. Computer Vision and Image Understanding: CVIU, 73(2):258– 268, February 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0739/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0739/production/pdf;

http://www.idealibrary.
com/links/artid/cviu.1998.
0739/production/ref.

Wu:2021:OCA

[WLFL21] Jhih-Ciang Wu, Sherman Lu, Chiou-Shann Fuh, and Tyng-Luh Liu. One-class anomaly detection via novelty normalization. Computer Vision and Image Understanding: CVIU, 210 (??):??, September 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314221000709.

Wong:2008:MBA

[WLI08] Sam T. S. Wong, Howard Leung, and Horace H. S. Ip. Model-based analysis of Chinese calligraphy images. Computer Vision and Image Understanding: CVIU, 109(1):69–85, January 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2022:SST

[WLL^{+}22a] Haidong Wang, Zhiyong Li, Yaping Li, Ke Nai, and Ming Wen. STURE: Spatial-temporal mutual representation learning for robust data association in online multi-object tracking. Computer Vision and Image Understanding: CVIU, 220(??):??, July 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222000558.

Wu:2022:AGW

[WLL22b] Ran Wu, Huanyu Liu, and Jun-Bao Li. Adaptive gradients and weight projection based on quantized neural networks for efficient image classification. Computer Vision and Image Understanding: CVIU, 223(??): ??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001059.

Wolf:2014:EVA

[WLM^{+}14] Christian Wolf, Eric Lombardi, Julien Mille, Oya Celiktutan, Mingyuan Jiu, Emre Dogan, Gonen Eren, Moez Baccouche, Emmanuel Dellandréa, Charles-Edmond Bichot, Christophe Garcia, and Bülent Sankur. Evaluation of video activity localizations integrating quality and quantity measurements. Computer Vision and Image Understanding: CVIU, 127(??): 14–30, October 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001416.

Wu:2008:TSO

[WLMG08] Yiming Wu, Xiuwen Liu, Washington Mio, and K. A. Gallivan. Two-stage optimal component analysis. Computer Vision and Image Understanding: CVIU, 110(1): 91–101, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2018:RDB

[WLO^{+}18] Pichao Wang, Wanqing Li, Philip Ogunbona, Jun Wan, and Sergio Escalera. RGB-D-based human motion recognition with deep learning: a survey. Computer Vision and Image Understanding: CVIU, 171(??): 118–139, June 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300663.

Wang:2023:UDA

[WLSO23] Zhijie Wang, Xing Liu, Masanori Suganuma, and Takayuki Okatani. Unsupervised domain adaptation for semantic segmentation via cross-region alignment. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001236.

Wang:2016:RBN

[WLW^{+}16] Yi Wang, Bin Li, Yang Wang, Fang Chen, Bang Zhang, and Zhidong Li. Robust Bayesian non-parametric dictionary learning with heterogeneous Gaussian noise. Computer Vision and Image Understanding: CVIU, 150(??):31–43, September 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216300765.

Wang:2014:SFB

[WLX^{+}14] Xiaofeng Wang, Yong Liu, Bingchao Xu, Lu Li, and Jianru Xue. A statistical feature based approach to distinguish PRCG from photographs. Computer Vision and Image Understanding: CVIU, 128(??): 84–93, November 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421400160X.

Wang:2024:VFW

[WLXL24] Zhihao Wang, Lin Li, Zhongwei Xie, and Chuanbo Liu. Video frame-wise explanation driven contrastive learning for procedural text generation. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000353.

Wang:2023:LLI

[WLZ23] Manli Wang, Jiayue Li, and Changsen Zhang. Lowlight image enhancement by deep learning network for improved illumination map. Computer Vision and Image Understanding: CVIU, 232(??):??, July 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000619.

Wei:2020:SDA

[WLZM20] Haiyang Wei, Zhixin Li, Canlong Zhang, and Huifang Ma. The synergy of double attention: Combine sentence-level and wordlevel attention for image captioning. Computer Vision and Image Understanding: CVIU, 201 (??):Article 103068, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220301089.

Wang:2004:DCS

[WLZW04] Ze Wang, Chi-Sing Leung, Yi-Sheng Zhu, and Tien-Tsin Wong. Data compression with spherical wavelets and wavelets for the imagebased relighting. Computer Vision and Image Understanding: CVIU, 96(3):327– 344, December 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ward:2020:SLB

[WM20] Daniel Ward and Peyman Moghadam. Scalable learning for bridging the species gap in imagebased plant phenotyping. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 103009, August 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314220300746.

Wiliem:2012:SBD

[WMBY12] Arnold Wiliem, Vamsi Madasu, Wageeh Boles, and Prasad Yarlagadda. A suspicious behaviour detection using a context space model for smart surveillance systems. Computer Vision and Image Understanding: CVIU, 116(2):194–209, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314211002001.

Wang:2021:CFL

[WML21] Yiyang Wang, Long Ma, and Risheng Liu. A convergent framework with learnable feasibility for Hadamard-based image recovery. Computer Vision and Image Understanding: CVIU, 202(??):Article 103095, January 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220301259.

Wu:2023:CGS

[WMZY23] Jiahao Wu, Bo Ma, Yuping Zhang, and Xin Yi. Capturing geometric structure change through deformation aware correlation. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001649.

Wachter:1999:TPM

[WN99] S. Wachter and H.-H. Nagel.

Tracking persons in monocular image sequences. Computer Vision and Image Understanding: CVIU, 74(3): 174–192, June 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0758/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1999.
                                        0758/production/pdf.

Wong:2005:FAD

[WNH05] Andrew K. C. Wong, Peiyi Niu, and Xiang He. Fast acquisition of dense depth data by a new structured light scheme. Computer Vision and Image Understanding: CVIU, 98(3):398– 422, June 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2010:GED

[WO10] Hongzhi Wang and John Oliensis. Generalizing edge detection to contour detection for image segmentation. Computer Vision and Image Understanding: CVIU, 114(7):731–744, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Worthington:2005:RDS

[Wor05] Philip L. Worthington. Reillumination-driven shape from shading. Computer Vision and Image Understanding: CVIU, 98(2):325– 343, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wuhrer:2014:EHB

[WPB^{+}14] Stefanie Wuhrer, Leonid Pishchulin, Alan Brunton, Chang Shu, and Jochen Lang. Estimation of human body shape and posture under clothing. Computer Vision and Image Understanding: CVIU, 127(??): 31–42, October 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214001398.

Wang:2016:MPC

[WPI^{+}16] Bo Wang, Marcel Prastawa, Andrei Irimia, Avishek Saha, Wei Liu, S. Y. Matthew Goh, Paul M. Vespa, John D. Van Horn, and Guido Gerig. Modeling 4D pathological changes by leveraging normative models. Computer Vision and Image Understanding: CVIU, 151(??):3–13, October 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216000205.

Withey:2009:DET

[WPK09] D. J. Withey, W. Pedrycz, and Z. J. Koles. Dynamic edge tracing: Boundary identification in medical images. Computer Vision and Image Understanding: CVIU, 113(10):1039–1052, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2020:CMH

[WPQ20] Jiaze Wang, Xiaojiang Peng, and Yu Qiao. Cascade multi-head attention networks for action recognition. Computer Vision and Image Understanding: CVIU, 192(??):Article 102898, March 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421930178X.

Wang:2003:BSM

[WPS03] Yongmei Wang, Bradley S. Peterson, and Lawrence H. Staib. 3D brain surface matching based on geodesics and local geometry. Computer Vision and Image Understanding: CVIU, 89 (2–3):252–271, February/ March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2018:BID

[WPSL18] Hongyan Wang, Jinshan Pan, Zhixun Su, and Songxin Liang. Blind image deblurring using elastic-net based rank prior. Computer Vision and Image Understanding: CVIU, 168(??): 157–171, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217302060.

Wang:2016:SII

[WPZ^{+}16] Liang Wang, Ioannis Patras, Jian Zhang, Greg Mori, and Larry Davis. Special issue on individual and group activities in video event analysis. Computer Vision and Image Understanding: CVIU, 144(??): 1–2, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421600028X.

Wu:2018:OTO

[WPZ^{+}18] Feng Wu, Shaowu Peng, Jingkai Zhou, Qiong Liu, and Xiaojia Xie. Object tracking via Online Multiple Instance Learning with reliable components. Computer Vision and Image Understanding: CVIU, 172(??): 25–36, July 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421830050X.

Wang:2021:IEV

[WQY^{+}21] Yasi Wang, Yuankai Qi, Hongxun Yao, Dong Gong, and Qi Wu. Image editing with varying intensities of processing. Computer Vision and Image Understanding: CVIU, 211(??): ??, October 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221001041.

Worz:2008:PBE

[WR08] Stefan Wörz and Karl Rohr. Physics-based elastic registration using non-radial basis functions and including landmark localization uncertainties. Computer Vision and Image Understanding: CVIU, 111(3):263–274, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Weinland:2006:FVA

[WRB06] Daniel Weinland, Remi Ronfard, and Edmond Boyer. Free viewpoint action recognition using motion history volumes. Computer Vision and Image Understanding: CVIU, 104 (2–3):249–257, November/ December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Weinland:2011:SVB

[WRB11] Daniel Weinland, Remi Ronfard, and Edmond Boyer. A survey of visionbased methods for action representation, segmentation and recognition. Computer Vision and Image Understanding: CVIU, 115 (2):224–241, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Williams:1997:IDM

[WRH97] Christopher K. I. Williams, Michael Revow, and Geoffrey E. Hinton. Instantiating deformable models with a neural net. Computer Vision and Image Understanding: CVIU, 68(1):120– 126, October 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0540/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0540/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0540/production/ref.

Walther:2005:SVA

[WRKP05] Dirk Walther, Ueli Rutishauser, Christof Koch, and Pietro Perona. Selective visual attention enables learning and recognition of multiple objects in cluttered scenes. Computer Vision and Image Understanding: CVIU, 100(1–2):41–63, October/ November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wong:2006:TBM

[WS06] King Yuen Wong and Minas E. Spetsakis. Tracking based motion segmentation under relaxed statistical assumptions. Computer Vision and Image Understanding: CVIU, 101(1):45– 64, January 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2008:VLR

[WS08] Liang Wang and David Suter. Visual learning and recognition of sequential data manifolds with applications to human movement analysis. Computer Vision and Image Understanding: CVIU, 110(2):153– 172, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wientapper:2018:UCF

[WSFTK18] Folker Wientapper, Michael Schmitt, Matthieu Fraissinet-Tachet, and Arjan Kuijper. A universal, closedform approach for absolute pose problems. Computer Vision and Image Understanding: CVIU, 173(??): 57–75, August 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300572.

Workman:2015:SSE

[WSJ15] Scott Workman, Richard Souvenir, and Nathan Jacobs. Scene shape estimation from multiple partly cloudy days. Computer Vision and Image Understanding: CVIU, 134(??): 116–129, May 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002069.

Woo:2013:MSU

[WSKH13] Jonghye Woo, Piotr J.

Slomka, C.-C. Jay Kuo, and Byung-Woo Hong. Multiphase segmentation using an implicit dual shape prior: Application to detection of left ventricle in cardiac MRI. Computer Vision and Image Understanding: CVIU, 117(9):1084– 1094, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000684.

Worring:1996:PFB

[WSSD96] Marcel Worring, Arnold W. M. Smeulders, Lawrence H. Staib, and James S. Duncan. Parameterized feasible boundaries in gradient vector fields. Computer Vision and Image Understanding: CVIU, 63(1):135– 144, January 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0009/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0009/production/pdf.

Wang:2013:ASI

[WSSS13] Zhimin Wang, Qing Song, Yeng Chai Soh, and Kang Sim. An adaptive spatial information-theoretic fuzzy clustering algorithm for image segmentation. Computer Vision and Image Understanding: CVIU, 117(10):1412–1420, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314213000957.

Wang:2005:EEG

[WSV05] Jian-Gang Wang, Eric Sung, and Ronda Venkateswarlu. Estimating the eye gaze from one eye. Computer Vision and Image Understanding: CVIU, 98(1):83– 103, April 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2016:CEA

[WSY^{+}16] Peng Wang, Lifeng Sun, Shiqiang Yang, Alan F. Smeaton, and Cathal Gurrin. Characterizing everyday activities from visual lifelogs based on enhancing concept representation. Computer Vision and Image Understanding: CVIU, 148(??):181– 192, July 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421500209X.

Werghi:2015:LBP

[WTBdB15] Naoufel Werghi, Claudio Tortorici, Stefano Berretti, and Alberto del Bimbo. Local binary patterns on triangular meshes: Concept and applications. Computer Vision and Image Understanding: CVIU, 139(??): 161–177, October 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000843.

Wu:2017:VQA

[WTW^{+}17] Qi Wu, Damien Teney, Peng Wang, Chunhua Shen, Anthony Dick, and Anton van den Hengel. Visual question answering: a survey of methods and datasets. Computer Vision and Image Understanding: CVIU, 163(??):21–40, October 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217300772.

Wu:2018:DPE

[WTYC18] Po-Chen Wu, Hung-Yu Tseng, Ming-Hsuan Yang, and Shao-Yi Chien. Direct pose estimation for planar objects. Computer Vision and Image Understanding: CVIU, 172(??): 50–66, July 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218300419.

Wang:2021:DHP

[WTZ^{+}21] Jinbao Wang, Shujie Tan, Xiantong Zhen, Shuo Xu, Feng Zheng, Zhenyu He, and Ling Shao. Deep 3D human pose estimation: a review. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000692.

Wang:2024:GGF

[WTZ24] Xuan Wang, Hao Tang, and Zhigang Zhu. GMC: a general framework of multistage context learning and utilization for visual detection tasks. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000250.

Williams:1997:APV

[WW97] James Williams and Lawrence Wolff. Analysis of the pulmonary vascular tree using differential geometry based vector fields. Computer Vision and Image Understanding: CVIU, 65(2):226– 236, February 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0571/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0571/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0571/production/ref.

Wang:2016:POD

[WW16] Huiling Wang and Tinghuai Wang. Primary object discovery and segmentation in videos via graphbased transductive inference. Computer Vision and Image Understanding: CVIU, 143(??):159– 172, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002507.

Wang:2015:RNR

[WWCZ15] Gang Wang, Zhicheng Wang, Yufei Chen, and Weidong Zhao. A robust non-rigid point set registration method based on asymmetric Gaussian representation. Computer Vision and Image Understanding: CVIU, 141(??):67–80, December 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215001289.

Wu:2018:SDH

[WWG^{+}18] Lin Wu, Yang Wang, Zongyuan Ge, Qichang Hu, and Xue Li. Structured deep hashing with convolutional neural networks for fast person reidentification. Computer Vision and Image Understanding: CVIU, 167(??): 63–73, February 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217302011.

Wachsmuth:2007:CIV

[WWH07] Sven Wachsmuth, Sebastian Wrede, and Marc Hanheide. Coordinating interactive vision behaviors for cognitive assistance. Computer Vision and Image Understanding: CVIU, 108(1–2):135–149, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wu:2013:ADF

[WWJ13a] Yuwei Wu, Yuanquan Wang, and Yunde Jia. Adaptive diffusion flow active contours for image segmentation. Computer Vision and Image Understanding: CVIU, 117(10):1421–1435, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001033.

Wu:2013:SLV

[WWJ13b] Yuwei Wu, Yuanquan Wang, and Yunde Jia. Segmentation of the left ventricle in cardiac cine MRI using a shape-constrained snake model. Computer Vision and Image Understanding: CVIU, 117(9):990– 1003, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000787.

Wang:2016:CGS

[WWJ16] Shangfei Wang, Chongliang Wu, and Qiang Ji. Capturing global spatial patterns for distinguishing posed and spontaneous expressions. Computer Vision and Image Understanding: CVIU, 147 (??):69–76, June 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001824.

Weise:2011:OLC

[WWLV11] Thibaut Weise, Thomas Wismer, Bastian Leibe, and Luc Van Gool. Online loop closure for real-time interactive 3D scanning. Computer Vision and Image Understanding: CVIU, 115(5): 635–648, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wu:1995:RAF

[WWW95] Zhongquan Wu, Lide Wu, and Aicheng Wu. The robust algorithms for finding the center of an arc. Computer Vision and Image Understanding: CVIU, 62(3):269–278, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1995.
                                        1054/production;   http:

//www.idealibrary.com/
links/artid/cviu.1995.
1054/production/pdf.

Wu:2023:FPA

[WWWF23] Jiaxin Wu, Xiaopeng Wang, Tongyi Wei, and Chao Fang. Full-parameter adaptive fuzzy clustering for noise image segmentation based on non-local and local spatial information. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001455.

Wu:2024:SAM

[WWXK24] Cong Wu, Xiao-Jun Wu, Tianyang Xu, and Josef Kittler. Scene adaptive mechanism for action recognition. Computer Vision and Image Understanding: CVIU, 238 (??):??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002345.

Wang:2016:STT

[WX16] Jing Wang and Zhijie Xu. Spatio-temporal texture modelling for realtime crowd anomaly detection. Computer Vision and Image Understanding: CVIU, 144(??):177– 187, March 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421500185X.

Wang:2020:NAS

[WXC20] Ping Wang, Guili Xu, and Yuehua Cheng. A novel algebraic solution to the perspective-three-line pose problem. Computer Vision and Image Understanding: CVIU, 191(??):Article 102711, February 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218301929.

Wang:2018:ESP

[WXWC18] Ping Wang, Guili Xu, Zhengsheng Wang, and Yuehua Cheng. An efficient solution to the perspective-three-point pose problem. Computer Vision and Image Understanding: CVIU, 166(??):81–87, January 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314217301698.

Wu:2024:SGM

[WXZ24] Guangli Wu, Tongjie Xu, and Jing Zhang. Sparse graph matching network for temporal language localization in videos. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002886.

Wei:2018:SHR

[WXZG18] Xing Wei, Xiaobin Xu, Jiawei Zhang, and Yihong Gong. Specular highlight reduction with known surface geometry. Computer Vision and Image Understanding: CVIU, 168(??): 132–144, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421730173X.

Wang:2007:STM

[WY07] Jun Wang and Lijun Yin.

Static topographic modeling for facial expression recognition and analysis. Computer Vision and Image Understanding: CVIU, 108(1–2):19–34, October/ November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2021:TTM

[WY21] Yanze Wang and Junyong Ye. TMF: Temporal motion and fusion for action recognition. Computer Vision and Image Understanding: CVIU, 213(??): ??, December 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422100148X.

Wang:2015:MCA

[WYC15] Junyan Wang, Sai-Kit Yeung, and Kap Luk Chan. Matching-constrained active contours with affineinvariant shape prior. Computer Vision and Image Understanding: CVIU, 132 (??):39–55, March 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314214002227.

Wu:2022:AJH

[WYW^{+}22] Chunlei Wu, Ziyu Yuan, Shaohua Wan, Leiquan Wang, and Weishan Zhang. Anti-jamming heart rate estimation using a spatialtemporal fusion network. Computer Vision and Image Understanding: CVIU, 216 (??):??, February 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001673.

Wu:2016:MLT

[WYX^{+}16] Shiqian Wu, Lingxian Yang, Wangming Xu, Jinghong Zheng, Zhengguo Li, and Zhijun Fang. A mutual local-ternary-pattern based method for aligning differently exposed images. Computer Vision and Image Understanding: CVIU, 152(??):67–78, November 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301072.

Wang:2004:DIO

[WZ04] Yongmei Michelle Wang and Hongjiang Zhang. Detecting image orientation based on low-level visual content. Computer Vision and Image Understanding: CVIU, 93(3):328–346, March 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wan:2008:SVU

[WZ08] Dingrui Wan and Jie Zhou.

Stereo vision using two PTZ cameras. Computer Vision and Image Understanding: CVIU, 112(2):184–194, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2023:CUC

[WZ23] Xuan Wang and Zhigang Zhu. Context understanding in computer vision: a survey. Computer Vision and Image Understanding: CVIU, 229(??): ??, March 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223000267.

Wang:2021:EIQ

[WZC^{+}21] Zhizhong Wang, Lei Zhao, Haibo Chen, Zhiwen Zuo, Ailin Li, Wei Xing, and Dongming Lu. Evaluate and improve the quality of neural style transfer. Computer Vision and Image Understanding: CVIU, 207 (??):??, June 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221000473.

Wu:2024:JLF

[WZCG24] Qin Wu, Pengcheng Zhu, Zhilei Chai, and Guodong Guo. Joint learning of foreground, background and edge for salient object detection. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223002953.

Wu:2022:LWS

[WZCY22] Wen Wu, Kai Zhou, Xiao-Diao Chen, and Jun-Hai Yong. Light-weight shadow detection via GCNbased annotation strategy and knowledge distillation. Computer Vision and Image Understanding: CVIU, 216 (??):??, February 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001752.

Wang:2024:PRN

[WZF^{+}24] Yonghui Wang, Wengang Zhou, Hao Feng, Li Li, and Houqiang Li. Progressive Recurrent Network for shadow removal. Computer Vision and Image Understanding: CVIU, 238(??): ??, January 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002412.

Wang:2023:CTS

[WZFL23] Hao Wang, Libo Zhang, Heng Fan, and Tiejian Luo. Collaborative three-stream transformers for video captioning. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001790.

Wang:2024:GKK

[WZG^{+}24] Junhuang Wang, Weiwei Zhang, Yufeng Guo, Peng Liang, Ming Ji, Chenghui Zhen, and Hanmeng Wang. Global key knowledge distillation framework. Computer Vision and Image Understanding: CVIU, 239(??): ??, February 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223002825.

Wang:2021:VDS

[WZJ^{+}21] Tao Wang, Xiaoqin Zhang, Runhua Jiang, Li Zhao, Huiling Chen, and Wenhan Luo. Video deblurring via spatiotemporal pyramid network and adversarial gradient prior. Computer Vision and Image Understanding: CVIU, 203(??):Article 103135, February 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220301545.

Wang:2023:CDF

[WZQ^{+}23] Xiang Wang, Shiwei Zhang, Zhiwu Qing, Yiliang Lv, Changxin Gao, and Nong Sang. Cross-domain fewshot action recognition with unlabeled videos. Computer Vision and Image Understanding: CVIU, 233(??):

??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001170.

Wang:2024:MVC

[WZQ^{+}24] Shaowei Wang, Lingling Zhang, Tao Qin, Jun Liu, Yifei Li, Qianying Wang, and Qinghua Zheng. Multiview cognition with path search for one-shot part labeling. Computer Vision and Image Understanding: CVIU, 244(??): ??, July 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000961.

Wang:2013:MTP

[WZT13] Tao Wang, Zhigang Zhu, and Clark N. Taylor. A multimodal temporal panorama approach for moving vehicle detection, reconstruction and classification. Computer Vision and Image Understanding: CVIU, 117(12):1724–1735, December 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213001240.

Wang:2017:OCS

[WZW17] Shiping Wang, Handuo Zhang, and Han Wang. Object co-segmentation via weakly supervised data fusion. Computer Vision and Image Understanding: CVIU, 155(??):43–54, February 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301825.

Wu:2016:TML

[WZWH16] Fuchao Wu, Ming Zhang, Guanghui Wang, and Zhanyi Hu. Triangulation and metric of lines based on geometric error. Computer Vision and Image Understanding: CVIU, 145(??): 111–127, April 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000448.

Wang:2024:OTM

[WZWL24] Yingquan Wang, Pingping Zhang, Dong Wang, and Huchuan Lu. Other tokens matter: Exploring global and local features of Vision Transformers for Object Re-Identification. Computer Vision and Image Understanding: CVIU, 244 (??):??, July 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224001115.

Williams:1999:CSC

[WZWT99] Lance Williams, John Zweck, Tairan Wang, and Karvel Thornber. Computing stochastic completion fields in linear-time using a resolution pyramid. Computer Vision and Image Understanding: CVIU, 76(3):289– 297, December 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0800/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0800/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0800/production/ref.

Wang:2014:ITR

[WZX^{+}14] Jingdong Wang, Jiazhen Zhou, Hao Xu, Tao Mei, Xian-Sheng Hua, and Shipeng Li. Image tag refinement by regularized latent Dirichlet allocation. Computer Vision and Image Understanding: CVIU, 124(??): 61–70, July 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421400037X.

Wang:2024:TPO

[WZX^{+}24] Pei Wang, Yu Zhu, Danna Xue, Qingsen Yan, Jinqiu Sun, Sung eui Yoon, and Yanning Zhang. Take a prior from other tasks for severe blur removal. Computer Vision and Image Understanding: CVIU, 245 (??):??, August 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224001085.

Wang:2013:MSD

[WZY13] Qi Wang, Guokang Zhu, and Yuan Yuan. Multispectral dataset and its application in saliency detection. Computer Vision and Image Understanding: CVIU, 117(12):1748–1754, December 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S107731421300132X.

Wang:2014:SQS

[WZY14] Qi Wang, Guokang Zhu, and Yuan Yuan. Statistical quantization for similarity search. Computer Vision and Image Understanding: CVIU, 124(??): 22–30, July 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000538.

Xu:2007:OSU

[XAB07] Ning Xu, Narendra Ahuja, and Ravi Bansal. Object segmentation using graph cuts based active contours. Computer Vision and Image Understanding: CVIU, 107(3):210–224, September 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xia:2016:SME

[XFP^{+}16] Zhaoqiang Xia, Xiaoyi Feng, Jinye Peng, Xianlin Peng, and Guoying Zhao. Spontaneous micro-expression spotting via geometric deformation modeling. Computer Vision and Image Understanding: CVIU, 147 (??):87–94, June 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002702.

Xie:2013:RMI

[XFSC13] Jing Xie, Evan Fletcher, Baljeet Singh, and Owen Carmichael. Robust measurement of individual localized changes to the aging hippocampus. Computer Vision and Image Understanding: CVIU, 117(9): 1128–1137, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314213000696.

Xiang:2008:IAA

[XG08a] Tao Xiang and Shaogang Gong. Incremental and adaptive abnormal behaviour detection. Computer Vision and Image Understanding: CVIU, 111(1): 59–73, July 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xiang:2008:ODG

[XG08b] Tao Xiang and Shaogang Gong. Optimising dynamic graphical models for video content analysis. Computer Vision and Image Understanding: CVIU, 112(3): 310–323, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2022:CED

[XGT^{+}22] Han Xu, Meiqi Gong, Xin Tian, Jun Huang, and Jiayi Ma. CUFD: an encoderdecoder network for visible and infrared image fusion based on common and unique feature decomposition. Computer Vision and Image Understanding: CVIU, 218(??): ??, April 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000352.

Xu:2012:SST

[XHJF12] Yong Xu, Sibin Huang, Hui Ji, and Cornelia Fermüller. Scale-space texture description on SIFT-like textons. Computer Vision and Image Understanding: CVIU, 116(9):999– 1013, September 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000781.

Xiao:2009:GMG

[XHW09] Bai Xiao, Edwin R. Hancock, and Richard C. Wilson. A generative model for graph matching and embedding. Computer Vision and Image Understanding: CVIU, 113(7):777– 789, July 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xia:2019:GGS

[XHX^{+}19] Gui-Song Xia, Jin Huang, Nan Xue, Qikai Lu, and Xiaoxiang Zhu. GeoSay: a geometric saliency for extracting buildings in remote sensing images. Computer Vision and Image Understanding: CVIU, 186(??):37–47, September 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S1077314219300918.

Xu:2012:ISP

[XJK12] Li Xu, Jiaya Jia, and Sing Bing Kang. Improving sub-pixel correspondence through upsampling. Computer Vision and Image Understanding: CVIU, 116(2):250–261, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731421100258X.

Xiong:1998:ESC

[XL98] Wei Xiong and John Chung-Mong Lee. Efficient scene change detection and camera motion annotation for video classification. Computer Vision and Image Understanding: CVIU, 71(2): 166–181, August 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0711/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0711/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0711/production/ref.

Xu:2015:LFD

[XMN^{+}15] Yichao Xu, Kazuki Maeno, Hajime Nagahara, Atsushi Shimada, and Rin ichiro Taniguchi. Light field distortion feature for transparent object classification. Computer Vision and Image Understanding: CVIU, 139(??):122– 135, October 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000430.

Xu:2022:SCN

[XMT22] Sixiang Xu, Damien Muselet, and Alain Trémeau. Sparse coding and normalization for deep Fisher score representation. Computer Vision and Image Understanding: CVIU, 220(??): ??, July 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000571.

Xiao:2005:EPS

[XOF05] G. Xiao, S. H. Ong, and K. W. C. Foong. Efficient partial-surface registration for 3D objects. Computer Vision and Image Understanding: CVIU, 98(2):271– 293, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2011:SIC

[XP11] Mai Xu and Maria Petrou. 3D Scene interpretation by combining probability theory and logic: The tower of knowledge. Computer Vision and Image Understanding: CVIU, 115(11):1581– 1596, November 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314211001834.

Xing:2023:DTC

[XQZL23] Zhengyi Xing, Yulong Qiao, Yue Zhao, and Wenhui Liu. Dynamic texture classification based on bag-of-models with mixture of Student’s t-hidden Markov models. Computer Vision and Image Understanding: CVIU, 230(??):??, April 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223000334.

Xiong:1998:DSD

[XS98] Yalin Xiong and Steven A. Shafer. Dense structure from a dense optical flow sequence. Computer Vision and Image Understanding: CVIU, 69(2):222– 245, February 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/

links/artid/cviu.1997.
0601/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0601/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0601/production/ref.

Xiao:2004:TVM

[XS04] Jiangjian Xiao and Mubarak Shah. Tri-view morphing. Computer Vision and Image Understanding: CVIU, 96(3):345–366, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2012:TSA

[XSD12] Ziyue Xu, Punam K.

Saha, and Soura Dasgupta. Tensor scale: an analytic approach with efficient computation and applications. Computer Vision and Image Understanding: CVIU, 116(10):1060– 1075, October 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212000896.

Xu:2015:SML

[XSK15] Zezhong Xu, Bok-Suk Shin, and Reinhard Klette. A statistical method for line segment detection. Computer Vision and Image Understanding: CVIU, 138(??): 61–73, September 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215001174.

Xiang:2023:IRA

[XSL^{+}23] Wenzhao Xiang, Hang Su, Chang Liu, Yandong Guo, and Shibao Zheng. Improving the robustness of adversarial attacks using an affine-invariant gradient estimator. Computer Vision and Image Understanding: CVIU, 229(??): ??, March 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314223000279.

Xin:2024:TTF

[XSL^{+}24] Zewei Xin, Shalayiding Sirejiding, Yuxiang Lu, Yue Ding, Chunlin Wang, Tamam Alsarhan, and Hongtao Lu. TFUT: Task fusion upward transformer model for multi-task learning on dense prediction. Computer Vision and Image Understanding: CVIU, 244(??):??, July 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731422400095X.

Xu:2015:DSD

[XSQZ15] Yong Xu, Yuping Sun, Yuhui Quan, and Bo Zheng. Discriminative structured dictionary learning with hierarchical group sparsity. Computer Vision and Image Understanding: CVIU, 136 (??):59–68, July 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000211.

Xu:2004:DLM

[XST04] Yaowu Xu, Eli Saber, and A. Murat Tekalp. Dynamic learning from multiple examples for semantic object segmentation and search. Computer Vision and Image Understanding: CVIU, 95(3):334–353, September 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2020:LLM

[XSZ^{+}20] Wenjie Xu, Huihui Song, Kaihua Zhang, Qingshan Liu, and Jia Liu. Learning lightweight multi-scale feedback residual network for single image superresolution. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 103005, August 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300722.

Xiao:2018:BVD

[XTZ^{+}18] Jinsheng Xiao, Hong Tian, Yongqin Zhang, Yongqiang Zhou, and Junfeng Lei. Blind video denoising via texture-aware noise estimation. Computer Vision and Image Understanding: CVIU, 169(??): 1–13, April 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S107731421730200X.

Xie:2014:FAN

[XTZZ14] Lingxi Xie, Qi Tian, Wengang Zhou, and Bo Zhang. Fast and accurate nearduplicate image search with affinity propagation on the ImageWeb. Computer Vision and Image Understanding: CVIU, 124(??):31– 41, July 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213002427.

Xu:2016:NDL

[XW16] Mai Xu and Zulin Wang. A novel double-layer sparse representation approach for unsupervised dictionary learning. Computer Vision and Image Understanding: CVIU, 143(??):1–10, February 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002179.

Xing:2023:SCM

[XWC^{+}23a] Changda Xing, Meiling Wang, Yuhua Cong, Zhisheng Wang, Chaowei Duan, and Yiliu Liu. Sparse coding with morphology segmentation and multi-label fusion for hyperspectral image super-resolution. Computer Vision and Image Understanding: CVIU, 227(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001813.

Xu:2023:GGA

[XWC^{+}23b] Bangwu Xu, Qin Wu, Zhilei Chai, Xueliang Guo, and Jianbo Shi. GAMA: Geometric analysis based motion-aware architecture for moving object segmentation. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001315.

Xia:2023:SLC

[XWDL23] Ziheng Xia, Penghui Wang, Ganggang Dong, and Hongwei Liu. Spatial location constraint prototype loss for open set recognition. Computer Vision and Image Understanding: CVIU, 229 (??):??, March 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000310.

Xu:2023:EFF

[XWLY23] Leiyang Xu, Qiang Wang, Xiaotian Lin, and Lin Yuan. An efficient framework for few-shot skeletonbased temporal action segmentation. Computer Vision and Image Understanding: CVIU, 232(??): ??, July 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000875.

Xu:2015:RSS

[XXCR15] Jun Xu, Kui Xu, Ke Chen, and Jishou Ruan. Reweighted sparse subspace clustering. Computer Vision and Image Understanding: CVIU, 138(??):25–37, September 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000879.

Xiang:2024:TAF

[XYQE24] Xuezhi Xiang, Hang Yin, Yulong Qiao, and Abdulmotaleb El Saddik. Temporal adaptive feature pyramid network for action detection. Computer Vision and Image Understanding: CVIU, 240(??): ??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000262.

Xu:2017:DAE

[XYRS17] Dan Xu, Yan Yan, Elisa Ricci, and Nicu Sebe. Detecting anomalous events in videos by learning deep representations of appearance and motion. Computer Vision and Image Understanding: CVIU, 156(??): 117–127, March 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216301618.

Xu:2008:ASD

[XYW^{+}08] Shuchang Xu, Xiuzi Ye, Yin Wu, Franck Giron, Jean-Luc Leveque, and Bernard Querleux. Automatic skin decomposition based on single image. Computer Vision and Image Understanding: CVIU, 110(1):1– 6, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2011:LST

[XYW11] Jiang Xu, Junsong Yuan, and Ying Wu. Learning spatio-temporal dependency of local patches for complex motion segmentation. Computer Vision and Image Understanding: CVIU, 115 (3):334–351, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xie:2016:LOS

[XYZ16] Xiaohua Xie, Lingxiao Yang, and Wei-Shi Zheng. Learning object-specific DAGs for multi-label material recognition. Computer Vision and Image Understanding: CVIU, 143(??):183– 190, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002623.

Xiao:2011:LIS

[XYZH11] Bai Xiao, Song Yi-Zhe, and Peter Hall. Learning invariant structure for object identification by using graph methods. Computer Vision and Image Understanding: CVIU, 115(7): 1023–1031, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000841.

Xu:2021:ADN

[XZQJ21] Yong Xu, Ye Zhu, Yuhui Quan, and Hui Ji. Attentive deep network for blind motion deblurring on dynamic scenes. Computer Vision and Image Understanding: CVIU, 205(??):Article 103169, April 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000138.

Xu:2023:STT

[XZW^{+}23] Feifei Xu, Yitao Zhu, Chun Wang, Yangze Cao, Zheng Zhong, and Xiongmin Li. Spatio-temporal two-stage fusion for video question answering. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002011.

Xie:2021:CKO

[XZX^{+}21] Cheng Xie, Ting Zeng, Hongxin Xiang, Keqin Li, Yun Yang, and Qing Liu. Class knowledge overlay to visual feature learning for zero-shot image classification. Computer Vision and Image Understanding: CVIU, 207(??): ??, June 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314221000503.

Yang:2012:SRN

[YA12] Qingxiong Yang and Narendra Ahuja. Surface reflectance and normal estimation from photometric stereo. Computer Vision and Image Understanding: CVIU, 116(7):793– 802, July 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314212000483.

Yao:2008:ILR

[YAK^{+}08] Yi Yao, Besma R. Abidi, Nathan D. Kalka, Natalia A. Schmid, and Mongi A. Abidi. Improving long range and high magnification face recognition: Database acquisition, evaluation, and enhancement. Computer Vision and Image Understanding: CVIU, 111(2):111–125, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yuan:2020:CMS

[YARL^{+}20] Juefei Yuan, Hameed Abdul-Rashid, Bo Li, Yijuan Lu, Tobias Schreck, Song Bai, Xiang Bai, Ngoc-Minh Bui, Minh N. Do, Trong-Le Do, Anh-Duc Duong, Kai He, Xinwei He, Mike Holenderski, Dmitri Jarnikov, Tu-Khiem Le, Wenhui Li, Anan Liu, Xiaolong Liu, Vlado Menkovski, Khac-Tuan Nguyen, Thanh-An Nguyen, Vinh-Tiep Nguyen, Weizhi Nie, Van-Tu Ninh, Perez Rey, Yuting Su, Vinh Ton-That, Minh-Triet Tran, Tianyang Wang, Shu Xiang, Shandian Zhe, Heyu Zhou, Yang Zhou, and Zhichao Zhou. A comparison of methods for 3D scene shape retrieval. Computer Vision and Image Understanding: CVIU, 201 (??):Article 103070, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220301090.

Yu:1995:GSA

[YB95] Shan Yu and Marc Berthod.

A game strategy approach for image labeling. Computer Vision and Image Understanding: CVIU, 61(1): 32–37, January 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1003/production;   http:
//www.idealibrary.com/

                                        links/artid/cviu.1995.
                                        1003/production/pdf.

Yacoob:1999:PMR

[YB99] Yaser Yacoob and Michael J. Black. Parameterized modeling and recognition of activities. Computer Vision and Image Understanding: CVIU, 73(2):232– 247, February 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0726/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1998.
                                        0726/production/pdf;
                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0726/production/ref.

Yin:2001:GRF

[YB01] Lijun Yin and Anup Basu. Generating realistic facial expressions with wrinkles for model-based coding. Computer Vision and Image Understanding: CVIU, 84(2):201–240, November 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        cviu.2001.0949;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/cviu.
                                        2001.0949/pdf;   http:
                                        //www.idealibrary.com/

links/doi/10.1006/cviu.
2001.0949/ref.

Yan:2007:FAI

[YB07] Ping Yan and Kevin W.

Bowyer. A fast algorithm for ICP-based 3D shape biometrics. Computer Vision and Image Understanding: CVIU, 107(3):195–202, September 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yi:1998:MBO

[YC98] June Ho Yi and David M.

Chelberg. Model-based 3D object recognition using Bayesian indexing. Computer Vision and Image Understanding: CVIU, 69(1): 087–105, January 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0597/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0597/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0597/production/ref.

Yoo:2005:NNI

[YC05] Dong Hyun Yoo and Myung Jin Chung. A novel nonintrusive eye gaze estimation using cross-ratio under large head motion. Computer Vision and Image Understanding: CVIU, 98(1): 25–51, April 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yang:2010:ABF

[YCA^{+}10] Ming Yang, James Crenshaw, Bruce Augustine, Russell Mareachen, and Ying Wu. AdaBoost-based face detection for embedded systems. Computer Vision and Image Understanding: CVIU, 114(11):1116–1125, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yalcin:2007:BEU

[YCH07] Hulya Yalcin, Robert Collins, and Martial Hebert. Background estimation under rapid gain change in thermal imagery. Computer Vision and Image Understanding: CVIU, 106(2–3): 148–161, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yao:2010:AOC

[YCKA10] Yi Yao, Chung-Hao Chen, Andreas Koschan, and Mongi Abidi. Adaptive online camera coordination for multi-camera multi-target surveillance. Computer Vision and Image Understanding: CVIU, 114(4):463– 474, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yu:2007:NSM

[YCL07] Yuan-Hui Yu, Chin-Chen Chang, and Iuon-Chang Lin. A new steganographic method for color and grayscale image hiding. Computer Vision and Image Understanding: CVIU, 107(3):183–194, September 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yang:2023:BAR

[YCZ^{+}23] Min Yang, Guo Chen, Yin-Dong Zheng, Tong Lu, and Limin Wang. BasicTAD: an astounding RGB-only baseline for temporal action detection. Computer Vision and Image Understanding: CVIU, 232(??): ??, July 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000723.

Yan:2020:FGF

[YDP^{+}20] Yongzhe Yan, Stefan Duffner, Priyanka Phutane, Anthony Berthelier, Xavier Naturel, Christophe Blanc, Christophe Garcia, and Thierry Chateau. Finegrained facial landmark detection exploiting intermediate feature representations. Computer Vision and Image Understanding: CVIU, 200(??):Article 103036, November 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220300886.

Youssef:2017:SBD

[YFDA17] Bendraou Youssef, Essannouni Fedwa, Aboutajdine Driss, and Salam Ahmed. Shot boundary detection via adaptive low rank and svd-updating. Computer Vision and Image Understanding: CVIU, 161(??): 20–28, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301157.

Yang:2023:MLC

[YFF^{+}23] Chengzhuan Yang, Lincong Fang, Benjie Fei, Qian Yu, and Hui Wei. Multi-level contour combination features for shape recognition. Computer Vision and Image Understanding: CVIU, 229 (??):??, March 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000309.

Yang:2018:VSR

[YFX^{+}18] Wenhan Yang, Jiashi Feng, Guosen Xie, Jiaying Liu, Zongming Guo, and Shuicheng Yan. Video super-resolution based on spatial-temporal recurrent residual networks. Computer Vision and Image Understanding: CVIU, 168 (??):79–92, ???? 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301583.

Yuille:1998:IWS

[YFZ98] Alan L. Yuille, Mario Ferraro, and Tony Zhang. Image warping for shape recovery and recognition. Computer Vision and Image Understanding: CVIU, 72(3): 351–359, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0676/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0676/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0676/production/ref.

Yun:2016:HFD

[YG16] Yixiao Yun and Irene Yu-Hua Gu. Human fall detection in videos via boosting and fusing statistical features of appearance, shape and motion dynamics on Riemannian manifolds with applications to assisted living. Computer Vision and Image Understanding: CVIU, 148(??): 111–122, July 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215002659.

Yun:2017:RMV

[YG17] Yixiao Yun and Irene Yu-Hua Gu. Riemannian manifold-valued partbased features and geodesicinduced kernel machine for activity classification dedicated to assisted living. Computer Vision and Image Understanding: CVIU, 161(??):65– 76, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301078.

Yang:2013:TLA

[YGC13] Wanqi Yang, Yang Gao, and Longbing Cao. TRASMIL: a local anomaly detection framework based on trajectory segmentation and multi-instance learning. Computer Vision and Image Understanding: CVIU, 117(10):1273–1286, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001658.

Yin:2015:ISR

[YGC15] Ming Yin, Junbin Gao, and Shuting Cai. Image superresolution via 2D tensor regression learning. Computer Vision and Image Understanding: CVIU, 132(??): 12–23, March 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214002252.

Yang:2011:RLD

[YGH11] Kun Yang, Shuzhi Sam Ge, and Hongsheng He. Robust line detection using twoorthogonal direction image scanning. Computer Vision and Image Understanding: CVIU, 115(8):1207– 1222, August 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000981.

Yan:2020:VSP

[YGJ^{+}20] Xiyu Yan, Huihui Gong, Yong Jiang, Shu-Tao Xia, Feng Zheng, Xinge You, and Ling Shao. Video scene parsing: an overview of deep learning methods and datasets. Computer Vision and Image Understanding: CVIU, 201 (??):Article 103077, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314220301120.

Yang:2019:VSR

[YH19] Liang Yang and Haifeng Hu. Visual skeleton and reparative attention for part-of-speech image captioning system. Computer Vision and Image Understanding: CVIU, 189 (??):Article 102819, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219301286.

Yang:2011:SOD

[YHN11] Bo Yang, Chang Huang, and Ram Nevatia. Segmentation of objects in a detection window by Nonparametric Inhomogeneous CRFs. Computer Vision and Image Understanding: CVIU, 115(11):1473–1482, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314211001779.

Yang:2005:SBD

[YHR^{+}05] Allen Y. Yang, Kun Huang, Shankar Rao, Wei Hong, and Yi Ma. Symmetrybased 3-D reconstruction from perspective images. Computer Vision and Image Understanding: CVIU, 99(2):210–240, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yi:1995:OSL

[YHS95] Seungku K. Yi, Robert M.

Haralick, and Linda G. Shapiro. Optimal sensor and light source positioning for machine vision. Computer Vision and Image Understanding: CVIU, 61(1): 122–137, January 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1995.
1009/production;   http:
//www.idealibrary.com/
links/artid/cviu.1995.
1009/production/pdf.

Yang:2020:IVA

[YHS^{+}20] Lie Yang, Guanghua Hu, Yonghao Song, Guofeng Li, and Longhan Xie. Intelligent video analysis: a pedestrian trajectory extraction method for the whole indoor space without blind areas. Computer Vision and Image Understanding: CVIU, 196(??):Article 102968, July 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220300473.

Yang:2023:FSO

[YHW^{+}23] Xiaobao Yang, Yulong He, Junsheng Wu, Wei Sun, Tianyu Liu, and Sugang Ma. 3DF-FCOS: Small object detection with 3D features based on FCOS. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001674.

Yang:2016:TDA

[YJ16] Min Yang and Yunde Jia. Temporal dynamic appearance modeling for online multi-person tracking. Computer Vision and Image Understanding: CVIU, 153(??):16–28, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216300510.

Yla-Jaaski:1996:GSS

[YJA96] Antti Ylä-Jääski and Frank Ade. Grouping symmetrical structures for object segmentation and description. Computer Vision and Image Understanding: CVIU, 63 (3):399–417, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0031/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0031/production/pdf.

Yu:2009:ACC

[YJC^{+}09] Xinguo Yu, Nianjuan Jiang, Loong-Fah Cheong, Hon Wai Leong, and Xin Yan. Automatic camera calibration of broadcast tennis video with applications to 3D virtual content insertion and ball detection and tracking. Computer Vision and Image Understanding: CVIU, 113(5):643–652, May 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yoon:2008:DSM

[YK08] Kuk-Jin Yoon and In So Kweon. Distinctive Similarity Measure for stereo matching under point ambiguity. Computer Vision and Image Understanding: CVIU, 112(2):173–183, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yaganapu:2024:MLS

[YK24] Avinash Yaganapu and Mingon Kang. Multilayered self-attention mechanism for weakly supervised semantic segmentation. Computer Vision and Image Understanding: CVIU, 239(??):??, February 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002667.

Yang:2001:FDU

[YKA01] Ming-Hsuan Yang, David Kriegman, and Narendra Ahuja. Face detection using multimodal density models. Computer Vision and Image Understanding: CVIU, 84(2): 264–284, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0937;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0937/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0937/ref.

Yu:2009:SSS

[YLA09] Tianli Yu, Jiebo Luo, and Narendra Ahuja. Search strategies for shape regularized active contour. Computer Vision and Image Understanding: CVIU, 113 (10):1053–1063, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yoo:2023:PER

[YLK^{+}23] Hongsang Yoo, Haopeng Li, Qiuhong Ke, Liangchen Liu, and Rui Zhang. Precondition and effect reasoning for action recognition. Computer Vision and Image Understanding: CVIU, 232(??):??, July 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000711.

Yang:2018:TET

[YLLG18] Shuai Yang, Jiaying Liu, Zhouhui Lian, and Zongming Guo. Text effects transfer via distributionaware texture synthesis. Computer Vision and Image Understanding: CVIU, 174 (??):43–56, September 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314218301139.

Yang:2011:DSE

[YLM11] Peng Yang, Qingshan Liu, and Dimitris Metaxas. Dynamic soft encoded patterns for facial event analysis. Computer Vision and Image Understanding: CVIU, 115 (3):456–465, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yan:2017:GEL

[YLM^{+}17] Yan Yan, Jiwen Lu, Ajmal Mian, Arun Ross, Vittorio Murino, and Radu Horaud. Guest editorial: Language in vision. Computer Vision and Image Understanding: CVIU, 163(??): 1–2, October 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217301868.

Yu:2018:NPI

[YLX^{+}18] Qinghua Yu, Jie Liang, Junhao Xiao, Huimin Lu, and Zhiqiang Zheng. A novel perspective invariant feature transform for RGB-D images. Computer Vision and Image Understanding: CVIU, 167(??):109– 120, February 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217302242.

Yin:2011:HKP

[YNCO11] Shimin Yin, Jin Hee Na, Jin Young Choi, and Songhwai Oh. Hierarchical Kalman-particle filter with adaptation to motion changes for object tracking. Computer Vision and Image Understanding: CVIU, 115(6):885–900, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yan:2019:CMM

[YNZ^{+}19] Yichao Yan, Bingbing Ni, Wendong Zhang, Jun Tang, and Xiaokang Yang. Crossmodality motion parameterization for fine-grained video prediction. Computer Vision and Image Understanding: CVIU, 183(??): 11–19, June 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314219300426.

Yao:2011:FHD

[YO11] Jian Yao and Jean-Marc Odobez. Fast human detection from joint appearance and foreground feature subset covariances. Computer Vision and Image Understanding: CVIU, 115(10): 1414–1426, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314211001391.

Yan:2023:OTB

[YQL^{+}23] Shilei Yan, Yujuan Qi, Mengxue Liu, Yanjiang Wang, and Baodi Liu. Object tracking based on siamese network with 3D attention and multiple graph attention. Computer Vision and Image Understanding: CVIU, 235(??): ??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001662.

Yue:2006:VMP

[YR06] Shigang Yue and F. Claire Rind. Visual motion pattern extraction and fusion for collision detection in complex dynamic scenes. Computer Vision and Image Understanding: CVIU, 104(1): 48–60, October 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yilmaz:2006:MAP

[YS06] Alper Yilmaz and Mubarak Shah. Matching actions in presence of camera motion. Computer Vision and Image Understanding: CVIU, 104 (2–3):221–231, November/ December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yilmaz:2008:DGA

[YS08] Alper Yilmaz and Mubarak Shah. A differential geometric approach to representing the human actions. Computer Vision and Image Understanding: CVIU, 109 (3):335–351, March 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yang:2009:CGM

[YS09] Ruiduo Yang and Sudeep Sarkar. Coupled grouping and matching for sign and gesture recognition. Computer Vision and Image Understanding: CVIU, 113(6): 663–681, June 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yao:2024:MGB

[YS24] Guilin Yao and Anming Sun. Multi-guided-based image matting via boundary detection. Computer Vision and Image Understanding: CVIU, 243(??): ??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000791.

Yu:2003:MMA

[YSD03] Weichuan Yu, Gerald Sommer, and Kostas Daniilidis. Multiple motion analysis: in spatial or in spectral domain? Computer Vision and Image Understanding: CVIU, 90(2):129– 152, May 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yu:2024:EUJ

[YSKL24] Yeonguk Yu, Sungho Shin, Minhwan Ko, and Kyoobin Lee. Exploring using jigsaw puzzles for outof-distribution detection. Computer Vision and Image Understanding: CVIU, 241(??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000493.

Yu:2011:GOE

[YSL11] Chanki Yu, Yongduek Seo, and Sang Wook Lee. Global optimization for estimating a multiple-lobe analytical BRDF. Computer Vision and Image Understanding: CVIU, 115(12):1679– 1688, December 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211001615.

Yan:2014:GTY

[YSL^{+}14] Yan Yan, Haoquan Shen, Gaowen Liu, Zhigang Ma, Chenqiang Gao, and Nicu Sebe. GLocal tells you more: Coupling GLocal structural for feature selection with sparsity for image and video classification. Computer Vision and Image Understanding: CVIU, 124(??): 99–109, July 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000320.

Yoshinaga:2014:ODB

[YSNiT14] Satoshi Yoshinaga, Atsushi Shimada, Hajime Nagahara, and Rin ichiro Taniguchi. Object detection based on spatiotemporal background models. Computer Vision and Image Understanding: CVIU, 122(??):84– 91, May 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314213002245.

Yang:2014:SSL

[YSS^{+}14] Yi Yang, Nicu Sebe, Cees Snoek, Xian-Sheng Hua, and Yueting Zhuang. Special section on learning from multiple evidences for large scale multimedia analysis. Computer Vision and Image Understanding: CVIU, 118 (??):1, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001689.

Yoshikawa:2021:MMV

[YST21] Yuya Yoshikawa, Yutaro Shigeto, and Akikazu Takeuchi. MetaVD: a Meta Video Dataset for enhancing human action recognition datasets. Computer Vision and Image Understanding: CVIU, 212(??): ??, November 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422100120X.

Yuan:2019:EPH

[YSX^{+}19] Feiniu Yuan, Jinting Shi, Xue Xia, Lin Zhang, and Shuying Li. Encoding pairwise Hamming distances of local binary patterns for visual smoke recognition. Computer Vision and Image Understanding: CVIU, 178(??):43–53, January 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218304260.

Yu:2018:HSI

[YSY^{+}18] Wei Yu, Xiaoshuai Sun, Kuiyuan Yang, Yong Rui, and Hongxun Yao. Hierarchical semantic image matching using CNN feature pyramid. Computer Vision and Image Understanding: CVIU, 169(??): 40–51, April 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300018.

Yuan:2023:URI

[YSZ23] Nianzeng Yuan, Bangyong Sun, and Xiangtao Zheng. Unsupervised real image super-resolution via knowledge distillation network. Computer Vision and Image Understanding: CVIU, 234(??):??, September 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223001169.

Ye:1999:SPO

[YT99] Yiming Ye and John K.

Tsotsos. Sensor planning for 3D object search. Computer Vision and Image Understanding: CVIU, 73(2):145– 168, February 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0736/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0736/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0736/production/ref.

Yi:2013:TES

[YT13] Chucai Yi and Yingli Tian.

Text extraction from scene images by character appearance and structure modeling. Computer Vision and Image Understanding: CVIU, 117(2):182–194, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001701.

Yu:1999:HMV

[YW99] Hong Heather Yu and Wayne Wolf. A hierarchical multiresolution video shot transition detection scheme. Computer Vision and Image Understanding: CVIU, 75(1–2):196– 213, July/August 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1999.
0773/production;   http:
//www.idealibrary.com/
links/artid/cviu.1999.
0773/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1999.
0773/production/ref.

Yemez:2007:VFT

[YW07] Y. Yemez and C. J. Wetherilt. A volumetric fusion technique for surface reconstruction from silhouettes and range data. Computer Vision and Image Understanding: CVIU, 105(1):30– 41, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yu:2016:SPU

[YW16] Tianshu Yu and Ruisheng Wang. Scene parsing using graph matching on streetview data. Computer Vision and Image Understanding: CVIU, 145(??):70– 80, April 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216000175.

Yang:2023:CBO

[YWH^{+}23] Shiqi Yang, Yaxing Wang, Luis Herranz, Shangling Jui, and Joost van de Weijer. Casting a BAIT for offline and online source-free domain adaptation. Computer Vision and Image Understanding: CVIU, 234(??): ??, September 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001273.

Yan:2020:GRC

[YWL^{+}20] Qingsen Yan, Bo Wang, Peipei Li, Xianjun Li, Ao Zhang, Qinfeng Shi, Zheng You, Yu Zhu, Jinqiu Sun, and Yanning Zhang. Ghost removal via channel attention in exposure fusion. Computer Vision and Image Understanding: CVIU, 201(??):Article 103079, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314220301132.

Yang:2022:NFC

[YWL^{+}22] Jianhua Yang, Ke Wang, Ruifeng Li, Zhonghao Qin, and Petra Perner. A novel fast combine-and-conquer object detector based on only one-level feature map. Computer Vision and Image Understanding: CVIU, 224(??): ??, November 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001394.

Yi:2019:AAS

[YWM19] Jingru Yi, Pengxiang Wu, and Dimitris N. Metaxas. ASSD: Attentive single shot multibox detector. Computer Vision and Image Understanding: CVIU, 189 (??):Article 102827, December 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314219301328.

Yang:2008:USN

[YWMS08] Allen Y. Yang, John Wright, Yi Ma, and S. Shankar Sastry. Unsupervised segmentation of natural images via lossy data compression.

Computer Vision and Image Understanding: CVIU, 110(2):212–225, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yang:2016:IMS

[YWY^{+}16] Jianyu Yang, Hongxing Wang, Junsong Yuan, Youfu Li, and Jianyang Liu. Invariant multi-scale descriptor for shape representation, matching and retrieval. Computer Vision and Image Understanding: CVIU, 145 (??):43–58, April 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216000187.

Yan:2011:RRT

[YWZ11] Chao Yan, Yuangqing Wang, and Zhaoyang Zhang. Robust real-time multi-user pupil detection and tracking under various illumination and large-scale head motion. Computer Vision and Image Understanding: CVIU, 115(8):1223– 1238, August 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000890.

Yang:2024:SCL

[YXW^{+}24] Shuwen Yang, Luwei Xiao, Xingjiao Wu, Junjie Xu, Linlin Wang, and Liang He. Simple contrastive learning in a self-supervised manner for robust visual question answering. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000572.

Ye:2023:CCL

[YYH^{+}23] Shuo Ye, Shujian Yu, Wenjin Hou, Yu Wang, and Xinge You. Coping with change: Learning invariant and minimum sufficient representations for fine-grained visual categorization. Computer Vision and Image Understanding: CVIU, 237(??):??, December 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002175.

Yamamoto:1996:AFV

[YYL96] Hiroyuki Yamamoto, Yehezkel Yeshurun, and Martin D. Levine. An active foveated vision system: Attentional mechanisms and scan path convergence measures. Computer Vision and Image Understanding: CVIU, 63(1):50–65, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0004/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0004/production/pdf.

Yeung:1998:SVC

[YYL98] Minerva Yeung, Boon-Lock Yeo, and Bede Liu. Segmentation of video by clustering and graph analysis. Computer Vision and Image Understanding: CVIU, 71 (1):94–109, July 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0628/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0628/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0628/production/ref.

Yang:2019:SHM

[YYZL19] Jing Yang, Xu Yang, Zhang-Bing Zhou, and Zhi-Yong Liu. Sub-hypergraph matching based on adjacency tensor. Computer Vision and Image Understanding: CVIU, 183(??): 1–10, June 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219300323.

Yao:2006:HSD

[YZ06] Jian Yao and Zhonefei (Mark) Zhang. Hierarchical shadow detection for color aerial images. Computer Vision and Image Understanding: CVIU, 102(1): 60–69, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yang:2016:ROT

[YZL16] Jing Yang, Kaihua Zhang, and Qingshan Liu. Robust object tracking by online Fisher discrimination boosting feature selection. Computer Vision and Image Understanding: CVIU, 153(??):100–108, December 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314216000485.

Yu:2021:DIB

[YZL^{+}21] Cuican Yu, Zihui Zhang, Huibin Li, Jian Sun, and Zongben Xu. A distribution independence based method for 3D face shape decomposition. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221000886.

Yan:2013:GSC

[YZT^{+}13] Pingkun Yan, Wuxia Zhang, Baris Turkbey, Peter L. Choyke, and Xuelong Li. Global structure constrained local shape prior estimation for medical image segmentation. Computer Vision and Image Understanding: CVIU, 117(9):1017– 1026, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731421300074X.

Yang:2017:RCS

[YZX^{+}17] Jiaqi Yang, Qian Zhang, Ke Xian, Yang Xiao, and Zhiguo Cao. Rotational contour signatures for both real-valued and binary feature representations of 3D local shape. Computer Vision and Image Understanding: CVIU, 160(??):133– 147, July 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217300322.

Yuan:2020:IDB

[YZX^{+}20] Feiniu Yuan, Yu Zhou, Xue Xia, Jinting Shi, Yuming Fang, and Xueming Qian. Image dehazing based on a transmission fusion strategy by automatic image matting. Computer Vision and Image Understanding: CVIU, 194(??):Article 102933, May 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300217.

Yu:2022:HRM

[YZX^{+}22] Simin Yu, Kuntian Zhang, Chuan Xiao, Joshua Zhexue Huang, Mark Junjie Li, and Makoto Onizuka. HS-GAN: Reducing mode collapse in GANs by the latent code distance of homogeneous samples. Computer Vision and Image Understanding: CVIU, 214(??): ??, January 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314221001570.

Yang:2011:UCA

[YZY11] Lei Yang, Nanning Zheng, and Jie Yang. A unified context assessing model for object categorization. Computer Vision and Image Understanding: CVIU, 115(3): 310–322, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2022:REU

[ZA22] Xiaowei Zhang and Daniel Aliaga. RFCNet: Enhancing urban segmentation using regularization, fusion, and completion. Computer Vision and Image Understanding: CVIU, 220(??): ??, July 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S107731422200056X.

Zach:2018:GFM

[Zac18] Christopher Zach. Generalized fusion moves for continuous label optimization. Computer Vision and Image Understanding: CVIU, 173(??):47– 56, August 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300559.

Zhang:2015:ESB

[ZBDP15] Hao Zhang, J. Ross Beveridge, Bruce A. Draper, and P. Jonathon Phillips. On the effectiveness of soft biometrics for increasing face verification rates. Computer Vision and Image Understanding: CVIU, 137(??):50– 62, August 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314215000491.

Zhang:2019:HLP

[ZC19] Haoruo Zhang and Qixin Cao. Holistic and local patch framework for 6D object pose estimation in RGB-D images. Computer Vision and Image Understanding: CVIU, 180(??): 59–73, March 2019. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314219300050.

Zhou:2013:NBS

[ZCF13] Hong Zhou, Yiru Chen, and Rong Feng. A novel background subtraction method based on color invariants. Computer Vision and Image Understanding: CVIU, 117(11):1589–1597, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314213001380.

Zivkovic:2009:ABM

[ZCK09] Zoran Zivkovic, Ali Taylan Cemgil, and Ben Kröse. Approximate Bayesian methods for kernel-based object tracking. Computer Vision and Image Understanding: CVIU, 113(6):743– 749, June 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhu:1999:ICE

[ZCL99] Hui Zhu, Francis H. Y.

Chan, and F. K. Lam. Image contrast enhancement by constrained local histogram equalization. Computer Vision and Image Understanding: CVIU, 73(2):281–290, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0723/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0723/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0723/production/ref.

Zhao:2020:QEC

[ZCLX20] Bao Zhao, Xiaobo Chen, Xinyi Le, and Juntong Xi. A quantitative evaluation of comprehensive 3D local descriptors generated with spatial and geometrical features. Computer Vision and Image Understanding: CVIU, 190(??):Article 102842, January 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218302698.

Zheng:2023:MEM

[ZCW^{+}23] Ziyang Zheng, Zhixiang Chen, Shuqi Wang, Wenpeng Wang, and Hui Wang. Memory-efficient multi-scale residual dense network for single image rain removal. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001467.

Zhou:2024:SHP

[ZCW24] Lu Zhou, Yingying Chen, and Jinqiao Wang. Slow-FastFormer for 3D human pose estimation. Computer Vision and Image Understanding: CVIU, 243(??): ??, June 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314224000730.

Zou:2023:LRI

[ZCWH23] Yuliang Zou, Jinwoo Choi, Qitong Wang, and Jia-Bin Huang. Learning representational invariances for data-efficient action recognition. Computer Vision and Image Understanding: CVIU, 227(??): ??, January 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314222001758.

Ziou:2001:DDE

[ZD01] Djemel Ziou and François Deschenes. Depth from defocus estimation in spatial domain. Computer Vision and Image Understanding: CVIU, 81(2): 143–165, February 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0899;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0899/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0899/ref.

Zhang:2018:SPI

[ZD18] Mingli Zhang and Christian Desrosiers. Structure preserving image denoising based on low-rank reconstruction and gradient histograms. Computer Vision and Image Understanding: CVIU, 171(??): 48–60, June 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300730.

Zara:2024:SOS

[ZdCR^{+}24] Giacomo Zara, Victor Guilherme Turrisi da Costa, Subhankar Roy, Paolo Rota, and Elisa Ricci. Simplifying open-set video domain adaptation with contrastive learning. Computer Vision and Image Understanding: CVIU, 241(??): ??, April 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314224000341.

Zhu:2010:KSR

[ZDF10] Youding Zhu, Behzad Dariush, and Kikuo Fujimura. Kinematic self retargeting: a framework for human pose estimation. Computer Vision and Image Understanding: CVIU, 114(12):1362– 1375, December 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zappella:2013:JES

[ZDLS13] Luca Zappella, Alessio Del Bue, Xavier Lladó, and Joaquim Salvi. Joint estimation of segmentation and structure from motion. Computer Vision and Image Understanding: CVIU, 117(2):113–129, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314212001312.

Zhou:2023:EEF

[ZDZ^{+}23] Yan Zhou, Zhaolong Dang, Huaidong Zhang, Xuemiao Xu, Jing Qin, Wenjun Li, Fanzhi Zeng, and Xiangyu Liu. EFSCNN: Encoded Feature Sphere Convolution Neural Network for fast nonrigid 3D models classification and retrieval. Computer Vision and Image Understanding: CVIU, 233 (??):??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001042.

Zhang:2015:FVO

[ZEGEJ15] Haopeng Zhang, Tarek El-Gaaly, Ahmed Elgammal, and Zhiguo Jiang. Factorization of view-object manifolds for joint object recognition and pose estimation. Computer Vision and Image Understanding: CVIU, 139(??):89–103, October 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215000715.

Zhang:2008:ISE

[ZFG08] Hui Zhang, Jason E. Fritts, and Sally A. Goldman. Image segmentation evaluation: a survey of unsupervised methods. Computer Vision and Image Understanding: CVIU, 110(2): 260–280, May 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhou:2022:UAC

[ZFG^{+}22] Qianyu Zhou, Zhengyang Feng, Qiqi Gu, Guangliang Cheng, Xuequan Lu, Jianping Shi, and Lizhuang Ma. Uncertainty-aware consistency regularization for cross-domain semantic segmentation. Computer Vision and Image Understanding: CVIU, 221(??): ??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222000625.

Zhou:2024:SAT

[ZFW^{+}24] Kai Zhou, Jing-Long Fang, Wen Wu, Yan-Li Shao, Xing-Qi Wang, and Dan Wei. Semantic-aware transformer for shadow detection. Computer Vision and Image Understanding: CVIU, 240 (??):??, March 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000225.

Zagorchev:2006:PLR

[ZG06] Lyubomir Zagorchev and Ardeshir Goshtasby. A paintbrush laser range scanner. Computer Vision and Image Understanding: CVIU, 101(2):65–86, February 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2010:ACS

[ZG10] Jianguo Zhang and Shaogang Gong. Action categorization by structural probabilistic latent semantic analysis. Computer Vision and Image Understanding: CVIU, 114(8):857–864, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zolna:2020:CAS

[ZGC20] Konrad Zolna, Krzysztof J.

Geras, and Kyunghyun Cho. Classifier-agnostic saliency map extraction. Computer Vision and Image Understanding: CVIU, 196(??): Article 102969, July 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314220300461.

Zhu:2004:LLA

[ZH04] Zhigang Zhu and Allen R.

Hanson. LAMP: 3D layered, adaptive-resolution, and multi-perspective panorama— a new scene representation. Computer Vision and Image Understanding: CVIU, 96(3):294–326, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2018:EBC

[ZH18] Junfeng Zhang and Haifeng Hu. Exemplar-based Cascaded Stacked Auto-Encoder Networks for robust face alignment. Computer Vision and Image Understanding: CVIU, 171(??):95– 103, June 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218300687.

Zhang:1997:TLB

[Zha97]  Zhengyou Zhang.  A

tighter lower bound on the Spetsakis-Aloimonos trilinear constraints. Computer Vision and Image Understanding: CVIU, 67(2):202– 204, August 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1997.
0526/production;   http:
//www.idealibrary.com/
links/artid/cviu.1997.
0526/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1997.
0526/production/ref.

Zhao:2020:AMI

[ZHL^{+}20] Xiaole Zhao, Xiafei Hu, Ying Liao, Tian He, Tao Zhang, Xueming Zou, and Jinsha Tian. Accurate MR image super-resolution via lightweight lateral inhibition network. Computer Vision and Image Understanding: CVIU, 201 (??):Article 103075, December 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220301119.

Zhang:2024:MFB

[ZHS^{+}24] Zhihao Zhang, Jie He, Mouquan Shen, Jiantao Shi, and Xianqiang Yang. Multimodel fore-/background alignment for seam-based parallax-tolerant image stitching. Computer Vision and Image Understanding: CVIU, 240(??):??, March 2024. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223002928.

Zhu:2017:ALE

[ZHZ17] Ming-Zhu Zhu, Bing-Wei He, and Li-Wei Zhang. Atmospheric light estimation in hazy images based on color-plane model. Computer Vision and Image Understanding: CVIU, 165 (??):33–42, December 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314217301637.

Zhang:2022:DAS

[ZhZFL22] Zhi Zhang, Sheng hua Zhong, Ahmed Fares, and Yan Liu. Detecting abnormality with separated foreground and background: Mutual generative adversarial networks for video abnormal event detection. Computer Vision and Image Understanding: CVIU, 219(??):??, June 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422200042X.

Zheng:2013:CFI

[ZIT^{+}13] Bo Zheng, Ryo Ishikawa, Jun Takamatsu, Takeshi Oishi, and Katsushi Ikeuchi. A coarse-to-fine IP-driven registration for pose estimation from single ultrasound image. Computer Vision and Image Understanding: CVIU, 117(12):1647– 1658, December 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213001252.

Zivkovic:2010:WSC

[Ziv10] Zoran Zivkovic. Wireless smart camera network for real-time human 3D pose reconstruction. Computer Vision and Image Understanding: CVIU, 114(11): 1215–1222, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhu:2005:RRT

[ZJ05] Zhiwei Zhu and Qiang Ji.

Robust real-time eye detection and tracking under variable lighting conditions and various face orientations. Computer Vision and Image Understanding: CVIU, 98(1):124– 154, April 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2022:AFO

[ZJJ22] Tao Zhang, Bo Jin, and Wenjing Jia. An anchorfree object detector based on soften optimized bidirectional FPN. Computer Vision and Image Understanding: CVIU, 218(??): ??, April 2022. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314222000388.

Zhou:2023:SYS

[ZJL23] Huayi Zhou, Fei Jiang, and Hongtao Lu. SSDA-YOLO: Semi-supervised domain adaptive YOLO for cross-domain object detection. Computer Vision and Image Understanding: CVIU, 229(??): ??, March 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223000292.

Zhu:2015:SFP

[ZJW15] Xiaolong Zhu, Xuhui Jia, and Kwan-Yee K. Wong. Structured forests for pixellevel hand detection and hand part labelling. Computer Vision and Image Understanding: CVIU, 141(??):95–107, December 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314215001617.

Zeng:2016:AMS

[ZJZY16] Zhi Zeng, Jianyuan Jia, Zhaofei Zhu, and Dalin Yu. Adaptive maintenance scheme for codebook-based dynamic background subtraction. Computer Vision and Image Understanding: CVIU, 152(??):58–66, November 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314216301175.

Zagoruyko:2017:DCS

[ZK17] S. Zagoruyko and N. Komodakis. Deep compare: a study on using convolutional neural networks to compare image patches. Computer Vision and Image Understanding: CVIU, 164(??): 38–55, November 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314217301716.

Zhou:2003:PRH

[ZKC03] Shaohua Zhou, Volker Krueger, and Rama Chellappa. Probabilistic recognition of human faces from video. Computer Vision and Image Understanding: CVIU, 91(1–2):214– 245, July/August 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2023:MSS

[ZKLZ23] Yunzuo Zhang, Weili Kang, Yameng Liu, and Pengfei Zhu. Multi-scale semantic and detail extraction network for lightweight person re-identification. Computer Vision and Image Understanding: CVIU, 236(??): ??, November 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001935.

Zhu:2004:DMC

[ZKRH04] Zhigang Zhu, Deepak R. Karuppiah, Edward M. Riseman, and Allen R. Hanson. Dynamic mutual calibration and view planning for cooperative mobile robots with panoramic virtual stereo vision. Computer Vision and Image Understanding: CVIU, 95(3): 261–286, September 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zadrija:2018:SWS

[ZKSV18] Valentina Zadrija, Josip Krapac, Sinisa Segvić, and Jakob Verbeek. Sparse weakly supervised models for object localization in road environment. Computer Vision and Image Understanding: CVIU, 176– 177(??):9–21, November/ December 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314218304211.

Zhang:2001:EFM

[ZL01] Zhengyou Zhang and Charles Loop. Estimating the fundamental matrix by transforming image points in projective space. Computer Vision and Image Understanding: CVIU, 82(2): 174–180, May 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2001.0909;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0909/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2001.0909/ref.

Zhang:2023:JCD

[ZLFH23] Chengfang Zhang, Haoyue Li, Ziliang Feng, and Sidi He. Joint coupled dictionaries-based visibleinfrared image fusion method via texture preservation structure in sparse domain. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001613.

Zhou:2018:MVI

[ZLHJ18] Shiwei Zhou, Zhengyang Lou, Yu Hen Hu, and Hongrui Jiang. Multiple view image denoising using 3D focus image stacks. Computer Vision and Image Understanding: CVIU, 171(??):

34–47, June 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314218300729.

Zhang:2014:ICN

[ZLL^{+}14] Chunjie Zhang, Jing Liu, Chao Liang, Zhe Xue, Junbiao Pang, and Qingming Huang. Image classification by non-negative sparse coding, correlation constrained low-rank and sparse decomposition. Computer Vision and Image Understanding: CVIU, 123(??): 14–22, June 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314214000393.

Zhu:2024:UDC

[ZLL^{+}24] Hangyan Zhu, Shaohui Liu, Ming Liu, Zifei Yan, and Wangmeng Zuo. De^{2}Net: Under-display camera image restoration with feature deconvolution and kernel decomposition. Computer Vision and Image Understanding: CVIU, 244 (??):??, July 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224001097.

Zhan:2021:DCE

[ZLLP21] Hongjian Zhan, Shujing Lyu, Yue Lu, and Umapada Pal. DenseNet-CTC: an end-to-end RNN-free architecture for context-free string recognition. Computer Vision and Image Understanding: CVIU, 204 (??):Article 103168, March 2021. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314221000126.

Zhou:2013:MSB

[ZLS^{+}13] Huiyu Zhou, Xuelong Li, Gerald Schaefer, M. Emre Celebi, and Paul Miller. Mean shift based gradient vector flow for image segmentation. Computer Vision and Image Understanding: CVIU, 117(9):1004– 1016, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000751.

Zhao:2020:PSS

[ZLY^{+}20] Hao Zhao, Ming Lu, Anbang Yao, Yiwen Guo, Yurong Chen, and Li Zhang. Pointly-supervised scene parsing with uncertainty mixture. Computer Vision and Image Understanding: CVIU, 200(??):Article 103040, November 2020.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1077314220300904.

Zhang:2017:SOD

[ZLZH17] Qiang Zhang, Yi Liu, Siyang Zhu, and Jungong Han. Salient object detection based on super-pixel clustering and unified low-rank representation. Computer Vision and Image Understanding: CVIU, 161(??): 51–64, August 2017. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314217300759.

Zhao:1996:GTD

[ZM96] ChangSheng Zhao and Roger Mohr. Global threedimensional surface reconstruction from occluding contours. Computer Vision and Image Understanding: CVIU, 64(1): 62–96, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0046/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0046/production/pdf.

Zapater:2005:GAS

[ZMCA05] V. Zapater, L. Martínez-Costa, and G. Ayala. A granulometric analysis of specular microscopy images of human corneal endothelia. Computer Vision and Image Understanding: CVIU, 97(3):297– 314, March 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2015:AFP

[ZMJ^{+}15] Li Zhang, Kamlesh Mistry, Ming Jiang, Siew Chin Neoh, and Mohammed Alamgir Hossain. Adaptive facial point detection and emotion recognition for a humanoid robot. Computer Vision and Image Understanding: CVIU, 140(??):93–114, November 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215001605.

Zhang:2022:DAM

[ZMM^{+}22] Yuhang Zhang, Zhenwei Miao, Tiebin Mi, Jie Li, and Robert C. Qiu. Dual adversarial model: Exploring lowdimensional space features for point clouds generating and completing. Computer Vision and Image Understanding: CVIU, 223(??): ??, October 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222001291.

Zhang:2008:EPT

[ZN08] Hongsheng Zhang and Shahriar Negahdaripour. Epiflow

— a paradigm for tracking stereo correspondences. Computer Vision and Image Understanding: CVIU, 111(3):307–328, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhu:2013:MMO

[ZNG^{+}13] Y. Zhu, N. Nayak, U. Gaur, B. Song, and A. Roy-Chowdhury. Modeling multi-object interactions using “string of feature graphs”. Computer Vision and Image Understanding: CVIU, 117(10):1313– 1328, October 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314212001646.

Zhao:2000:INS

[ZOMK00] Hong-Kai Zhao, Stanley Osher, Barry Merriman, and Myungjoo Kang. Implicit and nonparametric shape reconstruction from unorganized data using a variational level set method. Computer Vision and Image Understanding: CVIU, 80(3): 295–314, December 2000.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
cviu.2000.0875;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0875/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/cviu.
2000.0875/ref.

Zafeiriou:2011:EGM

[ZP11] Stefanos Zafeiriou and Maria Petrou. 2.5D Elastic graph matching. Computer Vision and Image Understanding: CVIU, 115(7): 1062–1072, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000762.

Zhang:2018:CRG

[ZP18] Tong Zhang and Fatih Porikli. Cascade residuals guided nonlinear dictionary learning. Computer Vision and Image Understanding: CVIU, 173(??): 86–97, August 2018. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314218300560.

Zhao:2011:GAE

[ZRKZ^{+}11] Haifeng Zhao, Antonio Robles-Kelly, Jun Zhou, Jianfeng Lu, and Jing-Yu Yang. Graph attribute embedding via Riemannian submersion learning. Computer Vision and Image Understanding: CVIU, 115(7): 962–975, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314211000737.

Zhang:2011:KFS

[ZRL^{+}11] Nan Zhang, Su Ruan, Stéphane Lebonvallet, Qingmin Liao, and Yuemin Zhu. Kernel feature selection to fuse multi-spectral MRI images for brain tumor segmentation. Computer Vision and Image Understanding: CVIU, 115(2):256–269, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zaheer:2018:SVR

[ZRRK18] Aamer Zaheer, Maheen Rashid, Muhammad Ahmed Riaz, and Sohaib Khan. Single-view reconstruction using orthogonal line-pairs. Computer Vision and Image Understanding: CVIU, 172(??):107–123, July 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314217302059.

Zhang:2024:ECI

[ZRZ^{+}24] Songyang Zhang, Ge Ren, Xiaoxi Zeng, Liang Zhang, Kailun Du, Gege Liu, and Hong Lin. Efficient crossinformation fusion decoder for semantic segmentation. Computer Vision and Image Understanding: CVIU, 240 (??):??, March 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002989.

Zheng:2011:ART

[ZS11] Liying Zheng and Daming Shi. Advanced Radon transform using generalized interpolated Fourier method for straight line detection. Computer Vision and Image Understanding: CVIU, 115 (2):152–160, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zimmermann:2019:FTM

[ZS19] Roland S. Zimmermann and Julien N. Siems. Faster training of Mask R-CNN by focusing on instance boundaries. Computer Vision and Image Understanding: CVIU, 188(??):Article 102795, November 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S1077314218303059.

Zhang:2023:AIS

[ZSC^{+}23] Qiao Zhang, Xiaoxiao Sun, Yurui Chen, Yanliang Ge, and Hongbo Bi. Attentioninduced semantic and boundary interaction network for camouflaged object detection. Computer Vision and Image Understanding: CVIU, 233(??): ??, August 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000991.

Zeng:2008:TCN

[ZSCP08] Yun Zeng, Dimitris Samaras, Wei Chen, and Qunsheng Peng. Topology cuts: a novel min-cut/max-flow algorithm for topology preserving segmentation in N– D images. Computer Vision and Image Understanding: CVIU, 112(1):81–90, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhan:2019:DIM

[ZSDK19] Huijing Zhan, Boxin Shi, Ling-Yu Duan, and Alex C. Kot. DeepShoe: an improved multi-task viewinvariant CNN for street-to-shop shoe retrieval. Computer Vision and Image Understanding: CVIU, 180 (??):23–33, March 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314219300013.

Zhao:2020:RLI

[ZSG^{+}20] Lin Zhao, Meimei Shang, Fei Gao, Rongsheng Li, Fei Huang, and Jun Yu. Representation learning of image composition for aesthetic prediction. Computer Vision and Image Understanding: CVIU, 199(??):Article 103024, October 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300801.

Zamorski:2023:CLP

[ZSK^{+}23] Maciej Zamorski, Michał Stypułkowski, Konrad Karanowski, Tomasz Trzciński, and Maciej Zieba. Continual learning on 3D point clouds with random compressed rehearsal. Computer Vision and Image Understanding: CVIU, 228(??): ??, February 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000012.

Zhang:2016:RDE

[ZSL^{+}16] Shuo Zhang, Hao Sheng, Chao Li, Jun Zhang, and Zhang Xiong. Robust depth estimation for light field via spinning parallelogram operator. Computer Vision and Image Understanding: CVIU, 145(??):148– 159, April 2016. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215002714.

Zhang:2024:RAS

[ZSL^{+}24] Zhanjie Zhang, Jiakai Sun, Guangyuan Li, Lei Zhao, Quanwei Zhang, Zehua Lan, Haolin Yin, Wei Xing, Huaizhong Lin, and Zhiwen Zuo. Rethink arbitrary style transfer with transformer and contrastive learning. Computer Vision and Image Understanding: CVIU, 241(??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000328.

Zhang:2016:GGG

[ZSSF16] Wenhao Zhang, Melvyn L. Smith, Lyndon N. Smith, and Abdul Farooq. Gender and gaze gesture recognition for human-computer interaction. Computer Vision and Image Understanding: CVIU, 149(??):32– 50, August 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300078.

Zhang:2019:VGM

[ZSY^{+}19] Quanshi Zhang, Xuan Song, Yu Yang, Haotian Ma, and Ryosuke Shibasaki. Visual graph mining for graph matching. Computer Vision and Image Understanding: CVIU, 178(??):16–29, January 2019. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218304442.

Zheng:1998:GDP

[ZT98] Jiang Yu Zheng and Saburo Tsuji. Generating dynamic projection images for scene representation and understanding. Computer Vision and Image Understanding: CVIU, 72(3):237– 256, December 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1998.
0678/production;   http:
//www.idealibrary.com/
links/artid/cviu.1998.
0678/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1998.
0678/production/ref.

Zhao:2009:MOR

[ZT09] Qi Zhao and Hai Tao. A motion observable representation using color correlogram and its applications to tracking. Computer Vision and Image Understanding: CVIU, 113(2):273–290, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2015:HFD

[ZT15] Chenyang Zhang and Yingli Tian. Histogram of 3D facets: a depth descriptor for human action and hand gesture recognition. Computer Vision and Image Understanding: CVIU, 139(??):29–39, October 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314215001216.

Zhang:2020:AER

[ZTB20] Bowen Zhang, Benedetta Tondi, and Mauro Barni. Adversarial examples for replay attacks against CNNbased face recognition with anti-spoofing capability. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 102988, August 2020. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314220300606.

Zhang:2018:DDA

[ZTGL18] Chenyang Zhang, Yingli Tian, Xiaojie Guo, and Jingen Liu. DAAL: Deep activation-based attribute learning for action recognition in depth videos. Computer Vision and Image Understanding: CVIU, 167(??):37–49, February 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301996.

Zhang:2011:MSS

[ZTH^{+}11] Shiliang Zhang, Qi Tian, Gang Hua, Wengang Zhou, Qingming Huang, Houqiang Li, and Wen Gao. Modeling spatial and semantic cues for large-scale nearduplicated image retrieval. Computer Vision and Image Understanding: CVIU, 115 (3):403–414, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2014:OTS

[ZTH^{+}14] Shiliang Zhang, Qi Tian, Gang Hua, Qingming Huang, and Wen Gao. ObjectPatch-Net: Towards scalable and semantic image annotation and retrieval. Computer Vision and Image Understanding: CVIU, 118(??): 16–29, January 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314213001574.

Zhang:1996:P

[ZTS96] Ruo Zhang, Ping-Sing Tsai, and Mubarak Shah. Photomotion. Computer Vision and Image Understanding: CVIU, 63(2):221– 231, March 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0016/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/cviu.1996.
                                        0016/production/pdf.

Zhuge:2009:ISS

[ZU09] Ying Zhuge and Jayaram K. Udupa. Intensity standardization simplifies brain MR image segmentation. Computer Vision and Image Understanding: CVIU, 113(10):1095–1103, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zunic:2003:DTC

[Žun03] Joviša Žunić. On discrete triangles characterization. Computer Vision and Image Understanding: CVIU, 90(2):169–189, May 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhuge:2006:VSB

[ZUS06] Ying Zhuge, Jayaram K.

Udupa, and Punam K. Saha. Vectorial scale-based fuzzy-connected image segmentation. Computer Vision and Image Understanding: CVIU, 101(3):177– 193, March 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zabrodsky:1997:UBS

[ZW97] Hagit Zabrodsky and Daphna Weinshall. Using bilateral symmetry to improve 3D reconstruction from image sequences. Computer Vision and Image Understanding: CVIU, 67(1): 48–57, July 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.idealibrary.com/
links/artid/cviu.1996.
0506/production;   http:
//www.idealibrary.com/
links/artid/cviu.1996.
0506/production/pdf;
http://www.idealibrary.
com/links/artid/cviu.1996.
0506/production/ref.

Zhang:2022:COD

[ZWB^{+}22] Cong Zhang, Kang Wang, Hongbo Bi, Ziqi Liu, and Lina Yang. Camouflaged object detection via neighbor connection and hierarchical information transfer. Computer Vision and Image Understanding: CVIU, 221 (??):??, August 2022. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314222000637.

Zhu:2016:CBE

[ZWL16] Guibo Zhu, Jinqiao Wang, and Hanqing Lu. Clustering based ensemble correlation tracking. Computer Vision and Image Understanding: CVIU, 153(??): 55–63, December 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314216300534.

Zhang:2024:RFE

[ZWLH24] Jianyang Zhang, Wei Wang, Xiangyu Li, and Yanjiang Han. Recognizing facial expressions based on pyramid multi-head grid and spatial attention network. Computer Vision and Image Understanding: CVIU, 244(??):??, July 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314224000912.

Zhu:2014:CER

[ZWN14] Shiai Zhu, Xiao-Yong Wei, and Chong-Wah Ngo. Collaborative error reduction for hierarchical classification. Computer Vision and Image Understanding: CVIU, 124(??):79– 90, July 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314214000769.

Zhang:2014:FFS

[ZWT^{+}14] Yong-Dong Zhang, Yu Wang, Sheng Tang, Steven C. H. Hoi, and Jin-Tao Li. FSpH: Fitted spectral hashing for efficient similarity search. Computer Vision and Image Understanding: CVIU, 124 (??):3–11, July 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000241.

Zhang:2020:PCB

[ZWW^{+}20] Xiaoqin Zhang, Tao Wang, Jinxin Wang, Guiying Tang, and Li Zhao. Pyramid channel-based feature attention network for image dehazing. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 103003, August 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300709.

Zhu:2014:TSC

[ZWY14] Guokang Zhu, Qi Wang, and Yuan Yuan. Tag-Saliency: Combining bottom-up and top-down information for saliency detection. Computer Vision and Image Understanding: CVIU, 118 (??):40–49, January 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314213001586.

Zhao:2016:LWP

[ZWZ^{+}16] Chaoyang Zhao, Jinqiao Wang, Guibo Zhu, Yi Wu, and Hanqing Lu. Learning weighted part models for object tracking. Computer Vision and Image Understanding: CVIU, 143(??):173– 182, February 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215002143.

Zhang:2018:MDO

[ZWZZ18] Quanshi Zhang, Ying Nian Wu, Hao Zhang, and Song-Chun Zhu. Mining deep and-or object structures via cost-sensitive questionanswer-based active annotations. Computer Vision and Image Understanding: CVIU, 176– 177(??):33–44, November/ December 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314218302972.

Zeng:2020:HIR

[ZXC^{+}20] Haijin Zeng, Xiaozhen Xie, Haojie Cui, Yuan Zhao, and Jifeng Ning. Hyperspectral image restoration via CNN denoiser prior regularized low-rank tensor recovery. Computer Vision and Image Understanding: CVIU, 197–198(??):Article 103004, August 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314220300710.

Zhu:2002:RTA

[ZXK02] Yuanxin Zhu, Guangyou Xu, and David J. Kriegman. A real-time approach to the spotting, representation, and recognition of hand gestures for human– computer interaction. Computer Vision and Image Understanding: CVIU, 85(3): 189–208, March 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zamalieva:2014:BSM

[ZY14] Daniya Zamalieva and Alper Yilmaz. Background subtraction for the moving camera: a geometric approach. Computer Vision and Image Understanding: CVIU, 127(??):73–85, October 2014. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314214001349.

Zhang:2023:IFS

[ZYQ^{+}23] Lu Zhang, Xu Yang, Lu Qi, Shaofeng Zeng, and Zhiyong Liu. Incremental fewshot object detection with scale- and centerness-aware weight generation. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314223001546.

Zhou:2009:OTU

[ZYS09] Huiyu Zhou, Yuan Yuan, and Chunmei Shi. Object tracking using SIFT features and mean shift. Computer Vision and Image Understanding: CVIU, 113(3): 345–352, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zheng:2010:DOB

[ZYT10] Songfeng Zheng, Alan Yuille, and Zhuowen Tu. Detecting object boundaries using low-, mid-, and highlevel information. Computer Vision and Image Understanding: CVIU, 114(10):

1055–1067, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2013:RIR

[ZYXZ13] Jun Zhang, Lei Ye, Yang Xiang, and Wanlei Zhou. Robust image retrieval with hidden classes. Computer Vision and Image Understanding: CVIU, 117(6): 670–679, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000520.

Zhang:2024:SSM

[ZYZ24] Heng Zhang, Yi-Jun Yang, and Wei Zeng. Selfsupervised multi-scale semantic consistency regularization for unsupervised image-to-image translation. Computer Vision and Image Understanding: CVIU, 241(??):??, April 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224000316.

Zhang:2007:JDZ

[ZZ07] Lei Zhang and David Zhang.

A joint demosaicking–zooming scheme for single chip digital color cameras. Computer Vision and Image Understanding: CVIU, 107(1– 2):14–25, July/August 2007.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhao:2010:SAL

[ZZ10] Keke Zhao and Zhenyue Zhang. Successively alternate least square for lowrank matrix factorization with bounded missing data. Computer Vision and Image Understanding: CVIU, 114(10):1084–1096, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2020:PLF

[ZZ20] Ming Zhang and Bing Zeng. A progressive learning framework based on single-instance annotation for weakly supervised object detection. Computer Vision and Image Understanding: CVIU, 193(??):Article 102903, April 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314220300011.

Zhang:2013:ASA

[ZZC^{+}13] Shaoting Zhang, Yiqiang Zhan, Xinyi Cui, Mingchen Gao, Junzhou Huang, and Dimitris Metaxas. 3D anatomical shape atlas construction using mesh quality preserved deformable models. Computer Vision and Image Understanding: CVIU, 117(9):1061– 1071, September 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314213000799.

Zeng:2023:FIE

[ZZC^{+}23] Hao Zeng, Wei Zhang, Keyu Chen, Zhimeng Zhang, Lincheng Li, and Yu Ding. Face identity and expression consistency for game character face swapping. Computer Vision and Image Understanding: CVIU, 236(??):??, November 2023. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314223001868.

Zhao:2014:SLB

[ZZCL14] Mingbo Zhao, Zhao Zhang, Tommy W. S. Chow, and Bing Li. Soft label based Linear Discriminant Analysis for image recognition and retrieval. Computer Vision and Image Understanding: CVIU, 121(??): 86–99, April 2014. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314214000150.

Zhou:2024:CCI

[ZZG^{+}24] Dazheng Zhou, Mingliang Zhang, Xianjie Gao, Youmei Zhang, and Bin Li. Complete contextual information extraction for selfsupervised monocular depth estimation. Computer Vision and Image Understanding: CVIU, 245(??): ??, August 2024. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314224001139.

Zeng:2023:RWE

[ZZH23] Guotian Zeng, Bi Zeng, and Huiting Hu. Realworld efficient fall detection: Balancing performance and complexity with FDGA workflow. Computer Vision and Image Understanding: CVIU, 237(??): ??, December 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223002126.

Zhao:2023:USI

[ZZHZ23] Boyu Zhao, Qian Zhou, Lijun Huang, and Qiang Zhang. Unpaired sonar image denoising with simultaneous contrastive learning. Computer Vision and Image Understanding: CVIU, 235 (??):??, October 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223001637.

Zeppelzauer:2018:STD

[ZZJS18] Matthias Zeppelzauer, Bartosz Zieliński, Mateusz Juda, and Markus Seidl. A study on topological descriptors for the analysis of 3D surface texture. Computer Vision and Image Understanding: CVIU, 167(??):74–88, February 2018. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1077314217301753.

Zamorski:2020:AAC

[ZZK^{+}20] Maciej Zamorski, Maciej Zieba, Piotr Klukowski, Rafał Nowak, Karol Kurach, Wojciech Stokowiec, and Tomasz Trzciński. Adversarial autoencoders for compact representations of 3D point clouds. Computer Vision and Image Understanding: CVIU, 193(??):Article 102921, April 2020. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S107731422030014X.

Zhang:2013:CRI

[ZZL13] Jun Zhang, Heng Zhao, and Jimin Liang. Continuous rotation invariant local descriptors for texton dictionary-based texture classification. Computer Vision and Image Understanding: CVIU, 117(1): 56–75, January 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314212001348.

Zhou:2012:HOS

[ZZP12] Bingyin Zhou, Fan Zhang, and Lizhong Peng. Higherorder SVD analysis for crowd density estimation. Computer Vision and Image Understanding: CVIU, 116(9):1014–1021, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1077314212000884.

Zafeiriou:2016:ESI

[ZZP^{+}16] Stefanos Zafeiriou, Guoying Zhao, Matti Pietikainen, Rama Chellappa, Irene Kotsia, and Jeffrey Cohn. Editorial of special issue on spontaneous facial behaviour analysis. Computer Vision and Image Understanding: CVIU, 147(??): 50–51, June 2016. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314216300376.

Zhang:2015:OLE

[ZZRC15] Shu Zhang, Yingying Zhu, and Amit Roy-Chowdhury. An online learned elementary grouping model for multi-target tracking. Computer Vision and Image Understanding: CVIU, 134 (??):64–73, May 2015. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314215000168.

Zhang:2023:MYI

[ZZS^{+}23] Yichi Zhang, Zijian Zhu, Hang Su, Jun Zhu, Shibao Zheng, Yuan He, and Hui Xue. To make yourself invisible with Adversarial Semantic Contours. Computer Vision and Image Understanding: CVIU, 230 (??):??, April 2023. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1077314223000395.

Zhang:2021:SSA

[ZZSD21] Xiao-Yu Zhang, Yaru Zhang, Haichao Shi, and Jing Dong. SAPS: Self-Attentive Pathway Search for weaklysupervised action localization with background-action augmentation. Computer Vision and Image Understanding: CVIU, 210(??): ??, September 2021. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1077314221001004.

Zhang:2006:AMI

[ZZZ06] Yongjun Zhang, Zuxun Zhang, and Jianqing Zhang. Automatic measurement of industrial sheetmetal parts with CAD data and nonmetric image sequence. Computer Vision and Image Understanding: CVIU, 102(1):52–59, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zafeiriou:2015:SFD

[ZZZ15] Stefanos Zafeiriou, Cha Zhang, and Zhengyou Zhang. A survey on face detection in the wild: Past, present and future. Computer Vision and Image Understanding: CVIU, 138(??):1–24, September 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1077314215000727.

Zhao:2009:OSP

[ZZZP09] Y. Zhao, L. Zhang, D. Zhang, and Q. Pan. Object separation by polarimetric and spectral imagery fusion. Computer Vision and Image Understanding: CVIU, 113 (8):855–866, August 2009. CODEN CVIUF4. ISSN REFERENCES 692

1077-3142 (print), 1090-235X (electronic).